A Complete Bibliography of Publications in the Bulletin of Mathematical Biology

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

08 February 2024 Version 1.24

Title word cross-reference

(1 + 1) [2578]. (M, R) [20, 1, 614]. (n − 1) [368, 404]. + [2271]. 1 [3341]. $1000 [1001]. $130.00 [1648]. $140.00 [1840]. 2 [2083, 2024, 2564, 2127, 3082, 1021, 3571, 3128]. 28 [1889]. $29.95 [1646]. 2n = 40 [2954]. 3 [3537, 2564, 2379, 3221]. $34.95 [1899]. $39.50 [1647]. 4 [714]. $40.00 [1920]. $42.95 [1900]. $45.00 [1930]. $69.95 [1831]. $75.00 [2143]. $79.95 [1802, 1940]. $80.00US [1819]. $89.95 [1818]. $95.00 [1877, 2129]. ^{+} [3303, 2785, 1926, 3054, 2266]. ^{1} [759]. ^{2+} [1328, 1434, 2282, 1562, 1667, 2392, 1942, 1573, 1941]. _{1} [2109]. _{2} [93, 1810]. _{3} [2835, 1942]. _{4}^{+} [3341]. _{6} [2598]. _{60} [3010]. _{W} [15]. A [1378]. A_{1} [1876]. Aβ [1767]. A ⇀ ↽ B [1130]. β [3523, 2515, 923, 2921, 1344, 3249, 2071]. Ca^{2+} [3503]. CD4^{+}T [3091]. d [3609]. δ = 2μ [645]. V̇_{A}/Q̇ [247]. V̇ O_{2} [2286]. ϵ [2716]. G [2933, 3625]. G_{1} [2103]. Gα_{i} [2933, 3625]. γ [2515]. Gβγ [2933, 3625]. K [626, 3467, 3494, 738]. κ [2972]. N [1764, 2343, 110, 169, 357, 425, 2640, 75, 526, 549]. N^{6} [1876]. o [642].

O(N^{2} log N) [1191]. p [75]. R [3557, 80]. R_{0} [2141, 2930]. S [3557, 771]. T [102]. τ [3557]. V_{t} [501]. φ [512]. X_{a} [501].

-Based [2735]. -Ca [1942]. -cell [2071]. -Cells [3249, 923, 1344]. -compartment [110, 169, 357, 549]. -compartmental [425, 626]. -Consensus [771]. -current [1378]. -cycles [2083]. -deoxyglucose [1021]. -Dimensional [2578]. -Dynamics [2392]. -gon [1889]. -Leaping [3557, 3557]. -Mediated [3341]. -mer [3467]. -patch [1764]. -player [3609]. -Secretase [2515]. -site [714]. -system [102]. -systems [20, 1, 614]. -taxon [2343]. -trees [526]. -Tuple-Site [3494]. -Uniqueness [642].

/S [2103].

0 [1646, 1900, 1802, 1819, 1647, 1877, 1847, 1648, 1840]. 0-12-256781-1 [1819]. 0-262-07226-2 [1900]. 0-387-95228-4 [1940]. 0-387-95354-X [1802]. 0-387-98992-7 [1818]. 0-521-45705-X [1646]. 0-521-52586-1 [1920]. 0-521-62171-2 [1847]. 0-521-78579-0 [1840]. 0-521-80038-2 [1877]. 0-521-81980-6 [1920]. 0-521-82700-0 [2129]. 0-521-84150-X [2143]. 0-691-00541-9 [1647]. 0-8018-7403-3 [1930]. 0-817-65792-4 [1648].

1 [122, 1724, 3262, 1258, 2017, 2952, 1229, 1819, 2898, 2692, 3508, 2302, 2881, 2096, 1788, 2354, 2186, 1831, 1890, 1957, 1899]. 1-85233-536-X [1899]. 1-85233-560-2 [1831]. 15 [1362]. 1977 [302]. 1983 [746]. 1986 [861]. 1988 [3628]. 1st [1101].

2 [122, 3290, 1258, 1900, 2357, 2693, 1877, 1831, 1847]. 2001 [1818]. 2N [2952]. 2N-1 [2952]. 2nd [1818].

3 [1257, 1107, 2863]. 3’-dideoxycytidine [1257, 1107]. 3/8 [2863]. 3rd [1159].

4 [1648].

5 [1876, 3628]. 5-Trisphosphate [2081, 1573].

60.00/$90.00 [1847]. 60th [1123].

7 [3279, 3410, 2672].

8 [2863].

9 [1647]. 9th [505].

A/H3N2 [3345]. Abdul [1802]. Aberrant [2505]. abnormal [163].

Abrams [3616]. abrin [598]. Abrupt [1425]. Absolute [2879]. absorbing [1213, 409]. Absorption [2102, 401, 1465, 650, 1869]. abstract [675, 265]. abstractions [226]. Abundance [2907, 1831]. abundances [1469]. Abuse [3195, 2764]. Academic [1819]. Accelerate [2268]. Accelerated [3557]. Acceleration [3340, 3127, 772, 781]. accelerations [694]. Access [2506, 1171]. accommodation [1734]. Accomplishing [2732, 2731]. According [3205]. account [1863]. Accounting [3380]. Accumulation [2727, 157, 494]. Accuracy [3555, 3354]. Accurate [3131, 3372, 3503, 3476, 1929, 1106]. acervorum [1094, 1095]. Acid [3262, 3290, 1947, 1301, 1820, 39, 1361, 1342, 723, 756, 40]. Acids [2124]. Acinar [3503, 2088, 3476]. Acini [2256, 2291]. Ackers [2298]. Acoustic [2174, 1869]. Acrosome [3415]. Across [3515, 2898, 3519, 667, 102, 1301, 686, 101]. Actin [2911, 2908, 2596, 1539, 1306, 1542, 1518, 1519, 1252, 1510, 1487]. Acting [3583, 3146, 1519]. Action [3196, 2539, 2324, 3055, 3295, 507, 2697, 1426, 1744, 1203, 1236, 1323, 237, 274, 1407]. Action-angle [507]. Actions [2323, 2636, 3204]. Activate [2298]. Activated [3380, 1763]. Activation [3265, 2030, 2707, 3244, 2406, 178, 1249, 2033, 1130, 501, 403, 855]. Activator [2732, 2731]. Activators [2298]. Active [3606, 2007, 2345, 2494, 2354, 2703, 479, 706]. Activin [2323]. Activities [2593, 262, 991]. Activity [2449, 3159, 2163, 2024, 2280, 2266, 1081, 2372, 53, 179, 1354, 1067, 507, 459, 965, 1866]. activity-dependent [1866]. Acto [3082]. Acto-Cytosolic [3082]. Actomyosin [2974, 1539]. acts [424]. Acute [2663, 3320, 3070, 3246, 2031]. Adaptability [893]. Adaptation [3359, 2044, 2428, 2367, 2466, 394, 471, 1234, 984]. adaption [1684]. Adaptive [2896, 3507, 2638, 718, 2576, 2858, 3557, 2618, 1441, 3250, 1699, 1777, 1934]. adaptiveness [1769]. addendum [203]. Adding [2329]. Addition [2702]. Additional [2603]. Additive [3370, 1011, 1049]. address [303]. adduction [1444]. Adenosine [2109, 1876]. adenosine-5’-uronamides [1876]. Adhesion [2329, 3532, 2754, 1652, 232, 1477, 1587]. adipose [1726]. adjacent [1525]. adjoint [50]. adjustable [1470]. Adjustment [1723]. adjuvant [1948]. Administration [3569, 1257, 1107]. Admit [3596]. Admixed [3468]. Adoptive [3363]. adsorbed [80]. adsorption [16, 264]. Adult [2668, 2799, 2900, 1489]. adults [1056]. Advance [2377, 1632]. Advantage [2749, 3614, 1672]. Advective [2292, 2467, 2762, 3248]. Adverse [2243]. Ae4 [3328]. Aedes [2078, 2284, 1945]. Aegypti [2078, 2284, 1945]. AER [2244]. aerial [1129]. aerobic [733, 727]. aerobic/anaerobic [727]. aerosol [442, 361]. Aerosolised [2116, 2101]. aeruginosa [3251, 1732, 1426, 1670]. Affect [2571]. Affecting [2387]. Affects [2577, 3228, 3378, 3586, 1924]. afferent [197]. Affine [2808, 888]. Affinity [3480, 2716, 1613]. Affinity-Dependent [3480]. Africa [2764]. African [2506, 2724]. Africanized [1974, 1936]. After [3395, 3398, 2280, 1273, 1439, 1821, 2075, 474, 657]. Against [2906, 3534, 2189, 2680, 2678, 3084, 2697, 1972, 1253, 1769]. Age [2506, 2169, 3260, 3399, 2963, 3062, 2134, 2667, 3382, 2680, 2100, 2930, 2187, 2218, 2352, 2694, 793, 719, 334, 1676, 1391, 691, 722, 1450, 699, 977, 1317, 737, 337, 1464]. Age- [2169, 691, 722]. age-dependent [719, 334, 337]. Age-Specific [2680]. Age-Stratified [2930]. Age-Structured [2506, 3062, 3382, 2218, 1450, 699, 977, 737, 1464]. Agent [3202, 2573, 2258, 3434, 3024]. Agent-Based [3202, 2258, 3434, 3024]. Agents [2702, 346, 1897, 1024, 1203, 1236, 1323, 1708]. aggreagation [443]. Aggregate [2203]. Aggregates [2112, 3361, 3565, 2344, 1096]. Aggregation [2951, 3523, 2366, 2064, 2099, 444, 1763, 1265, 1442, 243, 1887, 412, 1366]. Aggressive [3351]. Aging [2928, 3048]. Agiogenesis [2227]. Agonist [1491, 1817, 1876]. Agonist-induced [1491]. Agonists [2109]. Agroecosystems [3017]. aided [504]. AIDS [2015, 2402, 2630, 1838, 1401, 1318, 2187, 2580, 2261]. Aims [27, 28]. Air [3446, 141, 246]. Airway [2161, 2132, 1723, 190]. airways [548, 822, 514, 593, 442, 475, 282, 361]. Alcohol [2724]. Alcoholic [3296]. Alexei [1819]. Algae [2396, 1365, 1987]. algal [1561]. Algebra [2875, 2993, 2676, 838, 1708]. Algebraic [3461, 2834, 2640, 3106, 3036, 2104, 2568, 3463, 2537, 20, 85, 1719]. algebraic-combinatorial [1719]. algebras [262]. Algorithm [2445, 3555, 3617, 2619, 3468, 3557, 2065, 2410, 3597, 1013, 1049, 1807, 1756, 1191, 738, 1595, 1188, 1829]. Algorithmic [3150, 2546, 10]. Algorithms [1005, 3233, 3305, 2386, 752, 3561, 220, 3556, 1266, 902, 375, 1312]. Alien [2185, 3074]. aligned [1330]. Alignment [2542, 888, 1246, 754, 1079, 1227, 1543, 1756, 1188, 1252, 1896, 1487, 1386, 749, 1312, 1006]. alignments [1088, 1929]. all-cause [1728]. Allee [3383, 2201, 2423, 3171, 2565, 3136, 3502, 2100, 1950, 2243, 2287, 1970, 3344]. allergy [1435]. Allman [1920]. Allocation [2316, 3515, 2885, 3118, 2461, 1718, 1912, 1813, 1223, 1266]. Allochthonous [3623]. allometric [1356]. allometry [1986, 1545, 1597, 1917]. allosteric [1619]. allowing [1079]. Almost [249]. alone [1656]. along [630, 93, 1152, 1300, 1425]. alpha [583]. Alter [2982]. Alternate [2938]. Alternating [2581]. Alternative [3004, 3252, 3124, 3388, 1533, 3281, 1372, 1342, 1021]. Alters [2762]. Altman [2058]. Altruistic [2618]. Aluru [2222]. alveolar [550]. Alzheimer [3523, 1767, 1808]. amacrine [1209]. Amalgamation [3409]. ambiguous [219]. ameboid [1839]. America [850, 2963, 3084]. American [3583, 2151]. Amino [2124, 723, 756]. aminoacyl [1468, 1889]. aminoacyl-tRNA [1468, 1889]. ammonium [1428]. amoebae [412]. Among [2704, 2015, 2419, 2316, 2188, 3227, 371, 1448, 1523, 800, 2073, 1460]. AMP [2043]. amphibian [1047]. Amphibians [3125]. amphiphiles [1442]. amplification [1294]. Amplify [3393]. Amyloid [3523, 2515]. Amyloid- [3523, 2515]. Anabolic [3204]. Anaemia [2125]. Anaerobic [2286, 727]. analog [1012, 400]. Analogies [2525, 2526]. analogue [1832]. analogue-modelling [1832]. analogy [1430, 79]. Analyses [2369, 2677, 2214, 3220, 535, 1377, 168]. Analysing [3039]. Analysis [2927, 2004, 2944, 2906, 3523, 651, 3383, 3330, 2589, 3364, 3391, 2247, 2661, 3198, 3264, 3580, 2999, 3618, 671, 2833, 1145, 2534, 3555, 2283, 3528, 2392, 2209, 3303, 2688, 750, 2116, 3229, 2486, 1020, 2949, 3350, 3447, 1383, 3261, 1560, 3585, 3319, 1147, 2376, 2298, 2588, 2074, 3559, 3517, 2750, 3348, 2884, 2540, 2636, 2913, 3258, 3289, 268, 2568, 3601, 2674, 1836, 3468, 3504, 1314, 1780, 3007, 1024, 3527, 2641, 3217, 2350, 2591, 2950, 2806, 2542, 2405, 3605, 398, 2333, 3037, 2149, 2433, 1164, 3002, 3338, 3164, 3356, 2768, 3150, 2872, 3299, 2981, 2221, 2143, 2814, 3166, 2443, 2590]. Analysis [2483, 3429, 3143, 422, 423, 2545, 3510, 2280, 3310, 323, 2286, 3445, 2549, 628, 3405, 3595, 2059, 2172, 2297, 1096, 1189, 2533, 1859, 3569, 2413, 2248, 3471, 930, 3063, 3421, 2427, 3120, 626, 433, 783, 1263, 371, 630, 421, 458, 158, 205, 286, 1436, 1154, 1550, 1054, 1881, 690, 1175, 1536, 487, 444, 144, 346, 1334, 63, 1017, 308, 236, 1640, 1823, 299, 1329, 295, 1588, 379, 1963, 270, 834, 948, 172, 829, 926, 1156, 538, 1308, 557, 909, 1727, 30, 658, 798, 1429, 1743, 1292, 740, 1590, 567, 1801, 1362, 1735, 455, 1728, 70, 1161, 700, 29, 1877]. analysis [1984, 554, 366, 208, 543, 269, 453, 785, 1490, 1625, 945, 730, 801, 921, 1140, 1601, 1322, 871, 1910, 1700, 1840, 373, 3626]. Analytic [1589, 2273, 2924, 1057, 291]. Analytical [786, 1273, 2702, 2247, 2908, 441, 593, 181, 2124, 327, 692, 2005, 2281, 292, 2289, 335, 3573, 3569, 1504, 1729, 907, 2453, 1882]. analytically [191]. Analyzing [2834, 2980, 2487]. anaphylaxis [1284]. anatomical [431, 442]. Anatomically [3372]. anatomy [955]. Ancestor [2851]. Ancestral [3292, 3465, 3312, 3354, 2487]. ancient [955]. Androgenic [2883]. Anemia [3313, 434]. anesthetic [1607]. aneurysms [343]. Angiogenesis [2825, 2451, 2970, 2220, 2775, 1534, 993, 1355, 1703, 1741, 1753, 1449, 1914, 1937]. angiogenic [1795, 2691]. Angiostatin [2775]. Angiotensin [3355]. angle [507]. angles [288]. Anhui [3413]. Animal [1117, 2333, 2548, 3211, 3182, 1414, 1583, 1196, 680, 1225, 1831]. Animals [3458, 2350, 1835, 1129, 715, 852]. Anisotropic [2992, 3371, 1488, 52, 78]. Anisotropy [3478, 2075]. Annihilates [3603]. annotated [760]. announce [1001]. Announcement [26, 43, 44, 710, 734, 1044, 1233, 894]. Announcements [452, 584, 597, 632, 697, 777, 803, 904, 981, 1085, 1197, 951]. announces [125, 149]. Annual [505, 562, 850, 1075, 303, 1001]. Anode [2675]. anoxia [629]. Ant [3026, 1718, 1094, 1095, 1223, 1708, 1716]. Antagonism [2659, 1757]. Antagonistic [2188]. Anterior [1992]. Anthony [140]. Anthrax [3458, 3211]. Anti [3398, 2659, 2970, 1316, 2466, 2173, 1588]. Anti-Angiogenesis [2970]. Anti-Cancer [2173]. Anti-HIV [2466]. Anti-phase [1316]. anti-predatory [1588]. Anti-tumour [2659]. Antibiotic [3175, 2709, 2678, 3207, 1426]. Antibiotic-Resistant [2678]. Antibiotics [2196, 3389, 1821]. Antibody [3480, 3592, 3574, 2466, 1276, 444, 443, 21]. Antibody-Based [3574]. Antibody-Mediated [3592]. anticancer [1816, 870, 871]. Anticipate [3206]. Antigen [3544, 444, 443, 21, 1849]. antigen-antibody [444, 443]. antigen-like [21]. Antigenic [3253, 2741, 2532, 1479]. Antimalarial [3457]. Antimicrobial [3201]. Antimitotic [2170]. Antiresorptive [3204]. Antiretroviral [3303, 2354, 2186, 1788]. Antiretrovirals [2506]. Antiviral [3010, 2241, 2337, 3118, 2416]. Ants [2847, 1708, 3011, 1406, 1571, 1266]. Any [2289]. aorta [716]. aortic [380, 128]. aperiodic [1608, 1316]. aphakic [499]. Apical [3564]. Apoptosis [2590]. apparatus [378, 1752]. Apparent [2384]. appearance [1860]. Applicability [2359]. applicable [1147]. Application [2824, 255, 2661, 157, 3283, 2473, 3162, 143, 216, 1329, 1567, 2911, 2025, 2157, 2102, 3241, 2669, 2807, 3352, 3258, 3077, 3098, 3088, 2713, 215, 2124, 2078, 3152, 2796, 2352, 1704, 2161, 3174, 1275, 3471, 3064, 3275, 3421, 1718, 390, 1828, 1065, 201, 563, 108, 168, 1405, 507, 1910, 475, 2290, 633, 3337]. Applications [2802, 3290, 2162, 2122, 1802, 2839, 3530, 2701, 668, 2641, 3560, 2375, 791, 3129, 2948, 2152, 122, 750, 417, 672, 673, 1607, 220, 34, 677, 2026, 627, 1039]. Applied [3433, 1101, 3480, 3552, 573, 485, 2492, 2672, 764, 333, 354, 1585, 1363, 413]. Approach [3145, 2837, 2150, 2875, 3273, 3367, 3265, 2236, 2959, 2831, 2808, 3594, 2456, 2907, 3554, 2640, 2647, 3443, 2588, 3477, 2399, 2659, 2326, 3055, 3493, 2104, 3575, 2641, 2583, 3223, 3380, 2047, 222, 2406, 3164, 3483, 2208, 2143, 2891, 3612, 3444, 3208, 3296, 3525, 2181, 3470, 2002, 2356, 3064, 626, 1439, 381, 411, 1659, 1725, 1634, 342, 1909, 1180, 1530, 1500, 695, 551, 1745, 85, 827, 1812, 1028, 1162, 684, 234, 937, 504, 547, 576, 507, 1291, 345, 521, 385, 1708, 846, 908, 524, 113, 1281, 661, 221, 1893]. Approaches [2999, 1802, 3348, 3608, 3150, 3553, 2338, 2294, 2295, 3017, 3282, 35, 1538, 396]. Appropriate [2863, 3039, 1685, 1742, 876]. Approximate [2837, 946, 3605, 1004, 372]. Approximating [2840, 3030, 3108]. Approximation [2141, 3067, 3550, 2914, 2113, 2431, 2410, 3129, 1388, 822, 1638, 392, 1731, 830, 1106, 9, 244, 1283, 917]. Approximations [3072, 3559, 3499, 2304, 3349, 3137, 1951, 1694, 1202, 1960, 1440]. aquatic [1129, 1904, 1428, 1429, 1267]. Aqueous [1992, 324]. Arabidopsis [3564, 3460]. arabinose [3268]. Arbitrary [3229, 2839, 132, 193]. Arbuscular [2599]. ARC [84]. Archaea [2345]. Archiving [2720]. arcoid [1745]. Arctic [1910]. area [377, 1799, 9]. Argentine [1716]. Arid [2674]. Arise [2514]. Arising [2604, 2338, 2753, 1773, 467]. Armitage [3346]. army [1406, 1571]. Array [2992, 1854]. Arrays [2948]. arrival [1280]. arrivals [154]. Art [2048, 3184]. Arterial [781, 22, 315, 578, 943, 849, 1417, 70, 1103, 588, 907, 316, 610]. arterial-venous [1417]. arteries [809, 816, 1392, 1485, 916, 82, 139, 52, 78, 914, 694]. arterioles [809]. arteriosclerotic [826]. artery [820, 915, 652, 496, 466]. Arthropods [97]. Arthur [1818]. article [573]. Articular [2869]. artificial [1071, 945]. Arvind [1847]. Ascaris [3138, 3427]. Ascending [1298]. Asexual [2395, 2643]. aspect [1116, 1796, 289]. Aspects [152, 55, 257, 1641, 1658, 1236, 1911, 1120]. Assay [3273, 2192, 3349, 1854]. Assays [3238, 1147]. Assembly [2539, 3508, 131, 1731]. Assessing [3575, 2655, 3227, 2187, 2208, 3400, 1888, 1972]. Assessment [2729, 2563, 2017, 1853, 1450]. assimilation [1363, 1864]. assistance [128, 391]. Associated [2116, 3042, 2420, 2447, 2830, 2095, 2450, 3288, 2012, 1011, 1288, 1704, 587]. association [1985]. Associations [2954, 1015]. assumed [357]. assuming [297]. assumption [995]. assumptions [1030, 443]. assurance [1758, 1961]. aster [1752]. Asters [2041]. Asthmatic [3276]. Asymmetric [3292, 2734, 935, 660, 190, 1316]. Asymmetrical [3416, 580, 413]. asymmetries [1541]. Asymmetry [2516, 2146, 1201]. Asymptomatic [2973, 2229, 362]. Asymptotic [3260, 2247, 397, 2870, 3387, 60, 3346, 538, 2267, 3007, 1743, 3190, 3150, 2872, 468, 1490, 3513, 2046, 371, 63, 1964, 1416, 88]. Asymptotics [2510]. Asynchronous [2806, 3505, 2804, 2803]. Asynchrony [2296, 2673, 3268]. At-Risk [2425]. Athero [2923]. Athero-protective [2923]. Atherosclerosis [2923, 3042, 2481, 3591, 1725]. Atherosclerotic [3325, 82, 139]. Atlanta [590]. Atlantic [2667]. ATP [2371, 1668, 2836, 1520]. ATP-producing [1520]. atria [1897]. Atrial [2285]. atrioventricular [1771]. Attached [2192, 2682]. attachment [1539]. Attack [2715, 1467]. attacks [362]. Attention [2525, 1081, 2526]. Attenuant [2083]. Attenuated [2266]. attitude [39]. attractant [1147]. Attraction [3057]. Attractor [3505, 2525, 2526]. attractors [1060, 1603, 1221, 1798]. attributive [1879]. auditory [263]. Augmentation [3180]. aureus [2692, 2693]. Auto [2468, 434]. Auto-Correlation [2468]. auto-immune [434]. Autocatalysis [2979]. Autocatalytic [1424, 2886, 1124, 1387, 1240, 1241, 1087, 1132, 1333, 1661, 1952, 938, 1544]. Autocorrelation [2024, 1875]. Autocrine [2319, 1141]. Autoimmunity [2509, 1326]. Automata [138, 1019, 2330, 3221, 1161, 1941, 614]. Automated [3162]. Automatic [2580]. Automaton [2812, 3037, 1579]. Autonomous [2191, 2090, 3374, 2621, 1050]. autoreceptors [1582]. Autoregulated [2370]. autoregulation [1299, 1298]. Aux [2474]. Aux/IAA [2474]. Auxiliary [3447]. Auxin [3114]. Availability [3221, 2990, 1747]. Available [2840]. Avascular [2378]. Average [2844]. Avert [2488]. Averting [3457]. Avian [2973, 2065, 3605, 1807, 1632]. Avoidance [3250, 1875]. Avoiding [3041]. Award [1001]. Awareness [3282]. axial [946, 1218]. axially [166]. Axis [2883, 1231]. axisymmetric [166]. Axon [2094, 242, 45, 244]. Axonal [212, 213, 2628]. axons [398, 494]. Aziz [1802].

B [573, 1828, 1369, 1819, 2397, 1118, 1375, 609, 2598, 2770, 1199, 839, 2096, 2972]. B.C. [1123]. B7 [2642]. Ba [458, 421]. Ba-treated [458, 421]. Babesiosis [2882, 3029]. Bacillus [2513, 1843]. back [495]. backbone [1453].

Background [2744, 2945, 3032, 2525, 2526]. backpropagation [1407]. Backward [3526, 2458, 1926, 3603]. Bacteria [3474, 3262, 3290, 2422, 2608, 2982, 2348, 2678, 2728, 3060, 3545, 1477, 1478, 1553, 1347, 1397, 1398, 713, 359, 692, 733, 727]. Bacterial [2112, 2196, 2828, 2514, 3507, 3280, 2205, 2379, 2713, 3221, 2276, 3167, 3130, 2294, 2295, 2463, 2735, 2975, 1905, 843, 1959, 378, 1928, 1147, 1478, 1244, 743, 1166, 1359, 645, 211]. Bacteriocin [3262, 3290]. Bacteriocin-Producing [3262, 3290]. Bacteriophage [2608, 1346]. Bactria [2887]. bad [1956]. Badger [3486]. balance [918, 1729, 413]. Balanced [3418, 3235, 2619]. Balancing [2543]. balloon [128]. balls [840]. band [1905, 971]. bande [48, 33]. bandpass [48, 33]. Bands [2975, 843, 713]. Banister [3520]. Bank [2949]. banks [886]. Bards [2600]. Barley [2622]. barn [1814]. Baroreflex [3071]. Barr [2581]. Barrier [2196, 487, 686]. barrier-limited [487]. barriers [714, 1130, 1554, 1649]. Bart [1646]. Bartholomay [140]. Basal [2421]. Base [3110, 512]. Based [2004, 3202, 2951, 3462, 3424, 2837, 2670, 2060, 2999, 2052, 3594, 2940, 3107, 3303, 2825, 3448, 2423, 3335, 2749, 3389, 2573, 2399, 2630, 3401, 2610, 2183, 3493, 2568, 2258, 3468, 2365, 2701, 3463, 2170, 3434, 2713, 2224, 3574, 2663, 2350, 3380, 3221, 3227, 3147, 2302, 2312, 2491, 3333, 3024, 2490, 3483, 3368, 3135, 2778, 2114, 3556, 3129, 2945, 2173, 2482, 2735, 2214, 3390, 1726, 3144, 3564, 1892, 1536, 1951, 2550, 1217, 1689, 892, 2451, 1397, 1343, 1815, 245, 2438, 660, 1832, 1644, 1790, 1405, 712, 1605, 1661, 1358, 1854, 1582, 1399, 3485, 3032, 1512]. Bases [2023]. Basic [2605, 2141, 2403, 3305, 2016, 3271, 236, 443, 1874, 532]. Basis [1069, 4, 3627]. batch [1398]. Batesian [159]. Bay [2092]. Bayes [1543, 1565]. Bayes-optimal [1543]. Bayesian [3273, 3475, 3453, 3234, 2185, 2131, 2240, 1553, 901, 900, 2031, 1704, 2143]. BCG [2178]. Bcl [2227]. Bcl-2 [2227]. Be [3206, 1685, 357, 1956, 1096, 3077]. Be-CoDiS [3077]. bead [232]. Beating [2278]. because [2257]. Beds [2384]. Bee [3260, 3619, 3070, 3246, 2533, 1936, 1974, 1511]. Beehives [2919]. Bees [2174]. Beetle [3062, 2440, 2841, 1467]. before [1273]. Behavior [3015, 2201, 3387, 2710, 2016, 3468, 3295, 3347, 2226, 3586, 2533, 1617, 63, 163, 1259, 1260, 1676, 1588, 836, 714, 60, 702, 477, 825, 593, 1591, 360, 208, 18, 323, 347, 1528]. Behavioral [3247, 1646, 259, 585]. Behaviors [3566, 3023, 948]. Behaviour [3191, 2779, 2556, 2055, 3135, 3188, 182, 1412, 1563, 1094, 330, 1095, 680, 343, 1832, 422, 423, 1708, 1927, 1346, 1345]. Behavioural [2146, 1123]. Behaviours [2505, 2972, 1293]. beige [1766]. bell [1399]. bell-shaped [1399]. bend [1752]. beneficial [1851]. Benefits [2017, 2215]. Bengal [2092]. benign [1849]. Benzene [2438]. bequeathal [1178]. Bernard [792, 515]. Bernoulli [1878]. Best [1001]. Bet [3359]. Better [3041, 865]. Between [3010, 2877, 2951, 3345, 3364, 3262, 3290, 3153, 2508, 2424, 2422, 3606, 3566, 2045, 2594, 2153, 3122, 3077, 3360, 3384, 3530, 2230, 2037, 3380, 3047, 2990, 1647, 2643, 2079, 2847, 2032, 2493, 1484, 457, 493, 1781, 704, 2394, 784, 1529, 1871, 2244, 1619, 558, 95, 276, 196, 1298, 1106, 360, 1376, 45, 1078, 2359, 1330, 2360, 1985, 459, 1277, 211, 1365, 1466, 1859, 1582]. Betweenness [3577]. Bevacizumab [3398]. Beyond [2914, 765]. Bi [1679, 11]. Bi-ionic [11]. Bi-trophic [1679]. BIAcore [2019, 2045]. Bias [3191, 2534, 1565]. Biased [3622, 3129, 1139]. bibliography [760]. bicarbonate [1300]. biconcave [1559]. Bidirectional [2358, 2935, 1539]. bidomain [1488]. Biflagellates [2650]. Bifurcating [699, 1146]. Bifurcation [2927, 158, 205, 286, 3526, 3323, 2458, 569, 3161, 1334, 3229, 376, 2074, 1268, 172, 3603, 2584, 3166, 1412, 430, 1926, 1368, 1292, 1360, 698, 288, 373, 1429]. Bifurcations [2437, 3290, 1804, 2867, 1374, 637, 1612, 1096, 1087]. bilayer [640, 483, 1841]. Bimodal [2249]. Bimolecular [2915, 3067]. Binary [2833, 2566, 2066, 252, 1442, 636, 1405, 422, 423, 819, 805, 858, 1201, 801, 921, 1096, 765, 1544, 1042]. Binding [3193, 2855, 3067, 3147, 3437, 3570, 2184, 3573, 1539, 457, 493, 1903, 844, 817, 599, 696, 598, 267]. Binets [3242]. Binomial [3540, 1010]. bio [277]. bio-environmental [277]. biobarrier [1585]. Biochemical [2175, 2978, 3479, 3547, 2688, 3531, 3493, 3484, 3550, 2604, 3299, 2718, 3558, 3096, 3556, 488, 1810, 1719, 735, 601, 1127, 1128, 1528, 1840]. Biochemistry [2040]. Biocide [2697]. Biocontrol [3169, 3203, 145]. Bioconvection [3058, 1628]. Biodiversity [2721, 1799]. biodynamic [503]. biodynamics [469]. bioeconomic [1172]. Bioeconomics [1118, 1169]. Biofilm [2207, 2128, 2262, 2767, 2601, 2761, 3434, 2880, 3093, 3095, 2697, 3232, 1824, 1274]. Biofilms [3389, 3221, 2682, 2463, 1959, 1870]. biogeography [1157]. Bioinspired [3456]. Biologica [329]. Biological [2140, 3552, 2538, 2388, 2800, 3528, 2573, 2886, 624, 3168, 2344, 3147, 3165, 2433, 2364, 3608, 3505, 2503, 2435, 161, 2303, 3033, 3255, 2603, 2414, 2549, 2064, 984, 3048, 2780, 959, 371, 389, 451, 640, 488, 357, 322, 162, 448, 253, 510, 187, 573, 485, 299, 1329, 1280, 1561, 2660, 1286, 312, 741, 1409, 1060, 126, 1491, 268, 437, 557, 1727, 103, 321, 768, 1818, 170, 224, 277, 276, 454, 6, 386, 1793, 551, 1146, 516, 1611, 1068, 622, 317, 1796, 1950, 130, 305, 352, 374, 774, 208, 507, 1510, 601, 602, 603, 1346, 1345, 846, 1850]. biological [616, 440, 1345]. Biologically [2520]. biologii [124]. biologique [126]. Biologist [2879]. biologists [813]. Biology [3202, 149, 861, 961, 1101, 1255, 2061, 2909, 3106, 2155, 2537, 3307, 3040, 3039, 3296, 3045, 2753, 3516, 2517, 122, 2062, 1385, 1891, 1380, 2, 264, 303, 852, 1858, 3628, 124, 791, 13, 226, 517, 313, 648, 1001, 1920, 2129, 1940, 2222, 2221, 1899, 2035]. Biomass [2772, 654, 695, 775]. Biomat [2850]. BioMath [2730]. Biomathematical [2759, 864]. Biomathematics [125, 961, 1400, 225, 374, 204, 861]. Biomathematiques [787]. Biomechanical [2182]. Biomechanics [3185]. Biomedical [2819, 505]. Biomedicine [2253]. Biomembrane [2494]. Biomembranes [2007]. Biometrics [1400]. Biometry [1930]. Biomodels [3442]. Biomolecular [3318, 740, 795]. Biomolecules [2826]. Biomphalaria [69, 2308]. Bionomic [3164]. Biophysical [2896, 2155, 12]. biophysics [236]. biopolymers [375]. bioreactions [1265]. Bioreactor [2872]. bioreactors [1257, 1107]. Bioregulatory [2545]. Biosensors [3287, 2948]. biosystems [978]. Biotic [2842]. Bipeds [2589]. Biphasic [2677]. birch [1036]. Bird [3111]. Birds [2011, 3060, 1971, 1180]. birhythmicity [1573]. Birkhäuser [1648]. Birth [3479, 3315, 3386, 3430, 107, 691, 722, 318, 712, 521, 1798]. Birthday [1123]. Bistability [2774, 3491, 3268, 407, 3231, 2271, 2590, 3044]. Bistable [2863, 3459, 2780]. Bivalent [2954]. bivalves [1745]. Biventricular [497]. Black [2723, 1728]. black-white [1728]. Bladder [2178, 2317]. blastomere [1942]. blips [1788]. Blocks [3242, 1155]. Blood [165, 2586, 2996, 2966, 2634, 3359, 2056, 3066, 2391, 3143, 2200, 2623, 772, 1257, 1516, 1536, 314, 164, 1037, 56, 550, 826, 820, 915, 1299, 189, 686, 914, 975, 652, 1135, 1751, 496, 466, 1107, 274, 694, 908, 3, 572, 571]. blood-flow [1536]. blood-retina [686]. blooms [1842, 1717, 1873]. Blower [1802]. Bluetongue [2956]. Bobisud [350]. Body [2361, 1684, 3027, 1634, 134, 1778, 14, 1097, 627, 772, 781]. Bogoliubov [333]. BOLD [2623]. Boltzmann [1969, 2975]. Bone [3362, 3009, 2620, 3574, 3204, 15, 1615, 938]. Book [1558, 1272, 949, 1666, 1721, 973, 574, 137, 1577, 1616, 1646, 1402, 1313, 1760, 1702, 1532, 1108, 1158, 1285, 1549, 980, 992, 1043, 1455, 1900, 1636, 1740, 1920, 2129, 1395, 1524, 1195, 1486, 1693, 1774, 1092, 1168, 1635, 1802, 1819, 1540, 1657, 1930, 1626, 1022, 1475, 1061, 1000, 1142, 1052, 1568, 1133, 1711, 1505, 1818, 1074, 1357, 194, 933, 1364, 427, 1596, 2253, 1675, 1647, 1122, 1606, 1254, 1447, 1877, 1420, 1295, 1245, 2222, 2221, 464, 1084, 1219, 1325, 1831, 2143, 1940, 1150, 1847, 1331, 1515, 1759, 1495, 1750, 1730, 1899, 1637, 1586, 897, 1100, 1720, 1648, 1207]. Book [1340, 967, 1411, 1840]. boolean [619, 2288, 2805, 2476, 3268, 1921, 1968, 2995]. Borchers [1831]. Borne [3622, 2141, 3391, 2106, 3212, 3606, 3604, 3277, 3525]. Both [2169]. Bottlenecks [2606]. Bottom [1873]. Bottom-up [1873]. Bound [2992, 1013, 1595]. Boundaries [3194, 2201, 2230, 3286, 1367, 1880]. Boundary [2936, 2050, 3532, 3064, 92, 830, 38, 917]. Bounded [2863, 2184, 2497]. boundedness [606]. bounding [786]. Bounds [229, 118, 132, 239, 646, 860, 919, 960, 1227, 1026]. Bovine [2882, 3029]. Bowel [2827]. Bradford [1900]. Brain [2896, 3162, 3372, 3512, 2522, 2677, 2623, 2280, 3371, 2002, 3571, 1767, 1380, 701, 817, 991, 1433]. Brake [3624]. Brake-Driven [3624]. Branch [3173, 1550, 195]. branched [937, 152]. Branches [2575, 558]. Branching [2332, 2264, 2540, 2321, 2372, 2551, 334, 796, 201, 926, 563, 1294, 849, 415, 534, 568, 797, 677, 588, 730, 805, 858, 195, 610, 653]. BrdUrd [1737]. Breaking [2741, 2974, 1868]. Breast [3496, 2300, 2629, 2823, 3279, 2531, 2672, 1483]. breathing [246]. breeding [1961, 1180]. Bridging [3557]. Brief [2508]. bright [255]. Brine [2426]. British [1123]. Britton [1899]. Brizalina [1054]. Broad [3543]. bronchial [558, 739]. Brownian [1995, 1659, 2560]. Browsing [3228]. Broyden [3597]. brucei [2905]. Brusselator [673]. Bucket [2730]. Buckland [1831]. Buckling [3332]. Bud [2244, 2312, 1860]. Budding [2182]. budgets [928, 1813]. Budworm [2260]. build [1732]. build-up [1732]. Building [2232, 1373, 1354, 3242, 307, 19]. Buildup [2410]. Bulletin [1858, 2035].

Bulmer [1930]. bundling [1928]. burden [1767]. Burgers’ [1491]. burn [1732]. burn-wound [1732]. burster [1743]. bursters [1548, 1598, 1316, 1683]. Bursting [2313, 2779, 2848, 3594, 1573, 2057, 2226, 2561, 1353, 1892, 779, 1761, 1642]. bush [1451]. butterflies [1212]. Butterfly [3431, 852].

C [3010, 2766, 1900, 2714, 2143, 1648, 1567, 3504, 2096, 3578]. C-MAPK [1567]. C. [1865]. Ca [1328, 1434, 1562, 2392, 1942, 1573, 1941, 2282, 1667]. CA1 [2524]. CA3 [1376, 1514]. Cable [1550, 365, 809, 1131, 1376, 1770]. Cables [3323, 1194]. caderas [2279]. Caenorhabditis [2974, 3423]. calcification [1525]. Calcium [2495, 2614, 293, 2081, 2691, 1237, 2088, 3249, 1892, 1206, 1491, 1834, 1580, 938, 1165]. calcium-based [1892]. calcium-induced [1237, 1165]. calculated [514]. Calculating [3271, 2152, 1967, 1021]. Calculation [2924, 1618]. calculus [707, 1067, 576]. Calibration [3474, 3397, 3234, 2306]. Cambridge [1646, 1920, 2129, 1877, 2143, 1847]. Came [2039]. cAMP [3625, 2933]. Campaigns [3593]. Campbell [609]. Can [2346, 2257, 3624, 3324, 3224, 3206, 2982, 2941, 3114, 3567, 3351, 3437, 2499, 3216, 2268, 3343, 3495, 3163, 1685, 1742, 3393, 1232, 779, 3136, 3013, 1697, 1656, 2028, 1956, 1096]. Canada [1776]. Canalyzing [2658]. Cancer [3362, 3105, 3496, 2169, 3364, 2436, 3420, 3579, 2178, 2317, 2001, 2752, 3448, 3153, 2508, 2300, 2629, 3524, 3006, 3365, 2769, 3521, 3384, 3377, 3018, 3530, 2530, 2823, 2063, 3136, 3380, 3279, 2686, 3151, 2531, 2646, 2986, 3208, 3532, 2672, 2985, 872, 3375, 3595, 2173, 2947, 3254, 3376, 3080, 1633, 1849, 144, 346, 390, 1862, 876, 865, 193, 396]. cancerous [954]. Candidate [2978, 2420, 2447]. Canine [2081]. Cannibalism [207, 769, 1464, 1352, 127, 737, 350, 1910]. Cannibalistic [3166]. canonical [1054]. Capability [2429]. Capacity [3228, 3195, 2286, 1320, 948, 248, 2026, 1337]. Capillaries [2729]. capillarized [768]. Capillary [728, 828, 2384, 2654, 2586, 2996, 2966, 2500, 2391, 487, 919, 960, 920, 946, 1703, 1741, 972, 908]. capillary-tissue [487, 972]. Caps [2867]. Capsid [2539, 3508]. Capture [2668, 2613, 2520, 2655, 2180, 931]. Captures [3204]. Capturing [3377]. Capybaras [2279]. carbohydrates [1336]. Carbon [2004]. Carcinogenesis [3346, 3209, 867, 289]. Carcinoma [3348, 3366, 3401, 1091]. Cardiac [2824, 2247, 3451, 2895, 2475, 2638, 2114, 2195, 56, 1918, 1437, 1155, 705, 128, 391, 1811, 1580, 785, 37, 522, 1165]. Cardiovascular [3065, 3099, 2931]. Care [2630, 2367]. Cargo [1995]. Carlo [3552, 784, 190, 554, 2598, 1211]. Carlos [1802]. Carriage [3081, 3289]. Carrier [3147]. Carrier-Mediated [3147]. Carrying [2730]. Cartilage [2869]. Cascade [2059]. Cascades [3265, 2640]. cascading [1047]. Case [2140, 2318, 2408, 3528, 2840, 2669, 3106, 2228, 3436, 3284, 3569, 1571, 1629, 1678, 870, 408, 368, 404, 419, 613, 1786, 511, 495, 1267, 765]. Case-Study [2840]. Cases [3145, 3065, 1048]. Caspase [3341, 2863]. Caspase- [3341].

Castillo [1802]. Catalog [2957]. catalysed [1339]. Catalysis [2633, 761, 1339, 1333, 703]. catalysts [185]. Catalytic [2743, 1735, 1319, 366]. catalytically [383]. Catastrophe [417, 393, 382, 573, 342, 485, 715]. catastrophes [1221, 311]. catastrophy [852]. catatonic [65, 257]. Catch [3624, 1174]. Categorical [707, 708]. categories [643, 614]. category [1718]. catenary [725]. Catherine [1900]. Cation [457, 493, 483]. cation-phospholipid-induced [483]. cations [457, 493]. Cats [1971]. CatSper [2495]. Cattle [2905, 3029]. Cause [3136, 1981, 1728]. Caused [2408, 474, 1663]. Causes [3172]. Causing [2944]. cautious [313]. Caveolae [2030]. cavities [1213]. Cayley [1276]. CCN1 [3052]. CD [3341]. CD4 [3303, 1926, 3054, 2810]. CD8 [2785, 3544]. Cdc42 [3498]. Celebration [1123]. Cell [2103, 2509, 2776, 2877, 2998, 1916, 1931, 562, 681, 3455, 2446, 2511, 3564, 3496, 2169, 2661, 3191, 3420, 2485, 2799, 3273, 2835, 1991, 2001, 2770, 2535, 2722, 2825, 2742, 2785, 2452, 806, 869, 2369, 2634, 3153, 3298, 2300, 2629, 2716, 2707, 2642, 2908, 2807, 2366, 2903, 2155, 3566, 2045, 2167, 3401, 3218, 2581, 2898, 2451, 2442, 3374, 3136, 3533, 2204, 2192, 2224, 3140, 3574, 3293, 2969, 2950, 2374, 2991, 3279, 3407, 2000, 3039, 2759, 3349, 1253, 2494, 3037, 2675, 2937, 3259, 1494, 3544, 3333, 2375, 3057, 2943, 2857, 3185, 2646, 2352, 2431, 2443, 2691, 3532, 2672]. Cell [2360, 2703, 2125, 2754, 2145, 2294, 2745, 2947, 2381, 3376, 3314, 2413, 2237, 3232, 646, 173, 1975, 1958, 676, 1773, 1139, 346, 1306, 163, 1548, 1598, 2071, 1037, 969, 1527, 1229, 1371, 1477, 1478, 1090, 682, 1265, 669, 3565, 902, 1446, 1493, 1314, 1787, 309, 237, 1271, 1317, 1535, 1591, 13, 243, 495, 1785, 1250, 339, 545, 1310, 1587, 1829, 1592, 1375, 1584, 1040, 1242, 1698, 572, 1189, 1601, 1888, 598, 1939, 957]. Cell-Based [2825, 3564, 2451]. Cell-Driven [2947]. Cell-Mediated [2991]. cell-target [1139]. Cells [3105, 3364, 2799, 3091, 3303, 3392, 2716, 3361, 2769, 3054, 3360, 3384, 3565, 3238, 2442, 2988, 2192, 2336, 2933, 2397, 2827, 3380, 3407, 2345, 2686, 2457, 3503, 2233, 3076, 2614, 2081, 2970, 3208, 2110, 2226, 2689, 2341, 2211, 2088, 3476, 3341, 2720, 2846, 3249, 1633, 421, 458, 1257, 83, 1542, 866, 1567, 1766, 35, 1672, 1926, 1839, 1369, 1686, 1491, 821, 1982, 1238, 923, 1472, 1362, 1837, 1344, 1860, 742, 1696, 652, 1612, 1677, 1209, 119, 1064, 1284, 1587, 1316, 1283, 1107, 1487, 1578, 908, 571, 1165, 703, 1700, 3625, 2159]. Cellular [2571, 3451, 2330, 1579, 2066, 2812, 2902, 2025, 3394, 2578, 3360, 2679, 2615, 3221, 3037, 2614, 3370, 1077, 389, 530, 856, 1763, 1924, 964, 1500, 289, 297, 1242, 1459, 382, 2157]. Center [125, 2387, 1837, 1613]. Centering [2041]. Central [149, 3407, 1776, 1326]. Centrality [3577]. Centrifugal [1042]. Centrifugal-order [1042]. cephalopods [955]. Cereal [2622]. Cerebellar [2280]. Cerebellum [3140]. Cerebral [2525, 2526]. cerevisiae [3069]. Certain [2844, 238, 733, 587, 19]. Cervical [3105]. CFSE [2511, 2038, 2397, 2652]. CFSE-Labeling [2652]. Chagas [3514]. Chain [3555, 3561, 3484, 2293, 1826, 658, 1429, 1679, 266, 766, 1941, 1442]. Chains [3417, 3560, 2764, 439, 1529, 1508, 680, 937, 879]. Challenge [3406]. Challenges [2122]. Chamber [1992, 324, 3]. Chambers [3095]. Chance [2925]. Change [3177, 2934, 3186, 2016, 2597, 3101, 2176, 479]. Changes [3520, 3444, 2381, 376, 102, 839, 687, 483, 1886, 161, 1425]. changing [1388, 1232, 1231]. Channel [3496, 2495, 2293, 714, 1393, 9, 105]. channelling [1361]. Channelrhodopsin [2860]. Channelrhodopsin-2 [2860]. Channels [3170, 3305, 2058, 580, 1941, 1580]. Chaos [3272, 1382, 265, 1260, 1461, 1383, 1919, 1573, 481, 1221, 985, 1211, 1605]. Chaotic [2883, 2556, 833, 2434, 807, 3437, 1383, 1796]. Chaparral [2866]. character [1405]. characteristic [1270, 1345]. Characteristics [3153, 2860, 1091, 764, 578, 1639, 583, 1478]. Characterization [998, 3043, 2845, 2806, 2164, 3558, 3431, 3318, 906, 181]. Characterize [3543, 1704]. Characterizing [1903, 774, 3148]. Characters [2566, 252, 1627]. Charge [3170, 947, 304, 339]. Charged [3237, 282]. Chavez [1802]. Checkpoint [1954, 3244]. Cheliped [2097]. Chemical [2496, 2329, 2837, 3131, 2147, 3497, 3491, 2358, 2117, 406, 2326, 2922, 2305, 2971, 867, 3295, 2683, 2477, 2864, 459, 3408, 1069, 2211, 2995, 2172, 2306, 420, 465, 188, 376, 430, 471, 33, 479, 1796, 130, 289, 1847, 612, 3627, 205, 48]. Chemically [2621, 453]. Chemo [3595]. chemoattractant [1327]. Chemokine [2770]. chemokinesis [1283]. Chemomechanical [2684]. Chemoprophylaxis [2277]. chemosensory [1327]. Chemostat [2887, 2239, 806, 1529, 2335]. Chemotactic [1808, 2735, 843, 1147, 1347, 713, 1496, 1632, 359, 692, 727]. Chemotaxis [3210, 3350, 2893, 2127, 2374, 2375, 1283, 2294, 2295, 2072, 2735, 2579, 2975, 1916, 1931, 1645, 2934, 1780, 1146, 1164, 645]. Chemotherapeutic [1448, 346]. Chemotherapy [3317, 2659, 2991, 2646, 3370, 144, 346, 1471, 1948, 1401, 1753, 1404, 269, 872, 396]. Cherry [2845]. Cheyne [3027]. Chicago [1075]. chicken [278]. chickenpox [1600]. chicks [1180]. chief [226]. Child [2187, 3584, 1861]. Childhood [3098]. Children [3457]. chimiques [48, 33]. China [2015, 3486, 2690, 2771, 3421]. Chinese [3486]. chiral [1315]. Chlamydia [2840, 1996]. Chlamydial [2862, 1854]. chloride [309]. Chloroplast [3076]. Choice [3311, 3118, 3368, 1954]. choices [1812]. Cholera [3419, 2958, 2498, 2814, 2631, 2738, 3282]. Cholesterol [3042, 1103]. Choosing [3039]. choroid [14]. choroidal [525]. Chromatin [2855]. Chromium [2830]. Chromosimal [104]. chromosimique [104]. Chromosomal [2941]. chromosome [220, 1453]. Chromosomes [2613, 1154]. Chronic [3364, 2486, 2634, 2492, 2258, 2314, 2553, 2352]. chronically [1981]. Chronification [3614]. Chyuan [537]. Chyuan-Yih [537]. Cicadas [3492]. Cichlid [3103]. Cilia [2161]. Ciliary [2107, 2158, 53, 14, 1660]. ciliate [1397, 1398]. circadian [1462, 1341, 1682, 1224, 360, 355]. Circuit [2805, 2597, 1865, 871]. Circuitry [3373]. circuits [259]. Circular [3005, 3031, 2908, 2124, 1504, 1748, 49, 406, 1438]. Circulating [2235]. circulation [525, 3]. circulatory [30, 1440, 1566]. Circumcision [2208]. circumnutating [1315]. Cities [3606, 3022]. City [2740, 2078]. Clark [1123, 1175]. Class [3154, 3543, 3319, 2386, 2572, 3531, 2601, 2784, 2497, 2532, 169, 577, 348, 699, 441, 883]. Class-Dependent [3154]. Classes [3075, 3119, 1468]. classic [1118]. Classical [3520, 1552]. Classics [3628]. Classification [877, 850, 2978, 2550, 2050, 1353, 1814, 2062, 1553, 220, 1359, 1987]. Clausius [152]. Clearance [2515, 2111, 3595, 919, 960, 1048]. cleavage [1752]. cleft [1311]. Climate [2346, 3177, 3186, 2078, 1880, 3101]. Clinical [841, 3046, 1849, 1472, 168]. Cliques [2600]. clock [161, 735]. clocks [355]. clonable [1710, 1083]. Clonal [2101, 1792, 2073]. Clone [3190, 1154, 1707]. clones [1205]. cloning [1707]. closed [1428, 1429, 1831, 879, 1622, 1507]. Closing [2236]. Clostridium [3288, 3201]. Closure [2309, 3036, 2687, 3281, 2181, 1638, 1960]. Closures [2701]. clumped [769]. Cluster [3362]. Clustered [2637, 1484]. Clustering [3424, 2648, 3291, 2272, 2778, 2801, 2198, 1011, 1845, 556, 1856]. Clusters [2560, 1652, 1791]. clutch [1184]. clutch-size [1184]. CMF [2531]. CML [3538]. CNP [1864]. Co [3280, 3039, 2734, 3338]. Co-colonization [3280]. Co-culture [3039]. Co-existence [2734]. Co-infection [3338]. Coalescence [2715]. Coalescent [3321, 3469, 2197, 3467, 3466, 2407]. Coarse [2768]. Coasting [2215]. coaxial [62]. coccidiosis [1891]. cod [1910]. Code [3534, 2456, 2104, 2568, 3425, 2124, 1504, 1988, 1879, 1895, 1153, 1921, 1968, 723, 1889]. Coded [2124]. Codes [2834, 3269, 3048, 1035]. Coding [1407, 1484, 1504, 354, 750, 1705, 1125, 64, 1008]. CoDiS [3077]. Codon [3237, 1468, 1008, 1273]. coefficient [129, 729]. Coefficients [2003, 569, 1210, 1619, 392, 279]. Coexistence [2887, 720, 3160, 2395, 3536, 2904, 2651, 2581, 3116, 1872, 3500, 2502, 410, 2726, 1779, 1983]. Coexpression [2871]. Cognate [3237]. Cognition [2518]. Coherence [2057]. Coherent [2521]. cohort [1258, 976]. Coil [1995, 1297]. Coiled [1995]. Coinfection [2402, 2339]. Cold [3069]. coli [2318, 2369, 2560, 2934, 3350, 1397, 1398, 3268, 3434, 2271, 2891]. Colin [1123]. Collagen [3009, 2727, 1489, 1501]. Collapse [3619, 3286, 3019]. collapsible [1412, 1432]. collapsible-tube [1432]. Collaterally [3390]. collections [1612]. Collective [2877, 3191, 3566, 3349, 1812]. Collectives [2163]. Colon [3407, 3376]. Colonic [3333]. Colonies [3141, 3246, 1223]. Colonization [2112, 2587, 1439, 3280, 1787]. Colony [3260, 3619, 3623, 1708]. Colorectal [2182]. Colored [2407, 1349]. Columbia [1123]. Column [1993, 232]. Comb [2174]. Combat [3142]. Combination [2215, 3003, 1788, 1506]. Combinations [2840, 3370]. combinatorial [1570, 1719]. Combinatorics [3334, 2225]. combinatory [1454]. Combined [3163, 2985, 1203, 1236, 1323, 1410, 657]. Combining [2705, 3289, 3561]. Comma [3269]. Comma-Free [3269]. Commensal [2678]. Commensalism [733]. Comment [775, 516, 340, 350]. Comments [662, 3352, 790, 812, 832, 403, 522]. Commercial [2016, 1120]. Common [2851, 2124, 581, 590, 1119]. common-property [1119]. Commons [2788]. Communication [3218, 2043]. Communities [2721, 2818, 2724, 2734, 1469, 800, 1494, 1221, 1031, 698, 846, 221].

Community [3038, 3288, 2772, 1397, 1398, 1457]. Community-Structured [3038]. Comorosan [184]. compaction [856, 964]. Comparaison [129]. Comparative [2906, 3198, 3264, 1555, 199, 2143, 1322, 1251, 1853, 846, 1483, 19]. compare [1790]. compared [1307]. Comparing [1694, 2772]. Comparison [2915, 1258, 1915, 624, 591, 2013, 3271, 2957, 2598, 3556, 1098, 2293, 3498, 3084, 1082, 959, 1958, 444, 1190, 1396, 1321, 172, 660, 963, 196, 1191, 747, 1243, 129, 1975]. Comparisons [2374, 1202, 1858, 753]. compartiment [1107]. Compartment [2999, 3075, 23, 523, 281, 2196, 8, 487, 110, 167, 169, 241, 357, 834, 546, 408, 556, 549, 604, 117, 154, 515, 25, 930]. Compartment-Based [2999]. Compartmental [3424, 1210, 659, 408, 556, 2372, 626, 167, 169, 241, 390, 425, 1789, 476, 235, 792, 924, 799, 1820, 217, 416, 883, 534, 568, 171, 234, 527, 661]. compartments [523, 626, 110, 94, 117, 154, 25, 325, 290, 341]. Compatibility [3233, 252]. Compensation [2577, 3249]. Compensatory [3620, 1912]. Competent [3604, 1859]. Competing [3585, 2092, 2689, 720, 1056, 1557, 298]. Competition [2721, 2704, 2951, 3435, 2202, 1439, 3175, 457, 2865, 2587, 2045, 3417, 3213, 2937, 2622, 1862, 2734, 2243, 3023, 2168, 3248, 3412, 2335, 3064, 493, 1579, 1187, 490, 892, 1517, 1799, 418, 1793, 1830, 638, 1241, 1880, 1277, 1492, 524, 2469, 2302]. Competition-Colonization [2587]. Competitive [2766, 2115, 3122, 2696, 2591, 1779, 2113, 2762, 1917, 3063, 100, 338, 1430, 1404, 1622, 1514, 1663, 186, 1593]. competitiveness [1176]. Competitor [2377]. Competitors [2502]. Complementary [365, 450, 533, 1504, 844]. Complete [2486, 458, 724]. completely [585, 476]. Complex [3418, 2853, 2121, 2409, 2673, 2543, 2946, 1749, 2573, 1192, 2747, 3188, 1293, 1373, 1020, 899, 1828, 1529, 1627, 2253]. Complexity [1279, 2258, 1647, 693, 929, 1588, 570, 757, 867]. Compliance [2194, 22]. Component [3287, 3471, 1679, 361]. Components [3165, 1184, 75, 152]. Composite [3483, 152]. Composition [3588, 3369, 1097]. compound [1202]. Compounds [3010, 978]. compressed [1752]. Compression [3006, 2822]. computability [131]. Computation [2433, 674, 1434, 808, 594]. Computational [3273, 2681, 3049, 2392, 929, 2993, 1013, 3443, 3348, 3055, 701, 2521, 2933, 3625, 1595, 2256, 2291, 2222, 3612, 2938, 2860, 3375, 2338, 3266, 965, 1514, 2822, 2278, 1841, 1840]. computations [660]. Computed [3309, 685]. Computer [125, 3159, 1139, 739, 2221, 1077, 874, 873, 245, 1886, 504]. computer-aided [504]. Computing [3511, 510, 752]. concave [963]. concealed [1771]. Concentrating [2135, 2065, 2429, 1302, 1815, 1807, 1152, 1303]. Concentration [1897, 729, 2975, 806, 1347, 1317, 652, 1305, 1631, 339]. Concentration-dependent [729]. Concentrations [3596, 2932, 246, 309, 1827]. Concept [170, 454, 1270, 757, 1323, 224, 627, 723, 1345]. conception [107]. concepts [791, 478]. Conceptual [2344]. concerted [1619]. Concomitant [1370].

Condition [3520, 2864, 2006, 72, 645, 112, 1127, 1128]. Conditional [2404, 2889, 1469, 674]. conditioning [1552]. Conditions [2799, 2191, 2090, 2353, 1089, 3276, 2434, 2243, 2796, 2256, 2291, 2497, 689, 319, 651, 1717, 438, 1709, 624, 266, 272, 1786, 327, 774, 1269, 560]. Conductance [3594, 1124, 687, 398, 835, 522]. Conductance-Based [3594]. Conducted [3012]. conducting [322, 514]. Conduction [2638, 2510, 533, 860, 1065, 1771, 477, 1770]. Conductivities [2400]. conductivity [830]. conductor [1550, 1634]. Cones [3160, 1059]. Conference [138, 301, 877, 861, 941, 961, 1032, 1101, 1159, 876, 505, 681, 1123]. Configurations [3465, 2540]. Confined [2916, 2917, 3445, 1834]. Confinement [2997]. Conflicting [1471, 2028]. conflicts [1629]. conformation [1278]. conformational [839]. Conifer [2025, 2157]. conjugacy [478]. Conjugate [2610]. Conjugation [3114]. Conjunctive [2476]. Connecting [3100]. connection [1374, 1808, 211]. connectivities [1533, 423]. Connectivity [2340, 200, 1104]. conscious [991]. Consensus [1217, 526, 1144, 1256, 771, 749, 794]. Consequence [3080]. Consequences [3538, 3186, 2929, 2362, 2243, 2359, 297, 2984, 2484, 1476]. Conservation [2316, 2129, 1522, 2603, 1678, 1170]. conservative [327]. conserved [7]. Considerations [2425, 548, 441]. considered [33, 1096]. considerées [33]. Considering [2469]. Consistency [1088]. Consistently [2359]. consisting [880]. constancy [254]. Constant [2743, 378, 210, 366]. Constants [2688, 2120, 1194, 1680]. Constrained [2920, 2702, 1246, 901, 900]. constraint [429]. Constraints [3613, 1423, 3487, 901, 437, 900]. constrictions [82, 139]. Constructing [2309, 2932, 1913]. Construction [3004, 2174, 3611, 168, 191, 2359, 541]. Consumer [3166, 2427, 3120, 1041]. Consumers [2788, 2044, 1864]. Consumption [3018, 2724, 385, 1021]. Contact [2571, 3200, 3334, 3566, 2365, 3057, 1996, 1655]. Contact-dependent [1996]. Contacts [3334, 389]. Contain [3144]. containing [866]. Contamination [3619, 1712]. Content [3449, 3052, 2713, 2964, 188, 1988, 469]. Contest [3295]. Contests [2536]. Context [1015, 2524, 713]. Context-dependent [1015]. Context-Sensitive [2524]. Continuing [125]. Continuous [1534, 3260, 2501, 2394, 2376, 72, 2564, 1470, 2255, 2013, 2415, 3560, 680, 2348, 2375, 3302, 2753, 927, 1145, 1012, 1749, 666, 235, 1294, 1317, 703]. Continuous-time [680]. Continuously [3102]. Continuum [2207, 2128, 2611, 3475, 3448, 3377, 3349, 2733, 2664, 3082, 3179, 3026, 2064, 2625, 2029, 1839, 1500]. Contraceptives [3227]. Contractile [3342, 543, 433]. Contractility [2822, 385]. Contracting [3158]. Contraction [2645, 1829, 1893]. contractions [1560, 349]. Contribute [2868]. Contribution [3605, 1883, 1134]. Contributions [1369, 1111, 1112, 1113, 3547]. Control [2906, 3433, 3202, 2140, 2596, 2671, 3201, 2458, 3267, 2890, 3593, 2388, 2001, 2052, 2940, 3091, 3458, 3515, 2162, 2216, 3241, 2929, 2117, 3272, 3606, 2883, 2839, 2798, 3352, 2610, 2905, 3601, 3374, 2958, 1864, 2609, 2364, 2838, 3002, 2503, 3351, 2435, 2208, 2287, 3413, 2981, 3436, 2876, 3143, 2529, 2200, 2603, 2414, 2260, 3188, 3545, 2242, 2737, 3376, 3180, 2308, 2152, 2931, 1421, 1257, 1972, 179, 1783, 763, 1942, 98, 821, 1428, 875, 6, 1351, 1271, 1348, 1449, 168, 1547, 970, 119, 503, 621, 297, 292, 445, 1620, 385, 269, 625, 872, 1346, 1127, 1128, 332, 1888, 1582]. control [1893]. Controllability [1673]. Controlled [677, 2677, 429]. Controlling [1461, 3366, 3410, 1605, 259]. Controls [2702, 2840, 3320]. Convection [2019, 2102, 291, 213]. convection-diffusion [291]. Convective [2879, 582, 101, 552]. convective-diffusion [101]. Convergence [3008, 1634, 3149, 2651, 892, 2051, 3429, 3263]. Convergent [2889]. Convict [3103]. Convolution [2865]. Cooperation [3610, 2618, 3023, 3000, 3128]. Cooperative [264, 2935, 844, 16, 80]. Cooperativity [3393, 1127]. Cooperators [2618]. Coordinate [2174]. coordination [1262]. Copies [2777]. Coping [1383]. copper [328]. Coral [3124]. Cord [2351, 3099]. cords [1816]. Core [3540]. Coreceptor [2607, 2096]. Cores [2759]. cornea [198, 1575, 910]. Corneal [2451]. Corneum [2180]. coronary [1886, 1338]. Coronavirus [2420, 2447]. Corralling [2783]. correct [1565]. Correction [3626, 3427, 3485, 3428, 3625, 2941, 3337, 696]. correlated [901, 900]. correlates [1248]. Correlation [2951, 2468, 1125, 1162, 459]. Correlations [1344, 671]. Correspondence [2732, 3055, 1468]. Corresponding [2394, 3429]. Cortex [1993, 2525, 2526]. Cortical [2992, 3179, 2523, 2625, 1497, 965, 1081]. cortical-like [965, 1081]. Cost [2817, 3351, 2849, 2981]. Cost-Effectiveness [2981]. costs [888, 1416]. Could [3615]. Counter [2087]. Countercurrent [46, 1302, 29, 18]. counterdiffusion [294]. Counterintuitive [3360]. Countries [2963, 3300, 3077]. Counts [2634, 1608]. couple [483, 975]. Coupled [3433, 2389, 1951, 3447, 2621, 2349, 2457, 2411, 3573, 2195, 2270, 783, 1462, 1374, 1817, 1548, 1598, 539, 1561, 502, 518, 177, 383, 480, 1982, 1855, 272, 413, 1387, 1941, 1216, 360, 1612, 1901, 33, 1316]. couplées [33]. Coupling [3156, 2782, 2475, 3517, 779, 2659, 2723, 62, 180, 130, 1376, 1683]. Couplings [3551]. Course [1075, 2387, 1136, 1766, 1819]. Couvet [2129]. Cover [3403]. Cover-Encodings [3403]. Coverage [2655, 1997]. Covering [2647]. Cowan [507]. Crab [3583, 2097]. Craters [3026]. crayfish [1682, 1224]. Crazy [3449]. creativity [34]. Crest [2661]. Crisis [3441]. Criteria [3016, 3516, 2546, 1472, 391, 606]. criterion [996, 1619]. Critical [993, 1717, 2789, 2249, 3385, 3108, 459, 3556, 3376, 1886, 249, 865]. critique [1017, 373]. crocodilians [1167]. croisements [129]. Cronin [596, 663]. Crop [3582, 2165, 1689]. Crops [2388]. Cross [3435, 3345, 2692, 2693, 3360, 306, 377, 2794, 2441, 3093, 2656, 1569, 9, 1399]. Cross-Diffusion [3093]. Cross-Feeding [2794]. Cross-Immunity [3345]. cross-linking [1399]. Cross-reactive [2441]. cross-regulation [1569]. Cross-sectional [306, 377, 9]. Cross-Strain [2692, 2693]. Cross-Talk [3360]. Cross-Talking [2656]. Crossflow [2681]. Crossing [1649]. crossings [129]. Crossover [2250, 1728, 1754, 1775]. Crosstalk [2183]. Crowded [3304]. Crowding [3430]. Crustacea [1443]. cruzi [2812]. Cryoprotective [2702]. Crypt [3332, 2182]. Crypts [3407, 3333, 1222]. Crystal [2764]. crystallography [31]. CSF [2752, 1767, 2200]. CSF/ [3193]. CSTR [2682]. CTL [2444, 2715, 2663, 2582]. cube [1468]. Cubic [3331, 447]. cubic-polynomial [447]. cuckoo [1769]. Culling [3203, 2484]. culmination [1736]. Cultural [3426]. Culture [2898, 1308, 2810, 3093, 3444, 3039, 1317, 598, 703]. Cultures [2348, 1145]. cumulative [1553, 1618]. Cure [3538]. Current [3598, 3184, 3443, 3444, 3255, 2012, 2087, 1065, 1811, 1925, 1378, 522]. Currents [2744, 436, 802, 1706, 1488, 477]. Curvature [2139, 3478, 3437, 101]. Curvature-Sensitive [3437]. Curve [3543, 2653, 1897]. curved [836, 52, 78, 261]. Curves [3324, 1903, 41, 1289, 1473, 1786, 1506, 1048, 1379]. Cut [2208, 1607]. cut-off [1607]. cutoff [1222]. Cuttrees [730]. CXCL8 [2227]. cyanide [835]. cyanide-treated [835]. Cycle [681, 3062, 2629, 2807, 2224, 2809, 2982, 3259, 2646, 2206, 2133, 2672, 1440, 2145, 3539, 651, 676, 346, 1783, 1259, 1090, 1692, 1428, 1591, 422, 423, 772]. Cycle-time [1440]. Cycles [2552, 2565, 420, 465, 1509, 156, 1857, 2083, 312, 544, 1080, 1173, 1656, 1663]. Cyclic [2721, 3156, 2556, 2530, 2872, 3216, 2382, 2925, 1542, 2043, 762, 1938]. Cyclin [2024]. cycling [330, 1664, 18]. cylindrical [1599, 1452, 52, 78, 388, 971]. Cytochrome [2598]. cytogel [1252]. Cytokinesis [3342]. Cytokinin [2024]. cytometric [1975, 1958, 969]. Cytometry [2302]. Cyton [2397]. Cytosolic [3082, 1763]. Cytotoxic [2642, 2532, 3254, 1205, 1893]. cytotoxicants [657].

D [1900, 2129, 2025, 2157, 1831, 1940, 3537, 2835, 2024, 2564, 2379, 2127, 3221, 3082, 3571, 3128]. Dairy [2900]. Damage [2105, 3370, 1408, 1419, 1809]. Damped [488]. Dan [3547]. dans [126]. Danziger [65, 215]. Daphnia [2118, 1832, 3302]. Daphnicle [1832]. Darby [2081]. D’Arcy [116, 203]. Darwinian [160, 1390, 1641, 1658]. Data [2668, 2119, 2511, 3397, 2940, 3069, 2038, 2993, 3467, 2829, 3494, 2655, 3098, 3527, 2224, 3213, 3217, 2350, 2930, 2652, 3347, 2499, 3519, 3558, 2221, 2143, 2050, 2932, 2841, 3048, 3516, 2413, 1077, 1154, 1975, 1958, 328, 336, 1706, 750, 1383, 969, 1846, 901, 1049, 1431, 1619, 247, 214, 760, 1727, 886, 1363, 900, 660, 1984, 1405, 1545, 1597, 1565]. Data-Based [2350]. Data-Driven [3519, 3558]. database [751]. dataresolution [247]. Date [3315]. Day [3181]. Day-to-Day [3181]. Dayhoff [746, 758]. days [15]. Dead [2607]. Dead-End [2607]. Deakin [573]. Dealing [1998]. Death [3191, 3479, 2066, 2038, 3395, 2224, 3430, 3341, 1773, 379, 691, 722, 472, 986, 712, 474, 521]. Death-Movement [3430]. debris [969]. debt [1967]. Decades [2963]. Decay [3149, 133, 931]. decaying [195]. deciduous [928]. decision [259]. decision-making [259]. decisions [1181]. Decomposition [2519, 1627]. decompression [1810]. Deconstruction [2807]. Decontamination [3201]. decrease [1956]. decrement [613]. Dedifferentiation [3151]. deduced [31]. Deduction [1289, 1379]. deductions [1921]. Deductive [313]. Deer [2234, 918, 1943]. Default [2523]. Defect [3065]. Defectors [2618]. Defence [2137, 882, 988, 1480, 1769]. Defense [3581]. Deficiency [2496, 3196, 3511, 3493]. Deficiency-Based [3493]. deficit [1967]. Defined [2673, 2244]. defining [1416, 572]. Definition [3225]. deformable [102, 101]. deformation [1213, 236, 196, 1335]. deformations [577, 1885]. Degeneracy [2383]. Degeneracy-Driven [2383]. degenerate [1801]. Degradable [2869]. Degradation [3448, 2602, 3053]. degree [2051]. degrees [1230, 75]. dehydrogenase [267]. Delay [2241, 2315, 3383, 2741, 2126, 2817, 2465, 3258, 2479, 2892, 3118, 3500, 3127, 3123, 3539, 1541, 397, 802, 518, 1050, 1820, 689]. delay-differential [1050]. Delayed [3392, 3238, 3388, 2602, 2274, 2098, 1175, 881]. Delays [3323, 1994, 2504, 2615, 2505, 3510, 2804, 2803, 3396, 987, 927, 1247, 1373, 502, 655, 567, 912]. Deleted [3584]. Deletion [2713]. Delivery [3454, 3118, 3570, 920, 1433]. demand [1967]. demand/deficit/debt [1967]. demi [129]. Demographic [2467, 2187]. demography [1117, 1196]. demolition [616]. Dendritic [2001, 2372, 1138, 906, 1761, 1131, 1209, 1407]. Dendrograms [2648, 1640]. Dendron [3049]. Deneubourg [1648]. Dengue [2906, 3621, 2191, 2090, 3142, 3181, 2798, 3097, 2930, 3180, 3421]. Denise [1802]. Denominator [3441]. densities [1481]. Density [2577, 3580, 3626, 2201, 3594, 2923, 2949, 2903, 2855, 3336, 2747, 1413, 3000, 2010, 1366, 2327, 1617, 1187, 379, 1200, 1450, 1872, 304, 1674, 748, 1688, 1440, 1545, 1597, 1663]. Density-Compensation [2577]. Density-Dependent [3580, 3626, 2949, 2747, 3000, 1413, 2010, 1617, 379, 1200, 1450, 1872, 1688]. deoxyglucose [1021]. dependant [1849]. Dependence [2528, 3336, 2223, 3315, 3544, 3355, 835, 2708, 640, 1824, 599, 696, 1269]. dependency [285]. Dependent [3551, 2355, 3480, 3191, 3580, 3626, 2201, 2600, 3154, 3292, 2949, 2024, 2347, 3122, 2696, 2194, 3425, 3550, 2597, 2838, 2747, 3000, 3600, 2623, 3344, 2372, 3243, 3396, 2792, 1867, 1617, 793, 719, 110, 167, 169, 241, 1817, 334, 1334, 1187, 1676, 425, 379, 1261, 1200, 999, 691, 722, 1450, 1551, 392, 1872, 1089, 1015, 729, 970, 318, 1413, 337, 1622, 1587, 1688, 2010, 1627, 325, 1300, 858, 1866, 1996, 1663, 979, 1582]. depending [410]. Depends [3118, 2675]. Depensatory [2962]. depletion [793, 532, 633]. depolarization [770]. Depolarized [802]. Deposition [2452, 237, 442, 353, 79, 475, 361]. Depression [3179, 2625, 129, 129]. Depth [2658, 1824]. Derek [2221]. Derivation [187, 2473, 2108, 2725, 3034, 1372, 513, 956]. Derivative [2181, 6]. Derived [2704, 2024, 3499, 3464, 2235]. Deriving [2400]. Dermal [2687, 1489, 1135]. descending [46, 1152]. Describe [3213, 1218]. Described [3484, 2794, 1317, 1290]. Describing [2191, 2090, 2874, 1805, 418, 215, 954, 412, 1293]. Description [2967, 1036, 1393, 3430, 3456, 1434, 548, 1065, 441, 2067, 422]. Descriptions [2446]. descriptor [1941]. desensitization [1594]. Design [2670, 2902, 2599, 1791, 937, 1520]. Designed [3206]. Designing [1533]. Desirable [3526]. Destroy [2518]. destructor [3070, 3246]. Detached [2192]. detachment [1599]. Detailed [2543, 1283]. Detecting [3272, 2884]. Detection [3490, 308, 2934, 2075, 4, 159, 592]. Detections [2342]. Determinant [3376]. Determinants [2011, 2570]. Determination [2742, 1194, 1437, 2930, 1575, 599, 696, 2627, 2171, 56, 72, 1103, 479, 1911, 910, 382]. Determinative [3165]. determine [1594]. Determined [2855, 1871, 1698]. determines [1837]. Determining [3144, 2538, 2283, 2919, 759, 829, 1500, 375, 1520]. determinism [1467]. Deterministic [2394, 2572, 3460, 2594, 2745, 3498, 1947, 1090, 60, 1551, 362, 1605, 1937, 524]. deterrence [635]. Detrain [1648]. deuterated [1738]. Develop [3357]. Developing [2611, 2733, 3608]. Development [3568, 2207, 2128, 3021, 138, 2121, 1689, 2444, 2699, 2082, 3140, 2821, 2489, 2042, 3567, 3572, 2256, 2291, 2303, 3296, 2801, 3375, 2754, 1996, 2850, 1609, 1659, 856, 1828, 1347, 1662, 1703, 1741, 964, 529, 10, 827, 1494, 622, 1632, 399, 629, 1047, 1251]. Developmental [3506, 1993, 3045, 2472, 763, 1915, 902]. developments [224]. deviation [1792]. Deviations [2332, 1010, 1849]. Dextran [2019, 1297]. di- [457]. di-and [493]. Diabetes [625]. Diabetic [2492, 1998, 685]. Diagnosis [3401, 55]. diagnostic [55, 1863]. diagram [376, 430, 1368]. diagrams [569, 847, 758]. Dialect [3426]. dialysis [945]. diameter [558]. diameters [288]. diastolic [210]. Diatom [2494]. Dicarboxylic [1361]. Dichotomy [3254]. Dictyostelium [2043, 2574, 1736, 1040]. dideoxycytidine [1257, 1107]. Dieckmann [2129]. dielectric [540]. diet [1182]. Difference [2395, 2136, 3158, 3380, 2405, 3024, 481, 19]. Differences [2004, 2394, 2214, 494, 45, 880]. Different [2757, 2059, 3119, 2173, 2772, 1374, 1654, 619, 1036, 838, 423]. Differential [2241, 2282, 2808, 2074, 3561, 2641, 3574, 2375, 2727, 3363, 2482, 3180, 1541, 1884, 381, 397, 1321, 675, 1203, 1236, 1864, 1050, 761, 366, 1587, 1275, 19, 319]. Differentiated [2341, 3333]. Differentiation [2571, 2799, 197, 1527, 3054, 2943, 3073, 173, 131, 1569, 1446, 1457, 1277, 1375, 382, 1939]. difficile [3288, 3201]. Diffuse [3565, 1667, 1834]. diffused [317]. diffusible [1773]. Diffusion [1091, 2915, 2998, 630, 3490, 2036, 1239, 1652, 2916, 2831, 2800, 2865, 2863, 3229, 3392, 3322, 2909, 2911, 2904, 2136, 1029, 1973, 2669, 2884, 1265, 2913, 2578, 3561, 392, 2674, 1695, 2679, 2426, 2889, 2204, 2736, 2936, 250, 1731, 3304, 222, 2003, 2989, 3093, 2505, 3371, 887, 589, 3129, 3092, 2912, 3513, 3586, 2046, 3516, 2975, 3123, 649, 450, 786, 158, 205, 286, 173, 814, 911, 927, 569, 291, 157, 231, 1602, 1755, 678, 134, 724, 513, 7, 111, 172, 682, 1206, 1343, 762, 386, 10, 516, 1860, 190, 1804, 2867, 729, 700, 9, 45, 1026, 384, 495, 118, 132]. diffusion [155, 185, 273, 372, 101, 275, 1366, 176, 414, 859, 628, 587, 661, 373, 2917, 212, 2374]. Diffusion-enhanced [1265]. diffusion-governed [173]. diffusion-reaction [7]. diffusional [810, 579, 983]. diffusionally [502, 518].

Diffusive [2211, 3220, 1517, 1852]. Diffusively [1548, 1598]. Diffusivity [2384]. digest [1291]. digestive [783]. digestive-tract [783]. Dilemma [2190, 2123, 2405, 1183]. Dilemmas [3610]. dilute [3130, 58]. dilution [1289, 1379, 1511]. Dimension [2138, 931, 586, 991, 1162]. Dimensional [2824, 2514, 2767, 2669, 3517, 3054, 2578, 2568, 3291, 3342, 3094, 3222, 2912, 3095, 3376, 3232, 3496, 1572, 2485, 500, 958, 1018, 809, 430, 1488, 1206, 1437, 1220, 166, 766, 387, 132, 176, 938, 1303]. Dimensionality [2780, 1358]. Dimensions [1995, 3462, 3030, 3519, 3445, 1139, 1006]. Dimerisation [2910]. dimerised [3573]. Dimers [2735]. Dimorphism [3239, 1653, 1853]. d’intérêt [126]. Dioxide [2004]. Dipole [446, 68, 436, 322, 477, 477]. dipoles [180, 42]. Direct [2924, 3605, 3548, 1789]. Direct- [2924]. Directed [2698, 2491, 2360, 3182, 738]. direction [1231, 1232]. Directional [3191, 1991, 3326, 1916, 1931, 2726]. Directions [3184, 873, 1620]. directly [1513]. Disaggregating [3347]. Disc [3217]. Discerning [1706]. discoideum [2574, 1736, 1040]. discontinuities [1311]. discontinuity [84]. discontinuous [672]. Discovering [1094]. Discovery [3568, 3213, 374]. Discrete [3131, 3265, 3210, 2108, 2394, 2839, 3036, 2013, 3336, 3502, 3349, 2733, 2930, 2020, 2857, 2584, 2545, 2161, 2753, 3602, 1534, 958, 438, 1749, 1909, 1470, 977, 1793, 799, 689]. Discrete-Time [3336, 3602, 2013, 438]. discretization [1905]. discrimination [761]. Discussion [1110, 1066, 1068, 1070, 1118, 1064, 1114, 17, 304]. Disease [3622, 2811, 3523, 3189, 2861, 2516, 3025, 3458, 3041, 3395, 2162, 2106, 2839, 2153, 2325, 3077, 3098, 2179, 2717, 2816, 2827, 2021, 3277, 3457, 3288, 2956, 3351, 2435, 3115, 2218, 2570, 3296, 3514, 2125, 2480, 3284, 3112, 3602, 1589, 1845, 1835, 362, 1808, 434, 1882]. Disease-Induced [3395, 3457]. Diseased [3285, 3428]. Diseases [2141, 2960, 2963, 3300, 1802, 3606, 2229, 3077, 2363, 2924, 2554, 2415, 3113, 3604, 2737, 1247, 970, 1691]. Disinfection [2262, 2767]. Disjunctive [2476]. disk [441]. disk-sphere [441]. disorder [451]. Disorders [3197, 1418]. Dispersal [2199, 3090, 2831, 2712, 2946, 2651, 2404, 2873, 3116, 2889, 2892, 2166, 3500, 2100, 2020, 2747, 3074, 2984, 2841, 2414, 3182, 1421, 1845, 1461, 1936, 1974, 1688, 1115, 1945, 932]. Dispersion [2384, 2654, 2102, 2882, 2468, 2975, 605, 1481]. dispersive [582, 552]. Display [2983]. Disposition [3150]. Disseminated [64, 1633]. Dissemination [3443]. dissimilar [220]. dissimilarity [929]. dissipation [88]. Dissipative [587, 1422, 85, 729, 218, 331, 971]. Dissociation [2688, 21]. distal [190]. Distance [2980, 2244, 3074, 2410, 2627, 3470, 683, 759, 788, 1895, 1049, 1405]. Distances [2359, 2032, 2213, 2844, 1913]. Distancing [2849]. distemper [1523]. Distinct [1890]. Distinguished [750]. distinguishes [1484]. Distinguishing [2360, 2226, 1511]. distorted [1129]. distorting [166]. Distributed [354, 2162, 2852, 3388, 2065, 2401, 254, 1210, 1463, 931, 966, 1015, 1376, 234, 358, 1580]. Distribution [2751, 3622, 1539, 2409, 2109, 7, 2011, 607, 2305, 782, 2252, 3190, 3172, 2164, 1078, 2359, 2593, 2027, 3571, 195, 1997, 2237, 3063, 2306, 1434, 1484, 1010, 978, 1842, 242, 165, 425, 1037, 969, 1639, 2660, 2051, 321, 2073, 245, 695, 181, 356, 1969, 1317, 1106, 1163, 142, 1788, 1051, 359, 1135, 339, 1137, 422, 1618, 840, 58, 1222, 553, 575, 221, 983, 349, 463]. Distributions [2920, 3163, 2496, 3079, 3196, 2865, 2249, 2839, 2701, 1130, 2971, 2161, 2356, 3284, 2953, 1518, 1519, 1147, 901, 900, 1417, 399, 692, 727, 848, 997, 1242, 90, 1042]. Disturbance [2949, 2772]. Disturbances [2266]. disturbed [1439]. Diurnal [2011]. Divergence [2252]. Diverse [1633, 1698]. diversification [1446]. Diversifies [2779]. Diversity [3163, 2339, 2946, 384, 3426, 2772, 1469, 1479, 1391, 654, 1644, 775]. Division [125, 2038, 2750, 3565, 2397, 1531, 24, 297, 211]. Division- [2750]. DNA [1976, 1512, 1484, 1105, 1077, 1154, 2589, 1748, 2833, 1499, 1007, 2023, 969, 1444, 844, 1199, 2588, 948, 1933, 2139, 2636, 2855, 909, 3425, 839, 1953, 2899, 2321, 1125, 1955, 1106, 2069, 1710, 153, 1188, 1896, 1083, 857, 1997, 31]. DNA-based [1512]. Do [2174, 3173, 3567, 3013, 355]. dodecanedioic [1947]. Does [2518, 2571, 1244, 2276, 3207, 1861]. Dog [3486]. Doll [3346]. Domain [2036, 3342, 2736, 2415, 3372, 2079, 3108, 1239, 1755, 118, 339]. Domains [2446, 3453, 3161, 2127, 3291, 2615, 2505, 1328, 1602, 1787]. domesticus [2954]. Dominance [1583, 3050, 3155, 849, 1993, 1073, 491]. dominated [1591]. Don [2025, 2157]. Done [3040]. donor [1547]. Dormancy [2060, 2276]. Dose [2920, 2269, 3590, 2708, 15, 1816, 1203, 1473, 545, 657, 1506]. Dose-Dependence [2708]. dose-effect [1203, 657]. dose-response [1473, 545, 1506]. doses [1704]. dosimetric [328]. Dosimetry [2098]. Dosing [3569]. Double [1852, 1953, 494]. Double-jump [1852]. double-layer [494]. double-stranded [1953]. Doubling [2742]. Down [2770, 2744, 3286]. Down-Regulation [2770]. Doxorubicin [2531, 1925]. Drag [3583, 3400]. Drainage [2765, 3056]. Drastic [2176]. Dreitlein [229, 239]. Driessche [1802]. Drift [3253]. Drifted [3345]. Drinkers [3247]. Drive [3624]. Driven [2383, 2036, 2800, 3624, 3244, 3109, 3565, 2921, 3342, 3372, 2781, 3519, 3558, 2947, 1239, 856, 1302]. Driving [3341]. Droop [2335]. droplet [1716]. Drosophila [3217, 647]. Drug [2942, 3568, 2015, 3294, 3454, 3153, 2102, 3575, 2530, 2215, 3003, 3457, 2881, 3567, 3195, 3572, 2764, 3150, 2186, 2714, 3570, 2478, 1957, 3571, 1433, 3119, 1999, 866, 1471, 1448, 1871, 870, 1474, 1956, 1890, 1985, 559, 560, 611, 1481, 1980]. Drug-resistant [1999, 866]. Drug-Sparing [2215]. Drug-Supply [2764]. Drugs [2061, 3390, 1744, 871]. Dryland [3536]. dsDNA [1438]. Dual [3424, 2513, 1734]. Dual-Layer [3424]. dual-mode [1734]. Due [2155, 3619, 2990, 3259, 2441, 53, 1733, 394, 1739, 637, 1277, 945, 3027]. Duncan [1647]. Duplex [3310, 1748]. Duplexing [3425]. Duration [3273, 1962, 1794, 1985]. During [2877, 3455, 2681, 2825, 2963, 2369, 2451, 3434, 3140, 3369, 2809, 2982, 3544, 2271, 2553, 2186, 3209, 2754, 1723, 989, 918, 1884, 1439, 1746, 1489, 141, 246, 1810, 1766, 839, 441, 525, 1361, 42]. Dustbathing [114]. Dwarf [2622]. Dynamic [2193, 3247, 3424, 2316, 2395, 2357, 3477, 2699, 2583, 551, 3047, 2597, 128, 2522, 2120, 2406, 3164, 2497, 3143, 3296, 2110, 1985, 2261, 1954, 326, 278, 232, 1900, 1784, 1299, 75, 214, 1734, 1304, 1813, 1116, 1768, 1178, 2009, 1467, 2159]. Dynamical [2148, 2330, 3265, 2160, 2807, 3109, 2165, 2326, 3141, 1476, 3307, 1801, 2284, 735, 3012, 352, 2891, 366, 2367, 1346, 1345, 3188, 770, 50, 1563, 1382, 668, 906, 825, 434, 1611, 251, 970, 218, 331, 1, 260, 1927, 3563]. Dynamics [3253, 2234, 2518, 3622, 2748, 2824, 2944, 2469, 3015, 3021, 1159, 3105, 3501, 2383, 1995, 2557, 3455, 2250, 3189, 2315, 2449, 2511, 2853, 2389, 2419, 3260, 2563, 2458, 2485, 1374, 2080, 2259, 1944, 3618, 2388, 3609, 2962, 2681, 3399, 2685, 3317, 2552, 2160, 2076, 2392, 3257, 69, 2191, 2090, 3193, 2515, 2473, 2785, 2673, 2902, 2177, 2911, 2562, 3171, 3068, 3138, 3427, 2376, 3030, 3083, 102, 2210, 2134, 3606, 3398, 2883, 2167, 2053, 2594, 3036, 2556, 2229, 2818, 3521, 3486, 2898, 3124, 2476, 2939, 2520, 2363, 3377, 3459, 2442, 2701, 3097, 3581, 2958, 2612, 2349, 2615, 2641, 3574, 2479, 2739, 3078, 2223]. Dynamics [2794, 3492, 2809, 3407, 3227, 2100, 3512, 3349, 2868, 2763, 2457, 2856, 3053, 2142, 2622, 2364, 1647, 1198, 2086, 2976, 3002, 3333, 3338, 2764, 2956, 2495, 1613, 2078, 3503, 2618, 3172, 2964, 3437, 3034, 2614, 2756, 481, 3270, 2416, 2728, 3549, 2784, 3578, 3386, 2081, 3240, 2431, 3135, 2814, 2585, 2570, 3101, 2926, 2529, 3525, 3086, 2200, 3532, 2342, 1087, 1333, 1661, 2797, 2382, 2322, 2987, 3426, 2154, 2125, 2418, 2480, 2297, 3542, 3119, 3412, 3396, 1998, 2947, 1459, 3573, 3607, 703, 2525, 2526, 3282, 2653, 3232, 3452, 1483, 2723, 2832, 1894, 1949, 987, 1764, 1276, 764]. dynamics [1943, 333, 397, 1881, 1516, 1536, 566, 1406, 1099, 1555, 1306, 1977, 1056, 1020, 1556, 1846, 1561, 881, 1749, 1909, 1529, 1835, 966, 1603, 1838, 691, 722, 1262, 1458, 338, 1687, 1679, 1701, 1874, 1192, 1476, 1530, 939, 432, 1292, 1220, 1317, 1811, 1327, 688, 1161, 807, 1832, 1978, 1796, 1901, 944, 318, 506, 1120, 1031, 1441, 1580, 1510, 1885, 1592, 1945, 1907, 1716, 1869, 938, 1698, 1290, 1544, 1358, 1157, 1922, 1649, 3121]. Dynein [2722]. Dysplastic [3333]. Dysregulation [3259, 3080]. Dystrophy [3589].

E. [2560, 1397, 1398, 3434, 2891]. each [1130]. Early [2923, 3009, 3366, 2974, 2481, 2797, 3325, 3591, 2089, 3376, 1609, 1986, 474]. Eavesdropping [2149]. Ebenman [1056]. Eberhart [1840]. Ebola [3077, 3284]. Echinocytes [2471]. echo [1944]. eclosion [647]. ECM [3361]. Eco [2315, 2853, 3488, 3617, 338]. eco-dynamics [338]. Eco-Epidemiological [2315, 2853, 3488]. Eco-evolutionary [3617]. Ecohydrology [3017]. Ecological [3015, 2334, 2377, 248, 2297, 3586, 2773, 2270, 1469, 1396, 800, 1638, 1965, 570, 695, 327, 273]. Ecologists [3357, 1646]. Ecology [2946, 2595, 3239, 3236, 3007, 2342, 160, 1123, 1951, 299, 1678, 1117, 1196, 1116, 654, 775, 1605, 603, 587, 1104]. economic [1119, 1174]. economics [1121, 1169]. Ecosystem [2364, 347, 1825, 784, 617, 1428, 1429, 1522, 893, 1211]. Ecosystem-like [347]. Ecosystems [2066, 2473, 3187, 2497, 2176, 254, 1919, 449, 426, 249, 1267, 1114, 936]. ectoderm [1457]. ed [2222]. Eden [2488]. Edge [3424, 3584, 244].

Edge-Based [3424]. Edge-Deleted [3584]. edition [1818]. Editor [1898, 790, 812, 832]. Editorial [744, 778, 789, 804, 811, 823, 824, 831, 842, 851, 862, 895, 913, 922, 934, 942, 953, 994, 1002, 1014, 1023, 1033, 1045, 1053, 1062, 1076, 1086, 1093, 1102, 1109, 2909, 3192, 3104, 3440, 1722, 1989, 2093, 2507, 894, 905, 952, 1046, 1034]. Editors [1648, 2129, 863]. eds [2253]. Education [2730, 17, 125]. EEG [2526, 1143, 991, 2525]. Effect [2557, 3383, 2895, 2231, 2409, 2201, 2616, 2030, 2767, 3590, 1415, 2515, 3577, 2423, 2904, 2992, 2710, 2565, 3359, 2393, 2055, 3484, 1870, 3136, 2988, 2554, 2892, 2374, 2305, 3132, 3502, 2100, 3103, 2956, 2243, 3587, 2287, 3570, 3439, 1135, 2286, 2382, 1152, 3344, 3250, 1358, 2775, 2306, 1944, 1667, 1485, 1680, 1654, 197, 1639, 617, 1185, 189, 1607, 1203, 923, 1705, 1472, 837, 1457, 1410, 525, 652, 1950, 38, 1880, 1776, 657, 1580, 1970, 887, 560, 1303, 559]. Effective [2538, 3130, 2580, 165, 976]. effectively [594]. Effectiveness [3601, 2981, 2624, 2994, 2792, 3183]. effectors [1902]. Effects [2469, 3501, 2282, 2861, 2563, 1842, 3323, 2196, 2080, 2236, 2923, 3257, 3588, 2061, 1548, 1598, 816, 3237, 2019, 2300, 2629, 3287, 3171, 3329, 2334, 820, 2637, 3052, 3360, 2696, 3384, 2467, 3050, 3003, 2085, 3502, 2285, 1220, 830, 2881, 344, 582, 3179, 2964, 3152, 3115, 3070, 2354, 2646, 495, 273, 2092, 2268, 2970, 3308, 496, 466, 552, 605, 1557, 2985, 694, 3430, 2414, 3027, 3254, 2975, 1513, 3084, 2472, 2653, 3180, 3411, 2099, 1747, 184, 1845, 165, 346, 1448, 1948, 1444, 1835, 431, 870, 540, 362, 1615, 1744, 1323, 454, 1604, 742, 1978, 647, 1600, 1890, 1957, 1081]. effects [1511, 571, 186, 983]. Efficacy [2906, 3575, 2700, 1817, 1871, 1925]. Efficiency [3329, 2817, 2503, 227, 310, 1735, 1837]. Efficient [3233, 2358, 1227, 2884, 2453, 3558, 1829]. efflux [1686]. effort [1758, 1833]. EGFR [3537]. Egg [3044, 1167]. eggs [1752]. Egypt [2786]. Eigen [644, 3387, 564, 268, 2512, 2643, 3146, 1140]. Eigensolutions [586]. eigenvalue [678]. eight [613]. eight-parameter [613]. Either [3114]. elastance [455]. Elastic [3528, 182, 577, 1427, 565, 52, 78, 196, 261]. elasticity [1575]. Elastohydrodynamics [2238]. Electric [640, 2475, 2620, 356, 388, 387, 2195, 1875, 785, 400]. Electric-Field [2475]. Electrical [2848, 2285, 1064, 2510, 272]. Electrically [68, 1982, 387]. electrocardiographic [55]. electrocardiography [830, 38]. electrocardiology [1634]. electrolyte [615]. Electromagnetic [579, 59]. Electromigration [2654]. Electron [62, 687]. Electron-phonon [62]. Electrophoresis [2654, 2500]. Electrophysiological [2475]. Electrophysiology [3451, 2895]. electropotentials [615]. Electrostatic [639, 3400]. Electrotonic [1376, 1209]. elegans [1865, 2974, 3423]. Element [2998, 3496, 3030, 6, 1664, 853]. element/level [1934]. Elementary [2257, 3511, 342, 1784]. Elevated [3367, 1148]. Eleventh [125]. Elimination [3569, 1481]. Elizabeth [1920]. Elmergreen [65, 215]. elongating [1238, 1314]. elongation [1826, 92]. Elovich [133, 187]. elucidate [1886]. Elucidating [2037]. Embedded [3110, 2912]. embedding [1149].

Embracing [3497]. Embryo [2974, 3226, 3495, 1609, 856, 964, 1251]. Embryonic [2661, 2161, 1384]. embryos [964, 1047]. Emergence [3004, 3518, 1527, 2950, 2618, 1547, 2186, 1275, 2198, 1177, 1777, 1058]. Emergent [2155, 2442, 1982]. Emerging [1802, 2737]. Emission [2052, 3295, 133]. Emotion [2518]. Emotion-Cognition [2518]. Empathy [1772]. Emphasis [152, 29]. Empirical [2638, 1436, 164, 260]. Employment [150]. emption [245]. Enable [3548]. enables [1434]. encoders [369]. encoding [1082, 1942]. Encodings [3403]. encounter [1139]. encounters [1650]. Encouragement [2788]. End [2607, 956, 195]. end-points [195]. end-systole [956]. Endangered [2435]. Endemic [2854]. Endemicity [3604, 3603, 1112, 1113]. Endocrine [2883, 2319]. Endocytosis [2327]. Endogenous [3569]. endosymbiosis [1275]. Endothelial [2825, 2451, 3360, 2691, 1542, 1311, 1487]. ends [714]. Enemies [2858]. Enemy [2858]. Energetically [2156]. energetics [705, 37]. energies [454]. Energy [301, 2189, 2393, 601, 602, 603, 2002, 3048, 2461, 690, 62, 315, 856, 429, 687, 170, 224, 454, 1130, 1813, 1664, 479, 543, 1278]. Engineered [3545]. Engineering [505, 2781, 1847, 1310, 1877]. Enhanced [3484, 363, 1265]. Enhancement [2700]. Enrichment [2499, 2842, 882]. ensemble [338, 1312]. Enterocutaneous [2236]. Enterotoxigenic [2318]. Entity [1850]. entrained [741]. entrainment [998, 272]. Entropy [928, 918, 2403, 2361, 2189, 1143, 1162, 1790, 936]. Entry [3061]. Enumeration [2539, 2267, 54]. environ [834]. Environment [3177, 2066, 2674, 2591, 3047, 3304, 2461, 3182, 3063, 1555, 721, 673, 2459, 2073, 1404, 283, 1025]. Environmental [3201, 2577, 1177, 1391, 2011, 3619, 3605, 2243, 3197, 1915, 75, 98, 175, 277, 276, 217, 1269, 1467, 1425]. Environmentally [2554]. Environments [2121, 2311, 2712, 2657, 2439, 2464, 2698, 2276, 2233, 2762, 3248, 1439, 1797, 679, 1579, 407, 1665, 1768, 1691, 1223]. Enzootic [2959]. Enzymatic [2743, 3319, 2636, 702, 762, 1339, 937]. Enzyme [2290, 2060, 83, 3448, 3447, 3322, 3499, 2489, 2113, 2593, 1631, 2054, 2059, 363, 178, 401, 1415, 364, 1249, 480, 454, 761, 1339, 726, 1622, 995, 1059, 459, 86, 1952, 1827, 1281, 703]. Enzyme-Based [3448]. enzyme-catalysed [1339]. enzyme-substrate [454]. enzymes [459]. Epidemic [2605, 2811, 2015, 2506, 3424, 3535, 3090, 2437, 3134, 2307, 3481, 3198, 3264, 2231, 2409, 2408, 2191, 2090, 3515, 3028, 3395, 2528, 2829, 3068, 2249, 2839, 1431, 2637, 2465, 3077, 2701, 2331, 3088, 2816, 3078, 2021, 2888, 3605, 2725, 3271, 2838, 2310, 2784, 2008, 2849, 3510, 2322, 3343, 3188, 3061, 3487, 3602, 1110, 1845, 135, 1846, 1200, 60, 1872, 634, 2031, 1956, 1546, 1490, 1625]. Epidemics [3435, 3189, 2412, 2918, 3618, 2408, 2616, 2076, 3181, 2647, 3072, 2399, 2997, 3195, 3216, 2631, 3035, 3600, 3445, 2690, 2771, 3183, 1800, 1794, 1694, 528, 1111, 1112, 1113, 1960, 581, 590, 1507]. Epidemiological [2315, 2162, 3212, 2053, 3038, 2018, 3174, 3405, 2853, 3488, 3089, 1714, 1513]. Epidemiology [3598, 2982, 3597, 2212]. epilepsy [1134]. epileptic [66, 67]. epimorphic [996]. episode [1962]. Episodic [2313, 1389]. epistatic [1851].

Epithelial [2256, 2291, 2081, 1189, 1601, 1888]. Epithelium [2570]. Epizootic [2959, 620, 1639]. Epstein [2581]. Equation [2837, 3131, 3237, 2102, 2326, 3219, 2641, 3388, 2794, 2727, 3363, 3180, 435, 381, 644, 291, 326, 187, 69, 518, 1491, 1508, 1559, 536, 795, 1830, 9, 481, 234, 511, 1250, 872, 628, 845]. Equations [2199, 3177, 2395, 3252, 2136, 2074, 2053, 3561, 2462, 2304, 3158, 3024, 3108, 101, 3035, 2338, 2579, 1541, 158, 205, 286, 233, 397, 569, 631, 231, 675, 135, 586, 1268, 172, 1964, 538, 535, 268, 563, 1855, 418, 1050, 215, 368, 404, 419, 766, 327, 623, 118, 132, 155, 372, 447, 706, 366, 1283, 292, 275, 531, 176, 279, 400, 587, 19, 373]. Equilibria [2688, 2265, 3493, 3066, 3224, 2849, 3487, 3263, 854, 228, 815]. Equilibrium [3322, 1803, 1614, 2761, 3088, 3500, 615, 3147, 2989, 944, 2427, 3120, 3079, 903, 2562, 1848, 1904, 222, 142, 153, 273, 698, 601, 602, 87, 1459]. equimolar [414]. equiprobably [738]. equitability [245]. Equivalence [2424, 563]. equivalent [15, 1634, 3465, 1131, 311]. Era [2354]. Eradicate [3458]. Eradication [2235, 2795]. Erlang [1210]. Errata [287, 151]. Erratum [1456, 1975, 3031, 3264, 2917, 1658, 2157, 3121, 2447, 1976, 404, 1974, 1379, 2159, 2804, 3032, 2994, 2526, 2771, 3120]. Error [2068, 1227, 1662, 624, 1240, 1241]. error-minimizing [624]. error-prone [1240, 1241]. Errors [3131, 2941]. erythrocyte [71, 441]. Erythrocytes [2471, 1559, 937, 925]. escaping [1649]. Escherichia [2318, 2369, 2934, 3350, 3268, 2271]. Esophageal [3401]. ESS [2144, 2034]. essay [1118]. Essential [3577, 912, 1899]. Establishment [3362, 2946]. Estimates [2668, 15, 3432, 156, 3470, 678]. Estimating [3145, 1715, 2038, 1766, 1672, 2564, 2321, 2888, 3347, 1786, 1710, 3087, 1810, 1608, 545, 657, 1831]. Estimation [2119, 3345, 2193, 2511, 2460, 2408, 3069, 2644, 2165, 3463, 2302, 3572, 2120, 3578, 2718, 3600, 3597, 1222, 2773, 2261, 179, 1143, 390, 1897, 1292, 1465, 1163, 1162, 1342, 1624, 921]. estimator [1828]. Eukaryotic [3549, 2110, 2159, 1780, 1283]. Europe [3414]. European [1101, 1814, 3300]. Eusocial [2146]. Eutacticity [2251]. Eutrophication [3085]. Evaluating [2117, 2792, 2729, 908]. Evaluation [1136, 3189, 3397, 2907, 2165, 304, 2652, 1138, 1789, 199, 830, 773, 442, 1526, 459]. Evaluations [3284]. Evaporation [3310]. evasion [502]. Even [3172]. Event [3558]. Events [2166, 2171, 1539, 1903, 763, 592]. Every [3110]. evidence [1861]. evidenced [1539]. Evo [3089]. Evo-epidemiological [3089]. Evolution [2748, 2469, 2790, 2355, 2202, 2373, 2757, 2137, 1629, 3610, 2600, 3283, 2552, 2048, 3257, 2587, 1126, 940, 2562, 3489, 2968, 3235, 2789, 3212, 2404, 2619, 617, 2581, 2082, 3565, 2873, 3116, 3478, 2713, 3369, 2251, 2044, 3047, 2124, 2142, 3608, 2490, 1574, 2020, 2747, 3172, 737, 1443, 3000, 1072, 2198, 2368, 1894, 1949, 160, 1105, 1867, 411, 1394, 776, 1390, 1389, 1641, 1658, 5, 796, 1392, 1671, 299, 539, 480, 1458, 1359, 1604, 1169, 937, 848, 350, 601, 602, 603, 1699, 1935, 158, 776]. Evolutionarily [2889, 3117, 3612]. Evolutionary [2380, 2559, 1255, 2077, 3089, 2388, 3609, 2865, 2361, 1668, 2376, 2651, 2594, 2556, 1904, 2568, 3611, 3078, 2362, 3231, 2618, 3172, 3614, 3240, 2987, 3426, 2154, 1907, 3616, 1894, 1949, 1504, 207, 3617, 752, 1709, 701, 338, 1687, 997, 1281, 1649, 2129]. evolutive [899]. Evolvability [1627]. Evolve [2044, 3207]. Evolving [3453]. Exact [2811, 1430, 2971, 3560, 2252, 3035, 279, 298, 433, 631, 945, 653, 810]. Exactly [2760, 251]. examine [1985]. Examining [2099, 2198]. Example [594, 3517, 2214, 1491]. exceptional [1048]. Excess [2964]. Exchange [2677, 2087, 246, 946, 29, 46, 972, 358]. Exchangeability [2131]. excitability [1064]. Excitable [2824, 3323, 3043, 2364, 68, 1873, 477, 12, 1216, 1612, 1673, 1316, 1237, 1324]. Excitation [2247, 2510, 206, 240, 1065, 199, 16, 477, 536, 1378, 42, 59, 553, 575]. excitations [180]. Excitatory [2791, 566]. Excluding [2546]. Exclusion [2766, 2661, 2601, 2591, 100, 338, 1779, 1087]. excretion [725]. exercise [1967]. exhaustible [1121]. exhaustion [1792]. exhibiting [1642]. Exhibits [2972]. Existence [126, 3219, 418, 848, 698, 2734, 815, 461, 126]. Exit [3061]. exocytotic [1766, 1582]. Exosomes [3365]. Expanding [3158, 3230]. Expansion [2140, 3186, 3615, 795, 2794, 3111, 1893, 1766, 785]. expansions [1638]. expected [1409, 1083]. Expenditure [2393]. experience [1013, 1595]. Experiment [3324, 2037, 2652, 2598, 829, 12, 1064]. Experimental [3397, 2902, 2045, 1103, 2050, 2599, 1223, 2801, 2820, 3516, 2822, 336, 1737, 1594, 444, 1727, 660]. Experiments [2670, 2574, 2980, 2165, 3039, 2544, 3325, 3373, 530, 608, 23, 1506, 1514, 461, 2302, 609]. expiration [141]. Explain [3437, 3023]. explanation [184, 402, 428, 1370]. Explicit [435, 3522, 919, 960, 3309, 2351, 1951, 1689]. Exploitation [2639, 1383, 1170]. Exploited [2468]. Exploiter [2137]. Exploration [2387, 3389, 2428]. Explore [2998, 3537, 3273]. Exploring [3404, 3508, 3277]. Explosive [985]. explosivity [3418]. Exponential [3072, 3284, 931, 836, 1130]. Exponentially [2162, 966]. exponents [1516]. Exportations [3113]. expose [1674]. Exposed [2169]. Exposure [2193, 3569, 657]. Exposures [2829]. express [1674]. Expression [2779, 3393, 2629, 1994, 2878, 2504, 2635, 2936, 2602, 2777, 2505, 3423, 1273, 1942, 1854]. Expressions [2059, 1004]. Extended [2684, 2104, 3623, 936]. Extending [1388, 3506, 3168, 2796]. Extension [1995, 3067, 809]. Extensions [3034, 792, 1908]. Extent [2488, 1222, 1407]. External [2770, 2669, 1308, 237, 344, 818]. Externally [1302]. Extinction [1764, 3383, 721, 2528, 3139, 2434, 3313, 521, 1507, 2356, 3602, 649, 1617, 1639, 1409, 1551, 3002, 797, 944, 112]. Extracellular [2371, 2825, 938]. Extracellular-Matrix-Guided [2825]. extraction [1483]. Extracts [3044]. extrapair [1758]. Extremal [2806, 996]. Extreme [2774, 2377, 3240]. Extremum [182, 533, 1492]. Extrinsic [3479, 3115, 2268]. Eye [3160, 1992, 3272, 182, 179, 134, 324, 1228, 499, 74]. Eyelid [2238].

f [2598, 2596, 1306, 3219]. F-actin [2596, 1306]. Fabricius [1094, 1095]. Faces [2619]. Facets [3235]. Facilitation [2419, 1579, 1206]. Factor [2324, 2724, 2687, 2781, 1773, 501, 1416, 1937]. factor-induced [1937]. Factorization [3162]. Factors [2752, 2387, 3042, 3336, 572, 849, 1141, 1271, 1591]. Factors-Mediated [2752]. facts [1858]. Fahraeus [837]. Failure [2364, 2414, 1155]. fairness [1772]. falciparum [2028, 2441, 3607]. Fallopia [3169]. Family [2183, 366, 1966]. Far [2543, 87]. Farmers [3414]. Fast [2831, 3149, 1612, 1214, 1020, 779, 1582, 1643]. Fastest [2986]. Fatal [2774]. Fatality [2228]. fate [1924, 660]. fatty [1342]. FBLM [3429]. Fc [2716]. Fear [3250]. Feasibility [2795, 624]. feasible [426]. Features [3469, 2155, 378]. Featuring [2536]. Fechner [187]. Fecundity [3526, 3117, 1458]. Federation [877]. Feed [2244]. Feed-Forward [2244]. Feedback [3015, 1247, 3393, 3593, 3193, 2863, 3507, 3156, 2929, 2244, 3244, 3484, 2037, 3136, 2349, 2417, 2990, 2285, 2042, 2474, 1901, 2925, 1742, 1270, 394, 2033, 1493, 266, 1449, 1298, 621, 445, 1267, 1345, 1127, 1128, 1938, 1643]. Feedback-mediated [1901]. Feeding [2744, 2794, 2585, 1229, 1371, 1185, 1476, 1660, 1483, 1229]. FEM [3429]. females [1983]. Ferment [3213]. Ferret [3486]. Ferrière [2129]. Ferris [1899]. Fertility [1758, 1961, 600, 880]. fertilization [1309, 1758]. Fetal [3099, 1489]. Fever [2906, 2959, 3142, 2786, 2798, 2930, 3320, 3174, 3180]. Few [3531]. fewer [1923]. Fewnomial [3491]. Fiber [2393, 2822, 1489, 102, 210, 31]. Fibers [2012, 1216, 388]. Fibonacci [2254]. fibre [945]. Fibres [3532, 2195, 539, 1155]. Fibrin [3009]. fibro [1418]. fibro-proliferative [1418]. Fibroblast [2324, 3073, 1489, 1501]. fibroblasts [1885]. Fibrocontractive [2687]. fibrosis [474]. Fibrous [3218, 1525]. Fiddler [2097]. Field [2668, 2475, 540, 2272, 2869, 2237, 640, 320, 35, 1129, 1500, 237, 356, 218, 331, 340, 59, 274, 90]. Field-induced [540]. Fields [2770, 470, 173, 750, 197, 406, 1660, 785]. Fighting [2144]. figures [1858]. filament [1542]. Filaments [2824, 1518, 1519, 2070, 1252, 1487]. file [758]. Film [2557, 2055, 3310, 589, 2723, 2832]. Filopodia [2911]. Filtering [2896, 357, 672, 1887]. filters [48, 33]. Filtration [2681]. Filtres [48, 33]. Final [2605, 2412, 3481, 3395, 2021, 2725]. Find [2855]. Finding [2688, 3133, 1278, 220]. fingerprint [1154]. Fingerprints [2061, 1905, 2062, 1891]. Finite [2998, 2077, 2250, 3496, 2916, 2917, 2376, 2136, 3030, 2924, 3529, 3505, 3022, 3240, 3146, 3464, 87, 365, 1194, 577, 1707, 1183, 853, 1907, 1934, 1235]. Finite-Difference [2136]. finite-element [1934]. finite-element/level-set [1934]. Finite-Population [2924]. Finite-Size [2916, 2917]. Finkelstein [1819]. Fire [2791, 1834, 2050, 1667, 1630]. Fire-diffuse-fire [1834, 1667]. firing [1378]. First [877, 2290, 2705, 2613, 3007, 2333, 1963, 1480, 218]. first-line [1480]. first-order [1963]. First-Passage [2290]. First-Passage-Time [2613]. Fish [3580, 2681, 2644, 2609, 2763, 3117, 1842, 1875, 1611, 1690, 3626]. Fisher [435, 2425, 2643, 447]. fisheries [1169, 1120, 1171]. fishery [1119]. Fission [3332, 2182, 3498]. Fistulae [2236]. Fit [2356, 825]. Fitch [2487]. Fitness [3154, 3280, 2749, 2068, 2898, 2873, 3403, 3611, 2276, 2302, 2233, 3146, 2551, 1894, 1949, 1148, 5, 1126, 1871, 1627, 478, 1649]. fittest [1280]. Fitting [2668, 1975, 1958, 1897, 214]. Five [3300, 1349, 1140]. Fix [607]. Fixation [3163, 2077, 2577, 1851, 3609, 2749, 1944, 2577]. Fixed [3144, 2288, 1531, 3072, 2415, 1629, 1222]. Flagellar [92, 3549, 1928, 2070]. flagellates [1365]. Flat [2674, 1126]. Flexibility [3472, 2878, 2961]. Flexible [2666, 3336, 306, 377, 1218]. Flies [2305, 2306]. Flight [2746]. Fligner [1646]. FLIM [2980]. flip [477]. flip-flop [477]. Flocculation [2259]. Flocs [2682]. flop [477]. Flow [233, 1213, 53, 2538, 690, 3107, 2586, 2996, 2966, 1385, 134, 1992, 2908, 3359, 826, 2744, 2045, 2608, 3066, 2821, 2374, 108, 2302, 3082, 2391, 2876, 3143, 2677, 2945, 2762, 2463, 3232, 989, 1412, 1975, 1958, 1645, 1516, 1536, 165, 164, 969, 324, 467, 550, 820, 915, 1299, 558, 943, 227, 310, 1664, 1417, 1660, 1753, 166, 565, 368, 404, 419, 914, 81, 190, 70, 975, 684, 652, 171, 1311, 520, 1135, 1751, 496, 466, 576, 694, 772, 781, 58, 908, 261, 1104, 845, 316, 3032, 2381]. Flow-Controlled [2677]. Flow-Cytometry-Based [2302]. Flow-induced [1213]. flow-rate [1417]. Flowering [3460]. Flowers [2815]. Flows [2514, 2939, 2650, 56, 11, 1432]. Flu [3605, 1776]. Fluctuating [2939, 1509, 268, 407, 1608, 1025]. Fluctuation [453, 903]. Fluctuations [2175, 2577, 3479, 2049, 2768, 2154, 806, 1391, 818, 1655, 87]. Fluid [2091, 3367, 2328, 2681, 2262, 2514, 3107, 1992, 2765, 3082, 2876, 2478, 2677, 2463, 3476, 3027, 2723, 2832, 182, 233, 1560, 837, 108, 1487, 938]. fluid-bone [938]. fluids [1385]. fluminensis [3245, 3587]. Fluorescence [2075, 1214, 1903]. Fluorescent [3452]. Flurothyl [1465]. Flurothyl-induced [1465]. Flutter [2285]. Flux [3511, 233, 810, 380, 479]. Fluxes [2071, 870, 429, 623, 1361]. fly [647]. Flytrap [2739]. focal [1652]. Focusing [3622, 3272, 1287]. Fokker [511, 1250]. Fold [2934]. Fold-Change [2934]. folding [1012, 1279, 1803]. folding-unfolding [1803]. follicle [157, 1500]. Follicular [2091, 2387]. Follistatin [2323]. follow [390]. follow-up [390]. Following [3340, 3320, 1816, 141, 629, 1762]. Food [2536, 3210, 3083, 3121, 1529, 3417, 2144, 2859, 2964, 492, 2603, 2099, 439, 736, 1429, 1508, 1679, 766, 879]. food-chain [1429]. food-chains [879]. Foot [2570]. Foot-and-Mouth [2570]. footprints [1821]. Forager [3246]. Foragers [2144]. Foraging [2419, 3620, 1416, 1650, 1179]. foraminifera [1054]. Force [1995, 2369, 4, 1739, 1674, 576, 1856]. Forced [2685, 49, 1432]. Forces [3293, 540, 57, 627]. Forcing [2861]. Forecast [3021]. Forecasting [3382, 3600, 3597, 1318]. Forerster [69]. Forest [3062, 2151]. Foreword [664, 745, 1003, 863]. Form [2496, 3196, 3495, 135, 102, 1986, 1068, 327, 1622, 1048]. Formaggia [2253]. Formal [138, 948, 1987]. formalism [936]. formalization [794]. Formation [2245, 2329, 2571, 3125, 3161, 3506, 3009, 2163, 3554, 1994, 2669, 2884, 2674, 3155, 3434, 2504, 2450, 3217, 2350, 3221, 2990, 3502, 2914, 2731, 2364, 2079, 2203, 1993, 3437, 2628, 2256, 2584, 2268, 3308, 2841, 3092, 3045, 2211, 2168, 3423, 3095, 3623, 1572, 1733, 1585, 1602, 1755, 1489, 1389, 1463, 1615, 1703, 1741, 1745, 529, 1860, 1496, 1587, 859, 2732]. Formations [2344]. formed [1240, 1241]. Formic [1301]. Forms [2717]. Formula [2021]. formulae [1789]. Formulation [2365, 2612, 3388, 2078, 1865, 1323, 1038]. Formulations [3305]. forum [896, 950]. Forward [2244]. Foundation [2572, 1064]. founding [2660]. Four [3054, 1607, 725, 930]. four-compartment [930]. Four-Dimensional [3054]. Four-urn [725]. Fourier [1627]. FPT [3233]. Fractal [2833, 2280, 1060, 1886, 991, 985]. Fractals [1381, 798]. fraction [1710, 423, 1083]. Fractional [3126, 2199, 3305, 3392, 931, 3092, 2059, 1390]. Fractional-in-Space [3092]. Fractionated [3369]. fractions [1344]. Fracture [3009]. Fragmentation [3477, 2382, 1518, 1519]. fragmented [1445]. Framework [3475, 3051, 3068, 3524, 2573, 3054, 3244, 3565, 2094, 3611, 2058, 3039, 1581, 1815]. France [1159, 3088]. Francis [140, 1930]. Francisco [976]. Frankenhaeuser [706]. FRAP [3217]. Fraser [1173]. Free [3269, 2651, 2929, 2610, 2857, 3548, 3326, 3048, 1849, 2073, 479, 1278]. Free-Swimming [3326]. freedom [1230, 75]. French [104, 129, 123, 48, 33, 126]. Frequencies [2254]. Frequency [2600, 2969, 455, 2544, 2527, 3249, 1942, 578, 847, 1984, 818]. Frequency-Dependent [2600]. Frequency-Sensitive [3249]. Frequent [2866]. frère [129]. frère-sœur [129]. Freshwater [3122, 3220, 733]. FRET [2980]. Friendship [2039]. FRiND [3589]. Frog [3311, 3044, 1581]. Frogs [2668]. Front [2378, 1228]. Fronts [2395, 2377, 1517, 1535]. Frost [2310, 1863]. Fruit [2305, 2306, 647, 979]. fruits [1356]. FT4 [2929]. Fuel [2675, 2937]. Fujita [847]. Fully [2296, 2113, 1622]. Function [2927, 2083, 2813, 2109, 2164, 2778, 3143, 3510, 2002, 1797, 291, 1771, 821, 718, 215, 1218, 29, 1167, 1618, 1514, 1399]. Functional [2617, 253, 2240, 3043, 2565, 2333, 2614, 2846, 2198, 484, 717, 885, 582, 1209]. Functionality [2805]. Functions [3079, 3551, 2018, 3096, 3318, 890, 959, 786, 1525, 441, 963, 1256, 1132, 1545, 1597]. Fundamental [3242, 536, 279]. Funding [2316]. Fungal [2121, 3133, 3216, 1797, 620, 1156]. Fungi [2599]. Further [1175, 548, 3274, 709, 1113]. Future [3184, 873, 865, 1620]. Fuxman [1372]. Fuzzy [1906, 1359, 2022, 2310, 3164].

G [1817, 2030, 3193, 2200, 3573]. G-CSF [3193, 2200]. G-CSF/ [3193]. G-protein [1817, 2030]. G2 [3590]. G2/M [3590]. Galerkin [853, 372]. Galerkin-Ritz [372]. Galton [1930, 3480, 1089]. Game [2077, 2376, 2594, 2699, 2190, 2123, 3530, 3419, 3155, 2405, 2149, 3295, 2142, 2849, 3240, 2987, 2154, 2533, 2034, 1908, 1604, 791, 147, 1907]. Game-Theoretic [3419, 2149, 3295, 2034, 1604, 791]. Game-Theoretical [3155]. Games [3609, 3530, 1482, 1979]. Gametocytes [2418]. Gamma [3336]. Ganglia [2421, 575, 553]. Gap [2895, 2475, 888, 752, 1376, 1386]. gap-junction [1376]. Gap-Junctional [2475]. gap-penalty [1386]. gapped [1929]. gaps [1079]. Gas [3037, 141, 246, 548, 1810, 247, 190, 414, 589]. Gastric [2646, 1820, 908]. Gastro [2318]. Gastro-Intestinal [2318]. gastrointestinal [651]. GATA1 [2597]. Gatherers [3414]. Gating [2796, 2293, 436, 477, 1393]. Gause [143, 2565]. Gaussian [1638, 2768].

GBM [2970]. GC [2713]. Gels [2303]. gelsolin [1519]. gender [1814]. Gene [2355, 3469, 2779, 3393, 3069, 2993, 3465, 3051, 2386, 1994, 2813, 2710, 3624, 2711, 3258, 2878, 2504, 2936, 1472, 2124, 2602, 2777, 2348, 2164, 2505, 2910, 2656, 2871, 2198, 3249, 2488, 3080, 1867, 1942, 1294, 886, 818, 1620, 1856, 1938, 1854]. Genealogies [2663]. General [2550, 2448, 2167, 103, 565, 2856, 3190, 2676, 3222, 747, 644, 167, 241, 425, 1323, 883, 368, 404, 726, 1181, 208, 443, 325]. Generalised [2169, 2390, 3060]. Generalist [2044]. Generality [2021]. generalization [766, 588, 633]. Generalizations [2224, 623]. Generalized [2506, 2485, 3610, 2115, 3237, 1286, 3055, 1727, 3219, 2058, 2627, 255, 624, 1496, 1131, 511, 774, 239, 985, 1917, 478]. Generalizing [3548]. generate [779, 738]. Generated [2392, 2088, 911, 2514]. generates [1857]. Generating [2209, 421, 458, 1462, 1739]. Generation [3251, 3564, 3357, 2091, 3397, 2369, 1080, 3140, 2401, 1785, 173, 82, 139, 462, 1146, 1696, 547, 1869, 1854, 74]. Generations [3230]. generative [948]. generator [1634]. Generic [1454, 3263]. Genes [2600, 3577, 2420, 2447, 2986, 1504, 1674]. Genetic [2380, 3163, 3534, 763, 3257, 3617, 2456, 3269, 3186, 2667, 1879, 1895, 2104, 2568, 3425, 3230, 85, 3209, 2695, 262, 1391, 1988, 96, 1153, 1921, 1968, 748, 204, 1889, 1844]. Genetically [2125]. Genetics [2813, 760, 1176, 670, 1070]. Genistein [2098]. genoma [512]. Genome [3449, 2968, 2205, 2969, 1997, 2254, 1009, 2488]. Genomes [2713, 2871, 3137, 1707]. genomic [1710, 1083]. Genomics [3449]. Genotope [2219]. genotype [848]. genotypes [880]. Genotypic [3613]. genotyping [1905]. Geographic [2279]. Geometric [3270, 3096, 2410, 2293, 1321, 849, 497]. Geometrical [2452, 2104, 906]. Geometries [2916, 925, 753, 2917]. Geometry [2269, 2833, 2558, 2393, 3218, 3425, 3463, 2675, 2628, 2410, 2627, 2482, 683, 759, 788, 739, 1452, 1238, 1203, 1818, 1453, 761, 481, 311, 610]. German [776]. Germinal [2387, 1837, 1613]. germination [1156, 743]. Germline [2689]. Get [2257]. GH3 [3114]. GH3-Mediated [3114]. giant [398, 494]. Gierer [2504, 1016]. Gill [2681]. Gillespie [3547, 3548]. Given [2027, 1915]. glabrata [69, 2308]. Gland [3328]. glass [232]. glass-bead [232]. Glazier [3199]. Glia [3179]. Glioblastoma [2759]. Glioma [3052, 3372]. Global [2670, 927, 3391, 2076, 3008, 509, 2265, 881, 348, 3517, 2465, 3289, 1874, 2018, 2179, 2331, 3100, 2641, 2479, 3336, 368, 419, 879, 2497, 2532, 2235, 2546, 490, 1292, 689, 880, 88, 1140, 404]. globular [808]. glomerular [1297]. glucagon [314]. gluconeogenesis [1361]. Glucose [3157, 2686, 2271, 3355, 314, 1738, 1021]. glutamate [267]. glycemia [685]. Glycine [3159]. glycolysis [498, 937]. glycoprotein [1215, 1288]. Glycoproteins [3279]. GM [2752]. GnRH [2848, 3049]. Go [2970, 112]. Goal [2521]. Goals [27, 28, 865]. Goldbeter [3541]. Goldfish [2371]. Gompertz [1946, 41, 966]. Gompertzian [2170, 635]. gon [1889]. Gonadotrophs [2323]. gonadotropin [621, 445]. Goodness [2356]. Goodwin [292]. Gordon [861, 961, 1118]. Got [2257]. Gottman [1900].

Governed [2442, 173, 118, 132]. Governing [3333, 5]. Gradient [2610, 1674]. Gradients [2669, 2230, 2095, 2450, 1147, 1347, 1535, 1425, 983]. Grafts [2116, 2101]. Grained [2768]. Grammar [2142, 757, 1343, 1850]. grammars [1779]. grammatical [1850]. grammatically [1861]. grand [338]. Grandmothering [3283]. Graner [3199]. Granule [3140, 2105, 1408, 1419]. Granulopoiesis [3194, 3193]. Graph [2837, 2647, 1570, 606, 2152]. Graph-Based [2837]. Graph-theoretical [1570]. graphical [299, 459]. Graphs [2806, 2856, 3035, 693, 561, 815, 857]. grass [1923]. Grazer [3501, 3613, 2964, 3620, 1664]. grazers [1923]. Green [2668, 291]. Gregory [2143]. Grid [2155]. Grid-Cell [2155]. Gromov [3339]. Group [3462, 2539, 3610, 2163, 2749, 2442, 3463, 3347, 2847, 3176, 2987, 2289, 940, 882, 988, 1835, 61]. Group-Based [3462, 3463]. Group-Mass [2847]. Groups [3146, 1354, 459]. Grow [2970]. Grow-or-Go [2970]. Growing [2446, 2918, 3161, 3554, 3478, 2615, 1993, 2872, 2505, 2926, 1824, 1602, 2051, 1343, 1787, 1804, 1403, 1752]. grown [724, 858]. Growth [2103, 3163, 3501, 2403, 2036, 2596, 1797, 3543, 3001, 3234, 2473, 3582, 2025, 2157, 2324, 2396, 2301, 3398, 2398, 2769, 3477, 3565, 3324, 2170, 3478, 2239, 2991, 2888, 3372, 2302, 1271, 2312, 2687, 2271, 2079, 3190, 2151, 2377, 3076, 2964, 3215, 2781, 3270, 3549, 3023, 3019, 2957, 2970, 2985, 2220, 3375, 3087, 2351, 2695, 805, 2947, 3376, 2653, 1091, 1633, 641, 1883, 1356, 1946, 530, 679, 1805, 1138, 334, 1755, 378, 1187, 163, 898, 295, 1689, 1912, 1244, 926, 1308, 1343, 798, 472, 1870, 1089, 1924, 1799, 713, 1813, 250, 1166, 529, 1307, 1955, 1141, 1348, 1449]. growth [1218, 1591, 415, 954, 13, 1841, 774, 912, 337, 297, 1132, 410, 1416, 1937, 193, 211, 635, 1545, 1597, 1565, 932, 1290, 478, 703, 1859, 1503, 653]. growth-limiting [703]. GTPases [2183]. Guangdong [3421]. guaranteed [1227]. Guest [863]. Guidance [1991, 2746, 2094]. Guide [2879, 3516]. Guided [2825, 2301]. guinea [1897]. Gut [3306, 2661, 3570]. gynamical [1247]. Gypsy [2242]. Gyrotactic [3058, 1628]. Gyrotaxis [1287, 1365].

H [759, 1952, 1810, 759]. H. [1118]. H1 [2642]. H1N1 [3088, 2529]. H3N2 [3345]. HAART [1884, 1981]. Habitat [2201, 3158, 3094, 3102, 2382, 1799, 1880]. Habitats [2404, 1509, 1603, 1445]. Hadamard [2343]. haematopoiesis [677, 1762]. Haemophilic [3397]. Hair [3309]. Hairpins [2870]. half [129, 793]. half-life [793]. Hall [3472]. Hall-Type [3472]. HAM [2444]. HAM/TSP [2444]. Hamilton [641, 216]. Hamiltonian [469]. Hamming [1895]. Handbook [2222]. Handedness [2097]. handling [1651]. Hans [3214]. Hantavirus [2014, 2231, 2439, 1800]. Haploid [2027, 818]. Haptotactic [2000]. Haptotaxis [2781]. harbour [1523]. hard [1278]. hardback [1920, 1877, 1840]. Hardcover [1831, 2143, 1818, 1647]. hardness [1747]. hare [1588]. Harmless [2774, 655]. harmonic [413]. Harnessing [3175]. Harold [1118]. Harvested [2644]. Harvesting [2757, 2177, 3024, 3164, 337, 1498, 174, 1282, 186]. Harvesting-Induced [2757]. hasard [129]. having [425]. hazard [408].

HCV [3145]. HDL [2923]. head [533, 860]. Healing [3178, 2617, 3009, 3100, 2192, 2727, 2338, 1489, 1418]. Health [125, 3593, 2017, 3206, 504]. Healthcare [3201, 3288]. Healthcare-Associated [3288]. Healthy [3160, 3397]. hearing [540]. Heart [3065, 2510, 852, 782, 385, 3]. Heat [462, 103, 321, 3423, 533, 278, 860, 768, 715, 853, 79]. heated [230]. Heating [2105]. Heavy [2199]. Hebbian [2871]. Hedging [3359]. Height [2897, 3587, 2034]. held [3628]. Helical [2564, 1230, 1231, 1232]. Help [3357, 3173, 2570, 1683]. helper [1685, 1742, 1472]. Hematocrit [637, 1289, 1379]. Hematological [3041]. Hematoma [3009]. Hematopoietic [2352]. hemispheres [1804]. Hemodynamical [2793]. hemolytic [434]. Henle [2417, 1152]. Henri [2743]. Henry [302]. hepatic [1481]. Hepatitis [2766, 3504, 3578, 2714, 1318]. Hepatocellular [3366]. Hepatocytes [2371, 2327]. HER2 [2300, 2629]. Herbivore [3228, 2555, 1263, 1923, 1864]. Herbivorous [3458]. Herd [2900]. Heredity [1930]. hermaphoroditic [460]. Hermaphrodites [3422, 1983]. Herpes [2357]. Hes1 [2703, 2910]. Hessenberg [1128]. Heterodimerization [2115]. Heterogeneity [2942, 3317, 3329, 3615, 2334, 3394, 3003, 2085, 3209, 2984, 3586, 3080, 911, 1737, 1548, 1598, 1705, 1886, 1545, 1597, 1683]. Heterogeneous [2207, 2128, 884, 3420, 2307, 3399, 3432, 2392, 3613, 3085, 2404, 3054, 2939, 3221, 2120, 3057, 2502, 2266, 3182, 1797, 1007, 999, 1870, 634, 2073, 1146, 1058, 1271, 1612, 1691, 727, 1137]. heterogeneously [1580]. heterogeneously-distributed [1580]. Heterosexual [2724]. Heterosexuals [3227]. Hidden [3483, 2006, 1887]. hierarchal [1850]. Hierarchical [2538, 899, 2648, 825, 449, 331]. Hierarchically [1462]. Hierarchies [3050, 1011, 1414, 1248, 1144]. Hierarchy [3155, 2350, 1527, 491, 1593]. Higgins [498]. High [3393, 2923, 2716, 3517, 2630, 309, 1631, 1827]. High-Dimensional [3517]. Higher [3445, 1661]. higher-order [1661]. Highlights [3360]. Highly [2354]. Hill [1619, 335]. Hindered [2821]. Hindmarsh [1806]. Hippocampal [2524, 1376]. hippocampus [1514]. Hirundo [1814]. Histories [2397]. History [2757, 3274, 1883, 1714, 1813, 318]. history-dependent [318]. hit [1773, 1362]. HIT-15 [1362]. HIV [2119, 3010, 3455, 2015, 2506, 2202, 1881, 2080, 2330, 1724, 1714, 3091, 3303, 1821, 1258, 2017, 2707, 2210, 2084, 2630, 1838, 2016, 2898, 1906, 2610, 1981, 1474, 2215, 2715, 2663, 3508, 1472, 2724, 3227, 2302, 2881, 2096, 2187, 3338, 3410, 1788, 2208, 2354, 2186, 976, 2580, 1956, 1038, 1890, 1957, 1985, 1057, 2797, 3119, 2720, 2466, 2493, 2795, 2792, 3183, 2402, 2261]. HIV-1 [1724, 1258, 2017, 2898, 3508, 2302, 2881, 2096, 1788, 2354, 2186, 1890, 1957]. HIV-2 [1258]. HIV-Infected [3410]. HIV/ [2402, 2261]. HIV/AIDS [2015, 2630, 1838, 2187, 2580]. Hodgkin [3252, 1268, 1964, 1855, 2796, 1925, 1064, 1378, 1642]. Hoffmann [1017]. Hogeweg [3199]. Holling [2565]. Hollow [2256, 945]. hollow-fibre [945]. Holography [955]. Homage [2804, 2803]. Home [1733, 2630, 2008, 1973].

Home-Based [2630]. Homeostasis [1902, 2381, 1864, 1459, 1601, 1888]. Homeostatic [3563, 3374, 3299, 3073, 3249]. homeothermic [715, 852]. Homo [1097]. Homodimers [3573]. Homogeneous [2692, 2924, 2266, 111, 1430, 541, 11, 146]. Homogenisation [3582]. Homogenization [2595, 2400, 2475, 3182]. Homology [813, 3529, 886, 773]. Homoplasies [2213, 2171]. Honey [3260, 3619, 2533, 1936, 1974, 1511]. honey-bee [1511]. Honeybee [2746, 3070, 3246]. Honeycomb [2966]. Hopf [3323, 3161, 2584]. Hopfield [2068]. Hopfield-type [2068]. Hopkins [1930]. Hormonal [3227, 1783]. Hormonal-Based [3227]. hormone [1615, 209, 445]. hornet [1354]. Hornets [2174]. Horseshoe [3583]. Hospitalization [2630]. Hospitals [3207]. Host [2234, 2202, 3175, 3621, 2665, 2337, 3212, 2581, 1458, 2991, 2809, 3604, 2874, 2982, 2411, 3115, 3216, 3174, 3542, 2787, 3607, 2217, 2493, 3411, 719, 1845, 1555, 1857, 1477, 1478, 1765, 1185, 2905, 1476, 2479, 2441, 1181, 1557, 1441]. Host-parasite [1458]. host-parasitoid [1555, 1476, 1441]. Hostile [2434]. Hosts [2469, 3604, 2720]. Hotel [1075]. Hotelling [1118]. Hotels [2568]. Household [2053]. Houssine [2804, 2803]. Houston [125]. Hox [1674]. HPV [2944, 3105, 2265, 2320, 3599]. HSV [2357]. HSV-2 [2357]. HTLV [1926, 2444, 2635, 2582, 1592]. HTLV-I [1926, 2444, 2635, 2582, 1592]. Hua [1847]. hudsonicus [2151]. Human [2942, 3015, 2169, 3475, 2091, 3300, 1992, 3486, 2219, 2581, 3077, 2471, 3274, 2809, 3259, 2961, 1762, 2111, 3607, 2254, 1723, 182, 533, 1154, 1975, 1958, 1536, 1521, 136, 860, 822, 739, 1262, 1771, 1734, 1480, 1684, 1050, 1820, 441, 593, 660, 307, 613, 991, 1575, 853, 168, 910, 855, 475, 282, 1483]. Humans [3152, 2214, 22]. humidity [822]. humile [1716]. humoral [1801]. Humour [1992]. Hungarica [329]. Hungry [1180]. Hunter [3414]. Hunter-Gatherers [3414]. Hutchinson [1114]. Huxley [3252, 1268, 1964, 1855, 2796, 1064, 1378, 706, 1642]. Hyatt [1075]. Hybrid [2779, 3555, 2893, 3559, 3377, 2823, 3382, 3492, 2733, 2691, 2213, 2370, 1343]. Hybridization [2845, 2171, 153]. hydranth [656]. Hydraulic [3308, 829]. hydrocephalus [1452, 1590]. Hydrochaeris [2279]. Hydrodynamic [2422, 3326, 2975, 324]. hydrodynamical [520]. Hydrodynamically [2457]. Hydrodynamically-Coupled [2457]. hydrodynamics [115]. hydromechanics [1452]. hydrostatic [1581]. hygroscopic [593]. Hyperbaric [2492]. Hyperbolic [3420, 2470, 2579]. hypercomplex [1896]. Hypercube [2598]. hypercycle [1140]. hypercycles [366]. hyperemic [1338]. hypergeometric [553, 575]. Hypergraphs [3134]. Hyperinfectivity [2738]. hyperphagia [1912]. hyperplasia [1849]. hyperpolarization [706, 835]. hypertrophied [497]. Hyphae [2599]. Hypnosis [1332]. hypotension [1886]. Hypothalamic [209, 445]. Hypothalamus [2929]. hypotheses [653]. Hypothesis [2766, 1583, 210, 576, 1713]. hypothetical [1265, 819]. Hypoxia [2752, 2087, 1584, 1672]. Hypoxic [2759]. Hysteresis [2711, 183].

IAA [2474]. iconic [354]. Ideal [2749, 2651, 2073]. ideality [1303]. ideas [1633, 1067]. Identifiability [3467, 969, 2210, 3531, 3466, 3075, 2649, 2580, 3174, 3405, 2261]. Identifiable [2499]. Identification [3453, 2131, 1823, 2993, 2902, 528, 1005, 357, 1270, 1553, 1719, 1027]. Identify [2570]. Identifying [3469, 3577, 2011, 3073, 260]. Identity [152]. Idiotypic [1375, 1436, 1018, 1017, 1030, 1198, 1277, 1293, 1399, 1337]. IEEE [125]. If [3624, 3144]. IgE [2716]. IgG [269]. II [1949, 987, 458, 55, 828, 566, 169, 444, 213, 1231, 1260, 1519, 430, 1371, 102, 56, 100, 901, 1095, 172, 1398, 1112, 321, 1429, 964, 419, 67, 13, 2291, 284, 1968, 633, 423, 590, 445, 1333, 602, 25, 1346, 560, 1128, 2295, 91, 2708, 2306]. III [205, 1232, 2565, 1113, 1132, 603, 325, 611, 1996]. IL-7 [3410]. Illuminates [3047]. illumination [525]. illus [1831]. illustration [1814]. IMA [1802]. Image [2833, 3107, 3162, 2301, 901, 900]. Image-Based [3107]. Image-Guided [2301]. images [220]. Imaging [3572, 3452]. Imatinib [2258, 2314, 2553]. Imatinib-Treated [2314]. imipenem [1426]. Imitation [2624, 2994]. immanent [1067]. Immature [3621, 3508]. Immersed [3065]. immigrant [568]. Immigration [2712]. Immobilization [3592]. immobilized [726]. Immune [3364, 2861, 2741, 3621, 2828, 2688, 1259, 1260, 258, 2508, 2632, 2769, 3459, 2933, 3625, 2921, 2991, 3380, 2441, 284, 2532, 3254, 3607, 1276, 308, 1380, 1369, 1792, 1801, 434, 1161, 1193, 1651, 2028, 1526, 1326, 1902, 1320, 1859]. Immune-Suppressed [3607]. Immune-Virus [2632]. Immunity [3435, 3345, 2755, 2437, 2616, 2854, 2631, 2738, 3266, 719, 1458, 1513, 1346]. Immuno [3212, 3174]. immunochemical [1029]. Immunodominance [2576]. immunogenic [1292]. immunoglobulin [1922]. Immunological [3538, 2363, 2958, 1890, 1957, 1698]. Immunology [3184, 1019, 407]. Immunotherapy [2178, 2317, 2001, 2659, 3521, 3363, 3595, 2700, 1435]. Impact [2942, 2766, 2506, 2962, 2017, 2903, 2153, 2325, 2698, 3097, 3227, 2100, 2937, 2187, 3185, 3400, 3510, 2260, 3542, 2235, 2720, 2624, 3607, 3111, 3282, 3183, 1767, 1988, 2994]. Impacts [2315, 2419, 2177, 3446, 3500, 2842]. impedance [1706]. Imperfect [2017, 3261, 2973, 2084, 2900, 2814, 2418]. implant [1525]. implants [1525]. Implementation [2980]. Implementing [3470]. Implications [3443, 2929, 2444, 3369, 2759, 1753, 1860, 1613, 2677, 19, 640, 1714, 1030, 1838, 210, 1813, 1278, 461, 407]. Implicit [2773, 1541]. Implies [2817]. Importance [3218, 2601, 2989, 2033, 1820]. Important [2283, 3316, 1120]. Importations [3113]. Improved [3520, 1756]. Improving [2993, 3217, 3571]. Impulse [2012, 326, 531, 611]. impulses [1964]. Impulsive [2388, 2126, 2736, 2881, 2503, 3180, 1890, 1957]. In-host [2479]. in-series [128]. Inaccurate [2453]. inactivation [1786]. Inactive [3380, 3221]. inbreeding [104, 600]. Incidence [3145, 2465, 2179, 3078, 2982, 1318, 1600]. incisions [1036]. includes [1389]. Including [3051, 3065, 2991, 3029, 1967, 1139, 1988, 805]. Inclusive [2551]. incoming [806]. incompatibility [966]. Inconsistency [3494].

Inconsistent [3223, 2965]. Incorporating [3501, 1995, 3193, 2101, 1948, 2016, 2862, 2687, 2964, 3430, 719, 1334, 1771, 1887]. incorporation [1737, 1103]. Increase [3163, 2276]. Increased [3526, 1723]. Incubation [2315, 3490, 3243, 1639, 1167]. independence [662]. Independent [2394, 2713, 752, 1323, 154]. independently [424]. Index [3368, 794, 1853, 491]. indicator [1525, 1289, 1379]. Indicators [3385, 3019]. Indices [2052, 771, 1222]. indirect [1465, 908]. Indirectly [2363]. Indiscriminate [1833]. Individual [2469, 2951, 3588, 2630, 2717, 3088, 3380, 3221, 3037, 3347, 1644, 3152, 2599, 2289, 2945, 1358, 2034, 1951, 1659, 2423, 1689, 1915, 966, 1605, 1397, 3032]. Individual-and [1644]. Individual-Based [2951, 3380, 3221, 2945, 1951, 1689, 1605, 3032, 1397]. Individual-Level [2717]. Individuals [2704, 2536, 2163, 2637, 3172, 2008, 1714, 1114]. induce [1970]. Induced [2004, 3383, 2774, 2799, 2757, 2616, 2770, 3395, 2973, 2139, 3581, 2554, 3457, 2049, 2105, 2057, 68, 1534, 770, 1213, 444, 296, 351, 1560, 443, 1976, 1933, 1491, 540, 1695, 1465, 483, 434, 1753, 1335, 1252, 1237, 1937, 1408, 1419, 1809, 1869, 1165]. Inducible [2752]. Induction [2489, 3197, 999]. industry [1838]. Inequalities [3472]. Inequality [2897, 1013, 1595]. inert [247]. Inertia [2194]. Infected [3455, 2707, 2988, 3410, 2720, 1906, 1981, 1788]. Infection [2014, 2766, 3145, 3105, 3455, 3526, 2080, 2840, 3091, 3303, 2385, 3395, 2216, 2444, 2973, 3446, 2229, 2997, 3504, 2641, 3276, 2582, 3546, 2809, 2862, 3259, 2881, 3127, 3320, 3544, 2797, 3445, 2626, 3119, 2089, 2466, 1800, 1881, 1724, 1714, 1478, 1909, 1926, 1981, 1474, 1732, 1472, 3338, 2028, 1890, 1957, 1592, 767, 1490, 1625, 1863]. Infections [2318, 3593, 3432, 2785, 2501, 3138, 2581, 3097, 3288, 2930, 2854, 3614, 3086, 528, 1513, 3427]. Infectious [3072, 1802, 2229, 2363, 3098, 2179, 2331, 2717, 2415, 3113, 2021, 3313, 2737, 362, 1691]. infective [634]. infectivity [1057]. Inference [3321, 2185, 2993, 2240, 3402, 902, 3217, 2575, 2490, 1704]. Inferential [1853]. Inferring [1846, 3068, 3373, 929]. Infested [3246]. Infinite [2250, 830, 3240, 3513, 1707, 118]. Inflammation [2617, 2955, 2448, 3254]. Inflammatory [3256, 3406, 2827, 2481]. Influence [2705, 3395, 3361, 2883, 2855, 2679, 3478, 2504, 2821, 3221, 2854, 2503, 2703, 767, 2463, 246, 1261, 1909, 1782, 1068, 818]. Influences [2946, 3385, 3525, 1525]. Influenza [3253, 2241, 3345, 2263, 3261, 2973, 3088, 2641, 2680, 3118, 2809, 2416, 2529]. influx [1686]. Infomax [2521]. Inform [3046]. Information [2538, 3269, 2122, 2789, 2462, 97, 2305, 2026, 3516, 2306, 188, 808, 667, 1955, 807, 280, 64, 152, 1968, 984, 1104, 1648]. Informational [2048, 1028]. infrared [62]. infrastructure [876]. infrequent [642]. Infrequently [2484]. Infusion [2677, 235]. inhaled [660, 475]. inhaling [474]. inherent [1889]. inherited [2709]. Inhibin [2323]. Inhibition [2752, 3566, 2692, 2693, 1579, 263, 266, 1288, 403, 1890, 1643, 74]. Inhibitor [2515, 2732, 2731, 2714, 1387]. Inhibitors [2024, 2881, 1795, 1339].

Inhibitory [2024, 2183, 2775, 394]. Inhomogeneous [2693, 3139, 3315, 1239, 1634]. Initial [3198, 3264, 2888, 3351, 2975, 571]. Initiation [2513, 2063, 1216, 1843]. Injecting [2015]. injection [559, 560]. Injections [3410]. Injury [3297, 2280]. Innate [3459]. Inner [2429, 1302, 46]. Inners [145]. Inositol [2081, 1573]. inotropic [1897]. input [1706, 560]. Insect [2193, 2907, 3141, 2368, 2472, 1531, 620, 719, 112]. insect-pathogen [719]. Insecticide [2905]. Insecticide-Treated [2905]. Insects [2203, 1648]. Insertion [2713]. Insight [2982]. Insights [2716, 3384, 2333, 2433, 2703, 2792]. Insoluble [2111, 474]. inspiration [141]. Inspired [3392, 3007, 3339]. Instabilities [2378, 2585]. Instability [2632, 3477, 2003, 2879, 3623, 206, 240, 1239, 1782, 1852, 1442, 1127, 1128]. instantaneously [1941]. Instigation [2625]. Institute [1900]. Institutions [3015]. instruction [1742]. Insulin [3152, 1695, 591, 923, 1362, 1344, 685]. intact [782, 1438]. Integral [2949, 86, 563, 234, 628]. Integrate [2791, 1630, 2050]. Integrate-and-Fire [2791, 1630]. Integrated [1927, 3256]. Integrating [2377]. Integration [2059, 197, 761, 223]. Integrative [2278]. Integrin [2033, 2000]. Integrin-mediated [2000]. Integrity [3443]. Integro [2395, 3158]. Integrodifference [3177, 2304, 3108, 1445, 2217]. intense [1762]. Intensity [2289, 59, 1407]. Intensive [2221]. Inter [2877, 2423, 1820]. inter-compartmental [1820]. Inter-individual [2423]. Inter-Synchronization [2877]. Interacting [2449, 2347, 2997, 1572, 35, 99, 100, 1409, 75, 238, 1393, 191, 399, 424]. Interaction [2518, 3364, 3262, 3290, 1942, 2423, 2642, 2422, 3384, 2674, 2655, 2459, 2991, 81, 3356, 2987, 1365, 2726, 2493, 3411, 3080, 1539, 1099, 1334, 1821, 1671, 143, 1610, 111, 1308, 80, 1024, 1457, 347, 1582, 930]. Interactions [3010, 371, 3327, 2861, 3170, 2955, 2508, 2632, 3538, 988, 2136, 3122, 3486, 2188, 3309, 2823, 2679, 3574, 2809, 2990, 1298, 2079, 2662, 2935, 3250, 3586, 2089, 2146, 1436, 998, 1851, 1555, 721, 580, 1835, 1761, 582, 260, 552, 446, 1663, 979, 1399, 1398]. interactive [542, 566, 875]. Interbreeding [3102]. Intercellular [2826, 2089, 630, 2043, 1311]. interconnected [957]. intercoupled [651, 567]. interest [126, 130]. Interface [3565, 938]. interfaces [540, 830]. interfacial [494]. interference [800, 379, 704, 1799]. Interferences [2503]. Interfilament [57]. intergral [70]. interior [738]. interleukin [1410]. interleukin-2 [1410]. interleukin-4 [1410]. Interlocking [1348]. intermediate [1235]. Intermediates [3540, 3263, 1803]. Intermittency [2270]. Intermittent [2491, 3347, 2985, 1962, 3457]. internal [431, 830, 818]. Internalization [3193, 856, 964, 1578]. International [301, 681, 841, 877, 941, 1159, 1208]. Internodons [3004]. interphase [1453]. Interplay [2877, 3566, 2847, 1448]. interpopulation [418]. interpretation [1894, 1949, 847, 1473, 1209]. Interpreting [2397]. interrelated [1727]. Interruption [2119]. Interruptions [2553, 1884]. Interspecies [2937]. Interspecific [2704, 1782, 3411]. Interstitial [3367, 3056]. intertwined [1889]. Interval [2285, 857, 561]. Intervened [2659]. Intervention [2498, 3351].

Interventions [3284, 2792, 3183]. Intestinal [3332, 2318, 3407, 272, 413, 1222]. Intra [2877, 3175, 1187, 3615, 2441, 2493, 1560, 128, 1511]. Intra- [2877]. intra-aortic [128]. Intra-Host [3175, 2493, 2441]. intra-nest [1511]. Intra-specific [1187]. Intra-tumour [3615]. intra-uterine [1560]. Intracellular [3551, 471, 2398, 2830, 2479, 3553, 1996, 3373, 3275, 3337, 1562, 1478, 1444, 1573, 1206, 1834, 1941, 1026]. Intraguild [2140, 3581, 3412, 2719]. Intrahost [2715, 2663]. intramolecular [759]. intraocular [1575]. Intrapredatory [2503]. Intraspecific [3417, 2243, 3412]. Intratrophic [1537]. Intratumoral [3359, 2227]. intravascular [1481]. Intrinsic [2834, 1329, 2139, 3115, 2268, 2957, 2154]. Introduce [2941]. Introduction [3562, 665, 2609, 1296, 2435, 77, 2130, 1802, 2221]. introductory [17]. Invading [2021, 2360]. invariance [814]. Invariant [2863, 673, 408]. Invasibility [2815, 1880]. Invasion [3475, 2395, 2828, 2831, 2185, 2812, 3361, 1871, 3046, 2939, 3581, 3546, 3372, 2377, 2857, 2287, 3074, 3532, 2341, 2660, 1732, 1697, 1950, 1137, 1934]. Invasions [2311, 2304, 3187, 2401, 3286]. Invasive [2140, 3169, 3348, 3245, 3122, 2698, 2435, 3215, 2970, 2152]. Invasiveness [2378]. Inverse [1436, 3609, 3098, 1634, 1807, 1059]. Inversion [2796, 924]. inversions [1188]. Invertebrate [2861]. Invertebrate-Disease [2861]. Investigate [2959, 2817, 3289, 2426]. Investigating [2724]. Investigation [2234, 2877, 3434, 2620, 2541, 1878, 1925, 2801, 2793, 2822, 1854, 19]. investment [1269]. Invitation [3576]. Involved [2290, 260, 1038]. involving [1249, 1803, 472, 638]. inward [522]. iodide [157]. iodine [793]. Ion [3170, 3305, 2058, 1335, 2293, 1667, 971, 90, 91]. Ionic [339, 1706, 356, 11, 623, 242]. Ionically [2114]. ionizing [742]. Ions [3523, 2007, 2345, 2494]. IP [1942]. IPT [3457]. irradiations [454]. Irreducibility [2432]. Irreversible [2054, 1946, 110, 425, 549, 89, 1339, 218]. irreversible-type [1339]. Irrigation [3017, 1309]. irritability [2062]. Irritation [2248, 2708]. ISBN [1646, 1900, 1920, 2129, 1802, 1819, 1930, 1818, 1647, 1877, 1831, 2143, 1940, 1847, 1899, 1648, 1840]. Ischemia [3297, 1338]. Ising [1199]. island [1764, 2009, 1157]. islet [923]. Islets [2282, 2926, 1892, 1344]. isobaric [294]. isobole [1024, 1236]. isoclinal [846]. isoeffect [1098]. Isoforms [2282]. isolated [1947, 1137, 1885]. Isolation [3416]. Isologous [1446]. isomerase [1735]. isometric [335]. Isometry [3146]. Isotherms [80]. isotype [1922]. Issue [2035, 2909, 3033]. Issues [3450, 2357, 3174, 993]. Iterated [2405, 1183]. Iteroparous [1701]. ITIS [3256]. IV [808, 256]. Ixodes [3111].

J [1646, 1647, 1648]. J. [1940]. Jacobi [216]. Jacobian [3597, 1128]. Jacques [1648]. James [1900]. jamming [1875]. Jane [663, 596]. japonica [3169, 2926]. Jellyfish [2744, 2585]. John [1900, 1920]. Johns [1930]. Joining [2619, 3468, 1935]. Jr [302]. July [1075, 1123, 3628]. Jump [2473, 3386, 3060, 1852]. Jump-Growth [2473]. junction [1311, 1376].

Junctional [2475]. Junctions [2895]. June [1159]. Justification [3067]. justify [1244]. Juvenile [2668, 2231, 3029, 1187]. juveniles [1056]. Juxtacrine [2268, 1535, 1621].

K12 [3434]. Kagna [595]. Katchalsky [233, 623]. Kauffman [422, 423]. KCl [309]. Kedem [233, 623]. Keep [2346]. Keizer [2172]. Keller [2039, 1283]. Keratinocyte [2101]. Keratocyte [3082]. Keratocytes [2047]. Kermack [1110, 135]. Kernels [2199]. Kerner [73]. Key [3328]. Kick [3124]. Kidney [3157, 3298, 2135, 2081, 2876, 3143, 323, 945, 2429]. Kill [3236]. killer [1785]. killing [545]. Kinase [2024, 1567]. Kinds [2211]. Kinematics [1230]. Kinesin [1995, 3437]. Kinesins [2684]. Kinetic [1737, 3210, 3447, 2398, 1763, 870, 2302, 2597, 1339, 2499, 1520, 113, 229, 580, 1619, 1156, 89, 142, 1832, 543, 1300, 1338, 983]. Kinetics [3196, 562, 2060, 2835, 157, 2371, 133, 3193, 3499, 294, 3484, 1387, 237, 2687, 2203, 3553, 2286, 635, 1512, 1633, 450, 786, 363, 487, 83, 1773, 1706, 364, 1390, 869, 1810, 1286, 919, 960, 1249, 682, 669, 1205, 1591, 317, 726, 1026, 13, 243, 146, 265, 2067, 1631, 995, 1059, 1967, 628, 1827, 855, 1281, 598, 1251]. Kinetochore [3244]. Kinetochore-Driven [3244]. King [2058]. kingdom [1225]. Kirshner [1802]. Klebsiella [2259]. Kleptoparasites [2144]. Kleptoparasitic [2536]. kleptoparasitism [1908]. knockout [1629]. knots [1570]. knowledge [121, 32]. Kolmogorov [1143]. Koshland [3541]. KPP [3219]. Kristin [1900]. Krogh [2269]. Krogh-Type [2269]. Krylov [333]. Kunming [2015].

L [2253, 1831, 1648, 3268]. l-arabinose [3268]. Label [2750]. Label-Structured [2750]. labeled [738, 636]. Labeling [2224, 2652, 1738]. labelled [1578]. labelling [1222]. labor [1531]. lac [2271]. Lactate [2686]. Lactic [3262, 3290]. Lactose [2271]. Lag [334, 1244]. Lagged [2234]. Lagrangians [622]. lags [704, 490]. Lake [3132, 2287]. lakes [1561]. lambda [1346]. laminar [166]. Landmarks [3103, 1054]. Landscape [2946, 1894, 1949, 1126, 767]. Landscapes [3154, 2048, 3403, 2502, 3146, 3308, 3220, 2009, 1627]. Langevin [1250]. Language [2600, 1861, 3426, 948]. Languages [138, 869, 1987]. Laplace [853]. Laplacian [2856, 3053]. Large [2332, 236, 578, 2595, 2334, 2520, 2717, 2575, 2649, 3255, 3556, 1010, 1823, 35, 924, 196, 101, 1885, 1337]. Large-Scale [2595, 2717, 3255]. larvae [998]. Larvivorous [2609]. Laser [2105, 1408, 1419, 1809, 1869]. Laser-Induced [2105, 1408, 1419, 1809]. Lassoing [2783]. Last [2963]. Latent [2339, 2415]. Latently [3455, 2707, 2720]. Lateral [2998, 678, 3316, 2593, 2488, 263, 74]. lateralization [1497]. Latin [2598]. Lattice [2311, 3560, 3037, 2411, 3483, 1914, 2857, 2975, 1951, 648]. Lattice-Free [2857]. lattice-valued [648]. Lattices [2434, 648]. Laurence [302]. Law [2592, 3340, 129, 643, 187, 402, 428, 837]. laws [218]. Layer [3424, 2019, 2055, 2107, 2158, 989, 1680, 494, 579, 496, 917]. Layers [2640]. lazy [1526]. LDL [2327]. leaching [1309]. Lead [3216]. Leadership [2998]. Leading [3367, 2426, 3385, 988, 244, 1557]. Leads [3566, 1772]. leaf [1986, 1036]. leakage [1630]. Leaky [2730, 2050]. Leaping [3552, 3557]. Learn [3324, 1861]. Learners [2965]. Learning [3475, 3223, 2965, 2405, 2142, 2871, 1671, 701, 1861]. Least [315, 2718, 2359, 2410, 1935]. least-squares [1935]. Leaves [2032, 928, 1036]. leaving [523]. Lecture [225, 374]. lectures [1819]. Lee [537]. left [210, 455, 196, 344, 782, 168, 1039]. Leibich [1229]. Leishmaniasis [3241, 3352]. Length [3580, 3626, 2684, 2713, 1082, 1194, 1518, 1519, 1707, 752, 2070, 1106, 2028, 1222, 195, 1235]. Length-Structured [3580, 3626]. Lengths [3173, 422, 1235]. lens [1228]. lepidoptera [1587]. Leptothorax [1094, 1095]. Lesion [3512]. Leslie [63, 489, 192, 456, 1351, 780]. less [1140, 1935]. Lesson [2371]. LET [1704]. Letter [1898, 595, 663, 596]. Leukaemia [2634]. Leukemia [3364, 2486, 2258, 2314, 2553, 2352]. leukocyte [1327]. Level [2469, 2577, 3451, 2895, 2840, 3257, 3588, 3162, 3402, 2977, 2717, 3546, 2768, 2354, 2623, 3004, 1094, 3168]. Level-1 [3402, 2977]. level-set [1934]. Levels [2056, 2772, 889, 959, 784, 1956, 559, 560, 611]. Levenberg [3597]. Levenshtein [732]. Levins [1654]. lexical [1687]. Liapunov [270]. libraries [1707]. License [1171]. Lie [3464]. Liebig [1371]. Life [3004, 125, 2757, 2552, 2361, 1385, 1813, 3274, 2982, 3034, 77, 2662, 2206, 2133, 793, 672, 1380, 1692, 668, 1656, 1264, 256, 283, 148]. Life-History [2757]. lifespans [966]. Lifetime [2290, 545]. Lifetimes [1319]. Lift [3583]. ligaments [1745]. Ligand [2045, 3147, 3573, 844, 1757]. ligands [931]. ligation [1661, 1544]. Light [3615, 296, 351, 133, 1050, 409, 650, 647, 410]. light-induced [296, 351]. Like [2115, 2647, 2032, 1151, 1417, 21, 347, 965, 1081]. Likelihood [2424, 3463, 2734, 3470, 1828, 1163, 1162, 1466]. Limb [2245, 2244, 2312, 1860, 46]. limbs [1152]. Limit [2245, 3175, 2565, 544, 651, 156, 1259, 312, 422]. limit-cycle [651]. Limitation [3216, 1970]. limitations [1143, 1229, 1371]. Limited [3199, 3195, 1650, 487, 1857, 1029, 919, 960, 1041, 283, 1366]. Limiting [1781, 1525, 3587, 418, 703]. Limits [510, 233, 1055]. Limulus [3583]. Lin [516]. Lindeman [1114]. Lindqvist [837]. Line [2998, 3279, 2672, 753, 2723, 2832, 62, 1480]. Lineages [3374, 3293, 2715, 902]. Linear [2380, 2175, 2875, 3555, 2996, 3292, 2018, 3550, 2914, 3075, 937, 2849, 337, 2995, 3573, 3220, 783, 959, 641, 262, 662, 672, 1900, 1789, 76, 80, 624, 1024, 1236, 792, 1050, 799, 196, 914, 1015, 582, 700, 251, 674, 1786, 152, 366, 712, 176, 279, 559, 560, 611, 1027, 1844]. linearization [73]. Linearized [939, 2717]. Linepithema [1716]. Lines [2742, 2205, 1975, 1958, 1477, 1478]. Link [2594, 3047]. linkage [885]. Linked [2527, 1748, 838, 1413, 706]. Linkers [2684]. Linking [2843, 2712, 1664, 2841, 1829, 1399]. Links [2636, 1466]. Lipid [2055, 2593, 640, 483, 90, 91]. Lipids [2557]. Lipoprotein [2148, 1725, 2327]. Lipoproteins [2923]. liquefaciens [1747]. Liquid [2161, 2132]. Listeria [3262, 3290]. Live [2716]. Liver [2366, 2438, 3614, 3296, 1342, 358, 3]. Liver-based [2438]. livers [1947]. Living [2122, 2789, 2630, 2620, 3022, 1696, 283]. Load [3544, 1479, 1300]. Loading [3501, 2872, 706]. loads [1724]. Local [1654, 2594, 3100, 3355, 2340, 243, 3426, 2012, 3148, 2168, 3571, 610, 3420, 1694, 368, 404, 419, 1163, 1041, 3356, 1188, 1386]. Localized [2392, 542, 566]. Locally [890]. Locate [3173]. Location [2919]. Loci [2559, 838]. locking [369]. Locomotion [3566]. locus [1594]. Loewenstein [187]. Logic [2522, 2804, 2803]. Logical [2802, 1270, 3165, 2143, 2804, 2803, 262, 1823, 1067]. Logistic [3327, 3238, 3388, 2181, 932, 518, 1508, 1218, 774, 976, 478]. Long [731, 3091, 2457, 2575, 3351, 2233, 2584, 3074, 421, 458, 1461, 1079, 9, 773, 105, 1481]. Long-Distance [3074]. long-range [1461]. long-tailed [1481]. Long-Term [3351, 2584, 731]. Long-Time [2457]. longest [1106]. Longevity [3423, 1354]. Longevity-Related [3423]. longitudinal [684]. Loop [2863, 2244, 2417, 2065, 2474, 1270, 330, 815, 18, 1267, 1345]. loop-characteristic [1345]. Loops [3541, 2042, 1345]. Loser [3050, 3155, 2350, 1604]. Loss [2755, 2616, 1855, 2854, 3246, 2986, 8, 1513]. Lotka [903, 231, 216, 439, 3417, 617, 1517, 985, 887, 223, 1358]. Loudness [2428]. Low [3551, 3590, 2326, 1692, 2327, 1724, 1018, 578, 102, 454, 101, 1378, 1704, 59]. low-dimensional [1018]. Low-Dose [3590]. low-LET [1704]. low-molecular [102, 101]. Low-Rank [2326]. Lower [3351, 2912, 1026]. Lower-Dimensional [2912]. lowest [1278]. LPS [2921]. LPS-Driven [2921]. Lucas [302, 1889]. Łukasiewicz [262]. lumbricoides [3138, 3427]. Lumen [2291, 157]. Luminal [2381]. luminescence [1480]. Lumped [2194, 1881]. Lung [141, 2111, 291, 93, 246, 548, 660, 506, 442, 288, 475, 282, 361]. lungs [1581]. luteinizing [209, 445]. Lyapunov [3079, 1516, 2018]. Lymph [2330, 3107]. Lymphangiogenesis [3178, 3226]. Lymphatic [2765, 3056, 2303]. Lymphocyte [2038, 2042, 2532, 1828, 1205, 598, 1893]. Lymphocytes [2810, 232, 1410, 1738]. lymphoid [1193]. Lymphoma [2387, 2894, 1925]. Lysis [2570]. Lytic [2887, 3539].

M [1646, 573, 1900, 1648, 3590]. MacArthur [3623]. Macdonald [2274]. Machinery [3251, 735, 1582]. macromolecular [618, 177, 383, 405, 480, 947, 603]. macromolecules [1082, 759, 437, 429, 557, 907]. Macrophage [2955, 3589, 3052, 3276, 1998, 2173]. Macrophage-Based [2173]. Macrophages [2617, 2466]. Macrophytes [3132]. Macroscopic [2446, 2400, 321]. Maculinea [3411]. made [1685]. Madin [2081]. magic [807, 1104]. magna [3302]. magnetic [759, 274, 785]. magnitude [1328]. Maiden [2667]. mainland [1764, 2009]. mainland-island [2009]. Maintaining [2721, 2156]. Maintenance [3564, 2959, 2339, 1999, 1391, 1269, 1980]. Making [1825, 1410, 2561, 259]. mal [2279]. Malaria [2942, 2751, 2534, 2283, 2685, 2454, 3601, 2969, 2609, 2874, 2441, 2086, 2976, 2274, 2418, 2626, 3243, 3545, 2493, 2028]. Male [3274, 2208, 1758, 1833]. malignant [821, 677]. Mammalian [2495, 3415, 856, 964, 742, 442, 115]. Mammals [2011, 1097]. mammillary [15, 106, 217, 290, 341]. Man [1032, 15, 793, 920]. Management [301, 2150, 3181, 2468, 2763, 2603, 1123, 1120, 1927]. Maneuver [2607]. Manifolds [3478]. mansoni [2308]. Many [2389, 3224, 109]. Map [3336, 2434, 2411, 1497, 1951, 1191]. MAPK [1942, 1567]. Mapping [2114, 1154, 1719, 1291]. mappings [619, 1879]. Maps [3200, 3334, 2623, 1454, 857]. Margaret [746]. Marine [2962, 2473, 1965, 2011, 3236, 1120]. Mark [2668]. marker [1686]. marking [1733]. Markov [626, 2061, 3561, 238, 658, 3560, 680, 1941, 415, 3483, 1887, 2006, 2676, 3222, 3464, 2027, 3314, 2237]. Markovian [1821, 1393]. Marquardt [3597]. marriage [1900]. MARS [2641]. Marsh [3220]. Martin [2224, 2397]. Marx [2832]. Mass [2324, 3055, 2847, 3196, 806, 13, 1416, 589]. Massachusetts [1900]. Massimo [1847]. Mast [2716, 1766, 1284]. Master [2837, 3131, 2326, 795]. matched [61]. matches [1078]. Matching [3465, 867, 1460, 2181, 1905, 784, 1106, 1004, 1330]. Mate [3311, 3133, 3216, 1970]. Matematicheskie [124]. Material [1428, 1429, 690, 3628]. Mathematica [2143]. Mathematical [1263, 2014, 3358, 149, 841, 941, 1001, 1159, 2035, 3598, 3523, 2611, 17, 3175, 3357, 2323, 2282, 107, 2799, 1414, 600, 2060, 3579, 2843, 3125, 2178, 2317, 1355, 3160, 2236, 3406, 2052, 3454, 2371, 2752, 3001, 2283, 2685, 2940, 144, 346, 1609, 3590, 3193, 3091, 3303, 257, 2061, 63, 2742, 2785, 3262, 3290, 2454, 2101, 2071, 3392, 232, 3506, 2560, 2955, 3153, 3350, 2508, 2632, 3184, 2300, 2629, 3261, 2746, 35, 197, 2492, 3297, 3329, 3524, 3042, 2642, 1477, 2919, 2216, 3241, 2929, 2588, 3169, 1678, 2366, 2084, 1686, 2573, 3054, 3589, 2659, 3352, 3244]. Mathematical [3077, 2513, 3046, 3052, 3360, 3081, 3384, 2183, 2194, 760, 30, 3139, 2530, 3468, 3504, 3100, 2094, 2830, 817, 1787, 2471, 2537, 3276, 3342, 3140, 2450, 1703, 1741, 2582, 2969, 501, 2827, 2950, 3274, 2991, 1472, 3223, 1362, 2000, 2862, 2965, 2429, 1753, 2733, 2686, 2652, 2474, 3259, 3599, 3127, 1169, 2428, 3338, 2989, 2901, 2943, 3437, 2628, 2961, 3152, 1473, 3215, 3167, 3355, 2872, 2781, 3368, 3176, 1911, 3070, 3246, 3151, 2531, 2646, 2081, 374, 504, 3204, 3211, 972, 2891, 2555, 1038, 3444, 3101, 2097, 2483, 3208, 2879, 2672, 1899, 2932, 2266, 1945, 2793, 2418, 2820]. Mathematical [2689, 3266, 2294, 2295, 1966, 3045, 3595, 3571, 3328, 3476, 329, 3099, 3284, 3341, 3516, 2466, 3112, 2327, 1621, 2947, 3226, 2700, 2493, 2308, 2792, 396, 2931, 1421, 1534, 1883, 259, 55, 1737, 1972, 542, 1406, 1725, 278, 1767, 2062, 1501, 258, 1891, 1670, 1681, 899, 1567, 1147, 370, 1341, 1948, 1444, 1897, 312, 741, 1071, 1826, 201, 486, 1302, 1771, 303, 380, 1111, 1112, 1113, 484, 717, 1732, 1224, 499, 1858, 837, 1807, 441, 1307, 827, 986, 1141, 1215, 1288, 1936, 1974, 1410, 1554, 1862, 1103, 1418, 1404, 161, 355, 293, 284, 1925, 185, 289]. mathematical [907, 647, 956, 1754, 1775, 865, 1829, 412, 1885, 414, 602, 1040, 587, 984, 1854, 1048, 661, 221, 79, 1920, 2459, 124, 1940, 1802].

Mathematical-Modeling [2891]. Mathematically [3611]. Mathematicians [3213]. Mathematics [3568, 1032, 1101, 3041, 1802, 3013, 3040, 3567, 1264, 77, 417, 24, 1900, 1377, 225]. Mating [3416, 3274, 3422, 2027, 2533, 1999, 1980]. MATLAB [2136, 3030]. Matrices [3137, 1436, 929, 1149, 1879, 1895, 456, 780, 1128]. Matrix [2403, 3454, 3588, 2456, 2825, 3162, 1450, 171, 336, 1482, 63, 192, 1687, 180, 1351]. Mats [3587]. Matsuda [3616]. Matters [3362]. Maturation [3480, 2757, 2550, 2387, 2451, 3123, 1613]. Mature [3621]. Maturity [2468]. Maximal [2361, 2528]. maximization [5]. maximizing [1182]. Maximum [2288, 2566, 2569, 2424, 3494, 1828, 3312, 3463, 2429, 1163, 3470, 3148, 1469, 1848, 1162, 1026, 1466, 936]. Maximum-likelihood [1163]. may [1674]. Mayr [1716]. McCulloch [1066, 2517]. MCF [3279, 2672]. MCF-7 [3279, 2672]. McKendrick [1110, 135]. MCMC [1887]. Mdm2 [2703]. Me [3624]. Mean [2290, 2705, 2394, 1789, 5, 1529, 1437, 171]. Meandering [3316]. meaning [17, 1921]. Means [2920, 1923, 133, 128, 1967]. Measles [3098]. Measure [2811, 3601, 3116, 889, 1653, 1480, 991, 1405, 1201]. measured [1136, 267]. measurement [22, 1680, 395, 711]. Measurements [2620, 2120, 1973, 759, 1727, 479, 616]. Measures [2840, 3272, 2981, 3209, 1011, 752, 1853, 2070]. Measuring [84, 230, 234]. Mechanical [2877, 2908, 3218, 3293, 343, 2342, 182, 1430, 885]. Mechanical-Statistical [2342]. Mechanically [2266, 1870]. Mechanics [2819, 3448, 1992, 3361, 2072, 641, 1500, 1841, 1252, 37]. Mechanism [2313, 3518, 2209, 2041, 2515, 3536, 3272, 3366, 2135, 2065, 3071, 3296, 2523, 2524, 2168, 1942, 1265, 1302, 540, 947, 14, 1815, 687, 114, 1155, 6, 1362, 1807, 1016, 1152, 1303, 1165, 1713]. Mechanisms [2004, 3367, 2722, 2369, 3606, 2037, 2679, 3508, 2481, 3299, 2005, 2281, 2088, 3073, 471, 267, 1173, 1886, 1298, 1888, 1281]. Mechanistic [2108, 2423, 3382, 1654, 1234]. Mechano [2285, 2195]. Mechano-Electric [2195]. Mechanobiology [2820]. Mechanochemical [2433, 2687, 3456]. mechanoelectrical [1493]. mechanogram [335]. Media [2824, 3043, 2578, 2467, 3510, 1585, 1324]. Median [3137]. mediate [1284]. Mediated [2571, 2752, 2769, 3592, 2679, 2349, 2991, 2489, 3147, 2495, 3150, 3266, 3341, 1763, 1205, 2000, 1901, 579, 2389, 3114]. mediates [1983]. mediator [1215]. Medical [3487, 17]. Medical-Resource [3487]. Medicine [1101, 2061, 3041, 17, 122, 2062, 1891, 1264, 841]. Medium [2039, 935, 1747, 724, 409, 830]. Medulla [2429, 330, 1302, 46]. medullary [1152]. Meeting [149, 562, 850, 1075, 3628]. Meinhardt [3214, 2504, 1016]. Melanin [2105, 1408, 1419]. Melanoma [3475]. Melanosomes [2105]. Melees [2536]. mellitus [625]. membered [1140]. members [1659]. Membrane [2722, 2071, 2716, 2648, 1599, 2095, 2450, 2058, 2703, 2413, 646, 802, 667, 639, 1037, 1784, 1065, 1124, 1301, 304, 817, 180, 885, 222, 368, 404, 419, 1811, 709, 12, 11, 42, 1885, 400, 1303]. Membrane-Associated [2095, 2450]. membrane-biophysical [12]. Membranes [2237, 68, 365, 810, 457, 493, 640, 1746, 102, 101, 612, 1366, 90, 91]. memories [1015]. Memory [3323, 3547, 2666, 2785, 1018, 3484, 2114, 1277, 1337, 731, 1369, 1124, 1698]. Mendelian [1072]. Meniscal [2832]. menstrual [1783]. Menten [450, 786, 2743, 487, 3237, 682, 538, 535, 3484, 1026, 1952, 2054, 628, 1827]. Menten-style [1952]. mer [3467]. Meristem [3564]. mesenchyme [1860]. Mesendoderm [3125]. mesoderm [1457]. Mesophyll [3076]. message [1082]. Messenger [2782]. messengers [1763]. Metabolic [328, 2361, 2071, 2358, 3434, 2406, 2961, 3048, 3099, 1576, 1020, 1668, 1791, 870, 1684, 1877, 45, 1856, 329, 1502]. Metabolically [3221]. Metabolism [2148, 3157, 2967, 2686, 3152, 2002, 1533, 289, 871]. Metabolite [2932]. Metabolites [2438, 1924]. Metabolizing [2489, 2593]. metacontrast [354]. Metaecoepidemic [2547]. Metal [3523]. Metapopulation [3522, 3327, 2106, 2020, 2618, 1439, 1654, 1538, 1564, 1574, 1623, 1688, 1498, 1282]. Metapopulations [3171, 2190, 2123, 2100, 2168, 1764, 1461, 1765, 1614, 2009, 2010]. Metastability [3536, 979]. Metastable [1649]. Metastases [3362]. Metastasis [3340, 993, 1404]. Metastatic [3362, 3524, 2775]. Metazoan [2082, 1778]. Meth [2764]. Methane [2052, 733]. methane-producing [733]. Method [2670, 3543, 3162, 2610, 3098, 3168, 2058, 3578, 3558, 2778, 3096, 2932, 3548, 2176, 2487, 2152, 1905, 764, 1954, 333, 1594, 822, 1274, 270, 1897, 56, 1465, 1304, 413, 70, 73, 260, 1790, 468, 819, 628, 1243, 1934]. Methods [3552, 2343, 1646, 3461, 2802, 2424, 3030, 2122, 3289, 2537, 3271, 2221, 3255, 2046, 3275, 3337, 1190, 1585, 1396, 794, 1227, 2130, 1620, 747, 1027]. metrazol [770]. metrazol-induced [770]. Metric [3210, 3529, 3146, 754]. Metrics [1497, 3381, 3339, 732, 2022]. MGB [1027]. MI [1853]. Mice [2234, 2438]. Michael [1930]. Michaelis [2743, 450, 786, 487, 3237, 682, 538, 535, 3484, 1026, 1952, 2054, 628, 1827]. Michell [1581]. Micro [2459, 875, 1287, 1854]. micro- [875]. micro-array [1854]. Micro-environment [2459]. micro-organisms [1287]. Microarray [3069]. microbe [334]. Microbial [2880, 2675, 2937, 720, 1712, 1192, 1317, 1608, 1786, 1211, 521, 1267, 1502]. Microbiota [3306]. microcirculation [993, 975]. Microcolonies [3221]. Microcontinuum [684]. microcurie [15]. microcurie-days [15]. Microdomain [2716]. Microenvironment [3365, 2170, 2775]. Microevolution [1054]. microexovesicles [1599]. Microfibril [3185]. Microfluidic [3496, 3095]. microglia [1808]. microorganism [105]. microorganisms [986, 410, 1232]. microparasite [1857]. Micropharmacology [3575]. Microscopic [2446, 2400, 983]. microsphere [24]. Microstructure [3361, 189]. Microtubule [2041, 2992, 3477, 3437, 2953, 1731, 1290]. Microtubule-Bound [2992]. Microtubules [2613]. Microvascular [3066, 3224, 2391, 2793, 181, 1289, 1379]. middle [714]. Miekisz [340]. Migrating [3037]. Migration [2776, 2877, 2998, 2446, 2516, 2661, 1991, 2812, 2825, 2560, 2904, 2451, 3291, 2192, 2000, 3349, 2375, 3135, 3208, 2206, 2133, 2754, 3111, 1489, 1147, 449, 1852, 687, 1253, 582, 853, 1574, 1677, 1413, 1250, 2010, 1040]. Migrations [2177]. Migratory [2770, 2325]. mild [496]. milestones [1915]. milk [1483]. Milling [2344]. Milner [1118]. mimic [658]. mimicry [159]. minicomputers [875]. Minimal [3089, 3235, 2462, 2736, 385, 2972, 3123, 258, 1517, 54, 576]. minimization [856]. Minimize [2670]. minimizing [624]. Minimum [1082, 1309, 2592, 2619, 2490, 1935, 1790]. minus [807]. Miocene [1054, 1443]. Mis [3237]. Mis-Charged [3237]. Mis-Reading [3237]. Misactivation [2758]. mismatch [748]. Mistletoes [2583]. mite [1334, 979]. Mitochondria [2836, 62, 297]. mitochondrial [1361]. Mitosis [2613, 3199]. Mitotic [3244, 84, 1222, 1752]. Mitropolskii [333]. Mixed [1650, 2080, 3518, 2144, 2275, 3093, 1689]. Mixed-Culture [3093]. Mixed-Mechanism [3518]. Mixed-up [2275]. Mixing [2307, 2924, 2365, 141, 30, 946]. Mixture [2649]. Mixtures [3462, 2275, 1653, 1024, 1442, 589]. Mobility [2721, 2601, 1782, 1226]. Mode [783, 2036, 567, 2768, 2737, 1734]. Model [2668, 3253, 2004, 2245, 3622, 2241, 2509, 2998, 3321, 3332, 2207, 2128, 2944, 3126, 3021, 2329, 3501, 3523, 1995, 2611, 2148, 3327, 2506, 2355, 3256, 2449, 3469, 3564, 3496, 2036, 3306, 3383, 2169, 2853, 3488, 2670, 3175, 2112, 3330, 2596, 2671, 2323, 2269, 3364, 2755, 2277, 2247, 3089, 2437, 2741, 2485, 843, 2080, 3537, 2060, 2896, 3198, 3264, 2665, 2178, 2317, 2330, 3580, 3626, 3160, 2409, 2001, 2828, 3406, 314, 2052, 3454, 2534, 2770, 2752, 2848, 3001, 3049, 3170, 3305, 3399, 2066, 2617, 2283, 2685, 2940, 3210, 2262, 2767, 2923, 3590, 3107, 3193, 3091, 3303, 3458]. Model [3617, 2061, 3387, 2473, 2863, 2825, 2996, 3028, 3262, 3290, 2454, 3448, 2452, 3392, 2486, 2560, 2634, 3062, 2955, 2949, 2934, 2019, 2632, 3536, 2300, 2629, 2265, 2320, 3261, 3301, 2746, 3205, 2968, 2492, 3319, 2904, 724, 2882, 3042, 2106, 2324, 2642, 2126, 2068, 2298, 2710, 3624, 3239, 2444, 2565, 3169, 2074, 2366, 2903, 2973, 2084, 2564, 2422, 2817, 2475, 2638, 3398, 2765, 3056, 2448, 2167, 2056, 2053, 3038, 3346, 3348, 3366, 3245, 3589, 2016, 2379, 2578, 2699, 2684, 2165, 3077, 2692, 2693, 2451, 3046, 3052, 2610, 3081, 3289, 2696, 3268, 2183, 3377, 2194, 3414, 3601, 3499]. Model [3224, 2674, 3309, 2314, 2459, 2823, 2365, 3504, 3419, 2426, 2958, 3098, 3136, 3236, 2170, 2974, 3100, 3155, 2740, 3382, 3088, 923, 3434, 2336, 3199, 2349, 2713, 2224, 3065, 3276, 3342, 3140, 2450, 3416, 2239, 2663, 1990, 2440, 3213, 2479, 2582, 3078, 3546, 2892, 2969, 2455, 3508, 3336, 2827, 2991, 2107, 2866, 2489, 2344, 3221, 3492, 3279, 3619, 2645, 3500, 3277, 2000, 2862, 2874, 2438, 2065, 2429, 2759, 3147, 3231, 2285, 2682, 3520, 3059, 2733, 2345, 2494, 2930, 2401, 2686, 3295, 2597, 2124, 2664, 3259, 2881, 2411, 2441, 2180, 2622, 3127, 2364, 2187, 2687, 2512, 2643, 2976, 2120, 3082, 3333, 3203]. Model [3179, 2156, 2495, 3024, 1993, 2078, 2284, 2481, 3503, 3057, 3164, 3356, 2768, 2020, 3076, 3150, 2901, 2943, 2964, 3437, 2961, 3152, 3215, 3483, 3094, 3026, 3355, 2872, 2781, 3368, 3270, 3413, 2416, 2796, 3093, 2728, 3070, 3246, 2057, 2008, 3519, 3023, 2531, 2646, 2352, 2081, 3204, 2274, 3211, 2050, 3135, 2814, 3096, 2555, 2970, 2114, 2691, 2727, 3146, 2590, 3102, 2497, 2672, 3415, 2181, 2758, 2110, 2159, 3600, 2132, 2158, 3255, 2985, 2932, 2226, 2841, 1487, 2676, 3222, 3371, 3426, 2125, 2972, 2351, 1716, 3325, 3591, 2689, 3092, 2480, 2294, 2295, 2064, 2341, 2027, 2945, 3032, 3595]. Model [925, 3571, 3328, 3476, 2372, 3616, 3188, 3048, 2787, 3061, 3099, 3119, 3243, 3341, 3396, 3487, 3516, 2466, 3112, 2327, 2947, 2760, 2773, 3226, 3095, 2335, 3027, 1996, 2700, 3607, 1165, 2217, 3569, 2270, 2493, 2735, 3498, 2975, 2278, 2625, 2214, 2869, 3182, 2719, 2098, 2212, 2308, 3063, 3123, 3539, 3421, 2931, 433, 1263, 1421, 68, 1764, 365, 533, 1276, 421, 458, 1412, 1105, 1504, 1867, 262, 1946, 259, 55, 1539, 1426, 15, 281, 328, 508, 1892, 498, 363, 104, 530, 1414, 1881, 1175, 600, 793, 1972, 1718, 542, 1805, 656, 1406, 1571, 1652, 620, 719, 487, 1099, 291, 167]. model [169, 241, 1145, 1499, 242, 500, 1609, 179, 1783, 357, 866, 278, 1334, 401, 1172, 1667, 1767, 2062, 63, 489, 1056, 1501, 164, 93, 898, 258, 1670, 1671, 1777, 232, 1681, 860, 1373, 143, 899, 1567, 1712, 425, 376, 430, 1654, 229, 1147, 1561, 324, 370, 1274, 1689, 330, 1444, 844, 1199, 1729, 580, 99, 100, 1926, 655, 1839, 1090, 1244, 1095, 1488, 60, 431, 1826, 449, 1299, 1517, 1573, 75, 564, 486, 1397, 1398, 1302, 1450, 230, 557, 1771, 658, 235, 607, 546, 362, 472, 380, 484, 717, 1734, 1401, 906, 1234, 1429, 1732, 1192, 715]. model [1761, 127, 1224, 499, 1363, 1815, 1437, 743, 634, 792, 740, 795, 6, 245, 501, 272, 250, 1453, 1166, 761, 885, 1472, 1351, 1146, 1393, 1807, 1820, 593, 217, 986, 549, 604, 1886, 1604, 1141, 1215, 1288, 455, 497, 1811, 613, 1327, 190, 1535, 709, 1410, 638, 1611, 1554, 1496, 1862, 1878, 1403, 742, 1103, 1240, 1241, 1591, 700, 1719, 1297, 1418, 1632, 975, 1404, 534, 168, 1311, 130, 271, 1131, 1376, 1178, 515, 954, 293, 1806, 1841, 284, 1925, 511, 2067, 907, 1785, 985, 647, 366, 712, 474, 545, 1310, 358, 1283, 297, 581, 1642, 1546]. model [1623, 1688, 1624, 292, 507, 957, 1440, 1890, 1957, 1985, 880, 725, 1107, 412, 1510, 112, 1885, 1967, 1937, 193, 625, 859, 917, 1016, 1152, 1300, 1408, 1419, 1952, 1809, 1869, 1287, 938, 347, 1490, 1625, 1466, 1040, 1242, 335, 223, 1917, 1140, 1189, 1275, 1320, 1481, 316, 1910, 1854, 598, 979, 1338, 1157, 1047, 79, 1293, 3353, 1503, 1863, 583, 19, 2305, 2310]. Model-Based [2004, 3333]. model-building [19]. model-mimic [658]. Modèle [104]. Modeled [3418, 2050, 3137]. Modeler [3482]. Modeling [3010, 2819, 2469, 3145, 3441, 1208, 1541, 2015, 2202, 3247, 3535, 3294, 3311, 2460, 2264, 2402, 2709, 2080, 3593, 2236, 2959, 2328, 2371, 1849, 3157, 3317, 1542, 1959, 2076, 2061, 2515, 3300, 2949, 2163, 2911, 800, 3489, 3507, 1435, 3297, 3329, 2357, 2106, 2786, 2136, 2929, 2782, 2301, 3398, 2883, 2398, 3477, 2711, 3521, 3486, 2898, 2227, 3360, 3565, 3459, 3139, 2530, 2258, 943, 2567, 2094, 3533, 2224, 2680, 2921, 2415, 2583, 3113, 3003, 2900, 3132, 3223, 2118, 3407, 2965, 2880, 3304, 2675, 2937, 2302, 2498, 3599, 2003, 2164, 2628, 1193, 1651, 2208, 2614, 3614, 3151, 1738, 2354, 2714, 2836, 2891]. Modeling [3570, 3439, 1600, 3525, 1620, 2342, 2860, 3514, 1592, 2286, 2111, 2826, 2266, 3266, 2260, 2002, 2463, 3344, 1866, 3250, 3545, 2089, 1601, 1888, 2381, 1997, 3254, 1700, 3376, 3282, 440, 2248, 2368, 2472, 2370, 2237, 2697, 2690, 3180, 1247, 2062, 390, 1477, 741, 864, 1686, 1906, 701, 1493, 874, 1530, 1744, 817, 1703, 1741, 416, 876, 23, 865, 1290, 1513, 2105, 2771]. Modelling [2103, 841, 941, 1214, 3424, 1958, 3474, 1645, 3194, 3621, 2536, 3125, 2895, 2091, 1991, 2831, 2962, 3283, 2802, 1817, 2030, 1306, 3234, 1714, 1977, 1523, 2116, 2101, 2182, 3051, 3322, 2909, 2025, 2157, 3524, 2102, 2519, 2024, 3446, 2630, 1838, 2659, 2798, 2513, 2453, 3384, 2905, 3050, 3611, 2830, 2635, 1050, 3274, 2609, 3380, 1736, 2047, 1735, 3039, 2941, 2474, 2433, 3608, 3195, 2764, 2151, 2956, 2435, 2989, 3167, 1209, 2894, 2443, 2570, 2599, 3101, 2097, 3208, 2529, 1751, 2478, 3532, 2801, 3033, 2161, 3553, 2220, 2703, 3375, 1578, 2754, 2145, 1938, 2173, 2195, 1939, 2723]. Modelling [74, 1502, 783, 1883, 651, 1737, 1640, 1821, 1891, 1810, 1948, 1332, 413, 2031, 1753, 1068, 1832, 1925, 1708, 1621, 1975, 1095]. Models [2927, 2915, 2014, 2776, 3202, 2704, 3551, 2811, 173, 2403, 3490, 122, 3178, 3090, 3391, 3420, 3534, 2196, 2875, 3579, 2307, 3025, 2843, 3475, 2311, 2408, 3543, 2808, 3397, 2131, 2812, 2138, 2785, 2337, 3252, 3153, 1518, 2508, 3184, 1519, 2108, 2423, 2501, 2162, 1920, 2386, 2893, 2595, 2919, 2273, 2572, 2210, 2155, 2396, 2390, 2400, 2347, 1384, 3517, 2750, 3460, 2839, 2594, 2573, 2465, 2476, 2568, 3141, 3561, 3530, 2039, 2655, 2258, 2601, 2761, 2063, 2018, 2179, 2331, 3463, 2717, 3382, 2816, 2889, 2471, 2397, 2641, 2299, 2013, 2736, 2350, 2591, 1315, 1552, 2965, 2058, 3315]. Models [2007, 3037, 3075, 2652, 2602, 2433, 3505, 2491, 2377, 3351, 2375, 3620, 2857, 2105, 2499, 2784, 3578, 2649, 1691, 3302, 3029, 2580, 3436, 3166, 2957, 2505, 3296, 2738, 2510, 2006, 2532, 3464, 3363, 2820, 3174, 3405, 3045, 2753, 2995, 524, 2547, 3284, 2482, 3573, 3370, 2293, 2261, 2034, 2772, 1922, 475, 2214, 3623, 3275, 3337, 2427, 3602, 1726, 649, 1534, 131, 1923, 884, 927, 1617, 107, 333, 1589, 1825, 690, 1951, 1355, 1138, 1579, 1555, 1585, 1007, 145, 1396, 1755, 509, 958, 1676, 1210, 1563, 1380, 659, 1630, 1694, 1882, 438, 2029, 1749, 1638, 736, 1204, 270]. models [1915, 312, 348, 1828, 919, 960, 1538, 1368, 1200, 1470, 2051, 834, 1525, 1784, 1431, 926, 214, 1873, 1308, 669, 1397, 437, 383, 1551, 95, 238, 1294, 1872, 1508, 1874, 852, 304, 591, 946, 1787, 1304, 977, 924, 799, 650, 1307, 408, 556, 39, 867, 680, 1058, 1860, 307, 89, 142, 1804, 1941, 1613, 622, 674, 568, 161, 355, 1537, 1914, 797, 527, 243, 1887, 191, 374, 504, 848, 912, 2009, 588, 239, 1378, 1009, 442, 657, 689, 345, 1526, 1829, 1655, 1605, 323, 1945, 887, 1798, 1927, 589, 1966, 3, 1445, 805, 329, 1303, 1566, 478, 661]. models [930, 1844, 964, 1128, 3120, 1900]. modern [1169]. Modernization [3444]. Modes [2520, 3511, 2173]. Modification [2059, 1644, 459]. Modified [3149, 3466, 2645, 459]. Modularity [1160]. Modulated [2593, 2266, 2656]. Modulation [2103, 2933, 2876, 1573, 3625]. module [1780]. modules [731]. modulus [1575, 910]. mold [1242]. Molecular [2843, 2037, 687, 2491, 13, 119, 2222, 3296, 3400, 3373, 2413, 3275, 3337, 1641, 1658, 813, 102, 760, 432, 1256, 101, 1713]. Molecule [2513, 1739, 893]. molecules [693, 268]. molluscan [580, 116, 203]. Moment [1638, 2053, 3036, 2701, 3281, 3034, 3033, 2181, 3430, 1960]. Moments [1969, 797, 674]. Monitor [2174]. Monitoring [2002, 1387, 908]. mono [457, 493]. mono- [457, 493]. Monocyclic [2059]. Monod [1244, 1166]. Monogamy [2749]. Monolayer [2801]. Monomial [3596]. Monostationarity [3541]. monotherapy [1474]. Monotone [1132]. Monotrichous [2379]. Monovalent [3084]. Monte [3552, 784, 190, 554, 2598, 1211]. Morbidelli [1847]. Morphed [2471]. Morphoelasticity [3563]. Morphogen [2669, 2095, 2615, 2450, 2936, 529]. Morphogen-Regulated [2615]. morphogénèse [123]. Morphogenesis [935, 2312, 2867, 2040, 2072, 1069, 1916, 1931, 123, 367, 1350, 1736, 1068, 300, 859, 3627]. Morphogenetic [2230, 173, 406, 61, 1674]. morphology [1209]. morphometric [1400, 514]. Morphostatic [2245]. mortal [1180]. Mortality [2963, 3300, 3457, 3115, 1563, 1872, 613, 1728]. Mosquito [3622, 3546, 3396]. Mosquito-Borne [3622]. Mosquitoe [3622]. Mosquitoes [3189, 2685, 1181]. Most [2963, 3300, 3165, 220]. Moth [2242]. Mother [2187]. Mother-to-Child [2187]. mothers [1178]. Motif [2399]. Motif-Based [2399]. Motile [2379, 733]. Motility [1991, 2127, 2360, 2351, 1747, 1367, 1839]. Motion [3496, 2589, 1746, 2535, 3129, 2413, 433, 1139, 1659, 1230, 1231, 1232, 1560, 368, 404, 419, 81]. motions [115]. Motivated [3217]. Motor [2246, 2612, 2491, 3400, 2953, 1550, 1682, 402, 428]. Motor-Based [2491]. motor-nerve [1550]. Motors [3275, 3337]. Mound [2232]. Mound-Building [2232]. Mountain [3062, 2440, 2841, 1467]. Mouse [3140, 1766, 817]. Mouth [2570]. Move [2008]. Movement [2942, 3522, 3588, 3582, 3181, 2595, 2698, 2426, 2739, 3380, 2862, 2047, 2333, 3430, 2145, 3182, 659, 1835, 1327, 1624, 1511]. Movements [2564, 3606]. Moves [3404]. Moving [3194, 3082, 3094, 3532, 1385, 92, 1880]. Moving-Habitat [3094]. MPC [15]. MPSC [1136]. MRI [3162]. MRSA [3081, 3289]. MS2 [512]. much [1861]. Mucin [3570]. Muco [2107, 2158]. Muco-Ciliary [2158]. Mucus [3592]. Multi [3004, 3441, 3480, 3391, 2661, 2196, 1482, 2388, 2632, 2320, 3565, 2905, 2740, 3168, 3604, 2814, 3086, 2293, 3314, 273, 74, 25]. Multi-cell [3565]. Multi-City [2740]. Multi-compartment [2196, 25]. Multi-Host [3604, 2905]. Multi-Instability [2632]. Multi-level [3004, 3168]. multi-neurone [74]. Multi-Objective [2388]. Multi-player [1482]. Multi-scale [3441]. Multi-species [2661, 273]. Multi-Stability [2632]. Multi-stage [3391, 3086, 3314]. Multi-State [2293]. Multi-strain [2814]. Multi-Type [2320, 3480]. Multibreath [506]. multicell [1253, 1677]. Multicellular [3332, 3125, 3361, 2145, 1307]. Multiciliary [2278]. Multicompartment [2438, 2947, 611]. multicompartmental [671, 662]. multicomponent [589]. Multidimensional [2207, 2128, 2712, 2675, 384]. Multidrug [2665, 2903, 2787, 1215, 1288]. Multifractal [3431]. Multifunctionality [1791]. multifurcations [801, 1096]. Multigeneration [2637]. Multigroup [3436]. Multilayer [2970]. Multilayered [3564]. Multilevel [3554]. multilimb [1262]. multilocus [848]. Multimodal [3376, 611]. Multimodality [2878]. Multiparametric [2002]. Multipatch [2816]. Multiphase [3095, 1805]. Multiple [2776, 2559, 3451, 2536, 1783, 3515, 2742, 3582, 3287, 3301, 2565, 854, 1603, 3066, 338, 2331, 2582, 3279, 2777, 954, 1221, 2483, 3375, 1798, 2480, 3248, 2088, 2461, 3119, 3487, 3569, 2975, 2948, 1905, 389, 1190, 1139, 1177, 1227, 1803, 1872, 1679, 408, 556, 1460, 269, 1035, 87, 1243]. Multiple-Component [3287]. Multiple-Relaxation-Time [2975]. Multiple-Target [2483]. Multiplex [3424, 2588]. multiplication [1302]. Multiplicity [2005, 2281]. multiplier [18]. Multipole [1634, 785]. Multiscale [3537, 2328, 2911, 3559, 2638, 3377, 3561, 3504, 3434, 3065, 2620, 2047, 3304, 3215, 3578, 2691, 3208, 2478, 3532, 3375, 2482, 3318]. Multisite [3008, 2852]. Multispecies [3321, 2207, 2128, 3039, 1274, 605]. Multistability [2444]. Multistable [2349]. Multistage [3028, 3601]. Multistationarity [3541, 3596, 3509, 3491, 2852, 3540, 2864]. Multistationary [2978]. Multistrain [2516, 3068]. Multitype [2652, 2551, 1089, 797]. Multivariate [2519, 228, 1443, 1984]. multivibrating [146]. Murine [2611, 2617]. Murphy [2645]. Murray [1900, 1940]. Mus [2954]. Muscle [2729, 2645, 2614, 2087, 433, 764, 49, 102, 1488, 431, 1982, 57, 1457, 388, 1284, 1829, 335, 37, 522, 349, 463]. Muscular [3589]. Mussels [3122]. mutant [1672, 997, 1584]. mutants [161]. Mutation [2748, 2559, 2202, 3474, 3089, 401, 2068, 3315, 2124, 2431, 2407, 3225, 1148, 1848, 268, 1768, 1669, 1333]. Mutation-absorption [401]. mutation-selection [1848]. Mutations [2355, 3534, 2657, 2903, 2695, 1273, 1867, 1851, 564]. mutual [800, 379, 704]. Mutualism [2074, 1275]. Mutualisms [3011]. Mutualistic [2868, 582]. mycelia [1797]. mycelial [1308]. Mycobacterium [2089]. Mycorrhizal [2599]. Myelogenous [3364, 2486, 2634, 2258, 2314, 2553]. Myeloid [2352]. myeloma [269]. Myocardial [2400, 1886]. Myocardium [2266, 497, 543, 627]. Myocytes [2475, 1580]. myogenic [1298]. Myomas [3001]. myosin [1739]. myotonia [293]. Myrmica [3411]. Mystery [2205, 2428]. Myxobacteria [2040, 2029].

N [2952, 1877]. N-site [2952]. NADH [1668, 267]. NADH-binding [267]. nanobeads [1746]. Nanomechanical [2589]. Nanotubes [2826]. narrow [164, 9]. Nasal [3081, 3289, 353]. Nash [3088]. National [3426]. Natural [473, 3340, 3530, 2958, 2858, 1714, 1071, 1862, 1785]. Naturally [2471]. nature [10, 148]. Navier [565]. Navigation [2548]. Near [2249, 2832]. Near-Critical [2249]. nearest [80]. Necessary [1709, 2864, 112, 2291]. Neck [2684]. necrosis [1934]. Necrotic [2759]. Nectarless [2815]. need [1861, 1970]. Needed [2566]. Negative [2824, 3393, 3193, 3336, 2474, 2925, 1897]. Negatively [2370]. negligible [495]. negotiations [51]. Neighbor [2619, 3468, 80, 1935]. Neighbor-Joining [3468, 1935]. Neighborhoods [3005, 1005, 3031]. Neighboring [2230]. Neighbour [3191]. Neighbour-Dependent [3191]. neo [1934]. neo-vascularization [1934]. Neolithic [3414]. neoplasia [258, 284]. neoplastic [295]. Nephron [2756]. Nephrons [2389, 2349, 1901].

Nernst [3170, 3305, 400]. Nerve [3323, 180, 2012, 1434, 206, 240, 470, 365, 421, 458, 1550, 242, 326, 802, 62, 4, 539, 35, 1964, 1065, 199, 14, 536, 687, 356, 244, 42, 59, 531, 1770, 36, 446]. nervous [179, 1067, 225]. nest [1511]. Nested [2658, 3396]. Nestedness [2868]. nests [1094, 1095]. Net [2152]. nets [1066, 1576, 585, 85, 541, 688, 765, 555, 891]. Network [3469, 3424, 3198, 3264, 2966, 2296, 2993, 2791, 3173, 2399, 3218, 2513, 3404, 2476, 2696, 2520, 3511, 2977, 2365, 2658, 2868, 2768, 2756, 3022, 2057, 2303, 3540, 3436, 3409, 2523, 2910, 3548, 3408, 2735, 3080, 1276, 2099, 1019, 1436, 1030, 1259, 1260, 1556, 1151, 2051, 762, 827, 1941, 1240, 1241, 146, 195, 1320, 1293]. Network-Based [2365]. Networks [3144, 3015, 2175, 2496, 3079, 3196, 3418, 1208, 2288, 3541, 2121, 2802, 3399, 2115, 2840, 2805, 2983, 2666, 3069, 2586, 2996, 2673, 2979, 3509, 3051, 3497, 3149, 3491, 2257, 2647, 2386, 3335, 3278, 2710, 2807, 2633, 2309, 3559, 2573, 2852, 2711, 3402, 3218, 3531, 2922, 3224, 3374, 2971, 3550, 3560, 3165, 2868, 2604, 1647, 3584, 2391, 2784, 3299, 2340, 2718, 3540, 3096, 3110, 3086, 2477, 2545, 2864, 2676, 2987, 2793, 3556, 2549, 2072, 2524, 2995, 2032, 2213, 2430, 2606, 2844, 3242, 1018, 1017, 1823, 1791, 619, 1963, 471, 1424, 702, 1332, 1417, 1753, 1198, 1319]. networks [1454, 1669, 937, 422, 423, 1620, 1087, 1132, 1333, 1661, 200, 1277, 1346, 1345, 1104, 1938, 1699, 1844, 1337, 74, 3485]. Neural [1208, 2661, 2834, 2296, 2791, 2696, 2520, 2521, 3176, 2860, 259, 1374, 542, 566, 1066, 750, 394, 1151, 1865, 1332, 113, 965, 1081, 891]. Neuregulin [2671]. Neuritic [1290]. Neuro [3046]. Neuro-Oncology [3046]. Neuroblastoma [3398]. Neurochemical [2421]. Neurodevelopmental [3197]. neuroelectric [507]. Neurological [2519, 1247]. Neuron [2848, 3049, 35, 702, 762, 1761, 45, 66, 67, 1407]. Neuronal [2673, 1343, 2625, 1743, 706, 1235]. neurone [74]. neurones [1640]. Neurons [3594, 2319, 2517, 197, 580, 1705, 1131, 1376, 1806, 957, 1027]. neurophysiology [666]. neuroreceptor [930, 983]. Neuroscience [3517]. Neurotransmitter [1206, 2037, 1328, 1136, 1582, 1643, 1713]. Neurovascular [2782, 3517]. Neutral [2273, 2390, 3230, 1454, 2340, 2772, 1669, 1699, 1649]. Neutrophil [3193, 2955]. Neutrophils [2200]. Newly [2021]. Newman [3]. news [1956]. Newtonian [108, 466]. Next [3357]. Neyman [607]. NF [2972]. NF- [2972]. Niche [2689, 384, 492]. Nicholas [1899]. night [918]. Nile [1977, 2013, 1990, 2458, 1972, 2385, 2353, 3353]. nitrate [1428]. Nitric [2590, 1751]. Nitrification [2880]. nitrogen [1944, 1309, 1428]. NMDA [2596, 2671]. NMR [1911]. No [2731, 1018, 297, 2732, 1361]. nodal [1771]. Node [2330, 2161, 835]. Nodes [3107, 738]. Noise [2036, 3383, 2896, 1875, 2616, 2907, 3497, 3550, 2914, 2943, 2057, 2268, 2656, 2154, 2370, 1743, 1806, 818]. Noise-Induced [3383, 2616, 2057]. Noisy [3393, 2120]. Nombre [2234]. Non [3079, 3196, 3418, 641, 903, 3420, 2709, 3618, 1585, 2191, 2090, 672, 3465, 2508, 2562, 2162, 2018, 3382, 2095, 2450, 792, 1241, 3356, 3023, 3284, 3487, 783, 1484, 959, 262, 720, 1748, 754, 1274, 72, 48, 768, 1530, 1430, 1050, 1423, 914, 700, 251, 1914, 1925, 366, 1330, 466, 414, 1386, 1303, 461, 616, 1027, 900]. non-aligned [1330]. Non-Autonomous [2191]. non-coding [1484]. non-demolition [616]. Non-equilibrium [3079, 903, 2562]. non-equimolar [414]. Non-equivalent [3465]. Non-explosivity [3418]. Non-exponential [3284]. non-Hodgkin [1925]. non-homogeneous [1430]. non-ideality [1303]. Non-inherited [2709]. non-lattice [1914]. Non-linear [641, 672, 792, 783, 959, 262, 1050, 914, 700, 251, 366, 1027]. non-linked [1748]. Non-local [3420, 3356, 1386]. Non-mass [3196]. Non-mechanistic [3382]. non-metric [754]. Non-periodic [3618, 48]. Non-phenomenological [3023]. non-predator [720]. Non-receptors [2450]. non-repeatable [461]. non-sessile [1423]. Non-smooth [3487]. Non-spatial [2508]. Non-standard [1585]. non-stationary [768]. non-steady-state [1274]. non-stochastic [1530]. Non-uniformities [1241]. Non-valued [900]. Nonautonomous [2336, 2212]. nonelectrolyte [368, 404, 419]. nonelectrolytes [102, 101]. Nonequilibrium [2789]. nonexistence [460]. Nonlinear [2518, 1255, 1516, 3454, 2500, 2884, 2465, 2674, 1953, 2179, 1292, 2739, 2936, 3078, 386, 1318, 3223, 623, 517, 507, 2677, 1211, 2098, 3220, 1934, 3275, 3337, 890, 889, 533, 646, 158, 286, 420, 465, 1881, 1355, 860, 143, 1630, 586, 834, 172, 829, 557, 418, 516, 481, 155, 1107, 275, 917, 628, 1910]. Nonlinearities [2042]. nonlinearity [912]. Nonlocal [3327, 2041, 3057, 2064, 3341]. Nonmyelinated [2012]. Nonnegative [3162]. Nonoverlapping [3230]. Nonparametric [3600]. Nonrandomness [2069]. nonsteady [234]. Nonuniform [3533, 1755]. Norm [2550]. Normal [163, 3065, 3276, 2768, 2928, 2359, 2430, 295, 1653, 1590, 497, 677]. normalization [685]. North [850, 2151, 3084]. Northeast [505, 1910]. Nosocomial [2196, 3025, 2567]. Notch [2449]. Note [2875, 120, 2166, 2725, 2012, 644, 631, 1499, 322, 24, 76, 1964, 71, 837, 217, 491, 515, 13, 318, 773, 32, 461, 661]. notes [225, 374]. Notification [3145]. Notocarinovalva [1443]. Notz [1646]. Novel [3579, 3543, 2907, 2309, 2568, 1960, 3608, 2529, 1767, 1339, 1438]. November [1840]. NP [1278]. NP-hard [1278]. Nuclear [2703, 2910, 1903, 759]. Nucleations [2992]. Nucleic [40, 39]. Nucleotide [2254, 1848, 1359, 1078, 1790, 2026, 1035]. nucleotides [762, 1125]. Nucleus [2722, 3361]. Number [2288, 3522, 2141, 2403, 3306, 3260, 2566, 2870, 2023, 3465, 2016, 2845, 3271, 2206, 2133, 920, 807, 96, 1078, 118, 1896, 459, 2605]. Numbering [636, 819]. Numbers [2924, 2254, 35, 1104]. Numerical [2877, 1355, 2586, 2386, 2453, 3463, 1660, 1305, 2391, 3166, 1432, 2382, 2793, 2059, 223, 1566, 2775, 3232, 2427, 1572, 1585, 172, 1304, 924, 1745, 3120]. Nutrient [2887, 3582, 3221, 2901, 2964, 2599, 1747, 1099, 806, 1244, 1041]. nutrient-limited [1041]. Nutritional [2946].

O [1877, 1840, 93]. Oakley [746]. Obituary [1932, 3214]. Objective [2388]. objectives [1471, 867]. Obligate [3617]. observable [1618]. observables [260]. Observation [305, 784]. Observations [2564, 115, 672]. observed [214, 1788]. Obstacles [3430]. Obtained [2397, 1975, 1958, 1786]. occasional [801, 1096]. Occasionally [3172]. occlusion [629]. Occlusions [3099]. Occupancy [2100, 3572, 2184, 2689, 1997, 1538, 599, 696]. occur [801]. occurs [545]. Ocean [1324]. Ocular [1993, 2723, 2062]. ODE [2923, 3319, 3504, 2481]. odour [592]. Off [2587, 2336, 704, 1871, 1607, 2772]. Official [3145]. Offs [2556]. Offspring [3163, 2206, 2133]. Oleg [1819]. olfactory [1715, 1151, 1705, 1407]. olfactory-like [1151]. Oligocene [1443]. On-site [2850]. Oncogenic [2944]. Oncology [3358, 3046]. Oncolytic [3384, 3539]. One [2767, 2138, 1155, 3313, 766, 117, 3445, 1713, 8, 1773, 958, 1463, 467, 931, 1678, 586, 546, 1437, 1220, 408, 556, 1317, 1613, 3597, 424, 176, 698, 661, 1752, 283]. one-compartment [546, 408, 556]. one-compartmental [661]. One-Dimensional [2767, 766, 958, 1437, 1220, 176]. one-hit [1773]. one-pass [1613]. one-predator [698]. one-predator-two-prey [1678]. One-vesicle [1713]. One-way [1155]. Onset [2387, 2614]. ontogeny [1682, 1986, 1224]. open [76, 154]. Operation [3059]. Operator [3331, 3597]. operators [728, 828]. Operon [2271, 2005, 2281, 3268, 2891]. opiate [817]. Opioid [3535]. Opportunistic [1981]. optic [14]. Optical [3287, 2948]. Optima [3148]. Optimal [888, 2702, 2458, 2316, 3273, 2890, 3593, 2232, 2001, 2940, 2840, 1822, 3091, 3458, 2454, 3142, 1795, 1588, 2528, 2919, 1079, 2610, 3419, 2680, 1351, 2498, 1768, 3203, 3024, 2678, 1167, 2218, 2981, 3135, 2891, 3612, 3422, 1498, 269, 872, 2626, 1282, 1269, 1336, 332, 2461, 2242, 2034, 2795, 3390, 890, 1005, 1650, 1257, 1088, 1049, 1228, 95, 1428, 1522, 1476, 1543, 1205, 970, 503, 588, 385, 625, 1893]. Optimality [3016, 2619, 98, 1812, 1988, 624]. Optimization [2920, 3202, 2436, 2388, 2189, 3205, 2885, 2065, 2406, 288, 1281, 3064, 1469, 1668, 1791, 1735, 128, 391, 1877, 997, 1184, 396]. Optimized [886]. Optimizing [2075, 718]. Optimum [920]. Oral [2102, 2570]. Orbits [3433, 2539, 2808]. Order [3126, 2174, 3366, 1647, 451, 133, 1390, 1963, 1861, 797, 481, 1661, 858, 88, 1922, 1042]. order-dependent [858]. ordered [1154, 1281]. Ordering [2992]. Orderings [3005, 3031]. orders [1583]. Ordinary [2359, 3363, 1935]. organ [884]. Organisation [2147, 2679, 2505]. Organismic [34, 148, 473]. Organisms [2800, 3326, 1160, 1230, 1729, 1423, 1287]. Organization [2041, 2716, 3141, 2352, 1497, 470, 389, 1583, 1160, 776, 1415, 1389, 1280, 2043, 1610, 432, 119, 366, 603]. Organizing [2163, 1414, 471]. Organoids [3376]. organs [56, 1193]. orient [1232]. Oriental [1354]. orientalis [998, 486, 1354]. Orientation [1230, 1231, 1232, 2650, 1501]. orientational [1911]. Orientations [3185]. Oriented [3267]. Origin [2189, 3438, 866, 1380, 1451, 1587]. origin-dependent [1587]. originating [1573]. Origins [2793]. orthopedic [1525]. oscilations [1259]. Oscillating [2464, 928, 488]. Oscillation [2925, 2525, 2526, 3249, 1938]. Oscillations [2282, 2616, 2770, 3518, 2966, 2732, 2043, 3066, 2582, 2731, 2604, 2441, 2421, 2057, 2523, 2793, 3044, 2088, 2546, 3498, 205, 1353, 1145, 1260, 312, 1573, 48, 939, 120, 266, 499, 1050, 517, 1432, 1316, 1642, 612, 1528]. oscillator [216, 779, 84]. Oscillators [2334, 2621, 783, 651, 1462, 1374, 741, 272, 413, 567, 360, 507]. Oscillatory [2978, 2634, 3043, 915, 2056, 3258, 2239, 3114, 2728, 261, 1512, 206, 240, 229, 1491, 1695, 1362]. Osmolarity [2723]. Osmotic [2989, 990, 989, 441]. Osteoblastic [2835]. Osteoblasts [2822]. Ostracoda [1443]. Other [3490, 224, 164, 750, 759, 1853, 309, 424]. our [3444]. Outbreak [3062, 2637, 2816, 2440, 2342, 2260, 3421, 979]. Outbreak-Recovery [3062]. Outbreaks [2516, 2959, 2385, 3062, 3395, 2353, 2737]. outbursts [1608]. Outcome [3369, 2894, 2762]. Outcrossing [3422]. Output [2670, 56, 1437]. Ovarian [2091, 2883]. Over-Consumers [2788]. Over-Expression [2629]. Overall [2092, 1870]. overelastic [849]. Overexpression [3537, 2300]. Overlapping [2008]. Overview [2757, 2294, 2295]. ovum [271]. oxalacetate [1361]. Oxford [3628]. oxidation [1725, 1342]. Oxide [2590, 1751]. Oxygen [2729, 2492, 1729, 2204, 198, 3101, 724, 370, 682, 920, 484, 717, 359, 727, 1967, 385, 572]. oxygen-chemotactic [727]. Oxygenation [2623, 571].

P [2143, 3279, 1215, 1288, 956, 1039]. P-glycoprotein [1215, 1288]. P-Glycoproteins [3279]. p53 [1672, 1584, 2703]. Pace [2346]. Pacific [1172]. packaging [1821]. packets [1314]. Packing [3565, 248]. page [150]. pages [1930, 2143, 1940, 1899, 1840]. Pair [3390]. Pairs [2604]. Pairwise [1830, 2784, 1896]. Palette [3506]. Panama [1973]. Pancreatic [2282, 2071, 3365, 3521, 2088, 3249, 1892]. Pandemic [2241, 2263, 2680, 3118, 2529]. Paper [1001, 516, 350]. Paperback [1840, 1646, 1920]. papers [1118, 393, 1114]. Papillomavirus [3259]. Papovaviridae [1966]. Parabolic [3420, 279, 1492]. Paracrine [3512, 1141]. Paradigms [3449]. Paradox [2769, 2172, 882]. Paradoxes [2592]. paradoxical [560]. Parallel [1434, 2670, 3301, 2331, 1506, 810, 1595]. Paralysis [3070, 3246]. Parameter [1975, 1958, 2460, 2445, 3453, 2408, 2828, 3069, 2705, 1292, 3217, 970, 3578, 2580, 3597, 921, 2261, 1881, 390, 546, 613, 161, 1510]. Parameter-dependent [970]. Parameterise [3475]. Parameterised [3309]. Parameterization [2306]. Parameterized [3325]. Parameters [2193, 2538, 2283, 3467, 3205, 2165, 2302, 2124, 2804, 2803, 2773, 179, 357, 1973, 1564, 514, 1786, 599, 696, 1038, 1513]. Parametric [2569, 2540, 3258, 1559, 2899, 2542, 1312, 126, 1847]. Parametric-equation [1559]. paramétriques [126]. Parametrization [3596, 1545, 1597]. Parametrizing [3493]. Parasite [2859, 3542, 3607, 3411, 1458]. Parasites [3133, 3216, 1611]. parasitic [1176]. parasitism [1769, 1557]. Parasitoid [2411, 2217, 1555, 1476, 1441]. parasitoids [1185]. Parathyroid [1615]. Pardee [1415]. Parental [2367, 1758, 1833]. parents [1180]. parietal [908].

Parkinsonian [1556]. Parotid [3503, 3476]. paroxysmal [770]. Parsimonious [2569, 3292, 3034, 1640]. parsimoniously [1349]. Parsimony [2566, 2424, 3494, 3312, 3354, 3148, 1466]. Part [122, 3262, 3290, 1229, 2692, 2693, 1397, 1398, 2628]. Partial [3435, 3561, 451, 1865, 1732, 1291]. partial-thickness [1732]. Partially [3582, 2325, 1763, 1855, 879]. Particle [2716, 58, 2413, 1746, 165, 593, 1660, 118, 475]. Particles [2916, 2917, 2862, 2111, 660, 534, 568, 474, 282]. particular [1097]. particulate [444, 164, 443]. Partition [2131]. Partitioning [2817, 1336]. pass [1613]. Passage [2290, 3323, 2705, 2613, 3007, 2333]. passe [48, 33]. passe-bande [48, 33]. Passerine [2325]. Passive [102, 3606, 196, 101, 1407]. Past [2046, 865]. Pasteels [1648]. Patch [3588, 2892, 3500, 2100, 3513, 1764, 1654, 1538, 1137, 3353]. Patch-Matrix [3588]. patch-occupancy [1538]. Patches [3522, 2153, 3313, 2073]. Patchiness [2217, 1690]. Patchy [2311, 2698, 3074, 3061, 1589, 1579, 1555, 1882, 1919, 1547]. Path [2607, 2528, 1468]. Pathogen [3406, 3212, 3277, 3525, 2480, 620, 719, 1872]. Pathogenesis [3589, 1590, 1592]. Pathogens [2774, 3280, 2117, 2362, 2678]. pathological [717]. paths [798, 1649]. Pathway [3265, 1817, 2934, 3350, 2324, 3106, 2227, 2758, 2972, 93, 1567, 870, 1533, 1591, 1877]. Pathway-dependent [1817]. Pathways [2449, 2539, 3301, 2358, 2782, 2406, 2703, 2656, 2480, 3569, 810, 1415, 1668, 1719, 937, 1342]. Patient [3401, 3046, 3368, 3371]. Patient-Specific [3401, 3046, 3368, 3371]. Patients [3410, 2714, 1981, 1788, 685]. Pattern [2245, 814, 2571, 3161, 2897, 1755, 3506, 3554, 1463, 1994, 2732, 2669, 2884, 1262, 2674, 2504, 3217, 2874, 2990, 3502, 2733, 2914, 2731, 2364, 1496, 2079, 2584, 2268, 1587, 748, 3308, 1330, 2841, 859, 3092, 3045, 2211, 2168, 749, 3423, 1572, 911, 219, 1602, 1174, 338, 418, 1745, 1146, 529, 1860, 1068, 1164, 547, 1055, 971]. Patterning [2953, 1621]. Patterns [2264, 2999, 2514, 2163, 3536, 2190, 2123, 2082, 3478, 2615, 3230, 2434, 2433, 2867, 2662, 2555, 2797, 3495, 3220, 1773, 430, 406, 1256, 730, 653]. Pau [1159]. Pauline [1802]. PBPK [2098]. PCR [1186]. PDE [2314, 3504]. PDGF [3372, 3512]. PDGF-Driven [3372]. Peak [2549, 559, 560, 611]. Pearl [872]. Pecten [1228]. Pediatric [2567]. Peer [3444]. penalties [752]. penalty [1386]. PENDISC [2932]. Penetration [1478, 1703, 1741, 409]. pentose [1533]. People [2630]. peptide [1826]. peptides [642]. perception [1650, 219, 1055]. Perceptual [2696, 643]. Perennial [3216]. Performance [2117, 2675, 2937, 391, 2891, 1875, 1039]. performed [1143]. perfused [103, 321]. perfusion [111, 230, 1886]. periaxonal [494]. peridic [257]. Peridomestic [2439]. perifused [1362]. Period [2437, 3062, 2415, 3243, 202]. Periodic [3433, 2887, 2141, 2403, 2665, 3618, 2808, 231, 2767, 2337, 1676, 2710, 2698, 2582, 434, 3271, 971, 2787, 2335, 1783, 65, 376, 430, 1341, 1603, 48, 392, 699, 694, 781, 1938, 319]. Periodical [3492, 1528]. Periodically [2685, 2233, 544, 1404]. Periodically-Forced [2685]. periodicity [1384]. periodiques [48]. Periods [3490, 3072]. peripheral [496, 1902]. peristaltic [227, 310, 81, 468]. Permanence [3236, 2719, 1087]. permeabilities [1152]. permeability [990, 919, 960, 829, 1301, 885, 356, 495]. permeable [102, 826, 108, 185]. Permeation [2058, 829]. Permutation [2539]. Permutations [3449]. Permutoassociahedron [3235]. Persistence [649, 3522, 2318, 3432, 3588, 1576, 1261, 3280, 439, 736, 3417, 2581, 3316, 1430, 977, 2635, 3158, 2374, 2085, 3277, 2682, 2859, 2854, 2243, 3094, 2706, 880, 2984, 3426, 3011, 3602, 1931, 1177, 800, 988, 1564, 1557, 1445, 2467]. Persistent [3566, 3529, 3318, 1528]. Persisters [2767]. Personal [3419]. Perspective [3202]. Perspectives [3568, 303]. persuasion [51]. Perturbation [2743, 1341, 3373, 143, 1459]. Perturbations [2292, 2705, 1762]. perturbed [544, 1368, 1221]. Perverse [2484]. Pest [2140, 2150, 2388, 2907, 2946, 2216, 2858, 2342, 2603, 2484, 1927]. Pest-Management [2150]. Pesticide [2858, 3619, 1871]. Pesticides [2193]. pests [332]. PET [3572, 2120]. pH [1849]. pH- [1849]. Phage [2887, 2958]. phagocyte [1480]. Phantom [2313, 1892]. pharmacodynamic [1426]. Pharmacodynamics [3370]. Pharmacokinetic [2438, 3569, 2214, 1726, 1373, 527, 559, 560, 611]. pharmacokinetics [868, 1521, 1948, 1744, 579]. Pharmacostatics [3576]. Phase [3327, 2169, 3198, 3264, 535, 3082, 2012, 1467, 2869, 68, 640, 334, 1766, 369, 1316, 1507, 855, 1503, 1528]. Phase-Field [2869]. Phase-plane [535]. phase-shift [1528]. phases [1948]. Phenological [3378]. Phenologically [2223]. Phenologically-Structured [2223]. Phenology [2368, 2472]. Phenomena [2632, 264, 833, 565, 1211, 1081]. Phenomenological [2336, 1353, 707, 3023]. Phenomenon [3526, 1094, 1095, 16]. Phenotype [2657, 3366, 2885, 2921, 96]. Phenotypes [3451]. Phenotypic [2790, 2667, 3003, 3333, 1985]. phenylcarbamoyl [1876]. Pheromone [2758]. pheromones [1511]. Phloem [2989]. phocine [1523]. phonon [62]. phosphate [1533]. Phospholipid [2593, 483]. Phosphorus [3501]. Phosphorylation [3008, 3518, 2952, 2852]. photo [363]. photo-enhanced [363]. Photobleaching [2075]. Photocurrent [2860]. Photoelectric [486]. Photoinhibition [2818]. Photon [178, 62]. Photoreceptor [1773]. photosynthesis [1214, 1234]. Photosynthetic [3228, 133, 410]. Phototaxis [2299]. Phototransduction [1059]. PhrA [2513]. phyllotactic [1868]. Phyllotaxis [1350, 847]. Phylogenetic [3462, 3233, 2343, 2983, 2558, 3467, 3335, 3485, 3173, 3278, 2309, 3381, 2783, 3472, 3404, 2977, 3339, 2542, 2275, 2490, 2649, 2359, 3110, 3409, 2676, 3222, 3473, 2032, 2213, 2430, 2844, 2487, 3242]. Phylogenetics [3461, 2424, 3471]. Phylogenies [2845, 3466, 2575, 929, 819]. Phylogeny [2566, 1405]. phylograms [54]. Physical [3520, 2143, 1154, 242, 1385, 1561, 16, 991, 1674, 152, 1869]. physical-biological [1561]. Physician [2535]. physico [130]. physics [1819, 519, 776, 225]. Physik [776]. Physio [289]. Physio-chemical [289]. Physiological [868, 3205, 2260, 1959, 16, 567, 385, 589, 1048, 987].

Physiologically [1726, 3166, 2214]. Physiology [841, 3047, 467, 519]. Phytoplankton [1842, 2818, 2092, 1944, 1717, 1873, 1234, 1428]. Picking [2845]. Picture [3336]. Piebaldism [3506]. Piecewise [2808, 2849, 2995, 1844]. piecewise-linear [1844]. Piezoelectricity [4]. pig [1897]. Piglets [2318]. Pigmented [2105]. pigmentosa [2890]. pigs [1810]. Pine [3062, 2440, 2841, 1467]. Pinning [2745]. Pinus [2025, 2157]. Pioneer [1930]. Pitts [1066, 2517]. Pituitary [2323, 2929, 2883, 2226, 209, 445]. plan [1778]. Planar [2443, 1562]. Planck [3170, 3305, 511, 1250, 400]. plane [535, 1752]. Plankton [2092, 3101, 3123, 1099, 1563, 1324]. plankton-nutrient [1099]. planning [865]. Plant [1923, 2897, 2949, 2992, 3133, 3228, 2674, 2990, 2401, 2622, 2867, 2901, 3185, 3216, 3386, 2555, 1031, 2726, 2815, 2034, 1263, 928, 1589, 1882, 1986, 1265, 1238, 1314, 2073, 1350, 1804, 1917, 221, 1593]. Plant-herbivore [1923, 1263]. Plants [2264, 1845, 1315, 1218, 1269, 1336]. Plaques [3325, 1808, 1751]. plasma [1746, 1767, 667, 637]. plasminogen [855]. Plasmodium [2028, 2441, 3607]. Plasticity [2667, 3589, 2653]. Plateau [2561, 1344]. Plateaus [2226]. Platelet [2821]. Plausible [2156]. player [1482, 3609, 1979]. pleased [1001]. plexus [14]. plume [1287]. plumes [1628]. plurality [1217]. plus [368, 404, 419, 807, 2130]. pneumoniae [2259]. pneumonitis [474]. Podocytes [3355]. Poets [2600]. Poikilotherm [2552]. Point [3534, 2908, 2304, 3482, 3034, 3012, 666]. Points [2288, 195]. Pointwise [646, 860]. Poiseuille [81]. Poisson [3170, 3305, 1202, 154]. Pol [272, 413]. Polar [2557]. Polarisation [3420, 2914, 2745]. Polarity [2950, 2443, 173]. Polarization [2974, 2047]. polarized [339]. Policy [3206]. Polio [2235]. Pollen [1970]. Pollination [3617, 3011]. Pollinator [2555, 2726, 2815]. Polluted [2591, 3063]. Pollution [3446]. Pólya [508]. Pólya-type [508]. Polyandry [1212, 1833]. polyethism [1223, 1266]. polyhedra [1226]. Polyhedral [2558, 2636]. polymer [1148, 1453, 524]. Polymerase [3330]. Polymeric [3454]. Polymerization [3456, 1518, 1519]. polymers [80]. Polymorphism [2667]. Polynomial [3096, 447]. Polynucleotide [2790, 796, 2022]. polyphemus [3583]. Polytope [2619]. Polytopes [3235]. Populated [2963, 3300]. Population [2668, 2234, 3622, 3163, 2469, 1159, 1255, 3522, 2141, 2403, 1617, 679, 2843, 2962, 3283, 2897, 3594, 2840, 2131, 3257, 2191, 2090, 3588, 2138, 2742, 2177, 2108, 3181, 3301, 3395, 2083, 2813, 2126, 2572, 2644, 2750, 3036, 2556, 691, 722, 2325, 1117, 1196, 2520, 2905, 2924, 2467, 3382, 3546, 2866, 3492, 3599, 2364, 2086, 2976, 3190, 2078, 3356, 3172, 2243, 1070, 2218, 3023, 2354, 3302, 2431, 2842, 3439, 2360, 1627, 2984, 2382, 2125, 2027, 2260, 3396, 1663, 3111, 2270, 2034, 2653, 3183, 987, 333, 397, 1406, 1659, 1509, 47, 63, 202, 489, 509, 958, 1056, 898, 1676, 1857, 1563, 1147]. population [1749, 2660, 1909, 1973, 544, 1409, 1470, 192, 1603, 1906, 1450, 472, 1089, 1701, 1476, 1530, 939, 127, 713, 634, 1173, 977, 1351, 670, 1183, 1161, 1644, 1978, 1051, 737, 944, 132, 774, 912, 318, 337, 818, 689, 112, 1798, 932, 1698, 1140, 1425]. population-based [1644]. Population-Level [2840, 3257]. Populations [3021, 2077, 2250, 2193, 2169, 3420, 3267, 2918, 3610, 3432, 3585, 2376, 2468, 2347, 2425, 3139, 2144, 2276, 3230, 2880, 3037, 3002, 3333, 3057, 3216, 3240, 2092, 3211, 2706, 3166, 3102, 2295, 2242, 1943, 720, 1426, 769, 600, 542, 566, 1571, 69, 3515, 1329, 1391, 1654, 229, 1229, 1371, 1204, 99, 100, 1409, 1653, 1080, 1397, 1450, 238, 1679, 1220, 1271, 1611, 1608, 1413, 1120, 1557, 1115, 957, 424, 1907, 204, 1324, 1072, 1593, 1831]. pore [1766]. Pores [2049]. Poroelasticity [3367]. Porous [2127, 1213, 1585, 550]. Portfolios [2137]. Position [3449, 3027]. Positive [3493, 3484, 3336, 1449, 1267, 2033]. Possible [3176, 640, 184, 162, 1942, 540, 839, 947, 220, 1713]. Post [3098]. Post-vaccination [3098]. Postconditioning [3297]. posterior [324]. Posttranslationally [3149]. potassium [802, 821, 398, 494, 1378]. potency [1817]. Potential [3079, 2506, 2325, 3097, 3227, 3545, 421, 458, 2099, 242, 1567, 1488, 494, 388, 387, 90]. potentials [639, 400, 1407]. power [510, 837]. pp [1646, 1900, 1819, 1818, 1647, 1831, 1648]. Practical [2210, 3174, 3405, 1345, 1201]. Practice [3046, 683, 788]. Pre [3098, 3002, 3573, 277, 245]. pre-biological [277]. Pre-dimerised [3573]. pre-emption [245]. Pre-extinction [3002]. Pre-vaccination [3098]. prebiotic [1389, 480, 524]. Precautions [2196]. Precipitation [282, 120]. Precise [3432, 1434]. Precision [2548, 1814]. precisions [1186]. Precursor [2341]. Predation [2140, 3581, 2362, 3542, 3412, 2719, 1697, 1537, 521]. Predator [2419, 2757, 3257, 2473, 2177, 3171, 3085, 882, 2136, 3030, 2240, 2565, 2347, 3581, 2223, 3500, 3502, 3203, 3024, 3164, 3356, 2584, 3135, 2842, 2603, 2322, 2414, 3344, 3250, 3220, 3623, 1512, 160, 720, 381, 411, 1571, 631, 1334, 721, 156, 143, 299, 379, 704, 988, 1261, 270, 348, 502, 1368, 1678, 1610, 854, 617, 1904, 1551, 392, 658, 1836, 1422, 1352, 1363, 699, 2003, 1950, 1537, 1690, 1547, 1221, 605, 345, 880, 606, 815, 1498, 698, 1464, 298, 88, 482, 1358, 979, 1137]. Predator-Induced [3581]. Predator-Prey [2177, 2223, 2419, 882, 1512, 721, 1610, 854, 1904, 1836, 1422, 1363, 1950, 1358]. Predators [2419, 2231, 3285, 2188, 2435, 2603, 3422, 3250, 256, 332, 3428]. predatory [1588]. Predict [3579, 3077, 3520, 245]. Predictability [2076, 1849]. Predicted [2215]. Predicting [1884, 2808, 3062, 2829, 2083, 2109, 3368, 1915]. Prediction [2119, 296, 351, 2710, 312, 3366, 2733, 525, 3483, 2778, 3571, 1027, 755]. Predictions [2841, 2111, 1095]. Predictive [3234, 3398, 1553]. Predictor [2361]. predistance [1149]. Predominance [3341]. Preface [1063]. Preference [3603, 1654, 1315, 1008]. Preferences [3416]. Preferential [2365, 1428]. Preferentially [3422]. Pregnane [2489]. prère [129]. prère-sœur [129]. Prescription [3535]. Presence [3523, 2516, 2907, 2117, 2744, 3286, 3195, 3338, 2764, 3164, 2678, 2505, 2590, 3422, 3011, 2466, 2953, 1851, 1367, 1387, 1339, 332]. Present [3426, 2046]. presented [3628]. Preserve [2435]. presidential [303]. Press [1646, 1900, 1920, 2129, 1819, 1930, 1647, 1877, 2143, 1847]. Pressure [3367, 2969, 316, 431, 1590, 1417, 344, 1575, 1584, 1869]. Pressure-flow [316]. pressure-volume [344]. pressures [344]. pressurized [1213]. prestressed [1427, 916]. pretreatment [923]. Prevalence [2234, 3145, 2630, 2874]. prevascular [1091]. Preventing [2788, 2087]. Prevention [2940, 2444, 2758, 2792, 3183]. Preventive [3198, 3264, 3457]. Prey [2757, 3257, 2473, 2177, 3171, 3085, 2136, 3030, 2240, 2565, 2347, 2255, 2223, 3500, 3502, 2362, 3203, 2435, 3024, 3164, 3356, 1690, 2584, 3135, 2842, 2585, 2603, 2322, 2414, 3344, 3616, 3250, 2719, 3220, 3623, 1512, 160, 720, 381, 411, 2419, 631, 1334, 721, 156, 143, 299, 379, 704, 882, 988, 1261, 270, 348, 502, 1368, 1678, 1610, 854, 617, 1904, 1551, 392, 1836, 1422, 1352, 1363, 699, 2003, 1697, 1950, 1537, 1547, 256, 1221, 1137, 605, 345, 880, 606, 815, 1498, 698, 1464, 298, 88, 482, 1358, 979]. Preytaxis [3488]. Primary [3340, 2442, 3544, 592, 1801, 375, 1008, 1490, 1625]. Primer [3529]. Priming [2861]. primitive [1632]. Primitivity [456, 780]. Primordium [2998]. Princeton [1647]. Principal [3471]. Principle [2380, 2559, 100, 2462, 2601, 2591, 5, 1848, 705, 1026, 1790, 385, 1492]. principles [182, 365, 450, 533, 1394, 197, 1520]. Priori [2701]. Priority [2316]. prisoner [1183, 2190, 2123, 2405]. Pro [2691]. Pro-angiogenic [2691]. probabiliste [104]. Probabilistic [1946, 104, 2875, 2808, 1810, 2376, 2069, 1009, 908]. Probabilities [3321, 3609, 3558, 2706, 110, 673, 425, 521, 858, 653]. Probability [2577, 2816, 2971, 3277, 3313, 1284, 2695, 1545, 1597, 2356, 1273, 1130, 106, 1608, 399, 848]. probe [230]. Problem [3441, 3247, 2150, 3609, 2610, 3565, 3482, 2410, 2627, 646, 814, 1257, 489, 1634, 35, 467, 1149, 1112, 1113, 768, 1428, 92, 244, 532, 633, 576, 1059, 628, 1278, 482]. Problems [2920, 3443, 3007, 3516, 450, 786, 667, 236, 124]. problemy [124]. Procedure [2580, 1021]. Procedures [1186, 372]. Proceed [2927]. Process [3126, 2343, 2052, 2851, 2612, 2110, 2159, 2407, 3456, 3314, 523, 55, 334, 1229, 1371, 1478, 1202, 1803, 1739, 30, 1731, 10, 1655, 1708, 1917, 661]. Process-Based [2052]. Processes [3551, 3480, 2661, 3479, 2843, 2550, 3617, 2071, 2902, 2909, 2423, 2913, 2321, 3034, 2105, 3386, 3033, 3430, 3060, 3129, 2912, 2624, 2994, 2551, 122, 488, 357, 796, 666, 701, 839, 563, 1089, 906, 266, 415, 797, 677, 712, 601]. Processing [97, 1648, 1151, 901, 900, 807]. Processive [3008, 3400]. Produce [2442, 1130]. produced [912]. Producer [3501, 3613, 2964, 3620, 1664]. Producing [3262, 3290, 2092, 1668, 733, 1520]. Product [2496, 3196, 456, 1339]. Product-Form [2496, 3196]. Production [2052, 3133, 2462, 2836, 3569, 928, 1822, 1665]. Productivity [2761]. products [780]. Profile [1995, 2656, 242, 1228, 441, 652, 1386]. Profiles [3265, 2588, 3571, 1305, 339, 685]. Program [1077, 754, 759, 303, 875, 871]. programmable [594]. Programming [2919, 3047, 1954]. Programs [3282, 1600]. Progress [3523]. Progression [2629, 2084, 2063, 2672, 1584]. Progressive [2778]. Projected [2160]. Projection [2949]. Prokaryotic [2871, 1681]. Proliferation [2511, 3191, 3273, 2066, 2300, 2451, 3046, 3238, 3533, 2204, 2224, 1793, 2781, 3368, 3314, 1277]. proliferative [1418]. proline [14]. prolong [2028]. Promiscuity [2749]. promote [1692]. Promotes [3172]. Promoting [3457, 3254]. prone [1240, 1241]. Proof [1278, 171]. proofreading [1763]. Propagating [2012, 1390, 1164, 1305]. Propagation [2175, 3134, 3479, 3394, 2625, 1641, 1658, 1155, 447, 1237]. Properties [2445, 732, 3149, 2257, 192, 2179, 2331, 3130, 3308, 3408, 2430, 20, 136, 1060, 2051, 1570, 1784, 201, 486, 702, 189, 1982, 1874, 1780, 709, 1442, 1454, 447, 1754, 1775, 1349, 858]. property [1119]. Prophylactic [2017]. Proportion [2228]. proportionality [197]. proposals [37]. proposed [1173, 1362]. prostaglandin [1952]. Prostate [3362, 1849, 1849]. prostitutes [1258]. protandry [1212]. Protease [3010, 2881, 2714, 1890]. protecting [1971]. Protection [2751, 3419]. protective [2923]. Protein [3237, 3149, 3200, 3334, 3109, 2655, 2612, 2933, 3625, 3059, 2768, 3483, 2728, 2778, 2627, 2110, 2159, 3423, 3573, 1504, 1817, 2030, 1321, 1681, 1567, 1279, 1803, 757, 1543, 886, 1125, 751, 1911, 554, 642, 758, 292, 1008, 1278, 1006, 1819]. Protein-16.2 [3423]. protein-coding [1125, 1008]. Protein-Coupled [3573]. Proteins [2246, 2648, 2237, 1652, 808, 1903, 759, 1619, 1582]. prothrombin [501]. protocell [917]. Protocells [2904]. Protocellular [2156]. protocol [1948, 1925]. Protocols [3384, 3368]. prototype [2043]. prototypes [265]. Province [3486, 3421, 3413]. Provisioning [3206]. Proximal [3298, 2381, 1301, 108, 1300, 845]. Proximate [2445, 1746]. proximities [759]. PSA [1849]. PSC [1136]. Pseudo [82, 2226, 2561, 1570]. pseudo-knots [1570]. Pseudo-Plateau [2561]. Pseudo-sound [82]. Pseudochaotic [2355]. Pseudoknots [2267, 2225]. Pseudomonas [3251, 1426, 1670, 1732]. pseudoobscura [647]. Pseudopodia [2776]. Pseuod [139]. Pseuod-sound [139]. psoriatic [1751]. PTH [3574]. Ptitsyn [1819]. PU1 [2597]. PU1-GATA1 [2597]. Public [3593, 2017, 3206]. Publication [3444]. pulmonary [550, 1437, 190, 1289, 1379, 474]. Pulsatile [550, 2393, 2931, 2043, 684]. Pulse [3251, 539, 916, 52, 78, 716, 2838, 1546, 2795, 1875, 816, 1392, 849, 369]. Pulsed [2317, 1404]. pulses [1798]. Pump [2266, 821, 1215, 1288, 835]. Pumping [2393, 2194, 274, 128, 1483]. pumps [1667]. Punishment [2788]. pupil [1050]. pure [567, 1604, 3059]. purine [1273]. purine/pyrimidine [1273]. purines [1484]. purines/pyrimidines [1484]. Purkinje [3392, 1155]. Pursuit [502]. Pursuit-evasion [502]. Putative [2513]. PXR [2489]. Pyragas [3433]. pyramidal [1376]. Pyrimethamine [2751, 2454]. pyrimidine [1273]. pyrimidines [1484]. pyroelectricity [4]. Pyroptosis [3341].

QSAR [1876, 2109]. Qualitative [421, 458, 770, 3319, 2298, 668, 1493, 1728, 360, 3084, 1844, 1843, 1224, 252]. Quality [3534, 410]. Quantal [1488]. Quantification [3234, 1897, 3382, 1796]. Quantified [3131]. Quantify [2648]. Quantifying [2263, 2734, 2488, 1757, 810, 1057]. Quantitative [1747, 2489, 3014, 3572, 1186, 3325, 3591, 1625, 3017, 1065, 1854]. Quantity [2603]. Quantum [36, 199, 1430, 885, 616]. Quarantine [3261, 1776]. Quarnet [3402]. Quarteroni [2253]. Quartets [3469, 2977]. Quasi [3552, 3499, 3067, 2491, 2113, 2168, 3275, 3337, 1617, 1388, 1676]. Quasi-Local [2168]. Quasi-Monte [3552]. quasi-periodic [1676]. quasi-stationary [1617]. Quasi-steady [2491, 1388]. Quasi-Steady-State [3499, 3337]. Quasispecies [3154, 3387, 2714, 3146, 1140]. Quasistationary [2172]. que [33]. queen [1511]. queenless [1354]. Queueing [2967]. Queues [3200]. Quorum [3251, 2843, 3301, 3507, 3156, 3389, 2513, 2692, 2693, 1732, 3221, 3167, 2945, 3032, 2463, 3232, 1824, 1670]. Quorum-Sensing [2513].

R [1900, 792, 515, 2130]. R. [2129, 609]. R/S [2130]. R/S-Plus [2130]. R5 [2607]. rabbit [1521, 1209]. Rabies [3486, 2690, 2771, 1639]. race [1859]. Radial [2025, 2157, 2109]. radiata [2025, 2157]. Radiation [2269, 2460, 3590, 3329, 3370, 1377, 742, 1704, 474]. radioactive [474]. radioligand [930, 983]. radioligand-neuroreceptor [930]. Radioresistance [3151]. Radiotherapy [2920, 3369, 3368, 1795, 1098]. Rainbow [3379]. Rakers [2681]. ramets [2073]. Random [2332, 2131, 2209, 721, 2657, 570, 3566, 2127, 2453, 2997, 2892, 2954, 3500, 1297, 712, 1115, 3445, 2695, 2027, 2760, 2237, 1273, 381, 411, 129, 679, 1210, 673, 619, 585, 564, 426, 1106, 1460, 271, 797, 1330, 200, 347, 932, 765]. Random-coil [1297]. random-interaction [347]. randomly [931, 957]. Randomness [1647]. Range [2140, 2169, 3615, 2972, 3111, 1733, 1461, 1973, 1705, 1535, 1880, 1081]. Ranges [2929, 2008]. Rank [2326, 3597]. Rank-one [3597]. Rapid [2431, 3452, 1744]. Rapid-Mutation [2431]. Rare [3558]. Rat [3298, 2135, 2429, 2876, 3143, 1947, 1301, 923, 1815, 1300]. Rate [2748, 3163, 3551, 2403, 2743, 2193, 3543, 2361, 2465, 2321, 3078, 1182, 2120, 3096, 1624, 2985, 3048, 3225, 2152, 644, 1842, 357, 1210, 1680, 1229, 1371, 2660, 1799, 852, 1465, 408, 1417, 976, 318, 193, 559, 560, 1290, 572, 571]. Rates [2559, 2755, 3089, 2038, 3149, 1990, 2888, 3600, 3087, 1148, 1641, 1658, 1258, 379, 1174, 691, 722, 268, 838, 1344, 1768, 415, 1378, 410]. Ratio [2347, 3344, 1073, 1261, 1828, 1551, 1176, 479, 1167]. Ratio-Dependent [2347, 3344, 1261, 1551]. Rationality [3311]. Ratios [2637, 3274, 1849, 926]. Rats [2248, 2708, 2098]. Ray [2855, 3309, 31]. RdCVF [3160]. Re [2729, 1802]. Re-emerging [1802]. Re-evaluating [2729]. reabsorption [1300]. reaching [106]. react [1339]. reacting [453]. Reaction [2915, 2998, 2496, 3079, 3418, 2290, 2743, 2036, 2831, 3555, 2550, 2865, 1602, 2863, 2979, 3229, 3509, 2019, 3322, 3497, 2909, 2911, 3491, 2136, 2669, 2633, 3559, 2884, 2913, 3531, 2922, 3493, 3561, 2679, 2889, 2736, 2936, 2971, 3560, 2604, 2003, 1804, 2867, 2683, 2505, 2477, 3415, 2864, 3556, 2549, 3408, 2995, 2172, 2912, 3586, 2046, 3516, 649, 158, 205, 286, 814, 911, 569, 133, 1755, 376, 430, 1963, 7, 172, 538, 535, 762, 120, 1130, 1387, 10, 1860, 228, 1339, 1028, 700, 118, 132, 155, 185, 372, 1087, 1132, 1333, 275, 859, 587, 223, 373]. Reaction-Norm [2550]. Reactions [3447, 3287, 3319, 3484, 2113, 2948, 33, 363, 364, 33, 1205, 1837, 479, 1622, 1661]. Reactive [3286, 2297, 2441, 1338]. reactivity [1193]. Reactor [2608, 2880, 1894, 1949]. Reading [2535, 3237]. Real [2940, 2732, 2731]. Realistic [2150, 2520, 2874, 2687, 1526]. realistic-size [1526]. realizability [277]. Realization [285, 1]. Rearguard [3295]. Rearrangement [3470]. Rearrangements [3449]. reasonable [1030]. Rebecca [1900]. rebound [1378]. Recapture [2668, 2655]. received [280]. receptive [197]. Receptor [2596, 3159, 2770, 2716, 2679, 1705, 2489, 3573, 1680, 1763, 1327, 599, 696, 1407]. receptor-mediated [1763]. Receptors [2671, 2109, 2045, 2095, 885, 1817, 4, 1876, 1763, 2450, 280]. Reciprocating [3394]. reciprocity [1423]. recognition [308, 1543, 761, 748, 1055, 749]. recognizing [852]. Recolonization [2151]. recombinant [948]. Recombination [2189, 1499, 1781, 838, 1627, 91]. Recombinations [2760]. Recommendations [3303]. Reconciling [1538]. Reconstitution [336]. reconstruct [1049]. Reconstructibility [3466]. Reconstructing [3381, 2977, 3584, 2487]. Reconstruction [2569, 3292, 2358, 3312, 3354, 3503, 3476, 2032, 2213, 2844, 1668, 1405]. Recover [2566]. Recovery [3126, 2755, 3062, 2075, 1459, 2708, 1954, 1903, 1888]. Recruitment [3362, 2847, 3607, 3183, 1842, 1175, 1450, 1690]. recta [46]. recta-descending [46]. rectangular [560]. recuperation [1448]. Recurrence [3605, 2854, 1404]. Recurrent [3618, 2408]. Recursions [2006]. recycling [1428, 1429]. Red [2151, 1257, 1037, 652, 1178, 495, 1107, 572, 571]. Reduce [3163, 3458, 2707]. Reduced [2396, 3560, 2057, 421]. Reduces [3603]. Reducing [2258, 2006]. Reduction [2380, 2559, 3330, 2830, 3165, 2491, 3096, 3255, 2985, 3222, 193, 2780, 2370, 2152, 1964, 637]. Reed [2310, 1863]. Reefs [3124]. Reentry [3438, 1216]. Reference [2929, 651, 417, 1097]. Refined [2768]. reflections [148]. reflex [1541, 1050]. Reflexive [3135]. refraction [1129]. refractive [1662]. Refractory [3594]. Refuge [3132, 3344, 1334]. Refuges [2411]. regarding [224]. Regency [1075]. Regeneration [2943, 2097, 2869, 656, 996]. Regime [3498, 230]. Regimens [2215]. Regimes [1950]. Region [125, 2863, 93, 1821, 1815, 1222, 1514]. region-based [1815]. Regional [149, 171]. Regions [2316, 2230, 1504, 132, 1008]. Regression [1816]. regrowth [1816]. Regular [3200, 3055, 1004]. Regularity [2246]. Regularization [2445]. Regulated [2615, 1091]. Regulation [2060, 1991, 2770, 2848, 3298, 3258, 2576, 3374, 3136, 2974, 2921, 2348, 2164, 3076, 2525, 2526, 3249, 814, 1531, 314, 315, 1569, 471, 718, 715, 209, 1751, 445, 1375, 329, 1917, 1140, 1189]. Regulations [2319]. Regulatory [2509, 3015, 2288, 2802, 3069, 3051, 2386, 2710, 2807, 2711, 2442, 2005, 2281, 2910, 2846, 114, 1346, 1345, 1938, 1844]. Rehabilitation [3195]. Reinfection [3231]. reinforcement [793, 532, 633]. reinforcement-depletion [793, 532]. Rejection [3556]. Rejection-Based [3556]. Related [3423, 1066, 197, 102, 926, 1756, 1655, 19]. Relation [2605, 558, 2643, 2860, 206, 240, 1094, 1431, 344, 797, 956, 1039]. relations [13, 648]. Relationship [2104, 2753, 2772, 570, 1619, 95]. Relationships [1529, 381, 411, 1781, 657, 1098, 316]. Relative [2906, 2196, 3543, 3346, 3605, 1539, 822, 1871]. Relaxation [267, 2975]. relay [651, 1535]. Release [2392, 2560, 2037, 2858, 2180, 2088, 1328, 1136, 1667, 1766, 1206, 591, 923, 1344, 1237, 1165, 1582, 1643, 1713]. Relevance [3274, 2603, 527]. Relevant [2688]. Reliability [2548, 1160, 909]. Reliable [1360]. Remark [573, 595]. Remarks [2803, 432, 2804]. Remodeling [3204, 1652]. Removal [2702, 266, 358]. remove [840]. renal [330, 1299, 1302, 1304, 108, 29, 46, 1305, 520, 1152, 1303, 845]. renewable [720]. Renewal [2612, 3129, 3376]. Renewal-Reward [2612]. Renin [3355]. reorganization [1542]. Reorientation [42]. Reoxygenation [2269, 1584]. repair [1444, 909]. Repeatability [608, 609]. repeatable [461]. Repeated [2280, 1672]. Repeats [2069]. Repellent [3622, 3142]. Reperfusion [3297]. Repertoire [2785, 1715, 1785]. repertoires [1526, 2383]. Replacement [2202]. replicates [1700]. replicating [1544]. Replication [2790, 1567, 2321, 2809, 1669, 1859, 1333]. replication-competent [1859]. Replicative [2302, 1240, 1241]. Replicator [2748, 2594, 2462, 2794, 1830, 1661]. Reply [663, 2732, 537, 596, 609]. Reported [3145]. Representation [3451, 3401, 3176, 3470, 3314, 1559, 1153, 1896, 756, 86]. Representation-Based [3401]. Representation-Theoretic [3470]. Representations [614]. represented [420, 465, 815]. Representing [3589]. Repress [2298]. repressing [2777]. Repressors [2298]. Reproducibility [3441]. Reproducible [3450, 3442, 3444]. Reproducing [3612, 405, 429]. Reproduction [2605, 2141, 2403, 2199, 3306, 3260, 2232, 2562, 1229, 1371, 2637, 2016, 719, 1701, 1813, 307]. Reproductive [2924, 3416, 3271, 2976, 2146, 2152, 1470, 1185, 1322]. Repulsion [2988, 3057]. Require [2820]. Required [3522, 2488, 840]. Requirements [2007, 2]. Requires [3444]. resampling [2130]. rescue [1654]. Research [861, 3184, 3450, 3443, 3442, 3444, 876]. Resection [3340]. Reserves [2962, 1965]. Reservoir [3455, 2707, 2958, 1712]. Resetting [2226]. residence [15, 1210, 1789, 1440]. residence-time [1440]. Resident [2562]. residual [459]. Residue [2713, 2768]. Residue-Level [2768]. Resilience [3585, 1041, 1248]. Resistance [2942, 3175, 3294, 2709, 1908, 3153, 3389, 2903, 2194, 2858, 3003, 3457, 2881, 3152, 2186, 2714, 3207, 2822, 1723, 1959, 1471, 1448, 820, 1871, 1474, 181, 1215, 1288, 1370, 1956, 1957, 1985]. resistance-associated [1288]. Resistant [2117, 2678, 866, 1980, 1999]. Resolution [2955, 247]. Resolve [2209]. Resolved [3591, 1480]. Resonance [2174, 2899, 759]. Resonances [1806, 126, 126]. Resonant [1509, 2083]. resorption [1615]. Resource [1123, 3515, 1912, 3018, 2044, 2842, 3166, 2497, 2461, 3487, 2427, 3120, 720, 1676, 892, 1119, 1799, 1665, 1041]. resource-based [892]. Resource-Bounded [2497]. resource-consumer [1041]. resource-dependent [1676]. Resources [2639, 3022, 3623, 1121, 1170, 879]. Respective [2804, 2803]. Respiration [3027]. Respiratory [2408, 3446, 3276, 1723, 278, 93, 201, 2031, 79, 74]. Response [2269, 2460, 3621, 2828, 3406, 3317, 3588, 2688, 3069, 2785, 3507, 2240, 2444, 2565, 2442, 3434, 2933, 3625, 2615, 3276, 2663, 2582, 2969, 2991, 2866, 3380, 2333, 2441, 3544, 2481, 3510, 2532, 2466, 3254, 2248, 2708, 1685, 1742, 1531, 296, 351, 187, 308, 1369, 1801, 582, 1473, 545, 1967, 1506, 1338, 1859, 1893]. Responses [2292, 3574, 3206, 3114, 2549, 3099, 1884, 326, 394, 1651, 2028, 611, 1027, 2336]. Responsible [2004]. rest [1948]. Resting [3060, 242]. Restitution [2114, 2510]. restraints [1911]. Restricted [2578, 1535, 2606, 659, 1171, 1511]. Restricted-range [1535]. restriction [1191, 1644, 1291]. restriction-modification [1644]. restructuring [1866]. Result [3008, 3114, 1372, 1792, 563, 1430]. Resultant [2560, 179]. resulting [806]. Results [3175, 2196, 2273, 3346, 2305, 3075, 2069, 2289, 420, 465, 444, 409, 1925, 1620, 1057, 1358]. Rethinking [2034]. Reticulate [3584]. retina [296, 351, 370, 686, 1209, 1408, 1419, 1809, 1869]. Retinal [2611, 2733, 2105, 525]. Retinitis [2890]. retreat [1632]. Retrograde [2908]. Retrograde-Flow [2908]. retrospective [876]. REU [3357]. revealed [1737]. Revealing [2513, 3612]. Reveals [2515, 3437, 3376, 955]. Reversal [3624]. Reversals [2042]. Reverse [3042, 1310, 1697, 1890]. Reversible [3322, 3511, 3067, 167, 169, 241, 604, 142, 154, 325]. Review [2508, 2129, 1173, 2253, 3167, 1064, 2222, 2221, 2143, 3444, 3553, 1558, 1721, 574, 1646, 1760, 1549, 1900, 1636, 1740, 1920, 1774, 1635, 1819, 1540, 1657, 1930, 1711, 1505, 1818, 194, 1179, 1647, 1606, 1877, 1831, 1940, 1847, 1759, 1750, 1730, 1899, 1637, 1720, 1648, 1840]. Reviews [137, 1272, 949, 1666, 973, 1577, 1616, 1402, 1313, 1702, 1532, 1108, 1158, 1285, 980, 992, 1043, 1455, 1395, 1524, 1195, 1486, 1693, 1092, 1168, 1802, 1626, 1022, 1475, 1061, 1000, 1142, 1052, 1568, 1133, 1074, 1357, 933, 1364, 427, 1596, 1675, 1122, 1254, 1447, 1420, 1295, 1245, 464, 1084, 1219, 1325, 1150, 1331, 1515, 1495, 1586, 897, 1100, 1207, 1340, 967, 1411]. Revising [2721]. Revisited [2530, 2851, 3071, 2260, 3616, 3123, 1583, 1971, 882, 1166, 807, 119, 891]. Revisiting [3085, 2424]. Reward [2612]. Rewiring [3198, 3264]. RF [540, 957]. RF-cell [957]. Rheological [3130, 189]. Rho [2183]. Rho-Family [2183]. Rhodes [1920]. rhodesiense [2905]. Rhythm [2266, 1341, 1682, 1224, 647, 74, 583]. Rhythmic [998, 349]. Rhythmogenesis [2319]. rhythms [1462, 272, 360]. RI [2716]. Richardson [340]. Ricker [3383, 3336, 2270]. Rift [2959, 2786, 3174]. rigid [550, 826]. Rigorous [3611]. Ring [2834, 3342, 3437, 2624, 2994, 49]. Ring-Driven [3342]. rings [1315]. Rise [2985, 1214]. Rising [2012]. Risk [2944, 3042, 3446, 3346, 3077, 2425, 2724, 1179, 1810, 1038]. Risk-sensitive [1179]. Risk-Structured [2944]. risks [1704, 545]. Ritz [372]. Rivalry [2696]. River [3239, 2939, 1173, 2926]. Rivers [3316, 2085]. RLC [272].

RNA [2332, 3330, 2870, 2587, 1012, 1570, 2540, 2267, 2225, 2432, 2104, 3379, 1969, 1878, 1454, 153, 755]. Robber [2726]. Robust [3162, 2442, 2319, 1462, 1602, 1935]. Robustness [3534, 2452, 2807, 2490, 1791, 1494]. Rod [2379]. Rod-Shaped [2379]. rodent [1973]. Rodents [2014, 2231]. Rods [3160]. Roff [2221]. Rogue [2558]. Role [2721, 2819, 3568, 3441, 3537, 3160, 1489, 2515, 3507, 3365, 2216, 2229, 2513, 2639, 2135, 2921, 2635, 3457, 2880, 2401, 2411, 2003, 2208, 3151, 2303, 2631, 2206, 2133, 2603, 2220, 2910, 2087, 2418, 553, 575, 3045, 3328, 1407, 3545, 2822, 1999, 1943, 314, 864, 471, 839, 1352, 687, 1345, 1127, 1980]. Roles [2897, 2789, 2827, 2804, 2803, 619]. roller [1036]. Rook [1153]. Root [3309, 3270, 1336]. root/shoot [1336]. root/shoot-partitioning [1336]. Rooted [2783, 3404, 819, 1913]. Rose [1806]. Rosenzweig [3623]. Ross [2274]. Rotational [2374, 1232]. Rotavirus [2567, 3084]. Rotors [3438]. rough [1814]. Roughened [2377]. rounds [1672]. route [985]. Routes [2263, 3605]. Rule [2530, 1217]. Rules [2309, 3402, 3279, 1527, 1343]. Run [2374, 3129, 1645]. run-and-tumble [1645]. Run-Tumble [3129]. Runners [3205]. Running [3205, 47]. Rural [3021]. Russian [124]. rustica [1814]. RyR [1580].

S [1920, 1831, 2103]. S-Plus [2130]. S. [792, 515]. saccadic [179]. Saccharomyces [3069]. Saddle [2304]. Saddle-Point [2304]. saddles [1603]. safety [217]. Saharan [2724]. Salivary [3328]. Sally [1802]. Salmon [3313, 1173, 2667]. Salmonella [2070, 2900, 3064]. salt [714, 1152]. salts [309]. Same [3102]. Sampling [2663, 2598, 8, 1565, 1646]. San [976]. sapiens [1097]. sarcomere [1829]. SARS [2420, 2447]. SARS-Coronavirus [2447, 2420]. Satellite [3606]. Satisficing [3016]. saturability [1127]. saturable [266]. Saturated [2870, 3582]. Saturation [2470, 3484, 3368, 487, 617]. Save [3343]. Scaffold [3009, 2204, 2781, 2758]. Scaffolds [2869]. scalar [785]. Scale [3582, 2595, 2520, 2717, 3255, 3441, 1610, 1587]. Scales [2898, 2599, 3375, 3431]. Scaling [3307, 2936, 1225, 3022, 976, 1329]. scallop [1228]. scapularis [3111]. Scenario [3161]. Scenarios [1602]. Scent [2847, 1733]. Schaefer [1118]. Schedule [3329, 2795]. Scheduling [3390, 1795, 1309, 1948]. Scheme [3118]. Schemes [2136]. Schistosoma [2308]. Schistosome [1999, 1980]. Schistosomiasis [3413, 2926]. schizophrenia [65, 257]. School [841, 1208, 1255]. Science [125, 941, 2048, 3443, 2039, 3014, 1381]. Sciences [3034, 2143, 3444, 573, 485, 672, 668, 1264, 504, 125]. Scientific [3441, 34]. Scientists [77]. SCoP [875]. scores [1163]. Scott [1118]. Scratch [3238, 3349]. Screening [3599, 3061]. Scroll [2824]. Sea [2667, 2251, 1609, 1752]. Sea-Age [2667]. seals [1523]. Search [2358, 2613, 2453, 2491, 1469, 1191, 1970]. searches [886]. Searching [751, 2555, 813]. Seasonal [2861, 2685, 2829, 1909, 2939, 1656, 3070, 2690, 2771, 719, 1692]. Seasonality [2412, 2385, 2455, 2622, 2956, 2631, 1943]. Seasonally [1857, 1368, 1221]. seasons [1180]. secant [1127, 1128]. Second [133, 2151, 481, 218, 88]. Second-Growth [2151]. second-order [481, 88]. Secondary [2870, 2025, 2157, 2782, 3379, 808, 1369, 1801, 1969, 1878, 1193, 1454, 1490, 1625, 755, 2332]. Secretase [2515]. Secretion [3328, 1996, 1695, 1362, 1820, 209, 621, 445]. Secretory [2226, 6]. Section [842, 744, 778, 789, 804, 811, 823, 824, 831, 851, 862, 878, 895, 913, 922, 934, 942, 953, 994, 962, 968, 974, 982, 1002, 1014, 1023, 1033, 1045, 1053, 1062, 1076, 1086, 1093, 1102, 1109, 905]. sectional [306, 377, 9]. Seed [2949]. seeder [1837]. seeking [1181]. Segel [2039, 1283]. segment [1005, 1235]. segmental [805, 653]. Segmentation [3162, 1360]. Segregation [2183, 2941]. SEIR [3253, 2465, 2331, 3600]. SEIRS [2126, 2212]. seizures [1465]. Selbstorganization [776]. select [1742]. Selected [2971, 393]. Selecting [2920, 2678, 2146]. Selection [3480, 3474, 3311, 2080, 2197, 2600, 3613, 2177, 2710, 3058, 2769, 564, 429, 3530, 3172, 3240, 2206, 2133, 2987, 2289, 3516, 1544, 259, 644, 308, 940, 1781, 1848, 2068, 1071, 177, 268, 437, 383, 405, 557, 740, 795, 1182, 638, 1862, 797, 147, 1785, 712, 1016, 1584, 1663]. Selective [2383, 2969, 2105, 3542, 1672]. selectively [1700]. selectivity [263, 356]. SELEX [2483, 2938]. Self [2509, 2383, 389, 131, 2514, 2041, 2966, 2163, 1610, 2679, 2777, 2156, 3587, 3612, 2505, 358, 3376, 1414, 1583, 776, 1390, 1389, 471, 1573, 405, 429, 1343, 432, 1240, 1241, 1326, 603, 1287, 1917]. Self-assembly [131]. self-focusing [1287]. Self-generated [2514]. self-growing [1343]. Self-Maintaining [2156]. self-modulation [1573]. Self-organisation [2679, 2505]. Self-Organization [2041, 389, 1610, 1583, 776, 1389, 432, 603]. Self-Organizing [2163, 1414, 471]. self-propagating [1390]. self-regulation [471]. Self-Renewal [3376]. self-replicative [1240, 1241]. Self-repressing [2777]. Self-Reproducing [3612, 405, 429]. Self-Shading [3587]. Self-similar [358]. Self-Structuring [2383]. Self-sustained [2966]. self-thinning [1917]. Self-tolerance [2509]. selfreproducing [383]. semantic [827]. Semi [3333, 674, 3130, 626, 164, 102, 1909]. Semi-differentiated [3333]. Semi-dilute [3130]. semi-discrete [1909]. semi-empirical [164]. Semi-linear [674]. semi-Markov [626]. semi-permeable [102]. Semiarid [3092]. semiclassical [59]. Semiconductor [90, 91]. Semigroups [3464]. semipermeable [101]. Senegal [1258]. senile [1808]. sense [1410]. Sensing [2776, 2843, 3301, 3507, 3156, 3389, 3394, 2513, 2692, 2693, 3221, 3167, 3326, 2945, 3032, 2463, 3232, 1824, 1670, 1029, 1732, 1780]. Sensitive [2528, 3437, 2524, 3249, 3390, 1179, 1201]. Sensitivities [3100, 690]. Sensitivity [3383, 1881, 2283, 179, 3303, 3559, 3272, 3517, 3258, 3289, 2641, 3429, 2297, 2214, 2306, 1743, 1705, 168, 1847, 1322]. Sensory [2336, 2248, 2708, 354, 1705, 280]. Separation [2351, 86]. Septal [3065]. Septation [2369]. SeqA [2560]. Sequence [813, 3312, 3354, 3425, 2713, 963, 2069, 888, 959, 1105, 1794, 1190, 1246, 754, 1202, 1088, 1079, 1227, 1543, 886, 1163, 751, 1188, 1896, 1008, 747, 753, 1312, 1243, 1528].

Sequence-Dependent [3425]. sequence-structure [1543]. Sequences [2069, 1868, 1484, 1007, 732, 1217, 750, 1848, 757, 512, 1125, 1359, 1955, 1756, 1256, 1106, 1460, 1078, 1790, 2026, 554, 642, 773, 758, 748, 1330, 756, 1035, 749]. Sequencing [1997]. Sequential [3475, 2852, 926, 702, 2406, 64, 1785]. Sequestration [2560]. SERCA [2282]. Series [2519, 2794, 2932, 2413, 1143, 1846, 128, 1847, 1889]. Serious [3481]. Serological [2930, 1984]. seropositivity [976]. Serratia [1747]. Service [1032]. Servocontrol [852]. sessile [1423]. Set [3162, 2845, 1814, 1934]. setpoint [1684]. Sets [2209, 2886, 566, 34, 148, 1048]. Settings [3201, 2630, 2724]. seven [807]. Seventeenth [850]. Several [3075, 749]. severe [2031]. Severity [3062]. Sex [3189, 2015, 3283, 2265, 2016, 1176, 3546, 3274, 2187, 1838, 1167]. Sexes [2189]. Sexual [3311, 2395, 3133, 3439, 2146, 1653, 1853]. Sexually [2774, 3593, 838]. Shading [3587]. shallow [1147]. Shannon [152]. Shape [3564, 2409, 3528, 2550, 2374, 3103, 3026, 2703, 3376, 820, 1228, 306, 71, 483, 116, 203, 1443, 395, 971, 1040]. Shaped [2379, 3293, 1399]. Shapes [2581, 711]. Shaping [2897, 2772]. Shapley [3473]. Sharing [3102, 256]. Shark [2896]. Sharp [3386]. Shea [2298]. Shear [2374, 2650, 1645, 1485, 718, 1487]. Shed [3615]. Sheep [3099]. sheet [550]. shell [116, 203, 3583]. Shift [2936, 2614, 197, 1528]. shift-effect [197]. Shifting [2346]. Shifts [2270, 770]. shingles [1600]. Shock [3069, 3423]. Shoot [3564]. shoot-partitioning [1336]. Short [1075, 3091, 3303, 3272, 2417, 2166, 3074, 2359, 1081, 773]. Short- [3091]. Short-Distance [3074]. Short-range [1081]. Short-Sighted [3272]. Short-Term [3303]. Shotgun [1997]. Should [2927]. Shrimp [2426]. shrinking [102]. shuffling [1009]. Shuttling [2758]. SIAM [1075]. sibling [129]. sibling-sister [129]. Sick [2257]. Sickle [2969, 2125]. sickled [925]. sickness [1810]. Side [2061, 1448]. sieving [1297]. Sighted [3272]. Sightings [2356]. Sign [2533]. Signal [3394, 2513, 2417, 3147, 2656, 3373, 2735, 578, 1573, 1780, 159, 1696]. Signaling [2329, 2449, 3265, 2640, 2324, 2933, 3625, 3512, 2860, 3073, 1942, 1808]. Signalling [3350, 2972, 998, 1817, 1939]. Signals [2691, 2623, 1341, 1763, 446]. Signature [2083]. Signatures [3333]. Significance [959, 3407, 1328, 889, 17, 1725, 253, 1202, 645, 893]. Significant [2515]. silent [528]. Silico [3397, 2659, 3575]. similar [358]. similarities [360]. similarity [890, 889, 959, 1011, 813, 752, 705, 1163, 352, 773]. Simple [3021, 3025, 2617, 2061, 2062, 1891, 3392, 3153, 3072, 3030, 3083, 3121, 3200, 2594, 3046, 2761, 2239, 3059, 2069, 2184, 2932, 2547, 1749, 1897, 513, 431, 472, 768, 1429, 634, 986, 638, 1537, 2067, 712, 1624, 1466, 1987, 1518]. Simplex [2357]. Simplification [2324, 1556, 1625]. Simplified [2486, 3346, 433, 1145, 1869]. Simplifying [2606]. Simply [3409]. simulate [1526]. Simulated [2378, 706, 2198, 1211]. Simulating [3617, 2638, 3372, 1776, 2466, 1504, 1554]. Simulation [530, 2779, 3547, 3555, 2586, 163, 2071, 3554, 3489, 3450, 3389, 856, 2750, 3468, 3168, 964, 3557, 3560, 2880, 2391, 275, 612, 3556, 3128, 2775, 1521, 869, 1341, 784, 864, 739, 870, 873, 875, 245, 1660, 190, 168, 1305, 1432, 965, 871, 1358, 1934, 1844, 1843, 74]. Simulation-Based [3389, 3468]. Simulations [3451, 2387, 2127, 2857, 2928, 2598, 3208, 3548, 2195, 3232, 354, 1355, 1139, 172, 1566]. Simulator [3159]. Simultaneous [2081, 3422, 785, 3569, 56, 190, 945]. Sin [2234]. Since [2755]. Single [2250, 2828, 1139, 2716, 3361, 714, 2682, 2184, 2280, 2294, 2293, 649, 927, 1816, 1746, 438, 881, 1341, 330, 102, 919, 960, 1470, 1614, 1739, 1220, 419, 613, 1442, 515, 11, 183, 18, 772, 174]. Single-Crossover [2250]. single-dose [1816]. Single-Occupancy [2184]. Single-Parameter [2828]. Single-salt [714]. single-species [649, 927, 881, 1614, 1220]. Singularity [3239, 2161, 136]. Sink [3522]. Sinking [2112]. sinusoidal [1432]. Siphons [2477]. SIQR [3261]. SIR [3126, 3424, 3090, 2960, 2257, 2249, 2465, 528, 2018, 2331, 3088, 3035, 1546]. SIRC [3253]. SIRI [3231]. SIRS [2018, 2838]. SIS [3090, 3134, 3399, 3432, 2647, 2053, 3038, 3542]. SIS-Like [2647]. sister [129]. sister-in-law [129]. Site [3494, 3172, 2544, 2760, 2527, 2952, 714, 1291, 1466, 2850]. Site-Frequency [2527]. Site-Selection [3172]. Site-Specific [2760]. Sites [2392, 2855, 2527, 1539, 844]. situation [630]. Six [2963, 1230]. Size [3145, 2605, 3362, 2412, 3481, 3610, 2916, 2917, 2197, 2897, 2550, 2361, 2968, 2707, 2394, 2126, 2249, 2134, 2736, 3118, 2021, 3349, 3190, 3108, 2985, 2027, 3376, 1593, 2653, 814, 1715, 402, 428, 1739, 1089, 245, 1051, 1526, 1565, 1184]. Size-at-Age [2134]. Size-based [2550]. Size-Independent [2394]. size-specific [1565]. Size-Structured [2653]. Sizes [3395, 2725, 1654]. Skeletal [2245, 2669, 2087, 1284, 349]. skeleton [1599]. Skellam [1114]. skimming [637]. Skin [3475, 2116, 2101, 853, 1135]. Slc4a9 [3328]. Sleep [2004, 1723]. Sliding [1905, 3056, 2737, 433]. slime [1242]. Sloping [3308]. Slow [3323, 2831, 3286, 2523, 1020, 779, 1697, 522]. Slowly [2464]. slug [1040, 1242]. Small [2334, 2425, 1978, 1132, 1729, 759, 89, 142, 1319, 1618, 1087, 1333, 453, 1647]. Smith [2224, 2397, 1163]. Smoes [229, 239]. Smoking [2563]. Smooth [3124, 2645, 2614, 1488, 1982, 3487, 1425]. smoothing [1013, 1595]. Snail [47]. Snoussi [2804, 2803]. Social [3015, 3610, 3285, 3428, 3141, 2149, 2203, 2849, 1648, 573, 485, 1835, 1838, 551]. Societies [877, 1414, 1531, 1583, 1718]. Society [149, 562, 850, 1527, 303, 148, 1001, 27, 28]. sockeye [1173]. sodium [580, 821, 687, 706]. sodium-potassium [821]. sœur [129]. Soft [3565, 2195, 320, 577, 829, 955, 1335]. Software [744, 778, 789, 804, 811, 823, 824, 831, 842, 851, 862, 878, 895, 913, 922, 934, 942, 953, 994, 962, 968, 974, 982, 1002, 1014, 1023, 1033, 1045, 1053, 1062, 1076, 1086, 1093, 1102, 1109, 905]. Soil [3582, 3309, 2990, 3308, 2850]. Solely [3293]. Solid [3294, 3367, 3317, 2903, 3003, 2220, 1883, 133, 187]. Solitary [618, 1427, 1562, 1485]. solitions [947]. solitonic [839]. soluble [251].

Solute [2180, 2723, 330, 919, 960, 419, 358, 18, 945]. Solutes [2384, 368, 404]. Solution [2837, 110, 2045, 924, 2410, 176, 3569, 646, 810, 644, 759, 768, 946, 565, 726, 105, 358, 335, 1303]. Solutions [2247, 3131, 2908, 155, 3573, 435, 158, 397, 631, 1355, 1562, 231, 1783, 376, 919, 960, 172, 1452, 392, 699, 1146, 954, 372, 692, 907, 1622, 279, 86, 298, 319]. Solvable [2760, 191]. solve [1274]. Solving [2627, 2338]. soma [45]. somatic [1761, 1376]. somatic-dendritic [1761]. Some [20, 55, 420, 465, 3025, 2408, 561, 378, 438, 2257, 2, 201, 210, 437, 383, 432, 266, 409, 307, 2069, 3034, 148, 403, 313, 1120, 627, 1039, 174, 2213, 2171, 478, 616, 37, 522, 8, 675, 236, 1060, 2051, 1853, 238, 304, 317]. Somite [2329]. Sorting [2375, 1040]. sound [82, 139]. Source [3522, 3223, 2965, 901, 900, 581, 590]. sources [322, 408, 556]. South [3352, 2506, 2963, 3241, 2764, 3084]. Soybean [2388]. Space [3131, 3596, 2519, 3404, 2639, 2272, 3199, 3339, 3003, 3502, 2499, 2584, 1983, 3092, 72, 384, 422, 311]. Space- [3502, 2584]. space-and [311]. Space-Limited [3199]. Spaces [3529, 3146, 630, 1834, 2022, 492, 1627, 461]. Span [2361]. Sparing [2215]. Sparse [3401]. Spatial [2942, 2776, 2951, 2292, 3323, 2661, 2897, 2786, 3615, 3624, 1174, 3394, 2153, 2190, 2123, 2855, 2183, 2063, 2717, 1744, 2415, 2305, 2411, 2377, 2284, 3034, 2008, 2662, 3135, 2555, 2599, 3033, 3553, 1655, 3514, 3209, 3430, 3341, 3586, 1140, 2237, 3123, 3539, 2306, 1572, 1842, 136, 1461, 2508, 1846, 1638, 2660, 1835, 1610, 1780, 1745, 700, 1796, 118, 1222, 1407, 221, 583]. spatial-temporal [583]. Spatially [542, 566, 254, 3613, 3085, 2404, 2692, 2693, 2939, 2415, 3550, 3591, 3623, 1951, 1463, 1689, 1146, 1691, 727, 1137, 279, 1507]. Spatio [3178, 2185, 3348, 2081, 2703, 2972, 1694, 2043]. Spatio-Temporal [2972, 3178, 2703, 1694, 2043]. Spatiotemporal [1773, 2848, 3366, 2583, 2622, 2364, 2322, 3376]. Spawners [2667]. Spawning [3347]. speak [1861]. Special [2035, 2909, 3274, 3033, 435, 298]. Specialist [2949, 2044]. Speciation [2273, 2390, 338]. Species [2721, 3321, 2951, 2290, 2140, 3469, 2346, 2232, 2185, 3465, 3531, 3316, 3291, 3466, 2971, 3482, 2682, 3378, 3074, 3135, 3248, 2356, 2935, 2335, 2146, 3063, 2152, 1469, 649, 2099, 927, 2661, 1390, 1389, 1463, 438, 881, 99, 490, 1470, 7, 1614, 80, 109, 1430, 245, 1220, 582, 1240, 1241, 283, 248, 118, 132, 191, 273, 399, 447, 506, 552, 689, 174, 298]. Specific [1849, 619, 3401, 3046, 2680, 3368, 3371, 2466, 2760, 1187, 948, 16, 264, 193, 1565]. Specification [2927]. Specificity [2886, 1205]. spectra [267]. Spectral [2280, 410]. spectroscopy [162]. Spectrum [3379, 2544, 2527]. Speed [2812, 2736, 1860, 3117, 2762, 435, 1667, 1392, 1517]. Speeds [2464, 3158, 3378]. spent [1130]. Sperm [1367, 3415]. Spermatocytes [2954]. Spermatozoa [2495, 115]. sphere [1778, 441]. Spherical [2867, 1091, 724, 1599, 682, 964, 339]. Spheroids [2460, 2145, 724, 1307, 1253, 1677, 1578]. SPICE2 [871]. spike [1766]. spikers [1683]. Spine [2372]. Spines [2596, 2671]. Split [3321, 2269, 3235].

Split-Dose [2269]. Split-Facets [3235]. Spontaneous [2950, 2914, 1696, 2217]. Spore [3133, 743]. spores [1156]. Sporulation [2513, 1843]. Sporulation-Initiation [2513]. Spread [2811, 2318, 2330, 2231, 2283, 3524, 2786, 3133, 2301, 3245, 2153, 3077, 2698, 2997, 1990, 2415, 2085, 2279, 3457, 2982, 3115, 2008, 2502, 3439, 3514, 2984, 3371, 1589, 898, 1523, 1694, 1882, 1838, 362, 34, 976, 1776, 193, 767, 3353]. Spreading [2464, 3414, 2736, 3158, 3179, 2310, 3378, 2625, 2029]. Spring [3111]. Springer [1802, 1818, 1831, 1940, 1899]. Springer-Verlag [1802, 1818, 1831]. Spruce [2260]. Squamous [3401]. Squares [2718, 2359, 2410, 1935]. squid [398, 494]. Squirrel [2151, 1178]. Srinivas [2222]. SSC [2175]. Stabilises [2552]. Stabilité [123]. Stability [3330, 3364, 3260, 3089, 2437, 566, 719, 958, 2138, 2996, 1017, 2486, 2632, 295, 2265, 3563, 3085, 379, 270, 570, 1784, 1918, 2884, 2465, 405, 2230, 2674, 3374, 2018, 3236, 3187, 2658, 740, 3500, 2859, 3505, 3164, 1442, 2291, 727, 2497, 3429, 1688, 2532, 86, 1698, 2846, 3220, 371, 927, 569, 1562, 1334, 1415, 509, 123, 1248, 438, 881, 1527, 348, 490, 1610, 1614, 1979, 702, 272, 1041, 582, 1216, 273, 300, 879, 1413, 552, 605, 689, 880, 606, 887, 88, 1140, 1593, 376]. Stability-complexity [570]. Stabilization [3267, 2597, 1351]. Stabilized [2244]. Stabilizes [2516, 3172]. Stabilizing [2931, 1352]. Stable [3563, 3124, 2889, 2582, 3231, 3117, 3597, 1783, 854, 815, 698, 1322]. Stage [2839, 2663, 2763, 2976, 3074, 3439, 2168, 3396, 2694, 3391, 263, 1391, 1080, 127, 284, 3086, 1798, 1040, 1322, 1910, 3314]. Stage-Structured [2763, 2976, 3074, 3439, 3396, 1391, 1798, 1910]. Staged [2084]. Stages [2923, 3009, 2162, 2331, 2481, 2662, 3119, 3284, 3376]. Standard [2104, 2568, 3426, 1585, 135]. Stands [2151]. Staphylococcus [2692, 2693]. Star [3436, 1756]. starburst [1209]. Starvation [2800, 3116, 1097, 1094, 1684]. State [2702, 3131, 3292, 3184, 2519, 3055, 3499, 3067, 2806, 2491, 2838, 2113, 2499, 3408, 3513, 2293, 2792, 3275, 3337, 206, 240, 365, 158, 1388, 487, 1724, 133, 187, 1274, 72, 1180, 172, 1619, 1452, 484, 717, 1130, 1182, 1735, 479, 726, 183, 359, 727, 848, 995, 423, 1762, 1345, 400, 1601, 461, 373, 422]. state-variable [1180]. States [3262, 3563, 2952, 3124, 2806, 2683, 2545, 3060, 2753, 2487, 1259, 1260, 1823, 1270, 497, 1887, 601, 602, 87]. Static [3424, 3296, 2009]. statics [846]. Stationary [2496, 3079, 3196, 3453, 2908, 1051, 997, 2953, 3063, 1617, 1259, 1260, 724, 768, 400]. Statistical [1154, 1646, 1138, 3528, 2209, 1640, 3494, 51, 2342, 662, 1202, 1570, 1125, 1756, 1163, 1443, 1984, 936]. Statistics [3490, 2205, 1929, 797, 2130]. status [5]. Steady [487, 3262, 2952, 3055, 1452, 3499, 3067, 2113, 2683, 359, 2545, 3408, 2753, 1601, 3275, 3337, 373, 365, 158, 1388, 1724, 187, 1823, 1270, 1274, 72, 1619, 1735, 2491, 479, 726, 183, 727, 848, 995, 1762, 87, 172]. Steady-State [3067, 2113, 3408, 3275, 487, 1452, 359, 1601, 365, 187, 1619, 479, 726, 727, 848]. Stealing [2536]. steep [1147]. Stem [2769, 3374, 3136, 3407, 3333, 2943, 2352, 3208, 2689, 2947, 3376]. stenosed [915]. stenoses [70]. stenosis [820, 1886, 652, 496, 466]. step [1739, 267, 1785]. Stepping [2684]. steps [313]. Sterilizing [2774]. Stewardship [3201]. Stieltjes [761]. stiffness [1525, 196]. stimulated [1591]. stimulation [36]. Stimuli [2696, 1232, 1151]. stimulus [1407]. Stochastic [2915, 3551, 3126, 2811, 3455, 2355, 3383, 381, 411, 3552, 2589, 122, 2779, 3177, 2918, 2999, 3547, 3555, 2066, 1007, 1659, 2076, 2061, 898, 667, 2296, 1947, 2949, 390, 3554, 2909, 2911, 3489, 2386, 2394, 1204, 2240, 2572, 3359, 3559, 3460, 3446, 3417, 2398, 2056, 2594, 666, 2913, 2878, 3109, 1865, 669, 3331, 557, 1979, 3561, 3499, 2063, 2816, 3416, 3168, 2299, 3078, 3557, 2591, 2971, 2900, 3277, 3047, 670, 3605, 2941, 2914, 1058, 2602, 1327, 89, 142, 228, 2512, 3002, 2120, 2377, 2078, 2284, 534, 568, 2943, 94, 117, 154, 2928, 3549, 2718, 147, 2009, 2067, 2050]. Stochastic [2891, 3102, 2181, 3553, 2985, 2111, 2154, 2172, 3456, 2745, 2912, 3498, 3063, 523, 626, 1764, 1867, 1617, 671, 662, 110, 167, 169, 241, 357, 866, 378, 2062, 675, 1891, 1963, 736, 1538, 1879, 1895, 564, 235, 546, 1872, 668, 906, 1960, 1530, 743, 740, 245, 799, 1731, 1393, 650, 106, 217, 408, 556, 416, 549, 883, 307, 742, 1941, 1812, 674, 1978, 1806, 527, 944, 581, 345, 1440, 25, 325, 223, 290, 341, 1566, 661, 1503]. Stochastically [3418]. Stochasticity [1739, 2467, 3502, 3385, 2164, 254, 1229]. stochstic [604]. Stock [2644, 2468, 1910]. Stockpile [3118]. Stocks [2763]. Stoichiometric [3613, 1791, 3620]. Stoichiometry [1664, 1668, 153]. Stokes [565, 3027]. Stomatitis [2398]. Stomatocytes [2471]. stop [1697]. stops [1752]. storage [354, 64]. Straight [2205]. Strain [2202, 3175, 2339, 2692, 2693, 3118, 2720, 836, 2814]. Strains [3345, 1999, 1872, 1980]. Strand [2475]. stranded [1953]. Strategic [2607, 2553]. Strategies [2906, 2077, 2232, 2160, 2960, 3205, 2651, 2404, 2453, 3419, 874, 2889, 2680, 3274, 3231, 2498, 3599, 2747, 3117, 3436, 3612, 3422, 2626, 3180, 2308, 1972, 1177, 1588, 1470, 1458, 1701, 1753, 1769, 1031, 174, 1282]. Strategy [2800, 2940, 3359, 2468, 3228, 2838, 2987, 2461, 207, 1629, 658, 1401, 1546]. Stratified [2930]. Stratum [2180]. streak [1632]. Stream [2706]. Streaming [2574]. strength [1757]. Stress [3256, 836, 3272, 1252, 2822, 1485, 210, 718, 565, 975, 1487]. Stress-induced [1252]. Stress-strain [836]. stressed [1204]. stresses [497, 782]. stretch [1542]. stretched [1130]. striated [57, 335]. Strict [3233]. Strings [2875]. strips [836, 486]. stroma [1703, 1741]. Stromal [3362, 2823]. Strong [2197, 3269, 3502, 1791, 1248]. Stronger [2864]. Strongly [2577, 2675, 3129, 1612]. Structural [451, 2978, 1248, 2357, 3348, 3394, 3309, 2768, 300, 3174, 3405, 551, 937, 1520, 123]. Structural-Temporal [3348]. Structurally [3221]. Structure [3564, 3260, 3399, 2131, 2834, 3615, 2325, 3003, 2275, 2868, 2859, 3483, 2256, 2291, 3540, 2627, 2260, 2694, 3318, 1999, 2214, 783, 1484, 2099, 676, 808, 1321, 570, 1431, 1129, 757, 1870, 906, 1543, 551, 700, 1209, 1911, 331, 1921, 1968, 1031, 375, 1655, 602, 1008, 204, 971, 936, 40, 461, 31, 1980]. Structured [2668, 2944, 2506, 2169, 3267, 3580, 3626, 3509, 3062, 2501, 2973, 3212, 2750, 3038, 3382, 2440, 2223, 2930, 2763, 2187, 2976, 2020, 2218, 3074, 2352, 3302, 3166, 3439, 2168, 3396, 2653, 2427, 3120, 1884, 1617, 1676, 1373, 1391, 1765, 1080, 1450, 699, 977, 737, 1507, 1798, 1464, 1910]. structurelle [123]. Structures [2332, 3528, 2870, 3036, 2267, 2225, 2432, 3379, 3176, 2912, 473, 188, 1570, 759, 85, 1969, 729, 1878, 1403, 285, 96, 1454, 1510, 1966, 288, 801, 921, 587, 1850, 1322, 440, 755, 1006]. Structuring [2383]. struggle [256, 283]. Students [2426]. studied [1054, 1516]. Studies [2140, 3528, 2109, 1647, 2499, 314, 390, 1117, 1196, 870, 1113, 972, 1038, 983]. Study [2315, 2282, 793, 3367, 2330, 2840, 3153, 2216, 3241, 3106, 2084, 2638, 2045, 3352, 2933, 3625, 2880, 2675, 3599, 2256, 2291, 2005, 2281, 2938, 2382, 2418, 2176, 2723, 1883, 770, 1412, 764, 1572, 281, 1247, 600, 1448, 1928, 394, 1876, 1678, 1228, 1853, 512, 1203, 1236, 1350, 1745, 914, 1339, 520, 1443, 353, 1223, 1039, 424, 414, 1266, 938, 3, 349, 1483]. Studying [2456, 1850]. style [1952]. Sub [3515, 2724]. Sub-populations [3515]. Sub-Saharan [2724]. subalignment [889]. subalignments [890]. Subcellular [978]. subcutaneous [853]. Subject [3613, 2985, 772]. subjected [1542, 1487]. Subjects [3397]. Sublethal [2193]. Sublinear [1955, 1641, 1658]. Submerged [3132]. submicron [282]. Submodularity [3472]. Subnetworks [2604, 3584]. subpopulation [999, 1058]. Subpopulations [3153, 3068, 800]. substances [317]. Substitutability [3020]. substituted [1876]. substituted-phenylcarbamoyl [1876]. Substitution [2713, 1466]. Substrate [3484, 2593, 454, 1317, 703]. subsurface [1585]. subsystem [157, 1612]. subsystems [1777]. subtilis [2513, 1843]. Subtype [2096]. Subtypes [2101]. Success [2660, 2414, 1185]. Successful [3362]. succession [221]. successive [1360]. Sudan [3352, 3241]. Sudden [2270]. Sufficient [2006, 1709, 112, 2256]. Sugar [2989]. Suicide [2618, 3172, 3616, 1709]. Sulfadoxine [2751, 2454]. Sulfadoxine-Pyrimethamine [2454]. Summary [261, 585]. Summer [841, 1255, 3357]. Super [2309]. Super-Networks [2309]. supercoil [1438]. supercoiled [1199]. Superficial [2178, 2317]. superimposed [1241]. Superinfecting [2988]. superparasitism [1441]. Superspreaders [2541]. supertrees [1913]. Supply [2729, 2887, 2764, 1857]. support [1514, 2143]. Supports [3328, 1581]. Suppressed [3607]. Suppression [3006, 3340, 1018, 1985]. Suppressor [2986]. Surface [2019, 2045, 2674, 2161, 2132, 2723, 2948, 856, 304, 91]. Surfaces [3229, 1203]. Surpass [2965]. surprising [1349]. surveillance [258, 284]. Survey [842, 3255, 744, 778, 789, 804, 811, 823, 824, 831, 851, 862, 878, 895, 913, 922, 934, 942, 953, 994, 962, 968, 974, 982, 1002, 1014, 1023, 1033, 1045, 1053, 1062, 1076, 1086, 1093, 1102, 1109, 1190, 905, 1754, 1775]. Surveys [417]. Survival [2800, 2591, 1665, 3063, 1571, 721, 1329, 966, 607, 1220, 1097, 1786, 1690, 478, 1025]. Survivin [3151]. Susceptibilities [2054]. Susceptibility [2420, 2447]. Suspended [2682, 2880]. suspension [1347, 1660, 576, 58]. Suspensions [3130, 164]. Sustainability [3016, 3020, 3018, 3013, 3014]. Sustainable [3015, 3021, 3017]. Sustained [2616, 3288, 2966, 266]. swallow [1814]. Swanson [1900]. Swarming [2118, 1747, 1832]. Swarms [2746, 2664, 2029]. swelling [102, 309]. Swimming [2564, 2422, 1253, 2457, 2650, 3117, 3326, 105, 115]. Swine [2529]. Switch [2607, 2779, 2115]. Switch-Like [2115]. Switching [3149, 2657, 2817, 2096, 2233, 2846, 2719, 1262, 1836, 146, 1922]. Sylvan [2439]. symbionts [1765]. Symbiotic [3545]. Symbolic [3381, 2541, 2545, 1789]. Symmetric [3463, 2715, 1018, 1017, 476, 660, 688]. symmetrical [714]. Symmetry [2741, 188, 2974, 2146, 448, 724, 1868, 1889]. symmetry-breaking [1868]. sympathetic [553, 575]. Symposia [329]. Symposium [125]. symptomatic [1906]. synapse [1866]. Synaptic [2596, 2671]. Synchronisation [2155]. Synchronising [2552]. Synchronism [1623, 2010]. Synchronization [2877, 2791, 3068, 1094, 2334, 1095, 2621, 2614, 1692, 1855]. synchronize [1656]. Synchronous [2525, 2526, 926]. Synchrony [2246, 2296, 2673, 1919, 2319, 1961]. Synchrotron [3309]. syncytia [388]. Syncytial [2408, 387]. Syndrome [2928, 2031]. Synergism [2659]. Synergistic [2188]. syntactic [757]. synthase [1952]. Synthesis [3237, 3109, 3059, 3053, 2110, 2159, 1400, 1681, 923, 292]. synthetases [1468, 1889]. Synthetic [2621, 1554]. Syphilis [3338]. System [3145, 3433, 2315, 2036, 2389, 3364, 2395, 3265, 3518, 2191, 2090, 2177, 2508, 3322, 2240, 3212, 2765, 2769, 2659, 3459, 2674, 3581, 2740, 2336, 2921, 3065, 2936, 2858, 3508, 3132, 3502, 3059, 2777, 3505, 2003, 2838, 2156, 2961, 3355, 2287, 3070, 2584, 3525, 3510, 2322, 3344, 2935, 3513, 2370, 3411, 1512, 1723, 433, 1263, 458, 770, 720, 255, 1944, 620, 167, 241, 136, 594, 187, 315, 578, 1373, 1380, 379, 102, 892, 1409, 1678, 854, 476, 201, 75, 1398, 1904, 1343, 658, 1836, 484, 717, 61, 1734, 1480, 1352, 1192, 1430, 120, 277, 418, 707, 708, 106, 549]. system [604, 1164, 1041, 582, 1161, 46, 251, 168, 1950, 154, 515, 225, 1413, 1968, 706, 579, 1526, 1237, 1326, 1464, 298, 290, 341, 316, 461, 979, 1157, 1027]. System-Mediated [2769]. Systematic [2757]. systematically [1484]. systèmes [126]. systemic [578, 1350]. Systems [2915, 3015, 2175, 2292, 1255, 2290, 3552, 2978, 2538, 2757, 2137, 3453, 3547, 3555, 2160, 3008, 3596, 3229, 3613, 3085, 2952, 2394, 1994, 2732, 2122, 2789, 3156, 2669, 3106, 3394, 2884, 3258, 3055, 3493, 2521, 2717, 3307, 2504, 97, 3168, 2731, 2253, 2411, 3271, 2683, 2847, 3019, 3558, 3612, 2184, 3296, 3255, 2603, 3408, 2925, 2059, 2815, 3412, 3586, 2046, 2146, 2270, 2780, 3220, 3123, 523, 626, 206, 240, 371, 641, 421, 389, 814, 50, 20, 451, 903, 1572, 420, 465, 1247, 618, 488, 8, 671, 662, 110, 169, 448, 510, 133, 1942, 639, 1020, 1668]. systems [899, 673, 1383, 1463, 229, 704, 882, 988, 1261, 1789, 1286, 99, 502, 544, 655, 570, 1060, 1382, 1249, 7, 76, 98, 126, 177, 383, 405, 429, 480, 909, 902, 392, 947, 1422, 668, 624, 170, 175, 224, 277, 276, 454, 792, 699, 386, 1664, 1793, 825, 695, 551, 615, 516, 883, 89, 228, 1644, 96, 317, 479, 171, 1673, 1796, 1547, 117, 234, 726, 119, 527, 218, 331, 1, 155, 183, 273, 305, 372, 503, 152, 366, 1754, 1775, 605, 606, 815, 1520, 21, 25, 325, 453, 460, 559, 560, 611, 971, 113, 965]. systems [1081, 88, 1850, 1938, 614, 1544, 616, 186, 1025, 319, 1847, 1840]. systes [142]. systole [956].

T [3547, 1831, 2846, 2509, 1685, 1742, 3303, 2785, 2642, 1926, 3054, 2442, 2224, 1205, 2827, 1472, 3380, 2810, 3544, 1738, 1310, 2532, 1592, 1698, 3341, 1893]. T-Cell [2785, 1310, 1592]. T-lymphocyte [1205]. T-lymphocytes [1738]. T. [2812]. T1D [3127]. T4 [2466]. T8 [2466]. tags [1611]. Tail [2199, 3404]. tailed [918, 1481]. taking [1863]. Talk [3360]. Talking [2656]. Tamiasciurus [2151]. tangle [1499]. tank [1037]. tank-treading [1037]. tant [33]. tapered [52, 78]. tapering [1131]. Target [3267, 3572, 3150, 2483, 1139, 1377]. Target-Mediated [3150]. Target-Oriented [3267]. Targeted [2840, 3360]. Targeting [2173]. targets [1139]. Task [2817, 3141, 1223, 1718, 1266]. Tau [3552]. Tau-Leaping [3552]. Tax [2635]. Taxa [2558, 3468, 2359]. taxation [1522]. Taxis [2255]. taxon [2343]. taxonomic [252, 819]. Taylor [2384]. TCR [1757]. Tear [2557, 2055, 3310, 3452, 2723, 2832]. technique [1011, 1363, 853, 908]. Techniques [2272, 3033, 357, 307, 1854]. Technology [941, 1900]. teleological [643, 402, 428]. tell [355]. Temperate [2078]. Temperature [3122, 2223, 2623, 3243, 3396, 463, 1849, 1334, 4, 321, 525, 1167, 1656, 1135, 979, 349]. Temperature- [3122]. temperature-dependant [1849]. Temperature-Dependent [3243, 3396, 1334, 979]. temperatures [1692]. Template [2904]. Templates [1544]. Temporal [2776, 2185, 3359, 1060, 3348, 2845, 2374, 2405, 119, 2081, 2599, 2656, 2972, 3178, 1694, 2043, 1615, 1223, 2703, 1266, 221, 583]. temporally [1509]. Temporary [2454, 2738]. Tendency [3614]. Tendon [2820]. Tension [2824, 320, 198]. Tenth [681]. Term [3091, 3303, 3351, 2233, 2584, 731]. Terminal [2139, 2610, 1235, 1550, 1976, 1933, 738, 636, 805, 653]. terminals [1434]. Termites [2232]. Terms [3534, 2164, 2660]. Ternary [3381, 414]. Territorial [1178]. Territory [3103]. Tertiary [3483, 808]. Tessellation [2485]. Tessellations [2729]. Test [2359, 1048, 1828, 1174]. Tested [2059]. Testing [2580, 1510, 1202]. Testosterone [2056]. tetanic [706, 835]. Tether [1995]. Tetrahymena [1397, 1397, 1398]. tetrangle [1013, 1595]. Texas [125]. TGF [2389, 2921]. TGF- [2921]. TGF-Mediated [2389]. Th1 [1685, 1569]. Th1/Th2 [1569]. Th17 [2921]. Th2 [1685, 1569]. thalamic [1806]. thalamocortical [1134]. thaliana [3564, 3460]. Their [1019, 2596, 3534, 2209, 1431, 2965, 3165, 2120, 2844, 1706, 1232, 3395, 1789, 919, 960, 1060, 2244, 766, 1608, 797, 677, 1754, 1775, 819, 1927, 2059, 2235, 1048, 475, 755]. Theorem [2425, 3048, 1110, 928, 255, 1565]. theorems [116, 203, 171, 478]. Theoretic [3419, 2272, 2149, 3295, 3470, 2034, 715, 1604, 791]. Theoretical [861, 961, 2315, 2563, 2017, 1380, 2967, 2045, 1228, 2425, 95, 3530, 3155, 2620, 2405, 645, 442, 353, 657, 385, 2125, 2549, 938, 2533, 983, 1594, 444, 548, 1588, 394, 1538, 1570, 829, 1590, 3628, 660, 1370, 937, 1925, 1115, 606, 349, 1668, 603].

Théorie [104]. Theories [3615, 1019, 320, 1182]. Theory [2743, 122, 3367, 444, 165, 2574, 22, 2147, 2967, 2122, 3239, 2613, 443, 2594, 3324, 2997, 2037, 2576, 483, 3147, 2348, 153, 2407, 1266, 1582, 1643, 523, 626, 206, 240, 987, 646, 262, 1814, 104, 1908, 110, 219, 763, 62, 65, 573, 1389, 342, 417, 485, 809, 143, 216, 367, 35, 1280, 2, 1286, 1119, 1071, 683, 788, 948, 51, 1174, 199, 1662, 177, 1446, 1111, 1112, 1113, 668, 103, 1543, 477, 175, 309, 707, 708, 551, 670, 494, 1179, 196, 1068, 159, 12, 1832, 94, 117, 154, 218, 331, 147, 1064, 1181, 185, 289, 340, 352, 393, 532]. theory [292, 59, 1770, 25, 325, 984, 36, 90, 446, 290, 341, 91, 382, 361]. Therapeutic [2017, 2759]. Therapies [2169, 3579, 3448, 3360, 3574, 3003, 3204]. Therapy [2665, 2337, 2492, 3297, 3398, 2170, 2215, 2354, 2894, 2787, 2173, 3390, 1472, 1788]. There [2840, 109, 1808]. Thermal [2105, 208, 21, 296, 351, 986, 218, 1408, 1419, 1809]. thermocouple [230]. Thermodynamic [695, 2525, 2526, 1520]. Thermodynamical [1894, 1949, 1242]. Thermodynamics [2789, 218, 331, 340]. Thickness [2049, 1732]. thin [261, 90, 91]. thin-walled [261]. Thinning [3452, 1917]. Third [1032, 841]. Thompson [116, 203]. thorium [15]. thorium-232 [15]. thoughts [2, 616]. Three [2824, 1995, 2757, 2514, 2568, 3342, 166, 2664, 3376, 3063, 3232, 1555, 500, 1915, 1488, 1206, 1979, 127, 387, 579, 880, 1107, 1303, 1859, 930, 1006]. three- [930]. three-compartiment [1107]. Three-Dimensional [2824, 2514, 2568, 3342, 3376, 3232, 500, 1488, 1206, 387, 1303]. three-layer [579]. three-player [1979]. three-way [1859]. Three-Zone [2664]. Threshold [2191, 2090, 1564, 2353, 2363, 2958, 2440, 3546, 6, 2859, 3396, 1110, 1372, 1845, 136, 106, 1025]. Thresholds [3262, 2068, 1531, 1220]. Thrombi [2821]. Thrombin [2091, 3397]. Thyroid [2929, 157]. Thyrotropin [2929]. Thyroxine [2929, 120]. Tick [2106, 3277, 3111, 1943]. Tick-Borne [2106, 3277]. ticks [1909]. tide [1253]. Tiers [3404]. tiger [1451]. tight [1311]. Tik [2764]. Time [2234, 2751, 3551, 2290, 2355, 2755, 2409, 2197, 3432, 2189, 3292, 2108, 1994, 2519, 2126, 2613, 490, 999, 2817, 2855, 2610, 2696, 2194, 1294, 3007, 2504, 2892, 3336, 3560, 2276, 3502, 3315, 2333, 2597, 2457, 2020, 2584, 399, 912, 2986, 1622, 2505, 3600, 2932, 2549, 2804, 2803, 2975, 2413, 3602, 1136, 1867, 927, 1257, 397, 110, 167, 169, 241, 802, 334, 489, 1194, 1143, 1373, 1329, 425, 376, 430, 1766, 1846, 438, 1639, 704, 881, 1409, 1244, 513, 691, 722, 437, 392, 1480, 1437, 2013, 1130, 567, 799, 408, 680, 455, 171, 1651, 1618, 689, 1440, 840]. time [325, 36, 1459, 1481, 1854, 311, 1818]. time-constants [1194]. Time-continuous [1294]. time-delayed [881]. Time-Dependent [3551, 2696, 2194, 2597, 999, 1622, 110, 167, 169, 241, 425, 691, 722, 325]. Time-Dependent-Asymmetric-Linear-Parsimonious [3292]. Time-Discrete [2584]. time-equivalent [311]. Time-Lagged [2234]. time-periodic [376, 430]. time-resolved [1480]. Time-Series [2932, 1846]. Time-to-Peak [2549]. Time-Varying [2276, 455, 36]. Times [3490, 2577, 2742, 2705, 3395, 2273, 2401, 2252, 1571, 1210, 1789, 607, 1028, 797, 1507].

Timescale [3523]. Timing [2560, 2562, 2218]. Tip [2867]. tips [1804]. Tissue [2485, 3451, 2895, 3448, 3554, 3575, 2075, 3512, 2943, 2872, 2781, 2195, 1726, 487, 919, 960, 111, 829, 920, 380, 943, 103, 321, 768, 1314, 1155, 1335, 972, 1934]. Tissue-Level [3451, 2895]. Tissues [2375, 3152, 3185, 2105, 1356, 1213, 322, 320, 577, 462, 1525, 477, 853, 387, 208]. together [1519]. Toggle [2779]. tolbutamide [923]. Tolerance [555, 2509, 1579]. Tomograms [2855]. Tomography [3309]. Too [2575]. Tool [2834, 810, 1850]. tools [760]. Topical [2196]. Topographic [470]. Topological [1353, 2636, 2082, 3527, 858, 801, 1778, 512, 730, 805, 1201, 921, 1042, 366]. Topologically [1748]. Topology [3436, 1198]. Toric [2683]. Torres [1877]. torso [830, 38]. torus [1778, 985]. Total [3596, 2126, 2321, 2113, 1849, 202, 118, 132]. tour [1153]. toxicant [1204]. toxicant-stressed [1204]. Toxicity [2830, 2990, 3164, 1448, 1480, 867]. Toxin [2092]. Toxin-Producing [2092]. toxins [1444]. Toxoplasmosis [590]. Tracer [2499, 2132, 255, 23, 495, 506, 460]. tracers [1286]. Tracheobronchial [2111, 500]. Tracking [2716, 2413, 1746, 179, 2381]. tract [783, 278, 79]. Traction [2107, 2158]. Trade [2587, 2556, 2772, 704, 1871]. Trade-Off [2587, 2772, 704, 1871]. Trade-Offs [2556]. tradeoff [1904]. Tradeoffs [2757]. Tradescantia [3245, 3587]. Traditional [3150]. Tragedy [2788]. Trails [2847]. Training [2286]. trains [502]. Traits [3283]. Trajectories [2963, 3037, 1660]. Trans [2799]. Trans-differentiation [2799]. Transbilayer [2049]. transcapillary [495, 358]. transcriptase [1890]. Transcription [3330, 3051, 3109, 2656, 2067]. Transcriptional [2779, 3114, 1620]. Transcriptomic [3080]. transdetermination [382]. transducers [471]. Transduction [2417, 3147, 2656, 3373, 2735, 1763, 1780]. Transfection [2001]. Transfer [3294, 3153, 3279, 3143, 2826, 2488, 136, 278, 667, 102, 103, 321, 687, 101, 589, 1104, 79]. Transfers [276]. transform [924, 853]. Transformation [666, 2499, 126, 1428]. transformations [473, 352]. Transformed [3504, 598]. Transient [2292, 330, 686, 3114, 18, 855, 3586, 571, 630, 810, 296, 351, 197, 394, 1619, 1981, 1103, 853, 479, 323, 1338]. Transients [2584, 2270]. transit [513, 1437, 171, 1481]. Transition [2103, 3590, 3414, 2806, 2561, 336, 110, 425, 1199, 441, 1467]. Transitions [3327, 2036, 3385, 68, 640, 436, 1619, 970, 289, 1507, 1868]. transitive [152]. Translate [2234]. Translated [2922]. Translation [3330, 3051, 3408, 1424]. Translations [3511]. transmembrane [1652]. Transmissibility [3118]. Transmission [3622, 2380, 2263, 2277, 2563, 3201, 2458, 2196, 3025, 2534, 3399, 2320, 2357, 3280, 2439, 3606, 2229, 3486, 2798, 2997, 2958, 2018, 2567, 2724, 3227, 2874, 3605, 2622, 2187, 3338, 3413, 2416, 2218, 2274, 3211, 3439, 2926, 2529, 3600, 2322, 3445, 2418, 2480, 3243, 3112, 3421, 62, 1258, 1909, 849, 1038, 553, 575, 1511]. Transmitted [2774, 3593, 2960, 2363, 2924, 2125, 1765, 1513]. transmural [370].

Transport [2819, 2328, 3170, 2722, 93, 3287, 3042, 2821, 2107, 2007, 2345, 2494, 2180, 2989, 2901, 907, 2478, 2132, 2158, 3400, 2703, 2910, 3476, 3275, 3337, 1726, 357, 62, 212, 213, 548, 370, 1784, 947, 484, 717, 768, 1301, 14, 686, 368, 404, 419, 1554, 479, 271, 495, 706, 468, 1440, 1300, 971, 945, 1303, 90, 91]. transport-load-dependent [1300]. Transportation [2980]. Transporters [2135]. Transvascular [2677]. trapped [62]. trapping [947]. traps [931]. Travel [2153]. Travelers [3113]. Traveling [1800, 3210, 472, 3219, 713, 2255, 1990, 2736, 3386, 2341, 3044, 2975, 231, 1517]. Travelling [3488, 2116, 3392, 2893, 2192, 2857, 1945, 2579, 1562, 1667, 1535]. traverse [346]. treading [1037]. Treadmilling [2596]. Treated [2905, 2258, 2314, 2714, 421, 458, 1788, 835]. Treating [3041]. Treatment [2241, 2469, 3247, 2277, 2307, 2890, 136, 3303, 3538, 3329, 3401, 3384, 3457, 3338, 2553, 2416, 2531, 2186, 2985, 2626, 3119, 2850, 1884, 1816, 866, 1321, 1891, 390, 1871, 1981, 1401, 1474, 1925, 865, 1985, 917, 2119]. Treatments [2530, 2678, 1448, 1767]. Tree [2668, 3004, 3321, 3144, 3469, 3541, 2539, 2983, 3335, 3173, 3339, 3584, 2359, 3110, 3035, 2111, 1201, 2032, 2844, 1276, 500, 558, 739, 95, 718, 1417, 1756, 288, 801, 921, 440, 3485]. Tree-Average [2844]. Tree-Based [3335, 3144, 3485]. Tree-Like [2032, 1417]. Trees [3144, 2332, 3005, 3233, 3031, 2569, 3465, 2025, 2157, 3381, 3354, 2783, 2275, 3409, 2676, 3222, 3473, 2606, 2487, 420, 465, 561, 1049, 201, 526, 738, 636, 1349, 771, 805, 858, 1201, 1096, 979, 1042, 1235]. tremor [1556]. Trend [3322]. Trends [3598, 3255]. triads [1348]. Tribolium [1051]. Trichloroethylene [2214, 1726]. trichome [1502]. Triggered [3389]. triggering [1757]. Trimer [2735]. triosephosphate [1735]. tripartite [51]. Triphasic [3179]. Trisphosphate [2081, 1573]. trivalent [457, 493]. tRNA [3237, 1988, 1468, 1889]. Trophic [2859, 3166, 1679, 1116]. trophic-dynamic [1116]. trophoblast [1841]. Tropical [3471]. Tropism [2581, 2466]. true [1575]. Truncation [3131]. Trypanosoma [2905]. Tryptophan [2005, 2281, 2891]. Tsetse [2905]. TSH [2929]. TSP [2444]. tube [1412, 467, 108, 1432]. Tuberculosis [2373, 2277, 2402, 2563, 2940, 2455, 2981]. tubes [164, 1427, 306, 377, 684, 261]. tubular [783, 1966]. Tubularia [656]. Tubule [3298, 2381, 990, 989, 102, 1301, 108, 1300, 845]. tubules [53, 1304, 1305, 520]. Tubuloglomerular [2349, 2417, 1298]. Tumble [2374, 3129, 1645]. Tumor [2460, 2060, 2632, 3329, 2642, 3340, 2769, 3046, 3360, 3575, 2459, 2823, 2170, 3293, 2991, 1307, 1404, 2986, 2957, 2985, 3087, 2351, 3266, 2695, 3595, 2947, 3254, 2700, 3376, 2775, 1534, 993, 407, 1703, 1741, 250, 1141, 1449, 1677, 545, 1618, 1700, 1859, 1934]. tumor-induced [1534]. Tumor-Promoting [3254]. Tumorigenesis [3407]. Tumors [3294, 2328, 3317, 2903, 3003, 2378, 866, 999, 1292, 1058, 1348, 1433]. Tumour [2169, 2269, 3234, 3369, 3215, 2220, 3371, 2145, 2173, 1883, 1975, 1958, 1737, 1816, 1805, 1355, 1672, 3615, 2659, 1753, 1914, 1584, 1578]. tumour-induced [1753]. Tumourigenesis [3537]. Tumours [3367, 2478, 1135]. Tunable [3424]. Túngara [3311]. Tuning [2445].

Tunneling [2826]. Tuple [3494]. Turbidity [3122, 120]. Turbidity-Dependent [3122]. Turbulence [2824]. Turing [1572, 3453, 2999, 3161, 367, 1994, 1782, 2674, 3478, 2615, 2003, 1068, 2584, 2797, 3092, 2211, 3495, 3220, 3623]. turnover [234]. Tutorial [1010, 3528, 3234]. Twice [2983]. Twist [2139, 816, 1268, 1933, 1976]. Twist-Induced [2139]. Two [3496, 2282, 2485, 1718, 2232, 3283, 2262, 2514, 2189, 263, 2265, 3205, 2029, 3030, 2669, 99, 2205, 2309, 870, 3289, 230, 3530, 3291, 3546, 2892, 3500, 2457, 3313, 3082, 3094, 1468, 265, 3146, 3409, 3102, 3510, 2689, 2211, 1104, 2935, 3513, 3095, 2335, 1048, 3282, 3353, 523, 206, 240, 720, 1572, 810, 508, 1374, 487, 631, 167, 241, 1139, 334, 809, 784, 490, 919, 960, 1678, 456, 267, 1761, 1130, 1453, 838, 604, 1811, 709, 1106, 582, 807, 1216, 1901, 1376, 117, 154, 284, 256, 132, 1785, 1330, 552, 689, 1557, 424, 698, 938, 361, 1503, 430]. two-backbone [1453]. Two-category [1718]. two-compartment [523, 487, 167, 241, 604]. two-component [361]. two-current [1811]. Two-Dimensional [2514, 2669, 3291, 3094, 3095, 3496, 2485, 1572, 809, 132, 938]. Two-Fluid [2262]. two-membrane [709]. Two-Patch [2892, 3500, 3513, 3353]. Two-Phase [3082, 1503]. two-prey [698]. Two-regime [230]. Two-Runners [3205]. Two-Sex [3283, 2265, 3546]. Two-Species [3291, 99, 490, 689]. Two-stage [263, 284]. two-state [206, 240]. two-step [267, 1785]. Two-Valued [3146]. two-variable [1761]. Type [2269, 2320, 2565, 3472, 2272, 2397, 3480, 508, 1875, 1821, 2068, 392, 1837, 1339, 2357, 2565, 1996]. Type-2 [2357]. Types [2944, 2695, 1374, 334]. typhimurium [2070]. Tyson [1900].

U [2129]. UK [1920, 1877, 3628]. Ultra [2523]. Ultra-Slow [2523]. ultradian [1462]. ultrafiltration [945]. Ultrametric [3354, 2487]. Umbilical [3099]. Unbounded [2184]. Uncertain [2927]. uncertainties [1863]. Uncertainty [2670, 3234, 2357, 3382, 2542, 519, 1965, 1530, 1184]. uncorrelated [901, 900]. Under-Consumers [2788]. underdetermination [1019]. Undergoing [2193, 3247, 3019, 1903, 349]. Underlying [2037, 2164, 998]. Underpinning [2820]. Understand [2617]. Understanding [3015, 3041, 2574, 3054, 1474, 2272, 3026, 3296, 3045, 16, 865]. Undirected [2360]. Unearthing [2722]. unequal [1754, 1775]. unfolding [1803]. Unforced [2239]. Uni [2726]. Uni-directional [2726]. Uniaxial [3533]. Unidirectional [1315, 810, 479]. Unification [3543]. Unified [2348, 547]. Unifies [3240, 1543]. uniformities [1241]. uniformly [1787]. unifying [1662, 723, 2061]. Union [3300]. Unique [2032, 2171, 3225]. uniqueness [554, 642, 758, 758]. unit [489, 402, 428]. Units [2850, 819]. univariate [1653]. Universal [3490, 2142, 3190, 594, 1368, 1779]. University [1159, 1123, 1646, 1920, 2129, 1930, 1647, 1877, 2143, 1847, 125]. unlabeled [738]. unlimited [1089]. Unobserved [2829, 3068]. Unperturbed [2169, 1975, 1958]. unpleasantness [121, 32]. Unreasonable [1030]. Unrooted [3233, 3335, 3485, 3381, 3473, 819]. Unscreened [3599]. Unstable [3433, 2290, 1294]. Unsteady [2586]. unstirred [989]. unstressed [1870]. Unstructured [2501]. Unsustainable [3019]. unusual [1237]. Unvaccinated [2637]. update [1064]. Updates [3597]. Upon [2177, 2568, 245, 1220, 3544]. Upper [1026, 1723, 500, 822, 1128]. Uprooted [3278]. Upside [2744]. Uptake [2599, 1947, 682, 1967, 1021, 572]. Urban [3606]. Urchin [2251, 1609, 1752]. Urea [2135, 1152]. urinary [1302]. Urine [2135, 2065, 2429, 1815, 1807, 1152]. urn [281, 508, 793, 792, 532, 633, 725, 840]. uronamides [1876]. urothelium [1939]. Usage [3142]. Use [2729, 3622, 301, 2174, 2454, 2798, 2905, 1218, 479, 1565, 871, 2195, 1854, 79, 1508, 217, 1345, 808]. Used [2294, 2295, 272]. useful [1545, 1597]. useless [1908]. Users [2015]. Using [3145, 2511, 3357, 2566, 3543, 2831, 3397, 2617, 3162, 3153, 2648, 2967, 3068, 3030, 2109, 3559, 2564, 3046, 3404, 3511, 3601, 2655, 3382, 2321, 2272, 2806, 3442, 2180, 1611, 2598, 2545, 2410, 3600, 1526, 3325, 3284, 3516, 2214, 890, 888, 959, 1257, 15, 1247, 2993, 1147, 1897, 1024, 1203, 1437, 245, 1050, 924, 1351, 1393, 1410, 3024, 1896, 576, 1107, 872, 785, 1913, 1935, 1844, 2597]. UTB [2135]. Uterine [3001, 2645, 1560]. uterus [193]. utility [1019]. utilization [1244, 1428].

V [125, 1819, 1877, 124, 283, 956, 1039]. Vaccinated [2257]. Vaccination [2944, 2160, 2960, 2265, 2126, 2249, 1200, 2973, 3097, 2740, 3088, 2554, 2680, 3231, 3599, 3320, 2416, 3343, 2624, 2795, 3084, 3098, 1600, 1310, 1546, 2994]. Vaccine [3028, 2454, 2973, 2084, 2354, 2814, 2418, 2235, 2700]. Vaccine-Induced [2973]. Vaccines [2017, 2900]. Vacuum [2236]. vagueness [1396]. Validated [2991]. validating [307]. Validation [3234, 3366, 3077, 3519, 2248, 1689]. Validity [2059, 995, 1481]. Valley [2959, 2786, 3174]. Value [3473, 3064]. Valued [3146, 648, 901, 900]. Values [2050, 1510, 1171, 1322]. Valve [3056]. Valveless [2194]. Vancouver [1123]. Variability [1396, 3577, 2134, 3038, 3316, 2285, 515, 281, 253, 1177, 1185, 408, 556, 893]. Variable [3610, 2684, 1814, 1180, 1761, 1182, 1665, 1135, 1223, 507]. Variable-Length [2684]. variables [1388, 865]. Variance [3551, 2749, 3560, 2206, 2133, 806]. Variance-Reduced [3560]. Variances [2120]. variate [1054]. Variation [2380, 2741, 2897, 3257, 2705, 2667, 2855, 2583, 3378, 2532, 3027, 2472, 1915]. Variational [1394, 1525, 365, 450, 576]. Variations [1148, 3359]. varicella [1600]. Various [2108, 3204, 2695]. Varma [1847, 482]. Varroa [3246, 3070]. Varying [2755, 2437, 2126, 2276, 2982, 3512, 569, 721, 437, 166, 455, 415, 9, 1786, 279, 1152, 36]. Vasa [46]. Vascular [2269, 2756, 2691, 2478, 2072, 1916, 1931, 1805, 95, 718, 1753, 1403, 288]. Vascularising [2204]. vascularization [250, 1934]. Vasculature [2611, 2733]. Vasculogenesis [2220]. Vasoreactivity [3012]. Vector [2906, 2141, 3391, 3267, 2534, 2812, 2960, 3212, 3606, 3601, 2924, 3604, 2874, 2086, 2976, 3351, 3603, 3525, 3174, 756]. Vector-Bias [2534]. Vector-Borne [2141, 3391, 3212, 3606, 3604, 3525]. Vector-Control [2906]. Vector-Transmitted [2960, 2924]. Vegetation [3536, 2866, 2990, 3308, 3092]. Vegetative [2674]. VEGF [3398, 2227, 3360]. VEGF-Targeted [3360]. VEGFC [3495]. Veils [2514]. velocities [52, 78, 716]. Velocity [3210, 1562, 1037, 2510, 3060, 1232, 1500]. Veneziani [2253]. venous [1417]. ventricle [210, 455, 196, 344]. Ventricular [3065, 1493, 343, 782, 168, 1039]. Venus [2739]. veratridine [580]. Verhulst [392, 774, 872]. Verlag [1802, 1818, 1831, 1940, 1899]. Versatility [2617, 1502]. Version [2790, 3520, 2512]. Versus [2197, 2574, 3016, 2745, 3602, 2250, 1470, 1813]. Vertebrate [2245, 2244, 2312]. Vertex [1226, 926]. Vertical [2177, 3421]. Vertically [1765]. vertices [636]. vesicle [1713]. Vesicular [2398]. Vespa [998, 486, 1354]. Vessel [2303, 826]. Vessels [2901, 166, 565]. Via [2050, 68, 3490, 3267, 3593, 1817, 2160, 1567, 270, 1897, 3472, 1343, 3463, 2319, 567, 3113, 3347, 2375, 3612, 2466, 1934, 1683, 1649]. viability [600]. viable [1825]. vibration [684]. Vibrational [1821]. vibrations [49]. Victim [2137]. View [2908, 3482, 3012, 643, 1778]. Viewpoint [1034]. views [676]. villous [1301, 1841]. Viral [2469, 2539, 2785, 3489, 2216, 3276, 2479, 3127, 3320, 1724, 1822, 1788, 1985]. viremia [1962, 1981]. Virions [3592, 2988]. Virotherapy [2436, 3052, 3266, 3539]. Virtual [3090, 3159, 2810]. Virulence [2373, 2587]. Virus [2234, 2766, 2790, 1479, 2458, 2665, 2408, 2959, 2385, 2337, 2632, 2357, 2301, 2398, 2581, 3077, 3384, 2353, 3504, 2809, 3313, 2622, 3070, 3246, 3578, 3439, 2570, 2532, 3343, 2787, 2466, 2795, 1972, 1523, 1906, 1874, 1554, 598, 1700, 1859, 3353, 1977, 2013, 1990]. Viruses [3621, 2587, 2235]. Visceral [3241, 3352]. Viscoelastic [2107, 2158]. viscosity [165, 496]. Visual [97, 136, 197, 1129]. Vitamin [2835]. vitreous [134]. Vitro [2366, 2224, 2397, 2319, 2686, 3325, 1512, 742, 1640, 1426, 157]. Vivo [2319, 3508, 1521, 2635, 198, 1307, 1575, 907, 930, 983, 3087, 1342, 910]. Voit [1877, 1840]. Volatile [3295]. voltage [1582]. voltage-dependent [1582]. Volterra [641, 903, 927, 231, 143, 216, 439, 3417, 617, 1517, 3331, 392, 338, 511, 985, 815, 887, 223, 1358, 319]. Volterra-Verhulst [392]. Volume [3298, 2167, 514, 1550, 72, 344, 171, 13, 1189]. Volumes [1802, 8]. volvocine [1987]. Voronoi [2729, 2485]. Vortex [2118]. vs [1439, 1947, 2768, 1785]. VSD [3065]. vulnerable [1170].

W [1646, 1831]. Walk [3566, 1294, 271]. Walkers [2997, 3445]. Walks [2209]. Wall [2452, 2682, 990, 826, 718, 196, 497, 782, 1103, 907]. walled [261]. Walls [3185, 550]. Warburg [3359]. Wards [2567]. Warts [2944]. Warts-Causing [2944]. wasps [1176]. Wastewater [2850]. Water [2384, 3170, 2674, 2990, 2901, 47, 1712, 380, 943, 368, 419, 404]. Waterborne [2480, 3112]. Waterman [1163]. Watson [3480, 1089]. Watts [1647]. Wave [2824, 2341, 2745, 435, 1355, 1562, 231, 502, 586, 716, 1237, 1869]. Wave-Pinning [2745]. Wavelength [3058, 1016, 105]. Wavelengths [1628]. Wavelet [1536]. Wavelet-based [1536]. Waves [3488, 3210, 2116, 3392, 2893, 2500, 3414, 3219, 2192, 2255, 1990, 2736, 2759, 2857, 3386, 3044, 2088, 2579, 1800, 421, 458, 618, 1667, 816, 1427, 1485, 916, 2043, 1918, 1491, 472, 1953, 1834, 1315, 1945]. Way [2865, 2707, 1155, 1859]. Weak [1011, 1301, 1791, 1144]. weaker [1430]. Weakly [3511, 1422, 3220]. Weapon [3295]. web [736]. Weber [643, 187, 402, 428]. Webs [3083, 2099, 439, 492, 3121, 2859]. Weed [3169, 3245, 1421]. Wei [536]. Weibull [966]. weight [1684]. Weighted [2784]. weighting [963]. Werner [2928]. West [2458, 1972, 1977, 2385, 2353, 2013, 1990, 3353]. wetlands [733]. Wheat [2901]. where [1923]. which [357, 594, 1653, 706, 424]. white [918, 1728, 840]. white-tailed [918]. whiting [1172]. Whole [2934, 3228, 3065, 1997]. Whole-Pathway [2934]. Whole-Plant [3228]. Wide [3163]. Wild [3449]. Wildfires [2866]. wildlife [1450]. will [1752]. wind [1945]. Window [3159, 1161, 1905]. Winfree [1818]. Wing [3217, 1945]. Wings [3431]. Winner [3236, 3050, 3155, 2350, 1604]. winter [918]. Wiper [2238]. wiring [891]. wiring-in [891]. Withdrawal [2630]. Within [2469, 2202, 3451, 2665, 2337, 3575, 3221, 2008, 2787, 2720, 3607, 2723, 281, 1094, 1409, 1095, 1500, 1223, 3532, 2360]. Within-Host [2202, 2665, 2337, 2787]. Within-Human [3607]. Without [2997, 946, 684]. Witt [143]. Wnt [3106]. Wolbachia [2831, 2501, 2798, 3546, 2982]. wolves [1733]. wool [1500]. word [1106]. words [750]. Workers [2015]. works [417]. World [1032, 3449, 3159]. Worlds [1647]. Worst [2530]. Worth [2144]. Wound [3178, 2617, 3100, 2192, 2687, 2727, 2338, 1998, 1489, 1732, 1418]. Wounds [2492]. Wright [2643]. Writhe [1933, 2139, 1976]. Wu [1847].

X [2143, 1646, 1802, 2855, 3309, 2489, 1899, 31]. X-Ray [2855, 3309, 31]. X174 [512]. X4 [2607]. Xbra [1942]. Xenobiotic [2489, 1480]. Xenograft [3398]. Xenopus [1942, 1457, 835]. Xylem [2901].

Yakubu [1802]. Yangtze [2926]. Yates [1415]. Yeast [2807, 2699, 3213, 2758, 3128, 3498, 1145]. Yeasts [2348]. Yellow [2622, 3320, 486]. Yield [2499, 1529]. Yih [537]. York [1646, 1818]. Young [910].

Z [1199]. Zebrafish [3226, 3495]. Zeeman [531]. Zero [2496, 3196, 3511]. Zhejiang [3486]. Zhitomirsky [595]. Zika [3189, 3439, 3405, 3343]. Zone [2664]. zoom [1936, 1974]. Zoonoses [3391]. Zoonotic [3112]. Zooplankton [1563]. ZPA [2244]. Zucchini [1831]. Zygotic [838]. zymogen [1387]. zymogens [1249].

References

Rosen:1973:DRS

[1] Robert Rosen. On the dynamical realization of (M, R)-systems. Bulletin of Mathematical Biology, 35(1–2):1–9, February 1973. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02558788.

Fox:1973:STR

[2] Michael A. Fox. Some thoughts on the requirements for a theory of biology. Bulletin of Mathematical Biology, 35(1–2):11–17, February 1973. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02558789.

Valentinuzzi:1973:NCM

[3] M. Valentinuzzi, M.D., M.Sc. and M. E. Valentinuzzi, E.E., Ph.D. Newman’s chamber models in the study of blood circulation through the liver and heart. Bulletin of Mathematical Biology, 35(1–2):19– 29, February 1973. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02558790.

Cope:1973:PPB

[4] Freeman W. Cope. Piezoelectricity and pyroelectricity as a basis for force and temperature detection by nerve receptors. Bulletin of Mathematical Biology, 35(1–2):31–41, February 1973. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02558791.

Deakin:1973:SMF

[5] Michael A. B. Deakin. On the status of mean fitness maximization as a governing principle in evolution. Bulletin of Mathematical Biology, 35(1– 2):43–50, February 1973. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02558792.

Licko:1973:TSM

[6] Vojtech Ličko. Threshold secretory mechanism: a model of derivative element in biological control. Bulletin of Mathematical Biology, 35(1– 2):51–58, February 1973. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02558793.

Hearon:1973:DCS

[7] John Z. Hearon. Distribution of conserved species in diffusionreaction systems. Bulletin of Mathematical Biology, 35(1–2):59–67, February 1973. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02558794.

Bright:1973:VSC

[8] Peter B. Bright. The volumes of some compartment systems with sampling and loss from one compartment. Bulletin of Mathematical Biology, 35(1–2):69–79, February 1973. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02558795.

Patlak:1973:ADE

[9] Clifford S. Patlak. An approximation diffusion equation for a long narrow channel with varying cross-sectional area. Bulletin of Mathematical Biology, 35(1–2):81–86, February 1973. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02558796.

Martinez:1973:ANR

[10] Hugo M. Martinez and Robert M. Baer. The algorithmic nature of a reaction–diffusion development process. Bulletin of Mathematical Biology, 35(1–2):87–94, February 1973. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02558797.

Richardson:1973:BIF

[11] I. W. Richardson. Bi-ionic flows through a single homogeneous membrane. Bulletin of Mathematical Biology, 35(1–2):95–100, February 1973. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02558798.

Offner:1973:EMB

[12] Franklin F. Offner. The excitable membrane-biophysical theory and experiment. Bulletin of Mathematical Biology, 35(1–2):101–107, February 1973. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02558799.

Reiner:1973:MBK

[13] John M. Reiner. Molecular biology and the kinetics of cell growth. II. A note on mass–volume relations. Bulletin of Mathematical Biology, 35(1– 2):109–113, February 1973. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02558800.

Landahl:1973:MTP

[14] H. D. Landahl. On the mechanism of transport of proline in the ciliary body, optic nerve and choroid plexus. Bulletin of Mathematical Biology, 35(1–2):115–127, February 1973. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02558801.

Bernard:1973:EMD

[15] S. R. Bernard. Estimates of microcurie-days residence, bone dose equivalent, and (MPC)_{W} for thorium-232 in man using a mammillary model. Bulletin of Mathematical Biology, 35(1–2):129–147, February 1973. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02558802.

Karreman:1973:TPU

[16] George Karreman. Towards a physical understanding of physiological excitation as a cooperative specific adsorption phenomenon. Bulletin of Mathematical Biology, 35(1–2):149–171, February 1973. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02558803.

Bartholomay:1973:MMI

[17] Anthony F. Bartholomay. Mathematical medicine, an introductory discussion of its meaning and significance in medicine and medical education. Bulletin of Mathematical Biology, 35(1–2):173–182, February 1973. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02558804.

Stephenson:1973:TBS

[18] John L. Stephenson. Transient behavior of the single loop solute cycling countercurrent multiplier. Bulletin of Mathematical Biology, 35(1–2): 183–194, February 1973. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02558805.

vanderVaart:1973:CIC

[19] H. R. van der Vaart. A comparative investigation of certain difference equations and related differential equations: Implications for model-building. Bulletin of Mathematical Biology, 35(1–2):195–211, February 1973. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02558806.

Baianu:1973:SAP

[20] I. Băianu. Some algebraic properties of (M, R)-systems. Bulletin of Mathematical Biology, 35(1–2):213–217, February 1973. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02558807.

Sugerman:1973:TDA

[21] David Sugerman and Hiram E. Hart. Thermal dissociation of antigen– antibody–antigen-like systems. Bulletin of Mathematical Biology, 35(1– 2):219–235, February 1973. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02558808.

Defares:1973:TMA

[22] J. G. Defares and M. E. Wise. Theory of the measurement of arterial compliance in humans. Bulletin of Mathematical Biology, 35(1–2): 237–244, February 1973. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02558809.

Rescigno:1973:CMT

[23] Aldo Rescigno and Lester Michels. Compartment modeling from tracer experiments. Bulletin of Mathematical Biology, 35(1–2):245– 257, February 1973. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02558810.

Easton:1973:NMM

[24] Thomas A. Easton, Ph.D. A note on the mathematics of microsphere division. Bulletin of Mathematical Biology, 35(1–2):259– 262, February 1973. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02558811.

Thakur:1973:STC

[25] A. K. Thakur, A. Rescigno, and D. E. Schafer. On the stochastic theory of compartments: II. Multi-compartment systems. Bulletin of Mathematical Biology, 35(1–2):263–271, February 1973. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02558812.

Anonymous:1973:Aa

[26] Anonymous. Announcement. Bulletin of Mathematical Biology, 35(1– 2):273, February 1973. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02558813.

Anonymous:1973:AGSa

[27] Anonymous. Aims and goals of the Society. Bulletin of Mathematical Biology, 35(1–2):274, February 1973. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02558814.

Anonymous:1973:AGSb

[28] Anonymous. Aims and goals of the Society. Bulletin of Mathematical Biology, 35(1–2):274, February 1973. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02558814.

Palatt:1973:ACE

[29] P. J. Palatt and G. M. Saidel. An analysis of countercurrent exchange with emphasis on renal function. Bulletin of Mathematical Biology, 35 (3):275–286, June 1973. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458336.

Kamiya:1973:MAC

[30] Akira Kamiya, M.D. and Tatsuo Togawa, Ph.D. Mathematical analysis of circulatory mixing process. Bulletin of Mathematical Biology, 35 (3):287–300, June 1973. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458337.

Wu:1973:SDD

[31] Tai Te Wu. The structure of DNA as deduced from fiber X-ray crystallography. Bulletin of Mathematical Biology, 35(3):301–311, June 1973. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02458338.

Smets:1973:NKU

[32] M. E. S. Philippe Smets, M.D. A note on knowledge and unpleasantness. Bulletin of Mathematical Biology, 35(3):313–317, June 1973. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02458339.

Hyver:1973:RCC

[33] C. Hyver. Réactions chimiques couplées considerées en tant que filtres passe-bande. (French) [Coupled chemical reactions considered as bandpass filters]. Bulletin of Mathematical Biology, 35(3):319–337, June 1973. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02458340.

Rashevsky:1973:OSS

[34] N. Rashevsky. Organismic sets. Scientific creativity, its spread and applications. Bulletin of Mathematical Biology, 35(3):339–344, June 1973. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02458341.

Fischer:1973:NFT

[35] Burkhart Fischer. A neuron field theory: Mathematical approaches to the problem of large numbers of interacting nerve cells. Bulletin of Mathematical Biology, 35(3):345–357, June 1973. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02458342.

Wei:1973:QTT

[36] Ling Y. Wei. Quantum theory of time-varying stimulation in nerve. Bulletin of Mathematical Biology, 35(3):359–374, June 1973. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02458343.

Wong:1973:SPC

[37] Alan Y. K. Wong, Ph.D. Some proposals in cardiac muscle mechanics and energetics. Bulletin of Mathematical Biology, 35(3):375–399, June 1973. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02458344.

Plonsey:1973:ETB

[38] Robert Plonsey, Ph.D. The effect of the torso boundary in electrocardiography. Bulletin of Mathematical Biology, 35(3):401–405, June 1973. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02458345.

McGavin:1973:ANA

[39] Stewart McGavin. An attitude to nucleic acid models. Bulletin of Mathematical Biology, 35(3):407–409, June 1973. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02458346.

Wilson:1973:NAS

[40] H. R. Wilson. Nucleic acid structure. Bulletin of Mathematical Biology, 35(3):411–413, June 1973. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458347.

Deakin:1973:GC

[41] Michael A. B. Deakin. On Gompertz curves. Bulletin of Mathematical Biology, 35(3):415, June 1973. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458348.

Sherebrin:1973:RMD

[42] M. H. Sherebrin. Reorientation of membrane dipoles during nerve excitation. Bulletin of Mathematical Biology, 35(3):417–418, June 1973. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02458349.

Anonymous:1973:Ab

[43] Anonymous. Announcement. Bulletin of Mathematical Biology, 35 (3):419, June 1973. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458350.

Anonymous:1973:Ac

[44] Anonymous. Announcement. Bulletin of Mathematical Biology, 35 (3):419, June 1973. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458350.

Rashevsky:1973:DMD

[45] N. Rashevsky. The diffusion and metabolic differences between soma and axon of a neuron. Bulletin of Mathematical Biology, 35(4):421– 429, August 1973. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02575186.

Palatt:1973:CEI

[46] Paul J. Palatt and Gerald M. Saidel. Countercurrent exchange in the inner renal medulla: Vasa recta-descending limb system. Bulletin of Mathematical Biology, 35(4):431–447, August 1973. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02575187.

Coutinho:1973:SPR

[47] A. Bezerra Coutinho and F. A. Bezerra Coutinho. Snail population in running water. Bulletin of Mathematical Biology, 35(4):449– 458, August 1973. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02575188.

Hyver:1973:FPB

[48] C. Hyver. Filtres passe-bande chimiques et oscillations non periodiques. (French) [Chemical bandpass filters and non-periodic oscillations]. Bulletin of Mathematical Biology, 35(4):459–466, August 1973. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02575189.

Delsemme:1973:FVC

[49] Pierre Delsemme and H. Westcott Vayo. Forced vibrations of a circular muscle ring. Bulletin of Mathematical Biology, 35(4):467– 474, August 1973. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02575190.

Baianu:1973:ADS

[50] Ion Bâianu and Dragoş Scripcariu. On adjoint dynamical systems. Bulletin of Mathematical Biology, 35(4):475–486, August 1973. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02575191.

Henderson:1973:STP

[51] G. P. Henderson and V. B. Cervin. Statistical theory of persuasion in tripartite negotiations. Bulletin of Mathematical Biology, 35(4): 487–494, August 1973. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02575192.

Mirsky:1973:PVC

[52] I. Mirsky. Pulse velocities in cylindrical, tapered and curved anisotropic elastic arteries. Bulletin of Mathematical Biology, 35(4):495–511, August 1973. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02575193.

Blake:1973:FTD

[53] John Blake. Flow in tubules due to ciliary activity. Bulletin of Mathematical Biology, 35(4):513–523, August 1973. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02575194.

Nastansky:1973:EMP

[54] Ludwig Nastansky, Stanley M. Selkow, and Neil F. Stewart. The enumeration of minimal phylograms. Bulletin of Mathematical Biology, 35 (4):525–533, August 1973. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02575195.

Bartholomay:1973:SMA

[55] Anthony F. Bartholomay. Some mathematical aspects of the diagnostic process. II: a mathematical model of electrocardiographic diagnosis. Bulletin of Mathematical Biology, 35(4):535–547, August 1973. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02575196.

Ghista:1973:MSD

[56] Dhanjoo N. Ghista, Sung-Cheng Huang, and Meyer Resnikoff. A method for simultaneous determination of (I) the blood flows to the organs and (II) the cardiac output. Bulletin of Mathematical Biology, 35(5– 6):549–563, October 1973. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458361.

Moisescu:1973:IFS

[57] D. G. Moisescu. Interfilament forces in striated muscle. Bulletin of Mathematical Biology, 35(5–6):565–575, October 1973. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02458362.

Tesfagaber:1973:PDD

[58] Ahferom Tesfagaber and Marshall M. Lih. Particle distribution for dilute suspension in flow. Bulletin of Mathematical Biology, 35(5–6): 577–589, October 1973. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458363.

Spiegel:1973:STN

[59] Ronald J. Spiegel and William T. Joines. A semiclassical theory for nerve excitation by a low intensity electromagnetic field. Bulletin of Mathematical Biology, 35(5–6):591–605, October 1973. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02458364.

Hethcote:1973:ABD

[60] Herbert W. Hethcote. Asymptotic behavior in a deterministic epidemic model. Bulletin of Mathematical Biology, 35(5–6):607–614, October 1973. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02458365.

Kharasch:1973:MSM

[61] Jerome A. Kharasch. A morphogenetic system matched to a group. Bulletin of Mathematical Biology, 35(5–6):615–625, October 1973. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02458366.

Cope:1973:EPT

[62] Freeman W. Cope. Electron-phonon (trapped photon) coupling and infrared coaxial transmission line theory of energy transport in mitochondria and nerve. Bulletin of Mathematical Biology, 35(5–6):627– 644, October 1973. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458367.

Cull:1973:MAA

[63] Paul Cull and Andrew Vogt. Mathematical analysis of the asymptotic behavior of the Leslie population matrix model. Bulletin of Mathematical Biology, 35(5–6):645–661, October 1973. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02458368.

Ovenden:1973:DSI

[64] Michael W. Ovenden and Archie E. Roy. Disseminated storage of information with sequential coding. Bulletin of Mathematical Biology, 35(5– 6):663–688, October 1973. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458369.

Cronin:1973:DET

[65] Jane Cronin. The Danziger–Elmergreen theory of periodic catatonic schizophrenia. Bulletin of Mathematical Biology, 35(5–6):689– 707, October 1973. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458370.

Rashevsky:1973:EN

[66] N. Rashevsky. The epileptic neuron: I. Bulletin of Mathematical Biology, 35(5–6):709–714, October 1973. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02458371.

Rashevsky:1974:ENI

[67] N. Rashevsky. The epileptic neuron: II. Bulletin of Mathematical Biology, 36(1):1–16, February 1974. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02461186.

Almeida:1974:EIP

[68] Silverio P. Almeida, James D. Bond, and Thomas C. Ward. Electrically induced phase transitions via the dipole model in excitable membranes. Bulletin of Mathematical Biology, 36(1):17–28, February 1974. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02461187.

Coutinho:1974:DPB

[69] F. A. B. Coutinho and A. B. Coutinho. Dynamics of populations of Biomphalaria glabrata and the Von Forerster equation. Bulletin of Mathematical Biology, 36(1):29–37, February 1974. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02461188.

Morgan:1974:IMA

[70] B. E. Morgan and D. F. Young. An intergral method for the analysis of flow in arterial stenoses. Bulletin of Mathematical Biology, 36(1): 39–53, February 1974. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02461189.

Lenard:1974:NSE

[71] J. G. Lenard. A note on the shape of the erythrocyte. Bulletin of Mathematical Biology, 36(1):55–58, February 1974. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02461190.

Gomez-Povina:1974:CDV

[72] Oscar A. Gómez-Poviña, Carmen Sainz de Calatroni, Susana Orden de Puhl, and Mariano J. Guerrero. Continuous determination of the volume of a space in a non steady state condition. Bulletin of Mathematical Biology, 36(1):59–66, February 1974. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02461191.

Morris:1974:KLM

[73] Hedley C. Morris. The Kerner linearization method. Bulletin of Mathematical Biology, 36(1):67–76, February 1974. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02461192.

vPalmay:1974:SMN

[74] F. v. Palmay, E. J. Davison, and J. Duffin. The simulation of multineurone networks: Modelling of the lateral inhibition of the eye and the generation of respiratory rhythm. Bulletin of Mathematical Biology, 36 (1):77–89, February 1974. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02461193.

Huddleston:1974:DME

[75] John V. Huddleston, Charles G. DeWald, and Hanasoge N. Jagadeesh. A dynamic model of an environmental system with n interacting components and p degrees of freedom. Bulletin of Mathematical Biology, 36 (1):91–96, February 1974. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02461194.

Hearon:1974:NOL

[76] John Z. Hearon. A note on open linear systems. Bulletin of Mathematical Biology, 36(1):97–99, February 1974. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02461195.

Richardson:1974:IML

[77] I. W. Richardson. Introduction to mathematics for life scientists. Bulletin of Mathematical Biology, 36(1):101–102, February 1974. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02461196.

Mirsky:1974:PVC

[78] I. Mirsky. Pulse velocities in cylindrical, tapered and curved anisotropic elastic arteries. Bulletin of Mathematical Biology, 36(1):103, February 1974. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
BF02461197;   http://link.springer.com/content/pdf/10.1007/
BF02461197.pdf.

Yeh:1974:UHT

[79] Hsu-Chi Yeh. Use of a heat transfer analogy for a mathematical model of respiratory tract deposition. Bulletin of Mathematical Biology, 36 (2):105–116, April 1974. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458596.

Karreman:1974:ISA

[80] George Karreman and Lee Levitt. Isotherms for r species adsorbed at linear polymers with nearest neighbor cooperative interaction. Bulletin of Mathematical Biology, 36(2):117–126, April 1974. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02458597.

Mittra:1974:IPM

[81] T. K. Mittra and S. N. Prasad. Interaction of peristaltic motion with Poiseuille flow. Bulletin of Mathematical Biology, 36(2): 127–141, April 1974. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458598.

Fredberg:1974:PSG

[82] J. J. Fredberg. Pseudo-sound generation at atherosclerotic constrictions in arteries. Bulletin of Mathematical Biology, 36(2):143–155, April 1974. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02458599.

Bunow:1974:EKC

[83] Barry Bunow. Enzyme kinetics in cells. Bulletin of Mathematical Biology, 36(2):157–169, April 1974. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458600.

Kauffman:1974:MMO

[84] S. Kauffman. Measuring a mitotic oscillator: The ARC discontinuity. Bulletin of Mathematical Biology, 36(2):171–182, April 1974. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02458601.

Martinez:1974:GND

[85] Hugo M. Martinez and Gunnar Carlsson. Genetic nets and dissipative structures: an algebraic approach. Bulletin of Mathematical Biology, 36 (2):183–196, April 1974. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458602.

Thames:1974:SES

[86] Howard D. Thames, Jr. Stability and enzyme separation: Integral representation of the solutions. Bulletin of Mathematical Biology, 36(2): 197–203, April 1974. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458603.

Turner:1974:FFM

[87] Jack S. Turner. Finite fluctuations and multiple steady states far from equilibrium. Bulletin of Mathematical Biology, 36(2):205–213, April 1974. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458604.

Walter:1974:GAS

[88] Charles Walter. The global asymptotic stability of prey–predator systems with second-order dissipation. Bulletin of Mathematical Biology, 36(2):215–217, April 1974. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458605.

Mullooly:1974:SKM

[89] John P. Mullooly. Stochastic kinetic models for small irreversible systems. Bulletin of Mathematical Biology, 36(3):219–227, June 1974. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02461325.

Wei:1974:STI

[90] Ling Y. Wei and B. Y. Woo. Semiconductor theory of ion transport in thin lipid membranes: I. Potential and field distributions. Bulletin of Mathematical Biology, 36(3):229–246, June 1974. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02461326.

Woo:1974:STI

[91] B. Y. Woo and Ling Y. Wei. Semiconductor theory of ion transport in thin lipid membranes: II. Surface recombination. Bulletin of Mathematical Biology, 36(3):247–264, June 1974. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02461327.

Levy:1974:FEM

[92] Elinor Miller Levy. Flagellar elongation as a moving boundary problem. Bulletin of Mathematical Biology, 36(3):265–273, June 1974. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02461328.

Davidson:1974:TAM

[93] M. R. Davidson and J. M. Fitz-Gerald. Transport of O_{2} along a model pathway through the respiratory region of the lung. Bulletin of Mathematical Biology, 36(3):275–303, June 1974. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02461329.

Purdue:1974:STCa

[94] Peter Purdue. Stochastic theory of compartments. Bulletin of Mathematical Biology, 36(3):305–309, June 1974. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02461330.

Kamiya:1974:TRB

[95] Akira Kamiya, Tatsuo Togawa, and Atsuko Yamamoto. Theoretical relationship between the optimal models of the vascular tree. Bulletin of Mathematical Biology, 36(3):311–323, June 1974. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/article/10.1007/BF02461331.

Palmer:1974:NPS

[96] Edgar M. Palmer and Robert W. Robinson. On the number of phenotype structures in genetic systems. Bulletin of Mathematical Biology, 36(3):325–338, June 1974. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02461332.

Leibovic:1974:IPV

[97] K. N. Leibovic. Information processing in the visual systems of arthropods. Bulletin of Mathematical Biology, 36(3):339–340, June 1974. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02461333.

Huddleston:1974:OCE

[98] John V. Huddleston. Optimality in the control of environmental systems. Bulletin of Mathematical Biology, 36(4):341–345, August 1974. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02464612.

Gomatam:1974:NMIa

[99] Jagannathan Gomatam. A new model for interacting populations — I: Two-species systems. Bulletin of Mathematical Biology, 36(4): 347–353, August 1974. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02464613.

Gomatam:1974:NMIb

[100] Jagannathan Gomatam. A new model for interacting populations — II: Principle of competitive exclusion. Bulletin of Mathematical Biology, 36 (4):355–364, August 1974. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02464614.

Rubinstein:1974:PTL

[101] L. I. Rubinstein. Passive transfer of low-molecular nonelectrolytes across deformable semipermeable membranes — I: Equations of convectivediffusion transfer of nonelectrolytes across deformable membranes of large curvature. Bulletin of Mathematical Biology, 36(4):365–377, August 1974. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02464615.

Geiman:1974:PTL

[102] H. M. Geiman and L. I. Rubinstein. Passive transfer of low-molecular nonelectrolytes across deformable semi-permeable membranes — II: Dynamics of a single muscle fiber swelling and shrinking and related changes of the T -system tubule form. Bulletin of Mathematical Biology, 36(4): 379–401, August 1974. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02464616.

Klinger:1974:HTP

[103] H. G. Klinger. Heat transfer in perfused biological tissue — I: General theory. Bulletin of Mathematical Biology, 36(4):403–415, August 1974. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02464617.

Biemont:1974:TCI

[104] C. Biémont, A. R. Bouffette, and J. Bouffette. Théorie chromosimique de l’inbreeding: Modèle probabiliste. (French) [Chromosimal theory of inbreeding: Probabilistic model]. Bulletin of Mathematical Biology, 36 (4):417–434, August 1974. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02464618.

Shack:1974:LWS

[105] W. J. Shack and T. J. Lardner. A long wavelength solution for a microorganism swimming in a channel. Bulletin of Mathematical Biology, 36 (4):435–444, August 1974. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02464619.

Matis:1974:PRT

[106] J. H. Matis, M. Cardenas, and R. L. Kodell. On the probability of reaching a threshold in a stochastic mammillary system. Bulletin of Mathematical Biology, 36(4):445–454, August 1974. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02464620.

Blumenson:1974:MMC

[107] Leslie E. Blumenson. Mathematical models of conception and birth. Bulletin of Mathematical Biology, 36(4):455–456, August 1974. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02464621.

Marshall:1974:FNF

[108] E. A. Marshall and E. A. Trowbridge. Flow of a Newtonian fluid through a permeable tube: The application to the proximal renal tubule. Bulletin of Mathematical Biology, 36(5–6):457–476, October 1974. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02463260.

Kerner:1974:WTM

[109] Edward H. Kerner. Why are there so many species? Bulletin of Mathematical Biology, 36(5–6):477–488, October 1974. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/article/10.1007/BF02463261.

Cardenas:1974:STC

[110] M. Cardenas and J. H. Matis. On the stochastic theory of compartments: Solution for n-compartment systems with irreversible, time-dependent transition probabilities. Bulletin of Mathematical Biology, 36(5–6): 489–504, October 1974. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02463262.

Hennessy:1974:IDP

[111] T. R. Hennessy. The interaction of diffusion and perfusion in homogeneous tissue. Bulletin of Mathematical Biology, 36(5–6):505– 526, October 1974. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02463263.

Staley:1974:NSC

[112] David H. Staley, William R. Holt, and William N. Cannon, Jr. Necessary and sufficient condition for a model insect population to go to extinction. Bulletin of Mathematical Biology, 36(5–6):527–533, October 1974. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02463264.

Ventriglia:1974:KAN

[113] F. Ventriglia. Kinetic approach to neural systems: I. Bulletin of Mathematical Biology, 36(5–6):535–544, October 1974. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/article/10.1007/BF02463265.

Levine:1974:DRM

[114] Ralph L. Levine, John E. Hunter, and Peter L. Borchelt. Dustbathing as a regulatory mechanism. Bulletin of Mathematical Biology, 36(5–6): 545–553, October 1974. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02463266.

Shack:1974:OHS

[115] W. J. Shack, C. S. Fray, and T. J. Lardner. Observations on the hydrodynamics and swimming motions of mammalian spermatozoa. Bulletin of Mathematical Biology, 36(5–6):555–565, October 1974. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02463267.

Lovtrup:1974:DTT

[116] Søren Løvtrup and Björn von Sydow. D’Arcy Thompson’s theorems and the shape of the molluscan shell. Bulletin of Mathematical Biology, 36(5– 6):567–575, October 1974. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.

1007/BF02463268. See addendum [203].

Purdue:1974:STCb

[117] Peter Purdue. Stochastic theory of compartments: One and two compartment systems. Bulletin of Mathematical Biology, 36(5–6):577– 587, October 1974. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02463269.

Rosen:1974:BTP

[118] Gerald Rosen. Bounds on the total particle number for species governed by reaction–diffusion equations in the infinite spatial domain. Bulletin of Mathematical Biology, 36(5–6):589–593, October 1974. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02463270.

Reiner:1974:MCS

[119] John M. Reiner. Molecular control systems: I — temporal organization in cells revisited. Bulletin of Mathematical Biology, 36(5–6): 595–599, October 1974. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02463271.

Landahl:1974:NOT

[120] H. D. Landahl. Note on oscillations of turbidity in a precipitation reaction system of thyroxine. Bulletin of Mathematical Biology, 36(5–6): 601–604, October 1974. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02463272.

Haberman:1974:KU

[121] David Haberman. On knowledge and unpleasantness. Bulletin of Mathematical Biology, 36(5–6):605, October 1974. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02463273.

Bergner:1974:SPA

[122] Per-Erik E. Bergner. Stochastic processes and applications in biology and medicine. Part 1 — theory. Part 2 — models. Bulletin of Mathematical Biology, 36(5–6):607–610, October 1974. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02463274.

Deakin:1974:SSM

[123] Michael A. B. Deakin. Stabilité structurelle et morphogénèse. (French) [Structural stability and morphogenesis]. Bulletin of Mathematical Biology, 36(5–6):611–613, October 1974. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02463275.

Rapoport:1974:MPV

[124] Anatol Rapoport. Matematicheskie problemy v biologii. (Russian) [Mathematical problems of biology]. Bulletin of Mathematical Biology, 36(5–6):615–616, October 1974. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02463276.

Anonymous:1974:UTH

[125] Anonymous. The University of Texas Health Science Center at Houston, Division of Continuing Education and The Region V of IEEE announces Eleventh Symposium on Biomathematics and Computer Science in the Life Sciences. Bulletin of Mathematical Biology, 36(5– 6):617, October 1974. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02463277.

Hyver:1975:ERP

[126] C. Hyver. Existence de résonances paramétriques dans des systèmes de transformation d’intérêt biologique. (French) [Existence of parametric resonances in transformation systems of biological interest]. Bulletin of Mathematical Biology, 37(1):1–9, March 1975. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02463487.

Landahl:1975:TSP

[127] H. D. Landahl and B. D. Hansen. A three stage population model with cannibalism. Bulletin of Mathematical Biology, 37(1):11–17, March 1975. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02463488.

Min:1975:DOS

[128] B. Min, W. Welkowitz, S. Fich, D. Jaron, and A. Kantrowitz. Dynamic optimization of in-series cardiac assistance by means of intraaortic balloon pumping. Bulletin of Mathematical Biology, 37(1): 19–35, March 1975. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02463489.

Bouffette:1975:CCA

[129] A. R. Bouffette, C. Biémont, and J. Bouffette. Comparaison des croisements au hasard, prère-sœur et demi frère-sœur par le coefficient de dépression. (French) [Comparison of random crossings, sister-in-law and half sibling-sister by the coefficient of depression]. Bulletin of Mathematical Biology, 37(1):37–49, March 1975. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02463490.

Popescu:1975:MPC

[130] Aurel I. Popescu and Alexandra A. Popescu. A model of a physico– chemical coupling of biological interest. Bulletin of Mathematical Biology, 37(1):51–57, March 1975. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02463491.

Baer:1975:SAD

[131] Robert M. Baer. Self-assembly and differentiation as models of computability. Bulletin of Mathematical Biology, 37(1):59–69, March 1975. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02463492.

Rosen:1975:BTP

[132] Gerald Rosen and Richard G. Fizell. Bounds on the total population for species governed by reaction–diffusion equations in arbitrary two-dimensional regions. Bulletin of Mathematical Biology, 37 (1):71–78, March 1975. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02463493.

Cope:1975:KLE

[133] Freeman W. Cope. Kinetics of light emission by photosynthetic systems: Second order light decay kinetics means Elovich kinetics in a solid state reaction; 1.5 order light decay kinetics means second order reaction kinetics. Bulletin of Mathematical Biology, 37(1):79–83, March 1975. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02463494.

Fatt:1975:FDV

[134] Irving Fatt. Flow and diffusion in the vitreous body of the eye. Bulletin of Mathematical Biology, 37(1):85–90, March 1975. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02463495.

Deakin:1975:SFK

[135] Michael A. B. Deakin. A standard form for the Kermack–McKendrick epidemic equations. Bulletin of Mathematical Biology, 37(1):91–95, March 1975. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02463496.

Chiou:1975:TSS

[136] W. Chiou. Treatment of singularity for spatial transfer properties of human visual system at threshold. Bulletin of Mathematical Biology, 37 (1):97–100, March 1975. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02463497.

Blum:1975:BR

[137] J. J. Blum, Otto E. Rössler, and Robert Rosen. Book reviews. Bulletin of Mathematical Biology, 37(1):101–107, March 1975. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02463498.

Anonymous:1975:CFL

[138] Anonymous. Conference on formal languages, automata and development. Bulletin of Mathematical Biology, 37(1):109, March 1975. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02463499.

Fredberg:1975:PSG

[139] J. J. Fredberg. Pseuod-sound generation at atherosclerotic constrictions in arteries. Bulletin of Mathematical Biology, 37(1):111, March 1975. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02463500;
http://link.springer.com/content/pdf/10.1007/BF02463500.pdf.

Karreman:1975:AFB

[140] George Karreman. Anthony Francis Bartholomay. Bulletin of Mathematical Biology, 37(2):i–ii, April 1975. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02470617.

Davidson:1975:LGM

[141] M. R. Davidson. Lung gas mixing during expiration following an inspiration of air. Bulletin of Mathematical Biology, 37(2):113–126, April 1975. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02470618.

Mullooly:1975:SKM

[142] John P. Mullooly. Stochastic kinetic models for small reversible systes; the equilibrium distribution. Bulletin of Mathematical Biology, 37(2): 127–138, April 1975. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02470619.

Dutt:1975:APT

[143] Ranabir Dutt, P. K. Ghosh, and B. B. Karmakar. Application of perturbation theory to the nonlinear Volterra–Gause–Witt model for prey–predator interaction. Bulletin of Mathematical Biology, 37(2): 139–146, April 1975. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02470620.

Chuang:1975:MAC

[144] Shui-Nan Chuang and Harris H. Lloyd. Mathematical analysis of cancer chemotherapy. Bulletin of Mathematical Biology, 37(2):147– 160, April 1975. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02470621.

Clark:1975:IBM

[145] Michael R. Clark, V. V. Krishnan, and Lawrence Stark. Inners and biocontrol models. Bulletin of Mathematical Biology, 37(2): 161–180, April 1975. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02470622.

Rossler:1975:MSN

[146] Otto E. Rössler. A multivibrating switching network in homogeneous kinetics. Bulletin of Mathematical Biology, 37(2):181–192, April 1975. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02470623.

Richter:1975:STS

[147] Peter H. Richter. Stochastic theory of a selection game. Bulletin of Mathematical Biology, 37(2):193–213, April 1975. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02470624.

Rosen:1975:OSS

[148] Robert Rosen. Organismic sets: Some reflections on the nature of life and society. Bulletin of Mathematical Biology, 37(2):215–218, April 1975. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02470625.

Anonymous:1975:SMB

[149] Anonymous. The society for mathematical biology announces central regional meeting. Bulletin of Mathematical Biology, 37(2):219, April 1975. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02470626.

Anonymous:1975:EP

[150] Anonymous. Employment page. Bulletin of Mathematical Biology, 37(2):220, April 1975. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02470627.

Deakin:1975:E

[151] M. A. B. Deakin. Errata. Bulletin of Mathematical Biology, 37 (2):221, April 1975. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02470628;   http://link.springer.com/content/pdf/10.
1007/BF02470628.pdf.

Ryan:1975:ACS

[152] J. Philip Ryan. Aspects of the Clausius–Shannon identity: Emphasis on the components of transitive information in linear, branched and composite physical systems. Bulletin of Mathematical Biology, 37(3): 223–254, June 1975. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02461443.

Reiner:1975:TRD

[153] John M. Reiner. Theory of RNA–DNA hybridization — I. Stoichiometry at equilibrium. Bulletin of Mathematical Biology, 37(3): 255–268, June 1975. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02461444.

Purdue:1975:STC

[154] Peter Purdue. Stochastic theory of compartments: an open, two compartment, reversible system with independent Poisson arrivals. Bulletin of Mathematical Biology, 37(3):269–275, June 1975. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02461445.

Rosen:1975:SSN

[155] Gerald Rosen. Solutions to systems of nonlinear reaction–diffusion equations. Bulletin of Mathematical Biology, 37(3):277–289, June 1975. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02461446.

DeAngelis:1975:EPP

[156] Donald L. De Angelis. Estimates of predator–prey limit cycles. Bulletin of Mathematical Biology, 37(3):291–299, June 1975. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02461447.

Cantraine:1975:KTF

[157] F. R. L. Cantraine and B. Dewandre. Kinetics of the thyroid follicle lumen diffusion subsystem: Application to iodide accumulation by the thyroid follicle in vitro. Bulletin of Mathematical Biology, 37(3): 301–321, June 1975. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02461448.

Auchmuty:1975:BAN

[158] J. F. G. Auchmuty and G. Nicolis. Bifurcation analysis of nonlinear reaction–diffusion equations — I. Evolution equations and the steady state solutions. Bulletin of Mathematical Biology, 37(4):323– 365, August 1975. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459519.

Oaten:1975:BMS

[159] A. Oaten, C. E. M. Pearce, and M. E. B. Smyth. Batesian mimicry and signal detection theory. Bulletin of Mathematical Biology, 37(4): 367–387, August 1975. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459520.

Allen:1975:DEP

[160] P. M. Allen. Darwinian evolution and a predator–prey ecology. Bulletin of Mathematical Biology, 37(4):389–405, August 1975. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02459521.

Pavlidis:1975:BCM

[161] Theodosios Pavlidis. Biological clock mutants and parameter changes in mathematical models. Bulletin of Mathematical Biology, 37(4): 407–417, August 1975. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459522.

Comorosan:1975:PBS

[162] Sorin Comorosan. On a possible biological spectroscopy. Bulletin of Mathematical Biology, 37(4):419–425, August 1975. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/article/10.1007/BF02459523.

Davison:1975:SCB

[163] E. J. Davison. Simulation of cell behavior: Normal and abnormal growth. Bulletin of Mathematical Biology, 37(5):427–458, October 1975. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02459513.

Das:1975:SEM

[164] R. N. Das and V. Seshadri. A semi-empirical model for flow of blood and other particulate suspensions through narrow tubes. Bulletin of Mathematical Biology, 37(5):459–470, October 1975. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02459514.

Chow:1975:BFT

[165] Joseph C. F. Chow. Blood flow: Theory, effective viscosity and effects of particle distribution. Bulletin of Mathematical Biology, 37(5): 471–488, October 1975. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459515.

Melbin:1975:TDL

[166] Julius Melbin, R. Gopalakrishnan, and Abraham Noordergraaf. Three dimensional laminar flow in distorting, axisymmetric, axially varying vessels. Bulletin of Mathematical Biology, 37(5):489–504, October 1975. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02459516.

Cardenas:1975:TDRa

[167] M. Cardenas and J. H. Matis. On the time-dependent reversible stochastic compartmental model — I. The general two-compartment system. Bulletin of Mathematical Biology, 37(5):505–519, October 1975. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02459517.

Patil:1975:CSC

[168] Mothiram K. Patil and Dhanjoo N. Ghista. Construction, simulation, clinical application and sensitivity analyses of a human left ventricular control system model. Bulletin of Mathematical Biology, 37(5): 521–553, October 1975. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459518.

Cardenas:1975:TDRb

[169] M. Cardenas and J. H. Matis. On the time-dependent reversible stochastic compartmental model — II. A class of n-compartment systems. Bulletin of Mathematical Biology, 37(6):555–564, December 1975. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02459524.

Leguizamon:1975:CEB

[170] Carlos A. Leguizamón. Concept of energy in biological systems. Bulletin of Mathematical Biology, 37(6):565–572, December 1975. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02459525.

Perl:1975:MPF

[171] W. Perl, N. A. Lassen, and R. M. Effros. Matrix proof of flow, volume and mean transit time theorems for regional and compartmental systems. Bulletin of Mathematical Biology, 37(6):573–588, December 1975. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02459526.

Herschkowitz-Kaufman:1975:BAN

[172] M. Herschkowitz-Kaufman. Bifurcation analysis of nonlinear reaction– diffusion equations — II. Steady state solutions and comparison with numerical simulations. Bulletin of Mathematical Biology, 37(6):589– 636, December 1975. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459527.

Babloyantz:1975:MCD

[173] A. Babloyantz and J. Hiernaux. Models for cell differentiation and generation of polarity in diffusion-governed morphogenetic fields. Bulletin of Mathematical Biology, 37(6):637–657, December 1975. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02459528.

Swan:1975:SSH

[174] G. W. Swan. Some strategies for harvesting a single species. Bulletin of Mathematical Biology, 37(6):659–673, December 1975. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02459529.

Leguizamon:1975:TES

[175] Carlos A. Leguizamón. A theory for environmental systems. Bulletin of Mathematical Biology, 37(6):675–689, December 1975. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02459530.

Swan:1976:SLO

[176] G. W. Swan. Solution of linear one-dimensional diffusion equations. Bulletin of Mathematical Biology, 38(1):1–13, January 1976. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02459536.

Jones:1976:TSC

[177] B. L. Jones, R. H. Enns, and S. S. Rangnekar. On the theory of selection of coupled macromolecular systems. Bulletin of Mathematical Biology, 38(1):15–28, January 1976. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459537.

Biscar:1976:PEA

[178] J. P. Biscar. Photon enzyme activation. Bulletin of Mathematical Biology, 38(1):29–38, January 1976. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02459538.

Clark:1976:SCP

[179] Michael R. Clark and Lawrence Stark. Sensitivity of control parameters in a model of saccadic eye tracking and estimation of resultant nervous activity. Bulletin of Mathematical Biology, 38(1):39–57, January 1976. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02459539.

Lee:1976:NEC

[180] Chyuan-Yih Lee and Chun Chiang. Nerve excitations by the coupling of the dipoles and the membrane matrix. Bulletin of Mathematical Biology, 38(1):59–70, January 1976. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459540.

Mayrovitz:1976:ACM

[181] H. N. Mayrovitz, M. P. Wiedeman, and A. Noordergraaf. Analytical characterization of microvascular resistance distribution. Bulletin of Mathematical Biology, 38(1):71–82, January 1976. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02459541.

Anderson:1976:EPE

[182] N. Anderson and A. M. Arthurs. Extremum principles and the elastic and fluid mechanical behaviour of the human eye. Bulletin of Mathematical Biology, 38(1):83–86, January 1976. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02459542.

Rosen:1976:HSS

[183] Robert Rosen. Hysteresis in systems with a single steady state. Bulletin of Mathematical Biology, 38(1):87–92, January 1976. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02459543.

Biscar:1976:PEC

[184] J. P. Biscar. A possible explanation of the Comorosan effects. Bulletin of Mathematical Biology, 38(1):93–94, January 1976. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02459544.

Rosen:1976:MTD

[185] Gerald Rosen. The mathematical theory of diffusion and reaction in permeable catalysts. Bulletin of Mathematical Biology, 38(1):95– 96, January 1976. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459545.

Yodzis:1976:EHC

[186] P. Yodzis. The effects of harvesting on competitive systems. Bulletin of Mathematical Biology, 38(2):97–109, March 1976. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02471750.

Cope:1976:DWF

[187] Freeman W. Cope. Derivation of the Weber–Fechner law and the Loewenstein equation as the steady-state response of an Elovich solid state biological system. Bulletin of Mathematical Biology, 38(2):111– 118, March 1976. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02471751.

Bonchev:1976:SIC

[188] D. Bonchev, D. Kamenski, and V. Kamenska. Symmetry and information content of chemical structures. Bulletin of Mathematical Biology, 38 (2):119–133, March 1976. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02471752.

Kang:1976:EMR

[189] C. K. Kang and A. C. Eringen. The effect of microstructure on the rheological properties of blood. Bulletin of Mathematical Biology, 38 (2):135–159, March 1976. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02471753.

Mon:1976:MCS

[190] Eduardo Mon and James S. Ultman. Monte Carlo simulation of simultaneous gas flow and diffusion in an asymmetric distal pulmonary airway model. Bulletin of Mathematical Biology, 38(2):161–192, March 1976. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02471754.

Rosen:1976:CAS

[191] Gerald Rosen. Construction of analytically solvable models for interacting species. Bulletin of Mathematical Biology, 38(2):193–197, March 1976. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02471755.

Hearon:1976:PLP

[192] John Z. Hearon. Properties of the Leslie population matrix. Bulletin of Mathematical Biology, 38(2):199–203, March 1976. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02471756.

Swan:1976:RMS

[193] G. W. Swan. Reduction of model for the spread of cancer in the uterus; arbitrary specific growth rate. Bulletin of Mathematical Biology, 38(2): 205–207, March 1976. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02471757.

Martinez:1976:BR

[194] Hugo M. Martinez, Harold J. Morowitz, and Leslie E. Blumenson. Book review. Bulletin of Mathematical Biology, 38(2):209–217, March 1976. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02471758.

Warner:1976:DEP

[195] William H. Warner and Theodore A. Wilson. Distribution of end-points of a branching network with decaying branch length. Bulletin of Mathematical Biology, 38(3):219–237, May 1976. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02459556.

Mirsky:1976:PEW

[196] I. Mirsky, R. F. Janz, B. R. Kubert, B. Korecky, and G. C. Taichman. Passive elastic wall stiffness of the left ventricle: A comparison between linear theory and large deformation theory. Bulletin of Mathematical Biology, 38(3):239–251, May 1976. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02459557.

Fischer:1976:MPA

[197] Burkhart Fischer and Jürgen Krüger. Mathematical principles in afferent visual neurons: Differentiation, integration and transient proportionality related to receptive fields and shift-effect. Bulletin of Mathematical Biology, 38(3):253–267, May 1976. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02459558.

Lin:1976:OTV

[198] S. H. Lin. Oxygen tension in the in vivo cornea. Bulletin of Mathematical Biology, 38(3):269–275, May 1976. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02459559.

Hodson:1976:CEQ

[199] C. Hodson and L. Y. Wei. Comparative evaluation of quantum theory of nerve excitation. Bulletin of Mathematical Biology, 38(3): 277–293, May 1976. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459560.

Stubbs:1976:CRN

[200] Derek F. Stubbs and Phillip I. Good. Connectivity in random networks. Bulletin of Mathematical Biology, 38(3):295–304, May 1976. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02459561.

Horsfield:1976:SMP

[201] Keith Horsfield. Some mathematical properties of branching trees with application to the respiratory system. Bulletin of Mathematical Biology, 38(3):305–315, May 1976. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459562.

Cull:1976:PTP

[202] Paul Cull and Andrew Vogt. The period of total population. Bulletin of Mathematical Biology, 38(3):317–319, May 1976. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/article/10.1007/BF02459563.

Lovtrup:1976:ADT

[203] Søren Løvtrup and Björn von Sydow. An addendum to “D’Arcy Thompson’s theorems and the shape of the molluscan shell”. Bulletin of Mathematical Biology, 38(3):321–322, May 1976. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02459564. See [116].

Thompson:1976:GSP

[204] E. A. Thompson. The genetic structure of populations, biomathematics. Bulletin of Mathematical Biology, 38(3):323–324, May 1976. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02459565.

Auchmuty:1976:BAR

[205] J. F. G. Auchmuty and G. Nicolis. Bifurcation analysis of reaction– diffusion equations — III. Chemical oscillations. Bulletin of Mathematical Biology, 38(4):325–350, July 1976. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02462209.

Aizawa:1976:TOI

[206] Yöji Aizawa and Yonosuke Kobatake. A theory of oscillatory instability in two-state systems in relation to nerve excitation. Bulletin of Mathematical Biology, 38(4):351–358, July 1976. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02462210.

Bobisud:1976:CES

[207] L. E. Bobisud. Cannibalism as an evolutionary strategy. Bulletin of Mathematical Biology, 38(4):359–368, July 1976. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/article/10.1007/BF02462211. See comment [350].

Shitzer:1976:TBB

[208] A. Shitzer and M. K. Kleiner. Thermal behavior of biological tissues — a general analysis. Bulletin of Mathematical Biology, 38(4): 369–386, July 1976. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462212.

Melnyk:1976:HRP

[209] T. W. Melnyk, I. W. Richardson, A. A. Simpson, and William R. Smith. Hypothalamic regulation of pituitary secretion of luteinizing hormone. Bulletin of Mathematical Biology, 38(4):387–400, July 1976. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02462213.

Janz:1976:SIC

[210] R. F. Janz and R. J. Waldron. Some implications of a constant fiber stress hypothesis in the diastolic left ventricle. Bulletin of Mathematical Biology, 38(4):401–413, July 1976. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02462214.

Thingstad:1976:CBB

[211] Tron F. Thingstad. On the connection between bacterial growth and division. Bulletin of Mathematical Biology, 38(4):415–423, July 1976. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02462215.

Copeland:1976:ATD

[212] Arthur R. Copeland. Axonal transport — I. Diffusion. Bulletin of Mathematical Biology, 38(4):425–433, July 1976. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02462216.

Copeland:1976:ATI

[213] Arthur R. Copeland. Axonal transport — II. Convection. Bulletin of Mathematical Biology, 38(4):435–444, July 1976. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/article/10.1007/BF02462217.

Huddleston:1976:FDM

[214] John V. Huddleston. The fitting of dynamic models to observed data. Bulletin of Mathematical Biology, 38(4):445–452, July 1976. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02462218.

MacDonald:1976:ADF

[215] N. MacDonald. Application of the describing function to the Danziger– Elmergreen equations. Bulletin of Mathematical Biology, 38(4):453– 458, July 1976. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462219.

Dutt:1976:AHJ

[216] Ranabir Dutt. Application of Hamilton–Jacobi theory to the Lotka– Volterra oscillator. Bulletin of Mathematical Biology, 38(4):459–465, July 1976. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462220.

Matis:1976:NUS

[217] J. H. Matis, R. L. Kodell, and M. Cardenas. A note on the use of a stochastic mammillary compartmental model as an environmental safety model. Bulletin of Mathematical Biology, 38(5):467– 478, September 1976. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459546.

Richardson:1976:DFT

[218] I. W. Richardson and S. Miekisz. A dynamical field theory for dissipative thermal systems. The first and second laws of irreversible thermodynamics. Bulletin of Mathematical Biology, 38(5):479–496, September 1976. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02459547.

Chiang:1976:TAP

[219] Chun Chiang. A theory of ambiguous pattern perception. Bulletin of Mathematical Biology, 38(5):497–504, September 1976. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02459548.

Lee:1976:AFM

[220] E. T. Lee. Algorithms for finding most dissimilar images with possible applications to chromosome classification. Bulletin of Mathematical Biology, 38(5):505–516, September 1976. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02459549.

Yamamura:1976:MAS

[221] N. Yamamura. A mathematical approach to spatial distribution and temporal succession in plant communities. Bulletin of Mathematical Biology, 38(5):517–526, September 1976. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02459550.

Meyer:1976:DTM

[222] J. P. Meyer and M. D. Kostin. Diffusion through a membrane: Approach to equilibrium. Bulletin of Mathematical Biology, 38(5):527– 534, September 1976. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459551.

Walker:1976:NIS

[223] Janet Walker, Anne Van Nypelseer, and W. E. Langlois. Numerical integration of a stochastic model for the Volterra–Lotka reaction. Bulletin of Mathematical Biology, 38(5):535–546, September 1976. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02459552.

Leguizamon:1976:ODR

[224] Carlos A. Leguizamón. Other developments regarding the concept of energy in biological systems. Bulletin of Mathematical Biology, 38(5): 547–563, September 1976. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459553.

Rosen:1976:LNB

[225] Robert Rosen. Lecture notes in biomathematics 4. Physics and mathematics of the nervous system. Bulletin of Mathematical Biology, 38(5): 565–568, September 1976. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459554.

Rosen:1976:CAB

[226] Robert Rosen. The chief abstractions of biology. Bulletin of Mathematical Biology, 38(5):569–571, September 1976. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02459555.

Liron:1976:PFE

[227] Nadav Liron. On peristaltic flow and its efficiency. Bulletin of Mathematical Biology, 38(6):573–596, November 1976. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02458635.

Mullooly:1976:SEM

[228] John P. Mullooly. Stochastic equilibria for multivariate reaction systems. Bulletin of Mathematical Biology, 38(6):597–622, November 1976. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02458636.

Fizell:1976:BPD

[229] Richard G. Fizell and Paul E. Rubin. Bounds on the populations in the Dreitlein–Smoes model of oscillatory kinetic systems. Bulletin of Mathematical Biology, 38(6):623–631, November 1976. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02458637.

Johnson:1976:TRM

[230] Webster R. Johnson and Abdo H. Abdelmessih. Two-regime model of the perfusion measuring heated thermocouple probe. Bulletin of Mathematical Biology, 38(6):633–642, November 1976. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02458638.

Chow:1976:PTW

[231] P. L. Chow and W. C. Tam. Periodic and traveling wave solutions to Volterra–Lotka equations with diffusion. Bulletin of Mathematical Biology, 38(6):643–658, November 1976. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02458639.

Doroszewski:1976:MMD

[232] Jan Doroszewski, Jadwiga Jakubas, and Wiesław Szlenk. Mathematical model of dynamic adhesion of lymphocytes on a glass-bead column. Bulletin of Mathematical Biology, 38(6):659–669, November 1976. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02458640.

Axel:1976:FLK

[233] Leon Axel. Flow limits of Kedem–Katchalsky equations for fluid flux. Bulletin of Mathematical Biology, 38(6):671–677, November 1976. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02458641.

Radziuk:1976:IEA

[234] J. Radziuk. An integral equation approach to measuring turnover in nonsteady compartmental and distributed systems. Bulletin of Mathematical Biology, 38(6):679–693, November 1976. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02458642.

Kapadia:1976:SCM

[235] Asha Seth Kapadia and Bayliss C. McInnis. A stochastic compartmental model with continuous infusion. Bulletin of Mathematical Biology, 38(6): 695–700, November 1976. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458643.

Demiray:1976:LDA

[236] Hilmi Demiray. Large deformation analysis of some basic problems in biophysics. Bulletin of Mathematical Biology, 38(6):701– 712, November 1976. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458644.

Marmur:1976:KCD

[237] A. Marmur, W. N. Gill, and E. Ruckenstein. Kinetics of cell deposition under the action of an external field. Bulletin of Mathematical Biology, 38(6):713–721, November 1976. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02458645.

Kannan:1976:SMM

[238] D. Kannan. On some Markov models of certain interacting populations. Bulletin of Mathematical Biology, 38(6):723–738, November 1976. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02458646.

Rubin:1976:BGD

[239] Paul E. Rubin. Bounds for generalized Dreitlein–Smoes models. Bulletin of Mathematical Biology, 38(6):739–743, November 1976. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02458647.

Aizawa:1977:TOI

[240] Yoji Aizawa and Yonosuke Kobatake. A theory of oscillatory instability in two-state systems in relation to nerve excitation. Bulletin of Mathematical Biology, 39(1):i, January 1977. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02460688;   http://link.springer.com/
content/pdf/10.1007/BF02460688.pdf.

Cardenas:1977:TDR

[241] M. Cardenas and J. H. Matis. On the time-dependent reversible stochastic compartmental model — I. The general two-compartment system. Bulletin of Mathematical Biology, 39(1):ii, January 1977. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02460689;
http://link.springer.com/content/pdf/10.1007/BF02460689.pdf.

Chang:1977:PMN

[242] Donald C. Chang. A physical model of nerve axon — I. Ionic distribution, potential profile, and resting potential. Bulletin of Mathematical Biology, 39(1):1–22, January 1977. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460678.

Rogers:1977:LMC

[243] Thomas D. Rogers. Local models of cell aggregation kinetics. Bulletin of Mathematical Biology, 39(1):23–42, January 1977. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02460679.

Rissman:1977:LEA

[244] Paul Rissman. The leading edge approximation to the nerve axon problem. Bulletin of Mathematical Biology, 39(1):43–58, January 1977. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02460680.

Lieberman:1977:SMB

[245] Milton Lieberman. A stochastic model based upon computer simulation using pre-emption to predict size distribution and species equitability. Bulletin of Mathematical Biology, 39(1):59–72, January 1977. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02460681.

Davidson:1977:IGE

[246] M. R. Davidson. The influence of gas exchange on lung gas concentrations during air breathing. Bulletin of Mathematical Biology, 39(1): 73–86, January 1977. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460682.

Howard:1977:RIG

[247] R. S. Howard and H. Bradner. The V̇_{A}/Q̇ resolution of inert gas dataresolution of inert gas data. Bulletin of Mathematical Biology, 39 (1):87–98, January 1977. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460683.

Rogers:1977:ECS

[248] Thomas D. Rogers. Ecological capacity and species packing. Bulletin of Mathematical Biology, 39(1):99–107, January 1977. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02460684.

Morris:1977:ACE

[249] Hedley C. Morris. Almost critical ecosystems. Bulletin of Mathematical Biology, 39(1):109–116, January 1977. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460685.

Liotta:1977:DMT

[250] Lance A. Liotta, Gerald M. Saidel, and Jerome Kleinerman. Diffusion model of tumor vascularization and growth. Bulletin of Mathematical Biology, 39(1):117–128, January 1977. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460686.

Papini:1977:ESN

[251] G. Papini. An exactly soluble non-linear dynamical system. Bulletin of Mathematical Biology, 39(1):129–131, January 1977. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02460687.

McMorris:1977:CBQ

[252] F. R. McMorris. On the compatibility of binary qualitative taxonomic characters. Bulletin of Mathematical Biology, 39(2):133–138, March 1977. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02462853.

Conrad:1977:FSB

[253] Michael Conrad. Functional significance of biological variability. Bulletin of Mathematical Biology, 39(2):139–156, March 1977. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02462854.

Crowley:1977:SDS

[254] Philip H. Crowley. Spatially distributed stochasticity and the constancy of ecosystems. Bulletin of Mathematical Biology, 39(2):157– 166, March 1977. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462855.

Bergner:1977:ATB

[255] Per-Erik E. Bergner. Application of a theorem by bright to the generalized tracer system. Bulletin of Mathematical Biology, 39(2): 167–178, March 1977. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462856.

Rescigno:1977:SLI

[256] Aldo Rescigno. The struggle for life — IV two predators sharing a prey. Bulletin of Mathematical Biology, 39(2):179–185, March 1977. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02462857.

Cronin:1977:MAP

[257] Jane Cronin. Mathematical aspects of peridic catatonic schizophrenia. Bulletin of Mathematical Biology, 39(2):187–199, March 1977. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02462858.

DeLisi:1977:ISN

[258] Charles DeLisi and Aldo Rescigno. Immune surveillance and neoplasia — 1 a minimal mathematical model. Bulletin of Mathematical Biology, 39(2):201–221, March 1977. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462859.

Barnwell:1977:MMD

[259] G. M. Barnwell and F. S. Stafford. A mathematical model of decisionmaking neural circuits controlling behavioral model selection. Bulletin of Mathematical Biology, 39(2):223–237, March 1977. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02462860.

Rosen:1977:EMI

[260] Robert Rosen. On an empirical method for identifying the observables involved in dynamical interactions. Bulletin of Mathematical Biology, 39 (2):239–244, March 1977. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462861.

Vayo:1977:OFT

[261] H. Westcott Vayo, Dhanjoo N. Ghista, and K. B. Chandran. Oscillatory flow in thin-walled curved elastic tubes — summary. Bulletin of Mathematical Biology, 39(2):245–248, March 1977. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02462862.

Baianu:1977:LMG

[262] I. C. Bâianu, Ph.D., M.Inst.P. A logical model of genetic activities in Łukasiewicz algebras: The non-linear theory. Bulletin of Mathematical Biology, 39(2):249–258, March 1977. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02462863.

Deutsch:1977:TSL

[263] Sid Deutsch. Two-stage lateral inhibition for auditory selectivity. Bulletin of Mathematical Biology, 39(2):259–266, March 1977. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02462864.

Karreman:1977:CSA

[264] George Karreman. Cooperative specific adsorption phenomena in biology. Bulletin of Mathematical Biology, 39(2):267–273, March 1977. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02462865.

Rossler:1977:CAK

[265] Otto E. Rössler. Chaos in abstract kinetics: Two prototypes. Bulletin of Mathematical Biology, 39(2):275–289, March 1977. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02462866.

Landahl:1977:SCS

[266] H. D. Landahl. Some conditions for sustained oscillations in chain processes with feedback inhibition and saturable removal. Bulletin of Mathematical Biology, 39(2):291–296, March 1977. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02462867.

Kempfle:1977:RST

[267] Manfred Kempfle. Relaxation spectra of two-step mechanisms as measured by the NADH-binding to glutamate dehydrogenase. Bulletin of Mathematical Biology, 39(3):297–310, May 1977. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/article/10.1007/BF02462910.

Jones:1977:AEE

[268] B. L. Jones. Analysis of Eigen’s equations for selection of biological molecules with fluctuating mutation rates. Bulletin of Mathematical Biology, 39(3):311–316, May 1977. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02462911.

Swan:1977:OCA

[269] George W. Swan and Thomas L. Vincent. Optimal control analysis in the chemotherapy of IgG multiple myeloma. Bulletin of Mathematical Biology, 39(3):317–337, May 1977. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02462912.

Gatto:1977:SAP

[270] M. Gatto and S. Rinaldi. Stability analysis of predator–prey models via the Liapunov method. Bulletin of Mathematical Biology, 39 (3):339–347, May 1977. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462913.

Portnow:1977:RWM

[271] J. Portnow, A. Talo, and B. J. Hodgson. A random walk model of ovum transport. Bulletin of Mathematical Biology, 39(3):349–357, May 1977. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02462914.

Linkens:1977:SEC

[272] D. A. Linkens. The stability of entrainment conditions for RLC coupled van der Pol oscillators used as a model for intestinal electrical rhythms. Bulletin of Mathematical Biology, 39(3):359–372, May 1977. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02462915.

Rosen:1977:EDS

[273] Gerald Rosen. Effects of diffusion on the stability of the equilibrium in multi-species ecological systems. Bulletin of Mathematical Biology, 39(3):373–383, May 1977. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462916.

Sud:1977:PAB

[274] V. K. Sud, G. S. Sekhon, and R. K. Mishra. Pumping action on blood by a magnetic field. Bulletin of Mathematical Biology, 39(3): 385–390, May 1977. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462917.

Stetson:1977:SNR

[275] R. F. Stetson, J. B. McGuire, and W. A. Hogan. Simulation of nonlinear reaction–diffusion equations. Bulletin of Mathematical Biology, 39(3):391–396, May 1977. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462918.

Leguizamon:1977:TBB

[276] Carlos A. Leguizamón. Transfers between biological and environmental systems. Bulletin of Mathematical Biology, 39(4):397–406, July 1977. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02462919.

Leguizamon:1977:BES

[277] Carlos A. Leguizamón. The bio-environmental system and the realizability of pre-biological systems. Bulletin of Mathematical Biology, 39 (4):407–413, July 1977. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462920.

Cole:1977:MMD

[278] G. W. Cole and N. R. Scott. A mathematical model of the dynamic heat transfer from the respiratory tract of a chicken. Bulletin of Mathematical Biology, 39(4):415–433, July 1977. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02462921.

Swan:1977:EFS

[279] G. W. Swan. Exact fundamental solutions of linear parabolic equations with spatially varying coefficients. Bulletin of Mathematical Biology, 39 (4):435–451, July 1977. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462922.

Norwich:1977:IRS

[280] Kenneth H. Norwich. On the information received by sensory receptors. Bulletin of Mathematical Biology, 39(4):453–461, July 1977. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02462923.

Bernard:1977:UMS

[281] S. R. Bernard. An urn model study of variability within a compartment. Bulletin of Mathematical Biology, 39(4):463–470, July 1977. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02462924.

Yu:1977:PSC

[282] C. P. Yu and K. Chandra. Precipitation of submicron charged particles in human lung airways. Bulletin of Mathematical Biology, 39 (4):471–478, July 1977. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462925.

Rescigno:1977:SLV

[283] Aldo Rescigno. The struggle for life — V. One species living in a limited environment. Bulletin of Mathematical Biology, 39(4): 479–485, July 1977. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462926.

Rescigno:1977:ISN

[284] Aldo Rescigno and Charles DeLisi. Immune surveillance and neoplasia — II. A two-stage mathematical model. Bulletin of Mathematical Biology, 39(4):487–497, July 1977. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462927.

Nishimura:1977:RDS

[285] H. Nishimura. Realization of dependency structures. Bulletin of Mathematical Biology, 39(4):499–503, July 1977. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02462928.

Auchmuty:1977:BAN

[286] J. F. G. Auchmuty and G. Nicolis. Bifurcation analysis of nonlinear reaction–diffusion equations — I. Bulletin of Mathematical Biology, 39 (4):505–506, July 1977. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462929;   http://link.springer.com/content/pdf/10.
1007/BF02462929.pdf.

Anonymous:1977:E

[287] Anonymous. Errata. Bulletin of Mathematical Biology, 39(4):507, July 1977. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
BF02462930;   http://link.springer.com/content/pdf/10.1007/
BF02462930.pdf.

Uylings:1977:ODB

[288] H. B. M. Uylings. Optimization of diameters and bifurcation angles in lung and vascular tree structures. Bulletin of Mathematical Biology, 39(5):509–520, September 1977. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02461198.

Rosen:1977:PCM

[289] Gerald Rosen. Physio-chemical mathematical theory for transitions in cellular metabolism: an aspect of carcinogenesis. Bulletin of Mathematical Biology, 39(5):521–532, September 1977. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02461199.

Weiner:1977:STC

[290] D. L. Weiner and P. Purdue. The stochastic theory of compartments — a mammillary system. Bulletin of Mathematical Biology, 39(5):533– 542, September 1977. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02461200.

Butler:1977:GFC

[291] James P. Butler. The Green’s function for the convection-diffusion equation in an analytic lung model. Bulletin of Mathematical Biology, 39(5): 543–563, September 1977. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02461201.

Singh:1977:ATC

[292] Virendra Singh. Analytical theory of the control equations for protein synthesis in the Goodwin model. Bulletin of Mathematical Biology, 39(5): 565–575, September 1977. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02461202.

Radu:1977:CMM

[293] H. Radu, A. Da Ronch, and V. Caldesi-Valeri. Calcium and myotonia: a mathematical model. Bulletin of Mathematical Biology, 39(5): 577–585, September 1977. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02461203.

Karreman:1977:KIC

[294] George Karreman and Christian J. Lambertsen. Kinetics of isobaric counterdiffusion. Bulletin of Mathematical Biology, 39(5):587– 595, September 1977. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02461204.

Eisen:1977:SAN

[295] Martin Eisen and John Schiller. Stability analysis of normal and neoplastic growth. Bulletin of Mathematical Biology, 39(5):597–605, September 1977. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02461205.

Cohoon:1977:PLI

[296] David K. Cohoon and James B. Stewart. Prediction of light-induced transient thermal response of the retina. Bulletin of Mathematical Biology, 39(5):607–612, September 1977. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02461206.

Shonkwiler:1977:CGD

[297] R. Shonkwiler. Consequences of the growth and division model for mitochondria assuming no cellular control. Bulletin of Mathematical Biology, 39(5):613–618, September 1977. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02461207.

Varma:1977:ESS

[298] V. S. Varma. Exact solutions for a special prey–predator or competing species system. Bulletin of Mathematical Biology, 39(5):619– 622, September 1977. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02461208.

Dutt:1977:GAB

[299] Ranabir Dutt and Prabir Kumar Ghosh. A graphical analysis of biological evolution in a prey–predator ecology. Bulletin of Mathematical Biology, 39(5):623–627, September 1977. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02461209.

Rosen:1977:SSM

[300] Robert Rosen. Structural stability and morphogenesis. Bulletin of Mathematical Biology, 39(5):629–632, September 1977. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/article/10.1007/BF02461210.

Anonymous:1977:ICE

[301] Anonymous. International conference on energy use management. Bulletin of Mathematical Biology, 39(5):633, September 1977. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02461211.

Anonymous:1977:HLL

[302] Anonymous. Henry Laurence Lucas, Jr. 1916–1977. Bulletin of Mathematical Biology, 39(6):635–636, November 1977. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/article/10.1007/BF02461773.

Karreman:1977:PMB

[303] George Karreman. Perspectives of mathematical biology presidential address 1976 annual program of the society for mathematical biology. Bulletin of Mathematical Biology, 39(6):637–641, November 1977. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02461774.

Lakshminarayanaiah:1977:EMS

[304] N. Lakshminarayanaiah. Evaluation of membrane surface charge density: a discussion of some models. Bulletin of Mathematical Biology, 39(6): 643–662, November 1977. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02461775.

Rosen:1977:OBS

[305] Robert Rosen. Observation and biological systems. Bulletin of Mathematical Biology, 39(6):663–678, November 1977. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/article/10.1007/BF02461776.

Kresch:1977:CSS

[306] Edward Kresch. Cross-sectional shape of flexible tubes. Bulletin of Mathematical Biology, 39(6):679–691, November 1977. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02461777.

Mode:1977:STB

[307] Charles J. Mode. Some techniques in building and validating stochastic models of human reproduction. Bulletin of Mathematical Biology, 39(6): 693–704, November 1977. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02461778.

DeLisi:1977:DAR

[308] Charles DeLisi. Detection and analysis of recognition and selection in the immune response. Bulletin of Mathematical Biology, 39(6):705– 719, November 1977. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02461779.

Ling:1977:TCS

[309] Gilbert N. Ling and Kim Peterson. A theory of cell swelling in high concentrations of KCl and other chloride salts. Bulletin of Mathematical Biology, 39(6):721–741, November 1977. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02461780.

Liron:1977:PFE

[310] N. Liron. On peristaltic flow and its efficiency. Bulletin of Mathematical Biology, 39(6):743, November 1977. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02461781;   http://link.springer.com/
content/pdf/10.1007/BF02461781.pdf.

Woodcock:1978:GST

[311] A. E. R. Woodcock. On the geometry of space-and time-equivalent catastrophes. Bulletin of Mathematical Biology, 40(1):1–25, January 1978. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02463127.

Glass:1978:PLC

[312] Leon Glass and Joel S. Pasternack. Prediction of limit cycles in mathematical models of biological oscillations. Bulletin of Mathematical Biology, 40(1):27–44, January 1978. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02463128.

Rossler:1978:DBS

[313] Otto E. Rössler. Deductive biology — some cautious steps. Bulletin of Mathematical Biology, 40(1):45–58, January 1978. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02463129.

Celeste:1978:RGR

[314] Ruby Celeste, Eugene Ackerman, Laël C. Gatewood, Clayton Reynolds, and George D. Molnar. The role of glucagon in the regulation of blood glucose: Model studies. Bulletin of Mathematical Biology, 40 (1):59–77, January 1978. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02463130.

Doubek:1978:LER

[315] Erich Doubek, Jr. Least energy regulation of the arterial system. Bulletin of Mathematical Biology, 40(1):79–93, January 1978. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02463131.

Weygandt:1978:PFR

[316] C. N. Weygandt, R. H. Cox, G. Karreman, and M. L. Cole. Pressure-flow relationships in a model for the arterial system. Bulletin of Mathematical Biology, 40(1):95–105, January 1978. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02463132.

Pao:1978:KDS

[317] C. V. Pao. The kinetics of diffused substances in some biological systems. Bulletin of Mathematical Biology, 40(1):107–121, January 1978. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02463133.

Ross:1978:NDP

[318] George G. Ross. A note on dynamics of population with historydependent birth rate. Bulletin of Mathematical Biology, 40(1):123– 131, January 1978. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02463134.

vanderVaart:1978:CPS

[319] H. Robert van der Vaart. Conditions for periodic solutions of Volterra differential systems. Bulletin of Mathematical Biology, 40(2):133– 160, March 1978. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02461432.

Danielson:1978:TFT

[320] D. A. Danielson. Tension field theories for soft tissues. Bulletin of Mathematical Biology, 40(2):161–182, March 1978. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02461433.

Klinger:1978:HTP

[321] H. G. Klinger. Heat transfer in perfused biological tissue — II. The “macroscopic” temperature distribution in tissue. Bulletin of Mathematical Biology, 40(2):183–199, March 1978. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02461434.

Collin:1978:NDS

[322] R. E. Collin and R. Plonsey. A note on dipole sources in conducting biological tissues. Bulletin of Mathematical Biology, 40(2): 201–209, March 1978. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02461435.

Stephenson:1978:ATB

[323] John L. Stephenson. Analysis of the transient behavior of kidney models. Bulletin of Mathematical Biology, 40(2):211–221, March 1978. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02461436.

Friedland:1978:HMA

[324] Allan B. Friedland. A hydrodynamic model of aqueous flow in the posterior chamber of the eye. Bulletin of Mathematical Biology, 40(2): 223–235, March 1978. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02461437.

Thakur:1978:STC

[325] Ajit K. Thakur and Aldo Rescigno. On the stochastic theory of compartments: III. General time-dependent reversible systems. Bulletin of Mathematical Biology, 40(2):237–246, March 1978. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02461438.

Chiang:1978:NIE

[326] Chun Chiang. On the nerve impulse equation: The dynamic responses of nerve impulse. Bulletin of Mathematical Biology, 40(2):

247–255, March 1978. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02461439.

Porati:1978:ACC

[327] M. I. Granero Porati, A. Porati, and A. Vecli. Analytical conditions for the conservative form of the ecological equations. Bulletin of Mathematical Biology, 40(2):257–264, March 1978. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02461440.

Bernard:1978:MMD

[328] S. R. Bernard. Metabolic model and dosimetric data for copper. Bulletin of Mathematical Biology, 40(2):265–269, March 1978. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02461441.

Walter:1978:MMM

[329] Charles Walter. Mathematical models of metabolic regulation, Symposia Biologica Hungarica. Bulletin of Mathematical Biology, 40(2): 271–272, March 1978. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02461442.

Garner:1978:TBS

[330] Jackie B. Garner, Kenny S. Crump, and John L. Stephenson. Transient behaviour of the single loop solute cycling model of the renal medulla. Bulletin of Mathematical Biology, 40(3):273–300, May 1978. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02461602.

Richardson:1978:DFT

[331] I. W. Richardson and S. Miekisz. A dynamical field theory for dissipative systems: The hierarchical structure of field thermodynamics. Bulletin of Mathematical Biology, 40(3):301–318, May 1978. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/article/10.1007/BF02461603. See comment [340].

Walsh:1978:OCP

[332] G. R. Walsh. Optimal control of pests in the presence of predators. Bulletin of Mathematical Biology, 40(3):319–333, May 1978. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02461604.

Bojadziev:1978:KBM

[333] G. Bojadziev. The Krylov–Bogoliubov–Mitropolskii method applied to models of population dynamics. Bulletin of Mathematical Biology, 40 (3):335–345, May 1978. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02461605.

Christensen:1978:LTM

[334] Mark J. Christensen and R. Shonkwiler. Lag time in microbe growth as an age-dependent branching process with two phase types. Bulletin of Mathematical Biology, 40(3):347–358, May 1978. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/article/10.1007/BF02461606.

Vit:1978:ASI

[335] K. Vit. Analytical solution to isometric mechanogram of Hill’s model of striated muscle. Bulletin of Mathematical Biology, 40(3): 359–368, May 1978. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02461607.

Bertrand:1978:RTM

[336] P. Bertrand, G. Le Cardinal, and E. Walter. Reconstitution of the transition matrix from experimental data. Bulletin of Mathematical Biology, 40(3):369–376, May 1978. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02461608.

Sanchez:1978:LAD

[337] David A. Sánchez. Linear age-dependent population growth with harvesting. Bulletin of Mathematical Biology, 40(3):377–385, May 1978. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02461609.

Kerner:1978:MSC

[338] Edward H. Kerner. Multiple speciation, competitive exclusion, evolutionary pattern, and the grand ensemble in Volterra eco-dynamics. Bulletin of Mathematical Biology, 40(3):387–410, May 1978. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02461610.

Schoepfle:1978:ICP

[339] Gordon M. Schoepfle, Warren S. Rehm, and John H. Young. Ionic concentration profiles and charge distribution for the domain of a polarized spherical cell. Bulletin of Mathematical Biology, 40(3): 411–424, May 1978. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02461611.

Rosen:1978:CRM

[340] Robert Rosen. A comment on the Richardson–Miekisz field theory of thermodynamics. Bulletin of Mathematical Biology, 40(3): 425–426, May 1978. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02461612. See [331].

Weiner:1978:STC

[341] D. L. Weiner and P. Purdue. The stochastic theory of compartments — a mammillary system. Bulletin of Mathematical Biology, 40(3):427, May 1978. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02461613;   http://link.springer.com/content/pdf/10.
1007/BF02461613.pdf.

Deakin:1978:EAC

[342] Michael A. B. Deakin. An elementary approach to catastrophe theory. Bulletin of Mathematical Biology, 40(4):429–450, July 1978. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02460770.

Mirsky:1978:MBV

[343] I. Mirsky, P. L. McGill, and R. F. Janz. Mechanical behaviour of ventricular aneurysms. Bulletin of Mathematical Biology, 40(4): 451–464, July 1978. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460771.

Mirsky:1978:EEP

[344] Israel Mirsky. Effects of external pressures on the pressure-volume relation of the left ventricle. Bulletin of Mathematical Biology, 40(4): 465–482, July 1978. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460772.

Smeach:1978:SAP

[345] Stephen C. Smeach and Albert Rust. A stochastic approach to predator– prey models. Bulletin of Mathematical Biology, 40(4):483–498, July 1978. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02460773.

Chuang:1978:MAC

[346] Shui-Nan Chuang and T. T. Soong. Mathematical analysis of cancer chemotherapy: The effects of chemotherapeutic agents on the cell cycle traverse. Bulletin of Mathematical Biology, 40(4):499–512, July 1978. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02460774.

Tregonning:1978:ELB

[347] Ken Tregonning and Alan Roberts. Ecosystem-like behavior of a random-interaction model. I. Bulletin of Mathematical Biology, 40(4): 513–524, July 1978. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460775.

Goh:1978:GSC

[348] B. S. Goh. Global stability in a class of prey–predator models. Bulletin of Mathematical Biology, 40(4):525–533, July 1978. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02460776.

Zhitomirskii:1978:TST

[349] I. S. Zhitomirskii and A. A. Kagna. A theoretical study of the temperature distribution in skeletal muscle undergoing rhythmic contractions. Bulletin of Mathematical Biology, 40(4):535–540, July 1978. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02460777.

Stenseth:1978:CBP

[350] Nils Chr Stenseth and Jon Reed. A comment on Bobisud’s paper on evolution of cannibalism. Bulletin of Mathematical Biology, 40(4): 541–545, July 1978. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460778. See [207].

Cohoon:1978:PLI

[351] David K. Cohoon and James B. Stewart. Prediction of light-induced transient thermal response of the retina. Bulletin of Mathematical Biology, 40(4):547, July 1978. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460779;   http://link.springer.com/content/pdf/10.
1007/BF02460779.pdf.

Rosen:1978:DST

[352] Robert Rosen. Dynamical similarity and the theory of biological transformations. Bulletin of Mathematical Biology, 40(5):549– 579, September 1978. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460731.

Scott:1978:TSN

[353] W. R. Scott, D. B. Taulbee, and C. P. Yu. Theoretical study of nasal deposition. Bulletin of Mathematical Biology, 40(5):581– 604, September 1978. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460732.

Bridgeman:1978:DSC

[354] Bruce Bridgeman. Distributed sensory coding applied to simulations of iconic storage and metacontrast. Bulletin of Mathematical Biology, 40(5): 605–623, September 1978. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460733.

Pavlidis:1978:WDM

[355] Theo Pavlidis. What do mathematical models tell us about circadian clocks? Bulletin of Mathematical Biology, 40(5):625–635, September 1978. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460734.

McIlroy:1978:EFD

[356] Douglas K. McIlroy and Brian D. Hahn. Electric field distribution, ionic selectivity and permeability in nerve. Bulletin of Mathematical Biology, 40(5):637–649, September 1978. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460735.

Cobelli:1978:IFT

[357] C. Cobelli and L. M. Morato. On the identification by filtering techniques of a biological n-compartment model in which the transport rate parameters are assumed to be stochastic processes. Bulletin of Mathematical Biology, 40(5):651–660, September 1978. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460736.

Shen:1978:SSS

[358] H. Shen, K. E. Lonngren, and E. L. Forker. Self-similar solution of a distributed model for transcapillary exchange and solute removal by the liver. Bulletin of Mathematical Biology, 40(5):661–669, September 1978. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02460737.

Rosen:1978:SSD

[359] Gerald Rosen. Steady-state distribution of bacteria chemotactic toward oxygen. Bulletin of Mathematical Biology, 40(5):671–674, September 1978. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460738.

Pavlidis:1978:QSB

[360] Theo Pavlidis. Qualitative similarities between the behavior of coupled oscillators and circadian rhythms. Bulletin of Mathematical Biology, 40(6):675–692, November 1978. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460599.

Yu:1978:TCT

[361] C. P. Yu. A two-component theory of aerosol deposition in lung airways. Bulletin of Mathematical Biology, 40(6):693–706, November 1978. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02460600.

Kemper:1978:EAA

[362] John T. Kemper. The effects of asymptomatic attacks on the spread of infectious disease: a deterministic model. Bulletin of Mathematical Biology, 40(6):707–718, November 1978. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460601.

Bhaumik:1978:MKP

[363] Debajyoti Bhaumik, Kamales Bhaumik, Asim Kumar Ray, and Binayak Dutta-Roy. On a model and the kinetics of photo-enhanced enzyme reactions. Bulletin of Mathematical Biology, 40(6):719–726, November 1978. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02460602.

Darvey:1978:KER

[364] I. G. Darvey, A. H. Klotz, and M. Ritter. On the kinetics of enzyme reactions. Bulletin of Mathematical Biology, 40(6):727–734, November 1978. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02460603.

Anderson:1978:CVP

[365] N. Anderson and A. M. Arthurs. Complementary variational principles for the steady-state finite cable model of nerve membranes. Bulletin of Mathematical Biology, 40(6):735–742, November 1978. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02460604.

Schuster:1978:DSU

[366] P. Schuster, K. Sigmund, and R. Wolff. Dynamical systems under constant organization. I. Topological analysis of a family of non-linear differential equations — a model for catalytic hypercycles. Bulletin of Mathematical Biology, 40(6):743–769, November 1978. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/article/10.1007/BF02460605.

Erneux:1978:TTM

[367] T. Erneux, J. Hiernaux, and G. Nicolis. Turing’s theory in morphogenesis. Bulletin of Mathematical Biology, 40(6):771–789, November 1978. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02460606.

Mikulecky:1978:GFE

[368] D. C. Mikulecky. Global flow equations for membrane transport from local equations of motion: I. The general case for (n − 1) nonelectrolyte solutes plus water. Bulletin of Mathematical Biology, 40(6): 791–805, November 1978. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.

1007/BF02460607. See erratum [404].

Rescigno:1978:PLP

[369] Aldo Rescigno. On phase locking of pulse encoders. Bulletin of Mathematical Biology, 40(6):807–821, November 1978. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/article/10.1007/BF02460608.

Friedland:1978:MMT

[370] Allan B. Friedland. A mathematical model of transmural transport of oxygen to the retina. Bulletin of Mathematical Biology, 40(6):823– 837, November 1978. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460609.

Andrietti:1978:IAB

[371] Francesco Andrietti. Interactions among biological systems: an analysis of asymptotic stability. Bulletin of Mathematical Biology, 40(6): 839–851, November 1978. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460610.

Rosen:1978:GRP

[372] Gerald Rosen. Galerkin-Ritz procedures for approximate solutions to systems of reaction–diffusion equations. Bulletin of Mathematical Biology, 40(6):853–863, November 1978. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460611.

vanderWerff:1978:SSB

[373] Terry J. van der Werff and Horst E. Wilhelm. Steady state bifurcation analysis of reaction–diffusion equations — a critique. Bulletin of Mathematical Biology, 40(6):865–872, November 1978. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02460612.

Rosen:1978:MMB

[374] Robert Rosen. Mathematical models in biological discovery. Lecture notes in biomathematics. Bulletin of Mathematical Biology, 40(6):873– 875, November 1978. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460613.

Smetanic:1979:ADP

[375] Ya. S. Smetanič and R. V. Polozov. On the algorithms for determining the primary structure of biopolymers. Bulletin of Mathematical Biology, 41(1):1–20, January 1979. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02547921.

Erneux:1979:BDMa

[376] T. Erneux and M. Herschkowitz-Kaufman. Bifurcation diagram of a model chemical reaction — I. Stability changes of time-periodic solutions. Bulletin of Mathematical Biology, 41(1):21–38, January 1979. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02547922.

Kresch:1979:CSA

[377] Edward Kresch. Cross-sectional area of flexible tubes. Bulletin of Mathematical Biology, 41(1):39–52, January 1979. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02547923.

Crump:1979:SSF

[378] Kenny S. Crump and Wan-Shin C. O’Young. Some stochastic features of bacterial constant growth apparatus. Bulletin of Mathematical Biology, 41(1):53–66, January 1979. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02547924.

Freedman:1979:SAP

[379] H. I. Freedman. Stability analysis of a predator–prey system with mutual interference and density-dependent death rates. Bulletin of Mathematical Biology, 41(1):67–78, January 1979. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02547925.

Kenyon:1979:MMW

[380] Douglas E. Kenyon. A mathematical model of water flux through aortic tissue. Bulletin of Mathematical Biology, 41(1):79–90, January 1979. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02547926.

Becus:1979:SPPa

[381] Georges A. Bécus. Stochastic prey–predator relationships: a random differential equation approach. Bulletin of Mathematical Biology, 41(1): 91–100, January 1979. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02547927.

Woodcock:1979:CTC

[382] A. E. R. Woodcock. Catastrophe theory and cellular determination, transdetermination and differentiation. Bulletin of Mathematical Biology, 41(1):101–117, January 1979. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02547928.

Jones:1979:SSM

[383] B. L. Jones. Some selection models in catalytically coupled selfreproducing macromolecular systems. Bulletin of Mathematical Biology, 41 (1):119–127, January 1979. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02547929.

Rogers:1979:DDM

[384] Thomas D. Rogers. Diversity and diffusion in multidimensional niche space. Bulletin of Mathematical Biology, 41(2):129–138, March 1979. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02460873.

Suga:1979:MOC

[385] Hiroyuki Suga. Minimal oxygen consumption and optimal contractility of the heart: Theoretical approach to principle of physiological control of contractility. Bulletin of Mathematical Biology, 41(2): 139–150, March 1979. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460874.

Lin:1979:NDB

[386] S. H. Lin. Nonlinear diffusion in biological systems. Bulletin of Mathematical Biology, 41(2):151–162, March 1979. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02460875.

Peskoff:1979:EPT

[387] A. Peskoff. Electric potential in three-dimensional electrically syncytial tissues. Bulletin of Mathematical Biology, 41(2):163–181, March 1979. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02460876.

Peskoff:1979:EPC

[388] A. Peskoff. Electric potential in cylindrical syncytia and muscle fibers. Bulletin of Mathematical Biology, 41(2):183–192, March 1979. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02460877.

Babloyantz:1979:SOB

[389] A. Babloyantz and L. K. Kaczmarek. Self-organization in biological systems with multiple cellular contacts. Bulletin of Mathematical Biology, 41(2):193–201, March 1979. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460878.

El-Asfouri:1979:SCM

[390] S. El-Asfouri, B. C. McInnis, and A. S. Kapadia. Stochastic compartmental modeling and parameter estimation with application to cancer treatment follow-up studies. Bulletin of Mathematical Biology, 41(2): 203–215, March 1979. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460879.

Min:1979:PCO

[391] B. G. Min. Performance criteria for optimization of cardiac assistance. Bulletin of Mathematical Biology, 41(2):217–227, March 1979. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02460880.

Kannan:1979:VVP

[392] D. Kannan. Volterra-Verhulst prey–predator systems with time dependent coefficients: Diffusion type approximation and periodic solutions. Bulletin of Mathematical Biology, 41(2):229–251, March 1979. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02460881.

Rosen:1979:CTS

[393] Robert Rosen. Catastrophe theory: Selected papers 1972–1977. Bulletin of Mathematical Biology, 41(2):253–255, March 1979. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02460882.

Fohlmeister:1979:TSN

[394] Jurgen Fohlmeister. A theoretical study of neural adaptation and transient responses due to inhibitory feedback. Bulletin of Mathematical Biology, 41(3):257–282, May 1979. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460812.

Rogers:1979:MS

[395] Thomas D. Rogers and S. C. Trofanenko. On the measurement of shape. Bulletin of Mathematical Biology, 41(3):283–304, May 1979. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02460813.

Zietz:1979:MAO

[396] S. Zietz and C. Nicolini. Mathematical approaches to optimization of cancer chemotherapy. Bulletin of Mathematical Biology, 41(3): 305–324, May 1979. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460814.

Bojadziev:1979:ASD

[397] G. Bojadziev and S. Chan. Asymptotic solutions of differential equations with time delay in population dynamics. Bulletin of Mathematical Biology, 41(3):325–342, May 1979. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460815.

McIlroy:1979:APC

[398] Douglas K. McIlroy. Analysis of potassium conductance in squid giant axons. Bulletin of Mathematical Biology, 41(3):343–356, May 1979. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02460816.

Rosen:1979:TDP

[399] Geral Rosen. Time development of probability distributions for interacting species. Bulletin of Mathematical Biology, 41(3):357–364, May 1979. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02460817.

Vaidhyanathan:1979:NPA

[400] V. S. Vaidhyanathan. Nernst–Planck analog equations and stationary state membrane electric potentials. Bulletin of Mathematical Biology, 41(3):365–385, May 1979. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460818.

Conrad:1979:MAM

[401] Michael Conrad. Mutation-absorption model of the enzyme. Bulletin of Mathematical Biology, 41(3):387–405, May 1979. CODEN BMTBAP.

ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/article/10.1007/BF02460819.

Hatze:1979:TEWa

[402] Herbert Hatze. A teleological explanation of Weber’s law and the motor unit size law. Bulletin of Mathematical Biology, 41(3):407–425, May 1979. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02460820.

Rosen:1979:SCA

[403] Robert Rosen. Some comments on activation and inhibition. Bulletin of Mathematical Biology, 41(3):427–445, May 1979. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/article/10.1007/BF02460821.

Mikulecky:1979:EGF

[404] D. C. Mikulecky. Erratum: Global flow equations for membrane transport from local equations of motion: 1. The general case for (n − 1) nonelectrolyte solutes plus water. Bulletin of Mathematical Biology, 41(3):447, May 1979. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460822;   http://link.springer.com/content/pdf/10.
1007/BF02460822.pdf. See [368].

Jones:1979:SSSa

[405] B. L. Jones. Stability and selection in self-reproducing macromolecular systems. Bulletin of Mathematical Biology, 41(4):449–460, July 1979. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02458323.

Hiernaux:1979:CPC

[406] J. Hiernaux and T. Erneux. Chemical patterns in circular morphogenetic fields. Bulletin of Mathematical Biology, 41(4):461–468, July 1979. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02458324.

Lefever:1979:BFE

[407] R. Lefever and W. Horsthemke. Bistability in fluctuating environments. Implications in tumor immunology. Bulletin of Mathematical Biology, 41(4):469–490, July 1979. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458325.

Matis:1979:CMM

[408] James H. Matis and H. Dennis Tolley. Compartmental models with multiple sources of stochastic variability: The one-compartment, time invariant hazard rate case. Bulletin of Mathematical Biology, 41(4): 491–515, July 1979. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458326.

Mann:1979:SRL

[409] J. E. Mann. Some results on light penetration of an absorbing medium. Bulletin of Mathematical Biology, 41(4):517–523, July 1979. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02458327.

Stephanopoulos:1979:CPM

[410] G. Stephanopoulos and A. G. Fredrickson. Coexistence of photosynthetic microorganisms with growth rates depending on the spectral quality of light. Bulletin of Mathematical Biology, 41(4):525–542, July 1979. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02458328.

Becus:1979:SPPb

[411] Georges A. Bécus. Stochastic prey–predator relationships: a random evolution approach. Bulletin of Mathematical Biology, 41(4): 543–554, July 1979. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458329.

Sperb:1979:MMD

[412] René P. Sperb. On a mathematical model describing the aggregation of amoebae. Bulletin of Mathematical Biology, 41(4):555–571, July 1979. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02458330.

Linkens:1979:MHB

[413] D. A. Linkens. The method of harmonic balance applied to coupled asymmetrical van der Pol oscillators for intestinal modelling. Bulletin of Mathematical Biology, 41(4):573–589, July 1979. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02458331.

Tai:1979:MSN

[414] Ronald C. Tai and Hsin-Kang Chang. A mathematical study of nonequimolar ternary gas diffusion. Bulletin of Mathematical Biology, 41 (4):591–606, July 1979. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458332.

Parthasarathy:1979:MBP

[415] P. R. Parthasarathy and A. Krishnamoorthy. Markov branching processes with varying growth rates. Bulletin of Mathematical Biology, 41 (4):607–610, July 1979. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458333.

McInnis:1979:SCM

[416] B. C. McInnis, S. A. El-Asfouri, and S. A. Kapadia. On stochastic compartmental modeling. Bulletin of Mathematical Biology, 41(4): 611–613, July 1979. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458334.

Deakin:1979:CTA

[417] Michael A. B. Deakin. Catastrophe theory and its applications (surveys and reference works in mathematics 2). Bulletin of Mathematical Biology, 41(4):615–616, July 1979. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458335.

Levine:1979:ELP

[418] Daniel S. Levine. Existence of a limiting pattern for a system of nonlinear equations describing interpopulation competition. Bulletin of Mathematical Biology, 41(5):617–628, September 1979. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02462418.

Mikulecky:1979:GFE

[419] D. C. Mikulecky. Global flow equations for membrane transport from local equations of motion — II. The case of a single nonelectrolyte solute plus water. Bulletin of Mathematical Biology, 41(5):629– 640, September 1979. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462419.

Beretta:1979:SRA

[420] E. Beretta, F. Vetrano, F. Solimano, and C. Lazzari. Some results about nonlinear chemical systems represented by trees and cycles. Bulletin of Mathematical Biology, 41(5):641–664, September 1979. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02462420.

Argemi:1979:QAM

[421] José Argémi, Maurice Gola, and Hélène Chagneux. Qualitative analysis of a model generating long potential waves in Ba-treated nerve cells — I. Reduced systems. Bulletin of Mathematical Biology, 41(5): 665–686, September 1979. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462421.

Sherlock:1979:ABKa

[422] R. A. Sherlock. Analysis of the behaviour of Kauffman binary networks — I. State space description and the distribution of limit cycle lengths. Bulletin of Mathematical Biology, 41(5):687–705, September 1979. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02462422.

Sherlock:1979:ABKb

[423] R. A. Sherlock. Analysis of the behaviour of Kauffman binary networks — II. The state cycle fraction for networks of different connectivities. Bulletin of Mathematical Biology, 41(5):707–724, September 1979. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02462423.

Smith:1979:STI

[424] Truett L. Smith and Chris P. Tsokos. On the study of two interacting populations one of which acts independently of the other. Bulletin of Mathematical Biology, 41(5):725–735, September 1979. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02462424.

Epperson:1979:DGI

[425] J. O. Epperson and J. H. Matis. On the distribution of the general irreversible n-compartmental model having time-dependent transition probabilities. Bulletin of Mathematical Biology, 41(5):737– 749, September 1979. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462425.

Manton:1979:FRE

[426] M. J. Manton. On feasible random ecosystems. Bulletin of Mathematical Biology, 41(5):751–755, September 1979. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02462426.

Mikulecky:1979:BR

[427] Donald C. Mikulecky and Robert Rosen. Book reviews. Bulletin of Mathematical Biology, 41(5):757–759, September 1979. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02462427.

Hatze:1979:TEWb

[428] H. Hatze. A teleological explanation of Weber’s law and the motor unit size law. Bulletin of Mathematical Biology, 41(6):i, November 1979. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02462387;
http://link.springer.com/content/pdf/10.1007/BF02462387.pdf.

Jones:1979:SSSb

[429] B. L. Jones. Selection in systems of self-reproducing macromolecules under the constraint of controlled energy fluxes. Bulletin of Mathematical Biology, 41(6):761–766, November 1979. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02462374.

Erneux:1979:BDMb

[430] T. Erneux and M. Herschkowitz-Kaufman. The bifurcation diagram of a model chemical reaction — II. Two dimensional time-periodic patterns. Bulletin of Mathematical Biology, 41(6):767–790, November 1979. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02462375.

Heukelom:1979:SAM

[431] B. Heukelom, A. van der Stelt, and P. C. Diegenbach. A simple anatomical model of muscle, and the effects of internal pressure. Bulletin of Mathematical Biology, 41(6):791–802, November 1979. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02462376.

Kuppers:1979:SRD

[432] Bernd Küppers. Some remarks on the dynamics of molecular selforganization. Bulletin of Mathematical Biology, 41(6):803–812, November 1979. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462377.

Aizawa:1979:EAS

[433] Yoji Aizawa and Hiroshi Shimizu. An exact analysis of the sliding motion in contractile system: a simplified model of muscle system. Bulletin of Mathematical Biology, 41(6):813–828, November 1979. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02462378.

Mackey:1979:PAI

[434] Michael C. Mackey. Periodic auto-immune hemolytic anemia: an induced dynamical disease. Bulletin of Mathematical Biology, 41(6):829– 834, November 1979. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462379.

Ablowitz:1979:ESF

[435] Mark J. Ablowitz and Anthony Zeppetella. Explicit solutions of Fisher’s equation for a special wave speed. Bulletin of Mathematical Biology, 41(6):835–840, November 1979. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02462380.

Chiang:1979:GCD

[436] Chun Chiang. On the gating currents and the dipole transitions. Bulletin of Mathematical Biology, 41(6):841–848, November 1979. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02462381.

Jones:1979:SMS

[437] B. L. Jones. Some models for selection of biological macromolecules with time varying constraints. Bulletin of Mathematical Biology, 41(6): 849–859, November 1979. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462382.

Fisher:1979:SSC

[438] M. E. Fisher, B. S. Goh, and T. L. Vincent. Some stability conditions for discrete-time single species models. Bulletin of Mathematical Biology, 41(6):861–875, November 1979. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02462383.

Gard:1979:PFW

[439] Thomas C. Gard and Thomas G. Hallam. Persistence in food webs — I. Lotka–Volterra food chains. Bulletin of Mathematical Biology, 41(6): 877–891, November 1979. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462384.

Yeh:1979:MBT

[440] H. C. Yeh. Modeling of biological tree structures. Bulletin of Mathematical Biology, 41(6):893–898, November 1979. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02462385.

Martino:1979:ACN

[441] R. Martino. Analytical considerations on a new class of functions for the mathematical description of the human Erythrocyte’s profile during the osmotic, “disk-sphere” transition. Bulletin of Mathematical Biology, 41(6):899–904, November 1979. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02462386.

Schum:1980:TEA

[442] Michael Schum and Hsu-Chi Yeh. Theoretical evaluation of aerosol deposition in anatomical models of mammalian lung airways. Bulletin of Mathematical Biology, 42(1):1–15, January 1980. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/article/10.1007/BF02462363.

Hart:1980:TAA

[443] Hiram E. Hart and Ki-Chuen Chak. Theory of antigen-antibody induced particulate aggreagation — I. General assumptions and basic theory. Bulletin of Mathematical Biology, 42(1):17–36, January 1980. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02462364.

Chak:1980:TAA

[444] Ki-Chuen Chak and Hiram Hart. Theory of antigen-antibody induced particulate aggregation — II theoretical analysis and comparison with experimental results. Bulletin of Mathematical Biology, 42 (1):37–56, January 1980. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462365.

Smith:1980:HRP

[445] William R. Smith. Hypothalamic regulation of pituitary secretion of luteinizing hormone — II feedback control of gonadotropin secretion. Bulletin of Mathematical Biology, 42(1):57–78, January 1980. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02462366.

Wei:1980:DTI

[446] Ling Y. Wei. Dipole theory of interactions of nerve signals. Bulletin of Mathematical Biology, 42(1):79–94, January 1980. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02462367.

Rosen:1980:FCP

[447] Gerald Rosen. On the Fisher and the cubic-polynomial equations for the propagation of species properties. Bulletin of Mathematical Biology, 42 (1):95–106, January 1980. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462368.

Comorosan:1980:NSB

[448] S. Comorosan, M. Hristea, and P. Murgoci. On a new symmetry in biological systems. Bulletin of Mathematical Biology, 42(1):107– 117, January 1980. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462369.

Hirata:1980:MHE

[449] Hironori Hirata. A model of hierarchical ecosystems with migration. Bulletin of Mathematical Biology, 42(1):119–130, January 1980. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02462370.

Anderson:1980:CVP

[450] N. Anderson and A. M. Arthurs. Complementary variational principles for diffusion problems with Michaelis–Menten kinetics. Bulletin of Mathematical Biology, 42(1):131–135, January 1980. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02462371.

Baianu:1980:SOP

[451] Ion C. Baianu. Structural order and partial disorder in biological systems. Bulletin of Mathematical Biology, 42(1):137–141, January 1980. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02462372.

Anonymous:1980:A

[452] Anonymous. Announcements. Bulletin of Mathematical Biology, 42(1): 143–145, January 1980. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462373.

Thakur:1980:FAS

[453] Ajit K. Thakur and Charles DeLisi. Fluctuation analysis in small chemically reacting systems. Bulletin of Mathematical Biology, 42(2): 147–160, March 1980. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02464634.

Leguizamon:1980:CEB

[454] Carlos A. Leguizamón and Juan C. Giménez. Concept of energy in biological systems and the effects of irradiations of low energies on enzyme-substrate systems. Bulletin of Mathematical Biology, 42(2): 161–172, March 1980. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02464635.

Min:1980:FAT

[455] B. G. Min, W. Welkowitz, and S. Fich. Frequency analysis of timevarying elastance model of the left ventricle. Bulletin of Mathematical Biology, 42(2):173–180, March 1980. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02464636.

Keller:1980:PPT

[456] Edward L. Keller. Primitivity of the product of two Leslie matrices. Bulletin of Mathematical Biology, 42(2):181–189, March 1980. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02464637.

Bentz:1980:CBMa

[457] Joe Bentz and Shlomo Nir. Cation binding to membranes: Competition between mono-, di- and trivalent cations. Bulletin of Mathematical Biology, 42(2):191–220, March 1980. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02464638.

Argemi:1980:QAM

[458] José Argémi, Maurice Gola, and Hélène Chagneux. Qualitative analysis of a model generating long potential waves in ba-treated nerve cells. II. Complete system. Bulletin of Mathematical Biology, 42(2): 221–238, March 1980. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02464639.

Stevens:1980:CME

[459] Evelyn Stevens and Roberta F. Colman. Chemical modification of enzymes: Critical evaluation of the graphical correlation between residual enzyme activity and number of groups modified. Bulletin of Mathematical Biology, 42(2):239–255, March 1980. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02464640.

Thron:1980:NHT

[460] C. D. Thron. The nonexistence of “hermaphoroditic” tracer systems. Bulletin of Mathematical Biology, 42(2):257–265, March 1980. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02464641.

Witten:1980:NSS

[461] Matthew Witten. A note on the structure of system state spaces and its implications on the existence of non-repeatable experiments. Bulletin of Mathematical Biology, 42(2):267–272, March 1980. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02464642.

Gray:1980:HGT

[462] B. F. Gray. Heat generation in tissues. Bulletin of Mathematical Biology, 42(2):273–274, March 1980. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02464643.

Zhitomirskii:1980:TDM

[463] I. Zhitomirskii and A. Kagna. Temperature distribution in muscle. Bulletin of Mathematical Biology, 42(2):275, March 1980. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02464644. See remark [595]

and replies [596, 663].

Rosen:1980:BR

[464] Robert Rosen. Book reviews. Bulletin of Mathematical Biology, 42(2): 277–281, March 1980. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02464645.

Beretta:1980:SRA

[465] E. Beretta. Some results about nonlinear chemical systems represented by trees and cycles. Bulletin of Mathematical Biology, 42 (2):282, March 1980. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02464646;   http://link.springer.com/content/pdf/10.
1007/BF02464646.pdf.

Shukla:1980:ESN

[466] J. B. Shukla, R. S. Parihar, and B. R. P. Rao. Effects of stenosis on non–Newtonian flow of the blood in an artery. Bulletin of Mathematical Biology, 42(3):283–294, May 1980. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460787.

Garner:1980:OTF

[467] J. B. Garner and R. B. Kellogg. A one tube flow problem arising in physiology. Bulletin of Mathematical Biology, 42(3):295–304, May 1980. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02460788.

Shen:1980:AMP

[468] M. C. Shen, S. M. Shim, and A. M. Wu. Asymptotic method for peristaltic transport. Bulletin of Mathematical Biology, 42(3): 305–325, May 1980. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460789.

Lumsden:1980:HCB

[469] Charles J. Lumsden and L. E. H. Trainor. On the Hamiltonian content of biodynamics. Bulletin of Mathematical Biology, 42(3): 327–337, May 1980. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460790.

Amari:1980:TON

[470] Shun-Ichi Amari. Topographic organization of nerve fields. Bulletin of Mathematical Biology, 42(3):339–364, May 1980. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/article/10.1007/BF02460791.

Grossberg:1980:IMA

[471] Stephen Grossberg. Intracellular mechanisms of adaptation and selfregulation in self-organizing networks: The role of chemical transducers. Bulletin of Mathematical Biology, 42(3):365–396, May 1980. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02460792.

Kennedy:1980:TWS

[472] C. R. Kennedy and R. Aris. Traveling waves in a simple population model involving growth and death. Bulletin of Mathematical Biology, 42(3):397–429, May 1980. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460793.

Baianu:1980:NTO

[473] Ion C. Baianu. Natural transformations of organismic structures. Bulletin of Mathematical Biology, 42(3):431–446, May 1980. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02460794.

Scott:1980:MED

[474] B. R. Scott. A model for early death caused by radiation pneumonitis and pulmonary fibrosis after inhaling insoluble radioactive particles. Bulletin of Mathematical Biology, 42(3):447–459, May 1980. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02460795.

Yeh:1980:MHL

[475] Hsu-Chi Yeh and G. M. Schum. Models of human lung airways and their application to inhaled particle deposition. Bulletin of Mathematical Biology, 42(3):461–480, May 1980. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460796.

Hearon:1980:CSC

[476] J. Z. Hearon. On the completely symmetric compartmental system. Bulletin of Mathematical Biology, 42(3):481–488, May 1980. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02460797.

Lee:1980:DFF

[477] Chyuan-Yih Lee. The dipole flip-flop theory for the excitation and conduction of excitable tissues — I. Dipole’s behavior and gating currents. Bulletin of Mathematical Biology, 42(4):489–505, July 1980. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02460966. See reply [537].

Witten:1980:SGC

[478] Matthew Witten. Some generalized conjugacy theorems and the concepts of fitness and survival in logistic growth models. Bulletin of Mathematical Biology, 42(4):507–528, July 1980. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02460967.

Patlak:1980:UTS

[479] C. S. Patlak, K. D. Pettigrew, and S. I. Rapoport. Use of transient and steady-state measurements of the unidirectional flux ratio for the determination of the free energy change of chemical reactions and active transport systems. Bulletin of Mathematical Biology, 42 (4):529–537, July 1980. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460968.

Jones:1980:EPE

[480] B. L. Jones. On the evolution of prebiotic enzyme coupled macromolecular systems. Bulletin of Mathematical Biology, 42(4):539–549, July 1980. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02460969.

Pounder:1980:GCD

[481] J. R. Pounder and Thomas D. Rogers. The geometry of chaos: Dynamics of a nonlinear second-order difference equation. Bulletin of Mathematical Biology, 42(4):551–597, July 1980. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02460970.

Willson:1980:VPP

[482] A. J. Willson. On Varma’s prey–predator problem. Bulletin of Mathematical Biology, 42(4):599–600, July 1980. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02460971.

Lin:1980:TCP

[483] Gregory S. B. Lin. Theory of cation-phospholipid-induced shape changes in a lipid bilayer couple. Bulletin of Mathematical Biology, 42(5):601– 625, September 1980. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460983.

Khanin:1980:MMF

[484] M. A. Khanin and I. B. Bukharov. A mathematical model of the functional state of the oxygen transport system. Bulletin of Mathematical Biology, 42(5):627–645, September 1980. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460984.

Deakin:1980:ACT

[485] M. A. B. Deakin. Applied catastrophe theory in the social and biological sciences. Bulletin of Mathematical Biology, 42(5):647–679, September 1980. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/article/10.

1007/BF02460985. See remark [573].

Ishay:1980:PPY

[486] Jacob S. Ishay, Bice Perna, Yosef Hochberg, and Mina Goldstein. Photoelectric properties of the “yellow strips” in Vespa orientalis; a mathematical model. Bulletin of Mathematical Biology, 42(5):681– 689, September 1980. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460986.

Bucolo:1980:SSA

[487] Richard J. Bucolo and Vijai K. Tripathi. Steady-state analysis of a two-compartment barrier-limited capillary-tissue model with Michaelis– Menten saturation kinetics. Bulletin of Mathematical Biology, 42(5): 691–700, September 1980. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460987.

Bojadziev:1980:DOP

[488] G. N. Bojadziev. Damped oscillating processes in biological and biochemical systems. Bulletin of Mathematical Biology, 42(5):701– 718, September 1980. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460988.

Cull:1980:PTU

[489] Paul Cull. The problem of time unit in Leslie’s population model. Bulletin of Mathematical Biology, 42(5):719–728, September 1980. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02460989.

Gopalsamy:1980:TLG

[490] K. Gopalsamy. Time lags and global stability in two-species competition. Bulletin of Mathematical Biology, 42(5):729–737, September 1980. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02460990.

Morris:1980:NDH

[491] Max D. Morris and C. Alex McMahan. A note on the dominance hierarchy index. Bulletin of Mathematical Biology, 42(5):739– 746, September 1980. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460991.

Rosen:1980:FWN

[492] Robert Rosen. Food webs and niche spaces. Bulletin of Mathematical Biology, 42(5):747–748, September 1980. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460992.

Bentz:1980:CBMb

[493] J. Bentz and S. Nir. Cation binding to membranes; competition between mono-, di-and trivalent cations. Bulletin of Mathematical Biology, 42 (5):749, September 1980. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460993;   http://link.springer.com/content/pdf/10.
1007/BF02460993.pdf.

McIlroy:1980:TID

[494] Douglas K. McIlroy. On the theory of interfacial double-layer potential differences and periaxonal accumulation of potassium in squid giant axons. Bulletin of Mathematical Biology, 42(6):751–763, November 1980. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02461057.

Roselli:1980:ERC

[495] R. J. Roselli. Effects of red cell permeability on transcapillary tracer transport: The case of negligible back diffusion. Bulletin of Mathematical Biology, 42(6):765–795, November 1980. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02461058.

Shukla:1980:EPL

[496] J. B. Shukla, R. S. Parihar, and S. P. Gupta. Effects of peripheral layer viscosity on blood flow through the artery with mild stenosis. Bulletin of Mathematical Biology, 42(6):797–805, November 1980. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02461059.

Mirsky:1980:GMM

[497] I. Mirsky and M. M. Laks. A geometric model for the myocardium: Biventricular wall stresses in normal and hypertrophied states. Bulletin of Mathematical Biology, 42(6):807–828, November 1980. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02461060.

Bhargava:1980:HMG

[498] S. C. Bhargava. On the Higgins model of glycolysis. Bulletin of Mathematical Biology, 42(6):829–836, November 1980. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/article/10.1007/BF02461061.

Larsen:1980:MMO

[499] Jesper Larsen. A mathematical model of the oscillations in the aphakic eye. Bulletin of Mathematical Biology, 42(6):837–845, November 1980. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02461062.

Chen:1980:TDM

[500] Wen-Jia Russell Chen, David Shii-Pyng Shiah, and C. S. Wang. A three-dimensional model of the upper tracheobronchial tree. Bulletin of Mathematical Biology, 42(6):847–859, November 1980. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02461063.

Liniger:1980:MMA

[501] W. Liniger, G. Karreman, R. Rawala, and R. Colman. Mathematical model of the activation of prothrombin by factor X_{a} and factor V_{t}. Bulletin of Mathematical Biology, 42(6):861–870, November 1980. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02461064.

Gopalsamy:1980:PEW

[502] K. Gopalsamy. Pursuit-evasion wave trains in prey–predator systems with diffusionally coupled delays. Bulletin of Mathematical Biology, 42(6):871–887, November 1980. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02461065.

Rosen:1980:COC

[503] Robert Rosen. On control and optimal control in biodynamic systems. Bulletin of Mathematical Biology, 42(6):889–897, November 1980. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02461066.

Rosen:1980:MMH

[504] Robert Rosen. Mathematical models in the health sciences: a computer-aided approach. Bulletin of Mathematical Biology, 42(6):899– 900, November 1980. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02461067.

Anonymous:1980:ANB

[505] Anonymous. The 9th Annual Northeast Biomedical Engineering Conference. Bulletin of Mathematical Biology, 42(6):901, November 1980. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02461068.

Saidel:1981:MTS

[506] Gerald M. Saidel and Gerald M. Burma. Multibreath tracer species dynamics in the lung. Bulletin of Mathematical Biology, 43(1):1– 19, January 1981. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460935.

Singh:1981:NBO

[507] Virendra Singh. Nonlinear biological oscillators: Action-angle variable approach with an application to Cowan’s model of neuroelectric activity. Bulletin of Mathematical Biology, 43(1):21–32, January 1981. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02460936.

Bernard:1981:TUM

[508] S. R. Bernard, Milton Sobel, and V. R. R. Uppuluri. On a two urn model of Pólya-type. Bulletin of Mathematical Biology, 43(1): 33–45, January 1981. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460937.

Cull:1981:GSP

[509] Paul Cull. Global stability of population models. Bulletin of Mathematical Biology, 43(1):47–58, January 1981. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02460938.

Conrad:1981:LCP

[510] Michael Conrad and Arnon Rosenthal. Limits on the computing power of biological systems. Bulletin of Mathematical Biology, 43(1): 59–67, January 1981. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460939.

Roerdink:1981:GFP

[511] J. Roerdink and A. Weyland. A generalized Fokker–Planck equation in the case of the Volterra model. Bulletin of Mathematical Biology, 43 (1):69–79, January 1981. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460940.

Kovacs:1981:TSB

[512] A. L. Kovacs, C. Pacitti, F. Conti, and A. L. Segre. A topological study of the base sequences of MS2 and φ x174 genoma. Bulletin of Mathematical Biology, 43(1):81–88, January 1981. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460941.

Hardt:1981:DTT

[513] Shoshana L. Hardt. The diffusion transit time; a simple derivation. Bulletin of Mathematical Biology, 43(1):89–99, January 1981. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02460942.

Horsfield:1981:VCA

[514] Keith Horsfield and Alison Thurlbeck. Volume of the conducting airways calculated from morphometric parameters. Bulletin of Mathematical Biology, 43(1):101–109, January 1981. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02460943.

Purdue:1981:VSC

[515] Peter Purdue. Variability in a single compartment system: a note on S. R. Bernard’s model. Bulletin of Mathematical Biology, 43(1): 111–116, January 1981. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460944.

McElwain:1981:CLP

[516] D. L. S. McElwain. A comment on Lin’s paper on nonlinear diffusion in biological systems. Bulletin of Mathematical Biology, 43(1): 117–120, January 1981. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460945.

Rosen:1981:NOB

[517] Robert Rosen. Nonlinear oscillations in biology. Bulletin of Mathematical Biology, 43(1):121–123, January 1981. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460946.

Gopalsamy:1981:LED

[518] K. Gopalsamy and B. D. Aggarwala. The logistic equation with a diffusionally coupled delay. Bulletin of Mathematical Biology, 43(2): 125–140, March 1981. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459438.

Norwich:1981:UPP

[519] Kenneth H. Norwich. Uncertainty in physiology and physics. Bulletin of Mathematical Biology, 43(2):141–149, March 1981. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02459439.

Radhakrishnamacharya:1981:HSF

[520] G. Radhakrishnamacharya, Peeyush Chandra, and M. R. Kaimal. A hydrodynamical study of the flow in renal tubules. Bulletin of Mathematical Biology, 43(2):151–163, March 1981. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02459440.

Stephanopoulos:1981:EPM

[521] Gregory Stephanopoulos and A. G. Fredrickson. Extinction probabilities in microbial predation: a birth and death approach. Bulletin of Mathematical Biology, 43(2):165–181, March 1981. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02459441.

Wong:1981:SCC

[522] Alan Y. K. Wong. Some comments on the conductance of slow inward current in cardiac muscle. Bulletin of Mathematical Biology, 43 (2):183–199, March 1981. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459442.

Agrafiotis:1981:STC

[523] G. K. Agrafiotis. On the stochastic theory of compartments: The leaving process of the two-compartment systems. Bulletin of Mathematical Biology, 43(2):201–211, March 1981. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02459443.

Venieratos:1981:MPP

[524] Dionyssis Venieratos. Models on prebiotic polymer competition: a deterministic approach. Bulletin of Mathematical Biology, 43(2):213– 232, March 1981. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459444.

Moseley:1981:PEC

[525] H. Moseley and R. Strang. Prediction of the effect of the choroidal circulation on the retinal temperature during illumination. Bulletin of Mathematical Biology, 43(2):233–238, March 1981. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02459445.

Margush:1981:CT

[526] T. Margush and F. R. McMorris. Consensus n-trees. Bulletin of Mathematical Biology, 43(2):239–244, March 1981. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02459446.

Rescigno:1981:RSC

[527] Aldo Rescigno and James H. Matis. On the relevance of stochastic compartmental models to pharmacokinetic systems. Bulletin of Mathematical Biology, 43(2):245–247, March 1981. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02459447.

Kemper:1981:ISI

[528] John T. Kemper. Identification of silent infections in SIR epidemics. Bulletin of Mathematical Biology, 43(3):249–257, May 1981. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02462199.

Marek:1981:MPF

[529] M. Marek and M. Kubíĉek. Morphogen pattern formation and development in growth. Bulletin of Mathematical Biology, 43(3):259– 270, May 1981. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462200.

Boffey:1981:SEM

[530] T. B. Boffey and D. F. Yates. Simulation experiments with a model for cellular growth. Bulletin of Mathematical Biology, 43(3): 271–278, May 1981. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462201.

Stewart:1981:ZEN

[531] Ian Stewart and A. E. R. Woodcock. On Zeeman’s equations for the nerve impulse. Bulletin of Mathematical Biology, 43(3):279–325, May 1981. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02462202.

Shenton:1981:RDU

[532] L. R. Shenton. A reinforcement-depletion urn problem — I. Basic theory. Bulletin of Mathematical Biology, 43(3):327–340, May 1981. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02462203.

Anderson:1981:CEP

[533] N. Anderson and A. M. Arthurs. Complementary extremum principles for a nonlinear model of heat conduction in the human head. Bulletin of Mathematical Biology, 43(3):341–346, May 1981. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/article/10.1007/BF02462204.

Parthasarathy:1981:SCM

[534] P. R. Parthasarathy and P. Mayilswami. Stochastic compartmental model with branching particles. Bulletin of Mathematical Biology, 43 (3):347–360, May 1981. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462205.

Ignetik:1981:PPA

[535] Rainer Ignetik, Michael A. B. Deakin, and Chris Fandry. Phase-plane analyses of the Michaelis–Menten reaction equations. Bulletin of Mathematical Biology, 43(3):361–370, May 1981. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02462206.

Lee:1981:WFE

[536] C.-Y. Lee. On Wei’s fundamental equation for nerve excitation. Bulletin of Mathematical Biology, 43(3):371–372, May 1981. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02462207.

Wei:1981:RCY

[537] L. Y. Wei. Reply to Chyuan-Yih Lee. Bulletin of Mathematical Biology, 43(3):372–373, May 1981. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462208. See [477].

Ignetik:1981:AAM

[538] Rainer Ignetik and Michael A. B. Deakin. Asymptotic analysis of the Michaelis–Menten reaction equations. Bulletin of Mathematical Biology, 43(4):375–388, July 1981. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459428.

Eilbeck:1981:PEC

[539] J. C. Eilbeck, S. D. Luzader, and A. C. Scott. Pulse evolution on coupled nerve fibres. Bulletin of Mathematical Biology, 43(4):389–400, July 1981. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02459429.

Joines:1981:FIF

[540] William T. Joines and Blake S. Wilson. Field-induced forces at dielectric interfaces as a possible mechanism of RF hearing effects. Bulletin of Mathematical Biology, 43(4):401–413, July 1981. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/article/10.1007/BF02459430.

Muir:1981:CHN

[541] A. Muir. The construction of homogeneous nets. Bulletin of Mathematical Biology, 43(4):415–426, July 1981. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02459431.

Brannan:1981:SLI

[542] James R. Brannan and William E. Boyce. Spatially localized interactive neural populations — I. A mathematical model. Bulletin of Mathematical Biology, 43(4):427–446, July 1981. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02459432.

Shoucri:1981:CAK

[543] Rachad M. Shoucri. Contractile analysis and kinetic energy of the myocardium. Bulletin of Mathematical Biology, 43(4):447–461, July 1981. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02459433.

Gopalsamy:1981:LCP

[544] K. Gopalsamy. Limit cycles in periodically perturbed population systems. Bulletin of Mathematical Biology, 43(4):463–485, July 1981. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02459434.

Scott:1981:DRM

[545] B. R. Scott. A dose-response model for estimating lifetime tumor risks when cell killing occurs. Bulletin of Mathematical Biology, 43 (4):487–501, July 1981. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459435.

Karmeshu:1981:OCM

[546] Karmeshu and C. K. Gupta. A one-compartment model with stochastic parameter. Bulletin of Mathematical Biology, 43(4):503–512, July 1981.

CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02459436.

Rosen:1981:UAP

[547] Robert Rosen. On a unified approach to pattern generation. Bulletin of Mathematical Biology, 43(4):513–516, July 1981. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/article/10.1007/BF02459437.

Davidson:1981:FCT

[548] Malcolm M. Davidson. Further considerations in a theoretical description of gas transport in lung airways. Bulletin of Mathematical Biology, 43(5): 517–548, September 1981. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462311.

Mehata:1981:SMC

[549] K. M. Mehata and D. Deivamoney Selvam. A stochastic model for the n-compartment irreversible system. Bulletin of Mathematical Biology, 43(5):549–561, September 1981. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02462312.

Gopalan:1981:PBF

[550] N. P. Gopalan. Pulsatile blood flow in a rigid pulmonary alveolar sheet with porous walls. Bulletin of Mathematical Biology, 43(5):563– 577, September 1981. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462313.

MacFarlane:1981:DST

[551] A. I. MacFarlane. Dynamic structure theory: a structural approach to social and biological systems. Bulletin of Mathematical Biology, 43(5): 579–591, September 1981. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462314.

Shukla:1981:ECD

[552] J. B. Shukla and Sunita Verma. Effects of convective and dispersive interactions on the stability of two species. Bulletin of Mathematical Biology, 43(5):593–610, September 1981. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02462315.

Toth:1981:RHD

[553] T. Tóth. Role of the hypergeometric distribution in the transmission of excitation through sympathetic ganglia. Bulletin of Mathematical Biology, 43(5):611–618, September 1981. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02462316.

Saroff:1981:UPS

[554] H. A. Saroff and F. A. Kutyna. The uniqueness of protein sequences: a Monte Carlo analysis. Bulletin of Mathematical Biology, 43(6):619– 639, November 1981. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458413.

Warner:1981:TN

[555] M. W. Warner. Tolerance nets. Bulletin of Mathematical Biology, 43(6): 641–650, November 1981. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458414.

Matis:1981:CMM

[556] James H. Matis and Thomas E. Wehrly. Compartmental models with multiple sources of stochastic variability: The one-compartment models with clustering. Bulletin of Mathematical Biology, 43(6):651– 664, November 1981. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458415.

Jones:1981:SAN

[557] B. L. Jones and H. K. Leung. Stochastic analysis of a nonlinear model for selection of biological macromolecules. Bulletin of Mathematical Biology, 43(6):665–680, November 1981. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02458416.

Horsfield:1981:RBD

[558] Keith Horsfield and Alison Thurlbeck. Relation between diameter and flow in branches of the bronchial tree. Bulletin of Mathematical Biology, 43(6):681–691, November 1981. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02458417.

Thron:1981:PDLa

[559] C. D. Thron. Peak drug levels in linear pharmacokinetic systems — I. Effect of rate of injection. Bulletin of Mathematical Biology, 43(6): 693–703, November 1981. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458418.

Thron:1981:PDLb

[560] C. D. Thron. Peak drug levels in linear pharmacokinetic systems — II. Conditions for a paradoxical injection rate effect with rectangular input. Bulletin of Mathematical Biology, 43(6):705–715, November 1981. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02458419.

Cohen:1981:STI

[561] Joel E. Cohen. Some trees are not interval graphs. Bulletin of Mathematical Biology, 43(6):717, November 1981. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02458420.

Anonymous:1981:AMC

[562] Anonymous. Annual meeting of the cell kinetics society. Bulletin of Mathematical Biology, 43(6):719, November 1981. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02458421.

Kimmel:1982:ERI

[563] Marek Kimmel. An equivalence result for integral equations with application to branching processes. Bulletin of Mathematical Biology, 44 (1):1–15, January 1982. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459415.

Inagaki:1982:SUR

[564] H. Inagaki. Selection under random mutations in stochastic Eigen model. Bulletin of Mathematical Biology, 44(1):17–28, January 1982. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02459416.

Melbin:1982:GSF

[565] Julius Melbin and Abraham Noordergraaf. General solution for flow and stress phenomena in elastic vessels (Navier–Stokes). Bulletin of Mathematical Biology, 44(1):29–42, January 1982. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02459417.

Brannan:1982:SLI

[566] James R. Brannan. Spatially localized interactive neural populations — II: Stability and dynamics of excitatory sets. Bulletin of Mathematical Biology, 44(1):43–56, January 1982. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02459418.

Linkens:1982:MAP

[567] D. A. Linkens and R. I. Kitney. Mode analysis of physiological oscillators intercoupled via pure time delays. Bulletin of Mathematical Biology, 44 (1):57–74, January 1982. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459419.

Parthasarathy:1982:SCM

[568] P. R. Parthasarathy and P. Mayilswami. Stochastic compartmental models with branching and immigrant particles. Bulletin of Mathematical Biology, 44(1):75–86, January 1982. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02459420.

Brown:1982:BSD

[569] K. J. Brown and J. C. Eilbeck. Bifurcation, stability diagrams, and varying diffusion coefficients in reaction–diffusion equations. Bulletin of Mathematical Biology, 44(1):87–102, January 1982. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02459421.

Granero-Porati:1982:RES

[570] M. I. Granero-Porati, R. Kron-Morelli, and A. Porati. Random ecological systems with structure: Stability-complexity relationship. Bulletin of Mathematical Biology, 44(1):103–117, January 1982. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02459422.

Weingarden:1982:TEI

[571] M. Weingarden, H. Mizukami, and Stephan A. Rice. Transient effects on the initial rate of oxygenation of red blood cells. Bulletin of Mathematical Biology, 44(1):119–134, January 1982. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02459423.

Weingarden:1982:FDR

[572] M. Weingarden, H. Mizukami, and S. A. Rice. Factors defining the rate of oxygen uptake by the red blood cell. Bulletin of Mathematical Biology, 44 (1):135–147, January 1982. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459424.

Cronin:1982:RAA

[573] Jane Cronin. Remark on the article, “Applied catastrophe theory in the social and biological sciences” by M. A. B. Deakin (1980). Bulletin of Mathematical Biology, 44(1):149–150, January 1982. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02459425. See [485].

Bernard:1982:BR

[574] S. R. Bernard. Book review. Bulletin of Mathematical Biology, 44 (1):151, January 1982. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459426.

Toth:1982:RHD

[575] T. Tóth. Role of the hypergeometric distribution in the transmission of excitation through sympathetic ganglia. Bulletin of Mathematical Biology, 44(1):152, January 1982. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02459427;   http://link.springer.com/
content/pdf/10.1007/BF02459427.pdf.

Siddique:1982:VCA

[576] Mohsin R. Siddique, T. Leo van Winkle, David M. Wetzel, and Marshall M. Lih. A variational calculus approach to the suspension flow problem using the minimal force hypothesis. Bulletin of Mathematical Biology, 44(2):153–174, March 1982. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02463245.

Demiray:1982:CFD

[577] Hilmi Demiray and Mark Levinson. On a class of finite deformations of elastic soft tissues. Bulletin of Mathematical Biology, 44(2): 175–192, March 1982. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02463246.

Dujardin:1982:LSL

[578] Jean-Pierre Dujardin and Boudewijn Van Gelder. Large signal, low frequency characteristics of the systemic arterial system. Bulletin of Mathematical Biology, 44(2):193–213, March 1982. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02463247.

Seto:1982:EMP

[579] Y. J. Seto and S. T. Hsieh. Electromagnetic mediated pharmacokinetics in a three-layer diffusional system. Bulletin of Mathematical Biology, 44 (2):215–229, March 1982. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02463248.

Gola:1982:AKM

[580] Maurice Gola, Hélène Chagneux, and José Argémi. An asymmetrical kinetic model for veratridine interactions with sodium channels in molluscan neurons. Bulletin of Mathematical Biology, 44(2):231– 258, March 1982. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02463249.

Shonkwiler:1982:CSEa

[581] Ronald Shonkwiler and Maynard Thompson. Common source epidemics. I: a stochastic model. Bulletin of Mathematical Biology, 44 (2):259–269, March 1982. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02463250.

Nallaswamy:1982:ECD

[582] R. Nallaswamy and J. B. Shukla. Effects of convective and dispersive migration on the linear stability of a two species system with mutualistic interactions and functional response. Bulletin of Mathematical Biology, 44(2):271–282, March 1982. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02463251.

vanRotterdam:1982:MST

[583] A. van Rotterdam, F. H. Lopes da Silva, J. van den Ende, M. A. Viergever, and A. J. Hermans. A model of the spatial-temporal characteristics of the alpha rhythm. Bulletin of Mathematical Biology, 44 (2):283–305, March 1982. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02463252.

Anonymous:1982:Aa

[584] Anonymous. Announcements. Bulletin of Mathematical Biology, 44 (2):307, March 1982. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02463253.

Gelfand:1982:BSC

[585] Alan E. Gelfand. A behavioral summary for completely random nets. Bulletin of Mathematical Biology, 44(3):309–320, May 1982. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02462282.

Gray:1982:ENW

[586] B. F. Gray and M. E. Sherrington. Eigensolutions of nonlinear wave equations in one dimension. Bulletin of Mathematical Biology, 44(3): 321–337, May 1982. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462283.

Voorhees:1982:DSA

[587] B. H. Voorhees. Dissipative structures associated to certain reaction– diffusion equations in mathematical ecology. Bulletin of Mathematical Biology, 44(3):339–348, May 1982. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02462284.

Roy:1982:GOM

[588] André G. Roy and Michael J. Woldenberg. A generalization of the optimal models of arterial branching. Bulletin of Mathematical Biology, 44(3):349–360, May 1982. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462285.

Taylor:1982:FMM

[589] R. Taylor and R. Krishnamurthy. Film models for multicomponent mass transfer: Diffusion in physiological gas mixtures. Bulletin of Mathematical Biology, 44(3):361–376, May 1982. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02462286.

Shonkwiler:1982:CSEb

[590] Ronald Shonkwiler and Maynard Thompson. Common source epidemics II: Toxoplasmosis in Atlanta. Bulletin of Mathematical Biology, 44 (3):377–398, May 1982. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462287.

Landahl:1982:CMI

[591] Herbert D. Landahl and Gerold M. Grodsky. Comparison of models of insulin release. Bulletin of Mathematical Biology, 44(3):399–409, May 1982. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02462288.

vanDrongelen:1982:PEO

[592] Wim van Drongelen, Yves Pagnotte, and Mat H. Hendriks. Primary events in odour detection. Bulletin of Mathematical Biology, 44(3): 411–423, May 1982. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462289.

Martonen:1982:AMH

[593] T. B. Martonen. Analytical model of hygroscopic particle behavior in human airways. Bulletin of Mathematical Biology, 44(3): 425–442, May 1982. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462290.

Conrad:1982:ESW

[594] Michael Conrad and Otto Rössler. Example of a system which is computation universal but not effectively programmable. Bulletin of Mathematical Biology, 44(3):443–447, May 1982. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02462291.

Cronin:1982:RLZ

[595] Jane Cronin. Remark on the letter by Zhitomirsky and Kagna (1980). Bulletin of Mathematical Biology, 44(3):449, May 1982. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02462292. See [463] and

replies [596, 663].

Zhitomirsky:1982:RLJ

[596] I. S. Zhitomirsky and A. A. Kagna. Reply to letter by Jane Cronin (1982). Bulletin of Mathematical Biology, 44(3):450, May 1982. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02462293. See [595, 663].

Anonymous:1982:Ab

[597] Anonymous. Announcements. Bulletin of Mathematical Biology, 44 (3):451–452, May 1982. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462294.

Witten:1982:KMA

[598] Mathew Witten and David Siegel. A kinetics model of abrin binding in a virus transformed lymphocyte cell culture. Bulletin of Mathematical Biology, 44(4):453–476, July 1982. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02459403.

Rescigno:1982:DDB

[599] Aldo Rescigno, James S. Beck, and H. Joseph Goren. Determination of dependence of binding parameters on receptor occupancy. Bulletin of Mathematical Biology, 44(4):477–489, July 1982. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/article/10.1007/BF02459404. See correction [696].

Bouffette:1982:MMS

[600] A. R. Bouffette and J. Bouffette. Mathematical model for the study of fertility and viability in inbreeding populations. Bulletin of Mathematical Biology, 44(4):491–500, July 1982. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02459405.

Tanyi:1982:EBEa

[601] G. E. Tanyi. Energy and biological evolution — I. The equilibrium states of biochemical processes. Bulletin of Mathematical Biology, 44 (4):501–535, July 1982. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459406.

Tanyi:1982:EBEb

[602] G. E. Tanyi. Energy and biological evolution — II. The mathematical structure of equilibrium states. Bulletin of Mathematical Biology, 44 (4):537–547, July 1982. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459407.

Tanyi:1982:EBEc

[603] G. E. Tanyi. Energy and biological evolution — III. Theoretical ecology and macromolecular self-organization. Bulletin of Mathematical Biology, 44(4):549–555, July 1982. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459408.

Mehata:1982:SMT

[604] K. M. Mehata and D. Deivamoney Selvam. A stochstic model for a two-compartment reversible system. Bulletin of Mathematical Biology, 44(4):557–570, July 1982. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459409.

Shukla:1982:EDS

[605] V. P. Shukla and P. C. Das. Effects of dispersion on stability of multispecies prey–predator systems. Bulletin of Mathematical Biology, 44 (4):571–578, July 1982. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459410.

Solimano:1982:GTC

[606] F. Solimano and E. Beretta. Graph theoretical criteria for stability and boundedness of predator–prey systems. Bulletin of Mathematical Biology, 44(4):579–585, July 1982. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459411.

Kapadia:1982:DST

[607] Asha Seth Kapadia and Bayliss C. McInnis. Distribution of survival times in the fix and Neyman model. Bulletin of Mathematical Biology, 44(4):587–591, July 1982. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459412.

Campbell:1982:RE

[608] R. B. Campbell. Repeatability of experiments. Bulletin of Mathematical Biology, 44(4):593, July 1982. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02459413. See reply [609].

Witten:1982:RRE

[609] Matthew Witten. A reply to “Repeatability of experiments” by R. B. Campbell. Bulletin of Mathematical Biology, 44(4):594–595, July 1982. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02459414. See

[608].

Zamir:1982:LGA

[610] M. Zamir. Local geometry of arterial branching. Bulletin of Mathematical Biology, 44(5):597–607, September 1982. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02462270.

Thron:1982:PDL

[611] C. D. Thron. Peak drug levels in linear pharmacokinetic systems — III. Multimodal impulse responses in multicompartment systems. Bulletin of Mathematical Biology, 44(5):609–635, September 1982. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02462271.

Stetson:1982:SCO

[612] R. F. Stetson and W. A. Hogan. Simulation of chemical oscillations in membranes. Bulletin of Mathematical Biology, 44(5):637– 645, September 1982. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462272.

Mode:1982:EPM

[613] Charles J. Mode and Robert C. Busby. An eight-parameter model of human mortality — the single decrement case. Bulletin of Mathematical Biology, 44(5):647–659, September 1982. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02462273.

Warner:1982:RSC

[614] M. W. Warner. Representations of (M, R)-systems by categories of automata. Bulletin of Mathematical Biology, 44(5):661–668, September 1982. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462274.

Marshall:1982:EEE

[615] E. A. Marshall. Equilibrium electropotentials in electrolyte systems. Bulletin of Mathematical Biology, 44(5):669–687, September 1982. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02462275.

Witten:1982:STQ

[616] Matthew Witten. Some thoughts on quantum non-demolition measurements in biological systems. Bulletin of Mathematical Biology, 44(5): 689–696, September 1982. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462276.

Hirata:1982:EPP

[617] Hironori Hirata. Evolution of a predator–prey Volterra–Lotka ecosystem with saturation effect. Bulletin of Mathematical Biology, 44(5): 697–704, September 1982. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462277.

Bhaumik:1982:SWM

[618] Debajyoti Bhaumik, Binayak Dutta-Roy, and Avijit Lahiri. Solitary waves and macromolecular systems. Bulletin of Mathematical Biology, 44(5):705–713, September 1982. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02462278.

Fogelman-Soulie:1982:SRD

[619] F. Fogelman-Soulie, E. Goles-Chacc, and G. Weisbuch. Specific roles of the different boolean mappings in random networks. Bulletin of Mathematical Biology, 44(5):715–730, September 1982. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/article/10.1007/BF02462279.

Brown:1982:EMI

[620] Grayson C. Brown and Gerald L. Nordin. An epizootic model of an insect–fungal pathogen system. Bulletin of Mathematical Biology, 44(5): 731–739, September 1982. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462280.

Schenzle:1982:FCG

[621] Dieter Schenzle. On feedback control of gonadotropin secretion. Bulletin of Mathematical Biology, 44(5):741–748, September 1982. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02462281.

Paine:1982:DLB

[622] Gregory H. Paine. The development of Lagrangians for biological models. Bulletin of Mathematical Biology, 44(6):749–760, November 1982. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02465178.

Richardson:1982:NGK

[623] I. W. Richardson, E. A. D. Foster, and S. Miekisz. Nonlinear generalizations of the Kedem–Katchalsky equations for ionic fluxes. Bulletin of Mathematical Biology, 44(6):761–775, November 1982. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02465179.

Krishnan:1982:BOC

[624] V. V. Krishnan, E. I. Jury, and L. Stark. Biological optimality: Comparison of generalized feasibility and optimality conditions for linear, error-minimizing systems. Bulletin of Mathematical Biology, 44(6):777– 791, November 1982. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02465180.

Swan:1982:OCM

[625] G. W. Swan. An optimal control model of Diabetes mellitus. Bulletin of Mathematical Biology, 44(6):793–808, November 1982. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02465181.

Agrafiotis:1982:STC

[626] G. K. Agrafiotis. On the stochastic theory of compartments: a semi-Markov approach for the analysis of the K-compartmental systems. Bulletin of Mathematical Biology, 44(6):809–817, November 1982. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02465182.

Shoucri:1982:SAC

[627] Rachad M. Shoucri, Jean G. Dumesnil, Jacques Rouleau, and Gilles R. Dagenais. Some applications of the concept of body forces for the myocardium. Bulletin of Mathematical Biology, 44(6):819–840, November 1982. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02465183.

Tosaka:1982:AND

[628] Nobuyoshi Tosaka and Shuhei Miyake. Analysis of a nonlinear diffusion problem with Michaelis–Menten kinetics by an integral equation method. Bulletin of Mathematical Biology, 44(6):841–849, November 1982. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02465184.

Salathe:1982:DAF

[629] Eric P. Salathé and Tseng-Chan Wang. The development of anoxia following occlusion. Bulletin of Mathematical Biology, 44(6):851– 877, November 1982. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02465185.

Andrietti:1982:DAI

[630] F. Andrietti and R. Pezzotta. Diffusion along intercellular spaces: an analysis of a transient situation. Bulletin of Mathematical Biology, 44(6): 879–891, November 1982. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02465186.

Burnside:1982:NES

[631] R. R. Burnside. A note on exact solutions of two prey–predator equations. Bulletin of Mathematical Biology, 44(6):893–897, November 1982. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02465187.

Anonymous:1982:Ac

[632] Anonymous. Announcements. Bulletin of Mathematical Biology, 44(6): 899–900, November 1982. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02465188.

Shenton:1983:RDU

[633] L. R. Shenton. A reinforcement depletion urn problem — II. Application and generalization. Bulletin of Mathematical Biology, 45(1): 1–9, January 1983. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459382.

Lefevre:1983:SEM

[634] Claude Lefevre. On a simple epidemic model with a heterogeneous infective population. Bulletin of Mathematical Biology, 45(1): 11–20, January 1983. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459383.

Torkington:1983:KDG

[635] P. Torkington. Kinetics of deterrence of Gompertzian growth. Bulletin of Mathematical Biology, 45(1):21–31, January 1983. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02459384.

Rohlf:1983:NBT

[636] F. James Rohlf. Numbering binary trees with labeled terminal vertices. Bulletin of Mathematical Biology, 45(1):33–40, January 1983. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02459385.

Perkkio:1983:HRB

[637] J. Perkkiö and R. Keskinen. Hematocrit reduction in bifurcations due to plasma skimming. Bulletin of Mathematical Biology, 45(1): 41–50, January 1983. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459386.

Mosqueira:1983:SCM

[638] F. G. Mosqueira. A simple competition model involving selection. Bulletin of Mathematical Biology, 45(1):51–67, January 1983. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02459387.

Duniec:1983:EPM

[639] Jacek T. Duniec and Sidney W. Thorne. Electrostatic potentials in membrane systems. Bulletin of Mathematical Biology, 45(1):69– 90, January 1983. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459388.

Bhaumik:1983:EFD

[640] Debajyoti Bhaumik, Binaryak Dutta-Roy, Tarun Kumar Chaki, and Avijit Lahiri. Electric field dependence of phase transitions in bilayer lipid membranes and possible biological implications. Bulletin of Mathematical Biology, 45(1):91–101, January 1983. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02459389.

Antonelli:1983:NLG

[641] P. L. Antonelli and B. H. Voorhees. Non-linear growth mechanics — I. Volterra–Hamilton systems. Bulletin of Mathematical Biology, 45(1): 103–116, January 1983. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459390.

Saroff:1983:UPS

[642] H. A. Saroff and H. T. Pretorius. The uniqueness of protein sequences. ouniqueness and infrequent peptides. Bulletin of Mathematical Biology, 45 (1):117–138, January 1983. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459391.

Brill:1983:WLP

[643] Michael H. Brill. Weber’s law and perceptual categories: Another teleological view. Bulletin of Mathematical Biology, 45(1):139– 142, January 1983. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459392.

Bessho:1983:NMG

[644] Chikafusa Bessho and Naoki Kuroda. A note on a more general solution of Eigen’s rate equation for selection. Bulletin of Mathematical Biology, 45(1):143–149, January 1983. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02459393.

Rosen:1983:TSC

[645] Gerald Rosen. Theoretical significance of the condition δ = 2μ in bacterial chemotaxis. Bulletin of Mathematical Biology, 45(2):151– 153, March 1983. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462353.

Arthurs:1983:PBS

[646] A. M. Arthurs and W. M. Arthurs. Pointwise bounds for the solution of a nonlinear problem in cell membrane theory. Bulletin of Mathematical Biology, 45(2):155–168, March 1983. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02462354.

Schilling:1983:MME

[647] R. J. Schilling. A mathematical model of the effects of light on the eclosion rhythm of the fruit fly Drosophila pseudoobscura. Bulletin of Mathematical Biology, 45(2):169–192, March 1983. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02462355.

Warner:1983:LLV

[648] M. W. Warner. Lattices and lattice-valued relations in biology. Bulletin of Mathematical Biology, 45(2):193–207, March 1983. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02462356.

Allen:1983:PES

[649] Linda J. S. Allen. Persistence and extinction in single-species reaction– diffusion models. Bulletin of Mathematical Biology, 45(2):209–227, March 1983. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462357.

Mann:1983:SML

[650] J. E. Mann. On stochastic models of light absorption. Bulletin of Mathematical Biology, 45(2):229–237, March 1983. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02462358.

Atherton:1983:ALC

[651] D. P. Atherton and D. A. Linkens. Analysis of limit-cycle conditions in intercoupled relay oscillators with reference to gastrointestinal modelling. Bulletin of Mathematical Biology, 45(2):239–257, March 1983. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02462359.

Perkko:1983:ECP

[652] J. Perkko and R. Keskinen. On the effect of the concentration profile of red cells on blood flow in the artery with stenosis. Bulletin of Mathematical Biology, 45(2):259–267, March 1983. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02462360.

vanPelt:1983:EPB

[653] J. van Pelt and R. W. H. Verwer. The exact probabilities of branching patterns under terminal and segmental growth hypotheses. Bulletin of Mathematical Biology, 45(2):269–285, March 1983. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02462361.

Lurie:1983:BDE

[654] David Lurié and Jorge Wagensberg. On biomass diversity in ecology. Bulletin of Mathematical Biology, 45(2):287–293, March 1983. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02462362.

Gopalsamy:1983:HDM

[655] K. Gopalsamy. Harmless delays in model systems. Bulletin of Mathematical Biology, 45(3):295–309, May 1983. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02459394.

Britton:1983:MHR

[656] N. F. Britton, G. Joly, and M.-C. Duban. A model for hydranth regeneration in tubularia. Bulletin of Mathematical Biology, 45(3): 311–321, May 1983. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459395.

Scott:1983:TME

[657] B. R. Scott. Theoretical models for estimating dose-effect relationships after combined exposure to cytotoxicants. Bulletin of Mathematical Biology, 45(3):323–345, May 1983. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02459396.

Kannan:1983:MCA

[658] D. Kannan. A Markov chain analysis of predator strategy in a model-mimic system. Bulletin of Mathematical Biology, 45(3):347– 400, May 1983. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459397.

Faddy:1983:CMR

[659] M. J. Faddy. Compartmental models with restricted movement. Bulletin of Mathematical Biology, 45(3):401–408, May 1983. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02459398.

Martonen:1983:FIP

[660] T. Martonen. On the fate of inhaled particles in the human: a comparison of experimental data with theoretical computations based on a symmetric and asymmetric lung. Bulletin of Mathematical Biology, 45(3):409–424, May 1983. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459399.

Witten:1983:DPA

[661] Matthew Witten. The diffusion process approach to one-compartmental stochastic models: a mathematical note. Bulletin of Mathematical Biology, 45(3):425–430, May 1983. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02459400.

Capasso:1983:CSI

[662] V. Capasso and S. L. Paveri-Fontana. Comments on statistical independence for linear stochastic multicompartmental systems. Bulletin of Mathematical Biology, 45(3):431–435, May 1983. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/article/10.1007/BF02459401.

Deakin:1983:RLJ

[663] M. A. B. Deakin. Reply to letter by Jane Cronin (1982). Bulletin of Mathematical Biology, 45(3):436, May 1983. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02459402.

Anonymous:1983:F

[664] Anonymous. Foreword. Bulletin of Mathematical Biology, 45(4): 437, July 1983. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459580.

DeLisi:1983:I

[665] C. DeLisi, M. Iannelli, and G. Koch. Introduction. Bulletin of Mathematical Biology, 45(4):439–442, July 1983. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02459581.

Holden:1983:SPN

[666] A. V. Holden. Stochastic processes in neurophysiology: Transformation from point to continuous processes. Bulletin of Mathematical Biology, 45(4):443–465, July 1983. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459582.

DeLisi:1983:SPI

[667] Charles DeLisi. Stochastic problems in information transfer across the plasma membrane. Bulletin of Mathematical Biology, 45(4): 467–482, July 1983. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459583.

Kliemann:1983:QTS

[668] Wolfgang Kliemann. Qualitative theory of stochastic dynamical systems — applications to life sciences. Bulletin of Mathematical Biology, 45 (4):483–506, July 1983. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459584.

Jagers:1983:SMC

[669] Peter Jagers. Stochastic models for cell kinetics. Bulletin of Mathematical Biology, 45(4):507–519, July 1983. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02459585.

Maruyama:1983:STP

[670] Takeo Maruyama. Stochastic theory of population genetics. Bulletin of Mathematical Biology, 45(4):521–554, July 1983. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/article/10.1007/BF02459586.

Capasso:1983:ACS

[671] V. Capasso and S. L. Paveri-Fontana. Analysis of the correlations for stochastic multicompartmental systems. Bulletin of Mathematical Biology, 45(4):555–569, July 1983. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02459587.

DiMasi:1983:NLF

[672] Giovanni B. Di Masi and Wolfgang J. Runggaldier. Non-linear filtering with discontinuous observations and applications to life sciences. Bulletin of Mathematical Biology, 45(4):571–577, July 1983. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02459588.

Ehrhardt:1983:IPS

[673] Manfred Ehrhardt. Invariant probabilities for systems in a random environment — with applications to the Brusselator. Bulletin of Mathematical Biology, 45(4):579–590, July 1983. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02459589.

Pardoux:1983:SLS

[674] E. Pardoux. Semi-linear stochastic models: Computation of moments and conditional moments. Bulletin of Mathematical Biology, 45(4): 591–597, July 1983. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459590.

DaPrato:1983:SAS

[675] G. Da Prato. On some abstract stochastic differential equations. Bulletin of Mathematical Biology, 45(4):599–603, July 1983. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02459591.

Bertuzzi:1983:RVC

[676] A. Bertuzzi and A. Gandolfi. Recent views on the cell cycle structure. Bulletin of Mathematical Biology, 45(4):605–616, July 1983. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02459592.

Rittgen:1983:CBP

[677] W. Rittgen. Controlled branching processes and their applications to normal and malignant haematopoiesis. Bulletin of Mathematical Biology, 45(4):617–626, July 1983. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459593.

DelGrosso:1983:LDE

[678] G. Del Grosso, A. Gerardi, and F. Marchetti. Lateral diffusion and eigenvalue estimates. Bulletin of Mathematical Biology, 45(4): 627–634, July 1983. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459594.

Braumann:1983:PGR

[679] Carlos A. Braumann. Population growth in random environments. Bulletin of Mathematical Biology, 45(4):635–641, July 1983. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02459595.

Metz:1983:CTM

[680] Hans A. J. Metz, Herman Dienske, Gerrit de Jonge, and Felix A. Putters. Continuous-time Markov chains as models for animal behaviour. Bulletin of Mathematical Biology, 45(4):643–658, July 1983. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02459596.

Anonymous:1983:TIC

[681] Anonymous. Tenth International Cell Cycle Conference. Bulletin of Mathematical Biology, 45(4):659, July 1983. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02459597.

Hiltmann:1983:ODS

[682] P. Hiltmann and P. Lory. On oxygen diffusion in a spherical cell with Michaelis–Menten oxygen uptake kinetics. Bulletin of Mathematical Biology, 45(5):661–664, September 1983. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460043.

Havel:1983:TPD

[683] Timothy F. Havel, Irwin D. Kuntz, and Gordon M. Crippen. The theory and practice of distance geometry. Bulletin of Mathematical Biology, 45(5):665–720, September 1983. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460044.

Parvathamma:1983:MAP

[684] S. Parvathamma and Rathna Devanathan. Microcontinuum approach to the pulsatile flow in tubes with and without longitudinal vibration. Bulletin of Mathematical Biology, 45(5):721–737, September 1983. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02460045.

Youn:1983:CIP

[685] Jang H. Youn, Hee C. Kim, Hong K. Lee, and Byoung G. Min. Computed insulin profiles for normalization of glycemia in diabetic patients.

Bulletin of Mathematical Biology, 45(5):739–748, September 1983. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02460046.

Larsen:1983:TTA

[686] Jesper Larsen, Henrik Lund-Andersen, and Bent Krogsaa. Transient transport across the blood-retina barrier. Bulletin of Mathematical Biology, 45(5):749–758, September 1983. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460047.

Lee:1983:MMS

[687] Chuyan-Yih Lee. Molecular mechanism of sodium conductance changes in nerve: The role of electron transfer and energy migration. Bulletin of Mathematical Biology, 45(5):759–780, September 1983. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02460048.

Muir:1983:DSN

[688] A. Muir and M. W. Warner. The dynamics of symmetric nets. Bulletin of Mathematical Biology, 45(5):781–792, September 1983. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02460049.

Shukla:1983:CGS

[689] V. P. Shukla. Conditions for global stability of two-species population models with discrete time delay. Bulletin of Mathematical Biology, 45(5): 793–805, September 1983. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460050.

Bosserman:1983:FAS

[690] Robert W. Bosserman. Flow analysis sensitivities for models of energy or material flow. Bulletin of Mathematical Biology, 45(5):807– 826, September 1983. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460051.

Huddleston:1983:PDA

[691] John V. Huddleston. Population dynamics with age- and time-dependent birth and death rates. Bulletin of Mathematical Biology, 45(5):827– 836, September 1983. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460052.

Rosen:1983:ASD

[692] Gerald Rosen. Analytical solutions for distributions of chemotactic bacteria. Bulletin of Mathematical Biology, 45(5):837–847, September 1983. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02460053.

Bertz:1983:CGM

[693] Steven H. Bertz. On the complexity of graphs and molecules. Bulletin of Mathematical Biology, 45(5):849–855, September 1983. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02460054.

Sud:1983:EPA

[694] V. K. Sud, H. E. Von Gierke, I. Kaleps, and H. L. Oestreicher. Effects of periodic accelerations on blood flow in arteries. Bulletin of Mathematical Biology, 45(5):857–867, September 1983. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460055.

Lurie:1983:TAB

[695] David Lurié, Joaquim Valls, and Jorge Wagensberg. Thermodynamic approach to biomass distribution in ecological systems. Bulletin of Mathematical Biology, 45(5):869–872, September 1983. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/article/10.1007/BF02460056.

Rescigno:1983:DDB

[696] Aldo Rescigno, James S. Beck, and H. Joseph Goren. Determination of dependence of binding parameters on receptor occupancy — a correction. Bulletin of Mathematical Biology, 45(5):873–874, September 1983. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02460057. See [599].

Anonymous:1983:Aa

[697] Anonymous. Announcements. Bulletin of Mathematical Biology, 45(5): 875, September 1983. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460058.

Takeuchi:1983:EBS

[698] Yasuhiro Takeuchi and Norihiko Adachi. Existence and bifurcation of stable equilibrium in two-prey, one-predator communities. Bulletin of Mathematical Biology, 45(6):877–900, November 1983. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02458820.

Levine:1983:BPS

[699] Daniel S. Levine. Bifurcating periodic solutions for a class of agestructured predator–prey systems. Bulletin of Mathematical Biology, 45(6):901–915, November 1983. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02458821.

Ochoa:1983:NLA

[700] Francisco Lara Ochoa and J. D. Murray. A non-linear analysis for spatial structure in a reaction–diffusion model. Bulletin of Mathematical Biology, 45(6):917–930, November 1983. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02458822.

Kampfner:1983:CME

[701] Roberto R. Kampfner and Michael Conrad. Computational modeling of evolutionary learning processes in the brain. Bulletin of Mathematical Biology, 45(6):931–968, November 1983. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02458823.

Kampfner:1983:SBS

[702] Roberto R. Kampfner and Michael Conrad. Sequential behavior and stability properties of enzymatic neuron networks. Bulletin of Mathematical Biology, 45(6):969–980, November 1983. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02458824.

Worz-Busekros:1983:DCC

[703] Angelika Wörz-Busekros and Peter Lange. Dynamics of a continuous culture with catalysis of the growth-limiting substrate by an enzyme of the cells. Bulletin of Mathematical Biology, 45(6):981–990, November 1983. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02458825.

Freedman:1983:TBM

[704] H. L. Freedman and V. Sree Hari Rao. The trade-off between mutual interference and time lags in predator–prey systems. Bulletin of Mathematical Biology, 45(6):991–1004, November 1983. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02458826.

Li:1983:NSP

[705] John K.-J. Li. A new similarity principle for cardiac energetics. Bulletin of Mathematical Biology, 45(6):1005–1011, November 1983. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02458827.

Schoepfle:1983:STH

[706] Gordon M. Schoepfle, John T. Tarvin, and Richard M. Martin. Simulated tetanic hyperpolarization by sodium loading of a neuronal system in which active transport is linked to the Frankenhaeuser–Huxley equations. Bulletin of Mathematical Biology, 45(6):1013–1028, November 1983. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02458828.

Louie:1983:CSTa

[707] A. H. Louie. Categorical system theory and the phenomenological calculus. Bulletin of Mathematical Biology, 45(6):1029–1045, November 1983. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02458829.

Louie:1983:CSTb

[708] A. H. Louie. Categorical system theory. Bulletin of Mathematical Biology, 45(6):1047–1072, November 1983. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02458830.

Monticelli:1983:FPT

[709] G. Monticelli and F. Celentano. Further properties of the twomembrane model. Bulletin of Mathematical Biology, 45(6):1073– 1096, November 1983. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458831.

Anonymous:1983:Ab

[710] Anonymous. Announcement. Bulletin of Mathematical Biology, 45(6): 1097, November 1983. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458832.

Supnick:1984:MS

[711] Martin Supnick. On the measurement of shapes. Bulletin of Mathematical Biology, 46(1):1–10, January 1984. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02463719.

Schuster:1984:RSS

[712] P. Schuster and K. Sigmund. Random selection — a simple model based on linear birth and death processes. Bulletin of Mathematical Biology, 46(1):11–17, January 1984. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02463720.

Lauffenburger:1984:TBC

[713] D. Lauffenburger, C. R. Kennedy, and R. Aris. Traveling bands of chemotactic bacteria in the context of population growth. Bulletin of Mathematical Biology, 46(1):19–40, January 1984. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02463721.

Hagglund:1984:SSB

[714] Jarl V. Hägglund, George Eisenman, and John P. Sandblom. Single-salt behavior of a symmetrical 4-site channel with barriers at its middle and ends. Bulletin of Mathematical Biology, 46(1):41–80, January 1984. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02463722.

Kuyk:1984:CTM

[715] Willem Kuyk. A catastrophe theoretic model of heat regulation in homeothermic animals. Bulletin of Mathematical Biology, 46(1):81– 102, January 1984. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02463723.

Misra:1984:PWV

[716] J. C. Misra and S. I. Singh. Pulse wave velocities in the aorta. Bulletin of Mathematical Biology, 46(1):103–114, January 1984. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02463724.

Khanin:1984:MMP

[717] M. A. Khanin and I. B. Bukharov. A mathematical model of the pathological functional state of the oxygen transport system. Bulletin of Mathematical Biology, 46(1):115–125, January 1984. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02463725.

Kamiya:1984:ARW

[718] Akira Kamiya, Razaq Bukhari, and Tatsuo Togawa. Adaptive regulation of wall shear stress optimizing vascular tree function. Bulletin of Mathematical Biology, 46(1):127–137, January 1984. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02463726.

Brown:1984:SIP

[719] Grayson C. Brown. Stability in an insect-pathogen model incorporating age-dependent immunity and seasonal host reproduction. Bulletin of Mathematical Biology, 46(1):139–153, January 1984. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02463727.

Baltzis:1984:CTM

[720] Basil C. Baltzis and A. G. Fredrickson. Coexistence of two microbial populations competing for a renewable resource in a non-predator–prey system. Bulletin of Mathematical Biology, 46(1):155–174, January 1984. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02463728.

De:1984:RPP

[721] S. S. De. Random predator-prey interactions in a varying environment: Extinction or survival. Bulletin of Mathematical Biology, 46(1): 175–184, January 1984. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02463729.

Huddleston:1984:PDA

[722] J. V. Huddleston. Population dynamics with age- and time-dependent birth and death rates. Bulletin of Mathematical Biology, 46(1):185, January 1984. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
BF02463730;   https://link.springer.com/content/pdf/10.1007/
BF02463730.pdf.

Swanson:1984:UCA

[723] Rosemarie Swanson. A unifying concept for the amino acid code. Bulletin of Mathematical Biology, 46(2):187–203, March 1984. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02460068.

Franko:1984:MDO

[724] A. J. Franko and H. I. Freedman. Model of diffusion of oxygen to spheroids grown in stationary medium — I. Complete spherical symmetry. Bulletin of Mathematical Biology, 46(2):205–217, March 1984. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02460069.

Sobel:1984:FUC

[725] M. Sobel, S. R. Bernard, V. R. R. Uppuluri, and C. W. Nestor. Foururn catenary model for excretion. Bulletin of Mathematical Biology, 46 (2):219–227, March 1984. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460070.

Rees:1984:GSS

[726] D. C. Rees. A general solution for the steady-state kinetics of immobilized enzyme systems. Bulletin of Mathematical Biology, 46(2): 229–234, March 1984. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460071.

Rosen:1984:SSH

[727] Gerald Rosen. Stability of spatially heterogeneous steady-state distributions of oxygen-chemotactic aerobic/anaerobic bacteria. Bulletin of Mathematical Biology, 46(2):235–246, March 1984. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/article/10.1007/BF02460072.

Bateman:1984:CO

[728] J. M. Bateman and T. R. Harris. Capillary operators. Bulletin of Mathematical Biology, 46(2):247–268, March 1984. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02460073.

Nandapurkar:1984:CDD

[729] P. Nandapurkar and V. Hlavacek. Concentration-dependent diffusion coefficient and dissipative structures. Bulletin of Mathematical Biology, 46(2):269–282, March 1984. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460074.

VanPelt:1984:CTA

[730] J. Van Pelt and R. W. H. Verwer. Cuttrees in the topological analysis of branching patterns. Bulletin of Mathematical Biology, 46(2): 283–294, March 1984. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460075.

Case:1984:LTM

[731] James H. Case and Paul S. Fisher. Long-term memory modules. Bulletin of Mathematical Biology, 46(2):295–326, March 1984. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02460076.

Day:1984:PLM

[732] William H. E. Day. Properties of Levenshtein metrics on sequences. Bulletin of Mathematical Biology, 46(2):327–332, March 1984. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02460077.

Rosen:1984:CMP

[733] Gerald Rosen. Commensalism of methane-producing and motile aerobic bacteria in certain freshwater wetlands. Bulletin of Mathematical Biology, 46(2):333–336, March 1984. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460078.

Anonymous:1984:Aa

[734] Anonymous. Announcement. Bulletin of Mathematical Biology, 46 (2):337, March 1984. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460079.

Phillipson:1984:DMB

[735] Paul E. Phillipson, Peter Schuster, and Fritz Kemler. Dynamical machinery of a biochemical clock. Bulletin of Mathematical Biology, 46 (3):339–355, May 1984. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462010.

Gard:1984:PSF

[736] Thomas C. Gard. Persistence in stochastic food web models. Bulletin of Mathematical Biology, 46(3):357–370, May 1984. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/article/10.1007/BF02462011.

Reed:1984:ECA

[737] Jon Reed and Nils C. Stenseth. Evolution of cannibalism in an age-structured population. Bulletin of Mathematical Biology, 46(3): 371–377, May 1984. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462012.

Oden:1984:AEG

[738] Neal L. Oden and Kwang-Tsao Shao. An algorithm to equiprobably generate all directed trees with k labeled terminal nodes and unlabeled interior nodes. Bulletin of Mathematical Biology, 46(3):379–387, May 1984. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02462013.

Horsfield:1984:CSG

[739] Keith Horsfield and Alison Thurlbeck. Computer simulation of the geometry of the human bronchial tree. Bulletin of Mathematical Biology, 46(3):389–398, May 1984. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462014.

Leung:1984:SAS

[740] H. K. Leung. Stability analysis of a stochastic model for biomolecular selection. Bulletin of Mathematical Biology, 46(3):399–406, May 1984. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02462015.

Grasman:1984:MME

[741] J. Grasman. The mathematical modeling of entrained biological oscillators. Bulletin of Mathematical Biology, 46(3):407–422, May 1984. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02462016.

Nelson:1984:SME

[742] Sarah J. Nelson. A stochastic model of the effects of ionizing radiation on mammalian cells in vitro. Bulletin of Mathematical Biology, 46(3):423–446, May 1984. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462017.

Leblanc:1984:SMB

[743] Raymond Leblanc and Gerald M. Lefebvre. A stochastic model of bacterial spore germination. Bulletin of Mathematical Biology, 46(3): 447–460, May 1984. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462018.

Anonymous:1984:ESS

[744] Anonymous. Editorial: Software survey section. Bulletin of Mathematical Biology, 46(4):i–iv, July 1984. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02459512.

Martinez:1984:F

[745] Hugo M. Martinez. Foreword. Bulletin of Mathematical Biology, 46 (4):461–465, July 1984. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459496.

Anonymous:1984:MOD

[746] Anonymous. Margaret Oakley Dayhoff, 1925–1983. Bulletin of Mathematical Biology, 46(4):467–472, July 1984. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02459497.

Waterman:1984:GMS

[747] Michael S. Waterman. General methods of sequence comparison. Bulletin of Mathematical Biology, 46(4):473–500, July 1984. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02459498.

Sellers:1984:PRG

[748] Peter H. Sellers. Pattern recognition in genetic sequences by mismatch density. Bulletin of Mathematical Biology, 46(4):501–514, July 1984. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02459499.

Waterman:1984:PRS

[749] M. S. Waterman, R. Arratia, and D. J. Galas. Pattern recognition in several sequences: Consensus and alignment. Bulletin of Mathematical Biology, 46(4):515–527, July 1984. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02459500.

Dayhoff:1984:DWD

[750] Judith E. Dayhoff. Distinguished words in data sequences: Analysis and applications to neural coding and other fields. Bulletin of Mathematical Biology, 46(4):529–543, July 1984. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02459501.

Orcutt:1984:SPS

[751] Bruce C. Orcutt and Winona C. Barker. Searching the protein sequence database. Bulletin of Mathematical Biology, 46(4):545–552, July 1984. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02459502.

Fredman:1984:ACE

[752] M. L. Fredman. Algorithms for computing evolutionary similarity measures with length independent gap penalties. Bulletin of Mathematical Biology, 46(4):553–566, July 1984. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02459503.

Waterman:1984:LGS

[753] Michael S. Waterman and Marcela D. Perlwitz. Line geometries for sequence comparisons. Bulletin of Mathematical Biology, 46(4): 567–577, July 1984. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459504.

Davison:1984:NMS

[754] Dan Davison and Keith H. Thompson. A non-metric sequence alignment program. Bulletin of Mathematical Biology, 46(4):579–590, July 1984. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02459505.

Zuker:1984:RSS

[755] Michael Zuker and David Sankoff. RNA secondary structures and their prediction. Bulletin of Mathematical Biology, 46(4):591–621, July 1984. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02459506.

Swanson:1984:VRA

[756] Rosemarie Swanson. A vector representation for amino acid sequences. Bulletin of Mathematical Biology, 46(4):623–639, July 1984. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02459507.

Jimenez-Montano:1984:SSP

[757] M. A. Jiménez-Montaño. On the syntactic structure of protein sequences and the concept of grammar complexity. Bulletin of Mathematical Biology, 46(4):641–659, July 1984. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02459508.

Saroff:1984:UPS

[758] H. A. Saroff. The uniqueness of protein sequences. Uniqueness diagrams for the Dayhoff file — 1984. Bulletin of Mathematical Biology, 46(4):661–672, July 1984. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459509.

Havel:1984:DGP

[759] Timothy Havel and Kurt Wüthrich. A distance geometry program for determining the structures of small proteins and other macromolecules from nuclear magnetic resonance measurements of intramolecular ^{1}H– 1H proximities in solution. Bulletin of Mathematical Biology, 46(4): 673–698, July 1984. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459510.

Jungck:1984:MTM

[760] John R. Jungck and Robert M. Friedman. Mathematical tools for molecular genetics data: an annotated bibliography. Bulletin of Mathematical Biology, 46(4):699–744, July 1984. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02459511.

Louie:1984:SID

[761] A. H. Louie and R. L. Somorjai. Stieltjes integration and differential geometry: a model for enzyme recognition, discrimination, and catalysis. Bulletin of Mathematical Biology, 46(5–6):745–764, September 1984. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02462069.

Kirby:1984:ENR

[762] Kevin G. Kirby and Michael Conrad. The enzymatic neuron as a reaction–diffusion network of cyclic nucleotides. Bulletin of Mathematical Biology, 46(5–6):765–783, September 1984. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02462070.

Clowes:1984:GCT

[763] J. S. Clowes and G. D. Wassermann. Genetic control theory of developmental events. Bulletin of Mathematical Biology, 46(5–6):785– 825, September 1984. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462071.

Avanzolini:1984:CMA

[764] G. Avanzolini and A. Cappello. The characteristics method applied to the study of muscle dynamics. Bulletin of Mathematical Biology, 46(5–6): 827–844, September 1984. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462072.

Walker:1984:BBC

[765] C. C. Walker. Beyond the binary case in random nets. Bulletin of Mathematical Biology, 46(5–6):845–857, September 1984. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/article/10.1007/BF02462073.

Narita:1984:ODF

[766] Kazuaki Narita. One-dimensional food chain equations and their generalization. Bulletin of Mathematical Biology, 46(5–6):859–868, September 1984. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462074.

Timischl:1984:ILS

[767] W. Timischl. Influence of landscape on the spread of an infection. Bulletin of Mathematical Biology, 46(5–6):869–877, September 1984. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02462075.

Klinger:1984:SSN

[768] H. G. Klinger. A simple solution of the non-stationary heat transport problem in capillarized biological tissue. Bulletin of Mathematical Biology, 46(5–6):879–889, September 1984. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02462076.

Bobisud:1984:CCP

[769] L. E. Bobisud and C. J. Potratz. Cannibalism in clumped populations. Bulletin of Mathematical Biology, 46(5–6):891–901, September 1984. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02462077.

Argemi:1984:QSD

[770] José Argémi, Hélène Chagneux, Christian Ducreux, and Maurice Gola. Qualitative study of a dynamical system for metrazol-induced paroxysmal depolarization shifts. Bulletin of Mathematical Biology, 46(5–6): 903–922, September 1984. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462078.

Stinebrickner:1984:CTI

[771] Ralph Stinebrickner. s-consensus trees and indices. Bulletin of Mathematical Biology, 46(5–6):923–935, September 1984. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/article/10.1007/BF02462079.

Sud:1984:BFS

[772] V. K. Sud and G. S. Sekhon. Blood flow subject to a single cycle of body acceleration. Bulletin of Mathematical Biology, 46(5–6):937– 949, September 1984. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462080.

Saroff:1984:NES

[773] Harry A. Saroff. A note on the evaluation of similarity (homology) of short sequences with long sequences. Bulletin of Mathematical Biology, 46(5–6):951–961, September 1984. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02462081.

Rosen:1984:CCG

[774] Gerald Rosen. Characterizing conditions for generalized Verhulst logistic growth of a biological population. Bulletin of Mathematical Biology, 46(5–6):963–965, September 1984. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02462082.

Rosen:1984:CBD

[775] Gerald Rosen. Comment on biomass diversity in ecology. Bulletin of Mathematical Biology, 46(5–6):967–969, September 1984. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02462083.

Conrad:1984:PSE

[776] Michael Conrad. Physik der Selbstorganization und Evolution. (German) [Physics of self-organization and evolution]. Bulletin of Mathematical Biology, 46(5–6):971–972, September 1984. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02462084.

Anonymous:1984:Ab

[777] Anonymous. Announcements. Bulletin of Mathematical Biology, 46(5–6): 973–974, September 1984. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462085.

Anonymous:1985:ESSa

[778] Anonymous. Editorial: Software survey section. Bulletin of Mathematical Biology, 47(1):i–iv, January 1985. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02459653.

Honerkamp:1985:CSF

[779] J. Honerkamp, G. Mutschler, and R. Seitz. Coupling of a slow and a fast oscillator can generate bursting. Bulletin of Mathematical Biology, 47(1):1–21, January 1985. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459643.

Taylor:1985:PPL

[780] G. C. Taylor. Primitivity of products of Leslie matrices. Bulletin of Mathematical Biology, 47(1):23–34, January 1985. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02459644.

Sud:1985:AFU

[781] V. K. Sud and G. S. Sekhon. Arterial flow under periodic body acceleration. Bulletin of Mathematical Biology, 47(1):35–53, January 1985. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02459645.

Misra:1985:DSL

[782] J. C. Misra and S. I. Singh. Distribution of stresses in the left ventricular wall of the intact heart. Bulletin of Mathematical Biology, 47 (1):53–70, January 1985. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459646.

Alian:1985:MAT

[783] S. E. D. Alian and D. A. Linkens. Mode analysis of a tubular structure of coupled non-linear oscillators for digestive-tract modelling. Bulletin of Mathematical Biology, 47(1):71–110, January 1985. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02459647.

Giro:1985:MCS

[784] A. Giro, J. A. Padro, J. Valls, and J. Wagensberg. Monte Carlo simulation of an ecosystem: a matching between two levels of observation. Bulletin of Mathematical Biology, 47(1):111–122, January 1985. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02459648.

Titomir:1985:SAC

[785] Leonid I. Titomir and Peter Kneppo. Simultaneous analysis of the cardiac electric and magnetic fields using the scalar multipole expansion. Bulletin of Mathematical Biology, 47(1):123–143, January 1985. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02459649.

Anderson:1985:ABF

[786] N. Anderson and A. M. Arthurs. Analytical bounding functions for diffusion problems with Michaelis–Menten kinetics. Bulletin of Mathematical Biology, 47(1):145–153, January 1985. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02459650.

Derome:1985:B

[787] Jean-Robert Derome. Biomathematiques. Bulletin of Mathematical Biology, 47(1):155, January 1985. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02459651.

Havel:1985:TPD

[788] T. F. Havel, I. D. Kuntz, and B. Crippen. The theory and practice of distance geometry. Bulletin of Mathematical Biology, 47(1): 157, January 1985. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459652;   http://link.springer.com/content/pdf/10.
1007/BF02459652.pdf.

Anonymous:1985:ESSb

[789] Anonymous. Editorial: Software survey section. Bulletin of Mathematical Biology, 47(2):i–iv, March 1985. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460042.

Martinez:1985:ECa

[790] Hugo M. Martinez. Editor’s comments. Bulletin of Mathematical Biology, 47(2):159, March 1985. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460028.

Rapoport:1985:AGT

[791] Anatol Rapoport. Applications of game-theoretic concepts in biology. Bulletin of Mathematical Biology, 47(2):161–192, March 1985. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02460029.

Leitnaker:1985:NLC

[792] Mary G. Leitnaker and Peter Purdue. Non-linear compartmental systems: extensions of S. R. Bernard’s urn model. Bulletin of Mathematical Biology, 47(2):193–204, March 1985. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460030.

Bowman:1985:SAD

[793] K. O. Bowman, L. R. Shenton, and S. R. Bernard. Study of age dependent half-life of iodine in man: a reinforcement-depletion urn model. Bulletin of Mathematical Biology, 47(2):205–213, March 1985. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02460031.

Day:1985:FCI

[794] William H. E. Day and F. R. McMorris. A formalization of consensus index methods. Bulletin of Mathematical Biology, 47(2):215– 229, March 1985. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460032.

Leung:1985:EME

[795] H. K. Leung. Expansion of the master equation for a biomolecular selection model. Bulletin of Mathematical Biology, 47(2):231– 238, March 1985. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460033.

Demetrius:1985:PEB

[796] Lloyd Demetrius, Peter Schuster, and Karl Sigmund. Polynucleotide evolution and branching processes. Bulletin of Mathematical Biology, 47 (2):239–262, March 1985. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460034.

Potscher:1985:MOS

[797] B. M. Pötscher. Moments and order statistics of extinction times in multitype branching processes and their relation to random selection models. Bulletin of Mathematical Biology, 47(2):263–272, March 1985. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02460035.

Katz:1985:FAG

[798] Michael J. Katz and Edwin B. George. Fractals and the analysis of growth paths. Bulletin of Mathematical Biology, 47(2):273– 286, March 1985. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460036.

Malice:1985:LSC

[799] M.-P. Malice and C. Lefevre. On linear stochastic compartmental models in discrete time. Bulletin of Mathematical Biology, 47(2): 287–293, March 1985. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460037.

Erbe:1985:MPM

[800] L. H. Erbe and H. I. Freedman. Modeling persistence and mutual interference among subpopulations of ecological communities. Bulletin of Mathematical Biology, 47(2):295–304, March 1985. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02460038.

Verwer:1985:TAB

[801] Ronald W. H. Verwer and Jaap Van Pelt. Topological analysis of binary tree structures when occasional multifurcations occur. Bulletin of Mathematical Biology, 47(2):305–316, March 1985. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02460039.

Chiang:1985:DPC

[802] Chun Chiang. Depolarized potassium currents and time delay in nerve membrane. Bulletin of Mathematical Biology, 47(2):317–320, March 1985. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460040.

Anonymous:1985:A

[803] Anonymous. Announcements. Bulletin of Mathematical Biology, 47 (2):321, March 1985. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460041.

Anonymous:1985:ESSc

[804] Anonymous. Editorial: Software survey section. Bulletin of Mathematical Biology, 47(3):i–iv, May 1985. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02459925.

VanPelt:1985:GMI

[805] J. Van Pelt and R. W. H. Verwer. Growth models (including terminal and segmental branching) for topological binary trees. Bulletin of Mathematical Biology, 47(3):323–336, May 1985. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02459919.

Domach:1985:CMC

[806] M. M. Domach, R. Armstrong, and M. S. Jhon. Cell mass concentration variance resulting from fluctuations in the incoming nutrient concentration to a chemostat. Bulletin of Mathematical Biology, 47 (3):337–342, May 1985. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459920.

Nicolis:1985:CDI

[807] John S. Nicolis and Ichiro Tsuda. Chaotic dynamics of information processing: The “magic number seven plus–minus two” revisited. Bulletin of Mathematical Biology, 47(3):343–365, May 1985. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/article/10.1007/BF02459921.

Cariani:1985:CTS

[808] Peter Cariani and Narendra S. Goel. On the computation of the tertiary structure of globular proteins — IV. Use of secondary structure information. Bulletin of Mathematical Biology, 47(3):367–407, May 1985. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02459922.

Deakin:1985:ETD

[809] Michael A. B. Deakin, T. O. Neild, and R. G. Turner. The extension of two-dimensional cable theory to arteries and arterioles. Bulletin of Mathematical Biology, 47(3):409–424, May 1985. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02459923.

Bass:1985:UFT

[810] Ludvik Bass. The unidirectional flux transient as a tool for quantifying parallel diffusional pathways through membranes. Exact solution for two pathways. Bulletin of Mathematical Biology, 47(3):425–434, May 1985. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02459924.

Anonymous:1985:ESSd

[811] Anonymous. Editorial: Software survey section. Bulletin of Mathematical Biology, 47(4):i–iv, July 1985. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460016.

Martinez:1985:ECb

[812] Hugo M. Martinez. Editor’s comments. Bulletin of Mathematical Biology, 47(4):435, July 1985. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460005.

Davison:1985:SSH

[813] Dan Davison. Sequence similarity (‘homology’) searching for molecular biologists. Bulletin of Mathematical Biology, 47(4):437–474, July 1985.

CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02460006.

Babloyantz:1985:PRR

[814] A. Babloyantz and A. Bellemans. Pattern regulation in reaction–diffusion systems — the problem of size invariance. Bulletin of Mathematical Biology, 47(4):475–487, July 1985. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460007.

Solimano:1985:ESE

[815] Fortunata Solimano. The existence of stable equilibria in Volterra predator–prey systems represented by loop graphs. Bulletin of Mathematical Biology, 47(4):489–494, July 1985. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02460008.

Demiray:1985:ETP

[816] Hilmi Demiray. Effects of twist on pulse waves in arteries. Bulletin of Mathematical Biology, 47(4):495–502, July 1985. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/article/10.1007/BF02460009.

Landahl:1985:MMO

[817] Herbert D. Landahl, Javier Garzón, and Nancy M. Lee. Mathematical modeling of opiate binding to mouse brain membrane. Bulletin of Mathematical Biology, 47(4):503–512, July 1985. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02460010.

Schumaker:1985:IIF

[818] Mark F. Schumaker and W. Horsthemke. The influence of internal fluctuations and external noise on the gene frequency in a haploid population. Bulletin of Mathematical Biology, 47(4):513–533, July 1985. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02460011.

Sourdis:1985:MNB

[819] John Sourdis. A method for numbering binary rooted and unrooted phylogenies and their hypothetical taxonomic units. Bulletin of Mathematical Biology, 47(4):535–543, July 1985. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02460012.

Haldar:1985:ESS

[820] K. Haldar. Effects of the shape of stenosis on the resistance to blood flow through an artery. Bulletin of Mathematical Biology, 47 (4):545–550, July 1985. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460013.

Joshi:1985:CFS

[821] Rajani R. Joshi. On control of the function of the sodium-potassium pump in malignant cells. Bulletin of Mathematical Biology, 47(4): 551–564, July 1985. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460014.

Ferron:1985:MAR

[822] G. A. Ferron, B. Haider, and W. G. Kreyling. A method for the approximation of the relative humidity in the upper human airways. Bulletin of Mathematical Biology, 47(4):565–589, July 1985. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02460015.

Anonymous:1985:ESSe

[823] Anonymous. Editorial: Software survey section. Bulletin of Mathematical Biology, 47(5):i–iv, September 1985. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460134.

Anonymous:1985:ESSf

[824] Anonymous. Editorial: Software survey section. Bulletin of Mathematical Biology, 47(5):i–iv, September 1985. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460134.

Lumsden:1985:HBF

[825] Charles J. Lumsden. Hierarchical behavior in fit dynamical systems. Bulletin of Mathematical Biology, 47(5):591–612, September 1985. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02460128.

Guha:1985:FAB

[826] H. S. Guha and T. K. Chaudhury. Flow in an arteriosclerotic blood vessel with rigid permeable wall. Bulletin of Mathematical Biology, 47(5):

613–627, September 1985. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460129.

Masui:1985:MAS

[827] Sayuri Masui and Charles J. Lumsden. A mathematical approach to semantic network development. Bulletin of Mathematical Biology, 47(5): 629–650, September 1985. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460130.

Bateman:1985:COI

[828] J. M. Bateman. Capillary operators — II. Bulletin of Mathematical Biology, 47(5):651–668, September 1985. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460131.

Holmes:1985:TAD

[829] Mark H. Holmes. A theoretical analysis for determining the nonlinear hydraulic permeability of a soft tissue from a permeation experiment. Bulletin of Mathematical Biology, 47(5):669–683, September 1985. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02460132.

Messinger-Rapport:1985:ETB

[830] B. J. Messinger-Rapport and Y. Rudy. Effects of the torso boundary and internal conductivity interfaces in electrocardiography: an evaluation of the ‘infinite medium’ approximation. Bulletin of Mathematical Biology, 47(5):685–694, September 1985. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460133.

Anonymous:1985:ESSg

[831] Anonymous. Editorial: Software survey section. Bulletin of Mathematical Biology, 47(6):i–iv, November 1985. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02469308.

Martinez:1985:ECc

[832] Hugo M. Martinez. Editor’s comments. Bulletin of Mathematical Biology, 47(6):695, November 1985. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02469298.

Kloeden:1985:CP

[833] P. E. Kloeden and A. I. Mees. Chaotic phenomena. Bulletin of Mathematical Biology, 47(6):697–738, November 1985. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02469299.

Hallam:1985:EAN

[834] T. G. Hallam and M. N. Antonios. An environ analysis for nonlinear compartment models. Bulletin of Mathematical Biology, 47(6):739– 748, November 1985. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02469300.

Schoepfle:1985:DTH

[835] Gordon M. Schoepfle, John T. Tarvin, and Richard M. Martin. Dependence of tetanic hyperpolarization on pump conductance in the cyanidetreated Xenopus node. Bulletin of Mathematical Biology, 47(6):749– 755, November 1985. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02469301.

Greene:1985:SSB

[836] Peter R. Greene. Stress-strain behavior for curved exponential strips. Bulletin of Mathematical Biology, 47(6):757–764, November 1985. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02469302.

Majhi:1985:MNF

[837] S. N. Majhi and L. Usha. A mathematical note on the Fahraeus– Lindqvist effect in power law fluid. Bulletin of Mathematical Biology, 47(6):765–769, November 1985. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02469303.

Lopez-Sanchez:1985:ZAT

[838] Jesus Lopez-Sanchez and Alberto Perez de Vargas. Zygotic algebra for two linked loci with sexually different recombination rates. Bulletin of Mathematical Biology, 47(6):771–782, November 1985. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02469304.

Khan:1985:PRS

[839] Angshutosh Khan, Debajyoti Bhaumik, and Binayak Dutta-Roy. The possible role of solitonic processes during A to B conformational changes in DNA. Bulletin of Mathematical Biology, 47(6):783–789, November 1985. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02469305.

Sobel:1985:DTR

[840] Milton Sobel, S. R. Bernard, and V. R. R. Uppuluri. On the distribution of the time required to remove white balls from an URN. Bulletin of Mathematical Biology, 47(6):791–797, November 1985. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02469306.

Anonymous:1985:TIS

[841] Anonymous. Third International Summer School on Mathematical Modelling in Physiology and Clinical Medicine. Bulletin of Mathematical Biology, 47(6):799, November 1985. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02469307.

Anonymous:1986:ESSa

[842] Anonymous. Editorial: Software survey section. Bulletin of Mathematical Biology, 48(1):i–iv, January 1986. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460067.

Boon:1986:MCB

[843] Jean-Pierre Boon and Brigitte Herpigny. Model for chemotactic bacterial bands. Bulletin of Mathematical Biology, 48(1):1–19, January 1986. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02460059.

Ghosh:1986:MCL

[844] Subhendu Ghosh and Anjali Mookerjee. A model for cooperative ligand binding at complementary sites of DNA. Bulletin of Mathematical Biology, 48(1):21–27, January 1986. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460060.

Weinstein:1986:EFR

[845] Alan M. Weinstein. An equation for flow in the renal proximal tubule. Bulletin of Mathematical Biology, 48(1):29–57, January 1986. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02460061.

Thompson:1986:IAC

[846] Grant G. Thompson, W. Michael Booty, William J. Liss, and Charles E. Warren. An isoclinal approach to the comparative statics of biological communities. Bulletin of Mathematical Biology, 48(1):59– 75, January 1986. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460062.

Jean:1986:IFF

[847] Roger V. Jean. An interpretation of Fujita’s frequency diagrams in phyllotaxis. Bulletin of Mathematical Biology, 48(1):77–86, January 1986. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02460063.

Rosen:1986:ESS

[848] Gerald Rosen. Existence of steady-state probability distributions in multilocus models for genotype evolution. Bulletin of Mathematical Biology, 48(1):87–95, January 1986. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460064.

Li:1986:DGO

[849] John K.-J. Li. Dominance of geometric overelastic factors in pulse transmission through arterial branching. Bulletin of Mathematical Biology, 48 (1):97–103, January 1986. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460065.

Anonymous:1986:SAM

[850] Anonymous. Seventeenth Annual Meeting of the Classification Society of North America. Bulletin of Mathematical Biology, 48(1): 105, January 1986. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460066.

Anonymous:1986:ESSb

[851] Anonymous. Editorial: Software survey section. Bulletin of Mathematical Biology, 48(2):i–iv, March 1986. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460027.

Kuyk:1986:SHR

[852] Willem Kuyk. Servocontrol of the heart rate in homeothermic animals: On recognizing butterfly catastrophy models in biology. Bulletin of Mathematical Biology, 48(2):107–124, March 1986. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02460017.

Pal:1986:GFE

[853] S. Pal and D. S. Pal. Galerkin’s finite element–Laplace transform technique to transient heat migration in human skin and subcutaneous tissues. Bulletin of Mathematical Biology, 48(2):125–136, March 1986. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02460018.

Harrison:1986:MSE

[854] Gary W. Harrison. Multiple stable equilibria in a predator-prey system. Bulletin of Mathematical Biology, 48(2):137–148, March 1986. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02460019.

Varon:1986:TPK

[855] R. Varón, A. Román, F. García, and F. García Carmona. Transient phase kinetics of activation of human plasminogen. Bulletin of Mathematical Biology, 48(2):149–166, March 1986. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460020.

Goel:1986:SCC

[856] Narendra S. Goel, Carlos F. Doggenweiler, and Richard L. Thompson. Simulation of cellular compaction and internalization in mammalian embryo development as driven by minimization of surface energy. Bulletin of Mathematical Biology, 48(2):167–187, March 1986. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02460021.

Waterman:1986:IGM

[857] Michael S. Waterman and Jerrold R. Griggs. Interval graphs and maps of DNA. Bulletin of Mathematical Biology, 48(2):189–195, March 1986. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02460022.

VanPelt:1986:TPB

[858] J. Van Pelt and R. W. H. Verwer. Topological properties of binary trees grown with order-dependent branching probabilities. Bulletin of Mathematical Biology, 48(2):197–211, March 1986. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02460023.

Tapaswi:1986:PFM

[859] P. K. Tapaswi and A. K. Saha. Pattern formation and morphogenesis: a reaction–diffusion model. Bulletin of Mathematical Biology, 48 (2):213–228, March 1986. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460024.

Duggan:1986:PBN

[860] R. C. Duggan and A. M. Goodman. Pointwise bounds for a nonlinear heat conduction model of the human head. Bulletin of Mathematical Biology, 48(2):229–236, March 1986. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460025.

Anonymous:1986:GRC

[861] Anonymous. 1986 Gordon Research Conference on Theoretical Biology and Biomathematics. Bulletin of Mathematical Biology, 48 (2):237, March 1986. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460026.

Anonymous:1986:ESSc

[862] Anonymous. Editorial: Software survey section. Bulletin of Mathematical Biology, 48(3–4):i–iv, May 1986. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02459694.

Greco:1986:GEF

[863] William R. Greco and Michael C. Kohn. Guest Editors’ foreword. Bulletin of Mathematical Biology, 48(3–4):239–240, May 1986. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02459679.

Greco:1986:RSB

[864] W. R. Greco. The role of simulation in biomathematical modeling. Bulletin of Mathematical Biology, 48(3–4):241–251, May 1986. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02459680.

Skipper:1986:MMC

[865] Howard E. Skipper. On mathematical modeling of critical variables in cancer treatment (goals: Better understanding of the past and better planning in the future). Bulletin of Mathematical Biology, 48(3– 4):253–278, May 1986. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459681.

Coldman:1986:SMO

[866] A. J. Coldman and J. H. Goldie. A stochastic model for the origin and treatment of tumors containing drug-resistant cells. Bulletin of Mathematical Biology, 48(3–4):279–292, May 1986. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02459682.

Menzel:1986:CCT

[867] Daniel B. Menzel and Robert L. Wolpert. Chemical carcinogenesis and toxicity models: Matching complexity to objectives. Bulletin of Mathematical Biology, 48(3–4):293–307, May 1986. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02459683.

Bischoff:1986:PP

[868] Kenneth B. Bischoff. Physiological pharmacokinetics. Bulletin of Mathematical Biology, 48(3–4):309–322, May 1986. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02459684.

Donaghey:1986:CKS

[869] Charles E. Donaghey. Cell kinetics simulation languages. Bulletin of Mathematical Biology, 48(3–4):323–336, May 1986. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/article/10.1007/BF02459685.

Jackson:1986:KSA

[870] Robert C. Jackson. Kinetic simulation of anticancer drug effects on metabolic pathway fluxes: Two case studies. Bulletin of Mathematical Biology, 48(3–4):337–351, May 1986. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02459686.

White:1986:UCS

[871] J. Courtland White. Use of the circuit simulation program SPICE2 for analysis of the metabolism of anticancer drugs. Bulletin of Mathematical Biology, 48(3–4):353–380, May 1986. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02459687.

Swan:1986:CCO

[872] G. W. Swan. Cancer chemotherapy: Optimal control using the Verhulst–Pearl equation. Bulletin of Mathematical Biology, 48(3–4): 381–404, May 1986. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459688.

Kootsey:1986:FDC

[873] J. Mailen Kootsey. Future directions in computer simulation. Bulletin of Mathematical Biology, 48(3–4):405–415, May 1986. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02459689.

Kohn:1986:SCM

[874] Michael C. Kohn. Strategies for computer modeling. Bulletin of Mathematical Biology, 48(3–4):417–426, May 1986. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02459690.

Kootsey:1986:SIS

[875] J. Mailen Kootsey, Michael C. Kohn, Michael D. Feezor, Glen R. Mitchell, and Peter R. Fletcher. SCoP: an interactive simulation control program for micro- and minicomputers. Bulletin of Mathematical Biology, 48(3–4):427–441, May 1986. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02459691.

Newton:1986:CRA

[876] Carol M. Newton. Conference retrospective: an appropriate modeling infrastructure for cancer research. Bulletin of Mathematical Biology, 48 (3–4):443–452, May 1986. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459692.

Anonymous:1986:FCI

[877] Anonymous. First conference of the international federation of classification societies. Bulletin of Mathematical Biology, 48(3–4):453, May 1986. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02459693.

Anonymous:1986:SSS

[878] Anonymous. Software survey section. Bulletin of Mathematical Biology, 48(5–6):i–iv, September 1986. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02462332.

Roy:1986:GSP

[879] A. B. Roy and F. Solimano. Global stability of partially closed foodchains with resources. Bulletin of Mathematical Biology, 48(5–6):455– 468, September 1986. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462317.

So:1986:PGS

[880] Joseph W. H. So and H. I. Freedman. Persistence and global stability in a predator–prey model consisting of three prey genotypes with fertility differences. Bulletin of Mathematical Biology, 48(5–6):469– 484, September 1986. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462318.

Freedman:1986:GST

[881] H. I. Freedman and K. Gopalsamy. Global stability in time-delayed single-species dynamics. Bulletin of Mathematical Biology, 48(5–6):485– 492, September 1986. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462319.

Freedman:1986:PPS

[882] H. I. Freedman and G. S. K. Wolkowicz. Predator-prey systems with group defence: The paradox of enrichment revisited. Bulletin of Mathematical Biology, 48(5–6):493–508, September 1986. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/article/10.1007/BF02462320.

Mehata:1986:CGS

[883] K. M. Mehata and D. Deivamoney Selvam. A class of general stochastic compartmental systems. Bulletin of Mathematical Biology, 48(5–6): 509–523, September 1986. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462321.

Bateman:1986:HOM

[884] J. M. Bateman. Heterogeneous organ models. Bulletin of Mathematical Biology, 48(5–6):525–543, September 1986. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02462322.

Lumsden:1986:RFL

[885] Charles J. Lumsden. Receptors and functional linkage in membrane permeability: a quantum mechanical model. Bulletin of Mathematical Biology, 48(5–6):545–567, September 1986. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02462323.

Lawrence:1986:OHS

[886] Charles B. Lawrence, Daniel A. Goldman, and Robert T. Hood. Optimized homology searches of the gene and protein sequence data banks. Bulletin of Mathematical Biology, 48(5–6):569–583, September 1986. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02462324.

Takeuchi:1986:DES

[887] Yasuhiro Takeuchi. Diffusion effect on stability of Lotka–Volterra models. Bulletin of Mathematical Biology, 48(5–6):585–601, September 1986. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02462325.

Altschul:1986:OSA

[888] Stephen F. Altschul and Bruce W. Erickson. Optimal sequence alignment using affine gap costs. Bulletin of Mathematical Biology, 48(5–6):

603–616, September 1986. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462326.

Altschul:1986:NMS

[889] Stephen F. Altschul and Bruce W. Erickson. A nonlinear measure of subalignment similarity and its significance levels. Bulletin of Mathematical Biology, 48(5–6):617–632, September 1986. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02462327.

Altschul:1986:LOS

[890] Stephen F. Altschul and Bruce W. Erickson. Locally optimal subalignments using nonlinear similarity functions. Bulletin of Mathematical Biology, 48(5–6):633–660, September 1986. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02462328.

Wassermann:1986:WNN

[891] G. D. Wassermann. The wiring-in of neural nets revisited. Bulletin of Mathematical Biology, 48(5–6):661–680, September 1986. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02462329.

Gopalsamy:1986:CRB

[892] K. Gopalsamy. Convergence in a resource-based competition system. Bulletin of Mathematical Biology, 48(5–6):681–699, September 1986. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02462330.

Rosen:1986:ASV

[893] Robert Rosen. Adaptability: The significance of variability from molecule to ecosystem. Bulletin of Mathematical Biology, 48(5–6):701– 703, September 1986. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462331.

Segel:1987:EA

[894] Lee Segel. Editorial announcement. Bulletin of Mathematical Biology, 49(1):i–iii, January 1987. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459954.

Anonymous:1987:ESSa

[895] Anonymous. Editorial: Software survey section. Bulletin of Mathematical Biology, 49(1):i–iv, January 1987. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02459964.

Lapidus:1987:F

[896] I. Richard Lapidus. The forum. Bulletin of Mathematical Biology, 49 (1):iii–iv, January 1987. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459955.

Tyson:1987:BR

[897] John J. Tyson. Book reviews. Bulletin of Mathematical Biology, 49 (1):iv, January 1987. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459956.

De:1987:SMP

[898] S. S. De. Stochastic model of population growth and spread. Bulletin of Mathematical Biology, 49(1):1–11, January 1987. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/article/10.1007/BF02459957.

Ehresmann:1987:HES

[899] A. C. Ehresmann and J.-P. Vanbremeersch. Hierarchical evolutive systems: a mathematical model for complex systems. Bulletin of Mathematical Biology, 49(1):13–50, January 1987. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02459958.

Liang:1987:BIP

[900] Z. Liang and H. Hart. Bayesian image processing of data from constrained source distributions — I. Non-valued, uncorrelated and correlated constraints. Bulletin of Mathematical Biology, 49(1):51– 74, January 1987. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459959.

Hart:1987:BIP

[901] H. Hart and Z. Liang. Bayesian image processing of data from constrained source distributions — II. Valued, uncorrelated and correlated constraints. Bulletin of Mathematical Biology, 49(1):75–91, January 1987. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459960.

Jurgensen:1987:IAD

[902] Helmut Jürgensen and Aristid Lindenmayer. Inference algorithms for developmental systems with cell lineages. Bulletin of Mathematical Biology, 49(1):93–123, January 1987. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459961.

Baishya:1987:NEF

[903] M. C. Baishya and C. G. Chakraborti. Non-equilibrium fluctuation in Volterra–Lotka systems. Bulletin of Mathematical Biology, 49(1): 125–131, January 1987. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459962.

Anonymous:1987:A

[904] Anonymous. Announcements. Bulletin of Mathematical Biology, 49(1): 133–134, January 1987. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459963.

Segel:1987:ESS

[905] Lee A. Segel. Editorial: Software survey section. Bulletin of Mathematical Biology, 49(2):i–iv, March 1987. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02459702.

Kliemann:1987:SDM

[906] Wolfgang Kliemann. A stochastic dynamical model for the characterization of the geometrical structure of dendritic processes. Bulletin of Mathematical Biology, 49(2):135–152, March 1987. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02459695.

Saidel:1987:TMA

[907] Gerald M. Saidel, Evan D. Morris, and Guy M. Chisolm III. Transport of macromolecules in arterial wall in vivo: a mathematical model and analytical solutions. Bulletin of Mathematical Biology, 49(2): 153–169, March 1987. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459696.

Torraang:1987:PAE

[908] Ingrid Torrång, Allan Gut, Lena Holm-Rutili, and Karl Johan Åbrink. A probabilistic approach for evaluating a technique for an indirect monitoring of the capillary blood flow to the gastric parietal cells. Bulletin of Mathematical Biology, 49(2):171–185, March 1987. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02459697.

Joshi:1987:RAD

[909] Rajani R. Joshi. On reliability analysis of DNA repair systems. Bulletin of Mathematical Biology, 49(2):187–216, March 1987. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02459698.

Sjontoft:1987:VDY

[910] Erik Sjøntoft and Carsten Edmund. In vivo determination of Young’s modulus for the human cornea. Bulletin of Mathematical Biology, 49 (2):217–232, March 1987. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459699.

Berding:1987:HRD

[911] Christoph Berding. On the heterogeneity of reaction–diffusion generated pattern. Bulletin of Mathematical Biology, 49(2):233–252, March 1987. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02459700.

Rosen:1987:TDP

[912] Gerald Rosen. Time delays produced by essential nonlinearity in population growth models. Bulletin of Mathematical Biology, 49(2): 253–255, March 1987. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459701.

Anonymous:1987:ESSb

[913] Anonymous. Editorial: Software survey section. Bulletin of Mathematical Biology, 49(3):i–iv, May 1987. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460127.

Misra:1987:SNL

[914] J. C. Misra and S. I. Singh. A study on the non-linear flow of blood through arteries. Bulletin of Mathematical Biology, 49(3):

257–277, May 1987. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460119.

Haldar:1987:OFB

[915] K. Haldar. Oscillatory flow of blood in a stenosed artery. Bulletin of Mathematical Biology, 49(3):279–287, May 1987. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/article/10.1007/BF02460120.

Erbay:1987:PWP

[916] H. A. Erbay, S. Erbay, and H. Demiray. Pulse waves in prestressed arteries. Bulletin of Mathematical Biology, 49(3):289–305, May 1987. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02460121.

Tarumi:1987:NTP

[917] Kazuaki Tarumi and Helmut Schwegler. A nonlinear treatment of the protocell model by a boundary layer approximation. Bulletin of Mathematical Biology, 49(3):307–320, May 1987. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02460122.

Aoki:1987:EBW

[918] Ichiro Aoki. Entropy balance of white-tailed deer during a winter night. Bulletin of Mathematical Biology, 49(3):321–327, May 1987. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02460123.

Gosselin:1987:KSC

[919] Richard P. Gosselin and Robert E. Gosselin. The kinetics of solute clearance by a single tissue capillary of limited permeability: Explicit solutions of two models and their bounds. Bulletin of Mathematical Biology, 49(3):329–349, May 1987. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460124.

Kamiya:1987:OCN

[920] Akira Kamiya, Shinya Takeda, and Masahiro Shibata. Optimum capillary number for oxygen delivery to tissue in man. Bulletin of Mathematical Biology, 49(3):351–361, May 1987. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02460125.

Verwer:1987:PET

[921] Ronald W. H. Verwer, Jaap Van Pelt, and André J. Noest. Parameter estimation in topological analysis of binary tree structures. Bulletin of Mathematical Biology, 49(3):363–378, May 1987. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02460126.

Anonymous:1987:ESSc

[922] Anonymous. Editorial: Software survey section. Bulletin of Mathematical Biology, 49(4):i–iv, July 1987. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02458868.

Landahl:1987:MSR

[923] H. D. Landahl, G. Gold, and G. M. Grodsky. Model of synthesis and release of insulin from rat islet β-cells and the effect of pretreatment with tolbutamide. Bulletin of Mathematical Biology, 49(4):379–394, July 1987. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02458857.

Luxon:1987:SLC

[924] Bruce A. Luxon. Solution of large compartmental models using numerical transform inversion. Bulletin of Mathematical Biology, 49(4): 395–402, July 1987. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458858.

Vayo:1987:MGS

[925] H. Westcott Vayo. Model geometries for sickled erythrocytes. Bulletin of Mathematical Biology, 49(4):403–411, July 1987. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/article/10.1007/BF02458859.

Horsfield:1987:SSG

[926] K. Horsfield, M. J. Woldenberg, and C. L. Bowes. Sequential and synchronous growth models related to vertex analysis and branching ratios. Bulletin of Mathematical Biology, 49(4):413–429, July 1987. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02458860.

Beretta:1987:GSS

[927] E. Beretta and Y. Takeuchi. Global stability of single-species diffusion Volterra models with continuous time delays. Bulletin of Mathematical Biology, 49(4):431–448, July 1987. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02458861.

Aoki:1987:EBD

[928] Ichiro Aoki. Entropy budgets of deciduous plant leaves and a theorem of oscillating entropy production. Bulletin of Mathematical Biology, 49 (4):449–460, July 1987. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458862.

Day:1987:CCI

[929] William H. E. Day. Computational complexity of inferring phylogenies from dissimilarity matrices. Bulletin of Mathematical Biology, 49 (4):461–467, July 1987. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458863.

Zeeberg:1987:ATF

[930] Barry R. Zeeberg and Henry N. Wagner, Jr. Analysis of three- and fourcompartment models for in vivo radioligand-neuroreceptor interaction. Bulletin of Mathematical Biology, 49(4):469–486, July 1987. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02458864.

Geurts:1987:FED

[931] Bernard J. Geurts and Frederik W. Wiegel. Fractional exponential decay in the capture of ligands by randomly distributed traps in one dimension. Bulletin of Mathematical Biology, 49(4):487–494, July 1987. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02458865.

Tuckwell:1987:LPG

[932] Henry C. Tuckwell and James A. Koziol. Logistic population growth under random dispersal. Bulletin of Mathematical Biology, 49(4): 495–506, July 1987. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458866.

McIntosh:1987:BR

[933] Robert P. McIntosh, Glenn Webb, James P. Keener, and Joan L. Aron. Book reviews. Bulletin of Mathematical Biology, 49(4): 507–517, July 1987. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458867.

Anonymous:1987:ESSd

[934] Anonymous. Editorial: Software survey section. Bulletin of Mathematical Biology, 49(5):i–iv, September 1987. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460142.

Almirantis:1987:MAM

[935] Y. Almirantis and G. Nicolis. Morphogenesis in an asymmetric medium. Bulletin of Mathematical Biology, 49(5):519–530, September 1987. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02460135.

Wagensberg:1987:EME

[936] Jorge Wagensberg and Joaquim Valls. The [extended] maximum entropy formalism and the statistical structure of ecosystems. Bulletin of Mathematical Biology, 49(5):531–538, September 1987. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02460136.

Reinhart:1987:TAE

[937] Heinrich Reinhart, Holzhütter Hermann-Georg, and Schuster Stefan. A theoretical approach to the evolution and structural design of enzymatic networks; linear enzymatic chains, branched pathways and glycolysis of erythrocytes. Bulletin of Mathematical Biology, 49(5):539– 595, September 1987. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460137.

Tracqui:1987:TST

[938] P. Tracqui, A. M. Perault-Staub, G. Milhaud, and J. F. Staub. Theoretical study of a two-dimensional autocatalytic model for calcium dynamics at the extracellular fluid-bone interface. Bulletin of Mathematical Biology, 49(5):597–613, September 1987. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02460138.

Kulenovic:1987:LOP

[939] M. R. S. Kulenović and G. Ladas. Linearized oscillations in population dynamics. Bulletin of Mathematical Biology, 49(5):615–627, September 1987. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460139.

Edelstein-Keshet:1987:ETG

[940] Leah Edelstein-Keshet. Evolution through group selection. Bulletin of Mathematical Biology, 49(5):629–647, September 1987. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02460140.

Anonymous:1987:ICM

[941] Anonymous. International conference on mathematical modelling in science and technology. Bulletin of Mathematical Biology, 49(5): 649, September 1987. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460141.

Anonymous:1987:ESSe

[942] Anonymous. Editorial: Software survey section. Bulletin of Mathematical Biology, 49(6):i–iv, November 1987. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02481775.

Klanchar:1987:MWF

[943] M. Klanchar and J. M. Tarbell. Modeling water flow through arterial tissue. Bulletin of Mathematical Biology, 49(6):651–669, November 1987. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02481766.

Roozen:1987:EES

[944] H. Roozen. Equilibrium and extinction in stochastic population dynamics. Bulletin of Mathematical Biology, 49(6):671–696, November 1987. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02481767.

Tyagi:1987:EAS

[945] V. P. Tyagi and M. Abbas. An exact analysis for a solute transport, due to simultaneous dialysis and ultrafiltration, in a hollow-fibre artificial kidney. Bulletin of Mathematical Biology, 49(6):697–717, November 1987. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02481768.

Landahl:1987:ASC

[946] H. D. Landahl. Approximate solution for capillary exchange models with and without axial mixing. Bulletin of Mathematical Biology, 49(6): 719–727, November 1987. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02481769.

Khan:1987:CTS

[947] Angshutosh Khan, Debajyoti Bhaumik, and Binayak Dutta-Roy. Charge trapping by solitions — a possible transport mechanism in macromolecular systems. Bulletin of Mathematical Biology, 49(6):729– 735, November 1987. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02481770.

Head:1987:FLT

[948] Tom Head. Formal language theory and DNA: an analysis of the generative capacity of specific recombinant behaviors. Bulletin of Mathematical Biology, 49(6):737–759, November 1987. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02481771.

Anonymous:1987:BR

[949] Anonymous. Book reviews. Bulletin of Mathematical Biology, 49(6): 761–776, November 1987. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02481772.

Scanlon:1987:F

[950] Jane Cronin Scanlon. The forum. Bulletin of Mathematical Biology, 49 (6):777, November 1987. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02481773.

Carson:1987:A

[951] E. R. Carson. Announcements. Bulletin of Mathematical Biology, 49(6): 779–780, November 1987. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02481774.

Segel:1988:E

[952] Lee A. Segel and Hugo M. Martinez. Editorial. Bulletin of Mathematical Biology, 50(1):i–ii, January 1988. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02459973.

Anonymous:1988:ESSa

[953] Anonymous. Editorial: Software survey section. Bulletin of Mathematical Biology, 50(1):i–iv, January 1988. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02459982.

Qi:1988:MSM

[954] An-Shen Qi. Multiple solutions of a model describing cancerous growth. Bulletin of Mathematical Biology, 50(1):1–17, January 1988. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02459974.

Illert:1988:HRS

[955] Christopher Roy Illert and Daniela Reverberi. Holography reveals the soft anatomy of ancient cephalopods. Bulletin of Mathematical Biology, 50(1):19–34, January 1988. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459975.

Shoucri:1988:MDP

[956] R. M. Shoucri. A mathematical derivation of the P–V relation at endsystole. Bulletin of Mathematical Biology, 50(1):35–41, January 1988. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02459976.

Skondras:1988:RCM

[957] Michael Skondras. RF-cell: a model for populations of randomly interconnected neurons. Bulletin of Mathematical Biology, 50(1):43– 66, January 1988. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459977.

Cull:1988:SDO

[958] Paul Cull. Stability of discrete one-dimensional population models. Bulletin of Mathematical Biology, 50(1):67–75, January 1988. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02459978.

Altschul:1988:SLB

[959] Stephen F. Altschul and Bruce W. Erickson. Significance levels for biological sequence comparison using non-linear similarity functions. Bulletin of Mathematical Biology, 50(1):77–92, January 1988. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02459979.

Gosselin:1988:KSC

[960] R. P. Gosselin and R. E. Gosselin. The kinetics of solute clearance by a single tissue capillary of limited permeability: Explicit solutions of two models and their bounds. Bulletin of Mathematical Biology, 50 (1):93–94, January 1988. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459980;   http://link.springer.com/content/pdf/10.
1007/BF02459980.pdf.

Anonymous:1988:GCT

[961] Anonymous. 1988 Gordon conference on theoretical biology and biomathematics. Bulletin of Mathematical Biology, 50(1):95, January 1988. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02459981.

Anonymous:1988:SSSa

[962] Anonymous. Software survey section. Bulletin of Mathematical Biology, 50(2):i–iv, March 1988. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459953.

Miller:1988:SCC

[963] Webb Miller and Eugene W. Myers. Sequence comparison with concave weighting functions. Bulletin of Mathematical Biology, 50(2): 97–120, March 1988. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459948.

Lewis:1988:SCC

[964] Harold W. Lewis III, Narendra S. Goel, and Richard L. Thompson. Simulation of cellular compaction and internalization in mammalian embryo development — II. Models for spherical embryos. Bulletin of Mathematical Biology, 50(2):121–142, March 1988. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02459949.

Ventriglia:1988:CSA

[965] F. Ventriglia. Computational simulation of activity of cortical-like neural systems. Bulletin of Mathematical Biology, 50(2):143–185, March 1988.

CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02459950.

Guess:1988:IGW

[966] Frank Guess and Matthew Witten. On the incompatibility of Gompertz or Weibull survival dynamics with exponentially distributed individual lifespans. Bulletin of Mathematical Biology, 50(2):187–192, March 1988. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02459951.

Winfree:1988:BR

[967] A. T. Winfree, R. M. Feldman, Masatoshi Nei, and David G. Stork. Book reviews. Bulletin of Mathematical Biology, 50(2):193–207, March 1988. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02459952.

Anonymous:1988:SSSb

[968] Anonymous. Software survey section. Bulletin of Mathematical Biology, 50(3):i–iv, May 1988. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458886.

Ferrante:1988:IDD

[969] L. Ferrante and G. Koch. Identifiability of DNA distribution from flow cytometric data with cell debris. Bulletin of Mathematical Biology, 50 (3):209–225, May 1988. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458881.

Rapp:1988:PDT

[970] P. E. Rapp, R. A. Latta, and A. I. Mees. Parameter-dependent transitions and the optimal control of dynamical diseases. Bulletin of Mathematical Biology, 50(3):227–253, May 1988. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02458882.

Toko:1988:PBP

[971] K. Toko, M. Nosaka, T. Fujiyoshi, K. Yamafuji, and K. Ogata. Periodic band pattern as a dissipative structure in ion transport systems with cylindrical shape. Bulletin of Mathematical Biology, 50(3): 255–288, May 1988. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458883.

Salathe:1988:MSC

[972] Eric P. Salathe. Mathematical studies of capillary-tissue exchange. Bulletin of Mathematical Biology, 50(3):289–311, May 1988. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02458884.

Belair:1988:BR

[973] Jacques Bélair, Leon Glass, Donald C. Mikulecky, Carl Hirschie Johnson, C. Van den Broeck, and Lev Ginzburg. Book reviews. Bulletin of Mathematical Biology, 50(3):313–328, May 1988. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02458885.

Anonymous:1988:SSSc

[974] Anonymous. Software survey section. Bulletin of Mathematical Biology, 50(4):i–iv, July 1988. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459710.

Pal:1988:CSM

[975] Dulal Pal, N. Rudraiah, and Rathna Devanathan. A couple stress model of blood flow in the microcirculation. Bulletin of Mathematical Biology, 50(4):329–344, July 1988. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459703.

Rosen:1988:HSS

[976] Gerald Rosen. HIV spread in the San Francisco cohort: Scaling of the effective logistic rate for seropositivity. Bulletin of Mathematical Biology, 50(4):345–349, July 1988. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459704.

Li:1988:PDA

[977] Jia Li. Persistence in discrete age-structured population models. Bulletin of Mathematical Biology, 50(4):351–366, July 1988. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02459705.

Balaz:1988:SDC

[978] Š. Baláž, E. Šturdík, and J. Augustín. Subcellular distribution of compounds in biosystems. Bulletin of Mathematical Biology, 50(4):

367–378, July 1988. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459706.

Wollkind:1988:MTD

[979] David J. Wollkind, John B. Collings, and Jesse A. Logan. Metastability in a temperature-dependent model system for predator–prey mite outbreak interactions on fruit trees. Bulletin of Mathematical Biology, 50(4):379–409, July 1988. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459707.

Ermentrout:1988:BRa

[980] G. B. Ermentrout and Kamil A. Grajski. Book reviews. Bulletin of Mathematical Biology, 50(4):411–419, July 1988. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02459708.

Anonymous:1988:A

[981] Anonymous. Announcements. Bulletin of Mathematical Biology, 50 (4):421–422, July 1988. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459709.

Anonymous:1988:SSSd

[982] Anonymous. Software survey section. Bulletin of Mathematical Biology, 50(5):i–iv, September 1988. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458856.

Zeeberg:1988:TER

[983] Barry R. Zeebèrg, Ralph C. Reid, Katherine A. Murphy, and Richard C. Reba. Theoretical effects of radioligand diffusional gradients and microscopic neuroreceptor distribution in in vivo kinetic studies. Bulletin of Mathematical Biology, 50(5):423–444, September 1988. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02458845.

Wagensberg:1988:BAM

[984] J. Wagensberg, J. Valls, and J. Bermudez. Biological adaptation and the mathematical theory of information. Bulletin of Mathematical Biology, 50(5):445–464, September 1988. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02458846.

Samardzija:1988:ERC

[985] Nikola Samardzija and Larry D. Greller. Explosive route to chaos through a fractal torus in a generalized Lotka–Volterra model. Bulletin of Mathematical Biology, 50(5):465–491, September 1988. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02458847.

McKee:1988:SMM

[986] S. McKee and G. W. Gould. A simple mathematical model of the thermal death of microorganisms. Bulletin of Mathematical Biology, 50(5): 493–501, September 1988. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458848.

Akcakaya:1988:TPD

[987] H. R. Akçakaya, L. R. Ginzburg, D. Slice, and L. B. Slobodkin. The theory of population dynamics — II. Physiological delays. Bulletin of Mathematical Biology, 50(5):503–515, September 1988. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02458849.

Freedman:1988:ILP

[988] H. I. Freedman and Quan Hongshun. Interactions leading to persistence in predator–prey systems with group defence. Bulletin of Mathematical Biology, 50(5):517–530, September 1988. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02458850.

Aldis:1988:ULD

[989] Geoffrey K. Aldis. The unstirred layer during osmotic flow into a tubule. Bulletin of Mathematical Biology, 50(5):531–545, September 1988. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02458851.

Aldis:1988:OPT

[990] Geoffrey K. Aldis. The osmotic permeability of a tubule wall. Bulletin of Mathematical Biology, 50(5):547–558, September 1988. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02458852.

Nan:1988:FDE

[991] Xu Nan and Xu Jinghua. The fractal dimension of EEG as a physical measure of conscious human brain activities. Bulletin of Mathematical Biology, 50(5):559–565, September 1988. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02458853.

Ermentrout:1988:BRb

[992] G. B. Ermentrout. Book reviews. Bulletin of Mathematical Biology, 50(5):567–576, September 1988. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02458854.

Anonymous:1988:CIT

[993] Anonymous. Critical issues in tumor microcirculation, angiogenesis and metastasis. Bulletin of Mathematical Biology, 50(5):577, September 1988. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02458855.

Anonymous:1988:ESSb

[994] Anonymous. Editorial: Software survey section. Bulletin of Mathematical Biology, 50(6):i–iv, November 1988. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460099.

Segel:1988:VSS

[995] Lee A. Segel. On the validity of the steady state assumption of enzyme kinetics. Bulletin of Mathematical Biology, 50(6):579–593, November 1988. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02460092.

Clarke:1988:ECE

[996] Bertrand S. Clarke, Jay E. Mittenthal, and Phillip A. Arcuri. An extremal criterion for epimorphic regeneration. Bulletin of Mathematical Biology, 50(6):595–634, November 1988. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460093.

Schuster:1988:SMD

[997] Peter Schuster and Jörg Swetina. Stationary mutant distributions and evolutionary optimization. Bulletin of Mathematical Biology, 50(6):635– 660, November 1988. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460094.

Barenholz-Paniry:1988:RSE

[998] Vered Barenholz-Paniry, Jacob S. Ishay, and Zvi Grossman. Rhythmic signalling and entrainment in Vespa orientalis larvae: Characterization of the underlying interactions. Bulletin of Mathematical Biology, 50(6): 661–679, November 1988. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460095.

Gyori:1988:TDS

[999] Istvan Gyori, Seth Michelson, and John Leith. Time-dependent subpopulation induction in heterogeneous tumors. Bulletin of Mathematical Biology, 50(6):681–696, November 1988. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460096.

Jacquez:1988:BR

[1000] John A. Jacquez and J. H. M. Thornley. Book reviews. Bulletin of Mathematical Biology, 50(6):697–700, November 1988. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02460097.

Anonymous:1988:SMB

[1001] Anonymous. The Society for Mathematical Biology is pleased to announce an Annual Best Paper Award of $1000. Bulletin of Mathematical Biology, 50(6):701, November 1988. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460098.

Anonymous:1989:ESSa

[1002] Anonymous. Editorial: Software survey section. Bulletin of Mathematical Biology, 51(1):i–iv, January 1989. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02458844.

Waterman:1989:F

[1003] Michael S. Waterman. Foreword. Bulletin of Mathematical Biology, 51 (1):1–4, January 1989. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458833.

Myers:1989:AMR

[1004] Eugene W. Myers and Webb Miller. Approximate matching of regular expressions. Bulletin of Mathematical Biology, 51(1):5–37, January 1989.

CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02458834.

Auger:1989:AOI

[1005] Ivan E. Auger and Charles E. Lawrence. Algorithms for the optimal identification of segment neighborhoods. Bulletin of Mathematical Biology, 51(1):39–54, January 1989. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458835.

Zuker:1989:APS

[1006] M. Zuker and R. L. Somorjai. The alignment of protein structures in three dimensions. Bulletin of Mathematical Biology, 51(1): 55–78, January 1989. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458836.

Churchill:1989:SMH

[1007] Gary A. Churchill. Stochastic models for heterogeneous DNA sequences. Bulletin of Mathematical Biology, 51(1):79–94, January 1989. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02458837.

Tavare:1989:CPP

[1008] Simon Tavaré and Brenda Song. Codon preference and primary sequence structure in protein-coding regions. Bulletin of Mathematical Biology, 51 (1):95–115, January 1989. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458838.

Sankoff:1989:PMG

[1009] David Sankoff and Martin Goldstein. Probabilistic models of genome shuffling. Bulletin of Mathematical Biology, 51(1):117–124, January 1989. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02458839.

Arratia:1989:TLD

[1010] R. Arratia and L. Gordon. Tutorial on large deviations for the binomial distribution. Bulletin of Mathematical Biology, 51(1):125– 131, January 1989. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458840.

Bandelt:1989:WHA

[1011] H.-J. Bandelt and A. W. M. Dress. Weak hierarchies associated with similarity measures — an additive clustering technique. Bulletin of Mathematical Biology, 51(1):133–166, January 1989. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02458841.

Ferretti:1989:CAR

[1012] Vincent Ferretti and David Sankoff. A continuous analog for RNA folding. Bulletin of Mathematical Biology, 51(1):167–171, January 1989. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02458842.

Easthope:1989:CEA

[1013] Peter L. Easthope and Timothy F. Havel. Computational experience with an algorithm for tetrangle inequality bound smoothing. Bulletin of Mathematical Biology, 51(1):173–194, January 1989. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02458843.

Anonymous:1989:ESSb

[1014] Anonymous. Editorial: Software survey section. Bulletin of Mathematical Biology, 51(2):i–iv, March 1989. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02458449.

Mizraji:1989:CDA

[1015] Eduardo Mizraji. Context-dependent associations in linear distributed memories. Bulletin of Mathematical Biology, 51(2):195–205, March 1989. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02458441.

Tarumi:1989:WSM

[1016] Kazuaki Tarumi and Ekkehard Mueller. Wavelength selection mechanism in the Gierer–Meinhardt model. Bulletin of Mathematical Biology, 51 (2):207–216, March 1989. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458442.

DeBoer:1989:SSI

[1017] Rob J. De Boer and Pauline Hogeweg. Stability of symmetric idiotypic networks — a critique of Hoffmann’s analysis. Bulletin of Mathematical Biology, 51(2):217–222, March 1989. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02458443.

DeBoer:1989:MNS

[1018] Rob J. De Boer and Pauline Hogeweg. Memory but no suppression in low-dimensional symmetric idiotypic networks. Bulletin of Mathematical Biology, 51(2):223–246, March 1989. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02458444.

Atlan:1989:ANT

[1019] Henri Atlan. Automata network theories in immunology: Their utility and their underdetermination. Bulletin of Mathematical Biology, 51 (2):247–253, March 1989. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458445.

Dvorak:1989:AMS

[1020] I. Dvořák and J. Šiška. Analysis of metabolic systems with complex slow and fast dynamics. Bulletin of Mathematical Biology, 51(2): 255–274, March 1989. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458446.

Superber:1989:APC

[1021] Göran O. Superber. An alternative procedure for calculating glucose consumption from 2-deoxyglucose uptake. Bulletin of Mathematical Biology, 51(2):275–286, March 1989. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458447.

Hoppensteadt:1989:BR

[1022] F. C. Hoppensteadt, John Ross, and Bruce Wallace. Book reviews. Bulletin of Mathematical Biology, 51(2):287–291, March 1989. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02458448.

Anonymous:1989:ESSc

[1023] Anonymous. Editorial: Software survey section. Bulletin of Mathematical Biology, 51(3):i–iv, May 1989. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460118.

Lam:1989:AIM

[1024] Gabriel K. Y. Lam. Analysis of interaction for mixtures of agents using the linear isobole. Bulletin of Mathematical Biology, 51(3): 293–309, May 1989. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460109.

Zhien:1989:TSS

[1025] Ma Zhien, Song Baojun, and Thomas G. Hallam. The threshold of survival for systems in a fluctuating environment. Bulletin of Mathematical Biology, 51(3):311–323, May 1989. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460110.

Regalbuto:1989:ULB

[1026] Mónica C. Regalbuto, William Strieder, and Arvind Varma. Upper and lower bounds from the maximum principle. Intracellular diffusion with Michaelis–Menten kinetics. Bulletin of Mathematical Biology, 51 (3):325–335, May 1989. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460111.

Yeshurun:1989:PLN

[1027] Y. Yeshurun, Z. Wollberg, and N. Dyn. Prediction of linear and nonlinear responses of MGB neurons by system identification methods. Bulletin of Mathematical Biology, 51(3):337–346, May 1989. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02460112.

Norwich:1989:IAR

[1028] K. H. Norwich, C. N. L. Seburn, and E. Axelrad. An informational approach to reaction times. Bulletin of Mathematical Biology, 51(3): 347–358, May 1989. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460113.

Geurts:1989:DLI

[1029] Bernard J. Geurts. Diffusion limited immunochemical sensing. Bulletin of Mathematical Biology, 51(3):359–379, May 1989. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02460114.

DeBoer:1989:UIR

[1030] Rob J. De Boer and Pauline Hogeweg. Unreasonable implications of reasonable idiotypic network assumptions. Bulletin of Mathematical Biology, 51(3):381–408, May 1989. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460115.

Schaffer:1989:PSD

[1031] W. M. Schaffer and G. A. Fox. Plant strategies and the dynamics and structure of plant communities. Bulletin of Mathematical Biology, 51 (3):409–411, May 1989. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460116.

Anonymous:1989:TWC

[1032] Anonymous. The third world conference on mathematics at the service of man. Bulletin of Mathematical Biology, 51(3):413, May 1989. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02460117.

Anonymous:1989:ESSd

[1033] Anonymous. Editorial: Software survey section. Bulletin of Mathematical Biology, 51(4):i–iv, July 1989. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460091.

Segel:1989:EV

[1034] Lee A. Segel. Editorial: Viewpoint. Bulletin of Mathematical Biology, 51(4):415, July 1989. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460080.

Trifonov:1989:MCN

[1035] E. N. Trifonov. The multiple codes of nucleotide sequences. Bulletin of Mathematical Biology, 51(4):417–432, July 1989. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/article/10.1007/BF02460081.

Horvath:1989:DBL

[1036] Gábor Horváth. Description of the birch leaf roller’s incisions for different leaves. Bulletin of Mathematical Biology, 51(4):433–447, July 1989. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02460082.

Feng:1989:VDM

[1037] Si shen Feng, Richard Skalak, and Shu Chien. Velocity distribution on the membrane of a tank-treading red blood cell. Bulletin of Mathematical Biology, 51(4):449–465, July 1989. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460083.

Sandberg:1989:MFS

[1038] Sonja Sandberg and Tamara E. Awerbuch. Mathematical formulation and studies of the risk parameters involved in HIV transmission. Bulletin of Mathematical Biology, 51(4):467–474, July 1989. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02460084.

Shoucri:1989:SAP

[1039] Rachad M. Shoucri, Jacques R. Rouleau, John G. Kingma, Jr., Jean G. Dumesnil, and Gilles R. Dagenais. Some applications of the P–V relation to the study of left ventricular performance. Bulletin of Mathematical Biology, 51(4):475–484, July 1989. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02460085.

Umeda:1989:MMC

[1040] Tamiki Umeda. A mathematical model for cell sorting, migration and shape in the slug stage of Dictyostelium discoideum. Bulletin of Mathematical Biology, 51(4):485–500, July 1989. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02460086.

Nakajima:1989:RLS

[1041] H. Nakajima and D. L. DeAngelis. Resilience and local stability in a nutrient-limited resource-consumer system. Bulletin of Mathematical Biology, 51(4):501–510, July 1989. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460087.

vanPelt:1989:COD

[1042] J. van Pelt, R. W. H. Verwer, and H. B. M. Uylings. Centrifugal-order distributions in binary topological trees. Bulletin of Mathematical Biology, 51(4):511–536, July 1989. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460088.

Ermentrout:1989:BR

[1043] G. B. Ermentrout, Steven H. Strogatz, and Jeffrey C. Long. Book reviews. Bulletin of Mathematical Biology, 51(4):537–544, July 1989. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02460089.

Anonymous:1989:A

[1044] Anonymous. Announcement. Bulletin of Mathematical Biology, 51 (4):545–546, July 1989. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460090.

Anonymous:1989:ESSe

[1045] Anonymous. Editorial: Software survey section. Bulletin of Mathematical Biology, 51(5):i–iv, September 1989. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02459972.

Segel:1989:E

[1046] Lee A. Segel. Editorial. Bulletin of Mathematical Biology, 51(5): 547, September 1989. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459965.

Yates:1989:CDM

[1047] Kemble Yates and Edward Pate. A cascading development model for amphibian embryos. Bulletin of Mathematical Biology, 51(5):549– 578, September 1989. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459966.

Wise:1989:TES

[1048] M. E. Wise and G. J. J. M. Borsboom. Two exceptional sets of physiological clearance curves and their mathematical form: Test cases? Bulletin of Mathematical Biology, 51(5):579–596, September 1989. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02459967.

Hein:1989:OAR

[1049] Jotun J. Hein. An optimal algorithm to reconstruct trees from additive distance data. Bulletin of Mathematical Biology, 51(5):597– 603, September 1989. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459968.

Longtin:1989:MAO

[1050] André Longtin and John G. Milton. Modelling autonomous oscillations in the human pupil light reflex using non-linear delay-differential equations. Bulletin of Mathematical Biology, 51(5):605–624, September 1989. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02459969.

Peters:1989:SDP

[1051] Craig Steven Peters, Marc Mangel, and R. F. Costantino. Stationary distribution of population size in tribolium. Bulletin of Mathematical Biology, 51(5):625–638, September 1989. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02459970.

Kaplan:1989:BR

[1052] Norman Kaplan, Michael A. Savageau, Ethan Akin, John J. Tyson, Bruce Wallace, and Robert Rosen. Book reviews. Bulletin of Mathematical Biology, 51(5):639–655, September 1989. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02459971.

Anonymous:1989:ESSf

[1053] Anonymous. Editorial: Software survey section. Bulletin of Mathematical Biology, 51(6):i–iv, November 1989. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02459662.

Bookstein:1989:MMB

[1054] Fred L. Bookstein and Richard A. Reyment. Microevolution in Miocene Brizalina (foraminifera) studied by canonical variate analysis and analysis of landmarks. Bulletin of Mathematical Biology, 51(6):657– 679, November 1989. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459654.

Thompson:1989:LPP

[1055] Colin J. Thompson and J. B. McGuire. On limits to perception and pattern recognition. Bulletin of Mathematical Biology, 51(6):681– 686, November 1989. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459655.

Cushing:1989:EMD

[1056] J. M. Cushing and Jia Li. On Ebenman’s model for the dynamics of a population with competing juveniles and adults. Bulletin of Mathematical Biology, 51(6):687–713, November 1989. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02459656.

Spouge:1989:ARQ

[1057] John L. Spouge, Scott P. Layne, and Micah Dembo. Analytic results for quantifying HIV infectivity. Bulletin of Mathematical Biology, 51(6): 715–730, November 1989. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459657.

Michelson:1989:SMS

[1058] S. Michelson, K. Ito, H. T. Tran, and J. T. Leith. Stochastic models for subpopulation emergence in heterogeneous tumors. Bulletin of Mathematical Biology, 51(6):731–747, November 1989. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02459658.

Sneyd:1989:PCI

[1059] James Sneyd and Daniel Tranchina. Phototransduction in cones: an inverse problem in enzyme kinetics. Bulletin of Mathematical Biology, 51(6):749–784, November 1989. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02459659.

Gutierrez:1989:TPS

[1060] Emilia Gutierrez and Helena Almirall. Temporal properties of some biological systems and their fractal attractors. Bulletin of Mathematical Biology, 51(6):785–800, November 1989. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02459660.

Hunt:1989:BR

[1061] Fern Hunt, G. B. Ermentrout, Layne T. Watson, and Marek Kimmel. Book reviews. Bulletin of Mathematical Biology, 51(6):801– 809, November 1989. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459661.

Anonymous:1990:ESSa

[1062] Anonymous. Editorial: Software survey section. Bulletin of Mathematical Biology, 52(1–2):i–iv, January 1990. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02459579.

Mangel:1990:P

[1063] Marc Mangel. Preface. Bulletin of Mathematical Biology, 52(1– 2):1, January 1990. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459566.

Rinzel:1990:DEE

[1064] John Rinzel. Discussion: Electrical excitability of cells, theory and experiment: Review of the Hodgkin–Huxley foundation and an update. Bulletin of Mathematical Biology, 52(1–2):3–23, January 1990. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02459567.

Hodgkin:1990:QDM

[1065] A. L. Hodgkin and A. F. Huxley. A quantitative description of membrane current and its application to conduction and excitation in nerve. Bulletin of Mathematical Biology, 52(1–2):25–71, January 1990. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02459568.

Cowan:1990:DMP

[1066] Jack D. Cowan. Discussion: McCulloch–Pitts and related neural nets from 1943 to 1989. Bulletin of Mathematical Biology, 52(1–2): 73–97, January 1990. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459569.

McCulloch:1990:LCI

[1067] Warren S. McCulloch and Walter Pitts. A logical calculus of the ideas immanent in nervous activity. Bulletin of Mathematical Biology, 52(1– 2):99–115, January 1990. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459570.

Murray:1990:DTT

[1068] J. D. Murray. Discussion: Turing’s theory of morphogenesis — its influence on modelling biological pattern and form. Bulletin of Mathematical Biology, 52(1–2):117–152, January 1990. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02459571.

Turing:1990:CBM

[1069] A. M. Turing. The chemical basis of morphogenesis. Bulletin of Mathematical Biology, 52(1–2):153–197, January 1990. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/article/10.1007/BF02459572. Reprinted from Philo-

sophical Transactions of the Royal Society (Part B), Vol 237, Pg 37–72, 1953 [3627].

Provine:1990:DPG

[1070] William B. Provine. Discussion: Population genetics. Bulletin of Mathematical Biology, 52(1–2):199–207, January 1990. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/article/10.1007/BF02459573.

Haldane:1990:MTN

[1071] J. B. S. Haldane. A mathematical theory of natural and artificial selection — I. Bulletin of Mathematical Biology, 52(1–2):209–240, January 1990. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02459574.

Wright:1990:EMP

[1072] Sewall Wright. Evolution in Mendelian populations. Bulletin of Mathematical Biology, 52(1–2):241–295, January 1990. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/article/10.1007/BF02459575.

Fisher:1990:DR

[1073] R. A. Fisher. On the dominance ratio. Bulletin of Mathematical Biology, 52(1–2):297–318, January 1990. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02459576.

Lauffenburger:1990:BR

[1074] Douglas A. Lauffenburger, R. Allen White, Donald L. DeAngelis, George Kampis, Eberhard O. Voit, Steven Vogel, Marc Mangel, and Markus Loeffler. Book reviews. Bulletin of Mathematical Biology, 52(1–2): 319–334, January 1990. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459577.

Anonymous:1990:SAM

[1075] Anonymous. SIAM 1990 Annual Meeting and Short Course 15–20 July 1990 Hyatt Regency Hotel, Chicago, IL. Bulletin of Mathematical Biology, 52(1–2):335–337, January 1990. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02459578.

Anonymous:1990:ESSb

[1076] Anonymous. Editorial: Software survey section. Bulletin of Mathematical Biology, 52(3):i–iv, May 1990. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02458584.

Atlan:1990:CCD

[1077] Henri Atlan and Moshe Koppel. The cellular computer DNA: Program or data. Bulletin of Mathematical Biology, 52(3):335–348, May 1990. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02458575.

Rinsma:1990:DNM

[1078] Ingrid Rinsma, Michael Hendy, and David Penny. Distribution of the number of matches between nucleotide sequences. Bulletin of Mathematical Biology, 52(3):349–358, May 1990. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02458576.

Gotoh:1990:OSA

[1079] Osamu Gotoh. Optimal sequence alignment allowing for long gaps. Bulletin of Mathematical Biology, 52(3):359–373, May 1990. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02458577.

Jansen:1990:GCS

[1080] V. A. A. Jansen, R. M. Nisbet, and W. S. C. Gurney. Generation cycles in stage structured populations. Bulletin of Mathematical Biology, 52 (3):375–396, May 1990. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458578.

Ventriglia:1990:ACL

[1081] F. Ventriglia. Activity in cortical-like neural systems: Short-range effects and attention phenomena. Bulletin of Mathematical Biology, 52 (3):397–429, May 1990. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458579.

Allison:1990:MML

[1082] L. Allison and C. N. Yee. Minimum message length encoding and the comparison of macromolecules. Bulletin of Mathematical Biology, 52 (3):431–453, May 1990. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458580.

Tang:1990:EFC

[1083] Betty Tang and Michael S. Waterman. The expected fraction of clonable genomic DNA. Bulletin of Mathematical Biology, 52(3): 455–475, May 1990. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458581.

Rosen:1990:BR

[1084] Robert Rosen, W. Horsthemke, and Calvin J. Ribbens. Book reviews. Bulletin of Mathematical Biology, 52(3):477–482, May 1990. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02458582.

Anonymous:1990:A

[1085] Anonymous. Announcements. Bulletin of Mathematical Biology, 52 (3):483–484, May 1990. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458583.

Anonymous:1990:ESSc

[1086] Anonymous. Editorial: Software survey section. Bulletin of Mathematical Biology, 52(4):i–iv, July 1990. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02462269.

Stadler:1990:DSA

[1087] Peter F. Stadler and Peter Schuster. Dynamics of small autocatalytic reaction networks — I. Bifurcations, permanence and exclusion. Bulletin of Mathematical Biology, 52(4):485–508, July 1990. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02462263.

Gotoh:1990:COS

[1088] Osamu Gotoh. Consistency of optimal sequence alignments. Bulletin of Mathematical Biology, 52(4):509–525, July 1990. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/article/10.1007/BF02462264.

Klebaner:1990:CUG

[1089] Fima C. Klebaner. Conditions for the unlimited growth in multitype population size dependent Galton–Watson processes. Bulletin of Mathematical Biology, 52(4):527–534, July 1990. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02462265.

Grasman:1990:DMC

[1090] Johan Grasman. A deterministic model of the cell cycle. Bulletin of Mathematical Biology, 52(4):535–547, July 1990. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/article/10.1007/BF02462266.

Adam:1990:DRG

[1091] J. A. Adam and S. A. Maggelakis. Diffusion regulated growth characteristics of a spherical prevascular carcinoma. Bulletin of Mathematical Biology, 52(4):549–582, July 1990. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02462267.

Goodman:1990:BR

[1092] Daniel Goodman, William J. Reed, L. David Roper, and John J. Tyson. Book reviews. Bulletin of Mathematical Biology, 52(4): 583–596, July 1990. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462268.

Anonymous:1990:ESSd

[1093] Anonymous. Editorial: Software survey section. Bulletin of Mathematical Biology, 52(5):i–iv, September 1990. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02462108.

Franks:1990:SBW

[1094] Nigel R. Franks, Steve Bryant, Richard Griffiths, and Lia Hemerik. Synchronization of the behaviour within nests of the ant Leptothorax acervorum (Fabricius) — I. Discovering the phenomenon and its relation to the level of starvation. Bulletin of Mathematical Biology, 52(5):597– 612, September 1990. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462100.

Hemerik:1990:SBW

[1095] Lia Hemerik, Nick F. Britton, and N. R. Franks. Synchronization of the behaviour within nests of the ant Leptothorax acervorum (Fabricius)

— II. Modelling the phenomenon and predictions from the model. Bulletin of Mathematical Biology, 52(5):613–628, September 1990. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02462101.

Verwer:1990:ABT

[1096] Ronald W. H. Verwer and Jaap Van Pelt. Analysis of binary trees when occasional multifurcations can be considered as aggregates of bifurcations. Bulletin of Mathematical Biology, 52(5):629–641, September 1990. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02462102.

Oristland:1990:SSB

[1097] Nils A. Øristland. Starvation survival and body composition in mammals with particular reference to Homo sapiens. Bulletin of Mathematical Biology, 52(5):643–655, September 1990. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02462103.

Voit:1990:CIR

[1098] Eberhard O. Voit and Pon N. Yi. Comparison of isoeffect relationships in radiotherapy. Bulletin of Mathematical Biology, 52(5):657– 675, September 1990. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462104.

Busenberg:1990:DMP

[1099] Stavros Busenberg, S. Kishore Kumar, Paul Austin, and Graeme Wake. The dynamics of a model of a plankton-nutrient interaction. Bulletin of Mathematical Biology, 52(5):677–696, September 1990. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02462105.

Tyson:1990:BR

[1100] John J. Tyson, Heike Schuster, Martin Gerhardt, Prem Ponka, Greg Krukonis, Susanne Schwinning, and James P. Keener. Book reviews.

Bulletin of Mathematical Biology, 52(5):697–708, September 1990. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02462106.

Anonymous:1990:ECM

[1101] Anonymous. 1st European conference on mathematics applied to biology and medicine. Bulletin of Mathematical Biology, 52(5):709, September 1990. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462107.

Anonymous:1990:ESSe

[1102] Anonymous. Editorial: Software survey section. Bulletin of Mathematical Biology, 52(6):i–iv, November 1990. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460811.

Neumann:1990:EDM

[1103] Steven J. Neumann, Scott A. Berceli, Eva M. Sevick, A. Michael Lincoff, ViJay S. Warty, Arthur M. Brant, Ira M. Herman, and Harvey S. Borovetz. Experimental determination and mathematical model of the transient incorporation of cholesterol in the arterial wall. Bulletin of Mathematical Biology, 52(6):711–732, November 1990. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02460805.

Wagensberg:1990:CIT

[1104] Jorge Wagensberg, Ambrosio Garcia, and Ricard V. Sole. Connectivity and information transfer in flow networks: Two magic numbers in ecology? Bulletin of Mathematical Biology, 52(6):733–740, November 1990. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02460806.

Arques:1990:MDS

[1105] Didier G. Arquès and Christian J. Michel. A model of DNA sequence evolution. Bulletin of Mathematical Biology, 52(6):741–772, November 1990. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02460807.

Mott:1990:AAD

[1106] R. F. Mott, T. B. L. Kirkwood, and R. N. Curnow. An accurate approximation to the distribution of the length of the longest matching word between two random DNA sequences. Bulletin of Mathematical Biology, 52(6):773–784, November 1990. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460808.

Solimano:1990:NTC

[1107] F. Solimano, G. I. Bischi, M. Bianchi, L. Rossi, and M. Magnani. A nonlinear three-compartiment model for the administration of 2’,3’- dideoxycytidine by using red blood cells as bioreactors. Bulletin of Mathematical Biology, 52(6):785–796, November 1990. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02460809.

Edelstein-Keshet:1990:BR

[1108] Leah Edelstein-Keshet, Vassiliki Betty Smocovitis, and Joel Keizer. Book reviews. Bulletin of Mathematical Biology, 52(6):797–805, November 1990. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460810.

Anonymous:1991:ESS

[1109] Anonymous. Editorial: Software survey section. Bulletin of Mathematical Biology, 53(1–2):i–iii, March 1991. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02464436.

Anderson:1991:DKM

[1110] Roy M. Anderson. Discussion: The Kermack–McKendrick epidemic threshold theorem. Bulletin of Mathematical Biology, 53(1–2):1–32, March 1991. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02464422.

Kermack:1991:CMTa

[1111] W. O. Kermack and A. G. McKendrick. Contributions to the mathematical theory of epidemics — I. Bulletin of Mathematical Biology, 53 (1–2):33–55, March 1991. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02464423.

Kermack:1991:CMTb

[1112] W. O. Kermack and A. G. McKendrick. Contributions to the mathematical theory of epidemics — II. The problem of endemicity. Bulletin of Mathematical Biology, 53(1–2):57–87, March 1991. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02464424.

Kermack:1991:CMTc

[1113] W. O. Kermack and A. G. McKendrick. Contributions to the mathematical theory of epidemics — III. Further studies of the problem of endemicity. Bulletin of Mathematical Biology, 53(1–2):89–118, March 1991. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02464425.

Toft:1991:DIE

[1114] Catherine A. Toft and Marc Mangel. Discussion: From individuals to ecosystems; the papers of Skellam, Lindeman and Hutchinson. Bulletin of Mathematical Biology, 53(1–2):119–134, March 1991. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02464426.

Skellam:1991:RDT

[1115] J. G. Skellam. Random dispersal in theoretical populations. Bulletin of Mathematical Biology, 53(1–2):135–165, March 1991. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02464427.

Lindeman:1991:TDA

[1116] Raymond L. Lindeman. The trophic-dynamic aspect of ecology. Bulletin of Mathematical Biology, 53(1–2):167–191, March 1991. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02464428.

Hutchinson:1991:PSA

[1117] G. Evelyn Hutchinson. Population studies: Animal ecology and demography. Bulletin of Mathematical Biology, 53(1–2):193–213, March 1991. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02464429.

Reed:1991:DBE

[1118] William J. Reed. Discussion: Bioeconomics. An essay on the classic papers of H. Scott Gordon, Milner B. Schaefer and Harold Hotelling. Bulletin of Mathematical Biology, 53(1–2):215–229, March 1991. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02464430.

Gordon:1991:ETC

[1119] H. Scott Gordon. The economic theory of a common-property resource: The fishery. Bulletin of Mathematical Biology, 53(1–2):231– 252, March 1991. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02464431.

Schaefer:1991:SAD

[1120] Milner B. Schaefer. Some aspects of the dynamics of populations important to the management of the commercial marine fisheries. Bulletin of Mathematical Biology, 53(1–2):253–279, March 1991. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02464432.

Hotelling:1991:EER

[1121] Harold Hotelling. The economics of exhaustible resources. Bulletin of Mathematical Biology, 53(1–2):281–312, March 1991. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02464433.

Nemethy:1991:BR

[1122] George Némethy, Michael Conrad, Michael Ruse, Robert Rosen, and G. B. Ermentrout. Book reviews. Bulletin of Mathematical Biology, 53 (1–2):313–326, March 1991. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02464434.

Anonymous:1991:RMB

[1123] Anonymous. Resource management and behavioural ecology: a Conference in Celebration of the 60th Birthday of Colin Clark, 24–26 July 1991, University of British Columbia, Vancouver, B.C. Bulletin of Mathematical Biology, 53(1–2):327, March 1991. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02464435.

Jaffe:1991:AMC

[1124] Steven Jaffe. Autocatalytic membrane conductance and memory. Bulletin of Mathematical Biology, 53(3):327–343, May 1991. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02460721.

Luo:1991:SCN

[1125] Liaofu Luo and Hong Li. The statistical correlation of nucleotides in protein-coding DNA sequences. Bulletin of Mathematical Biology, 53 (3):345–353, May 1991. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460722.

Derrida:1991:EFF

[1126] Bernard Derrida and Luca Peliti. Evolution in a flat fitness landscape. Bulletin of Mathematical Biology, 53(3):355–382, May 1991. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02460723.

Thron:1991:SCIa

[1127] C. D. Thron. The secant condition for instability in biochemical feedback control — I. The role of cooperativity and saturability. Bulletin of Mathematical Biology, 53(3):383–401, May 1991. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02460724.

Thron:1991:SCIb

[1128] C. D. Thron. The secant condition for instability in biochemical feedback control — II. Models with upper Hessenberg Jacobian matrices. Bulletin of Mathematical Biology, 53(3):403–424, May 1991. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02460725.

Horvath:1991:SAV

[1129] Gábor Horváth and Dezsö Varjú. On the structure of the aerial visual field of aquatic animals distorted by refraction. Bulletin of Mathematical Biology, 53(3):425–441, May 1991. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460726.

Liebovitch:1991:DAE

[1130] Larry S. Liebovitch and Tibor I. Tóth. Distributions of activation energy barriers that produce stretched exponential probability distributions for the time spent in each state of the two state reaction A ⇀ ↽ B. Bulletin of Mathematical Biology, 53(3):443–455, May 1991. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02460727.

Poznanski:1991:GTE

[1131] Roman R. Poznanski. A generalized tapering equivalent cable model for dendritic neurons. Bulletin of Mathematical Biology, 53(3):457– 467, May 1991. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460728.

Stadler:1991:SAR

[1132] Bärbel M. R. Stadler and Peter F. Stadler. Small autocatalytic reaction networks — III. Monotone growth functions. Bulletin of Mathematical Biology, 53(3):469–485, May 1991. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460729.

Kimmel:1991:BR

[1133] Marek Kimmel, David Axelrod, G. B. Ermentrout, Frank Stewart, Khidir W. Hilu, Lee A. Segel, Joan L. Aron, and Norman C. Gilinsky. Book reviews. Bulletin of Mathematical Biology, 53(3): 487–503, May 1991. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460730.

Nowack:1991:TCE

[1134] William J. Nowack and George C. Theodoridis. The thalamocortical contribution to epilepsy. Bulletin of Mathematical Biology, 53(4): 505–523, July 1991. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458626.

Saxena:1991:EDT

[1135] V. P. Saxena and Kamal Raj Pardasani. Effect of dermal tumours on temperature distribution in skin with variable blood flow. Bulletin of Mathematical Biology, 53(4):525–536, July 1991. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02458627.

Aumann:1991:ETC

[1136] Y. Aumann and H. Parnas. Evaluation of the time course of neurotransmitter release from the measured PSC and MPSC. Bulletin of Mathematical Biology, 53(4):537–555, July 1991. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02458628.

Seno:1991:PII

[1137] Hiromi Seno. Predator’s invasion into an isolated patch with spatially heterogeneous prey distribution. Bulletin of Mathematical Biology, 53 (4):557–577, July 1991. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458629.

Carriquiry:1991:SED

[1138] Alicia L. Carriquiry, William P. Ireland, Wolfgang Kliemann, and Etsuro Uemura. Statistical evaluation of dendritic growth models. Bulletin of Mathematical Biology, 53(4):579–589, July 1991. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02458630.

Charnick:1991:CSC

[1139] S. B. Charnick, E. S. Fisher, and D. A. Lauffenburger. Computer simulations of cell-target encounter including biased cell motion toward targets: Single and multiple cell-target simulations in two dimensions. Bulletin of Mathematical Biology, 53(4):591–621, July 1991. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02458631.

Weinberger:1991:SSA

[1140] Edward D. Weinberger. Spatial stability analysis of Eigen’s quasispecies model and the less than five membered hypercycle under global population regulation. Bulletin of Mathematical Biology, 53(4): 623–638, July 1991. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458632.

Michelson:1991:APG

[1141] Seth Michelson and John Leith. Autocrine and paracrine growth factors in tumor growth: a mathematical model. Bulletin of Mathematical Biology, 53(4):639–656, July 1991. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02458633.

Jeka:1991:BR

[1142] John J. Jeka, J. A. S. Kelso, Raymond Kapral, William Troy, and Tim Sauer. Book reviews. Bulletin of Mathematical Biology, 53(4): 657–663, July 1991. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458634.

Dunki:1991:EKE

[1143] Rudolf M. Dünki. The estimation of the Kolmogorov entropy from a time series and its limitations when performed on EEG. Bulletin of Mathematical Biology, 53(5):665–678, September 1991. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02461547.

McMorris:1991:CWH

[1144] F. R. McMorris and R. C. Powers. Consensus weak hierarchies. Bulletin of Mathematical Biology, 53(5):679–684, September 1991. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02461548.

Cazzador:1991:AOY

[1145] Lorenzo Cazzador. Analysis of oscillations in yeast continuous cultures by a new simplified model. Bulletin of Mathematical Biology, 53(5): 685–700, September 1991. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02461549.

Maini:1991:BSH

[1146] P. K. Maini, M. R. Myerscough, K. H. Winter, and J. D. Murray. Bifurcating spatially heterogeneous solutions in a chemotaxis model for biological pattern generation. Bulletin of Mathematical Biology, 53(5): 701–719, September 1991. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02461550.

Ford:1991:ACB

[1147] Roseanne M. Ford and Douglas A. Lauffenburger. Analysis of chemotactic bacterial distributions in population migration assays using a mathematical model applicable to steep or shallow attractant gradients. Bulletin of Mathematical Biology, 53(5):721–749, September 1991. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02461551.

Davis:1991:VPF

[1148] Brian K. Davis. Variations in polymer fitness at elevated mutation rates. Bulletin of Mathematical Biology, 53(5):751–768, September 1991. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02461552.

Glunt:1991:EPP

[1149] W. Glunt, T. L. Hayden, and Wei-Min Liu. The embedding problem for predistance matrices. Bulletin of Mathematical Biology, 53(5):769– 796, September 1991. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02461553.

Sauer:1991:BR

[1150] Tim Sauer, Peter Richter, William J. Floyd, and Brian Storrie. Book reviews. Bulletin of Mathematical Biology, 53(5):797–804, September 1991. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02461554.

Getz:1991:NNP

[1151] Wayne M. Getz. A neural network for processing olfactory-like stimuli. Bulletin of Mathematical Biology, 53(6):805–823, November 1991. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02461485.

Thomas:1991:EVS

[1152] S. R. Thomas. Effect of varying salt and urea permeabilities along descending limbs of Henle in a model of the renal medullary urine concentrating mechanism. Bulletin of Mathematical Biology, 53(6):825– 843, November 1991. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02461486.

Rosen:1991:RTR

[1153] Gerald Rosen. Rook’s tour representation of the genetic code. Bulletin of Mathematical Biology, 53(6):845–851, November 1991. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02461487.

Balding:1991:SAD

[1154] David J. Balding and David C. Torney. Statistical analysis of DNA fingerprint data for ordered clone physical mapping of human chromosomes. Bulletin of Mathematical Biology, 53(6):853–879, November 1991. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02461488.

Lewis:1991:OWB

[1155] M. A. Lewis and P. Grindrod. One-way blocks in cardiac tissue: a mechanism for propagation failure in Purkinje fibres. Bulletin of Mathematical Biology, 53(6):881–899, November 1991. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02461489.

Hsu:1991:KAG

[1156] Jyh-Ping Hsu and Tsue-Hwa Chen. A kinetic analysis of the germination of fungal spores. Bulletin of Mathematical Biology, 53(6):901– 909, November 1991. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02461490.

Wu:1991:SDM

[1157] Jianguo Wu and John L. Vankat. A system dynamics model of island biogeography. Bulletin of Mathematical Biology, 53(6):911– 940, November 1991. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02461491.

Edelstein-Keshet:1991:BR

[1158] Leah Edelstein-Keshet, G. B. Ermentrout, and James R. Griesemer. Book reviews. Bulletin of Mathematical Biology, 53(6):941– 951, November 1991. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02461492.

Anonymous:1991:ICM

[1159] Anonymous. 3rd International Conference on Mathematical Population Dynamics 1–5 June 1992, University of Pau, France. Bulletin of Mathematical Biology, 53(6):953, November 1991. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02461493.

Clarke:1992:MRO

[1160] Bertrand S. Clarke and Jay E. Mittenthal. Modularity and reliability in the organization of organisms. Bulletin of Mathematical Biology, 54 (1):1–20, January 1992. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458617.

Neumann:1992:WAA

[1161] Avidan U. Neumann and Gérard Weisbuch. Window automata analysis of population dynamics in the immune system. Bulletin of Mathematical Biology, 54(1):21–44, January 1992. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02458618.

Olofsen:1992:MLA

[1162] E. Olofsen, J. Degoede, and R. Heijungs. A maximum likelihood approach to correlation dimension and entropy estimation. Bulletin of Mathematical Biology, 54(1):45–58, January 1992. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/article/10.1007/BF02458619.

Mott:1992:MLE

[1163] Richard Mott. Maximum-likelihood estimation of the statistical distribution of Smith–Waterman local sequence similarity scores. Bulletin of Mathematical Biology, 54(1):59–75, January 1992. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02458620.

Myerscough:1992:APP

[1164] M. R. Myerscough and J. D. Murray. Analysis of propagating pattern in a chemotaxis system. Bulletin of Mathematical Biology, 54(1): 77–94, January 1992. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458621.

Wong:1992:MCI

[1165] Alan Y. K. Wong, Alexandre Fabiato, and James B. Bassingthwaighthe. Model of calcium-induced calcium release mechanism in cardiac cells. Bulletin of Mathematical Biology, 54(1):95–116, January 1992. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02458622.

Lobry:1992:MBG

[1166] J. R. Lobry, J. P. Flandrois, G. Carret, and A. Pave. Monod’s bacterial growth model revisited. Bulletin of Mathematical Biology, 54(1): 117–122, January 1992. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458623.

Phelps:1992:OSR

[1167] F. M. Phelps IV. Optimal sex ratio as a function of EGG incubation temperature in the crocodilians. Bulletin of Mathematical Biology, 54 (1):123–148, January 1992. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458624.

Goodman:1992:BR

[1168] Daniel Goodman, Robert Rosen, and W. M. Schaffer. Book reviews. Bulletin of Mathematical Biology, 54(1):149–161, January 1992. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02458625.

Munro:1992:MBE

[1169] Gordon R. Munro. Mathematical bioeconomics and the evolution of modern fisheries economics. Bulletin of Mathematical Biology, 54(2– 3):163–184, March 1992. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02464828.

Reed:1992:CEV

[1170] William J. Reed and Hector Echavarría Heras. The conservation and exploitation of vulnerable resources. Bulletin of Mathematical Biology, 54 (2–3):185–207, March 1992. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02464829.

Stefanou:1992:LVR

[1171] Spiro E. Stefanou and James E. Wilen. License values in restricted access fisheries. Bulletin of Mathematical Biology, 54(2–3): 209–218, March 1992. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02464830.

Conrad:1992:BMP

[1172] Jon M. Conrad. A bioeconomic model of the Pacific whiting. Bulletin of Mathematical Biology, 54(2–3):219–239, March 1992. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02464831.

Levy:1992:RPM

[1173] David A. Levy and Chris C. Wood. Review of proposed mechanisms for sockeye salmon population cycles in the Fraser River. Bulletin of Mathematical Biology, 54(2–3):241–261, March 1992. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02464832.

Hilborn:1992:SPC

[1174] Ray Hilborn and Robert B. Kennedy. Spatial pattern in catch rates: a test of economic theory. Bulletin of Mathematical Biology, 54(2–3): 263–273, March 1992. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02464833.

Botsford:1992:FAC

[1175] Louis W. Botsford. Further analysis of Clark’s delayed recruitment model. Bulletin of Mathematical Biology, 54(2–3):275–293, March 1992. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02464834.

Kaitala:1992:SRG

[1176] Veijo Kaitala and Wayne M. Getz. Sex ratio genetics and the competitiveness of parasitic wasps. Bulletin of Mathematical Biology, 54(2– 3):295–311, March 1992. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02464835.

DeStasio:1992:EVP

[1177] Bart T. De Stasio, Jr. and Nelson G. Hairston, Jr. Environmental variability and the persistence of multiple emergence strategies. Bulletin of Mathematical Biology, 54(2–3):313–334, March 1992. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02464836.

Price:1992:TBR

[1178] Karen Price. Territorial bequeathal by red squirrel mothers: a dynamic model. Bulletin of Mathematical Biology, 54(2–3):335–354, March 1992. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02464837.

McNamara:1992:RSF

[1179] John M. McNamara and Alasdair I. Houston. Risk-sensitive foraging: a review of the theory. Bulletin of Mathematical Biology, 54(2– 3):355–378, March 1992. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02464838.

Goutis:1992:HCM

[1180] Constantinos Goutis and David W. Winkler. Hungry chicks and mortal parents: a state-variable approach to the breeding seasons of birds.

Bulletin of Mathematical Biology, 54(2–3):379–400, March 1992. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02464839.

Roitberg:1992:GTH

[1181] Bernard D. Roitberg and W. G. Friend. A general theory for host seeking decisions in mosquitoes. Bulletin of Mathematical Biology, 54(2– 3):401–412, March 1992. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02464840.

Mangel:1992:RMS

[1182] Marc Mangel. Rate maximizing and state variable theories of diet selection. Bulletin of Mathematical Biology, 54(2–3):413–422, March 1992. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02464841.

Mesterton-Gibbons:1992:IPD

[1183] Michael Mesterton-Gibbons. On the iterated prisoner’s dilemma in a finite population. Bulletin of Mathematical Biology, 54(2–3):423– 443, March 1992. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02464842.

Yoshimura:1992:CUC

[1184] Jin Yoshimura and William M. Shields. Components of uncertainty in clutch-size optimization. Bulletin of Mathematical Biology, 54(2–3): 445–464, March 1992. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02464843.

Houston:1992:EVH

[1185] A. I. Houston, J. M. McNamara, and H. C. J. Godfray. The effect of variability on host feeding and reproductive success in parasitoids. Bulletin of Mathematical Biology, 54(2–3):465–476, March 1992. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02464844.

Nedelman:1992:QPP

[1186] Jerry Nedelman, Patrick Heagerty, and Chip Lawrence. Quantitative PCR: Procedures and precisions. Bulletin of Mathematical Biology, 54 (4):477–502, July 1992. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459631.

Cushing:1992:ISC

[1187] J. M. Cushing and Jia Li. Intra-specific competition and density dependent juvenile growth. Bulletin of Mathematical Biology, 54(4): 503–519, July 1992. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459632.

Schoniger:1992:LAD

[1188] Michael Schöniger and Michael S. Waterman. A local algorithm for DNA sequence alignment with inversions. Bulletin of Mathematical Biology, 54(4):521–536, July 1992. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459633.

Weinstein:1992:AVR

[1189] Alan M. Weinstein. Analysis of volume regulation in an epithelial cell model. Bulletin of Mathematical Biology, 54(4):537–561, July 1992. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02459634.

Chan:1992:SMS

[1190] S. C. Chan, A. K. C. Wong, and D. K. Y. Chiu. A survey of multiple sequence comparison methods. Bulletin of Mathematical Biology, 54(4):563–598, July 1992. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459635.

Myers:1992:RMC

[1191] Eugene W. Myers and Xiaoqiu Huang. An O(N^{2} log N) restriction map comparison and search algorithm. Bulletin of Mathematical Biology, 54 (4):599–618, July 1992. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459636.

Kot:1992:CDM

[1192] Mark Kot, Gary S. Sayler, and Terry W. Schultz. Complex dynamics in a model microbial system. Bulletin of Mathematical Biology, 54 (4):619–648, July 1992. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459637.

Perelson:1992:MIR

[1193] Alan S. Perelson and Gérand Weisbuch. Modeling immune reactivity in secondary lymphoid organs. Bulletin of Mathematical Biology, 54 (4):649–672, July 1992. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459638.

Deakin:1992:DTC

[1194] Michael A. B. Deakin, Robert A. R. Bywater, and S. J. Redman. Determination of time-constants in cables of finite length. Bulletin of Mathematical Biology, 54(4):673–686, July 1992. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02459639.

Fox:1992:BR

[1195] Gordon A. Fox, Ronald F. Fox, and Colin W. Clark. Book reviews. Bulletin of Mathematical Biology, 54(4):687–694, July 1992. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02459640.

Hutchinson:1992:PSA

[1196] G. E. Hutchinson. Population studies: animal ecology and demography. Bulletin of Mathematical Biology, 54(4):695, July 1992. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02459641;
http://link.springer.com/content/pdf/10.1007/BF02459641.pdf.

Anonymous:1992:A

[1197] Anonymous. Announcements. Bulletin of Mathematical Biology, 54 (4):697–698, July 1992. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459642.

Neumann:1992:DTI

[1198] Avidan U. Neumann and Gérard Weisbuch. Dynamics and topology of idiotypic networks. Bulletin of Mathematical Biology, 54(5):699– 726, September 1992. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459926.

Ghosh:1992:IMB

[1199] Subhendu Ghosh. Ising model for B–Z transition in supercoiled DNA. Bulletin of Mathematical Biology, 54(5):727–732, September 1992. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02459927.

Greenhalgh:1992:VDD

[1200] David Greenhalgh. Vaccination in density-dependent epidemic models. Bulletin of Mathematical Biology, 54(5):733–758, September 1992. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02459928.

VanPelt:1992:TAS

[1201] Jaap Van Pelt, Harry B. M. Uylings, Ronald W. H. Verwer, Roberta J. Pentney, and Michael J. Woldenberg. Tree asymmetry — a sensitive and practical measure for binary topological trees. Bulletin of Mathematical Biology, 54(5):759–784, September 1992. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02459929.

Goldstein:1992:PCP

[1202] Larry Goldstein and Michael S. Waterman. Poisson, compound Poisson, and process approximations for testing statistical significance in sequence comparisons. Bulletin of Mathematical Biology, 54(5):785– 812, September 1992. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459930.

Lam:1992:SCA

[1203] Gabriel K. Y. Lam. The study of combined action of agents using differential geometry of dose-effect surfaces. Bulletin of Mathematical Biology, 54(5):813–826, September 1992. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02459931.

Gard:1992:SMT

[1204] Thomas C. Gard. Stochastic models for toxicant-stressed populations. Bulletin of Mathematical Biology, 54(5):827–837, September 1992. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02459932.

Lefever:1992:KOS

[1205] René Lefever, Jacques Hiernaux, Jacques Urbain, and Patricia Meyers. On the kinetics and optimal specificity of cytotoxic reactions mediated by T-lymphocyte clones. Bulletin of Mathematical Biology, 54(5):839– 873, September 1992. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459933.

Hovav:1992:NRF

[1206] G. Hovav, H. Parnas, and I. Parnas. Neurotransmitter release: Facilitation and three-dimensional diffusion of intracellular calcium. Bulletin of Mathematical Biology, 54(5):875–894, September 1992. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02459934.

Wells:1992:BR

[1207] D. M. Wells, J. A. S. Kelso, Lisa Fauci, and Jacques Bélair. Book reviews. Bulletin of Mathematical Biology, 54(5):895–901, September 1992. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02459935.

Anonymous:1992:ISN

[1208] Anonymous. International school on neural modeling and neural networks. Bulletin of Mathematical Biology, 54(5):903, September 1992. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02459936.

Poznanski:1992:MES

[1209] Roman R. Poznanski. Modelling the electrotonic structure of starburst amacrine cells in the rabbit retina: a functional interpretation of dendritic morphology. Bulletin of Mathematical Biology, 54(6):905– 928, November 1992. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460658.

Ebaseh-Onofa:1992:CME

[1210] B. O. Ebaseh-Onofa and J. H. Matis. Compartmental models with Erlang distributed residence times and random rate coefficients. Bulletin of Mathematical Biology, 54(6):929–938, November 1992. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02460659.

Sole:1992:NPC

[1211] Ricard V. Solé and Joaquim Valls. Nonlinear phenomena and chaos in a Monte Carlo simulated microbial ecosystem. Bulletin of Mathematical Biology, 54(6):939–955, November 1992. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460660.

Zonneveld:1992:PPB

[1212] C. Zonneveld. Polyandry and protandry in butterflies. Bulletin of Mathematical Biology, 54(6):957–976, November 1992. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/article/10.1007/BF02460661.

Barry:1992:FID

[1213] S. I. Barry and G. K. Aldis. Flow-induced deformation from pressurized cavities in absorbing porous tissues. Bulletin of Mathematical Biology, 54(6):977–997, November 1992. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460662.

Baake:1992:MFF

[1214] Ellen Baake and Johannes P. Schlöder. Modelling the fast fluorescence rise of photosynthesis. Bulletin of Mathematical Biology, 54(6):999– 1021, November 1992. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460663.

Michelson:1992:MMP

[1215] Seth Michelson and Doris Slate. A mathematical model of the Pglycoprotein pump as a mediator of multidrug resistance. Bulletin of Mathematical Biology, 54(6):1023–1038, November 1992. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02460664.

Palmer:1992:ISR

[1216] A. Palmer, J. Brindley, and A. V. Holden. Initiation and stability of reentry in two coupled excitable fibers. Bulletin of Mathematical Biology, 54(6):1039–1056, November 1992. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460665.

Day:1992:CSB

[1217] William H. E. Day and F. R. McMorris. Consensus sequences based on plurality rule. Bulletin of Mathematical Biology, 54(6):1057– 1068, November 1992. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460666.

Morris:1992:UFL

[1218] Anne Krislov Morris and Wendy Kuhn Silk. Use of a flexible logistic function to describe axial growth of plants. Bulletin of Mathematical Biology, 54(6):1069–1081, November 1992. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02460667.

Rosen:1992:BR

[1219] Robert Rosen, Steven H. Strogatz, Josesph M. Mahaffy, and Joseph M. Mahaffy. Book reviews. Bulletin of Mathematical Biology, 54(6):1083– 1094, November 1992. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460668.

Masutani:1993:EST

[1220] Keiichi Masutani. Effects of survival thresholds upon one-dimensional dynamics of single-species populations. Bulletin of Mathematical Biology, 55(1):1–13, January 1993. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460292.

Rinaldi:1993:MAC

[1221] Sergio Rinaldi, Simona Muratori, and Yuri Kuznetsov. Multiple attractors, catastrophes and chaos in seasonally perturbed predator–prey communities. Bulletin of Mathematical Biology, 55(1):15–35, January 1993. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02460293.

Totafurno:1993:EEC

[1222] John Totafurno, Matthew Bjerknes, and Hazel Cheng. Estimation of the extent of the cutoff region from the spatial distribution of labelling and mitotic indices of intestinal crypts of a fixed length. Bulletin of Mathematical Biology, 55(1):37–73, January 1993. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02460294.

Sendova-Franks:1993:TAA

[1223] Ana Sendova-Franks and Nigel R. Franks. Task allocation in ant colonies within variable environments (a study of temporal polyethism: Experimental). Bulletin of Mathematical Biology, 55(1):75–96, January 1993. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02460295.

Lara-Aparicio:1993:QMM

[1224] Miguel Lara-Aparicio, Santiago López de Medrano, Beatriz Fuentes-Pardo, and Enrique Moreno-Sáenz. A qualitative mathematical model of the ontogeny of a circadian rhythm in crayfish. Bulletin of Mathematical Biology, 55(1):97–110, January 1993. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02460296.

Pritchard:1993:SAK

[1225] W. G. Pritchard. Scaling in the animal kingdom. Bulletin of Mathematical Biology, 55(1):111–129, January 1993. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02460297.

Luo:1993:VMP

[1226] Xincai Luo, Keith Taylor, and Paul G. Mezey. Vertex mobility of polyhedra. Bulletin of Mathematical Biology, 55(1):131–140, January 1993. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02460298.

Gusfield:1993:EMM

[1227] Dan Gusfield. Efficient methods for multiple sequence alignment with guaranteed error bounds. Bulletin of Mathematical Biology, 55(1): 141–154, January 1993. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460299.

Horvath:1993:TSO

[1228] Gábor Horváth and Dezsö Varjú. Theoretical study of the optimal shape of the front profile of the lens in the eye of the scallop, pecten. Bulletin of Mathematical Biology, 55(1):155–174, January 1993. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02460300.

Fuxman:1993:RRF

[1229] Yakov L. Fuxman. Reproduction rate, feeding process, and Leibich limitations in cell populations — Part 1. Feeding stochasticity and reproduction rate. Bulletin of Mathematical Biology, 55(1):175–195, January 1993. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02460301.

Crenshaw:1993:OHMa

[1230] Hugh C. Crenshaw. Orientation by helical motion — I. Kinematics of the helical motion of organisms with up to six degrees of freedom. Bulletin of Mathematical Biology, 55(1):197–212, January 1993. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02460302.

Crenshaw:1993:OHMb

[1231] Hugh C. Crenshaw and Leah Edelstein-Keshet. Orientation by helical motion — II. Changing the direction of the axis of motion. Bulletin of Mathematical Biology, 55(1):213–230, January 1993. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02460303.

Crenshaw:1993:OHMc

[1232] Hugh C. Crenshaw. Orientation by helical motion — III. Microorganisms can orient to stimuli by changing the direction of their rotational velocity. Bulletin of Mathematical Biology, 55(1):231–255, January 1993. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02460304.

Anonymous:1993:A

[1233] Anonymous. Announcement. Bulletin of Mathematical Biology, 55 (1):257, January 1993. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460305.

Kmet:1993:MMA

[1234] T. Kmeť, M. Straškraba, and P. Mauersberger. A mechanistic model of the adaptation of phytoplankton photosynthesis. Bulletin of Mathematical Biology, 55(2):259–275, March 1993. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02460883.

vanVeen:1993:TIS

[1235] Mark P. van Veen and Jaap van Pelt. Terminal and intermediate segment lengths in neuronal trees with finite length. Bulletin of Mathematical Biology, 55(2):277–294, March 1993. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460884.

Lam:1993:DAL

[1236] Gabriel K. Y. Lam. The differential aspects of the linear isobole in the study of combined action of agents. Bulletin of Mathematical Biology, 55 (2):295–313, March 1993. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460885.

Sneyd:1993:CWP

[1237] James Sneyd, Steven Girard, and David Clapham. Calcium wave propagation by calcium-induced calcium release: an unusual excitable system. Bulletin of Mathematical Biology, 55(2):315–344, March 1993. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02460886.

Korn:1993:GEP

[1238] Robert W. Korn. The geometry of elongating plant cells. Bulletin of Mathematical Biology, 55(2):345–364, March 1993. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/article/10.1007/BF02460887.

Benson:1993:DDI

[1239] Debbie L. Benson, Jonathan A. Sherratt, and Philip K. Maini. Diffusion driven instability in an inhomogeneous domain. Bulletin of Mathematical Biology, 55(2):365–384, March 1993. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460888.

Nuno:1993:MAN

[1240] J. C. Nuño, M. A. Andrade, F. Morán, and F. Montero. A model of an autocatalytic network formed by error-prone self-replicative species. Bulletin of Mathematical Biology, 55(2):385–415, March 1993. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02460889.

Nuno:1993:NUS

[1241] J. C. Nuño, M. A. Andrade, and F. Montero. Non-uniformities and superimposed competition in a model of an autocatalytic network formed by error-prone self-replicative species. Bulletin of Mathematical Biology, 55(2):417–449, March 1993. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460890.

Umeda:1993:TMC

[1242] Tamiki Umeda. A thermodynamical model of cell distributions in the slug of cellular slime mold. Bulletin of Mathematical Biology, 55(2): 451–464, March 1993. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460891.

Wong:1993:MSC

[1243] A. K. C. Wong, S. C. Chan, and D. K. Y. Chiu. A multiple sequence comparison method. Bulletin of Mathematical Biology, 55(2): 465–486, March 1993. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460892.

Gyllenberg:1993:DTL

[1244] Mats Gyllenberg. Does time lag of nutrient utilization justify Monod’s model of bacterial growth? Bulletin of Mathematical Biology, 55(2): 487–489, March 1993. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460893.

Rohlf:1993:BR

[1245] F. James Rohlf, Herbert W. Hethcote, Donald C. Mikulecky, and J.- V. Höltje. Book reviews. Bulletin of Mathematical Biology, 55(2): 491–502, March 1993. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460894.

Chao:1993:CSA

[1246] Kun-Mao Chao, Ross C. Hardison, and Webb Miller. Constrained sequence alignment. Bulletin of Mathematical Biology, 55(3):503– 524, May 1993. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460648.

Beuter:1993:FDN

[1247] Anne Beuter, Jacques Bélair, Christiane Labrie, and Jacques Bélair. Feedback and delays in neurological diseases: a modeling study using gynamical systems. Bulletin of Mathematical Biology, 55(3): 525–541, May 1993. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460649.

Findlay:1993:SCS

[1248] Scott Findlay and Ligang Zheng. Structural correlates of stability and resilience in strong hierarchies. Bulletin of Mathematical Biology, 55 (3):543–560, May 1993. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460650.

Havsteen:1993:KES

[1249] B. H. Havsteen, M. Garcia-Moreno, E. Valero, M. C. Manjabacas, and R. Varón. The kinetics of enzyme systems involving activation of zymogens. Bulletin of Mathematical Biology, 55(3):561–583, May 1993. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02460651.

Schienbein:1993:LEF

[1250] M. Schienbein and H. Gruler. Langevin equation, Fokker–Planck equation and cell migration. Bulletin of Mathematical Biology, 55(3): 585–608, May 1993. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460652.

Zonneveld:1993:CKE

[1251] C. Zonneveld and S. A. L. M. Kooijman. Comparative kinetics of embryo development. Bulletin of Mathematical Biology, 55(3): 609–635, May 1993. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460653.

Sherratt:1993:SIA

[1252] Jonathan A. Sherratt and Julian Lewis. Stress-induced alignment of actin filaments and the mechanics of cytogel. Bulletin of Mathematical Biology, 55(3):637–654, May 1993. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460654.

McElwain:1993:CMM

[1253] D. L. S. McElwain and G. J. Pettet. Cell migration in multicell spheroids: Swimming against the tide. Bulletin of Mathematical Biology, 55(3): 655–674, May 1993. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460655.

Novak:1993:BR

[1254] Bela Novak, G. B. Ermentrout, Marc Mangel, F. W. Schneider, Keith A. Crandall, Kerry L. Shaw, Thomas Caraco, and Elaina M. Tuttle. Book reviews. Bulletin of Mathematical Biology, 55(3):675–691, May 1993. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02460656.

Anonymous:1993:SSN

[1255] Anonymous. Summer school on nonlinear systems in evolutionary and population biology. Bulletin of Mathematical Biology, 55 (3):693, May 1993. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460657.

Mirkin:1993:CFP

[1256] Boris Mirkin and Fred S. Roberts. Consensus functions and patterns in molecular sequences. Bulletin of Mathematical Biology, 55 (4):695–713, July 1993. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460669.

Beretta:1993:OTC

[1257] E. Beretta and F. Solimano. An optimal time control problem for the administration of 2’,3’-dideoxycytidine by using red blood cells as bioreactors. Bulletin of Mathematical Biology, 55(4):715–730, July 1993. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02460670.

Donnelly:1993:CTR

[1258] Christl Donnelly, Wendy Leisenring, Phyllis Kanki, Tamara Awerbuch, and Sonja Sandberg. Comparison of transmission rates of HIV-1 and HIV-2 in a cohort of prostitutes in Senegal. Bulletin of Mathematical Biology, 55(4):731–743, July 1993. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460671.

DeBoer:1993:INBa

[1259] Rob J. De Boer, Alan S. Perelson, and Ioannis G. Kevrekidis. Immune network behavior — I. From stationary states to limit cycle oscilations. Bulletin of Mathematical Biology, 55(4):745–780, July 1993. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02460672.

DeBoer:1993:INBb

[1260] Rob J. De Boer, Alan S. Perelson, and Ioannis G. Kevrekidis. Immune network behavior — II. From oscillations to chaos and stationary states. Bulletin of Mathematical Biology, 55(4):781–816, July 1993. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02460673.

Freedman:1993:PPP

[1261] H. I. Freedman and R. M. Mathsen. Persistence in predator–prey systems with ratio-dependent predator influence. Bulletin of Mathematical Biology, 55(4):817–827, July 1993. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460674.

Jeka:1993:PSH

[1262] John J. Jeka, J. A. S. Kelso, and Tim Kiemel. Pattern switching in human multilimb coordination dynamics. Bulletin of Mathematical Biology, 55 (4):829–845, July 1993. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460675.

Allen:1993:MAM

[1263] Linda J. S. Allen, Mary K. Hannigan, and Monty J. Strauss. Mathematical analysis of a model for a plant-herbivore system. Bulletin of Mathematical Biology, 55(4):847–864, July 1993. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02460676.

Reed:1993:MML

[1264] Michael C. Reed. Mathematics in medicine and the life sciences. Bulletin of Mathematical Biology, 55(4):865–867, July 1993. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02460677.

Hsu:1993:DEB

[1265] Jyh-Ping Hsu, Wen-Chang Hsu, and Heng-Kwong Tsao. Diffusionenhanced bioreactions: a hypothetical mechanism for plant cell aggregation. Bulletin of Mathematical Biology, 55(5):869–889, September 1993. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02460690.

Tofts:1993:ATA

[1266] Chris Tofts. Algorithms for task allocation in ants. (A study of temporal polyethism: Theory). Bulletin of Mathematical Biology, 55(5): 891–918, September 1993. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460691.

Stone:1993:PFA

[1267] Lewi Stone and Tom Berman. Positive feedback in aquatic ecosystems: The case of the microbial loop. Bulletin of Mathematical Biology, 55(5): 919–936, September 1993. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460692.

Guckenheimer:1993:BHH

[1268] J. Guckenheimer and J. S. Labouriau. Bifurcation of the Hodgkin and Huxley equations: a new twist. Bulletin of Mathematical Biology, 55(5): 937–952, September 1993. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460693.

Velten:1993:OMI

[1269] Kai Velten and Otto Richter. Optimal maintenance investment of plants and its dependence on environmental conditions. Bulletin of Mathematical Biology, 55(5):953–971, September 1993. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02460694.

ElSnoussi:1993:LIA

[1270] Houssine El Snoussi and Rene Thomas. Logical identification of all steady states: The concept of feedback loop characteristic states. Bulletin of Mathematical Biology, 55(5):973–991, September 1993. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02460695.

Michelson:1993:GFG

[1271] Seth Michelson and John T. Leith. Growth factors and growth control of heterogeneous cell populations. Bulletin of Mathematical Biology, 55(5): 993–1011, September 1993. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460696.

Agur:1993:BR

[1272] Zvia Agur, Michael LaBarbera, George F. Estabrook, and Charles J. Brokaw. Book reviews. Bulletin of Mathematical Biology, 55(5):1013– 1024, September 1993. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460697.

Arques:1993:AEP

[1273] Didier G. Arquès and Christian J. Michel. Analytical expression of the purine/pyrimidine codon probability after and before random mutations. Bulletin of Mathematical Biology, 55(6):1025–1038, November 1993. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02460698.

Gadani:1993:NMS

[1274] Vincent Gadani, Pierre Villon, Jacques Manem, and Bruce Rittmann. A new method to solve a non-steady-state multispecies biofilm model. Bulletin of Mathematical Biology, 55(6):1039–1061, November 1993. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02460699.

Weisbuch:1993:EMA

[1275] Gérard Weisbuch and Guillemette Duchateau. Emergence of mutualism: Application of a differential model to endosymbiosis. Bulletin of Mathematical Biology, 55(6):1063–1090, November 1993. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02460700.

Anderson:1993:CTI

[1276] Russell W. Anderson, Avidan U. Neumann, and Alan S. Perelson. A Cayley tree immune network model with antibody dynamics. Bulletin of Mathematical Biology, 55(6):1091–1131, November 1993. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02460701.

Sulzer:1993:MIN

[1277] Bernhard Sulzer, J. Leo van Hemmen, Avidan U. Neumann, and Ulrich Behn. Memory in idiotypic networks due to competition between proliferation and differentiation. Bulletin of Mathematical Biology, 55(6): 1133–1182, November 1993. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460702.

Unger:1993:FLF

[1278] Ron Unger and John Moult. Finding the lowest free energy conformation of a protein is an NP-hard problem: Proof and implications. Bulletin of Mathematical Biology, 55(6):1183–1198, November 1993. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02460703.

Fraenkel:1993:CPF

[1279] Aviezri S. Fraenkel. Complexity of protein folding. Bulletin of Mathematical Biology, 55(6):1199–1210, November 1993. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02460704.

Fontana:1994:AFT

[1280] Walter Fontana and Leo W. Buss. “The arrival of the fittest”: Toward a theory of biological organization. Bulletin of Mathematical Biology, 56(1):1–64, January 1994. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458289.

Wilhelm:1994:EAE

[1281] Thomas Wilhelm, Edda Hoffmann-Klipp, and Reinhart Heinrich. An evolutionary approach to enzyme kinetics: Optimization of ordered mechanisms. Bulletin of Mathematical Biology, 56(1):65–106, January 1994. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458290.

Tuck:1994:OHS

[1282] Geoffrey N. Tuck and Hugh P. Possingham. Optimal harvesting strategies for a metapopulation. Bulletin of Mathematical Biology, 56(1): 107–127, January 1994. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458291.

Sherratt:1994:CCE

[1283] Jonathan A. Sherratt. Chemotaxis and chemokinesis in eukaryotic cells: The Keller–Segel equations as an approximation to a detailed model. Bulletin of Mathematical Biology, 56(1):129–146, January 1994. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02458292.

Sanchez-Mejorada:1994:PMC

[1284] Gabriela Sánchez-Mejorada, Horacio Merchant-Larios, Francisco AlonsodeFlorida, and Luis B. Morales. Probability of mast cells to mediate anaphylaxis in skeletal muscle. Bulletin of Mathematical Biology, 56 (1):147–160, January 1994. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458293.

Edelstein-Keshet:1994:BR

[1285] Leah Edelstein-Keshet, Ronald M. Shymko, Werner Horsthemke, and R. Allen White. Book reviews. Bulletin of Mathematical Biology, 56 (1):161–170, January 1994. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458294.

Gitterman:1994:GTK

[1286] Moshe Gitterman and George H. Weiss. Generalized theory of the kinetics of tracers in biological systems. Bulletin of Mathematical Biology, 56 (2):171–186, March 1994. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460638.

Timm:1994:GPM

[1287] Uwe Timm and Akira Okubo. Gyrotaxis: a plume model for selffocusing micro-organisms. Bulletin of Mathematical Biology, 56(2): 187–206, March 1994. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460639.

Michelson:1994:MMI

[1288] Seth Michelson and Doris Slate. A mathematical model for the inhibition of the multidrug resistance-associated P-glycoprotein pump. Bulletin of Mathematical Biology, 56(2):207–223, March 1994. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02460640.

Overholser:1994:DPM

[1289] K. A. Overholser, N. A. Lomangino, T. R. Harris, J. D. Bradléy, and S. Bosan. Deduction of pulmonary microvascular hematocrit from indicator dilution curves. Bulletin of Mathematical Biology, 56(2): 225–247, March 1994. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.

1007/BF02460641. See erratum [1379].

VanVeen:1994:NGR

[1290] Mark P. Van Veen and Jaap Van Pelt. Neuritic growth rate described by modeling microtubule dynamics. Bulletin of Mathematical Biology, 56 (2):249–273, March 1994. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460642.

Skiena:1994:PDA

[1291] Steven S. Skiena and Gopalakrishnan Sundaram. A partial digest approach to restriction site mapping. Bulletin of Mathematical Biology, 56 (2):275–294, March 1994. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460643.

Kuznetsov:1994:NDI

[1292] Vladimir A. Kuznetsov, Iliya A. Makalkin, Mark A. Taylor, and Alan S. Perelson. Nonlinear dynamics of immunogenic tumors: Parameter estimation and global bifurcation analysis. Bulletin of Mathematical Biology, 56(2):295–321, March 1994. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460644.

Zhang:1994:CBM

[1293] Li-Bing Zhang, Chan-Ying Du, and An-Shen Qi. Complex behaviours of AB model describing idiotypic network. Bulletin of Mathematical Biology, 56(2):323–336, March 1994. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460645.

Kimmel:1994:TCB

[1294] Marek Kimmel and David N. Stivers. Time-continuous branching walk models of unstable gene amplification. Bulletin of Mathematical Biology, 56(2):337–357, March 1994. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460646.

Ribbens:1994:BR

[1295] Calvin Ribbens and Matthew Witten. Book reviews. Bulletin of Mathematical Biology, 56(2):359–363, March 1994. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02460647.

Marsh:1994:I

[1296] D. J. Marsh. Introduction. Bulletin of Mathematical Biology, 56(3): 365–368, May 1994. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460462.

Oliver:1994:RCM

[1297] James D. Oliver III and William M. Deen. Random-coil model for glomerular sieving of dextran. Bulletin of Mathematical Biology, 56 (3):369–389, May 1994. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460463.

Moore:1994:AMA

[1298] Leon C. Moore, Adam Rich, and Daniel Casellas. Ascending myogenic autoregulation: Interactions between tubuloglomerular feedback and myogenic mechanisms. Bulletin of Mathematical Biology, 56(3): 391–410, May 1994. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460464.

Holstein-Rathlou:1994:DMR

[1299] N.-H. Holstein-Rathlou and D. J. Marsh. A dynamic model of renal blood flow autoregulation. Bulletin of Mathematical Biology, 56 (3):411–429, May 1994. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460465.

Thomas:1994:KMR

[1300] S. Randall Thomas and George Dagher. A kinetic model of rat proximal tubule transport-load-dependent bicarbonate reabsorption along the tubule. Bulletin of Mathematical Biology, 56(3):431–458, May 1994. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02460466.

Krahn:1994:WAP

[1301] Thomas A. Krahn, Peter S. Aronson, and Alan M. Weinstein. Weak acid permeability of a villous membrane: Formic acid transport across rat proximal tubule. Bulletin of Mathematical Biology, 56(3): 459–490, May 1994. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460467.

Jen:1994:EDC

[1302] J. Frank Jen and John L. Stephenson. Externally driven countercurrent multiplication in a mathematical model of the urinary concentrating mechanism of the renal inner medulla. Bulletin of Mathematical Biology, 56(3):491–514, May 1994. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460468.

Wang:1994:ESN

[1303] Xianqun Wang, Anthony S. Wexler, and Donald J. Marsh. The effect of solution non-ideality on membrane transport in three-dimensional models of the renal concentrating mechanism. Bulletin of Mathematical Biology, 56(3):515–546, May 1994. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460469.

Layton:1994:DNM

[1304] H. E. Layton and E. Bruce Pitman. A dynamic numerical method for models of renal tubules. Bulletin of Mathematical Biology, 56(3): 547–565, May 1994. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460470.

Pitman:1994:NSP

[1305] E. Bruce Pitman, H. E. Layton, and Leon C. Moore. Numerical simulation of propagating concentration profiles in renal tubules. Bulletin of Mathematical Biology, 56(3):567–586, May 1994. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/article/10.1007/BF02460471.

Civelekoglu:1994:MDF

[1306] Gül Civelekoğlu and Leah Edelstein-Keshet. Modelling the dynamics of F-actin in the cell. Bulletin of Mathematical Biology, 56(4): 587–616, July 1994. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460713.

Marusic:1994:TGV

[1307] Miljenko Marušic, Željko Bajzer, Stanimir Vuk-Pavlovic, and James P. Freyer. Tumor growth in vivo and as multicellular spheroids compared by mathematical models. Bulletin of Mathematical Biology, 56 (4):617–631, July 1994. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460714.

Indermitte:1994:CAE

[1308] C. Indermitte, Th. M. Liebling, and H. Clémençon. Culture analysis and external interaction models of mycelial growth. Bulletin of Mathematical Biology, 56(4):633–664, July 1994. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02460715.

Falkovitz:1994:MLS

[1309] Meira S. Falkovitz and Eli Feinerman. Minimum leaching scheduling of nitrogen fertilization and irrigation. Bulletin of Mathematical Biology, 56(4):665–686, July 1994. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460716.

Segel:1994:REM

[1310] Lee A. Segel and Eva Jäger. Reverse engineering: a model for T-cell vaccination. Bulletin of Mathematical Biology, 56(4):687–721, July 1994. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02460717.

Phillips:1994:MFT

[1311] C. G. Phillips, K. H. Parker, and W. Wang. A model for flow through discontinuities in the tight junction of the endothelial intercellular cleft. Bulletin of Mathematical Biology, 56(4):723–741, July 1994. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02460718.

Waterman:1994:PES

[1312] Michael S. Waterman. Parametric and ensemble sequence alignment algorithms. Bulletin of Mathematical Biology, 56(4):743–767, July 1994. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02460719.

Conrad:1994:BR

[1313] Michael Conrad and Donald L. DeAngelis. Book reviews. Bulletin of Mathematical Biology, 56(4):769–774, July 1994. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/article/10.1007/BF02460720.

Korn:1994:ACP

[1314] Robert W. Korn. Analysis of cell packets in elongating plant tissue. Bulletin of Mathematical Biology, 56(5):775–794, September 1994. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02458267.

Lubkin:1994:UWR

[1315] Sharon Lubkin. Unidirectional waves on rings: Models for chiral preference of circumnutating plants. Bulletin of Mathematical Biology, 56(5): 795–810, September 1994. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458268.

Sherman:1994:APA

[1316] Arthur Sherman. Anti-phase, asymmetric and aperiodic oscillations in excitable cells — I. Coupled bursters. Bulletin of Mathematical Biology, 56(5):811–835, September 1994. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02458269.

Minkevich:1994:DCM

[1317] I. G. Minkevich and A. Yu. Abramychev. The dynamics of continuous microbial culture described by cell age distribution and concentration of one substrate. Bulletin of Mathematical Biology, 56(5):837– 862, September 1994. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458270.

Liu:1994:NFH

[1318] Wei min Liu. Nonlinear forecasting of hepatitis and AIDS incidence. Bulletin of Mathematical Biology, 56(5):863–873, September 1994. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02458271.

Nuno:1994:LSC

[1319] Juan Carlos Nuño and Pedro Tarazona. Lifetimes of small catalytic networks. Bulletin of Mathematical Biology, 56(5):875–898, September 1994. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02458272.

Weisbuch:1994:CMI

[1320] Gérard Weisbuch and Mihaela Oprea. Capacity of a model immune network. Bulletin of Mathematical Biology, 56(5):899–921, September 1994.

CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02458273.

Da-Fu:1994:DGT

[1321] Ding Da-Fu, Qian Jiang, and Feng Zu-Kang. A differential geometric treatment of protein structure comparison. Bulletin of Mathematical Biology, 56(5):923–943, September 1994. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02458274.

Wennergren:1994:CSA

[1322] Uno Wennergren, Per Weinerfelt, and Göran Forsling. Comparative sensitivity analysis of stable stage structures and reproductive values. Bulletin of Mathematical Biology, 56(5):945–957, September 1994. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02458275.

Lam:1994:GFC

[1323] Gabriel K. Y. Lam. A general formulation of the concept of independent action for the combined effects of agents. Bulletin of Mathematical Biology, 56(5):959–980, September 1994. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02458276.

Truscott:1994:OPP

[1324] J. E. Truscott and J. Brindley. Ocean plankton populations as excitable media. Bulletin of Mathematical Biology, 56(5):981–998, September 1994. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458277.

Rosen:1994:BR

[1325] Robert Rosen, G. B. Ermantrout, Steven H. Strogatz, and Michael Mesterton-Gibbons. Book reviews. Bulletin of Mathematical Biology, 56(5):999–1008, September 1994. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02458278.

Sulzer:1994:CIS

[1326] Bernhard Sulzer, J. Leo Van Hemmen, and Ulrich Behn. Central immune system, the self and autoimmunity. Bulletin of Mathematical Biology, 56(6):1009–1040, November 1994. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460286.

Moghe:1994:SMC

[1327] Prabhas V. Moghe and Robert T. Tranquillo. Stochastic model of chemoattractant receptor dynamics in leukocyte chemosensory movement. Bulletin of Mathematical Biology, 56(6):1041–1093, November 1994. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460287.

Aharon:1994:MSC

[1328] S. Aharon, H. Parnas, and I. Parnas. The magnitude and significance of ca^{2+} domains for release of neurotransmitter. Bulletin of Mathematical Biology, 56(6):1095–1119, November 1994. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02460288.

Eakin:1994:ITS

[1329] T. Eakin. Intrinsic time scaling in survival analysis: Application to biological populations. Bulletin of Mathematical Biology, 56(6):1121– 1141, November 1994. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460289.

Sheng:1994:PMB

[1330] Ke-Ning Sheng and Joseph I. Naus. Pattern matching between two nonaligned random sequences. Bulletin of Mathematical Biology, 56(6):1143– 1162, November 1994. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460290.

Sevella:1994:BR

[1331] Bela Sevella, Nicholas D. Stone, Stephen P. Ellner, Philip K. Maini, and William S. C. Gurney. Book reviews. Bulletin of Mathematical Biology, 56(6):1163–1172, November 1994. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02460291.

Kuzin:1995:HMN

[1332] I. A. Kuzin. Hypnosis modelling in neural networks. Bulletin of Mathematical Biology, 57(1):1–20, January 1995. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02458314.

Stadler:1995:DSA

[1333] Peter F. Stadler, Wolfgang Schnabl, Christian V. Forst, and Peter Schuster. Dynamics of small autocatalytic reaction networks — II. Replication, mutation and catalysis. Bulletin of Mathematical Biology, 57 (1):21–61, January 1995. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458315.

Collings:1995:BSA

[1334] John B. Collings. Bifurcation and stability analysis of a temperaturedependent mite predator–prey interaction model incorporating a prey refuge. Bulletin of Mathematical Biology, 57(1):63–76, January 1995. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02458316.

Myers:1995:IID

[1335] T. G. Myers, OCIAM, G. K. Aldis, and S. Naili. Ion induced deformation of soft tissue. Bulletin of Mathematical Biology, 57(1): 77–98, January 1995. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458317.

Velten:1995:ORS

[1336] Kai Velten and Otto Richter. Optimal root/shoot-partitioning of carbohydrates in plants. Bulletin of Mathematical Biology, 57(1):99– 107, January 1995. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458318.

deMonvel:1995:MCL

[1337] Jacques H. Boutet de Monvel and Olivier C. Martin. Memory capacity in large idiotypic networks. Bulletin of Mathematical Biology, 57(1): 109–136, January 1995. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458319.

Wong:1995:KMC

[1338] Alan Y. K. Wong. A kinetic model of coronary reactive hyperemic response to transient ischemia. Bulletin of Mathematical Biology, 57(1): 137–156, January 1995. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458320.

Navarro-Lozano:1995:KSE

[1339] M. J. Navarro-Lozano, E. Valero, R. Varon, and F. Garcia-Carmona. Kinetic study of an enzyme-catalysed reaction in the presence of novel irreversible-type inhibitors that react with the product of enzymatic catalysis. Bulletin of Mathematical Biology, 57(1):157–168, January 1995. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458321.

West:1995:BR

[1340] Bruce West and J. M. Cushing. Book reviews. Bulletin of Mathematical Biology, 57(1):169–173, January 1995. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02458322.

Fuentes-Pardo:1995:PCR

[1341] Beatriz Fuentes-Pardo, Miguel Lara-Aparicio, and Santiago López de Medrano. Perturbation of a circadian rhythm by single and periodic signals and its mathematical simulation. Bulletin of Mathematical Biology, 57(2):175–189, March 1995. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460614.

Rognstad:1995:EAP

[1342] Robert Rognstad. On the estimation of alternative pathways of fatty acid oxidation in the liver In vivo. Bulletin of Mathematical Biology, 57 (2):191–203, March 1995. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460615.

Kalay:1995:NGH

[1343] A. Kalay, H. Parnas, and E. Shamir. Neuronal growth via hybrid system of self-growing and diffusion based grammar rules: I. Bulletin of Mathematical Biology, 57(2):205–227, March 1995. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02460616.

Miura:1995:CRI

[1344] Robert M. Miura and Mark Pernarowski. Correlations of rates of insulin release from islets and plateau fractions for β-cells. Bulletin of Mathematical Biology, 57(2):229–246, March 1995. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02460617.

Thomas:1995:DBB

[1345] René Thomas, Denis Thieffry, and Marcelle Kaufman. Dynamical behaviour of biological regulatory networks — I. Biological role of feedback loops and practical use of the concept of the loop-characteristic state. Bulletin of Mathematical Biology, 57(2):247–276, March 1995. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02460618.

Thieffry:1995:DBB

[1346] Denis Thieffry and René Thomas. Dynamical behaviour of biological regulatory networks — II. Immunity control in bacteriophage lambda. Bulletin of Mathematical Biology, 57(2):277–297, March 1995. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02460619.

Hillesdon:1995:DCG

[1347] A. J. Hillesdon, T. J. Pedley, and J. O. Kessler. The development of concentration gradients in a suspension of chemotactic bacteria. Bulletin of Mathematical Biology, 57(2):299–344, March 1995. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02460620.

Michelson:1995:ITG

[1348] Seth Michelson and John T. Leith. Interlocking triads of growth control in tumors. Bulletin of Mathematical Biology, 57(2):345–366, March 1995. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02460621.

Steel:1995:FSP

[1349] Mike Steel and Mike Charleston. Five surprising properties of parsimoniously colored trees. Bulletin of Mathematical Biology, 57(2): 367–375, March 1995. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460622.

Lubkin:1995:PSS

[1350] Sharon Lubkin. Phyllotaxis: a systemic study in plant morphogenesis. Bulletin of Mathematical Biology, 57(2):377–379, March 1995. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02460623.

Maghsoodi:1995:OPS

[1351] Yoosef Maghsoodi and Eric P. M. Grist. Optimal population stabilization and control using the Leslie matrix model. Bulletin of Mathematical Biology, 57(3):381–399, May 1995. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460631.

Kohlmeier:1995:SRC

[1352] C. Kohlmeier and W. Ebenhöh. The stabilizing role of cannibalism in a predator–prey system. Bulletin of Mathematical Biology, 57(3): 401–411, May 1995. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460632.

Bertram:1995:TPC

[1353] Richard Bertram, Manish J. Butte, Tim Kiemel, and Arthur Sherman. Topological and phenomenological classification of bursting oscillations. Bulletin of Mathematical Biology, 57(3):413–439, May 1995. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02460633.

Ishay:1995:BAL

[1354] Jacob S. Ishay, Ada H. Abes, Harry L. Chernobrov, Vered Barenholz-Paniry, Yisrael Parmet, and Camil Fuchs. Building activity and longevity of queenless groups of the Oriental hornet, Vespa orientalis. Bulletin of Mathematical Biology, 57(3):441–460, May 1995. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/article/10.1007/BF02460634.

Byrne:1995:MMT

[1355] H. M. Byrne and M. A. J. Chaplain. Mathematical models for tumour angiogenesis: Numerical simulations and nonlinear wave solutions. Bulletin of Mathematical Biology, 57(3):461–486, May 1995. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02460635.

Barabe:1995:AGT

[1356] Denis Barabé and Roger V. Jean. On the allometric growth of tissues in fruits. Bulletin of Mathematical Biology, 57(3):487–498, May 1995. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02460636.

Mangel:1995:BR

[1357] Marc Mangel, George W. Swan, and Thomas G. Hallam. Book reviews. Bulletin of Mathematical Biology, 57(3):499–506, May 1995. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02460637.

Wilson:1995:EDL

[1358] W. G. Wilson, E. McCauley, and A. M. De Roos. Effect of dimensionality on Lotka–Volterra predator-prey dynamics: Individual based simulation results. Bulletin of Mathematical Biology, 57(4):507–526, July 1995. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02460780.

Luo:1995:FCN

[1359] Liaofu Luo, Fengmin Ji, and Hong Li. Fuzzy classification of nucleotide sequences and bacterial evolution. Bulletin of Mathematical Biology, 57 (4):527–537, July 1995. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460781.

Saunders:1995:RSS

[1360] Peter T. Saunders and Mae Wan Ho. Reliable segmentation by successive bifurcation. Bulletin of Mathematical Biology, 57(4):539–556, July 1995. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02460782.

Rognstad:1995:DAF

[1361] Robert Rognstad. Dicarboxylic acid fluxes during gluconeogenesis. No channelling of mitochondrial oxalacetate. Bulletin of Mathematical Biology, 57(4):557–568, July 1995. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460783.

Maki:1995:MAP

[1362] Laura Weiss Maki and Joel Keizer. Mathematical analysis of a proposed mechanism for oscillatory insulin secretion in perifused HIT-15 cells. Bulletin of Mathematical Biology, 57(4):569–591, July 1995. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02460784.

Lawson:1995:DAT

[1363] Linda M. Lawson, Yvette H. Spitz, Eileen E. Hofmann, and Robert Bryan Long. A data assimilation technique applied to a predator-prey model.

Bulletin of Mathematical Biology, 57(4):593–617, July 1995. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02460785.

Meinhardt:1995:BR

[1364] Hans Meinhardt, Paul Fife, Hal L. Smith, Diana E. Woodward, Alan Hastings, and Stuart G. Whittington. Book reviews. Bulletin of Mathematical Biology, 57(4):619–630, July 1995. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02460786.

Timm:1995:GIB

[1365] Uwe Timm and Akira Okubo. Gyrotaxis: Interaction between algae and flagellates. Bulletin of Mathematical Biology, 57(5):631– 650, September 1995. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02461844.

Stollberg:1995:DDL

[1366] Jes Stollberg. Density and diffusion limited aggregation in membranes. Bulletin of Mathematical Biology, 57(5):651–677, September 1995. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02461845.

Fauci:1995:SMP

[1367] Lisa J. Fauci and Amy McDonald. Sperm motility in the presence of boundaries. Bulletin of Mathematical Biology, 57(5):679– 699, September 1995. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02461846.

Gragnani:1995:UBD

[1368] Alessandra Gragnani and Sergio Rinaldi. A universal bifurcation diagram for seasonally perturbed predator–prey models. Bulletin of Mathematical Biology, 57(5):701–712, September 1995. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02461847.

Guan:1995:CMB

[1369] Shu-Guang Guan and An-Shen Qi. Contributions of memory b cells to secondary immune response. Bulletin of Mathematical Biology, 57(5): 713–731, September 1995. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02461848.

Michelson:1995:TEC

[1370] Seth Michelson and John T. Leith. A theoretical explanation of “concomitant resistance”. Bulletin of Mathematical Biology, 57(5):733– 747, September 1995. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02461849.

Fuxman:1995:RRF

[1371] Yakov L. Fuxman. Reproduction rate, feeding process, and Liebig limitations in cell populations — II. Bulletin of Mathematical Biology, 57(5): 749–782, September 1995. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02461850.

Anonymous:1995:ADF

[1372] Anonymous. An alternative derivation of Fuxman’s threshold result. Bulletin of Mathematical Biology, 57(5):783–786, September 1995. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02461851.

Durisova:1995:BSM

[1373] M. Ďurišová, L. Dedík, and M. Balan. Building a structured model of a complex pharmacokinetic system with time delays. Bulletin of Mathematical Biology, 57(6):787–808, November 1995. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/article/10.1007/BF02458295.

Borisyuk:1995:DBT

[1374] Galina N. Borisyuk, Roman M. Borisyuk, Alexander I. Khibnik, and Dirk Roose. Dynamics and bifurcations of two coupled neural oscillators with different connection types. Bulletin of Mathematical Biology, 57(6): 809–840, November 1995. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458296.

Sulzer:1995:IRB

[1375] Bernhard Sulzer and Gerard Weisbuch. Idiotypic regulation of b cell differentiation. Bulletin of Mathematical Biology, 57(6):841– 864, November 1995. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458297.

Poznanski:1995:ECB

[1376] R. R. Poznanski, W. G. Gibson, and M. R. Bennett. Electrotonic coupling between two CA3 hippocampal pyramidal neurons: a distributed cable model with somatic gap-junction. Bulletin of Mathematical Biology, 57(6):865–881, November 1995. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF03354909.

Kempner:1995:MRT

[1377] E. S. Kempner. The mathematics of radiation target analyses. Bulletin of Mathematical Biology, 57(6):883–898, November 1995. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02458298.

Rush:1995:PCL

[1378] Maureen E. Rush and John Rinzel. The potassium A-current, low firing rates and rebound excitation in Hodgkin–Huxley models. Bulletin of Mathematical Biology, 57(6):899–929, November 1995. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02458299.

Overholser:1995:EDP

[1379] K. A. Overholser, N. A. Lomangino, T. R. Harris, J. D. Bradley, and S. Bosan. Erratum to Deduction of pulmonary microvascular hematocrit from indicator dilution curves. Bulletin of Mathematical Biology, 57(6): 931–933, November 1995. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458300;   http://link.springer.com/content/pdf/10.
1007/BF02458300.pdf.

Ermentrout:1995:TMB

[1380] G. B. Ermentrout. Theoretical models in biology: The origin of life, the immune system, and the brain. Bulletin of Mathematical Biology, 57(6): 935–937, November 1995. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458301.

Dixon:1995:FS

[1381] Paul Dixon, Alistair Hobday, and George Sugihara. Fractals in science. Bulletin of Mathematical Biology, 57(6):939–941, November 1995. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02458302.

Hastings:1995:CDS

[1382] Alan Hastings. Chaos in dynamical systems. Bulletin of Mathematical Biology, 57(6):943–944, November 1995. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02458303.

Ellner:1995:CCA

[1383] Stephen Ellner. Coping with chaos: Analysis of chaotic data and the exploitation of chaotic systems. Bulletin of Mathematical Biology, 57(6): 945–946, November 1995. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458304.

Harrison:1995:MEP

[1384] Lionel G. Harrison. Models for embryonic periodicity. Bulletin of Mathematical Biology, 57(6):947–948, November 1995. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/article/10.1007/BF02458305.

Denny:1995:LMF

[1385] Mark W. Denny. Life in moving fluids: The physical biology of flow. Bulletin of Mathematical Biology, 57(6):949–951, November 1995. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02458306.

Taylor:1996:NLG

[1386] William R. Taylor. A non-local gap-penalty for profile alignment. Bulletin of Mathematical Biology, 58(1):1–18, January 1996. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02458279.

Manjabacas:1996:KAZ

[1387] M. C. Manjabacas, E. Valero, M. García-Moreno, C. Garrido, and R. Varón. Kinetics of an autocatalytic zymogen reaction in the presence of an inhibitor coupled to a monitoring reaction. Bulletin of Mathematical Biology, 58(1):19–41, January 1996. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02458280.

Borghans:1996:EQS

[1388] José A. M. Borghans, Rob J. de Boer, and Lee A. Segel. Extending the quasi-steady state approximation by changing variables. Bulletin of Mathematical Biology, 58(1):43–63, January 1996. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02458281.

Davis:1996:TEI

[1389] Brian K. Davis. A theory of evolution that includes prebiotic selforganization and episodic species formation. Bulletin of Mathematical Biology, 58(1):65–97, January 1996. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02458282.

Davis:1996:DES

[1390] Brian K. Davis. Darwinian evolution by self-propagating species under fractional order kinetics. Bulletin of Mathematical Biology, 58(1): 99–101, January 1996. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458283.

Ellner:1996:EFM

[1391] Stephen Ellner. Environmental fluctuations and the maintenance of genetic diversity in age or stage-structured populations. Bulletin of Mathematical Biology, 58(1):103–127, January 1996. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02458284.

Demiray:1996:EPS

[1392] Hilmi Demiray. An evolution of pulse speed in arteries. Bulletin of Mathematical Biology, 58(1):129–140, January 1996. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02458285.

Manivannan:1996:DIC

[1393] K. Manivannan, R. T. Mathias, and E. Gudowska-Nowak. Description of interacting channel gating using a stochastic Markovian model. Bulletin of Mathematical Biology, 58(1):141–174, January 1996. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02458286.

Behera:1996:VPE

[1394] Narayan Behera. Variational principles in evolution. Bulletin of Mathematical Biology, 58(1):175–202, January 1996. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02458287.

Fogel:1996:BR

[1395] David B. Fogel and Mark A. Lewis. Book reviews. Bulletin of Mathematical Biology, 58(1):203–206, January 1996. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02458288.

Cosner:1996:VVC

[1396] Chris Cosner. Variability, vagueness and comparison methods for ecological models. Bulletin of Mathematical Biology, 58(2):207–246, March 1996. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458307.

Jaworska:1996:CMCa

[1397] Joanna S. Jaworska, Thomas G. Hallam, and T. Wayne Schultz. A community model of ciliate tetrahymena and bacteria E. coli: Part I. Individual-based models of Tetrahymena and E. coli populations. Bulletin of Mathematical Biology, 58(2):247–264, March 1996. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02458308.

Jaworska:1996:CMCb

[1398] Joanna S. Jaworska, Thomas G. Hallam, and T. Wayne Schultz. A community model of ciliate Tetrahymena and bacteria E. coli: Part II. Interactions in a batch system. Bulletin of Mathematical Biology, 58 (2):265–283, March 1996. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458309.

deBoer:1996:NBS

[1399] Rob J. de Boer, Maarten C. Boerlijst, Bernhard Sulzer, and Alan S. Perelson. A new bell-shaped function for idiotypic interactions based on cross-linking. Bulletin of Mathematical Biology, 58(2):285–312, March 1996. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458310.

Bookstein:1996:BBM

[1400] Fred L. Bookstein. Biometrics, biomathematics and the morphometric synthesis. Bulletin of Mathematical Biology, 58(2):313–365, March 1996. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02458311.

Kirschner:1996:MTS

[1401] Denise Kirschner and G. F. Webb. A model for treatment strategy in the chemotherapy of AIDS. Bulletin of Mathematical Biology, 58(2): 367–390, March 1996. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458312.

Charlesworth:1996:BR

[1402] Brian Charlesworth, Michael Mesterton-Gibbons, Michael Neubert, Sarah Little, J. O. Kessler, J. Demongeot, Steven H. Strogatz, and Glenn T. Vickers. Book reviews. Bulletin of Mathematical Biology, 58 (2):391–407, March 1996. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458313.

Nekka:1996:MGV

[1403] F. Nekka, S. Kyriacos, C. Kerrigan, and L. Cartilier. A model of growing vascular structures. Bulletin of Mathematical Biology, 58(3): 409–424, May 1996. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460590.

Panetta:1996:MMP

[1404] John Carl Panetta. A mathematical model of periodically pulsed chemotherapy: Tumor recurrence and metastasis in a competitive environment. Bulletin of Mathematical Biology, 58(3):425–447, May 1996. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02460591.

Schmidt:1996:DMB

[1405] W. Schmidt and E.-Ch. Müller. A distance measure based on binary character data and its application to phylogeny reconstruction. Bulletin of Mathematical Biology, 58(3):449–469, May 1996. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02460592.

Britton:1996:MMP

[1406] N. F. Britton, L. W. Partridge, and N. R. Franks. A mathematical model for the population dynamics of army ants. Bulletin of Mathematical Biology, 58(3):471–492, May 1996. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02460593.

Vermeulen:1996:CSI

[1407] A. Vermeulen, J.-P. Rospars, P. Lánský, and H. C. Tuckwell. Coding of stimulus intensity in an olfactory receptor neuron: Role of neuron spatial extent and passive dendritic backpropagation of action potentials. Bulletin of Mathematical Biology, 58(3):493–512, May 1996. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02460594.

Thompson:1996:MGMa

[1408] C. R. Thompson, B. S. Gerstman, S. L. Jacques, and M. E. Rogers. Melanin granule model for laser-induced thermal damage in the retina. Bulletin of Mathematical Biology, 58(3):513–553, May 1996. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02460595.

Grasman:1996:EET

[1409] Johan Grasman. The expected extinction time of a population within a system of interacting biological populations. Bulletin of Mathematical Biology, 58(3):555–568, May 1996. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02460596.

Morel:1996:MSC

[1410] Benoît F. Morel, Meghan A. Burke, Jayant Kalagnanam, Susan A. Mc-Carthy, David J. Tweardy, and Penelope A. Morel. Making sense of the combined effect of interleukin-2 and interleukin-4 on lymphocytes using a mathematical model. Bulletin of Mathematical Biology, 58 (3):569–594, May 1996. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460597.

Wolkowicz:1996:BR

[1411] Gail S. K. Wolkowicz, Rick Durrett, Mercedes Pascual, Arthur L. Koch, Yu. A. Kuznetsov, and W. M. Schaffer. Book reviews. Bulletin of Mathematical Biology, 58(3):595–609, May 1996. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02460598.

Armitstead:1996:SBB

[1412] J. P. Armitstead, C. D. Bertram, and O. E. Jensen. A study of the bifurcation behaviour of a model of flow through a collapsible tube. Bulletin of Mathematical Biology, 58(4):611–641, July 1996. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02459476.

Ruxton:1996:DDM

[1413] Graeme D. Ruxton. Density-dependent migration and stability in a system of linked populations. Bulletin of Mathematical Biology, 58 (4):643–660, July 1996. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459477.

Bonabeau:1996:MMS

[1414] Eric Bonabeau, Guy Theraulaz, and Jean-Louis Deneubourg. Mathematical model of self-organizing hierarchies in animal societies. Bulletin of Mathematical Biology, 58(4):661–717, July 1996. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02459478.

Costalat:1996:EEO

[1415] Robert Costalat and Jacques Burger. Effect of enzyme organization on the stability of Yates–Pardee pathways. Bulletin of Mathematical Biology, 58(4):719–737, July 1996. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02459479.

Stocker:1996:FCD

[1416] Sabine Stöcker and Daniel Weihs. Foraging costs as a defining factor for growth and asymptotic mass. Bulletin of Mathematical Biology, 58 (4):739–751, July 1996. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459480.

Mayer:1996:PFR

[1417] Stefan Mayer. On the pressure and flow-rate distributions in tree-like and arterial-venous networks. Bulletin of Mathematical Biology, 58 (4):753–785, July 1996. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459481.

Olsen:1996:MMF

[1418] Luke Olsen, Jonathan A. Sherratt, and Philip K. Maini. A mathematical model for fibro-proliferative wound healing disorders. Bulletin of Mathematical Biology, 58(4):787–808, July 1996. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02459482.

Thompson:1996:MGMb

[1419] C. R. Thompson, B. S. Gerstman, S. L. Jacques, and M. E. Rogers. Melanin granule model for laser-induced thermal damage in the retina. Bulletin of Mathematical Biology, 58(4):809–810, July 1996. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02459483;
http://link.springer.com/content/pdf/10.1007/BF02459483.pdf.

Renshaw:1996:BR

[1420] Eric Renshaw, Jon Seger, and J. W. Stubblefield Cera. Book reviews. Bulletin of Mathematical Biology, 58(4):811–814, July 1996. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/article/10.1007/BF02459484.

Allen:1996:MMW

[1421] L. J. S. Allen, E. J. Allen, and S. Ponweera. A mathematical model for weed dispersal and control. Bulletin of Mathematical Biology, 58(5): 815–834, September 1996. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459485.

King:1996:WDP

[1422] A. A. King, W. M. Schaffer, C. Gordon, J. Treat, and M. Kot. Weakly dissipative predator-prey systems. Bulletin of Mathematical Biology, 58(5): 835–859, September 1996. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459486.

Mesterton-Gibbons:1996:CRN

[1423] Michael Mesterton-Gibbons and Michael J. Childress. Constraints on reciprocity for non-sessile organisms. Bulletin of Mathematical Biology, 58(5):861–875, September 1996. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02459487.

Happel:1996:ANT

[1424] Robert Happel, Robert Hecht, and Peter F. Stadler. Autocatalytic networks with translation. Bulletin of Mathematical Biology, 58(5):877– 905, September 1996. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459488.

Wilson:1996:APC

[1425] William G. Wilson, Roger M. Nisbet, Alex H. Ross, Carlos Robles, and Robert A. Desharnais. Abrupt population changes along smooth environmental gradients. Bulletin of Mathematical Biology, 58(5):907– 922, September 1996. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459489.

Berg:1996:PMA

[1426] Paul H. Berg, Eberhard O. Voit, and Roger L. White. A pharmacodynamic model for the action of the antibiotic imipenem on Pseudomonas aeruginosa populations in vitro. Bulletin of Mathematical Biology, 58(5): 923–938, September 1996. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459490.

Demiray:1996:SWP

[1427] Hilmi Demiray. Solitary waves in prestressed elastic tubes. Bulletin of Mathematical Biology, 58(5):939–955, September 1996. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02459491.

Kmet:1996:MRCa

[1428] Tibor Kmeť. Material recycling in a closed aquatic ecosystem. I. Nitrogen transformation cycle and preferential utilization of ammonium to nitrate by phytoplankton as an optimal control problem. Bulletin of Mathematical Biology, 58(5):957–982, September 1996. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02459492.

Kmet:1996:MRCb

[1429] Tibor Kmeť. Material recycling in a closed aquatic ecosystem. II. Bifurcation analysis of a simple food-chain model. Bulletin of Mathematical Biology, 58(5):983–1000, September 1996. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02459493.

Kuperman:1996:PWS

[1430] M. N. Kuperman, B. Von Haeften, and H. S. Wio. Persistence of the weaker species in a non-homogeneous competitive system: Exact result through a quantum mechanical analogy. Bulletin of Mathematical Biology, 58(5):1001–1018, September 1996. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02459494.

Hethcote:1996:EMT

[1431] Herbert W. Hethcote. Epidemic models: Their structure and relation to data. Bulletin of Mathematical Biology, 58(5):1019–1022, September 1996. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459495.

She:1996:NSC

[1432] J. She and C. D. Bertram. Numerical simulation of collapsible-tube flows with sinusoidal forced oscillations. Bulletin of Mathematical Biology, 58(6):1023–1046, November 1996. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02458382.

Walker:1996:DDB

[1433] Wynn L. Walker and Julian Cook. Drug delivery to brain tumors. Bulletin of Mathematical Biology, 58(6):1047–1074, November 1996. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02458383.

Aharon:1996:PCE

[1434] S. Aharon, M. Bercovier, and H. Parnas. Parallel computation enables precise description of ca^{2+} distribution in nerve terminals. Bulletin of Mathematical Biology, 58(6):1075–1097, November 1996. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02458384.

Fishman:1996:MIA

[1435] Michael A. Fishman and Lee A. Segel. Modeling immunotherapy for allergy. Bulletin of Mathematical Biology, 58(6):1099–1121, November 1996. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458385.

B-Rao:1996:IAE

[1436] Chandrika B-Rao and John Stewart. Inverse analysis of empirical matrices of idiotypic network interactions. Bulletin of Mathematical Biology, 58(6):1123–1153, November 1996. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02458386.

LeSech:1996:DPM

[1437] C. Le Sech and A. Capderou. Determination of pulmonary mean transit time and cardiac output using a one-dimensional model. Bulletin of Mathematical Biology, 58(6):1155–1170, November 1996. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/article/10.1007/BF02458387.

Wu:1996:NIC

[1438] Richard Wu and Tai Te Wu. A novel intact circular dsDNA supercoil. Bulletin of Mathematical Biology, 58(6):1171–1185, November 1996. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/BF02458388.

Barradas:1996:CDC

[1439] Ignacio Barradas, Hal Caswell, and Joel E. Cohen. Competition during colonization vs competition after colonization in disturbed environments: a metapopulation approach. Bulletin of Mathematical Biology, 58(6): 1187–1207, November 1996. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458389.

Smith:1997:CTR

[1440] Charles E. Smith, Petr Lánsky, and Te-Hsin Lung. Cycle-time and residence-time density approximations in a stochastic model for circulatory transport. Bulletin of Mathematical Biology, 59(1):1–22, January 1997. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
BF02459468;   https://link.springer.com/content/pdf/10.1007/
BF02459468.pdf.

Sirot:1997:ASH

[1441] E. Sirot and V. Křivan. Adaptive superparasitism and host-parasitoid dynamics. Bulletin of Mathematical Biology, 59(1):23–41, January 1997. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02459469;
https://link.springer.com/content/pdf/10.1007/BF02459469.pdf.

Popescu:1997:SIP

[1442] Dumitru Popescu, Liviu Movileanu, Gheorghe Victor, and Grigore Turcu. Stability and instability properties of aggregation of single chain amphiphiles into binary mixtures. Bulletin of Mathematical Biology, 59 (1):43–61, January 1997. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459470;   https://link.springer.com/content/pdf/10.
1007/BF02459470.pdf.

Reyment:1997:ESO

[1443] Richard A. Reyment. Evolution of shape in Oligocene and Miocene Notocarinovalva (Ostracoda, Crustacea): a multivariate statistical study. Bulletin of Mathematical Biology, 59(1):63–87, January 1997. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02459471;
https://link.springer.com/content/pdf/10.1007/BF02459471.pdf.

Gaver:1997:MMI

[1444] D. P. Gaver, P. A. Jacobs, R. L. Carpenter, and J. G. Burkhart. A mathematical model for intracellular effects of toxins on DNA adduction and repair. Bulletin of Mathematical Biology, 59(1):89–106, January 1997. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02459472;
https://link.springer.com/content/pdf/10.1007/BF02459472.pdf.

VanKirk:1997:IMP

[1445] R. W. Van Kirk and M. A. Lewis. Integrodifference models for persistence in fragmented habitats. Bulletin of Mathematical Biology, 59(1):107–137, January 1997. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
BF02459473;   https://link.springer.com/content/pdf/10.1007/
BF02459473.pdf.

Kaneko:1997:IDT

[1446] Kunihiko Kaneko and Tetsuya Yomo. Isologous diversification: a theory of cell differentiation. Bulletin of Mathematical Biology, 59(1):139–196, January 1997. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
BF02459474;   https://link.springer.com/content/pdf/10.1007/
BF02459474.pdf.

Olsen:1997:BR

[1447] Lars Folke Olsen and David B. Fogel. Book reviews. Bulletin of Mathematical Biology, 59(1):197–204, January 1997. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02459475;
https://link.springer.com/content/pdf/10.1007/BF02459475.pdf.

Costa:1997:CTS

[1448] M. I. S. Costa and J. L. Boldrini. Chemotherapeutic treatments: a study of the interplay among drug resistance, toxicity and recuperation from side effects. Bulletin of Mathematical Biology, 59(2): 205–232, March 1997. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462001;   http://link.springer.com/content/pdf/10.
1007/BF02462001.pdf.

Michelson:1997:PFA

[1449] Seth Michelson and John T. Leith. Positive feedback and angiogenesis in tumor growth control. Bulletin of Mathematical Biology, 59(2): 233–254, March 1997. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462002;   http://link.springer.com/content/pdf/10.
1007/BF02462002.pdf.

Jensen:1997:MPM

[1450] A. L. Jensen. Matrix population model with density-dependent recruitment for assessment of age-structured wildlife populations. Bulletin of Mathematical Biology, 59(2):255–262, March 1997. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02462003;
http://link.springer.com/content/pdf/10.1007/BF02462003.pdf.

Lefever:1997:OTB

[1451] R. Lefever and O. Lejeune. On the origin of tiger bush. Bulletin of Mathematical Biology, 59(2):263–294, March 1997. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02462004;   http://link.springer.com/
content/pdf/10.1007/BF02462004.pdf.

Kaczmarek:1997:HHS

[1452] Mariusz Kaczmarek, Ravi P. Subramaniam, and Samuel R. Neff. The hydromechanics of hydrocephalus: Steady-state solutions for cylindrical geometry. Bulletin of Mathematical Biology, 59(2):295–323, March 1997. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02462005;
http://link.springer.com/content/pdf/10.1007/BF02462005.pdf.

Liu:1997:TBP

[1453] Bing Liu and Rainer K. Sachs. A two-backbone polymer model for interphase chromosome geometry. Bulletin of Mathematical Biology, 59 (2):325–337, March 1997. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02462006;   http://link.springer.com/content/pdf/10.
1007/BF02462006.pdf.

Reidys:1997:GPC

[1454] Christian Reidys, Peter F. Stadler, and Peter Schuster. Generic properties of combinatory maps: Neutral networks of RNA secondary structures. Bulletin of Mathematical Biology, 59(2):339–397, March 1997. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02462007;
http://link.springer.com/content/pdf/10.1007/BF02462007.pdf.

Ermentrout:1997:BR

[1455] G. Bard Ermentrout and Svetlozar T. Rachev. Book reviews. Bulletin of Mathematical Biology, 59(2):399–406, March 1997. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02462008;
http://link.springer.com/content/pdf/10.1007/BF02462008.pdf.

Anonymous:1997:E

[1456] Anonymous. Erratum. Bulletin of Mathematical Biology, 59(2):407, March 1997. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
BF02462009;   http://link.springer.com/content/pdf/10.1007/
BF02462009.pdf.

Monk:1997:CEE

[1457] Nicholas A. M. Monk. The community effect and ectoderm–mesoderm interaction in Xenopus muscle differentiation. Bulletin of Mathematical Biology, 59(3):409–425, May 1997. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02459458;   http://link.springer.com/
content/pdf/10.1007/BF02459458.pdf.

Kaitala:1997:HPD

[1458] Veijo Kaitala, Mikko Heino, and Wayne M. Getz. Host-parasite dynamics and the evolution of host immunity and parasite fecundity strategies. Bulletin of Mathematical Biology, 59(3):427–450, May 1997. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02459459;
http://link.springer.com/content/pdf/10.1007/BF02459459.pdf.

Weinstein:1997:DCH

[1459] Alan M. Weinstein. Dynamics of cellular homeostasis: Recovery time for a perturbation from equilibrium. Bulletin of Mathematical Biology, 59(3):451–481, May 1997. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459460;   http://link.springer.com/content/pdf/10.
1007/BF02459460.pdf.

Naus:1997:MAM

[1460] Joseph I. Naus and Ke-Ning Sheng. Matching among multiple random sequences. Bulletin of Mathematical Biology, 59(3):483–496, May 1997. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02459461;
http://link.springer.com/content/pdf/10.1007/BF02459461.pdf.

Doebeli:1997:CSC

[1461] Michael Doebeli and Graeme D. Ruxton. Controlling spatial chaos in metapopulations with long-range dispersal. Bulletin of Mathematical Biology, 59(3):497–515, May 1997. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02459462;   http://link.springer.com/
content/pdf/10.1007/BF02459462.pdf.

Barrio:1997:HCU

[1462] Rafael A. Barrio, Limei Zhang, and Philip K. Maini. Hierarchically coupled ultradian oscillators generating robust circadian rhythms. Bulletin of Mathematical Biology, 59(3):517–532, May 1997. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02459463;
http://link.springer.com/content/pdf/10.1007/BF02459463.pdf.

Ermentrout:1997:PFS

[1463] Bard Ermentrout and Mark Lewis. Pattern formation in systems with one spatially distributed species. Bulletin of Mathematical Biology, 59 (3):533–549, May 1997. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459464;   http://link.springer.com/content/pdf/10.
1007/BF02459464.pdf.

VandenBosch:1997:CAS

[1464] Frank Van den Bosch and Wilfried Gabriel. Cannibalism in an agestructured predator–prey system. Bulletin of Mathematical Biology, 59 (3):551–567, May 1997. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459465;   http://link.springer.com/content/pdf/10.
1007/BF02459465.pdf.

Lansky:1997:IMA

[1465] Petr Lánský, Jana Velíšková, and Libor Velíšek. An indirect method for absorption rate estimation: Flurothyl-induced seizures. Bulletin of Mathematical Biology, 59(3):569–579, May 1997. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02459466;   http://link.springer.com/
content/pdf/10.1007/BF02459466.pdf.

Tuffley:1997:LBM

[1466] Chris Tuffley and Mike Steel. Links between maximum likelihood and maximum parsimony under a simple model of site substitution. Bulletin of Mathematical Biology, 59(3):581–607, May 1997. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02459467;
http://link.springer.com/content/pdf/10.1007/BF02459467.pdf.

White:1997:PTE

[1467] Peter White and James Powell. Phase transition from environmental to dynamic determinism in mountain pine beetle attack. Bulletin of Mathematical Biology, 59(4):609–643, July 1997. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02458422;
https://link.springer.com/content/pdf/10.1007/BF02458422.pdf.

Rakocevic:1997:TCA

[1468] Miloje M. Rakočević. Two classes of the aminoacyl-tRNA synthetases in correspondence with the codon path cube. Bulletin of Mathematical Biology, 59(4):645–648, July 1997. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02458423;   https://link.springer.com/content/
pdf/10.1007/BF02458423.pdf.

Alexeyev:1997:SMS

[1469] V. L. Alexeyev and A. P. Levich. A search for maximum species abundances in ecological communities under conditional diversity optimization. Bulletin of Mathematical Biology, 59(4):649–677, July 1997. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02458424;
https://link.springer.com/content/pdf/10.1007/BF02458424.pdf.

Gyllenberg:1997:CVD

[1470] Mats Gyllenberg, Ilkka Hanski, and Torsten Lindström. Continuous versus discrete single species population models with adjustable reproductive strategies. Bulletin of Mathematical Biology, 59(4):679–705, July 1997. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
BF02458425;   https://link.springer.com/content/pdf/10.1007/
BF02458425.pdf.

Costa:1997:COC

[1471] M. I. S. Costa and J. L. Boldrini. Conflicting objectives in chemotherapy with drug resistance. Bulletin of Mathematical Biology, 59(4):707– 724, July 1997. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
BF02458426;   https://link.springer.com/content/pdf/10.1007/
BF02458426.pdf.

Lund:1997:GTH

[1472] Ole Lund, Ole Søgaard Lund, Gregers Gram, Susanne Dam Nielsen, Kristian Schønning, Jens Ole Nielsen, John-Erik Stig Hansen, and Erik Mosekilde. Gene therapy of T helper cells in HIV infection: Mathematical model of the criteria for clinical effect. Bulletin of Mathematical Biology, 59(4):725–745, July 1997. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02458427;
https://link.springer.com/content/pdf/10.1007/BF02458427.pdf.

Peleg:1997:MID

[1473] Micha Peleg, Mark D. Normand, and Evi Damrau. Mathematical interpretation of dose-response curves. Bulletin of Mathematical Biology, 59(4):747–761, July 1997. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02458428;   https://link.springer.com/content/pdf/10.
1007/BF02458428.pdf.

Kirschner:1997:UDR

[1474] Denise E. Kirschner and G. F. Webb. Understanding drug resistance for monotherapy treatment of HIV infection. Bulletin of Mathematical Biology, 59(4):763–785, July 1997. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/BF02458429;   https://link.springer.com/content/
pdf/10.1007/BF02458429.pdf.

Hoppensteadt:1997:BR

[1475] Frank Hoppensteadt, Harold M. Hastings, Michael Mesterton-Gibbons, K. A. Jane White, Henry D. I. Abarbanel, Michael G. Neubert, and Mark Pernarowski. Book reviews. Bulletin of Mathematical Biology, 59(4):787– 807, July 1997. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
BF02458430;   https://link.springer.com/content/pdf/10.1007/
BF02458430.pdf.

Krivan:1997:DCO

[1476] Vlastimil Křivan. Dynamical consequences of optimal host feeding on host-parasitoid population dynamics. Bulletin of Mathematical Biology, 59(5):809–831, September 1997. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02459994;   http://link.springer.com/
content/pdf/10.1007/BF02459994.pdf.

Galvez:1997:MMA

[1477] Jesus Gálvez, Francisco Lajarin, and Pilar García-Peñarrubia. Mathematical modeling of adhesion of bacteria to host cell lines. Bulletin of Mathematical Biology, 59(5):833–856, September 1997. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02459995;
http://link.springer.com/content/pdf/10.1007/BF02459995.pdf.

Galvez:1997:PHC

[1478] Jesus Gálvez, Francisco Lajarin, and Pilar García-Peñarrubia. Penetration of host cell lines by bacteria. Characteristics of the process of intracellular bacterial infection. Bulletin of Mathematical Biology, 59(5): 857–879, September 1997. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459996;   http://link.springer.com/content/pdf/10.
1007/BF02459996.pdf.

Bittner:1997:VLA

[1479] Barbara Bittner, Sebastian Bonhoeffer, and Martin A. Nowak. Virus load and antigenic diversity. Bulletin of Mathematical Biology, 59(5): 881–896, September 1997. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.

1007/BF02459997;   http://link.springer.com/content/pdf/10.
1007/BF02459997.pdf.

Kochel:1997:NMX

[1480] Bonawentura Kochel and Waldemar Sajewicz. A new measure of xenobiotic toxicity to the first-line human defence system from the time-resolved phagocyte luminescence. Bulletin of Mathematical Biology, 59(5):897– 910, September 1997. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02459998;   http://link.springer.com/content/pdf/10.
1007/BF02459998.pdf.

Weiss:1997:VDM

[1481] Michael Weiss, Christina Stedtler, and Michael S. Roberts. On the validity of the dispersion model of hepatic drug elimination when intravascular transit time densities are long-tailed. Bulletin of Mathematical Biology, 59(5):911–929, September 1997. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02459999;   http://link.springer.com/
content/pdf/10.1007/BF02459999.pdf.

Broom:1997:MPM

[1482] M. Broom, C. Cannings, and G. T. Vickers. Multi-player matrix games. Bulletin of Mathematical Biology, 59(5):931–952, September 1997. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02460000;
http://link.springer.com/content/pdf/10.1007/BF02460000.pdf.

Zoppou:1997:DHM

[1483] C. Zoppou, S. I. Barry, and G. N. Mercer. Dynamics of human milk extraction: a comparative study of breast feeding and breast pumping. Bulletin of Mathematical Biology, 59(5):953–973, September 1997. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02460001;
http://link.springer.com/content/pdf/10.1007/BF02460001.pdf.

Almirantis:1997:CSP

[1484] Yannis Almirantis and Astero Provata. The “clustered structure” of the purines/pyrimidines distribution in DNA distinguishes systematically between coding and non-coding sequences. Bulletin of Mathematical Biology, 59(5):975–992, September 1997. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/BF02460002;   http://link.springer.com/
content/pdf/10.1007/BF02460002.pdf.

Demiray:1997:ESS

[1485] Hilmi Demiray. The effect of shear stress on solitary waves in arteries. Bulletin of Mathematical Biology, 59(5):993–1012, September 1997. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02460003;
http://link.springer.com/content/pdf/10.1007/BF02460003.pdf.

Fox:1997:BR

[1486] Gordon A. Fox, Gregg Hartvigsen, W. M. Schaffer, Jonathan A. Sherratt, R. V. O’Neill, and Athel Cornish-Bowden. Book reviews. Bulletin of Mathematical Biology, 59(5):1013–1028, September 1997. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02460004;
http://link.springer.com/content/pdf/10.1007/BF02460004.pdf.

Suciu:1997:MAA

[1487] A. Suciu, G. Civelekoğlu, Y. Tardy, and J.-J. Meister. Model for the alignment of actin filaments in endothelial cells subjected to fluid shear stress. Bulletin of Mathematical Biology, 59(6):1029–1046, November 1997. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
BF02460100;   https://link.springer.com/content/pdf/10.1007/
BF02460100.pdf.

Henery:1997:QCP

[1488] R. Henery, W. G. Gibson, and M. R. Bennett. Quantal currents and potential in the three-dimensional anisotropic bidomain model of smooth muscle. Bulletin of Mathematical Biology, 59(6):1047–1075, November 1997. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
BF02460101;   https://link.springer.com/content/pdf/10.1007/
BF02460101.pdf.

Dale:1997:RFM

[1489] Paul D. Dale, Jonathan A. Sherratt, and Philip K. Maini. Role of fibroblast migration in collagen fiber formation during fetal and adult dermal wound healing. Bulletin of Mathematical Biology, 59(6):1077–1100, November 1997. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/

BF02460102;   https://link.springer.com/content/pdf/10.1007/
BF02460102.pdf.

Truscott:1997:AAE

[1490] J. E. Truscott, C. R. Webb, and C. A. Gilligan. Asymptotic analysis of an epidemic model with primary and secondary infection. Bulletin of Mathematical Biology, 59(6):1101–1123, November 1997. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02460103;
https://link.springer.com/content/pdf/10.1007/BF02460103.pdf.

Jafri:1997:AIC

[1491] M. Saleet Jafri and Joel Keizer. Agonist-induced calcium waves in oscillatory cells: a biological example of Burgers’ equation. Bulletin of Mathematical Biology, 59(6):1125–1144, November 1997. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02460104;
https://link.springer.com/content/pdf/10.1007/BF02460104.pdf.

Varga:1997:EPP

[1492] Zoltán Varga and Eörs Szathmáry. An extremum principle for parabolic competition. Bulletin of Mathematical Biology, 59(6):1145–1154, November 1997. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
BF02460105;   https://link.springer.com/content/pdf/10.1007/
BF02460105.pdf.

Knudsen:1997:QMM

[1493] Zuzana Knudsen, A. V. Holden, and J. Brindley. Qualitative modeling of mechanoelectrical feedback in a ventricular cell. Bulletin of Mathematical Biology, 59(6):1155–1181, November 1997. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02460106;
https://link.springer.com/content/pdf/10.1007/BF02460106.pdf.

Monk:1997:CCR

[1494] Nicholas A. M. Monk. Cell communities and robustness in development. Bulletin of Mathematical Biology, 59(6):1183–1189, November 1997. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/BF02460107;
https://link.springer.com/content/pdf/10.1007/BF02460107.pdf.

Sneyd:1997:BR

[1495] James Sneyd, Mark A. J. Chaplain, Edward F. Connor, and Stephen Cantrell. Book reviews. Bulletin of Mathematical Biology, 59(6):1191– 1201, November 1997. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/BF02460108;   https://link.springer.com/content/pdf/10.
1007/BF02460108.pdf.

Myerscough:1998:PFG

[1496] M. R. Myerscough, P. K. Maini, and K. J. Painter. Pattern formation in a generalized chemotactic model. Bulletin of Mathematical Biology, 60(1): 1–26, January 1998. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1006/bulm.1997.0010.pdf.

Alvarez:1998:MCM

[1497] Sergio A. Alvarez, Svetlana Levitan, and James A. Reggia. Metrics for cortical map organization and lateralization. Bulletin of Mathematical Biology, 60(1):27–47, January 1998. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1006/bulm.1997.0002.pdf.

Supriatna:1998:OHP

[1498] Asep K. Supriatna and Hugh P. Possingham. Optimal harvesting for a predator–prey metapopulation. Bulletin of Mathematical Biology, 60(1): 49–65, January 1998. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1006/bulm.1997.0005.pdf.

Cerf:1998:NTM

[1499] Corinne Cerf. A note on the tangle model for DNA recombination. Bulletin of Mathematical Biology, 60(1):67–78, January 1998. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/content/pdf/10.1006/bulm.1997.0024.pdf.

Louie:1998:CMA

[1500] K. Louie, A. J. Hall, and P. W. Gandar. A continuum mechanics approach to determining the cellular velocity field within a wool follicle. Bulletin of Mathematical Biology, 60(1):79–100, January 1998. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/content/pdf/10.1006/bulm.1997.0025.pdf.

Dallon:1998:MMF

[1501] John C. Dallon and Jonathan A. Sherratt. A mathematical model for fibroblast and collagen orientation. Bulletin of Mathematical Biology, 60(1):101–129, January 1998. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1006/bulm.1997.0027.pdf.

vandenBerg:1998:MMV

[1502] Hugo Antonius van den Berg. Modelling the metabolic versatility of a microbial trichome. Bulletin of Mathematical Biology, 60(1):131–150, January 1998. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1006/bulm.1997.0026.pdf.

Zheng:1998:STP

[1503] Q. Zheng. A stochastic two-phase growth model. Bulletin of Mathematical Biology, 60(1):151–161, January 1998. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/content/pdf/10.1006/bulm.1997.0029.pdf.

Arques:1998:EAM

[1504] Didier G. Arqués, Jean-Paul Fallot, and Christian J. Michel. An evolutionary analytical model of a complementary circular code simulating the protein coding genes, the 5’ and 3’ regions. Bulletin of Mathematical Biology, 60(1):163–194, January 1998. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1006/bulm.1997.0033.pdf.

Lange:1998:BR

[1505] Kenneth Lange. Book review. Bulletin of Mathematical Biology, 60(1): 195–196, January 1998. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/content/
pdf/10.1006/bulm.1997.0018.pdf.

Suhnel:1998:PDR

[1506] Jürgen Sühnel. Parallel dose-response curves in combination experiments. Bulletin of Mathematical Biology, 60(2):197–213, March 1998. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/content/pdf/10.1006/
bulm.1997.0009.pdf.

Swinton:1998:ETP

[1507] Jonathan Swinton. Extinction times and phase transitions for spatially structured closed epidemics. Bulletin of Mathematical Biology, 60(2): 215–230, March 1998. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/content/
pdf/10.1006/bulm.1997.0014.pdf.

Kooi:1998:ULE

[1508] B. W. Kooi, M. P. Boer, and S. A. L. M. Kooijman. On the use of the logistic equation in models of food chains. Bulletin of Mathematical Biology, 60(2):231–246, March 1998. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/content/pdf/10.1006/bulm.1997.0016.pdf.

Costantino:1998:RPC

[1509] R. F. Costantino, J. M. Cushing, Brian Dennis, Robert A. Desharnais, and Shandelle M. Henson. Resonant population cycles in temporally fluctuating habitats. Bulletin of Mathematical Biology, 60(2): 247–273, March 1998. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/content/
pdf/10.1006/bulm.1997.0017.pdf.

Spiros:1998:TMD

[1510] Athan Spiros and Leah Edelstein-Keshet. Testing a model for the dynamics of actin structures with biological parameter values. Bulletin of Mathematical Biology, 60(2):275–305, March 1998. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.

springer.com/content/pdf/10.1006/bulm.1997.0022.pdf.

Watmough:1998:DED

[1511] James Watmough, Mark Winston, and Keith Slessor. Distinguishing the effects of dilution and restricted movement on the intra-nest transmission of honey-bee queen pheromones. Bulletin of Mathematical Biology, 60 (2):307–327, March 1998. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/content/
pdf/10.1006/bulm.1997.0006.pdf.

Ackermann:1998:VDB

[1512] J. Ackermann, B. Wlotzka, and J. S. McCaskill. In vitro DNA-based predator-prey system with oscillatory kinetics. Bulletin of Mathematical Biology, 60(2):329–354, March 1998. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/content/pdf/10.1006/bulm.1997.0001.pdf.

Yang:1998:LID

[1513] Hyun Mo Yang and Ana Silvia Barreiros Silveira. The loss of immunity in directly transmitted infections modeling: Effects on the epidemiological parameters. Bulletin of Mathematical Biology, 60(2):355– 372, March 1998. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL https://link.springer.com/content/pdf/
10.1006/bulm.1997.0031.pdf.

Ventriglia:1998:CES

[1514] F. Ventriglia. Computational experiments support a competitive function in the the CA3 region of the hippocampus. Bulletin of Mathematical Biology, 60(2):373–407, March 1998. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/content/pdf/10.1006/bulm.1997.0032.pdf.

Siegel:1998:BR

[1515] Ralph M. Siegel, Edward Beltrami, and Ross Cressman. Book reviews. Bulletin of Mathematical Biology, 60(2):409–415, March 1998. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https:

//link.springer.com/content/pdf/10.1006/bulm.1997.0019.pdf.

Bracic:1998:NDB

[1516] Maja Bračič and Aneta Stefanovska. Nonlinear dynamics of the blood flow studied by Lyapunov exponents. Bulletin of Mathematical Biology, 60(3):417–433, May 1998. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/content/
pdf/10.1006/bulm.1997.0007.pdf.

Hosono:1998:MST

[1517] Yuzo Hosono. The minimal speed of traveling fronts for a diffusive Lotka–Volterra competition model. Bulletin of Mathematical Biology, 60(3):435–448, May 1998. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/content/
pdf/10.1006/bulm.1997.0008.pdf.

Edelstein-Keshet:1998:MLD

[1518] Leah Edelstein-Keshet and G. Bard Ermentrout. Models for the length distributions of actin filaments: I. Simple polymerization and fragmentation. Bulletin of Mathematical Biology, 60(3):449–475, May 1998. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/content/pdf/10.1006/
bulm.1997.0011.pdf.

Ermentrout:1998:MLD

[1519] G. Bard Ermentrout and Leah Edelstein-Keshet. Models for the length distributions of actin filaments: II. Polymerization and fragmentation by gelsolin acting together. Bulletin of Mathematical Biology, 60(3): 477–503, May 1998. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL https://link.springer.com/content/pdf/
10.1006/bulm.1997.0012.pdf.

Stephani:1998:KTP

[1520] Amadeus Stephani and Reinhart Heinrich. Kinetic and thermodynamic principles determining the structural design of ATP-producing systems. Bulletin of Mathematical Biology, 60(3):505–543, May 1998. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https:

//link.springer.com/content/pdf/10.1006/bulm.1997.0028.pdf.

Bugnon:1998:VSH

[1521] Denis Bugnon, Gilles Potel, Jocelyne Caillon, Denis Baron, Henri B. Drugeon, Philippe Feigel, and Marie F. Kergueris. In vivo simulation of human pharmacokinetics in the rabbit. Bulletin of Mathematical Biology, 60(3):545–567, May 1998. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/content/
pdf/10.1006/bulm.1997.0013.pdf.

Krishna:1998:CET

[1522] S. V. Krishna, P. D. N. Srinivasu, and B. Kaymakcalan. Conservation of an ecosystem through optimal taxation. Bulletin of Mathematical Biology, 60(3):569–584, May 1998. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/content/pdf/10.1006/bulm.1997.0023.pdf.

DeKoeijer:1998:MSP

[1523] Aline De Koeijer, Odo Diekmann, and Peter Reijnders. Modelling the spread of phocine distemper virus among harbour seals. Bulletin of Mathematical Biology, 60(3):585–596, May 1998. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.

springer.com/content/pdf/10.1006/bulm.1997.0030.pdf.

Ford:1998:BR

[1524] E. David Ford, Gary Benson, Elizabeth Thompson, F. James Rohlf, Brian D. Sleemank Eric Renshaw, and Brian A. Maurer. Book reviews. Bulletin of Mathematical Biology, 60(3):597–613, May 1998. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https:

//link.springer.com/content/pdf/10.1006/bulm.1997.0004.pdf.

Harrigan:1998:LMC

[1525] Timothy P. Harrigan. Limiting models for calcification in fibrous tissues adjacent to orthopedic implants: Variational indicator functions and influences of implant stiffness. Bulletin of Mathematical Biology, 60(4): 615–646, July 1998. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1006/bulm.1997.0003.pdf.

Smith:1998:ULE

[1526] Derek J. Smith, Stephanie Forrest, David H. Ackley, and Alan S. Perelson. Using lazy evaluation to simulate realistic-size repertoires in models of the immune system. Bulletin of Mathematical Biology, 60(4):647– 658, July 1998. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1006/bulm.1997.0035.pdf.

Furusawa:1998:ERC

[1527] Chikara Furusawa and Kunihiko Kaneko. Emergence of rules in cell society: Differentiation, hierarchy, and stability. Bulletin of Mathematical Biology, 60(4):659–687, July 1998. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1006/bulm.1997.0034.pdf.

delaFuente:1998:PBP

[1528] I. Martinez de la Fuente, L. Martinez, N. Benitez, J. Veguillas, and J. M. Aguirregabiria. Persistent behavior in a phase-shift sequence of periodical biochemical oscillations. Bulletin of Mathematical Biology, 60(4): 689–702, July 1998. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1006/bulm.1997.0036.pdf.

Gragnani:1998:FCC

[1529] Alessandra Gragnani, Oscar De Feo, and Sergio Rinaldi. Food chains in the chemostat: Relationships between mean yield and complex dynamics. Bulletin of Mathematical Biology, 60(4):703–719, July 1998. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/content/pdf/10.1006/bulm.1997.0039.pdf.

Krivan:1998:NSA

[1530] Vlastimil Křivan and Giovanni Colombo. A non-stochastic approach for modeling uncertainty in population dynamics. Bulletin of Mathematical Biology, 60(4):721–751, July 1998. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1006/bulm.1998.0040.pdf.

Bonabeau:1998:FRT

[1531] Eric Bonabeau, Guy Theraulaz, and Jean-Louis Deneubourg. Fixed response thresholds and the regulation of division of labor in insect societies. Bulletin of Mathematical Biology, 60(4):753–807, July 1998. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/content/pdf/10.1006/bulm.1998.
0041.pdf.

Drake:1998:BR

[1532] James A. Drake and Ernesto Pérez-Chavela. Book reviews. Bulletin of Mathematical Biology, 60(4):809–814, July 1998. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/content/pdf/10.1006/bulm.1998.0048.pdf.

Mittenthal:1998:DMA

[1533] Jay E. Mittenthal, Ao Yuan, Bertrand Clarke, and Alexander Scheeline. Designing metabolism: Alternative connectivities for the pentose phosphate pathway. Bulletin of Mathematical Biology, 60(5):815–856, September 1998. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1006/bulm.1997.0043.pdf.

Anderson:1998:CDM

[1534] A. R. A. Anderson and M. A. J. Chaplain. Continuous and discrete mathematical models of tumor-induced angiogenesis. Bulletin of Mathematical Biology, 60(5):857–899, September 1998. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/content/pdf/10.1006/bulm.1998.0042.pdf.

Monk:1998:RRG

[1535] Nicholas A. M. Monk. Restricted-range gradients and travelling fronts in a model of juxtacrine cell relay. Bulletin of Mathematical Biology, 60(5):901–918, September 1998. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1006/bulm.1998.0046.pdf.

Bracic:1998:WBA

[1536] Maja Bračič and Aneta Stefanovska. Wavelet-based analysis of human blood-flow dynamics. Bulletin of Mathematical Biology, 60(5):919–935, September 1998. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1006/bulm.1998.0047.pdf.

Pitchford:1998:IPS

[1537] Jonathan Pitchford and John Brindley. Intratrophic predation in simple predator–prey models. Bulletin of Mathematical Biology, 60(5):937–953, September 1998. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1006/bulm.1998.0053.pdf.

Gosselin:1998:RTA

[1538] Frédéric Gosselin. Reconciling theoretical approaches to stochastic patch-occupancy metapopulation models. Bulletin of Mathematical Biology, 60(5):955–971, September 1998. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1006/bulm.1998.0052.pdf.

Bentil:1998:DAE

[1539] Daniel E. Bentil. Distribution of attachment events relative to actin binding sites as evidenced in a bidirectional actomyosin interaction model. Bulletin of Mathematical Biology, 60(5):973–995, September 1998. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/content/pdf/10.1006/bulm.1998.
0055.pdf.

Hernandez:1998:BR

[1540] Gerardo Hernandez. Book review. Bulletin of Mathematical Biology, 60(5):997–998, September 1998. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1006/bulm.1998.0065.pdf.

Atay:1998:MRA

[1541] Fatihcan M. Atay and John Mallet-Paret. Modeling reflex asymmetries with implicit delay differential equations. Bulletin of Mathematical Biology, 60(6):999–1015, November 1998. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/content/pdf/10.1006/S0092-8240%2898%2990000-3.pdf.

Civelekoglu:1998:MAF

[1542] G. Civelekoğlu, Y. Tardy, and J.-J. Meister. Modeling actin filament reorganization in endothelial cells subjected to cyclic stretch. Bulletin of Mathematical Biology, 60(6):1017–1037, November 1998.

CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/content/pdf/10.1006/
S0092-8240%2898%2990001-5.pdf.

Lathrop:1998:BOS

[1543] Richard H. Lathrop, Robert G. Rogers, Jr., Temple F. Smith, and James V. White. A Bayes-optimal sequence-structure theory that unifies protein sequence-structure recognition and alignment. Bulletin of Mathematical Biology, 60(6):1039–1071, November 1998. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/content/pdf/10.1006/
S0092-8240%2898%2990002-7.pdf.

Wills:1998:SDA

[1544] Peter R. Wills, Stuart A. Kauffman, Barbel M. R. Stadler, and Peter F. Stadler. Selection dynamics in autocatalytic systems: Templates replicating through binary ligation. Bulletin of Mathematical Biology, 60(6): 1073–1098, November 1998. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/content/
pdf/10.1006/S0092-8240%2898%2990003-9.pdf.

Troynikov:1998:PDF

[1545] Vladimir S. Troynikov. Probability density functions useful for parametrization of heterogeneity in growth and allometry data. Bulletin of Mathematical Biology, 60(6):1099–1122, November 1998. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https:

//link.springer.com/content/pdf/10.1006/bulm.1998.0058.pdf.

Shulgin:1998:PVS

[1546] Boris Shulgin, Lewi Stone, and Zvia Agur. Pulse vaccination strategy in the SIR epidemic model. Bulletin of Mathematical Biology, 60(6):1123– 1148, November 1998. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/content/
pdf/10.1006/S0092-8240%2898%2990005-2.pdf.

Poggiale:1998:EDC

[1547] J. Ch. Poggiale, J. Michalski, and R. Arditi. Emergence of donor control in patchy predator — prey systems. Bulletin of Mathematical Biology, 60(6):1149–1166, November 1998. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/content/pdf/10.1006/S0092-8240%2898%2990006-4.pdf.

DeVries:1998:DCB

[1548] Gerda De Vries, Arthur Sherman, and Hsiu-Rong Zhu. Diffusively coupled bursters: Effects of cell heterogeneity. Bulletin of Mathematical Biology, 60(6):1167–1200, November 1998. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/content/pdf/10.1006/bulm.1998.0057.pdf.

Eguiarte:1998:BR

[1549] Luis E. Eguiarte. Book review. Bulletin of Mathematical Biology, 60(6): 1201–1205, November 1998. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/content/
pdf/10.1006/S0092-8240%2898%2990008-8.pdf.

Bennett:1999:CAM

[1550] M. R. Bennett, L. Farnell, and W. G. Gibson. Cable analysis of a motornerve terminal branch in a volume conductor. Bulletin of Mathematical Biology, 61(1):1–17, January 1999. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1006/bulm.1998.0066.pdf.

Jost:1999:ADE

[1551] Christian Jost, Ovide Arino, and Roger Arditi. About deterministic extinction in ratio-dependent predator–prey models. Bulletin of Mathematical Biology, 61(1):19–32, January 1999. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/content/pdf/10.1006/bulm.1998.0072.pdf.

Malaka:1999:MCC

[1552] Rainer Malaka. Models of classical conditioning. Bulletin of Mathematical Biology, 61(1):33–83, January 1999. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/content/pdf/10.1006/bulm.1998.9998.pdf.

Gyllenberg:1999:BPI

[1553] Mats Gyllenberg, Timo Koski, Tatu Lund, and Helge G. Gyllenberg. Bayesian predictive identification and cumulative classification of bacteria. Bulletin of Mathematical Biology, 61(1):85–111, January 1999. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/content/pdf/10.1006/bulm.1998.
0076.pdf.

Myers:1999:MMS

[1554] Matthew R. Myers, C. David Lytle, and Licia B. Routson. A mathematical model for simulating virus transport through synthetic barriers. Bulletin of Mathematical Biology, 61(1):113–140, January 1999. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/content/pdf/10.1006/bulm.1998.0080.pdf.

Cavalieri:1999:CDT

[1555] Liebe F. Cavalieri and Huseyin Kocak. Comparative dynamics of three models for host-parasitoid interactions in a patchy environment. Bulletin of Mathematical Biology, 61(1):141–155, January 1999. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/content/pdf/10.1006/bulm.1999.0083.pdf.

Edwards:1999:PTS

[1556] Roderick Edwards, Anne Beuter, and Leon Glass. Parkinsonian tremor and simplification in network dynamics. Bulletin of Mathematical Biology, 61(1):157–177, January 1999. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1006/bulm.1998.0086.pdf.

Sikder:1999:EPT

[1557] Asim Sikder. Effects of parasitism over two competing host populations leading to persistence. Bulletin of Mathematical Biology, 61(1):179– 205, January 1999. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1006/bulm.1998.0087.pdf.

Adler:1999:BR

[1558] Fred Adler. Book review. Bulletin of Mathematical Biology, 61(1):207– 208, January 1999. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1006/bulm.1998.0071.pdf.

Kuchel:1999:PER

[1559] Philip W. Kuchel and Edward D. Fackerell. Parametric-equation representation of biconcave erythrocytes. Bulletin of Mathematical Biology, 61(2):209–220, March 1999. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1006/bulm.1998.0064.pdf.

Eytan:1999:AIU

[1560] Osnat Eytan and David Elad. Analysis of intra-uterine fluid motion induced by uterine contractions. Bulletin of Mathematical Biology, 61(2): 221–238, March 1999. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1006/bulm.1998.0069.pdf.

Franke:1999:PBC

[1561] Ulrich Franke, Kolumban Hutter, and Klaus Jöhnk. A physical-biological coupled model for algal dynamics in lakes. Bulletin of Mathematical Biology, 61(2):239–272, March 1999. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1006/bulm.1998.0075.pdf.

Chopra:1999:VSS

[1562] G. C. Chopra, B. D. Sleeman, J. Brindley, D. G. Knapp, and A. V. Holden. Velocity and stability of solitary planar travelling wave solutions of intracellular [ca]^{2+}. Bulletin of Mathematical Biology, 61(2): 273–301, March 1999. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1006/bulm.1998.0081.pdf.

Edwards:1999:ZMD

[1563] Andrew M. Edwards and John Brindley. Zooplankton mortality and the dynamical behaviour of plankton population models. Bulletin of Mathematical Biology, 61(2):303–339, March 1999. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/content/pdf/10.1006/bulm.1998.0082.pdf.

Hernandez-Suarez:1999:TPM

[1564] Carlos M. Hernández-Suárez, Pablo A. Marquet, and Jorge X. Velasco-Hernández. Threshold parameters and metapopulation persistence. Bulletin of Mathematical Biology, 61(2):341–353, March 1999. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/content/pdf/10.1006/bulm.1998.0084.pdf.

Troynikov:1999:UBT

[1565] Vladimir S. Troynikov. Use of Bayes theorem to correct size-specific sampling bias in growth data. Bulletin of Mathematical Biology, 61(2): 355–363, March 1999. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1006/bulm.1998.0088.pdf.

Wimmer:1999:NSS

[1566] Gejza Wimmer, Ladislav Dedík, Marek Michal, Andrea Mudríková, and Mária Ďurišová. Numerical simulations of stochastic circulatory models. Bulletin of Mathematical Biology, 61(2):365–377, March 1999. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/content/pdf/10.1006/bulm.1998.0094.pdf.

El-Masri:1999:RPC

[1567] Hisham A. El-Masri and Christopher J. Portier. Replication potential of cells via the protein kinase C-MAPK pathway: Application of a mathematical model. Bulletin of Mathematical Biology, 61(2):379– 398, March 1999. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1006/bulm.1998.0077.pdf.

Keitt:1999:BR

[1568] Timothy H. Keitt and Bruce J. West. Book reviews. Bulletin of Mathematical Biology, 61(2):399–401, March 1999. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/content/pdf/10.1006/bulm.1998.0079.pdf.

Fishman:1999:TTD

[1569] Michael A. Fishman and Alan S. Perelson. Th1/Th2 differentiation and cross-regulation. Bulletin of Mathematical Biology, 61(3):403–436, May 1999. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/content/pdf/10.1006/
bulm.1998.0074.pdf.

Haslinger:1999:RSP

[1570] Christian Haslinger and Peter F. Stadler. RNA structures with pseudoknots: Graph-theoretical, combinatorial, and statistical properties. Bulletin of Mathematical Biology, 61(3):437–467, May 1999. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://

link.springer.com/content/pdf/10.1006/bulm.1998.0085.pdf.

Britton:1999:MST

[1571] N. F. Britton, L. W. Partridge, and N. R. Franks. A model of survival times for predator populations: The case of the army ants. Bulletin of Mathematical Biology, 61(3):469–482, May 1999. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.

springer.com/content/pdf/10.1006/bulm.1998.0091.pdf.

Barrio:1999:TDN

[1572] R. A. Barrio, C. Varea, J. L. Aragón, and P. K. Maini. A twodimensional numerical study of spatial pattern formation in interacting Turing systems. Bulletin of Mathematical Biology, 61(3):483–505, May 1999. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/content/pdf/10.1006/
bulm.1998.0093.pdf.

Houart:1999:BCB

[1573] Gérald Houart, Geneviève Dupont, and Albert Goldbeter. Bursting, chaos and birhythmicity originating from self-modulation of the inositol 1,4,5-trisphosphate signal in a model for intracellular ca^{2+} oscillations. Bulletin of Mathematical Biology, 61(3):507–530, May 1999. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https:

//link.springer.com/content/pdf/10.1006/bulm.1999.0095.pdf.

Parvinen:1999:EMM

[1574] Kalle Parvinen. Evolution of migration in a metapopulation. Bulletin of Mathematical Biology, 61(3):531–550, May 1999. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.

springer.com/content/pdf/10.1006/bulm.1999.0100.pdf.

Orssengo:1999:DTI

[1575] Girma J. Orssengo and David C. Pye. Determination of the true intraocular pressure and modulus of elasticity of the human cornea in vivo. Bulletin of Mathematical Biology, 61(3):551–572, May 1999. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https:

//link.springer.com/content/pdf/10.1006/bulm.1999.0102.pdf.

DelaFuente:1999:PMN

[1576] I. M. De la Fuente, N. Benitez, A. Santamaria, J. M. Aguirregabiria, and J. Veguillas. Persistence in metabolic nets. Bulletin of Mathematical Biology, 61(3):573–595, May 1999. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/content/pdf/10.1006/bulm.1999.0103.pdf.

Bolker:1999:BR

[1577] Benjamin Bolker and Octavio Miramontes. Book reviews. Bulletin of Mathematical Biology, 61(3):597–600, May 1999. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.

springer.com/content/pdf/10.1006/bulm.1998.0092.pdf.

Thompson:1999:MIL

[1578] K. E. Thompson and H. M. Byrne. Modelling the internalization of labelled cells in tumour spheroids. Bulletin of Mathematical Biology, 61(4): 601–623, July 1999. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1006/bulm.1999.0089.pdf.

Caswell:1999:CAM

[1579] Hal Caswell and Ron Etter. Cellular automaton models for competition in patchy environments: Facilitation, inhibition, and tolerance. Bulletin of Mathematical Biology, 61(4):625–649, July 1999. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/content/pdf/10.1006/bulm.1999.0090.pdf.

Spiro:1999:EHD

[1580] Peter A. Spiro and Hans G. Othmer. The effect of heterogeneouslydistributed RyR channels on calcium dynamics in cardiac myocytes. Bulletin of Mathematical Biology, 61(4):651–681, July 1999. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/content/pdf/10.1006/bulm.1999.0101.pdf.

Lawry:1999:HMF

[1581] James Lawry. A hydrostatic Michell framework supports frog lungs. Bulletin of Mathematical Biology, 61(4):683–700, July 1999. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/content/pdf/10.1006/bulm.1999.0104.pdf.

Yusim:1999:TFN

[1582] K. Yusim, H. Parnas, and L. Segel. Theory of fast neurotransmitter release control based on voltage-dependent interaction between autoreceptors and proteins of the exocytotic machinery. Bulletin of Mathematical Biology, 61(4):701–725, July 1999. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/content/pdf/10.1006/bulm.1999.0107.pdf.

Bonabeau:1999:DOA

[1583] Eric Bonabeau, Guy Theraulaz, and Jean-Louis Deneubourg. Dominance orders in animal societies: The self-organization hypothesis revisited. Bulletin of Mathematical Biology, 61(4):727–757, July 1999. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/content/pdf/10.1006/bulm.1999.0108.pdf.

Thompson:1999:HRP

[1584] Kevin E. Thompson and Janice A. Royds. Hypoxia and reoxygenation: a pressure for mutant p53 cell selection and tumour progression. Bulletin of Mathematical Biology, 61(4):759–778, July 1999. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/content/pdf/10.1006/bulm.1999.0109.pdf.

Chen:1999:NSN

[1585] Benito M. Chen and Hristo V. Kojouharov. Non-standard numerical methods applied to subsurface biobarrier formation models in porous media. Bulletin of Mathematical Biology, 61(4):779–798, July 1999. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/content/pdf/10.1006/bulm.1999.0113.pdf.

Thieme:1999:BR

[1586] Horst R. Thieme, Luis Aguirre, and Lourdes Esteva-Peralta. Book reviews. Bulletin of Mathematical Biology, 61(4):799–805, July 1999. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/content/pdf/10.1006/bulm.1999.
0099.pdf.

Sekimura:1999:PFS

[1587] Toshio Sekimura, Mei Zhu, Julian Cook, Philip K. Maini, and James D. Murray. Pattern formation of scale cells in lepidoptera by differential origin-dependent cell adhesion. Bulletin of Mathematical Biology, 61(5): 807–828, September 1999. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/content/
pdf/10.1006/bulm.1998.0062.pdf.

Focardi:1999:OSC

[1588] Stefano Focardi and Maurizio Rizzotto. Optimal strategies and complexity: a theoretical analysis of the anti-predatory behavior of the hare. Bulletin of Mathematical Biology, 61(5):829–848, September 1999. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https:

//link.springer.com/content/pdf/10.1006/bulm.1999.0114.pdf.

Bolker:1999:AMP

[1589] Benjamin M. Bolker. Analytic models for the patchy spread of plant disease. Bulletin of Mathematical Biology, 61(5):849–874, September 1999. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/content/pdf/10.1006/
bulm.1999.0115.pdf. See comment [1882].

Levine:1999:PNP

[1590] David N. Levine. The pathogenesis of normal pressure hydrocephalus: a theoretical analysis. Bulletin of Mathematical Biology, 61(5):875–916, September 1999. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL https://link.springer.com/content/pdf/
10.1006/bulm.1999.0116.pdf.

Obeyesekere:1999:MCC

[1591] Mandri N. Obeyesekere, Stuart O. Zimmerman, Edwin S. Tecarro, and Giles Auchmuty. A model of cell cycle behavior dominated by kinetics of a pathway stimulated by growth factors. Bulletin of Mathematical Biology, 61(5):917–934, September 1999. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/content/pdf/10.1006/bulm.1999.0118.pdf.

Stilianakis:1999:MCD

[1592] Nikolaos I. Stilianakis and Johannes Seydel. Modeling the T-cell dynamics and pathogenesis of HTLV-I infection. Bulletin of Mathematical Biology, 61(5):935–947, September 1999. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/content/pdf/10.1006/bulm.1999.0117.pdf.

Yokozawa:1999:SHS

[1593] Masayuki Yokozawa. Size hierarchy and stability in competitive plant populations. Bulletin of Mathematical Biology, 61(5):949–961, September 1999. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/content/pdf/10.1006/
bulm.1999.0120.pdf.

Buchman:1999:TEM

[1594] E. Buchman and H. Parnas. A theoretical experimental method to determine the locus of desensitization. Bulletin of Mathematical Biology, 61(5):963–986, September 1999. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/content/pdf/10.1006/bulm.1999.0121.pdf.

Rajan:1999:CEP

[1595] Kumar Rajan and Narsingh Deo. Computational experience with a parallel algorithm for tetrangle inequality bound smoothing. Bulletin of Mathematical Biology, 61(5):987–1008, September 1999. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://

link.springer.com/content/pdf/10.1006/bulm.1999.0123.pdf.

Miramontes:1999:BR

[1596] Pedro Miramontes and Jose Crossa. Book reviews. Bulletin of Mathematical Biology, 61(5):1009–1013, September 1999. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.

springer.com/content/pdf/10.1006/bulm.1999.0110.pdf.

Troynikov:1999:PDF

[1597] Vladimir S. Troynikov. Probability density functions useful for parametrization of heterogeneity in growth and allometry data. Bulletin of Mathematical Biology, 61(5):1015–1016, September 1999. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https:

//link.springer.com/content/pdf/10.1006/bulm.1999.0124.pdf.

DeVries:1999:DCB

[1598] Gerda De Vries, Arthur Sherman, and Hsiu-Rong Zhu. Diffusively coupled bursters: Effects of cell heterogeneity. Bulletin of Mathematical Biology, 61(5):1017, September 1999. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/content/pdf/10.1006/bulm.1999.0125.pdf.

Hagerstrand:1999:MSD

[1599] Henry Hägerstrand, Veronika Kralj-Iglič, Małgorzata Bobrowska-Hägerstrand, and Aleš Iglič. Membrane skeleton detachment in spherical and cylindrical microexovesicles. Bulletin of Mathematical Biology, 61(6):1019–1030, November 1999. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1006/bulm.1999.0128.pdf.

Schuette:1999:MEV

[1600] Matthew C. Schuette and Herbert W. Hethcote. Modeling the effects of varicella vaccination programs on the incidence of chickenpox and shingles. Bulletin of Mathematical Biology, 61(6):1031–1064, November 1999. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/content/pdf/10.1006/bulm.1999.
0126.pdf.

Weinstein:1999:MEC

[1601] Alan M. Weinstein. Modeling epithelial cell homeostasis: Steadystate analysis. Bulletin of Mathematical Biology, 61(6):1065–1091, November 1999. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1006/bulm.1999.0127.pdf.

Crampin:1999:RDG

[1602] Edmund J. Crampin, Eamonn A. Gaffney, and Philip K. Maini. Reaction and diffusion on growing domains: Scenarios for robust pattern formation. Bulletin of Mathematical Biology, 61(6):1093–1120, November 1999. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/content/pdf/10.1006/bulm.1999.
0131.pdf.

Henson:1999:MAS

[1603] Shandelle M. Henson, R. F. Costantino, J. M. Cushing, Brian Dennis, and Robert A. Desharnais. Multiple attractors, saddles, and population dynamics in periodic habitats. Bulletin of Mathematical Biology, 61(6):1121–1149, November 1999. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1006/bulm.1999.0136.pdf.

Mesterton-Gibbons:1999:EPW

[1604] Michael Mesterton-Gibbons. On the evolution of pure winner and loser effects: a game-theoretic model. Bulletin of Mathematical Biology, 61(6):1151–1186, November 1999. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1006/bulm.1999.0137.pdf.

Sole:1999:CCE

[1605] Ricard V. Solé, Javier G. P. Gamarra, Marta Ginovart, and Daniel López. Controlling chaos in ecology: From deterministic to individualbased models. Bulletin of Mathematical Biology, 61(6):1187–1207, November 1999. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1006/bulm.1999.0141.pdf.

Nisbet:1999:BR

[1606] Roger M. Nisbet. Book review. Bulletin of Mathematical Biology, 61(6):1209–1210, November 1999. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1006/bulm.1999.0135.pdf.

Katz:2000:FAA

[1607] Yehuda Katz and Itzhak Aharon. Four applications of the anesthetic cut-off effect. Bulletin of Mathematical Biology, 62(1):1–16, January 2000. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/content/pdf/10.1006/
bulm.1999.0098.pdf.

Peleg:2000:EPA

[1608] Micha Peleg and Joseph Horowitz. On estimating the probability of aperiodic outbursts of microbial populations from their fluctuating counts. Bulletin of Mathematical Biology, 62(1):17–35, January 2000. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https:

//link.springer.com/content/pdf/10.1006/bulm.1999.0112.pdf.

Ciliberto:2000:MME

[1609] Andrea Ciliberto and John J. Tyson. Mathematical model for early development of the sea urchin embryo. Bulletin of Mathematical Biology, 62(1):37–59, January 2000. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/content/
pdf/10.1006/bulm.1999.0129.pdf.

Gurney:2000:SOS

[1610] William S. C. Gurney and A. Roy Veitch. Self-organization, scale and stability in a spatial predator-prey interaction. Bulletin of Mathematical Biology, 62(1):61–86, January 2000. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/content/pdf/10.1006/bulm.1999.0130.pdf.

Mosquera:2000:UPB

[1611] J. Mosquera, M. Gómez-Gesteira, and V. Pérez-Villar. Using parasites as biological tags of fish populations: a dynamical model. Bulletin of Mathematical Biology, 62(1):87–99, January 2000. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.

springer.com/content/pdf/10.1006/bulm.1999.0142.pdf.

Pernarowski:2000:FSB

[1612] M. Pernarowski. Fast subsystem bifurcations in strongly coupled heterogeneous collections of excitable cells. Bulletin of Mathematical Biology, 62 (1):101–120, January 2000. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/content/
pdf/10.1006/bulm.1999.0143.pdf.

Oprea:2000:DOP

[1613] Mihaela Oprea, Erik Van Nimwegen, and Alan S. Perelson. Dynamics of one-pass germinal center models: Implications for affinity maturation. Bulletin of Mathematical Biology, 62(1):121–153, January 2000. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https:

//link.springer.com/content/pdf/10.1006/bulm.1999.0144.pdf.

Jang:2000:ESS

[1614] S. R.-J. Jang and Arun K. Mitra. Equilibrium stability of singlespecies metapopulations. Bulletin of Mathematical Biology, 62(1):155– 161, January 2000. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL https://link.springer.com/content/pdf/
10.1006/bulm.1999.0145.pdf.

Kroll:2000:PHT

[1615] Martin H. Kroll. Parathyroid hormone temporal effects on bone formation and resorption. Bulletin of Mathematical Biology, 62(1):163– 188, January 2000. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL https://link.springer.com/content/pdf/
10.1006/bulm.1999.0146.pdf.

Brauer:2000:BR

[1616] Fred Brauer, Graeme Ruxton, Stephen J. Merrill, Jonathan Heller, and Daniel E. Bentil. Book reviews. Bulletin of Mathematical Biology, 62 (1):189–197, January 2000. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/content/
pdf/10.1006/bulm.1999.0111.pdf.

Block:2000:PEQ

[1617] Garry L. Block and Linda J. S. Allen. Population extinction and quasistationary behavior in stochastic density-dependent structured models. Bulletin of Mathematical Biology, 62(2):199–228, March 2000. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https:

//link.springer.com/content/pdf/10.1006/bulm.1999.0147.pdf.

Sherman:2000:CCD

[1618] Claire D. Sherman and Christopher J. Portier. Calculation of the cumulative distribution function of the time to a small observable tumor. Bulletin of Mathematical Biology, 62(2):229–240, March 2000. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https:

//link.springer.com/content/pdf/10.1006/bulm.1999.0148.pdf.

Horovitz:2000:RBH

[1619] Amnon Horovitz and Ofer Yifrach. On the relationship between the Hill coefficients for steady-state and transient kinetic data: a criterion for concerted transitions in allosteric proteins. Bulletin of Mathematical Biology, 62(2):241–246, March 2000. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/content/
pdf/10.1006/bulm.1999.0150.pdf.

Smolen:2000:MTC

[1620] Paul Smolen, Douglas A. Baxter, and John H. Byrne. Modeling transcriptional control in gene networks — methods, recent results, and future directions. Bulletin of Mathematical Biology, 62(2):247–292, March 2000. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/content/pdf/10.1006/
bulm.1999.0155.pdf.

Wearing:2000:MMJ

[1621] H. J. Wearing, M. R. Owen, and J. A. Sherratt. Mathematical modelling of juxtacrine patterning. Bulletin of Mathematical Biology, 62 (2):293–320, March 2000. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/content/
pdf/10.1006/bulm.1999.0152.pdf.

Schnell:2000:TDC

[1622] S. Schnell and C. Mendoza. Time-dependent closed form solutions for fully competitive enzyme reactions. Bulletin of Mathematical Biology, 62 (2):321–336, March 2000. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/content/
pdf/10.1006/bulm.1999.0156.pdf.

Silva:2000:SMM

[1623] Jacques A. L. Silva, Manuela L. De Castro, and Dagoberto A. R. Justo. Synchronism in a metapopulation model. Bulletin of Mathematical Biology, 62(2):337–349, March 2000. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/content/pdf/10.1006/bulm.1999.0157.pdf.

Silverman:2000:RES

[1624] Emily Silverman and Mark Kot. Rate estimation for a simple movement model. Bulletin of Mathematical Biology, 62(2):351–375, March 2000. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/content/pdf/10.1006/
bulm.1999.0159.pdf.

Truscott:2000:QAM

[1625] J. E. Truscott, C. A. Gilligan, and C. R. Webb. Quantitative analysis and model simplification of an epidemic model with primary and secondary infection. Bulletin of Mathematical Biology, 62(2):377–393, March 2000. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL https://link.springer.com/content/pdf/
10.1006/bulm.1999.0158.pdf.

Holte:2000:BR

[1626] Sarah Holte and Christpher M. Kribs-Zaleta. Book reviews. Bulletin of Mathematical Biology, 62(2):395–398, March 2000. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.

springer.com/content/pdf/10.1006/bulm.1999.0105.pdf.

Stadler:2000:PDF

[1627] Peter F. Stadler, Rudi Seitz, and Günter P. Wagner. Population dependent Fourier decomposition of fitness landscapes over recombination spaces: Evolvability of complex characters. Bulletin of Mathematical Biology, 62(3):399–428, May 2000. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/content/pdf/10.1006/bulm.1999.0167.pdf.

Ghorai:2000:WGP

[1628] S. Ghorai and N. A. Hill. Wavelengths of gyrotactic plumes in bioconvection. Bulletin of Mathematical Biology, 62(3):429–450, May 2000. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/content/pdf/10.1006/
bulm.1999.0160.pdf.

Broom:2000:EKC

[1629] M. Broom, C. Cannings, and G. T. Vickers. Evolution in knockout conflicts: The fixed strategy case. Bulletin of Mathematical Biology, 62 (3):451–466, May 2000. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/content/
pdf/10.1006/bulm.1999.0161.pdf.

Feng:2000:IFM

[1630] Jianfeng Feng and David Brown. Integrate-and-fire models with nonlinear leakage. Bulletin of Mathematical Biology, 62(3):467–481, May 2000. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/content/pdf/10.1006/
bulm.1999.0162.pdf.

Schnell:2000:EKH

[1631] S. Schnell and P. K. Maini. Enzyme kinetics at high enzyme concentration. Bulletin of Mathematical Biology, 62(3):483–499, May 2000. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/content/pdf/10.1006/
bulm.1999.0163.pdf.

Painter:2000:CMA

[1632] K. J. Painter, P. K. Maini, and H. G. Othmer. A chemotactic model for the advance and retreat of the primitive streak in avian development. Bulletin of Mathematical Biology, 62(3):501–525, May 2000. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https:

//link.springer.com/content/pdf/10.1006/bulm.1999.0166.pdf.

Afenya:2000:DIG

[1633] Evans K. Afenya and Calixto P. Calderón. Diverse ideas on the growth kinetics of disseminated cancer cells. Bulletin of Mathematical Biology, 62(3):527–542, May 2000. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/content/
pdf/10.1006/bulm.1999.0165.pdf.

DeMuynck:2000:MAI

[1634] P. De Muynck, J. Cornelis, and L. I. Titomir. Multipole approach to the inverse problem in electrocardiology: Convergence of the multipole equivalent generator on the inhomogeneous body conductor. Bulletin of Mathematical Biology, 62(3):543–583, May 2000. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.

springer.com/content/pdf/10.1006/bulm.2000.0169.pdf.

Greenhalgh:2000:BR

[1635] David Greenhalgh. Book review. Bulletin of Mathematical Biology, 62 (3):585–588, May 2000. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/content/
pdf/10.1006/bulm.1999.0153.pdf.

Fernandez:2000:BR

[1636] Joaquín Delgado Fernandez. Book review. Bulletin of Mathematical Biology, 62(3):588–590, May 2000. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/content/pdf/10.1006/bulm.1999.0154.pdf.

Strogatz:2000:BR

[1637] Steven H. Strogatz. Book review. Bulletin of Mathematical Biology, 62 (3):591–593, May 2000. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/content/
pdf/10.1006/bulm.1999.0164.pdf.

Gandhi:2000:MES

[1638] Amar Gandhi, Simon Levin, and Steven Orszag. Moment expansions in spatial ecological models and moment closure through Gaussian approximation. Bulletin of Mathematical Biology, 62(4):595–632, July 2000. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/content/pdf/10.1006/
bulm.1999.0119.pdf.

Fowler:2000:EIT

[1639] A. C. Fowler. The effect of incubation time distribution on the extinction characteristics of a rabies epizootic. Bulletin of Mathematical Biology, 62(4):633–656, July 2000. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/content/
pdf/10.1006/bulm.1999.0170.pdf.

Devaud:2000:SAP

[1640] J. M. Devaud, B. Quenet, J. Gascuel, and C. Masson. Statistical analysis and parsimonious modelling of dendrograms of in vitro neurones. Bulletin of Mathematical Biology, 62(4):657–674, July 2000. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.

springer.com/content/pdf/10.1006/bulm.1999.0171.pdf.

Davis:2000:DAM

[1641] Brian K. Davis. Darwinian aspects of molecular evolution at sublinear propagation rates. Bulletin of Mathematical Biology, 62(4):675– 694, July 2000. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL https://link.springer.com/content/pdf/
10.1006/bulm.2000.0173.pdf. See erratum [1658].

Shorten:2000:HHM

[1642] Paul R. Shorten and David J. N. Wall. A Hodgkin–Huxley model exhibiting bursting oscillations. Bulletin of Mathematical Biology, 62(4): 695–715, July 2000. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL https://link.springer.com/content/pdf/
10.1006/bulm.2000.0172.pdf.

Yusim:2000:TFI

[1643] K. Yusim, H. Parnas, and L. A. Segel. Theory for the feedback inhibition of fast release of neurotransmitter. Bulletin of Mathematical Biology, 62 (4):717–757, July 2000. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/content/
pdf/10.1006/bulm.2000.0174.pdf.

Pagie:2000:IPB

[1644] Ludo Pagie and Paulien Hogeweg. Individual-and population-based diversity in restriction-modification systems. Bulletin of Mathematical Biology, 62(4):759–774, July 2000. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/content/pdf/10.1006/bulm.2000.0177.pdf.

Bearon:2000:MRT

[1645] R. N. Bearon and T. J. Pedley. Modelling run-and-tumble chemotaxis in a shear flow. Bulletin of Mathematical Biology, 62(4):775– 791, July 2000. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL https://link.springer.com/content/pdf/
10.1006/bulm.2000.0178.pdf.

Caraco:2000:BRS

[1646] Thomas Caraco and Stephan Glavanakov. Book review: Sampling and Statistical Methods for Behavioral Ecologists, by J. Bart, M. A. Fligner and W. I. Notz, Cambridge University Press, New York, 1998. $29.95 (paperback), 330 pp. ISBN 0-521-45705-X. Bulletin of Mathematical Biology, 62(4):793–794, July 2000. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/content/
pdf/10.1006/bulm.1999.0168.pdf.

Munoz-Martinez:2000:BRS

[1647] E. J. Muñoz-Martínez. Book review: Small Worlds: The Dynamics of Networks Between Order and Randomness, by Duncan J. Watts, (Princeton Studies in Complexity), Princeton University Press, 1999. $39.50 (hardcover), 262 pp. ISBN 0-691-00541-9. Bulletin of Mathematical Biology, 62(4):795–797, July 2000. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.com/

content/pdf/10.1006/bulm.1999.0175.pdf.

Wagner:2000:BRI

[1648] Israel A. Wagner. Book review: Information Processing in Social Insects, by C. Detrain, J. L. Deneubourg and Jacques M. Pasteels (Editors), Birkhäuser, 1999. $130.00, 415 pp. ISBN 0-817-65792-4. Bulletin of Mathematical Biology, 62(4):799–801, July 2000. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.

springer.com/content/pdf/10.1006/bulm.1999.0184.pdf.

vanNimwegen:2000:MED

[1649] Erik van Nimwegen and James P. Crutchfield. Metastable evolutionary dynamics: Crossing fitness barriers or escaping via neutral paths? Bulletin of Mathematical Biology, 62(5):799–848, September 2000. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/content/pdf/10.1006/bulm.2000.0180.pdf.

Berec:2000:MEL

[1650] Luděk Berec. Mixed encounters, limited perception and optimal foraging. Bulletin of Mathematical Biology, 62(5):849–868, September 2000. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/content/pdf/10.1006/bulm.2000.
0179.pdf.

Pilyugin:2000:MIR

[1651] Sergei S. Pilyugin and Rustom Antia. Modeling immune responses with handling time. Bulletin of Mathematical Biology, 62(5):869–890, September 2000. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1006/bulm.2000.0181.pdf.

Broday:2000:DCT

[1652] David M. Broday. Diffusion of clusters of transmembrane proteins as a model of focal adhesion remodeling. Bulletin of Mathematical Biology, 62(5):891–924, September 2000. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1006/bulm.2000.0183.pdf.

Ipina:2000:MSD

[1653] S. L. Ipiña and A. I. Durand. A measure of sexual dimorphism in populations which are univariate normal mixtures. Bulletin of Mathematical Biology, 62(5):925–941, September 2000. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1006/bulm.2000.0185.pdf.

Etienne:2000:LPD

[1654] Rampal S. Etienne. Local populations of different sizes, mechanistic rescue effect and patch preference in the Levins metapopulation model. Bulletin of Mathematical Biology, 62(5):943–958, September 2000. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/content/pdf/10.1006/bulm.2000.0186.pdf.

Snyder:2000:SSF

[1655] Robin E. Snyder and Roger M. Nisbet. Spatial structure and fluctuations in the contact process and related models. Bulletin of Mathematical Biology, 62(5):959–975, September 2000. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1006/bulm.2000.0191.pdf.

Powell:2000:STA

[1656] James A. Powell, Janette L. Jenkins, Jesse A. Logan, and Barbara J. Bentz. Seasonal temperature alone can synchronize life cycles. Bulletin of Mathematical Biology, 62(5):977–998, September 2000. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/content/pdf/10.1006/bulm.2000.0192.pdf.

Hernandez:2000:BR

[1657] Gerardo Hernández. Book review. Bulletin of Mathematical Biology, 62(5):999–1000, September 2000. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1006/bulm.2000.0176.pdf.

Davis:2000:EDA

[1658] Brian K. Davis. Erratum: Darwinian aspects of molecular evolution at sublinear propagation rates. Bulletin of Mathematical Biology, 62(5): 1001, September 2000. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/content/
pdf/10.1006/bulm.2000.0204.pdf. See [1641].

Clother:2000:SDI

[1659] D. R. Clother and J. Brindley. Stochastic development of individual members of a population: a Brownian motion approach. Bulletin of Mathematical Biology, 62(6):1003–1034, November 2000. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://

link.springer.com/content/pdf/10.1006/bulm.2000.0189.pdf.

Mayer:2000:NSF

[1660] Stefan Mayer. Numerical simulation of flow fields and particle trajectories in ciliary suspension feeding. Bulletin of Mathematical Biology, 62(6): 1035–1059, November 2000. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/content/
pdf/10.1006/bulm.2000.0190.pdf.

Stadler:2000:DAR

[1661] Bärbel M. R. Stadler, Peter F. Stadler, and Peter Schuster. Dynamics of autocatalytic replicator networks based on higher-order ligation reactions. Bulletin of Mathematical Biology, 62(6):1061–1086, November 2000. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/content/pdf/10.1006/
bulm.2000.0194.pdf.

Hung:2000:UTR

[1662] George K. Hung and Kenneth J. Ciuffreda. A unifying theory of refractive error development. Bulletin of Mathematical Biology, 62(6):1087– 1108, November 2000. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/content/
pdf/10.1006/bulm.2000.0199.pdf.

Witting:2000:PCC

[1663] Lars Witting. Population cycles caused by selection by density dependent competitive interactions. Bulletin of Mathematical Biology, 62(6):1109– 1136, November 2000. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/content/
pdf/10.1006/bulm.2000.0200.pdf.

Loladze:2000:SPG

[1664] Irakli Loladze, Yang Kuang, and James J. Elser. Stoichiometry in producer–grazer systems: Linking energy flow with element cycling. Bulletin of Mathematical Biology, 62(6):1137–1162, November 2000. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/content/pdf/10.1006/
bulm.2000.0201.pdf.

Muller:2000:SPV

[1665] Erik B. Muller and Roger M. Nisbet. Survival and production in variable resource environments. Bulletin of Mathematical Biology, 62(6):1163– 1189, November 2000. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/content/
pdf/10.1006/bulm.2000.0203.pdf.

Baer:2000:BR

[1666] Steven M. Baer, Warren J. Ewens, and C. M. Hernandez-Suarez. Book reviews. Bulletin of Mathematical Biology, 62(6):1191–1194, November 2000. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/content/pdf/10.1006/
bulm.1999.0195.pdf.

Coombes:2001:EIP

[1667] S. Coombes. The effect of ion pumps on the speed of travelling waves in the fire-diffuse-fire model of Ca^{2+} release. Bulletin of Mathematical Biology, 63(1):1–20, January 2001. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1006/bulm.2000.0193.pdf.

Ebenhoh:2001:EOM

[1668] Oliver Ebenhöh and Reinhart Heinrich. Evolutionary optimization of metabolic pathways. Theoretical reconstruction of the stoichiometry of ATP and NADH producing systems. Bulletin of Mathematical Biology, 63(1):21–55, January 2001. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1006/bulm.2000.0197.pdf.

Reidys:2001:RMN

[1669] Christian Reidys, Christian V. Forst, and Peter Schuster. Replication and mutation on neutral networks. Bulletin of Mathematical Biology, 63(1): 57–94, January 2001. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1006/bulm.2000.0206.pdf.

Dockery:2001:MMQ

[1670] Jack D. Dockery and James P. Keener. A mathematical model for quorum sensing in Pseudomonas aeruginosa. Bulletin of Mathematical Biology, 63(1):95–116, January 2001. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1006/bulm.2000.0205.pdf.

Dopazo:2001:MIL

[1671] H. Dopazo, M. B. Gordon, R. Perazzo, and S. Risau-Gusman. A model for the interaction of learning and evolution. Bulletin of Mathematical Biology, 63(1):117–134, January 2001. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1006/bulm.2000.0207.pdf.

Gammack:2001:ESA

[1672] D. Gammack, H. M. Byrne, and C. E. Lewis. Estimating the selective advantage of mutant p53 tumour cells to repeated rounds of hypoxia. Bulletin of Mathematical Biology, 63(1):135–166, January 2001. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/content/pdf/10.1006/bulm.2000.0210.pdf.

Pernarowski:2001:CES

[1673] M. Pernarowski. Controllability of excitable systems. Bulletin of Mathematical Biology, 63(1):167–184, January 2001. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/content/pdf/10.1006/bulm.2000.0212.pdf.

Papageorgiou:2001:PFM

[1674] Spyros Papageorgiou. A physical force may expose Hox genes to express in a morphogenetic density gradient. Bulletin of Mathematical Biology, 63(1):185–200, January 2001. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1006/bulm.2000.0211.pdf.

Moloney:2001:BR

[1675] Kirk A. Moloney, Lourdes Esteva-Peralta, and Craig Henriquez. Book reviews. Bulletin of Mathematical Biology, 63(1):201–205, January 2001. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/content/pdf/10.1006/bulm.2000.
0202.pdf.

Dilao:2001:PQP

[1676] Rui Dilão and Tiago Domingos. Periodic and quasi-periodic behavior in resource-dependent age structured population models. Bulletin of Mathematical Biology, 63(2):207–230, March 2001. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/content/pdf/10.1006/bulm.2000.0213.pdf.

Pettet:2001:MCM

[1677] G. J. Pettet, C. P. Please, M. J. Tindall, and D. L. S. McElwain. The migration of cells in multicell tumor spheroids. Bulletin of Mathematical Biology, 63(2):231–257, March 2001. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1006/bulm.2000.0217.pdf.

Grasman:2001:MCE

[1678] J. Grasman, F. van den Bosch, and O. A. van Herwaarden. Mathematical conservation ecology: a one-predator-two-prey system as case study. Bulletin of Mathematical Biology, 63(2):259–269, March 2001. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/content/pdf/10.1006/bulm.2000.0218.pdf.

Kooi:2001:BTF

[1679] B. W. Kooi and P. P. F. Hanegraaf. Bi-trophic food chain dynamics with multiple component populations. Bulletin of Mathematical Biology, 63(2):271–299, March 2001. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1006/bulm.2000.0219.pdf.

Edwards:2001:ERL

[1680] David A. Edwards. The effect of a receptor layer on the measurement of rate constants. Bulletin of Mathematical Biology, 63(2):301– 327, March 2001. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1006/bulm.2000.0224.pdf.

Drew:2001:MMP

[1681] Donald A. Drew. A mathematical model for prokaryotic protein synthesis. Bulletin of Mathematical Biology, 63(2):329–351, March 2001. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/content/pdf/10.1006/bulm.2000.
0225.pdf.

Fuentes-Pardo:2001:OMC

[1682] Beatriz Fuentes-Pardo, Miguel Lara-Aparicio, and Santiago López de Medrano. On the ontogeny of the motor circadian rhythm in crayfish. Bulletin of Mathematical Biology, 63(2):353–369, March 2001. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/content/pdf/10.1006/bulm.2001.0227.pdf.

deVries:2001:SBC

[1683] Gerda de Vries and Arthur Sherman. From spikers to bursters via coupling: Help from heterogeneity. Bulletin of Mathematical Biology, 63(2): 371–391, March 2001. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1006/bulm.2001.0228.pdf.

Kozusko:2001:BWS

[1684] Frank P. Kozusko. Body weight setpoint, metabolic adaption and human starvation. Bulletin of Mathematical Biology, 63(2):393–403, March 2001. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/content/pdf/10.1006/bulm.2001.
0229.pdf.

Bergmann:2001:TTH

[1685] Claudia Bergmann, J. Leo van Hemmen, and Lee A. Segel. Th1 or Th2: How an appropriate T helper response can be made. Bulletin of Mathematical Biology, 63(3):405–430, May 2001. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.

springer.com/content/pdf/10.1006/bulm.2000.0215.pdf.

Heckman:2001:MMM

[1686] C. A. Heckman, C. S. Runyeon, J. G. Wade, and S. Seubert. Mathematical modeling of marker influx and efflux in cells. Bulletin of Mathematical Biology, 63(3):431–449, May 2001. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/content/pdf/10.1006/bulm.2001.0216.pdf.

Komarova:2001:EDL

[1687] Natalia L. Komarova and Martin A. Nowak. The evolutionary dynamics of the lexical matrix. Bulletin of Mathematical Biology, 63(3):451– 484, May 2001. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL https://link.springer.com/content/pdf/
10.1006/bulm.2000.0222.pdf.

Silva:2001:SMM

[1688] Jacques A. L. Silva, Manuela L. De Castro, and Dagoberto A. R. Justo. Stability in a metapopulation model with density-dependent dispersal. Bulletin of Mathematical Biology, 63(3):485–505, May 2001. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https:

//link.springer.com/content/pdf/10.1006/bulm.2000.0221.pdf.

Garcia-Barrios:2001:DVS

[1689] Luis García-Barrios, David Mayer-Foulkes, Miguel Franco, Guillermo Urquijo-Vásquez, and Juan Franco-Pérez. Development and validation of a spatially explicit individual-based mixed crop growth model. Bulletin of Mathematical Biology, 63(3):507–526, May 2001. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.

springer.com/content/pdf/10.1006/bulm.2000.0226.pdf.

Pitchford:2001:PPP

[1690] Jonathan William Pitchford and John Brindley. Prey patchiness, predator survival and fish recruitment. Bulletin of Mathematical Biology, 63 (3):527–546, May 2001. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/content/
pdf/10.1006/bulm.2001.0230.pdf.

Rodriguez:2001:MID

[1691] Diego J. Rodríguez and Lourdes Torres-Sorando. Models of infectious diseases in spatially heterogeneous environments. Bulletin of Mathematical Biology, 63(3):547–571, May 2001. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/content/pdf/10.1006/bulm.2001.0231.pdf.

Jenkins:2001:LST

[1692] Janette L. Jenkins, James A. Powell, Jesse A. Logan, and Barbara J. Bentz. Low seasonal temperatures promote life cycle synchronization. Bulletin of Mathematical Biology, 63(3):573–595, May 2001. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https:

//link.springer.com/content/pdf/10.1006/bulm.2001.0237.pdf.

Gerrish:2001:BR

[1693] Philip Gerrish and Hong Qian. Book reviews. Bulletin of Mathematical Biology, 63(3):597–601, May 2001. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/content/pdf/10.1006/bulm.1999.0223.pdf.

Filipe:2001:CAS

[1694] J. A. N. Filipe and G. J. Gibson. Comparing approximations to spatiotemporal models for epidemics with local spread. Bulletin of Mathematical Biology, 63(4):603–624, July 2001. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/content/pdf/10.1006/bulm.2001.0234.pdf.

Keener:2001:DIO

[1695] James P. Keener. Diffusion induced oscillatory insulin secretion. Bulletin of Mathematical Biology, 63(4):625–641, July 2001. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.

springer.com/content/pdf/10.1006/bulm.2001.0235.pdf.

Oosawa:2001:SSG

[1696] Fumio Oosawa. Spontaneous signal generation in living cells. Bulletin of Mathematical Biology, 63(4):643–654, July 2001. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.

springer.com/content/pdf/10.1006/bulm.2001.0236.pdf.

Owen:2001:HPC

[1697] M. R. Owen and M. A. Lewis. How predation can slow, stop or reverse a prey invasion. Bulletin of Mathematical Biology, 63(4):655– 684, July 2001. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL https://link.springer.com/content/pdf/
10.1006/bulm.2001.0239.pdf.

Utzny:2001:SDI

[1698] C. Utzny and N. J. Burroughs. Stability of a diverse immunological memory is determined by T cell population dynamics. Bulletin of Mathematical Biology, 63(4):685–713, July 2001. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/content/pdf/10.1006/bulm.2001.0242.pdf.

Wilke:2001:AEN

[1699] Claus O. Wilke. Adaptive evolution on neutral networks. Bulletin of Mathematical Biology, 63(4):715–730, July 2001. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.

springer.com/content/pdf/10.1006/bulm.2001.0244.pdf.

Wu:2001:MAV

[1700] Joseph T. Wu, Helen M. Byrne, David H. Kirn, and Lawrence M. Wein. Modeling and analysis of a virus that replicates selectively in tumor cells. Bulletin of Mathematical Biology, 63(4):731–768, July 2001. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https:

//link.springer.com/content/pdf/10.1006/bulm.2001.0245.pdf.

Kooi:2001:IRS

[1701] B. W. Kooi, T. G. Hallam, F. D. L. Kelpin, C. M. Krohn, and S. A. L. M. Kooijman. Iteroparous reproduction strategies and population dynamics. Bulletin of Mathematical Biology, 63(4):769–794, July 2001. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https:

//link.springer.com/content/pdf/10.1006/bulm.2001.0246.pdf.

DeAngelis:2001:BR

[1702] Donald L. DeAngelis and Germinal Cocho. Book reviews. Bulletin of Mathematical Biology, 63(4):795–799, July 2001. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.

springer.com/content/pdf/10.1006/bulm.2001.0233.pdf.

Levine:2001:MMC

[1703] Howard A. Levine, Serdal Pamuk, Brian D. Sleeman, and Marit Nilsen-Hamilton. Mathematical modeling of capillary formation and development in tumor angiogenesis: Penetration into the stroma. Bulletin of Mathematical Biology, 63(5):801–863, September 2001. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/content/pdf/10.1006/bulm.2001.0240.pdf.

Schollnberger:2001:ABI

[1704] H. Schöllnberger, B. R. Scott, and T. E. Hanson. Application of Bayesian inference to characterize risks associated with low doses of low-LET radiation. Bulletin of Mathematical Biology, 63(5):865–883, September 2001. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/content/pdf/10.1006/bulm.2001.
0243.pdf.

Lansky:2001:RHE

[1705] Petr Lánský and Wayne M. Getz. Receptor heterogeneity and its effect on sensitivity and coding range in olfactory sensory neurons. Bulletin of Mathematical Biology, 63(5):885–908, September 2001. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/content/pdf/10.1006/bulm.2001.0249.pdf.

Cox:2001:DIC

[1706] Steven J. Cox and Lin Ji. Discerning ionic currents and their kinetics from input impedance data. Bulletin of Mathematical Biology, 63(5):909– 932, September 2001. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1006/bulm.2001.0250.pdf.

Fang:2001:CCL

[1707] Yi Fang and Susan R. Wilson. On cloning and clone libraries for finite and infinite length genomes. Bulletin of Mathematical Biology, 63(5):933– 950, September 2001. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1006/bulm.2001.0251.pdf.

Sumpter:2001:AAP

[1708] D. J. T. Sumpter, G. B. Blanchard, and D. S. Broomhead. Ants and agents: a process algebra approach to modelling ant colony behaviour. Bulletin of Mathematical Biology, 63(5):951–980, September 2001. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/content/pdf/10.1006/bulm.2001.
0252.pdf.

Gyllenberg:2001:NSC

[1709] Mats Gyllenberg and Kalle Parvinen. Necessary and sufficient conditions for evolutionary suicide. Bulletin of Mathematical Biology, 63(5):981– 993, September 2001. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1006/bulm.2001.0253.pdf.

Pounds:2001:EFC

[1710] Stan Pounds. Estimating the fraction of clonable genomic DNA. Bulletin of Mathematical Biology, 63(5):995–1002, September 2001. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/content/pdf/10.1006/bulm.2001.0254.pdf.

Kozusko:2001:BR

[1711] Frank Kozusko. Book review. Bulletin of Mathematical Biology, 63(5): 1003, September 2001. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/content/
pdf/10.1006/bulm.2001.0256.pdf.

Engel:2001:MMC

[1712] Robert Engel, Mark Normand, Joseph Horowitz, and Micha Peleg. A model of microbial contamination of a water reservoir. Bulletin of Mathematical Biology, 63(6):1005–1023, November 2001. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/content/pdf/10.1006/bulm.2001.0238.pdf.

Yusim:2001:OVH

[1713] K. Yusim, H. Parnas, and L. A. Segel. One-vesicle hypothesis for neurotransmitter release: a possible molecular mechanism. Bulletin of Mathematical Biology, 63(6):1025–1040, November 2001. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/content/pdf/10.1006/bulm.2001.0247.pdf.

Coutinho:2001:MNH

[1714] F. A. B. Coutinho, L. F. Lopez, M. N. Burattini, and E. Massad. Modelling the natural history of HIV infection in individuals and its epidemiological implications. Bulletin of Mathematical Biology, 63(6):1041–1062, November 2001. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1006/bulm.2001.0257.pdf.

Carmel:2001:ESO

[1715] Liran Carmel, David Harel, and Doron Lancet. Estimating the size of the olfactory repertoire. Bulletin of Mathematical Biology, 63(6):1063–1078, November 2001. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1006/bulm.2001.0259.pdf.

Theraulaz:2001:MDD

[1716] Guy Theraulaz, Eric Bonabeau, Christian Sauwens, Jean-Louis Deneubourg, Arnaud Lioni, François Libert, Luc Passera, and Ricard Solé. Model of droplet dynamics in the Argentine ant Linepithema humile (Mayr). Bulletin of Mathematical Biology, 63(6):1079–1093, November 2001. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1006/bulm.2001.0260.pdf.

Ebert:2001:CCP

[1717] Ute Ebert, Manuel Arrayás, Nico Temme, Ben Sommeijer, and Jef Huisman. Critical conditions for phytoplankton blooms. Bulletin of Mathematical Biology, 63(6):1095–1124, November 2001. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/content/pdf/10.1006/bulm.2001.0261.pdf.

Brandts:2001:TCM

[1718] Wendy A. M. Brandts, André Longtin, and Lynn E. H. Trainor. Twocategory model of task allocation with application to ant societies. Bulletin of Mathematical Biology, 63(6):1125–1161, November 2001. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/content/pdf/10.1006/bulm.2001.
0262.pdf.

Oliveira:2001:ACM

[1719] Joseph S. Oliveira, Colin G. Bailey, Janet B. Jones-Oliveira, and David A. Dixon. An algebraic-combinatorial model for the identification and mapping of biochemical pathways. Bulletin of Mathematical Biology, 63(6):1163–1196, November 2001. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/content/pdf/10.1006/bulm.2001.0263.pdf.

Voit:2001:BR

[1720] Eberhard O. Voit. Book review. Bulletin of Mathematical Biology, 63(6):1197–1198, November 2001. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1006/bulm.2001.0248.pdf.

Barahona:2001:BR

[1721] Ana Barahona. Book review. Bulletin of Mathematical Biology, 63(6): 1199–1200, November 2001. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/content/
pdf/10.1006/bulm.2001.0255.pdf.

Maini:2002:E

[1722] Philip K. Maini. Editorial. Bulletin of Mathematical Biology, 64(1): 1, January 2002. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1006/bulm.2001.0284.pdf.

Aittokallio:2002:AHR

[1723] T. Aittokallio, M. Gyllenberg, and O. Polo. Adjustment of the human respiratory system to increased upper airway resistance during sleep. Bulletin of Mathematical Biology, 64(1):3–28, January 2002. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/content/pdf/10.1006/bulm.2001.0264.pdf.

Callaway:2002:HIL

[1724] Duncan S. Callaway and Alan S. Perelson. HIV-1 infection and low steady state viral loads. Bulletin of Mathematical Biology, 64(1):29– 64, January 2002. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1006/bulm.2001.0266.pdf.

Cobbold:2002:LOS

[1725] C. A. Cobbold, J. A. Sherratt, and S. R. J. Maxwell. Lipoprotein oxidation and its significance for atherosclerosis: a mathematical approach. Bulletin of Mathematical Biology, 64(1):65–95, January 2002. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/content/pdf/10.1006/bulm.2001.0267.pdf.

Albanese:2002:PBP

[1726] R. A. Albanese, H. T. Banks, M. V. Evans, and L. K. Potter. Physiologically based pharmacokinetic models for the transport of trichloroethylene in adipose tissue. Bulletin of Mathematical Biology, 64(1):97–131, January 2002. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1006/bulm.2001.0268.pdf.

Kam:2002:GAE

[1727] Zvi Kam. Generalized analysis of experimental data for interrelated biological measurements. Bulletin of Mathematical Biology, 64(1):133– 145, January 2002. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1006/bulm.2001.0269.pdf.

Mohtashemi:2002:QAA

[1728] Mojdeh Mohtashemi and Richard Levins. Qualitative analysis of the all-cause black-white mortality crossover. Bulletin of Mathematical Biology, 64(1):147–173, January 2002. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1006/bulm.2001.0270.pdf.

Gielen:2002:OBS

[1729] J. L. W. Gielen and S. Kranenbarg. Oxygen balance for small organisms: an analytical model. Bulletin of Mathematical Biology, 64(1):175– 207, January 2002. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1006/bulm.2001.0275.pdf.

Sole:2002:BR

[1730] Ricard V. Sole and Javier G. P. Gamarra. Book review. Bulletin of Mathematical Biology, 64(1):209–212, January 2002. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/content/pdf/10.1006/bulm.2000.0208.pdf.

Maly:2002:DAS

[1731] Ivan V. Maly. Diffusion approximation of the stochastic process of microtubule assembly. Bulletin of Mathematical Biology, 64(2):213– 238, March 2002. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1006/bulm.2001.0265.pdf.

Koerber:2002:MMP

[1732] A. J. Koerber, J. R. King, J. P. Ward, P. Williams, J. M. Croft, and R. E. Sockett. A mathematical model of partial-thickness burn-wound infection by Pseudomonas aeruginosa: Quorum sensing and the build-up to invasion. Bulletin of Mathematical Biology, 64(2):239–259, March 2002. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/content/pdf/10.1006/bulm.2001.
0272.pdf.

Briscoe:2002:HRF

[1733] Brian K. Briscoe, Mark A. Lewis, and Stephen E. Parrish. Home range formation in wolves due to scent marking. Bulletin of Mathematical Biology, 64(2):261–284, March 2002. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1006/bulm.2001.0273.pdf.

Khosroyani:2002:DMD

[1734] Madjid Khosroyani and George K. Hung. A dual-mode dynamic model of the human accommodation system. Bulletin of Mathematical Biology, 64(2):285–299, March 2002. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1006/bulm.2001.0274.pdf.

Marin-Sanguino:2002:MSS

[1735] Alberto Marín-Sanguino and Néstor V. Torres. Modelling, steady state analysis and optimization of the catalytic efficiency of the triosephosphate isomerase. Bulletin of Mathematical Biology, 64(2):301–326, March 2002. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1006/bulm.2001.0276.pdf.

Maree:2002:MDD

[1736] Athanasius F. M. Marée and Paulien Hogeweg. Modelling Dictyostelium discoideum morphogenesis: The culmination. Bulletin of Mathematical Biology, 64(2):327–353, March 2002. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1006/bulm.2001.0277.pdf.

Bertuzzi:2002:KHE

[1737] Alessandro Bertuzzi, Mario Faretta, Alberto Gandolfi, Carmela Sinisgalli, Giuseppe Starace, Giorgio Valoti, and Paolo Ubezio. Kinetic heterogeneity of an experimental tumour revealed by BrdUrd incorporation and mathematical modelling. Bulletin of Mathematical Biology, 64(2): 355–384, March 2002. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1006/bulm.2001.0280.pdf.

Ribeiro:2002:MDG

[1738] Ruy M. Ribeiro, Hiroshi Mohri, David D. Ho, and Alan S. Perelson. Modeling deuterated glucose labeling of T-lymphocytes. Bulletin of Mathematical Biology, 64(2):385–405, March 2002. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/content/pdf/10.1006/bulm.2001.0282.pdf.

Kagawa:2002:SSS

[1739] Yuki Kagawa and Yoshimi Tsuchiya. Stochasticity of the step size in the force generating process due to a single myosin molecule. Bulletin of Mathematical Biology, 64(2):407–419, March 2002. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/content/pdf/10.1006/bulm.2002.0285.pdf.

Fister:2002:BR

[1740] Renee Fister. Book review. Bulletin of Mathematical Biology, 64(2):421– 422, March 2002. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1006/bulm.2000.0278.pdf.

Levine:2002:MMC

[1741] H. A. Levine, S. Pamuk, B. D. Sleeman, and M. Nilsen-Hamilton. Mathematical modeling of capillary formation and development in tumor angiogenesis: Penetration into the stroma. Bulletin of Mathematical Biology, 64(2):423, March 2002. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1006/bulm.2002.0294.pdf.

Bergmann:2002:HIF

[1742] Claudia Bergmann, J. Leo van Hemmen, and Lee A. Segel. How instruction and feedback can select the appropriate T helper response. Bulletin of Mathematical Biology, 64(3):425–446, May 2002. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/content/pdf/10.1006/bulm.2001.0258.pdf.

Kuske:2002:AAN

[1743] R. Kuske and S. M. Baer. Asymptotic analysis of noise sensitivity in a neuronal burster. Bulletin of Mathematical Biology, 64(3):447– 481, May 2002. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1006/bulm.2002.0279.pdf.

Lafrance:2002:SEM

[1744] Patrick Lafrance, Vincent Lemaire, France Varin, François Donati, Jacques Bélair, and Fahima Nekka. Spatial effects in modeling pharmacokinetics of rapid action drugs. Bulletin of Mathematical Biology, 64(3): 483–500, May 2002. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1006/bulm.2001.0281.pdf.

Madzvamuse:2002:NAS

[1745] Anotida Madzvamuse, Roger D. K. Thomas, Philip K. Maini, and Andrew J. Wathen. A numerical approach to the study of spatial pattern formation in the ligaments of arcoid bivalves. Bulletin of Mathematical Biology, 64(3):501–530, May 2002. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1006/bulm.2002.0283.pdf.

Broday:2002:MNP

[1746] David M. Broday. Motion of nanobeads proximate to plasma membranes during single particle tracking. Bulletin of Mathematical Biology, 64(3): 531–563, May 2002. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1006/bulm.2002.0289.pdf.

Bees:2002:QEM

[1747] M. A. Bees, P. Andresén, E. Mosekilde, and M. Givskov. Quantitative effects of medium hardness and nutrient availability on the swarming motility of Serratia liquefaciens. Bulletin of Mathematical Biology, 64(3): 565–587, May 2002. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1006/bulm.2002.0287.pdf.

Biegeleisen:2002:TNL

[1748] Ken Biegeleisen. Topologically non-linked circular duplex DNA. Bulletin of Mathematical Biology, 64(3):589–609, May 2002. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/content/pdf/10.1006/bulm.2002.0288.pdf.

Gamarra:2002:CDD

[1749] Javier G. P. Gamarra and Ricard V. Solé. Complex discrete dynamics from simple continuous population models. Bulletin of Mathematical Biology, 64(3):611–620, May 2002. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1006/bulm.2002.0286.pdf.

Soberon:2002:BR

[1750] Jorge Soberón. Book review. Bulletin of Mathematical Biology, 64(3): 621–622, May 2002. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1006/bulm.2002.0290.pdf.

Sherratt:2002:MBF

[1751] Jonathan A. Sherratt, Richard Weller, and Nicholas J. Savill. Modelling blood flow regulation by nitric oxide in psoriatic plaques. Bulletin of Mathematical Biology, 64(4):623–641, July 2002. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.

springer.com/content/pdf/10.1006/bulm.2001.0271.pdf.

Yoshigaki:2002:CPW

[1752] Tomoyoshi Yoshigaki. The cleavage plane will bend when one aster of the mitotic apparatus stops growing in compressed sea urchin eggs. Bulletin of Mathematical Biology, 64(4):643–672, July 2002. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.

springer.com/content/pdf/10.1006/bulm.2002.0298.pdf.

McDougall:2002:MMF

[1753] S. R. McDougall, A. R. A. Anderson, M. A. J. Chaplain, and J. A. Sherratt. Mathematical modelling of flow through vascular networks: Implications for tumour-induced angiogenesis and chemotherapy strategies. Bulletin of Mathematical Biology, 64(4):673–702, July 2002. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https:

//link.springer.com/content/pdf/10.1006/bulm.2002.0293.pdf.

Shpak:2002:SUCa

[1754] Max Shpak and Kevin Atteson. A survey of unequal crossover systems and their mathematical properties. Bulletin of Mathematical Biology, 64 (4):703–746, July 2002. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/content/
pdf/10.1006/bulm.2001.0299.pdf.

Crampin:2002:PFR

[1755] E. J. Crampin, W. W. Hackborn, and P. K. Maini. Pattern formation in reaction–diffusion models with nonuniform domain growth. Bulletin of Mathematical Biology, 64(4):747–769, July 2002. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.

springer.com/content/pdf/10.1006/bulm.2002.0295.pdf.

Miklos:2002:IAS

[1756] István Miklós. An improved algorithm for statistical alignment of sequences related by a star tree. Bulletin of Mathematical Biology, 64 (4):771–779, July 2002. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/content/
pdf/10.1006/bulm.2002.0300.pdf.

vandenBerg:2002:QSL

[1757] H. A. van den Berg, N. J. Burroughs, and D. A. Rand. Quantifying the strength of ligand antagonism in TCR triggering. Bulletin of Mathematical Biology, 64(4):781–808, July 2002. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.

springer.com/content/pdf/10.1006/bulm.2002.0302.pdf.

Fishman:2002:FAT

[1758] Michael A. Fishman and Lewi Stone. Fertility assurance through extrapair fertilization, and male parental effort. Bulletin of Mathematical Biology, 64(4):809–823, July 2002. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/content/pdf/10.1006/bulm.2002.0301.pdf.

Simonetti:2002:BR

[1759] Javier A. Simonetti. Book review. Bulletin of Mathematical Biology, 64 (4):825–826, July 2002. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/content/
pdf/10.1006/bulm.2002.0291.pdf.

Cushing:2002:BR

[1760] J. M. Cushing. Book review. Bulletin of Mathematical Biology, 64 (4):827–828, July 2002. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/content/
pdf/10.1006/bulm.2000.0292.pdf.

Laing:2002:TVM

[1761] Carlo R. Laing and André Longtin. A two-variable model of somaticdendritic interactions in a bursting neuron. Bulletin of Mathematical Biology, 64(5):829–860, September 2002. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1006/bulm.2002.0303.pdf.

Shochat:2002:HHS

[1762] Eliezer Shochat, Salomon M. Stemmer, and Lee Segel. Human haematopoiesis in steady state and following intense perturbations. Bulletin of Mathematical Biology, 64(5):861–886, September 2002. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/content/pdf/10.1006/bulm.2002.0305.pdf.

Hlavacek:2002:KPR

[1763] William S. Hlavacek, Antonio Redondo, Carla Wofsy, and Byron Goldstein. Kinetic proofreading in receptor-mediated transduction of cellular signals: Receptor aggregation, partially activated receptors, and cytosolic messengers. Bulletin of Mathematical Biology, 64(5):887–911, September 2002. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1006/bulm.2002.0306.pdf.

Alonso:2002:EDM

[1764] David Alonso and Alan McKane. Extinction dynamics in mainland– island metapopulations: an N-patch stochastic model. Bulletin of Mathematical Biology, 64(5):913–958, September 2002. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/content/pdf/10.1006/bulm.2002.0307.pdf.

Gyllenberg:2002:VTS

[1765] Mats Gyllenberg, Diana Preoteasa, and Kari Saikkonen. Vertically transmitted symbionts in structured host metapopulations. Bulletin of Mathematical Biology, 64(5):959–978, September 2002. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/content/pdf/10.1006/bulm.2002.0309.pdf.

Farrell:2002:ETC

[1766] Brenda Farrell and Steven J. Cox. Estimating the time course of pore expansion during the spike phase of exocytotic release in mast cells of the beige mouse. Bulletin of Mathematical Biology, 64(5):979–1010, September 2002. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1006/bulm.2002.0308.pdf.

Craft:2002:MMI

[1767] David L. Craft, Lawrence M. Wein, and Dennis J. Selkoe. A mathematical model of the impact of novel treatments on the Aβ burden in the Alzheimer’s brain, CSF and plasma. Bulletin of Mathematical Biology, 64(5):1011–1031, September 2002. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1006/bulm.2002.0304.pdf.

Nilsson:2002:OMR

[1768] Martin Nilsson and Nigel Snoad. Optimal mutation rates in dynamic environments. Bulletin of Mathematical Biology, 64(6):1033–1043, November 2002. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/content/pdf/10.1006/
bulm.2002.0314.pdf.

Planque:2002:ADS

[1769] Robert Planqué, Nicholas F. Britton, Nigel R. Franks, and Mark A. Peletier. The adaptiveness of defence strategies against cuckoo parasitism. Bulletin of Mathematical Biology, 64(6):1045–1068, November 2002. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/content/pdf/10.1006/
bulm.2002.0311.pdf.

Tasaki:2002:CTN

[1770] Ichiji Tasaki and Gen Matsumoto. On the cable theory of nerve conduction. Bulletin of Mathematical Biology, 64(6):1069–1082, November 2002. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/content/pdf/10.1006/
bulm.2002.0310.pdf.

Jorgensen:2002:MMH

[1771] Peter Jørgensen, Carsten Schäfer, Peter G. Guerra, Mario Talajic, Stanley Nattel, and Leon Glass. A mathematical model of human atrioventricular nodal function incorporating concealed conduction. Bulletin of Mathematical Biology, 64(6):1083–1099, November 2002. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://

link.springer.com/content/pdf/10.1006/bulm.2002.0313.pdf.

Page:2002:ELF

[1772] Karen M. Page and Martin A. Nowak. Empathy leads to fairness. Bulletin of Mathematical Biology, 64(6):1101–1116, November 2002. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/content/pdf/10.1006/
bulm.2002.0321.pdf.

Burns:2002:PDS

[1773] Jonathan Burns, Geoff Clarke, and Charles J. Lumsden. Photoreceptor death: Spatiotemporal patterns arising from one-hit death kinetics and a diffusible cell death factor. Bulletin of Mathematical Biology, 64(6): 1117–1145, November 2002. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/content/
pdf/10.1006/bulm.2002.0320.pdf.

Gerrish:2002:BR

[1774] Philip Gerrish. Book review. Bulletin of Mathematical Biology, 64(6): 1147, November 2002. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/content/
pdf/10.1006/bulm.2002.0318.pdf.

Shpak:2002:SUCb

[1775] M. Shpak and K. Atteson. A survey of unequal crossover systems and their mathematical properties. Bulletin of Mathematical Biology, 64(6):

1149, November 2002. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/content/
pdf/10.1006/bulm.2002.0312.pdf.

Sattenspiel:2003:SEQ

[1776] Lisa Sattenspiel and D. Ann Herring. Simulating the effect of quarantine on the spread of the 1918–19 flu in central Canada. Bulletin of Mathematical Biology, 65(1):1–26, January 2003. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/content/pdf/10.1006/bulm.2002.0317.pdf.

Dopazo:2003:MEA

[1777] H. Dopazo, M. B. Gordon, R. Perazzo, and S. Risau-Gusman. A model for the emergence of adaptive subsystems. Bulletin of Mathematical Biology, 65(1):27–56, January 2003. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1006/bulm.2002.0315.pdf.

Jockusch:2003:STT

[1778] Harald Jockusch and Andreas Dress. From sphere to torus: a topological view of the metazoan body plan. Bulletin of Mathematical Biology, 65(1): 57–65, January 2003. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1006/bulm.2002.0319.pdf.

Mitchener:2003:CEC

[1779] W. Garrett Mitchener and Martin A. Nowak. Competitive exclusion and coexistence of universal grammars. Bulletin of Mathematical Biology, 65(1):67–93, January 2003. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1006/bulm.2002.0322.pdf.

Krishnan:2003:AST

[1780] J. Krishnan and P. A. Iglesias. Analysis of the signal transduction properties of a module of spatial sensing in eukaryotic chemotaxis. Bulletin of Mathematical Biology, 65(1):95–128, January 2003. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://

link.springer.com/content/pdf/10.1006/bulm.2002.0323.pdf.

Fox:2003:LRB

[1781] Gordon A. Fox and Alan Hastings. Limiting relationships between selection and recombination. Bulletin of Mathematical Biology, 65(1):129– 141, January 2003. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1006/bulm.2002.0324.pdf.

Huang:2003:IIM

[1782] Yunxin Huang and Odo Diekmann. Interspecific influence on mobility and Turing instability. Bulletin of Mathematical Biology, 65(1):143– 156, January 2003. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1006/bulm.2002.0328.pdf.

Clark:2003:MSP

[1783] Leona Harris Clark, Paul M. Schlosser, and James F. Selgrade. Multiple stable periodic solutions in a model for hormonal control of the menstrual cycle. Bulletin of Mathematical Biology, 65(1):157–173, January 2003. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/content/pdf/10.1006/bulm.2002.
0326.pdf.

Hernandez:2003:SPE

[1784] Julio A. Hernández. Stability properties of elementary dynamic models of membrane transport. Bulletin of Mathematical Biology, 65(1):175– 197, January 2003. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1006/bulm.2002.0325.pdf.

Salmon-Divon:2003:GNK

[1785] Mali Salmon-Divon, Petter Höglund, and Ramit Mehr. Generation of the natural killer cell repertoire: The sequential vs. the two-step selection model. Bulletin of Mathematical Biology, 65(2):199–218, March 2003. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1016/S0092-8240%2802%2900093-9.pdf.

Peleg:2003:EMI

[1786] Micha Peleg, Mark D. Normand, and Osvaldo H. Campanella. Estimating microbial inactivation parameters from survival curves obtained under varying conditions — the linear case. Bulletin of Mathematical Biology, 65(2):219–234, March 2003. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1016/S0092-8240%2802%2900097-6.pdf.

Landman:2003:MMC

[1787] K. A. Landman, G. J. Pettet, and D. F. Newgreen. Mathematical models of cell colonization of uniformly growing domains. Bulletin of Mathematical Biology, 65(2):235–262, March 2003. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/content/pdf/10.1016/S0092-8240%2802%2900098-8.pdf.

Percus:2003:DVB

[1788] Jerome K. Percus, Ora E. Percus, Martin Markowitz, David D. Ho, Michele Di Mascio, and Alan S. Perelson. The distribution of viral blips observed in HIV-1 infected patients treated with combination antiretroviral therapy. Bulletin of Mathematical Biology, 65(2):263– 277, March 2003. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1016/S0092-8240%2802%2900095-2.pdf.

Garcia-Meseguer:2003:MRT

[1789] M. J. García-Meseguer, J. A. Vidal de Labra, M. García-Moreno, F. García-Cánovas, B. H. Havsteen, and R. Varón. Mean residence times in linear compartmental systems. Symbolic formulae for their direct evaluation. Bulletin of Mathematical Biology, 65(2):279–308, March 2003. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1016/S0092-8240%2802%2900096-4.pdf.

Sadovsky:2003:MCN

[1790] Michael G. Sadovsky. The method to compare nucleotide sequences based on the minimum entropy principle. Bulletin of Mathematical Biology, 65(2):309–322, March 2003. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1016/S0092-8240%2802%2900107-6.pdf.

Ebenhoh:2003:SDM

[1791] Oliver Ebenhöh and Reinhart Heinrich. Stoichiometric design of metabolic networks: Multifunctionality, clusters, optimization, weak and strong robustness. Bulletin of Mathematical Biology, 65(2):323– 357, March 2003. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1016/S0092-8240%2803%2900002-8.pdf.

Kesmir:2003:CER

[1792] Can Keşmir and Rob J. De Boer. Clonal exhaustion as a result of immune deviation. Bulletin of Mathematical Biology, 65(3):359–374, May 2003. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/content/pdf/10.1016/
S0092-8240%2803%2900009-0.pdf.

Louzoun:2003:PCD

[1793] Yoram Louzoun, Sorin Solomon, Henri Atlan, and Irun R. Cohen. Proliferation and competition in discrete biological systems. Bulletin of Mathematical Biology, 65(3):375–396, May 2003. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.

springer.com/content/pdf/10.1016/S0092-8240%2803%2900007-7.
pdf.

Barbour:2003:DSE

[1794] A. D. Barbour and Roman Kälin. The duration of a sequence of epidemics. Bulletin of Mathematical Biology, 65(3):397–405, May 2003. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/content/pdf/10.1016/
S0092-8240%2803%2900010-7.pdf.

Ergun:2003:OSR

[1795] Ayla Ergun, Kevin Camphausen, and Lawrence M. Wein. Optimal scheduling of radiotherapy and angiogenic inhibitors. Bulletin of Mathematical Biology, 65(3):407–424, May 2003. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.

springer.com/content/pdf/10.1016/S0092-8240%2803%2900006-5.
pdf.

Petrovskii:2003:QSA

[1796] Sergei Petrovskii, Bai-Lian Li, and Horst Malchow. Quantification of the spatial aspect of chaotic dynamics in biological and chemical systems. Bulletin of Mathematical Biology, 65(3):425–446, May 2003. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/content/pdf/10.1016/
S0092-8240%2803%2900004-1.pdf.

Boswell:2003:GFF

[1797] Graeme P. Boswell, Helen Jacobs, Fordyce A. Davidson, Geoffrey M. Gadd, and Karl Ritz. Growth and function of fungal mycelia in heterogeneous environments. Bulletin of Mathematical Biology, 65(3):447– 477, May 2003. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL https://link.springer.com/content/pdf/
10.1016/S0092-8240%2803%2900003-X.pdf.

Tang:2003:MAS

[1798] Sanyi Tang and Lansun Chen. Multiple attractors in stage-structured population models with birth pulses. Bulletin of Mathematical Biology, 65(3):479–495, May 2003. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/content/
pdf/10.1016/S0092-8240%2803%2900005-3.pdf.

Kuang:2003:BHA

[1799] Yang Kuang, William F. Fagan, and Irakli Loladze. Biodiversity, habitat area, resource growth rate and interference competition. Bulletin of Mathematical Biology, 65(3):497–518, May 2003. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.

springer.com/content/pdf/10.1016/S0092-8240%2803%2900008-9.
pdf.

Abramson:2003:TWI

[1800] G. Abramson, V. M. Kenkre, T. L. Yates, and R. R. Parmenter. Traveling waves of infection in the hantavirus epidemics. Bulletin of Mathematical Biology, 65(3):519–534, May 2003. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/content/pdf/10.1016/S0092-8240%2803%2900013-2.pdf.

Louzoun:2003:DAD

[1801] Yoram Louzoun, Martin Weigert, and Gyan Bhanot. Dynamical analysis of a degenerate primary and secondary humoral immune response. Bulletin of Mathematical Biology, 65(3):535–545, May 2003. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/content/pdf/10.1016/
S0092-8240%2803%2900023-5.pdf.

Gumel:2003:BRM

[1802] Abba Gumel and Lloyd Demetrius. Book reviews: Mathematical Approaches for Emerging and Re-emerging Infectious Diseases: an Introduction by Carlos Castillo Chavez, Sally Blower, Pauline van der Driessche, Denise Kirshner and Abdul Aziz Yakubu, 2002. IMA Volumes in Mathematics and its Applications 125, Springer-Verlag. $79.95, ISBN 0-387-95354-X. Bulletin of Mathematical Biology, 65(3):547–552, May 2003. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/content/pdf/10.1016/
S0092-8240%2803%2900011-9.pdf.

Hung:2003:EPF

[1803] Hui-Chih Hung, Yu-Hou Chen, Guang-Yaw Liu, Hwei-Jen Lee, and Gu-Gang Chang. Equilibrium protein folding-unfolding process involving multiple intermediates. Bulletin of Mathematical Biology, 65(4):553– 570, July 2003. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1016/S0092-8240%2803%2900024-7.pdf.

Nagata:2003:RDM

[1804] Wayne Nagata, Lionel G. Harrison, and Stephan Wehner. Reaction– diffusion models of growing plant tips: Bifurcations on hemispheres. Bulletin of Mathematical Biology, 65(4):571–607, July 2003. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1016/
S0092-8240%2803%2900025-9.pdf.

Breward:2003:MMD

[1805] Christopher J. W. Breward, Helen M. Byrne, and Claire E. Lewis. A multiphase model describing vascular tumour growth. Bulletin of Mathematical Biology, 65(4):609–640, July 2003. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/content/pdf/10.1016/S0092-8240%2803%2900027-2.pdf.

Reinker:2003:RNS

[1806] Stefan Reinker, Ernest Puil, and Robert M. Miura. Resonances and noise in a stochastic Hindmarsh–Rose model of thalamic neurons. Bulletin of Mathematical Biology, 65(4):641–663, July 2003. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/content/pdf/10.1016/S0092-8240%2803%2900026-0.
pdf.

Marcano-Velazquez:2003:IAM

[1807] M. Marcano-Velázquez and Harold E. Layton. An inverse algorithm for a mathematical model of an avian urine concentrating mechanism. Bulletin of Mathematical Biology, 65(4):665–691, July 2003. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1016/
S0092-8240%2803%2900029-6.pdf.

Luca:2003:CSM

[1808] Magdalena Luca, Alexandra Chavez-Ross, Leah Edelstein-Keshet, and Alex Mogilner. Chemotactic signaling, microglia, and Alzheimer’s disease senile plaques: Is there a connection? Bulletin of Mathematical Biology, 65(4):693–730, July 2003. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1016/S0092-8240%2803%2900030-2.pdf.

Till:2003:NML

[1809] S. J. Till, J. Till, P. K. Milsom, and G. Rowlands. A new model for laser-induced thermal damage in the retina. Bulletin of Mathematical Biology, 65(4):731–746, July 2003. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1016/S0092-8240%2803%2900028-4.pdf.

Fahlman:2003:PME

[1810] Andreas Fahlman and Susan R. Kayar. Probabilistic modelling for estimating gas kinetics and decompression sickness risk in pigs during H_{2} biochemical decompression. Bulletin of Mathematical Biology, 65(4): 747–766, July 2003. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1016/S0092-8240%2803%2900038-7.pdf.

Mitchell:2003:TCM

[1811] Colleen C. Mitchell and David G. Schaeffer. A two-current model for the dynamics of cardiac membrane. Bulletin of Mathematical Biology, 65(5): 767–793, September 2003. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/content/
pdf/10.1016/S0092-8240%2803%2900041-7.pdf.

Nicolis:2003:OCC

[1812] S. C. Nicolis, C. Detrain, D. Demolin, and J. L. Deneubourg. Optimality of collective choices: a stochastic approach. Bulletin of Mathematical Biology, 65(5):795–808, September 2003. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/content/pdf/10.1016/S0092-8240%2803%2900040-5.pdf.

Lika:2003:LHI

[1813] Konstadia Lika and Sebastiaan A. L. M. Kooijman. Life history implications of allocation to growth versus reproduction in dynamic energy budgets. Bulletin of Mathematical Biology, 65(5):809–834, September 2003. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/content/pdf/10.1016/
S0092-8240%2803%2900039-9.pdf.

Beynon:2003:IVP

[1814] M. J. Beynon and K. L. Buchanan. An illustration of variable precision rough set theory: The gender classification of the European barn swallow (Hirundo rustica). Bulletin of Mathematical Biology, 65(5):835– 858, September 2003. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/content/
pdf/10.1016/S0092-8240%2803%2900044-2.pdf.

Layton:2003:RBM

[1815] Anita T. Layton and Harold E. Layton. A region-based model framework for the rat urine concentrating mechanism. Bulletin of Mathematical Biology, 65(5):859–901, September 2003. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/content/pdf/10.1016/S0092-8240%2803%2900045-4.pdf.

Bertuzzi:2003:RRT

[1816] Alessandro Bertuzzi, Alberto D’Onofrio, Antonio Fasano, and Alberto Gandolfi. Regression and regrowth of tumour cords following singledose anticancer treatment. Bulletin of Mathematical Biology, 65(5):903– 931, September 2003. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/content/
pdf/10.1016/S0092-8240%2803%2900054-5.pdf.

Chen:2003:MSG

[1817] C. Y. Chen, Y. Cordeaux, S. J. Hill, and J. R. King. Modelling of signalling via G-protein coupled receptors: Pathway-dependent agonist potency and efficacy. Bulletin of Mathematical Biology, 65(5):933–958, September 2003. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL https://link.springer.com/content/pdf/
10.1016/S0092-8240%2803%2900055-7.pdf.

Lara-Aparicio:2003:BRG

[1818] Miguel Lara-Aparicio. Book review: The geometry of biological time by Arthur Winfree, Springer-Verlag, New York, 2nd edition, $89.95 (hardcover), 800 pp. ISBN 0-387-98992-7, 2001. Bulletin of Mathematical Biology, 65(5):959–960, September 2003. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/content/pdf/10.1016/S0092-8240%2803%2900042-9.pdf.

Hansmann:2003:BRP

[1819] Ulrich H. E. Hansmann. Book review: Protein physics — a course of lectures by Alexei V. Finkelstein and Oleg B. Ptitsyn, 2002. Academic Press. $80.00US, 354 pp., ISBN 0-12-256781-1. Bulletin of Mathematical Biology, 65(5):961–962, September 2003. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/content/pdf/10.1016/S0092-8240%2803%2900043-0.pdf.

Marino:2003:IIC

[1820] Simeone Marino, Suman Ganguli, Ian M. P. Joseph, and Denise E. Kirschner. The importance of an inter-compartmental delay in a model for human gastric acid secretion. Bulletin of Mathematical Biology, 65(6): 963–990, November 2003. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/content/
pdf/10.1016/S0092-8240%2803%2900046-6.pdf.

Diaz:2003:VMM

[1821] Humberto González Díaz, Ronal Ramos De Armas, and Reinaldo Molina. Vibrational Markovian modelling of footprints after the interaction of antibiotics with the packaging region of HIV type 1. Bulletin of Mathematical Biology, 65(6):991–1002, November 2003. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/content/pdf/10.1016/S0092-8240%2803%2900064-8.pdf.

Coombs:2003:OVP

[1822] Daniel Coombs, Michael A. Gilchrist, Jerome Percus, and Alan S. Perelson. Optimal viral production. Bulletin of Mathematical Biology, 65(6): 1003–1023, November 2003. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/content/
pdf/10.1016/S0092-8240%2803%2900056-9.pdf.

Devloo:2003:IAS

[1823] Vincent Devloo, Pierre Hansen, and Martine Labbé. Identification of all steady states in large networks by logical analysis. Bulletin of Mathematical Biology, 65(6):1025–1051, November 2003. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/content/pdf/10.1016/S0092-8240%2803%2900061-2.pdf.

Chopp:2003:DQS

[1824] D. L. Chopp, M. J. Kirisits, B. Moran, and M. R. Parsek. The dependence of quorum sensing on the depth of a growing biofilm. Bulletin of Mathematical Biology, 65(6):1053–1079, November 2003. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/content/pdf/10.1016/S0092-8240%2803%2900057-0.
pdf.

Bonneuil:2003:MEM

[1825] Noël Bonneuil. Making ecosystem models viable. Bulletin of Mathematical Biology, 65(6):1081–1094, November 2003. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/content/pdf/10.1016/S0092-8240%2803%2900060-0.pdf.

Heyd:2003:MME

[1826] Andrew Heyd and Donald A. Drew. A mathematical model for elongation of a peptide chain. Bulletin of Mathematical Biology, 65(6):1095–1109, November 2003. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1016/S0092-8240%2803%2900076-4.pdf.

Tzafriri:2003:MMK

[1827] A. R. Tzafriri. Michaelis–Menten kinetics at high enzyme concentrations. Bulletin of Mathematical Biology, 65(6):1111–1129, November 2003. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1016/S0092-8240%2803%2900059-4.pdf.

Gorfine:2003:MLE

[1828] Malka Gorfine, Laurence Freedman, Gitit Shahaf, and Ramit Mehr. Maximum likelihood estimator and likelihood ratio test in complex models: an application to b lymphocyte development. Bulletin of Mathematical Biology, 65(6):1131–1139, November 2003. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/content/pdf/10.1016/S0092-8240%2803%2900062-4.pdf.

Smith:2003:SCE

[1829] Nicolas P. Smith. From sarcomere to cell: an efficient algorithm for linking mathematical models of muscle contraction. Bulletin of Mathematical Biology, 65(6):1141–1162, November 2003. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/content/pdf/10.1016/S0092-8240%2803%2900063-6.pdf.

Morgan:2003:PCR

[1830] John Morgan and Ken Steiglitz. Pairwise competition and the replicator equation. Bulletin of Mathematical Biology, 65(6):1163–1172, November 2003. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1016/j.
bulm.2003.08.001.pdf.

Rowcliffe:2003:BRE

[1831] Marcus Rowcliffe. Book review: Estimating animal abundance — closed populations by D. L. Borchers, W. Zucchini and S. T. Buckland, 2002. Springer-Verlag. $69.95, 314 pp., 91 illus. Hardcover, ISBN 1-85233-560-2. Bulletin of Mathematical Biology, 65(6):1173–1174, November 2003. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1016/S0092-8240%2803%2900058-2.pdf.

Oien:2004:DDB

[1832] Alf H. Øien. Daphnicle dynamics based on kinetic theory: an analoguemodelling of swarming and behaviour of Daphnia. Bulletin of Mathematical Biology, 66(1):1–46, January 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/content/pdf/10.1016/S0092-8240%2803%2900065-X.pdf.

Fishman:2004:IPM

[1833] Michael A. Fishman and Lewi Stone. Indiscriminate polyandry and male parental effort. Bulletin of Mathematical Biology, 66(1):47–63, January 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1016/S0092-8240%2803%2900066-1.pdf.

Lemon:2004:FDF

[1834] Greg Lemon. Fire-diffuse-fire calcium waves in confined intracellular spaces. Bulletin of Mathematical Biology, 66(1):65–90, January 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1016/S0092-8240%2803%2900074-0.pdf.

Gudelj:2004:ESM

[1835] I. Gudelj, K. A. J. White, and N. F. Britton. The effects of spatial movement and group interactions on disease dynamics of social animals. Bulletin of Mathematical Biology, 66(1):91–108, January 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1016/S0092-8240%2803%2900075-2.pdf.

Khan:2004:APP

[1836] Q. J. A. Khan, E. Balakrishnan, and G. C. Wake. Analysis of a predatorprey system with predator switching. Bulletin of Mathematical Biology, 66(1):109–123, January 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1016/j.bulm.2003.08.005.pdf.

Meyer-Hermann:2004:TSC

[1837] Michael Meyer-Hermann and Tilo Beyer. The type of seeder cells determines the efficiency of germinal center reactions. Bulletin of Mathematical Biology, 66(1):125–141, January 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/content/pdf/10.1016/j.bulm.2003.08.003.pdf.

Hsieh:2004:MSD

[1838] Ying-Hen Hsieh and Chien Hsun Chen. Modelling the social dynamics of a sex industry: Its implications for spread of HIV/AIDS. Bulletin of Mathematical Biology, 66(1):143–166, January 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/content/pdf/10.1016/j.bulm.2003.08.004.pdf.

Gracheva:2004:CMM

[1839] Maria E. Gracheva and Hans G. Othmer. A continuum model of motility in ameboid cells. Bulletin of Mathematical Biology, 66(1):167–193, January 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1016/j.
bulm.2003.08.007.pdf.

delaFuente:2004:BRC

[1840] Alberto de la Fuente and Pedro Mendes. Book review: Computational analysis of biochemical systems, by Eberhart O. Voit, November 2000. Paperback (hardback), 544 pages, $140.00, ISBN 0-521-78579-0. Bulletin of Mathematical Biology, 66(1):195–197, January 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1016/
S0092-8240%2803%2900053-3.pdf.

Rejniak:2004:CMM

[1841] Katarzyna A. Rejniak, Harvey J. Kliman, and Lisa J. Fauci. A computational model of the mechanics of growth of the villous trophoblast bilayer. Bulletin of Mathematical Biology, 66(2):199–232, March 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1016/j.
bulm.2003.06.001.pdf.

Biktashev:2004:PBF

[1842] V. N. Biktashev and J. Brindley. Phytoplankton blooms and fish recruitment rate: Effects of spatial distribution. Bulletin of Mathematical Biology, 66(2):233–259, March 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1016/j.bulm.2003.08.008.pdf.

deJong:2004:QSI

[1843] Hidde de Jong, Johannes Geiselmann, Grégory Batt, Céline Hernandez, and Michel Page. Qualitative simulation of the initiation of sporulation in Bacillus subtilis. Bulletin of Mathematical Biology, 66(2):261– 299, March 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1016/j.bulm.2003.08.009.pdf.

deJong:2004:QSG

[1844] Hidde de Jong, Jean-Luc Gouzé, Céline Hernandez, Michel Page, Tewfik Sari, and Johannes Geiselmann. Qualitative simulation of genetic regulatory networks using piecewise-linear models. Bulletin of Mathematical Biology, 66(2):301–340, March 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/content/pdf/10.1016/j.bulm.2003.08.010.pdf.

Brown:2004:EDD

[1845] David H. Brown and Benjamin M. Bolker. The effects of disease dispersal and host clustering on the epidemic threshold in plants. Bulletin of Mathematical Biology, 66(2):341–371, March 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/content/pdf/10.1016/j.bulm.2003.08.006.pdf.

Filipe:2004:IDS

[1846] J. A. N. Filipe, W. Otten, G. J. Gibson, and C. A. Gilligan. Inferring the dynamics of a spatial epidemic from time-series data. Bulletin of Mathematical Biology, 66(2):373–391, March 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/content/pdf/10.1016/j.bulm.2003.09.002.pdf.

Schnell:2004:BRP

[1847] Santiago Schnell. Book review: Parametric sensitivity in chemical systems by Arvind Varma, Massimo Morbidelli and Hua Wu, 1999. Cambridge series in chemical engineering, Cambridge University Press. £60.00/$90.00, ISBN 0-521-62171-2. Bulletin of Mathematical Biology, 66(2):393–395, March 2004. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/content/
pdf/10.1016/j.bulm.2003.08.002.pdf.

Garske:2004:MPM

[1848] Tini Garske and Uwe Grimm. A maximum principle for the mutationselection equilibrium of nucleotide sequences. Bulletin of Mathematical Biology, 66(3):397–421, May 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1016/j.bulm.2003.08.013.pdf.

Chen:2004:MPT

[1849] Kuan-Chou Chen, Chiung-Huei Peng, Hui-Er Wang, Chiung-Chi Peng, and Robert Y. Peng. Modeling of the pH- and the temperaturedependant deviations of the free to total PSA (Prostate Specific Antigen) ratios for clinical predictability of prostate cancer and benign prostate hyperplasia. Bulletin of Mathematical Biology, 66(3):423–445, May 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1016/j.
bulm.2003.08.014.pdf.

Wang:2004:EGS

[1850] Yun Wang. Entity grammar systems: a grammatical tool for studying the hierarchal structures of biological systems. Bulletin of Mathematical Biology, 66(3):447–471, May 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1016/j.bulm.2003.08.011.pdf.

Campos:2004:FBM

[1851] Paulo R. A. Campos. Fixation of beneficial mutations in the presence of epistatic interactions. Bulletin of Mathematical Biology, 66(3):473–486, May 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1016/j.
bulm.2003.08.012.pdf.

Huang:2004:DJM

[1852] Yunxin Huang, Odo Diekmann, and Frank Van Den Bosch. Double-jump migration and diffusive instability. Bulletin of Mathematical Biology, 66 (3):487–504, May 2004. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/content/
pdf/10.1016/j.bulm.2003.09.004.pdf.

Ipina:2004:IAM

[1853] Santiago L. Ipiña and Ana I. Durand. Inferential assessment of the MI index of sexual dimorphism: a comparative study with some other sexual dimorphism measures. Bulletin of Mathematical Biology, 66(3):505–522, May 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1016/j.
bulm.2003.09.003.pdf.

Wilson:2004:UQG

[1854] D. P. Wilson, S. Mathews, C. Wan, A. N. Pettitt, and D. L. S. McElwain. Use of a quantitative gene expression assay based on micro-array techniques and a mathematical model for the investigation of chlamydial generation time. Bulletin of Mathematical Biology, 66(3):523–537, May 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1016/j.
bulm.2003.09.001.pdf.

Labouriau:2004:LSP

[1855] Isabel S. Labouriau and Carla Alves-Pinto. Loss of synchronization in partially coupled Hodgkin–Huxley equations. Bulletin of Mathematical Biology, 66(3):539–557, May 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1016/j.bulm.2003.09.006.pdf.

Svetic:2004:MFG

[1856] R. E. Svetic, C. R. MacCluer, C. O. Buckley, K. L. Smythe, and J. H. Jackson. A metabolic force for gene clustering. Bulletin of Mathematical Biology, 66(3):559–581, May 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1016/j.bulm.2003.09.008.pdf.

Dugaw:2004:SLH

[1857] Christopher J. Dugaw, Alan Hastings, Evan L. Preisser, and Donald R. Strong. Seasonally limited host supply generates microparasite population cycles. Bulletin of Mathematical Biology, 66(3):583–594, May 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1016/j.
bulm.2003.09.005.pdf.

Maini:2004:BMB

[1858] Philip K. Maini, Santiago Schnell, and Sara Jolliffe. Bulletin of mathematical biology — facts, figures and comparisons. Bulletin of Mathematical Biology, 66(4):595–603, July 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/content/pdf/10.1016/j.bulm.2004.03.003.pdf.

Wu:2004:ATW

[1859] Joseph T. Wu, David H. Kirn, and Lawrence M. Wein. Analysis of a three-way race between tumor growth, a replication-competent virus and an immune response. Bulletin of Mathematical Biology, 66(4):605–625, July 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1016/j.
bulm.2003.08.016.pdf.

Miura:2004:SPA

[1860] Takashi Miura and Philip K. Maini. Speed of pattern appearance in reaction–diffusion models: Implications in the pattern formation of limb bud mesenchyme cells. Bulletin of Mathematical Biology, 66(4):627–649, July 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1016/j.
bulm.2003.09.009.pdf.

Page:2004:LLH

[1861] Karen M. Page. Language learning: How much evidence does a child need in order to learn to speak grammatically? Bulletin of Mathematical Biology, 66(4):651–662, July 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1016/j.bulm.2003.09.007.pdf.

Nagy:2004:CNS

[1862] John D. Nagy. Competition and natural selection in a mathematical model of cancer. Bulletin of Mathematical Biology, 66(4):663–687, July 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1016/j.
bulm.2003.10.001.pdf.

deMenezes:2004:RFM

[1863] Renée X. de Menezes, Neli R. S. Ortega, and Eduardo Massad. A Reed–Frost model taking into account uncertainties in the diagnostic of the infection. Bulletin of Mathematical Biology, 66(4):689–706, July 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1016/j.
bulm.2003.10.003.pdf.

Logan:2004:CCH

[1864] J. David Logan, Anthony Joern, and William Wolesensky. Control of CNP homeostasis in herbivore consumers through differential assimilation. Bulletin of Mathematical Biology, 66(4):707–725, July 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1016/j.
bulm.2003.10.008.pdf.

Iwasaki:2004:SFP

[1865] Yuishi Iwasaki and Sohei Gomi. Stochastic formulation for a partial neural circuit of C. elegans. Bulletin of Mathematical Biology, 66(4): 727–743, July 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1016/j.bulm.2003.10.007.pdf.

Verzi:2004:MAD

[1866] D. W. Verzi. Modeling activity-dependent synapse restructuring. Bulletin of Mathematical Biology, 66(4):745–762, July 2004. CODEN BMT-BAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/content/pdf/10.1016/j.bulm.2003.10.005.
pdf.

Bahi:2004:SGE

[1867] Jacques M. Bahi and Christian J. Michel. A stochastic gene evolution model with time dependent mutations. Bulletin of Mathematical Biology, 66(4):763–778, July 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1016/j.bulm.2003.10.004.pdf.

Yamada:2004:SSB

[1868] Hiroyasu Yamada, Reiko Tanaka, and Toshiyuki Nakagaki. Sequences of symmetry-breaking in phyllotactic transitions. Bulletin of Mathematical Biology, 66(4):779–789, July 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1016/j.bulm.2003.10.006.pdf.

Till:2004:SPM

[1869] S. J. Till, P. K. Milsom, and G. Rowlands. A simplified physical model of pressure wave dynamics and acoustic wave generation induced by laser absorption in the retina. Bulletin of Mathematical Biology, 66(4): 791–808, July 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1016/j.bulm.2003.10.009.pdf.

Klapper:2004:EHS

[1870] I. Klapper. Effect of heterogeneous structure in mechanically unstressed biofilms on overall growth. Bulletin of Mathematical Biology, 66(4): 809–824, July 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1016/j.bulm.2003.11.008.pdf.

Hall:2004:IDP

[1871] Richard J. Hall, Simon Gubbins, and Christopher A. Gilligan. Invasion of drug and pesticide resistance is determined by a trade-off between treatment efficacy and relative fitness. Bulletin of Mathematical Biology, 66(4):825–840, July 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1016/j.bulm.2003.11.006.pdf.

Kirupaharan:2004:CMP

[1872] Nadarajah Kirupaharan and Linda J. S. Allen. Coexistence of multiple pathogen strains in stochastic epidemic models with densitydependent mortality. Bulletin of Mathematical Biology, 66(4):841–864, July 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1016/j.
bulm.2003.11.007.pdf.

Huppert:2004:BEM

[1873] Amit Huppert, Ronen Olinky, and Lewi Stone. Bottom-up excitable models of phytoplankton blooms. Bulletin of Mathematical Biology, 66 (4):865–878, July 2004. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/content/
pdf/10.1016/j.bulm.2004.01.003.pdf.

Korobeinikov:2004:GPB

[1874] Andrei Korobeinikov. Global properties of basic virus dynamics models. Bulletin of Mathematical Biology, 66(4):879–883, July 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/content/pdf/10.1016/j.bulm.2004.02.001.
pdf.

Capurro:2004:NAJ

[1875] Alberto Capurro and C. P. Malta. Noise autocorrelation and jamming avoidance performance in pulse type electric fish. Bulletin of Mathematical Biology, 66(4):885–905, July 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/content/pdf/10.1016/j.bulm.2004.02.002.pdf.

Gonzalez:2004:QSS

[1876] Maykel Pérez González and Maria del Carmen Terán Moldes. QSAR study of N^{6}-(substituted-phenylcarbamoyl) adenosine-5’-uronamides as agonist for A_{1} adenosine receptors. Bulletin of Mathematical Biology, 66 (4):907–920, July 2004. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/content/
pdf/10.1016/j.bulm.2004.04.002.pdf.

Puebla:2004:BRP

[1877] Hector Puebla. Book review: Pathway analysis and optimization in metabolic engineering by N. V. Torres and E. O. Voit, 2002, Cambridge University Press, Cambridge, UK, $95.00 (hardback), ISBN 0-521-80038-2. Bulletin of Mathematical Biology, 66(4):921–923, July 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1016/j.
bulm.2004.05.003.pdf.

Nebel:2004:IBM

[1878] Markus E. Nebel. Investigation of the Bernoulli model for RNA secondary structures. Bulletin of Mathematical Biology, 66(5):925–964, September 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1016/j.bulm.2003.08.015.pdf.

He:2004:GCA

[1879] Matthew He. Genetic code, attributive mappings and stochastic matrices. Bulletin of Mathematical Biology, 66(5):965–973, September 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1016/j.
bulm.2003.10.002.pdf.

Potapov:2004:CCE

[1880] A. B. Potapov and M. A. Lewis. Climate and competition: The effect of moving range boundaries on habitat invasibility. Bulletin of Mathematical Biology, 66(5):975–1008, September 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/content/pdf/10.1016/j.bulm.2003.10.010.pdf.

Bortz:2004:SAN

[1881] D. M. Bortz and P. W. Nelson. Sensitivity analysis of a nonlinear lumped parameter model of HIV infection dynamics. Bulletin of Mathematical Biology, 66(5):1009–1026, September 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/content/pdf/10.1016/j.bulm.2003.10.011.pdf.

Filipe:2004:AMP

[1882] J. A. N. Filipe, M. M. Maule, and C. A. Gilligan. On ‘Analytical models for the patchy spread of plant disease’. Bulletin of Mathematical Biology, 66(5):1027–1037, September 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1016/j.bulm.2003.11.001.pdf. See [1589].

Araujo:2004:HSS

[1883] R. P. Araujo and D. L. S. McElwain. A history of the study of solid tumour growth: The contribution of mathematical modelling. Bulletin of Mathematical Biology, 66(5):1039–1091, September 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1016/j.
bulm.2003.11.002.pdf.

Bajaria:2004:PDR

[1884] Seema H. Bajaria, Glenn Webb, and Denise E. Kirschner. Predicting differential responses to structured treatment interruptions during HAART. Bulletin of Mathematical Biology, 66(5):1093–1118, September 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1016/j.
bulm.2003.11.003.pdf.

Stephanou:2004:MMD

[1885] A. Stéphanou, M. A. J. Chaplain, and P. Tracqui. A mathematical model for the dynamics of large membrane deformations of isolated fibroblasts. Bulletin of Mathematical Biology, 66(5):1119–1154, September 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1016/j.
bulm.2003.11.004.pdf.

Meier:2004:CMF

[1886] Jens Meier, Martin Kleen, and Konrad Messmer. A computer model of fractal myocardial perfusion heterogeneity to elucidate mechanisms of changes in critical coronary stenosis and hypotension. Bulletin of Mathematical Biology, 66(5):1155–1171, September 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/content/pdf/10.1016/j.bulm.2003.11.005.pdf.

Rosales:2004:MHM

[1887] Rafael A. Rosales. MCMC for hidden Markov models incorporating aggregation of states and filtering. Bulletin of Mathematical Biology, 66(5):1173–1199, September 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1016/j.bulm.2003.12.001.pdf.

Weinstein:2004:MEC

[1888] Alan M. Weinstein. Modeling epithelial cell homeostasis: Assessing recovery and control mechanisms. Bulletin of Mathematical Biology, 66(5):1201–1240, September 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1016/j.bulm.2003.12.002.pdf.

Yang:2004:GSI

[1889] Chi Ming Yang. On the 28-gon symmetry inherent in the genetic code intertwined with aminoacyl-tRNA synthetases — the Lucas series. Bulletin of Mathematical Biology, 66(5):1241–1257, September 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1016/j.
bulm.2003.12.006.pdf.

Smith:2004:DEP

[1890] R. J. Smith and L. M. Wahl. Distinct effects of protease and reverse transcriptase inhibition in an immunological model of HIV-1 infection with impulsive drug effects. Bulletin of Mathematical Biology, 66(5):1259–1283, September 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1016/j.bulm.2003.12.004.pdf.

Diaz:2004:SSF

[1891] Humberto González Díaz, Iyusmila Bastida, Nilo Castañedo, Oslay Nasco, Ervelio Olazabal, Alcidez Morales, Hector S. Serrano, and Ronal Ramos De Armas. Simple stochastic fingerprints towards mathematical modelling in biology and medicine. 1. The treatment of coccidiosis. Bulletin of Mathematical Biology, 66(5):1285–1311, September 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1016/j.
bulm.2003.12.003.pdf.

Bertram:2004:CBP

[1892] Richard Bertram and Arthur Sherman. A calcium-based phantom bursting model for pancreatic islets. Bulletin of Mathematical Biology, 66(5):1313–1344, September 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1016/j.bulm.2003.12.005.pdf.

vandenBerg:2004:ECC

[1893] Hugo Antonius van den Berg and Yuri Nikolaevich Kiselëv. Expansion and contraction of the cytotoxic T lymphocyte response — an optimal control approach. Bulletin of Mathematical Biology, 66(5):1345–1369, September 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1016/j.bulm.2004.01.001.pdf.

Aita:2004:TIE

[1894] Takuyo Aita, Shunichi Morinaga, and Yuzuru Husimi. Thermodynamical interpretation of evolutionary dynamics on a fitness landscape in a evolution reactor, I. Bulletin of Mathematical Biology, 66(5):1371–1403, September 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1016/j.bulm.2004.01.004.pdf.

He:2004:GCH

[1895] Matthew X. He, Sergei V. Petoukhov, and Paolo E. Ricci. Genetic code, Hamming distance and stochastic matrices. Bulletin of Mathematical Biology, 66(5):1405–1421, September 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/content/pdf/10.1016/j.bulm.2004.01.002.pdf.

Shu:2004:PAD

[1896] Jian-Jun Shu and Li Shan Ouw. Pairwise alignment of the DNA sequence using hypercomplex number representation. Bulletin of Mathematical Biology, 66(5):1423–1438, September 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/content/pdf/10.1016/j.bulm.2004.01.005.pdf.

Gesztelyi:2004:CEC

[1897] Rudolf Gesztelyi, Judit Zsuga, Béla Juhász, Péter Dér, Miklós Vecsernyés, and A. József Szentmiklósi. Concentration estimation via curve fitting: Quantification of negative inotropic agents by using a simple mathematical method in guinea pig atria. Bulletin of Mathematical Biology, 66(5):1439–1453, September 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1016/j.bulm.2004.03.001.pdf.

Jockusch:2004:LE

[1898] Harald Jockusch and Andreas Dress. Letter to the Editor. Bulletin of Mathematical Biology, 66(5):1455, September 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.

springer.com/content/pdf/10.1016/j.bulm.2004.06.002.pdf.

Solis:2004:BRE

[1899] Francisco Solís. Book review: Essential Mathematical Biology by Nicholas Ferris Britton, Springer Verlag, 370 pages, $34.95, ISBN 1-85233-536-X. Bulletin of Mathematical Biology, 66(5):1457–1458, September 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL https://link.springer.com/content/pdf/
10.1016/j.bulm.2004.05.001.pdf.

Federico:2004:BRM

[1900] Paula Federico and Suzanne Lenhart. Book review: The mathematics of marriage: dynamic linear models by John M. Gottman, James D. Murray, Catherine C. Swanson, Rebecca Tyson, and Kristin R. Swanson, 2002. Bradford Book, Massachusetts Institute of Technology Press, pp. 403, $42.95, ISBN 0-262-07226-2. Bulletin of Mathematical Biology, 66(5):1459–1461, September 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/content/pdf/10.1016/j.bulm.2004.05.002.pdf.

Pitman:2004:FMD

[1901] E. Bruce Pitman, Roman M. Zaritski, Kevin J. Kesseler, Leon C. Moore, and Harold E. Layton. Feedback-mediated dynamics in two coupled nephrons. Bulletin of Mathematical Biology, 66(6):1463–1492, November 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1016/j.
bulm.2004.01.006.pdf.

Warrender:2004:HPI

[1902] Christina Warrender, Stephanie Forrest, and Lee Segel. Homeostasis of peripheral immune effectors. Bulletin of Mathematical Biology, 66(6):1493–1514, November 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1016/j.bulm.2004.02.003.pdf.

Carrero:2004:CFR

[1903] G. Carrero, E. Crawford, M. J. Hendzel, and G. de Vries. Characterizing fluorescence recovery curves for nuclear proteins undergoing binding events. Bulletin of Mathematical Biology, 66(6):1515–1545, November 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1016/j.
bulm.2004.02.005.pdf.

Jones:2004:ETE

[1904] Laura E. Jones and Stephen P. Ellner. Evolutionary tradeoff and equilibrium in an aquatic predator-prey system. Bulletin of Mathematical Biology, 66(6):1547–1573, November 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/content/pdf/10.1016/j.bulm.2004.02.006.pdf.

Austin:2004:SWD

[1905] Brian Austin, Peter Dawyndt, Mats Gyllenberg, Timo Koski, Tatu Lund, Jean Swings, and Fabiano L. Thompson. Sliding window discretization: a new method for multiple band matching of bacterial genotyping fingerprints. Bulletin of Mathematical Biology, 66(6):1575–1596, November 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1016/j.
bulm.2004.02.004.pdf.

Jafelice:2004:FMS

[1906] Rosana Motta Jafelice, Laécio Carvalho de Barros, Rodney Carlos Bassanezi, and Fernando Gomide. Fuzzy modeling in symptomatic HIV virus infected population. Bulletin of Mathematical Biology, 66(6):1597–1620, November 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1016/j.bulm.2004.03.002.pdf.

Taylor:2004:EGD

[1907] Christine Taylor, Drew Fudenberg, Akira Sasaki, and Martin A. Nowak. Evolutionary game dynamics in finite populations. Bulletin of Mathematical Biology, 66(6):1621–1644, November 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/content/pdf/10.1016/j.bulm.2004.03.004.pdf.

Broom:2004:RUE

[1908] Mark Broom, Roger M. Luther, and Graeme D. Ruxton. Resistance is useless? — Extensions to the game theory of kleptoparasitism. Bulletin of Mathematical Biology, 66(6):1645–1658, November 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1016/j.
bulm.2004.03.009.pdf.

Ghosh:2004:SPD

[1909] Mini Ghosh and Andrea Pugliese. Seasonal population dynamics of ticks, and its influence on infection transmission: a semi-discrete approach. Bulletin of Mathematical Biology, 66(6):1659–1684, November 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1016/j.
bulm.2004.03.007.pdf.

Wikan:2004:ANS

[1910] Arild Wikan and Arne Eide. An analysis of a nonlinear stagestructured cannibalism model with application to the Northeast Arctic cod stock. Bulletin of Mathematical Biology, 66(6):1685–1704, November 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1016/j.
bulm.2004.03.005.pdf.

Quine:2004:MAP

[1911] J. R. Quine, Timothy A. Cross, Michael S. Chapman, and Richard Bertram. Mathematical aspects of protein structure determination with NMR orientational restraints. Bulletin of Mathematical Biology, 66(6): 1705–1730, November 2004. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/content/
pdf/10.1016/j.bulm.2004.03.006.pdf.

Gurney:2004:RAH

[1912] William S. C. Gurney and Roger M. Nisbet. Resource allocation, hyperphagia and compensatory growth. Bulletin of Mathematical Biology, 66(6):1731–1753, November 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1016/j.bulm.2004.03.008.pdf.

Willson:2004:CRS

[1913] Stephen J. Willson. Constructing rooted supertrees using distances. Bulletin of Mathematical Biology, 66(6):1755–1783, November 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1016/j.
bulm.2004.04.006.pdf.

Plank:2004:LNL

[1914] M. J. Plank and B. D. Sleeman. Lattice and non-lattice models of tumour angiogenesis. Bulletin of Mathematical Biology, 66(6):1785–1819, November 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1016/j.bulm.2004.04.001.pdf.

Gilbert:2004:CTM

[1915] Estella Gilbert, James A. Powell, Jesse A. Logan, and Barbara J. Bentz. Comparison of three models predicting developmental milestones given environmental and individual variation. Bulletin of Mathematical Biology, 66(6):1821–1850, November 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1016/j.bulm.2004.04.003.pdf.

Ambrosi:2004:CDC

[1916] D. Ambrosi, A. Gamba, and G. Serini. Cell directional and chemotaxis in vascular morphogenesis. Bulletin of Mathematical Biology, 66(6):1851– 1873, November 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1016/j.bulm.2004.04.004.pdf.

Wang:2004:CRP

[1917] Gang Wang, Jianli Yuan, Xizhi Wang, Sa Xiao, and Wenbing Huang. Competitive regulation of plant allometry and a generalized model for the plant self-thinning process. Bulletin of Mathematical Biology, 66(6): 1875–1885, November 2004. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/content/
pdf/10.1016/j.bulm.2004.04.005.pdf.

Hinch:2004:SCW

[1918] R. Hinch. Stability of cardiac waves. Bulletin of Mathematical Biology, 66(6):1887–1908, November 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1016/j.bulm.2004.05.004.pdf.

Hillary:2004:SCP

[1919] R. M. Hillary and M. A. Bees. Synchrony & chaos in patchy ecosystems. Bulletin of Mathematical Biology, 66(6):1909–1931, November 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1016/j.
bulm.2004.05.006.pdf.

Fister:2004:BRE

[1920] K. Renee Fister. Book review: Elizabeth S. Allman and John A. Rhodes, Mathematical Models in Biology, Cambridge University Press, Cambridge, UK (2004). ISBN 0-521-81980-6 (hardback) or 0-521-52586-1 (paperback) $40.00 (paperback). Bulletin of Mathematical Biology, 66(6):1933–1934, November 2004. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1016/j.bulm.2004.07.002.pdf.

Sanchez:2005:GCBa

[1921] Robersy Sánchez, Eberto Morgado, and Ricardo Grau. A genetic code Boolean structure. I. The meaning of Boolean deductions. Bulletin of Mathematical Biology, 67(1):1–14, January 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/content/pdf/10.1016/j.bulm.2004.05.005.pdf.

Yaish:2005:MDO

[1922] Barak Yaish and Ramit Mehr. Models for the dynamics and order of immunoglobulin isotype switching. Bulletin of Mathematical Biology, 67(1): 15–32, January 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1016/j.bulm.2004.05.007.pdf.

Barnes:2005:PHM

[1923] Belinda Barnes and Harvinder Sidhu. Plant-herbivore models, where more grass means fewer grazers. Bulletin of Mathematical Biology, 67(1): 33–55, January 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1016/j.bulm.2004.06.001.pdf.

Kooijman:2005:HGA

[1924] S. A. L. M. Kooijman and L. A. Segel. How growth affects the fate of cellular metabolites. Bulletin of Mathematical Biology, 67(1):57–77, January 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1016/j.
bulm.2004.06.003.pdf.

Ribba:2005:MMD

[1925] B. Ribba, K. Marron, Z. Agur, T. Alarcón, and P. K. Maini. A mathematical model of Doxorubicin treatment efficacy for non-Hodgkin’s lymphoma: Investigation of the current protocol through theoretical modelling results. Bulletin of Mathematical Biology, 67(1):79–99, January 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1016/j.
bulm.2004.06.007.pdf.

Gomez-Acevedo:2005:BBM

[1926] Horacio Gómez-Acevedo and Michael Y. Li. Backward bifurcation in a model for HTLV-I infection of CD4^{+} T cells. Bulletin of Mathematical Biology, 67(1):101–114, January 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1016/j.bulm.2004.06.004.pdf.

Tang:2005:IPM

[1927] Sanyi Tang, Yanni Xiao, Lansun Chen, and Robert A. Cheke. Integrated pest management models and their dynamical behaviour. Bulletin of Mathematical Biology, 67(1):115–135, January 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/content/pdf/10.1016/j.bulm.2004.06.005.pdf.

Flores:2005:SBF

[1928] Heather Flores, Edgar Lobaton, Stefan Méndez-Diez, Svetlana Tlupova, and Ricardo Cortez. A study of bacterial flagellar bundling. Bulletin of Mathematical Biology, 67(1):137–168, January 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/content/pdf/10.1016/j.bulm.2004.06.006.pdf.

Kschischo:2005:TAS

[1929] Maik Kschischo, Michael Lässig, and Yi-Kuo Yuc. Toward an accurate statistics of gapped alignments. Bulletin of Mathematical Biology, 67(1):169–191, January 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1016/j.bulm.2004.07.001.pdf.

Hernandez:2005:BRM

[1930] Gerardo Hernández. Book review: Michael Bulmer, Francis Galton: Pioneer of Heredity and Biometry, Johns Hopkins University Press (2003), ISBN 0-8018-7403-3, 352 pages, $45.00. Bulletin of Mathematical Biology, 67(1):193–194, January 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1016/j.bulm.2004.08.008.pdf.

Ambrosi:2005:CDP

[1931] D. Ambrosi, A. Gamba, and G. Serini. Cell directional persistence and chemotaxis in vascular morphogenesis. Bulletin of Mathematical Biology, 67(1):195, January 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1016/j.bulm.2005.01.001.pdf.

Anonymous:2005:O

[1932] Anonymous. Obituary. Bulletin of Mathematical Biology, 67(2):iii, March 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/content/pdf/10.1016/
S0092-8240%2805%2900027-3.pdf.

Hu:2005:WDI

[1933] Kai Hu. Writhe of DNA induced by a terminal twist. Bulletin of Mathematical Biology, 67(2):197–209, March 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.

springer.com/content/pdf/10.1016/j.bulm.2004.05.008.pdf. See

erratum [1976].

Zheng:2005:NST

[1934] X. Zheng, S. M. Wise, and V. Cristini. Nonlinear simulation of tumor necrosis, neo-vascularization and tissue invasion via an adaptive finite-element/level-set method. Bulletin of Mathematical Biology, 67 (2):211–259, March 2005. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/content/
pdf/10.1016/j.bulm.2004.08.001.pdf.

Willson:2005:MEU

[1935] Stephen J. Willson. Minimum evolution using ordinary least-squares is less robust than neighbor-joining. Bulletin of Mathematical Biology, 67 (2):261–279, March 2005. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/content/
pdf/10.1016/j.bulm.2004.07.007.pdf.

Mistro:2005:AHB

[1936] Diomar Cristina Mistro, Luiz Alberto Díaz Rodrigues, and Wilson Castro Ferreira, Jr. The Africanized honey bee dispersal: a mathematical zoom. Bulletin of Mathematical Biology, 67(2):281–312, March 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/content/pdf/10.1016/j.
bulm.2004.07.006.pdf. See erratum [1974].

Sun:2005:DMG

[1937] Shuyu Sun, Mary F. Wheeler, Mandri Obeyesekere, and C. Patrick, Jr. A deterministic model of growth factor-induced angiogenesis. Bulletin of Mathematical Biology, 67(2):313–337, March 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.

springer.com/content/pdf/10.1016/j.bulm.2004.07.004.pdf.

Wang:2005:MPO

[1938] Ruiqi Wang, Zhujun Jing, and Luonan Chen. Modelling periodic oscillation in gene regulatory networks by cyclic feedback systems. Bulletin of Mathematical Biology, 67(2):339–367, March 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.

springer.com/content/pdf/10.1016/j.bulm.2004.07.005.pdf.

Woodroffe:2005:MCS

[1939] P. J. Woodroffe, J. R. King, C. L. Varley, and J. Southgate. Modelling cell signalling and differentiation in the urothelium. Bulletin of Mathematical Biology, 67(2):369–389, March 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.

springer.com/content/pdf/10.1016/j.bulm.2004.08.006.pdf.

Sanchez-Garduno:2005:BRJ

[1940] Faustino Sánchez-Garduño. Book review: J. D. Murray, Mathematical biology, Springer Verlag (2003). ISBN 0-387-95228-4, 736 pages $79.95. Bulletin of Mathematical Biology, 67(2):391–392, March 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/content/pdf/10.1016/j.
bulm.2004.09.001.pdf.

Nguyen:2005:SAN

[1941] Vien Nguyen, Roy Mathias, and Gregory D. Smith. A stochastic automata network descriptor for Markov chain models of instantaneously coupled intracellular ca^{2+} channels. Bulletin of Mathematical Biology, 67(3):393–432, May 2005. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL https://link.springer.com/content/
pdf/10.1016/j.bulm.2004.08.010.pdf.

Diaz:2005:IIC

[1942] José Díaz and Gustavo Martínez-Mekler. Interaction of the IP_{3}-Ca^{2+} and MAPK signaling systems in the Xenopus blastomere: A possible frequency encoding mechanism for the control of the Xbra gene expression. Bulletin of Mathematical Biology, 67(3):433–465, May 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/content/pdf/10.1016/j.
bulm.2004.08.002.pdf.

Awerbuch-Friedlander:2005:RSD

[1943] T. Awerbuch-Friedlander, R. Levins, and M. Predescu. The role of seasonality in the dynamics of deer tick populations. Bulletin of Mathematical Biology, 67(3):467–486, May 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/content/pdf/10.1016/j.bulm.2004.08.003.pdf.

Boushaba:2005:DEE

[1944] Khalid Boushaba and Mercedes Pascual. Dynamics of the ‘echo’ effect in a phytoplankton system with nitrogen fixation. Bulletin of Mathematical Biology, 67(3):487–507, May 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.springer.

com/content/pdf/10.1016/j.bulm.2004.08.004.pdf.

Takahashi:2005:MMA

[1945] Lucy Tiemi Takahashi, Norberto Anibal Maidana, Wilson Castro Ferreira, Jr., Petronio Pulino, and Hyun Mo Yang. Mathematical models for the Aedes aegypti dispersal dynamics: Travelling waves by wing and wind. Bulletin of Mathematical Biology, 67(3):509–528, May 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/content/pdf/10.1016/j.
bulm.2004.08.005.pdf.

Bardos:2005:PGM

[1946] D. C. Bardos. Probabilistic Gompertz model of irreversible growth. Bulletin of Mathematical Biology, 67(3):529–545, May 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https:

//link.springer.com/content/pdf/10.1016/j.bulm.2004.08.009.
pdf.

Ditlevsen:2005:SVD

[1947] Susanne Ditlevsen and Andrea de Gaetano. Stochastic vs. deterministic uptake of dodecanedioic acid by isolated rat livers. Bulletin of Mathematical Biology, 67(3):547–561, May 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.

springer.com/content/pdf/10.1016/j.bulm.2004.09.005.pdf.

Gaffney:2005:MMA

[1948] E. A. Gaffney. The mathematical modelling of adjuvant chemotherapy scheduling: Incorporating the effects of protocol rest phases and pharmacokinetics. Bulletin of Mathematical Biology, 67(3):563–611, May 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/content/pdf/10.1016/j.
bulm.2004.09.002.pdf.

Aita:2005:TIE

[1949] Takuyo Aita, Shunichi Morinaga, and Yuzuru Husimi. Thermodynamical interpretation of evolutionary dynamics on a fitness landscape in an evolution reactor, II. Bulletin of Mathematical Biology, 67(3):613–635, May 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL https://link.springer.com/content/pdf/10.1016/j.
bulm.2004.09.004.pdf.

Petrovskii:2005:RBI

[1950] Sergei Petrovskii, Andrew Morozov, and Bai-Lian Li. Regimes of biological invasion in a predator-prey system with the Allee effect. Bulletin of Mathematical Biology, 67(3):637–661, May 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.

springer.com/content/pdf/10.1016/j.bulm.2004.09.003.pdf.

Brannstrom:2005:CML

[1951] Å. Brännström and D. J. T. Sumpter. Coupled map lattice approximations for spatially explicit individual-based models of ecology. Bulletin of Mathematical Biology, 67(4):663–682, July 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/content/pdf/10.1016/j.bulm.2004.09.006.pdf.

Tien:2005:MMS

[1952] Joseph H. Tien, William D. Hazelton, Rachel Sparks, and Cornelia M. Ulrich. A Michaelis–Menten-style model for the autocatalytic enzyme prostaglandin h synthase. Bulletin of Mathematical Biology, 67(4):683– 700, July 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1016/j.bulm.2004.09.007.pdf.

Komarova:2005:NWD

[1953] Natalia L. Komarova and Avy Soffera. Nonlinear waves in doublestranded DNA. Bulletin of Mathematical Biology, 67(4):701–718, July 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1016/j.
bulm.2004.09.008.pdf.

Bax:2005:CMC

[1954] Eric Bax. Checkpoint method for choice recovery in dynamic programming. Bulletin of Mathematical Biology, 67(4):719–736, July 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1016/j.
bulm.2004.10.002.pdf.

Menconi:2005:SGI

[1955] Giulia Menconi. Sublinear growth of information in DNA sequences. Bulletin of Mathematical Biology, 67(4):737–759, July 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/content/pdf/10.1016/j.bulm.2004.10.005.
pdf.

Sanchez:2005:DDR

[1956] María S. Sánchez, Robert M. Grant, Travis C. Porco, Kimber L. Gross, and Wayne M. Getz. A decrease in drug resistance levels of the HIV epidemic can be bad news. Bulletin of Mathematical Biology, 67(4): 761–782, July 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1016/j.bulm.2004.10.001.pdf.

Smith:2005:DRI

[1957] R. J. Smith and L. M. Wahl. Drug resistance in an immunological model of HIV-1 infection with impulsive drug effects. Bulletin of Mathematical Biology, 67(4):783–813, July 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1016/j.bulm.2004.10.004.pdf.

Basse:2005:MFC

[1958] Britta Basse, Bruce C. Baguley, Elaine S. Marshall, Graeme C. Wake, and David J. N. Wall. Modelling the flow of cytometric data obtained from unperturbed human tumour cell lines: Parameter fitting and comparison. Bulletin of Mathematical Biology, 67(4):815–830, July 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1016/j.
bulm.2004.10.003.pdf. See erratum [1975].

Cogan:2005:MPR

[1959] N. G. Cogan, Ricardo Cortez, and Lisa Fauci. Modeling physiological resistance in bacterial biofilms. Bulletin of Mathematical Biology, 67(4): 831–853, July 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1016/j.bulm.2004.11.001.pdf.

Krishnarajah:2005:NMC

[1960] Isthrinayagy Krishnarajah, Alex Cook, Glenn Marion, and Gavin Gibson. Novel moment closure approximations in stochastic epidemics. Bulletin of Mathematical Biology, 67(4):855–873, July 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/content/pdf/10.1016/j.bulm.2004.11.002.
pdf.

Fishman:2005:FAB

[1961] Michael A. Fishman and Lewi Stone. Fertility assurance and breeding synchrony. Bulletin of Mathematical Biology, 67(4):875–883, July 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1016/j.
bulm.2004.11.004.pdf.

DiMascio:2005:DIE

[1962] Michele Di Mascio, Jerome K. Percus, Ora E. Percus, Martin Markowitz, David D. Ho, and Alan S. Perelson. Duration of an intermittent episode of viremia. Bulletin of Mathematical Biology, 67(4):885–900, July 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1016/j.
bulm.2004.11.003.pdf.

Gadgil:2005:SAF

[1963] Chetan Gadgil, Chang Hyeong Lee, and Hans G. Othmer. A stochastic analysis of first-order reaction networks. Bulletin of Mathematical Biology, 67(5):901–946, September 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1016/j.bulm.2004.09.009.pdf.

Hinch:2005:NAR

[1964] Robert Hinch. A note on the asymptotic reduction of the Hodgkin– Huxley equations for nerve impulses. Bulletin of Mathematical Biology, 67(5):947–955, September 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1016/j.bulm.2004.11.007.pdf.

Grafton:2005:MRE

[1965] R. Quentin Grafton, Tom Kompas, and David Lindenmayer. Marine reserves with ecological uncertainty. Bulletin of Mathematical Biology, 67(5):957–971, September 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1016/j.bulm.2004.11.006.pdf.

Twarock:2005:MMT

[1966] R. Twarock. Mathematical models for tubular structures in the family of Papovaviridae. Bulletin of Mathematical Biology, 67(5):973–987, September 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1016/j.bulm.2004.11.005.pdf.

Stirling:2005:MOU

[1967] J. R. Stirling, M. S. Zakynthinaki, and B. Saltin. A model of oxygen uptake kinetics in response to exercise: Including a means of calculating oxygen demand/deficit/debt. Bulletin of Mathematical Biology, 67(5):989–1015, September 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1016/j.bulm.2004.12.005.pdf.

Sanchez:2005:GCBb

[1968] Robersy Sanchez and Ricardo Grau. A genetic code Boolean structure. II. The genetic information system as a Boolean information system. Bulletin of Mathematical Biology, 67(5):1017–1029, September 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1016/j.
bulm.2004.12.004.pdf.

Miklos:2005:MBD

[1969] István Miklós, Irmtraud M. Meyer, and Borbála Nagy. Moments of the Boltzmann distribution for RNA secondary structures. Bulletin of Mathematical Biology, 67(5):1031–1047, September 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/content/pdf/10.1016/j.bulm.2004.12.003.pdf.

Stewart-Cox:2005:PLM

[1970] J. A. Stewart-Cox, N. F. Brittona, and M. Mogie. Pollen limitation or mate search need not induce an Allee effect. Bulletin of Mathematical Biology, 67(5):1049–1079, September 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/content/pdf/10.1016/j.bulm.2004.12.001.pdf.

Fan:2005:CPB

[1971] Meng Fan, Yang Kuang, and Zhilan Feng. Cats protecting birds revisited. Bulletin of Mathematical Biology, 67(5):1081–1106, September 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1016/j.
bulm.2004.12.002.pdf.

Bowman:2005:MMA

[1972] C. Bowman, A. B. Gumel, P. van den Driessche, J. Wu, and H. Zhu. A mathematical model for assessing control strategies against West Nile virus. Bulletin of Mathematical Biology, 67(5):1107–1133, September 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1016/j.
bulm.2005.01.002.pdf.

Giuggioli:2005:DHR

[1973] L. Giuggioli, G. Abramson, V. M. Kenkre, G. Suzán, E. Marcé, and T. L. Yates. Diffusion and home range parameters from rodent population measurements in Panama. Bulletin of Mathematical Biology, 67(5):1135–1149, September 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1016/j.bulm.2005.01.003.pdf.

Mistro:2005:EAH

[1974] Diomar Cristina Mistro, Luiz Alberto Díaz Rodrigues, and Wilson Castro Ferreira, Jr. Erratum: The Africanized honey bee dispersal: a mathematical zoom. Bulletin of Mathematical Biology, 67(5):1151, September 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1016/j.
bulm.2005.05.001.pdf. See [1936].

Basse:2005:EMF

[1975] Britta Basse, Bruce C. Baguley, Elaine S. Marshall, Graeme C. Wake, and David J. N. Wall. Erratum: Modelling the flow cytometric data obtained from unperturbed human tumour cell lines: Parameter fitting and comparison. Bulletin of Mathematical Biology, 67(5):1153, September 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1016/j.
bulm.2005.06.001.pdf. See [1958].

Hu:2005:EWD

[1976] Kai Hu. Erratum: Writhe of DNA induced by a terminal twist. Bulletin of Mathematical Biology, 67(5):1155, September 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/content/pdf/10.1016/j.bulm.2005.06.002.

pdf. See [1933].

Cruz-Pacheco:2005:MDW

[1977] Gustavo Cruz-Pacheco, Lourdes Esteva, Juan Antonio Montaõ-Hirose, and Cristobal Vargas. Modelling the dynamics of West Nile Virus. Bulletin of Mathematical Biology, 67(6):1157–1172, November 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1016/j.
bulm.2004.11.008.pdf.

Percus:2005:SPE

[1978] Jerome K. Percus. Small population effects in stochastic population dynamics. Bulletin of Mathematical Biology, 67(6):1173–1194, November 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1016/j.
bulm.2005.01.005.pdf.

Kaminski:2005:SST

[1979] Dominik Kamiński, Jacek Miekisz, and Marcin Zaborowski. Stochastic stability in three-player games. Bulletin of Mathematical Biology, 67(6):1195–1205, November 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1016/j.bulm.2005.01.004.pdf.

Xu:2005:RSM

[1980] Dashun Xu, Jason Curtis, Zhilan Feng, and Dennis J. Minchella. On the role of schistosome mating structure in the maintenance of drug resistant strains. Bulletin of Mathematical Biology, 67(6):1207–1226, November 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1016/j.
bulm.2005.01.007.pdf.

Jonesa:2005:OIC

[1981] Laura E. Jonesa and Alan S. Perelson. Opportunistic infection as a cause of transient viremia in chronically infected HIV patients under treatment with HAART. Bulletin of Mathematical Biology, 67(6):1227– 1251, November 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1016/j.bulm.2005.01.006.pdf.

Koenigsberger:2005:EPE

[1982] Michèle Koenigsberger, Roger Sauser, and Jean-Jacques Meister. Emergent properties of electrically coupled smooth muscle cells. Bulletin of Mathematical Biology, 67(6):1253–1272, November 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/content/pdf/10.1016/j.bulm.2005.02.001.
pdf.

Stewart-Cox:2005:SMC

[1983] J. A. Stewart-Cox, N. F. Britton, and M. Mogie. Space mediates coexistence of females and hermaphrodites. Bulletin of Mathematical Biology, 67(6):1273–1302, November 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

content/pdf/10.1016/j.bulm.2005.02.003.pdf.

Reyment:2005:SAM

[1984] Richard A. Reyment. The statistical analysis of multivariate serological frequency data. Bulletin of Mathematical Biology, 67(6):1303–1313, November 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1016/j.bulm.2005.02.002.pdf.

Snedecor:2005:DTM

[1985] Sonya J. Snedecor. Dynamic treatment model to examine the association between phenotypic drug resistance and duration of HIV viral suppression. Bulletin of Mathematical Biology, 67(6):1315–1332, November 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1016/j.
bulm.2005.02.004.pdf.

He:2005:ALF

[1986] Ji-Huan He. The allometry of leaf form in early plant ontogeny. Bulletin of Mathematical Biology, 67(6):1333–1337, November 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http:

//link.springer.com/content/pdf/10.1016/j.bulm.2005.02.005.
pdf.

Yoshida:2005:SCV

[1987] Hiroshi Yoshida, Takashi Yokomori, and Akira Suyama. A simple classification of the volvocine algae by formal languages. Bulletin of Mathematical Biology, 67(6):1339–1354, November 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/content/pdf/10.1016/j.bulm.2005.03.001.pdf.

Goodarzi:2005:IIT

[1988] Hani Goodarzi, Hamed Shateri Najafabadi, Hamed Ahmadi Nejad, and Noorossadat Torabi. The impact of including tRNA content on the optimality of the genetic code. Bulletin of Mathematical Biology, 67(6):1355– 1368, November 2005. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/content/pdf/10.
1016/j.bulm.2005.03.002.pdf.

Maini:2006:E

[1989] Philip K. Maini. Editorial. Bulletin of Mathematical Biology, 68 (1):1, January 2006. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-006-9081-0;   http://link.springer.com/content/
pdf/10.1007/s11538-006-9081-0.pdf.

Lewis:2006:TWS

[1990] Mark Lewis, Joanna Rencławowicz, and P. van den Driessche. Traveling waves and spread rates for a West Nile Virus model. Bulletin of Mathematical Biology, 68(1):3–23, January 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-005-9018-z;
 http://link.springer.com/content/pdf/10.1007/s11538-005-9018-
z.pdf.

Cai:2006:MDG

[1991] Anna Q. Cai, Kerry A. Landman, and Barry D. Hughes. Modelling directional guidance and motility regulation in cell migration. Bulletin of Mathematical Biology, 68(1):25–52, January 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
005-9028-x;   http://link.springer.com/content/pdf/10.1007/
s11538-005-9028-x.pdf.

Fitt:2006:FMH

[1992] A. D. Fitt and G. Gonzalez. Fluid mechanics of the human eye: Aqueous humour flow in the anterior chamber. Bulletin of Mathematical Biology, 68(1):53–71, January 2006. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-005-9015-2;   http://link.springer.com/content/
pdf/10.1007/s11538-005-9015-2.pdf.

Oster:2006:DMO

[1993] Andrew M. Oster and Paul C. Bressloff. A developmental model of ocular dominance column formation on a growing cortex. Bulletin of Mathematical Biology, 68(1):73–98, January 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-005-9055-7;
 http://link.springer.com/content/pdf/10.1007/s11538-005-9055-
7.pdf.

Gaffney:2006:GET

[1994] E. A. Gaffney and N. A. M. Monk. Gene expression time delays and Turing pattern formation systems. Bulletin of Mathematical Biology, 68 (1):99–130, January 2006. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-006-9066-z;   http://link.springer.com/content/
pdf/10.1007/s11538-006-9066-z.pdf.

Atzberger:2006:BDM

[1995] Paul J. Atzberger and Charles S. Peskin. A Brownian dynamics model of kinesin in three dimensions incorporating the force–extension profile of the coiled–coil cargo tether. Bulletin of Mathematical Biology, 68 (1):131–160, January 2006. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-005-9003-6;   http://link.springer.com/content/
pdf/10.1007/s11538-005-9003-6.pdf.

Wilson:2006:TIS

[1996] D. P. Wilson, P. Timms, D. L. S. Mcelwain, and P. M. Bavoil. Type III secretion, contact-dependent model for the intracellular development of Chlamydia. Bulletin of Mathematical Biology, 68(1):161–178, January 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
005-9024-1;   http://link.springer.com/content/pdf/10.1007/
s11538-005-9024-1.pdf.

Wendl:2006:OMC

[1997] Michael C. Wendl. Occupancy modeling of coverage distribution for whole genome shotgun DNA sequencing. Bulletin of Mathematical Biology, 68(1):179–196, January 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-005-9021-4;   http://link.springer.com/
content/pdf/10.1007/s11538-005-9021-4.pdf.

Waugh:2006:MDD

[1998] Helen V. Waugh and Jonathan A. Sherratt. Macrophage dynamics in diabetic wound dealing. Bulletin of Mathematical Biology, 68(1): 197–207, January 2006. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-005-9022-3;   http://link.springer.com/content/
pdf/10.1007/s11538-005-9022-3.pdf.

Xu:2006:RSM

[1999] Dashun Xu, Jason Curtis, Zhilan Feng, and Dennis J. Minchella. On the role of schistosome mating structure in the maintenance of drugresistant strains. Bulletin of Mathematical Biology, 68(2):209–229, February 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
005-9001-3;   http://link.springer.com/content/pdf/10.1007/
s11538-005-9001-3.pdf.

Mallet:2006:MMI

[2000] D. G. Mallet and G. J. Pettet. A mathematical model of integrinmediated haptotactic cell migration. Bulletin of Mathematical Biology, 68 (2):231–253, February 2006. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-005-9032-1;   http://link.springer.com/content/
pdf/10.1007/s11538-005-9032-1.pdf.

Castiglione:2006:OCM

[2001] Filippo Castiglione and Benedetto Piccoli. Optimal control in a model of dendritic cell transfection cancer immunotherapy. Bulletin of Mathematical Biology, 68(2):255–274, February 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-005-9014-3;
 http://link.springer.com/content/pdf/10.1007/s11538-005-9014-
3.pdf.

Vatov:2006:MBE

[2002] Larisa Vatov, Ziv Kizner, Eytan Ruppin, Sigal Meilin, Tamar Manor, and Avraham Mayevsky. Modeling brain energy metabolism and function: a multiparametric monitoring approach. Bulletin of Mathematical Biology, 68(2):275–291, February 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-005-9008-1;
 http://link.springer.com/content/pdf/10.1007/s11538-005-9008-
1.pdf.

Mukhopadhyay:2006:MRD

[2003] B. Mukhopadhyay and R. Bhattacharyya. Modeling the role of diffusion coefficients on Turing instability in a reaction–diffusion prey–predator system. Bulletin of Mathematical Biology, 68(2):293–313, February 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-

005-9007-2;   http://link.springer.com/content/pdf/10.1007/
s11538-005-9007-2.pdf.

Aittokallio:2006:MBA

[2004] T. Aittokallio, M. Gyllenberg, O. Polo, J. Toivonen, and A. Virkki. Model-based analysis of mechanisms responsible for sleep–induced carbon dioxide differences. Bulletin of Mathematical Biology, 68(2):315– 341, February 2006. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-005-9059-3;   http://link.springer.com/content/
pdf/10.1007/s11538-005-9059-3.pdf.

Santillan:2006:ASM

[2005] Moisés Santillán and Eduardo S. Zeronb. Analytical study of the multiplicity of regulatory mechanisms in the tryptophan operon. Bulletin of Mathematical Biology, 68(2):343–359, February 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
005-9025-0;   http://link.springer.com/content/pdf/10.1007/
s11538-005-9025-0.pdf.

Song:2006:SCR

[2006] Yun S. Song. A sufficient condition for reducing recursions in hidden Markov models. Bulletin of Mathematical Biology, 68(2):361–384, February 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
005-9045-9;   http://link.springer.com/content/pdf/10.1007/
s11538-005-9045-9.pdf.

Melkikh:2006:RMM

[2007] A. V. Melkikh and V. D. Seleznev. Requirements on models and models of active transport of ions in biomembranes. Bulletin of Mathematical Biology, 68(2):385–399, February 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-005-9035-y;
 http://link.springer.com/content/pdf/10.1007/s11538-005-9035-
y.pdf.

Reluga:2006:MSE

[2008] Timothy C. Reluga, Jan Medlock, and Alison P. Galvani. A model of spatial epidemic spread when individuals move within overlapping home ranges. Bulletin of Mathematical Biology, 68(2):401–416, February 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
005-9027-y;   http://link.springer.com/content/pdf/10.1007/
s11538-005-9027-y.pdf.

Ross:2006:SMM

[2009] J. V. Ross. Stochastic models for mainland-island metapopulations in static and dynamic landscapes. Bulletin of Mathematical Biology, 68(2): 417–449, February 2006. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-005-9043-y;   http://link.springer.com/content/
pdf/10.1007/s11538-005-9043-y.pdf.

Silva:2006:DDM

[2010] Jacques A. L. Silva and Flávia T. Giordani. Density-dependent migration and synchronism in metapopulations. Bulletin of Mathematical Biology, 68(2):451–466, February 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-005-9054-8;   http://link.springer.com/
content/pdf/10.1007/s11538-005-9054-8.pdf.

Henson:2006:IED

[2011] Shandelle M. Henson, James L. Hayward, and Smruti P. Damania. Identifying environmental determinants of diurnal distribution in marine birds and mammals. Bulletin of Mathematical Biology, 68(2): 467–482, February 2006. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-005-9009-0;   http://link.springer.com/content/
pdf/10.1007/s11538-005-9009-0.pdf.

Tasaki:2006:NLC

[2012] Ichiji Tasaki. A note on the local current associated with the rising phase of a propagating impulse in nonmyelinated nerve fibers. Bulletin of Mathematical Biology, 68(2):483–490, February 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
005-9012-5;   http://link.springer.com/content/pdf/10.1007/
s11538-005-9012-5.pdf.

Lewis:2006:CCD

[2013] Mark A. Lewis, Joanna Rencławowicz, P. van den Driessche, and Marjorie Wonham. A comparison of continuous and discrete-time West Nile Virus models. Bulletin of Mathematical Biology, 68(3):491–509, April 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
005-9039-7;   http://link.springer.com/content/pdf/10.1007/
s11538-005-9039-7.pdf.

Allen:2006:MMH

[2014] Linda J. S. Allen, Robert K. McCormack, and Colleen B. Jonsson. Mathematical models for hantavirus infection in rodents. Bulletin of Mathematical Biology, 68(3):511–524, April 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-005-9034-4;
 http://link.springer.com/content/pdf/10.1007/s11538-005-9034-
4.pdf.

Bacaer:2006:MHA

[2015] Nicolas Bacaër, Xamxinur Abdurahman, and Jianli Ye. Modeling the HIV/AIDS epidemic among injecting drug users and sex workers in Kunming, China. Bulletin of Mathematical Biology, 68(3):525–550, April 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
005-9051-y;   http://link.springer.com/content/pdf/10.1007/
s11538-005-9051-y.pdf.

Hsieh:2006:BRN

[2016] Ying-Hen Hsieh and Yun-Shih Wang. Basic reproduction number for HIV model incorporating commercial sex and behavior change. Bulletin of Mathematical Biology, 68(3):551–575, April 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
005-9050-z;   http://link.springer.com/content/pdf/10.1007/
s11538-005-9050-z.pdf.

Elbasha:2006:TAP

[2017] Elamin H. Elbasha and Abba B. Gumel. Theoretical assessment of public health impact of imperfect prophylactic HIV-1 vaccines with therapeutic benefits. Bulletin of Mathematical Biology, 68(3):577–614, April 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
005-9057-5;   http://link.springer.com/content/pdf/10.1007/
s11538-005-9057-5.pdf.

Korobeinikov:2006:LFG

[2018] Andrei Korobeinikov. Lyapunov functions and global stability for SIR and SIRS epidemiological models with non–linear transmission. Bulletin of Mathematical Biology, 68(3):615–626, April 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
005-9037-9;   http://link.springer.com/content/pdf/10.1007/
s11538-005-9037-9.pdf.

Edwards:2006:CEB

[2019] David A. Edwards. Convection effects in the BIAcore dextran layer: Surface reaction model. Bulletin of Mathematical Biology, 68(3):627–654, April 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
005-9023-2;   http://link.springer.com/content/pdf/10.1007/
s11538-005-9023-2.pdf.

Parvinen:2006:EDS

[2020] Kalle Parvinen. Evolution of dispersal in a structured metapopulation model in discrete time. Bulletin of Mathematical Biology, 68(3):655–678, April 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
005-9040-1;   http://link.springer.com/content/pdf/10.1007/
s11538-005-9040-1.pdf.

Ma:2006:GFS

[2021] Junling Ma and David J. D. Earn. Generality of the final size formula for an epidemic of a newly invading infectious disease. Bulletin of Mathematical Biology, 68(3):679–702, April 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-005-9047-7;
 http://link.springer.com/content/pdf/10.1007/s11538-005-9047-
7.pdf.

Nieto:2006:FPS

[2022] Juan J. Nieto, A. Torres, D. N. Georgiou, and T. E. Karakasidis. Fuzzy polynucleotide spaces and metrics. Bulletin of Mathematical Biology, 68 (3):703–725, April 2006. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-005-9020-5;   http://link.springer.com/content/
pdf/10.1007/s11538-005-9020-5.pdf.

Deng:2006:WND

[2023] Bo Deng. Why is the number of DNA bases 4? Bulletin of Mathematical Biology, 68(3):727–733, April 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-005-9019-y;   http://link.springer.com/
content/pdf/10.1007/s11538-005-9019-y.pdf.

Gonzalez:2006:AMI

[2024] Maykel Pérez González, Julio Caballero, Aliuska Morales Helguera, Miguel Garriga, Gerardo González, and Michael Fernández. 2D autocorrelation modelling of the inhibitory activity of cytokinin–derived cyclin–dependent kinase inhibitors. Bulletin of Mathematical Biology, 68(4):735–751, May 2006. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-005-9006-3;   http://link.springer.com/content/
pdf/10.1007/s11538-005-9006-3.pdf.

Forest:2006:CMS

[2025] Loïc Forest and Jacques Demongeot. Cellular modelling of secondary radial growth in conifer trees: Application to Pinus radiata (D. Don). Bulletin of Mathematical Biology, 68(4):753–784, May 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
005-9004-5;   http://link.springer.com/content/pdf/10.1007/
s11538-005-9004-5.pdf. See erratum [2157].

Sadovsky:2006:ICN

[2026] M. G. Sadovsky. Information capacity of nucleotide sequences and its applications. Bulletin of Mathematical Biology, 68(4):785–806, May 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
005-9017-0;   http://link.springer.com/content/pdf/10.1007/
s11538-005-9017-0.pdf.

Tyvand:2006:MMH

[2027] Peder A. Tyvand and Steinar Thorvaldsen. Markov model of haploid random mating with given distribution of population size. Bulletin of Mathematical Biology, 68(4):807–819, May 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-005-9026-z;
 http://link.springer.com/content/pdf/10.1007/s11538-005-9026-
z.pdf.

Recker:2006:CIR

[2028] M. Recker and S. Gupta. Conflicting immune responses can prolong the length of infection in plasmodium falciparum malaria. Bulletin of Mathematical Biology, 68(4):821–835, May 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-005-9041-0;
 http://link.springer.com/content/pdf/10.1007/s11538-005-9041-
0.pdf.

Gallegos:2006:TCM

[2029] Angela Gallegos, Barbara Mazzag, and Alex Mogilner. Two continuum models for the spreading of myxobacteria swarms. Bulletin of Mathematical Biology, 68(4):837–861, May 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-005-9031-2;
 http://link.springer.com/content/pdf/10.1007/s11538-005-9031-
2.pdf.

Chen:2006:MEC

[2030] C. Y. Chen and J. R. King. Modelling the effect of caveolae on Gprotein activation. Bulletin of Mathematical Biology, 68(4):863–888, May 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
005-9058-4;   http://link.springer.com/content/pdf/10.1007/
s11538-005-9058-4.pdf.

McBryde:2006:BME

[2031] E. S. McBryde, G. Gibson, A. N. Pettitt, Y. Zhang, B. Zhao, and D. L. S. McElwain. Bayesian modelling of an epidemic of severe acute respiratory syndrome. Bulletin of Mathematical Biology, 68(4):889–917, May 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
005-9005-4;   http://link.springer.com/content/pdf/10.1007/
s11538-005-9005-4.pdf.

Willson:2006:URT

[2032] Stephen J. Willson. Unique reconstruction of tree-like phylogenetic networks from distances between leaves. Bulletin of Mathematical Biology, 68(4):919–944, May 2006. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-005-9044-x;   http://link.springer.com/content/
pdf/10.1007/s11538-005-9044-x.pdf.

Iber:2006:IAI

[2033] Dagmar Iber and Iain D. Campbell. Integrin activation — the importance of a positive feedback. Bulletin of Mathematical Biology, 68(4):945–956, May 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
005-9049-5;   http://link.springer.com/content/pdf/10.1007/
s11538-005-9049-5.pdf.

Xiao:2006:EHP

[2034] Sa Xiao, Shu-Yan Chen, and Gang Wang. An ESS for the height of a plant population, or an optimal height for an individual? — rethinking game-theoretic models for plant height. Bulletin of Mathematical Biology, 68(4):957–967, May 2006. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-006-9073-0;   http://link.springer.com/content/
pdf/10.1007/s11538-006-9073-0.pdf.

Anonymous:2006:BMB

[2035] Anonymous. Bulletin of Mathematical Biology special issue. Bulletin of Mathematical Biology, 68(5):969–980, July 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

Barrass:2006:MTM

[2036] Iain Barrass, Edmund J. Crampin, and Philip K. Maini. Mode transitions in a model reaction–diffusion system driven by domain growth and noise. Bulletin of Mathematical Biology, 68(5):981–995, July 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
006-9106-8;   http://link.springer.com/content/pdf/10.1007/
s11538-006-9106-8.pdf.

Khanin:2006:FBT

[2037] Raya Khanin, Itzchak Parnas, and Hanna Parnas. On the feedback between theory and experiment in elucidating the molecular mechanisms underlying neurotransmitter release. Bulletin of Mathematical Biology, 68(5):997–1009, July 2006. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-006-9099-3;   http://link.springer.com/content/
pdf/10.1007/s11538-006-9099-3.pdf.

DeBoer:2006:ELD

[2038] Rob J. De Boer, Vitaly V. Ganusov, Dejan Milutinović, Philip D. Hodgkin, and Alan S. Perelson. Estimating lymphocyte division and death rates from CFSE data. Bulletin of Mathematical Biology, 68(5): 1011–1031, July 2006. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-006-9094-8;   http://link.springer.com/content/
pdf/10.1007/s11538-006-9094-8.pdf.

Keller:2006:SMF

[2039] Evelyn Fox Keller. Science as a medium for friendship: How the Keller– Segel models came about. Bulletin of Mathematical Biology, 68(5): 1033–1037, July 2006. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-006-9097-5;   http://link.springer.com/content/
pdf/10.1007/s11538-006-9097-5.pdf.

Sliusarenko:2006:BMM

[2040] Oleksii Sliusarenko, Jing Chen, and George Oster. From biochemistry to morphogenesis in myxobacteria. Bulletin of Mathematical Biology, 68 (5):1039–1051, July 2006. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-006-9113-9;   http://link.springer.com/content/
pdf/10.1007/s11538-006-9113-9.pdf.

Cytrynbaum:2006:NMS

[2041] E. N. Cytrynbaum, V. Rodionov, and A. Mogilner. Nonlocal mechanism of self-organization and centering of microtubule asters. Bulletin of Mathematical Biology, 68(5):1053–1072, July 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-006-9092-x;
 http://link.springer.com/content/pdf/10.1007/s11538-006-9092-
x.pdf.

Mehr:2006:FLR

[2042] Ramit Mehr. Feedback loops, reversals and nonlinearities in lymphocyte development. Bulletin of Mathematical Biology, 68(5):1073–1094, July 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
006-9124-6;   http://link.springer.com/content/pdf/10.1007/
s11538-006-9124-6.pdf.

Goldbeter:2006:OWC

[2043] Albert Goldbeter. Oscillations and waves of cyclic AMP in dictyostelium: a prototype for spatio-temporal organization and pulsatile intercellular communication. Bulletin of Mathematical Biology, 68(5):1095–1109, July 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
006-9090-z;   http://link.springer.com/content/pdf/10.1007/
s11538-006-9090-z.pdf.

Ma:2006:ERA

[2044] Junling Ma and Simon A. Levin. The evolution of resource adaptation: How generalist and specialist consumers evolve. Bulletin of Mathematical Biology, 68(5):1111–1123, July 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-006-9096-6;
 http://link.springer.com/content/pdf/10.1007/s11538-006-9096-
6.pdf.

He:2006:TES

[2045] Xiaoyi He, Daniel Coombs, David G. Myszka, and Byron Goldstein. A theoretical and experimental study of competition between solution and surface receptors for ligand in a Biacore flow cell. Bulletin of Mathematical Biology, 68(5):1125–1150, July 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
006-9093-9;   http://link.springer.com/content/pdf/10.1007/
s11538-006-9093-9.pdf.

Ward:2006:AMR

[2046] Michael J. Ward. Asymptotic methods for reaction–diffusion systems: Past and present. Bulletin of Mathematical Biology, 68(5):1151–1167, July 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
006-9091-y;   http://link.springer.com/content/pdf/10.1007/
s11538-006-9091-y.pdf.

Maree:2006:PMK

[2047] Athanasius F. M. Marée, Alexandra Jilkine, Adriana Dawes, Verônica A. Grieneisen, and Leah Edelstein-Keshet. Polarization and movement of keratocytes: a multiscale modelling approach. Bulletin of Mathematical Biology, 68(5):1169–1211, July 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-006-9131-7;
 http://link.springer.com/content/pdf/10.1007/s11538-006-9131-
7.pdf.

Cohen:2006:ILA

[2048] Irun R. Cohen. Informational landscapes in art, science, and evolution. Bulletin of Mathematical Biology, 68(5):1213–1229, July 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
006-9118-4;   http://link.springer.com/content/pdf/10.1007/
s11538-006-9118-4.pdf.

Movileanu:2006:TPI

[2049] Liviu Movileanu, Dumitru Popescu, Stelian Ion, and Aurel I. Popescu. Transbilayer pores induced by thickness fluctuations. Bulletin of Mathematical Biology, 68(6):1231–1255, August 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
006-9069-9;   http://link.springer.com/content/pdf/10.1007/
s11538-006-9069-9.pdf.

Sacerdote:2006:CED

[2050] L. Sacerdote, A. E. P. Villa, and C. Zucca. On the classification of experimental data modeled via a stochastic leaky integrate and fire model through boundary values. Bulletin of Mathematical Biology, 68(6): 1257–1274, August 2006. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-006-9107-7;   http://link.springer.com/content/
pdf/10.1007/s11538-006-9107-7.pdf.

Hagberg:2006:CPD

[2051] Oskar Hagberg and Carsten Wiuf. Convergence properties of the degree distribution of some growing network models. Bulletin of Mathematical Biology, 68(6):1275–1291, August 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-006-9085-9;
 http://link.springer.com/content/pdf/10.1007/s11538-006-9085-
9.pdf.

Chakraborty:2006:PBM

[2052] A. Chakraborty and D. K. Bhattacharaya. A process-based mathematical model on methane production with emission indices for control. Bulletin of Mathematical Biology, 68(6):1293–1314, August 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
006-9076-x;   http://link.springer.com/content/pdf/10.1007/
s11538-006-9076-x.pdf.

Hiebeler:2006:MED

[2053] David Hiebeler. Moment equations and dynamics of a household SIS epidemiological model. Bulletin of Mathematical Biology, 68(6):1315– 1333, August 2006. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-006-9080-1;   http://link.springer.com/content/
pdf/10.1007/s11538-006-9080-1.pdf.

Torralba:2006:SIM

[2054] Antonio S. Torralba and Yoel Rodríguez. Susceptibilities of an irreversible Michaelis–Menten enzyme. Bulletin of Mathematical Biology, 68(6):1335–1353, August 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-006-9111-y;   http://link.springer.com/
content/pdf/10.1007/s11538-006-9111-y.pdf.

Jones:2006:ELL

[2055] M. B. Jones, D. L. S. McElwain, G. R. Fulford, M. J. Collins, and A. P. Roberts. The effect of the lipid layer on tear film behaviour. Bulletin of Mathematical Biology, 68(6):1355–1381, August 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
006-9105-9;   http://link.springer.com/content/pdf/10.1007/
s11538-006-9105-9.pdf.

Heuett:2006:SMO

[2056] William J. Heuett and Hong Qian. A stochastic model of oscillatory blood testosterone levels. Bulletin of Mathematical Biology, 68(6): 1383–1399, August 2006. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-006-9098-4;   http://link.springer.com/content/
pdf/10.1007/s11538-006-9098-4.pdf.

Reinker:2006:NIC

[2057] Stefan Reinker, Yue-Xian Li, and Rachel Kuske. Noise-induced coherence and network oscillations in a reduced bursting model. Bulletin of Mathematical Biology, 68(6):1401–1427, August 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
006-9089-5;   http://link.springer.com/content/pdf/10.1007/
s11538-006-9089-5.pdf.

Mapes:2006:FMI

[2058] Eric J. Mapes and Mark F. Schumaker. Framework models of ion permeation through membrane channels and the generalized King–Altman method. Bulletin of Mathematical Biology, 68(7):1429–1460, October 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
005-9016-1;   http://link.springer.com/content/pdf/10.1007/
s11538-005-9016-1.pdf.

Varon:2006:EFM

[2059] R. Varón, E. Valero, M. Molina-Alarcón, F. García-Cánovas, F. García-Molina, M. E. Fuentes, and M. García-Moreno. Expressions for the fractional modification in different monocyclic enzyme cascade systems: Analysis of their validity tested by numerical integration. Bulletin of Mathematical Biology, 68(7):1461–1493, October 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
005-9010-7;   http://link.springer.com/content/pdf/10.1007/
s11538-005-9010-7.pdf.

Boushaba:2006:MMR

[2060] Khalid Boushaba, Howard A. Levinea, and Marit Nilsen-Hamilton. A mathematical model for the regulation of tumor dormancy based on enzyme kinetics. Bulletin of Mathematical Biology, 68(7):1495–1526, October 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
005-9042-z;   http://link.springer.com/content/pdf/10.1007/
s11538-005-9042-z.pdf.

Cruz-Monteagudo:2006:SSFa

[2061] Maykel Cruz-Monteagudo, Humberto González-Díaz, and Eugenio Uriarte. Simple stochastic fingerprints towards mathematical modeling in biology and medicine 2. Unifying Markov model for drugs side effects. Bulletin of Mathematical Biology, 68(7):1527–1554, October 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
005-9013-4;   http://link.springer.com/content/pdf/10.1007/
s11538-005-9013-4.pdf.

Cruz-Monteagudo:2006:SSFb

[2062] Maykel Cruz-Monteagudo, Humberto González-Díaz, Fernanda Borges, and Yenny González-Díaz. Simple stochastic fingerprints towards mathematical modeling in biology and medicine. 3. Ocular irritability classification model. Bulletin of Mathematical Biology, 68(7):1555–1572, October 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
006-9083-y;   http://link.springer.com/content/pdf/10.1007/
s11538-006-9083-y.pdf.

Komarova:2006:SSM

[2063] Natalia L. Komarova. Spatial stochastic models for cancer initiation and progression. Bulletin of Mathematical Biology, 68(7):1573–1599, October 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
005-9046-8;   http://link.springer.com/content/pdf/10.1007/
s11538-005-9046-8.pdf.

Topaz:2006:NCM

[2064] Chad M. Topaz, Andrea L. Bertozzi, and Mark A. Lewis. A nonlocal continuum model for biological aggregation. Bulletin of Mathematical Biology, 68(7):1601–1623, October 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-006-9088-6;
 http://link.springer.com/content/pdf/10.1007/s11538-006-9088-
6.pdf.

Marcano:2006:OAD

[2065] Mariano Marcano, Anita T. Layton, and Harold E. Layton. An optimization algorithm for a distributed–loop model of an avian urine concentrating mechanism. Bulletin of Mathematical Biology, 68(7):1625–1660, October 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
006-9087-1;   http://link.springer.com/content/pdf/10.1007/
s11538-006-9087-1.pdf.

Chignola:2006:PDB

[2066] R. Chignola, P. Dai Pra, L. M. Morato, and P. Siri. Proliferation and death in a binary environment: a stochastic model of cellular ecosystems. Bulletin of Mathematical Biology, 68(7):1661–1680, October 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
006-9078-8;   http://link.springer.com/content/pdf/10.1007/
s11538-006-9078-8.pdf.

Roussel:2006:SKD

[2067] Marc R. Roussel and Rui Zhu. Stochastic kinetics description of a simple transcription model. Bulletin of Mathematical Biology, 68(7): 1681–1713, October 2006. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-005-9048-6;   http://link.springer.com/content/
pdf/10.1007/s11538-005-9048-6.pdf.

Garske:2006:ETM

[2068] Tini Garske. Error thresholds in a mutation–selection model with Hopfield-type fitness. Bulletin of Mathematical Biology, 68(7):1715– 1746, October 2006. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-006-9072-1;   http://link.springer.com/content/
pdf/10.1007/s11538-006-9072-1.pdf.

Ndifon:2006:SPR

[2069] Wilfred Ndifon, Asamoah Nkwanta, and Dwayne Hill. Some probabilistic results on the nonrandomness of simple sequence repeats in DNA sequences. Bulletin of Mathematical Biology, 68(7):1747–1759, October 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
005-9029-9;   http://link.springer.com/content/pdf/10.1007/
s11538-005-9029-9.pdf.

Keener:2006:HST

[2070] J. P. Keener. How Salmonella typhimurium measures the length of flagellar filaments. Bulletin of Mathematical Biology, 68(7):1761–1778, October 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
005-9033-0;   http://link.springer.com/content/pdf/10.1007/
s11538-005-9033-0.pdf.

Diederichs:2006:MSM

[2071] Frank Diederichs. Mathematical simulation of membrane processes and metabolic fluxes of the pancreatic β-cell. Bulletin of Mathematical Biology, 68(7):1779–1818, October 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-005-9053-9;
 http://link.springer.com/content/pdf/10.1007/s11538-005-9053-
9.pdf.

Tosin:2006:MCM

[2072] A. Tosin, D. Ambrosi, and L. Preziosi. Mechanics and chemotaxis in the morphogenesis of vascular networks. Bulletin of Mathematical Biology, 68(7):1819–1836, October 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-006-9071-2;
 http://link.springer.com/content/pdf/10.1007/s11538-006-9071-
2.pdf.

Li:2006:IFD

[2073] Liang Li and Gang Wang. The ideal free distribution of clonal plant’s ramets among patches in a heterogeneous environment. Bulletin of Mathematical Biology, 68(8):1837–1850, November 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
005-9011-6;   http://link.springer.com/content/pdf/10.1007/
s11538-005-9011-6.pdf.

Graves:2006:BAD

[2074] Wendy Gruner Graves, Bruce Peckham, and John Pastor. A bifurcation analysis of a differential equations model for mutualism. Bulletin of Mathematical Biology, 68(8):1851–1872, November 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
006-9070-3;   http://link.springer.com/content/pdf/10.1007/
s11538-006-9070-3.pdf.

Lubkin:2006:ODT

[2075] S. R. Lubkin and X. Wan. Optimizing detection of tissue anisotropy by fluorescence recovery after photobleaching. Bulletin of Mathematical Biology, 68(8):1873–1891, November 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-006-9074-z;   http://link.springer.com/
content/pdf/10.1007/s11538-006-9074-z.pdf.

Colizza:2006:MGE

[2076] V. Colizza, A. Barrat, M. Barthélemy, and A. Vespignani. The modeling of global epidemics: Stochastic dynamics and predictability. Bulletin of Mathematical Biology, 68(8):1893–1921, November 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
006-9077-9;   http://link.springer.com/content/pdf/10.1007/
s11538-006-9077-9.pdf.

Antal:2006:FSE

[2077] Tibor Antal and István Scheuring. Fixation of strategies for an evolutionary game in finite populations. Bulletin of Mathematical Biology, 68(8): 1923–1944, November 2006. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-006-9061-4;   http://link.springer.com/content/
pdf/10.1007/s11538-006-9061-4.pdf.

Otero:2006:SPD

[2078] Marcelo Otero, Hernán G. Solari, and Nicolás Schweigmann. A stochastic population dynamics model for aedes aegypti: Formulation and application to a city with temperate climate. Bulletin of Mathematical Biology, 68(8):1945–1974, November 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
006-9067-y;   http://link.springer.com/content/pdf/10.1007/
s11538-006-9067-y.pdf.

Neville:2006:IBP

[2079] A. A. Neville, P. C. Matthews, and H. M. Byrne. Interactions between pattern formation and domain growth. Bulletin of Mathematical Biology, 68(8):1975–2003, November 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-006-9060-5;
 http://link.springer.com/content/pdf/10.1007/s11538-006-9060-
5.pdf.

Bortz:2006:MSM

[2080] D. M. Bortz and P. W. Nelson. Model selection and mixed–effects modeling of HIV infection dynamics. Bulletin of Mathematical Biology, 68(8): 2005–2025, November 2006. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-006-9084-x;   http://link.springer.com/content/
pdf/10.1007/s11538-006-9084-x.pdf.

Roose:2006:MMS

[2081] T. Roose, S. J. Chapman, and P. K. Maini. A mathematical model for simultaneous spatio–temporal dynamics of calcium and inositol 1,4,5-trisphosphate in Madin–Darby canine kidney epithelial cells. Bulletin of Mathematical Biology, 68(8):2027–2051, November 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-

006-9064-1;   http://link.springer.com/content/pdf/10.1007/
s11538-006-9064-1.pdf.

Isaeva:2006:TPM

[2082] Valeria Isaeva, Eugene Presnov, and Alexey Chernyshev. Topological patterns in metazoan evolution and development. Bulletin of Mathematical Biology, 68(8):2053–2067, November 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
006-9063-2;   http://link.springer.com/content/pdf/10.1007/
s11538-006-9063-2.pdf.

Franke:2006:SFP

[2083] John E. Franke and Abdul-Aziz Yakubu. Signature function for predicting resonant and attenuant population 2-cycles. Bulletin of Mathematical Biology, 68(8):2069–2104, November 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-006-9086-8;
 http://link.springer.com/content/pdf/10.1007/s11538-006-9086-
8.pdf.

Gumel:2006:MSS

[2084] A. B. Gumel, Connell C. McCluskey, and P. van den Driessche. Mathematical study of a staged–progression HIV model with imperfect vaccine. Bulletin of Mathematical Biology, 68(8):2105–2128, November 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
006-9095-7;   http://link.springer.com/content/pdf/10.1007/
s11538-006-9095-7.pdf.

Lutscher:2006:EHS

[2085] Frithjof Lutscher, Mark A. Lewis, and Edward McCauley. Effects of heterogeneity on spread and persistence in rivers. Bulletin of Mathematical Biology, 68(8):2129–2160, November 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
006-9100-1;   http://link.springer.com/content/pdf/10.1007/
s11538-006-9100-1.pdf.

Ngwa:2006:PDM

[2086] Gideon A. Ngwa. On the population dynamics of the malaria vector. Bulletin of Mathematical Biology, 68(8):2161–2189, November 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
006-9104-x;   http://link.springer.com/content/pdf/10.1007/
s11538-006-9104-x.pdf.

Teboh-Ewungkem:2006:RCC

[2087] Miranda I. Teboh-Ewungkem and Eric P. Salathe. The role of counter–current exchange in preventing hypoxia in skeletal muscle. Bulletin of Mathematical Biology, 68(8):2191–2204, November 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
006-9102-z;   http://link.springer.com/content/pdf/10.1007/
s11538-006-9102-z.pdf.

Ventura:2006:COW

[2088] Alejandra C. Ventura and James Sneyd. Calcium oscillations and waves generated by multiple release mechanisms in pancreatic acinar cells. Bulletin of Mathematical Biology, 68(8):2205–2231, November 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
006-9101-0;   http://link.springer.com/content/pdf/10.1007/
s11538-006-9101-0.pdf.

Warrender:2006:MII

[2089] Christina Warrender, Stephanie Forrest, and Frederick Koster. Modeling intercellular interactions in early mycobacterium infection. Bulletin of Mathematical Biology, 68(8):2233–2261, November 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
006-9103-y;   http://link.springer.com/content/pdf/10.1007/
s11538-006-9103-y.pdf.

Coutinhoa:2006:TCN

[2090] F. A. B. Coutinhoa, M. N. Burattinia, L. F. Lopeza, and E. Massada. Threshold conditions for a non–autonomous epidemic system describing the population dynamics of dengue. Bulletin of Mathematical Biology, 68(8):2263–2282, November 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-006-9108-6;
 http://link.springer.com/content/pdf/10.1007/s11538-006-9108-
6.pdf.

Bungay:2006:MTG

[2091] Sharene D. Bungay, Patricia A. Gentry, and Rodney D. Gentry. Modelling thrombin generation in human ovarian follicular fluid. Bulletin of Mathematical Biology, 68(8):2283–2302, November 2006. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
006-9115-7;   http://link.springer.com/content/pdf/10.1007/
s11538-006-9115-7.pdf.

Roy:2006:CET

[2092] Shovonlal Roy, S. Alam, and J. Chattopadhyay. Competing effects of toxin-producing phytoplankton on overall plankton populations in the Bay of Bengal. Bulletin of Mathematical Biology, 68(8):2303– 2320, November 2006. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-006-9109-5;   http://link.springer.com/content/
pdf/10.1007/s11538-006-9109-5.pdf.

Maini:2007:E

[2093] Philip K. Maini, Mark Chaplain, and Achi Dosanjh. Editorial. Bulletin of Mathematical Biology, 69(1):1, January 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
006-9191-8;   http://link.springer.com/content/pdf/10.1007/
s11538-006-9191-8.pdf.

Krottje:2007:MFM

[2094] Johannes K. Krottje and Arjen van Ooyen. A mathematical framework for modeling axon guidance. Bulletin of Mathematical Biology, 69(1):3–31, January 2007. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-006-9142-4;   http://link.springer.com/content/
pdf/10.1007/s11538-006-9142-4.pdf.

Lander:2007:MAN

[2095] A. D. Lander, Q. Nie, and F. Y. M. Wan. Membrane-associated non– receptors and morphogen gradients. Bulletin of Mathematical Biology, 69(1):33–54, January 2007. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-006-9152-2;   http://link.springer.com/content/
pdf/10.1007/s11538-006-9152-2.pdf.

Mugwagwa:2007:CSH

[2096] T. Mugwagwa and G. Witten. Coreceptor switching in HIV-1 Subtype B and Subtype C. Bulletin of Mathematical Biology, 69(1):55–75, January 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
006-9137-1;   http://link.springer.com/content/pdf/10.1007/
s11538-006-9137-1.pdf.

Seno:2007:MMC

[2097] Hiromi Seno and Mikiko Shigemoto. A mathematical modelling for the cheliped regeneration with handedness in fiddler crab. Bulletin of Mathematical Biology, 69(1):77–92, January 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-006-9155-z;
 http://link.springer.com/content/pdf/10.1007/s11538-006-9155-
z.pdf.

Zager:2007:DNP

[2098] Michael G. Zager, Paul M. Schlosser, and Hien T. Tran. A delayed nonlinear PBPK model for genistein dosimetry in rats. Bulletin of Mathematical Biology, 69(1):93–117, January 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-006-9068-x;
 http://link.springer.com/content/pdf/10.1007/s11538-006-9068-
x.pdf.

Arii:2007:EPE

[2099] Ken Arii, Raphaëlle Derome, and Lael Parrott. Examining the potential effects of species aggregation on the network structure of food webs. Bulletin of Mathematical Biology, 69(1):119–133, January 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
006-9065-0;   http://link.springer.com/content/pdf/10.1007/
s11538-006-9065-0.pdf.

Martcheva:2007:IAE

[2100] Maia Martcheva and Benjamin M. Bolker. The impact of the Allee effect in dispersal and patch–occupancy age on the dynamics of metapopulations. Bulletin of Mathematical Biology, 69(1):135–156, January 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
006-9075-y;   http://link.springer.com/content/pdf/10.1007/
s11538-006-9075-y.pdf.

Denman:2007:MMA

[2101] Paula K. Denman, D. L. Sean McElwain, Damien G. Harkin, and Zee Upton. Mathematical modelling of aerosolised skin grafts incorporating keratinocyte clonal subtypes. Bulletin of Mathematical Biology, 69 (1):157–179, January 2007. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-006-9082-z;   http://link.springer.com/content/
pdf/10.1007/s11538-006-9082-z.pdf.

Freijer:2007:ACD

[2102] Jan I. Freijer, Teun M. Post, Bart A. Ploeger, Joost DeJongh, and Meindert Danhof. Application of the convection–dispersion equation to modelling oral drug absorption. Bulletin of Mathematical Biology, 69 (1):181–195, January 2007. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-006-9122-8;   http://link.springer.com/content/
pdf/10.1007/s11538-006-9122-8.pdf.

Alarcon:2007:MCG

[2103] T. Alarcón and M. J. Tindall. Modelling cell growth and its modulation of the G_{1} /S transition. Bulletin of Mathematical Biology, 69 (1):197–214, January 2007. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-006-9154-0;   http://link.springer.com/content/
pdf/10.1007/s11538-006-9154-0.pdf.

Jose:2007:ERC

[2104] Marco V. José, Eberto R. Morgado, and Tzipe Govezensky. An extended RNA code and its relationship to the standard genetic code: an algebraic and geometrical approach. Bulletin of Mathematical Biology, 69 (1):215–243, January 2007. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-006-9119-3;   http://link.springer.com/content/
pdf/10.1007/s11538-006-9119-3.pdf.

Pustovalov:2007:MGM

[2105] V. K. Pustovalov and B. Jean. Melanin granule models for the processes of laser-induced thermal damage in pigmented retinal tissues. I. Modeling of laser-induced heating of melanosomes and selective thermal processes in retinal tissues. Bulletin of Mathematical Biology, 69 (1):245–263, January 2007. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-006-9123-7;   http://link.springer.com/content/
pdf/10.1007/s11538-006-9123-7.pdf.

Gaff:2007:MTB

[2106] Holly D. Gaff and Louis J. Gross. Modeling tick-borne disease: a metapopulation model. Bulletin of Mathematical Biology, 69(1):265– 288, January 2007. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-006-9125-5;   http://link.springer.com/content/
pdf/10.1007/s11538-006-9125-5.pdf.

Lubkin:2007:VTL

[2107] D. J. Lubkin, E. A. Gaffney, and J. R. Blake. A viscoelastic traction layer model of muco–ciliary transport. Bulletin of Mathematical Biology, 69 (1):289–327, January 2007. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-005-9036-x;   http://link.springer.com/content/
pdf/10.1007/s11538-005-9036-x.pdf.

Eskola:2007:MDV

[2108] Hanna T. M. Eskola and Stefan A. H. Geritz. On the mechanistic derivation of various discrete–time population models. Bulletin of Mathematical Biology, 69(1):329–346, January 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-006-9126-4;
 http://link.springer.com/content/pdf/10.1007/s11538-006-9126-
4.pdf.

Gonzalez:2007:QSU

[2109] Maykel Pérez González, Carmen Terán, Marta Teijeira, and Aliuska Morales Helguera. QSAR studies using radial distribution function for predicting A_{1} adenosine receptors agonists. Bulletin of Mathematical Biology, 69(1):347–359, January 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-006-9127-3;
 http://link.springer.com/content/pdf/10.1007/s11538-006-9127-
3.pdf.

Skjondal-Bara:2007:DMP

[2110] Nadav Skjøndal-Bara and David R. Morrisb. Dynamic model of the process of protein synthesis in eukaryotic cells. Bulletin of Mathematical Biology, 69(1):361–393, January 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-006-9128-2;
 http://link.springer.com/content/pdf/10.1007/s11538-006-9128-

2.pdf. See erratum [2159].

Sturm:2007:SMP

[2111] Robert Sturm and W. Hofmann. Stochastic modeling predictions for the clearance of insoluble particles from the tracheobronchial tree of the human lung. Bulletin of Mathematical Biology, 69(1):395–415, January 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
006-9143-3;   http://link.springer.com/content/pdf/10.1007/
s11538-006-9143-3.pdf.

Bearon:2007:MBC

[2112] R. N. Bearon. A model for bacterial colonization of sinking aggregates. Bulletin of Mathematical Biology, 69(1):417–431, January 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-005-9038-8;   https://link.springer.com/content/pdf/
10.1007/s11538-005-9038-8.pdf.

Pedersena:2007:TQS

[2113] Morten Gram Pedersena, Alberto M. Bersanib, and Enrico Bersanic. The total quasi–steady-state approximation for fully competitive enzyme reactions. Bulletin of Mathematical Biology, 69(1):433–457, January 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
006-9136-2;   http://link.springer.com/content/pdf/10.1007/
s11538-006-9136-2.pdf.

Schaeffer:2007:IBM

[2114] David G. Schaeffer, John W. Cain, Daniel J. Gauthier, Soma S. Kalb, Robert A. Oliver, Elena G. Tolkacheva, Wenjun Ying, and Wanda Krassowska. An ionically based mapping model with memory for cardiac restitution. Bulletin of Mathematical Biology, 69(2):459–482, February 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
006-9116-6;   http://link.springer.com/content/pdf/10.1007/
s11538-006-9116-6.pdf.

Cinquin:2007:GSL

[2115] Olivier Cinquin and Karen M. Page. Generalized, switch-like competitive heterodimerization networks. Bulletin of Mathematical Biology, 69(2): 483–494, February 2007. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-006-9114-8;   http://link.springer.com/content/
pdf/10.1007/s11538-006-9114-8.pdf.

Denman:2007:ATW

[2116] Paula K. Denman, D. L. Sean McElwain, and John Norbury. Analysis of travelling waves associated with the modelling of aerosolised skin grafts. Bulletin of Mathematical Biology, 69(2):495–523, February 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
006-9138-0;   http://link.springer.com/content/pdf/10.1007/
s11538-006-9138-0.pdf.

Hall:2007:EPC

[2117] Richard J. Hall, Simon Gubbins, and Christopher A. Gilligan. Evaluating the performance of chemical control in the presence of resistant pathogens. Bulletin of Mathematical Biology, 69(2):525–537, February 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
006-9139-z;   http://link.springer.com/content/pdf/10.1007/
s11538-006-9139-z.pdf.

Mach:2007:MVS

[2118] Robert Mach and Frank Schweitzer. Modeling vortex swarming in Daphnia. Bulletin of Mathematical Biology, 69(2):539–562, February 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
006-9135-3;   http://link.springer.com/content/pdf/10.1007/
s11538-006-9135-3.pdf.

Adams:2007:EPH

[2119] B. M. Adams, H. T. Banks, M. Davidian, and E. S. Rosenberg. Estimation and prediction with HIV–Treatment interruption data. Bulletin of Mathematical Biology, 69(2):563–584, February 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
006-9140-6;   http://link.springer.com/content/pdf/10.1007/
s11538-006-9140-6.pdf.

Niemi:2007:SDM

[2120] Jari Niemi, Ulla Ruotsalainen, Antti Saarinen, and Keijo Ruohonen. Stochastic dynamic model for estimation of rate constants and their variances from noisy and heterogeneous PET measurements. Bulletin of Mathematical Biology, 69(2):585–604, February 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
006-9150-4;   http://link.springer.com/content/pdf/10.1007/
s11538-006-9150-4.pdf.

Boswell:2007:DFN

[2121] Graeme P. Boswell, Helen Jacobs, Karl Ritz, Geoffrey M. Gadd, and Fordyce A. Davidson. The development of fungal networks in complex environments. Bulletin of Mathematical Biology, 69(2):605–634, February 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
005-9056-6;   http://link.springer.com/content/pdf/10.1007/
s11538-005-9056-6.pdf.

Gatenby:2007:ITL

[2122] Robert A. Gatenby and B. Roy Frieden. Information theory in living systems, methods, applications, and challenges. Bulletin of Mathematical Biology, 69(2):635–657, February 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-006-9141-5;
 http://link.springer.com/content/pdf/10.1007/s11538-006-9141-
5.pdf.

Huia:2007:SPP

[2123] C. Huia and M. A. McGeoch. Spatial patterns of Prisoner’s dilemma game in metapopulations. Bulletin of Mathematical Biology, 69(2): 659–676, February 2007. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-006-9145-1;   http://link.springer.com/content/
pdf/10.1007/s11538-006-9145-1.pdf.

Michel:2007:AMG

[2124] Christian J. Michel. An analytical model of gene evolution with 9 mutation parameters: an application to the amino acids coded by the common circular code. Bulletin of Mathematical Biology, 69(2):677–698, February 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-

006-9147-z;   http://link.springer.com/content/pdf/10.1007/
s11538-006-9147-z.pdf.

Tchuenche:2007:TPD

[2125] Jean M. Tchuenche. Theoretical population dynamics model of a genetically transmitted disease: Sickle–cell anaemia. Bulletin of Mathematical Biology, 69(2):699–730, February 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-006-9148-y;
 http://link.springer.com/content/pdf/10.1007/s11538-006-9148-
y.pdf.

Gao:2007:IVS

[2126] Shujing Gao, Lansun Chen, and Zhidong Teng. Impulsive vaccination of an SEIRS model with time delay and varying total population size. Bulletin of Mathematical Biology, 69(2):731–745, February 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
006-9149-x;   http://link.springer.com/content/pdf/10.1007/
s11538-006-9149-x.pdf.

Jabbarzadeh:2007:SCR

[2127] Ehsan Jabbarzadeh and Cameron F. Abrams. Simulations of chemotaxis and random motility in 2D random porous domains. Bulletin of Mathematical Biology, 69(2):747–764, February 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
006-9153-1;   http://link.springer.com/content/pdf/10.1007/
s11538-006-9153-1.pdf.

Alpkvista:2007:MMC

[2128] Erik Alpkvista and Isaac Klapper. A multidimensional multispecies continuum model for heterogeneous biofilm development. Bulletin of Mathematical Biology, 69(2):765–789, February 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
006-9168-7;   http://link.springer.com/content/pdf/10.1007/
s11538-006-9168-7.pdf.

Fleishman:2007:BRR

[2129] Erica Fleishman. Book review: R. Ferrière, U. Dieckmann, and D. Couvet, editors: Evolutionary Conservation Biology, Cambridge University Press, ISBN 0-521-82700-0, 2004, $95.00. Bulletin of Mathematical Biology, 69(2):791–793, February 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-006-9117-5;
 http://link.springer.com/content/pdf/10.1007/s11538-006-9117-
5.pdf.

Romero:2007:IST

[2130] Martín Romero. Introduction to statistics through resampling methods and R/S-Plus. Bulletin of Mathematical Biology, 69(2):795–796, February 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
006-9120-x;   http://link.springer.com/content/pdf/10.1007/
s11538-006-9120-x.pdf.

Corander:2007:RPM

[2131] Jukka Corander, Mats Gyllenberg, and Timo Koski. Random partition models and exchangeability for Bayesian identification of population structure. Bulletin of Mathematical Biology, 69(3):797–815, April 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
006-9161-1;   http://link.springer.com/content/pdf/10.1007/
s11538-006-9161-1.pdf.

Smith:2007:MTT

[2132] D. J. Smith, E. A. Gaffney, and J. R. Blake. A model of tracer transport in airway surface liquid. Bulletin of Mathematical Biology, 69(3):817–836, April 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
006-9163-z;   http://link.springer.com/content/pdf/10.1007/
s11538-006-9163-z.pdf.

Shpaka:2007:RLC

[2133] Max Shpaka and Stephen R. Proulx. The role of life cycle and migration in selection for variance in offspring number. Bulletin of Mathematical Biology, 69(3):837–860, April 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-006-9164-y;
 http://link.springer.com/content/pdf/10.1007/s11538-006-9164-
y.pdf.

Gurney:2007:DSA

[2134] William S. C. Gurney and A. Roy Veitch. The dynamics of size-at-age variability. Bulletin of Mathematical Biology, 69(3):861–885, April 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
006-9167-8;   http://link.springer.com/content/pdf/10.1007/
s11538-006-9167-8.pdf.

Layton:2007:RUU

[2135] Anita T. Layton. Role of UTB urea transporters in the urine concentrating mechanism of the rat kidney. Bulletin of Mathematical Biology, 69(3):887–929, April 2007. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-005-9030-3;   http://link.springer.com/content/
pdf/10.1007/s11538-005-9030-3.pdf.

Garvie:2007:FDS

[2136] Marcus R. Garvie. Finite-difference schemes for reaction–diffusion equations modeling predator–prey interactions in MATLAB. Bulletin of Mathematical Biology, 69(3):931–956, April 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-006-9062-3;
 http://link.springer.com/content/pdf/10.1007/s11538-006-9062-
3.pdf.

Britton:2007:EDP

[2137] N. F. Britton, R. Planqué, and N. R. Franks. Evolution of defence portfolios in exploiter–victim systems. Bulletin of Mathematical Biology, 69 (3):957–988, April 2007. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-006-9178-5;   http://link.springer.com/content/
pdf/10.1007/s11538-006-9178-5.pdf.

Cull:2007:PMS

[2138] P. Cull. Population models: Stability in one dimension. Bulletin of Mathematical Biology, 69(3):989–1017, April 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-006-9129-1;
 http://link.springer.com/content/pdf/10.1007/s11538-006-9129-
1.pdf.

Hu:2007:TTI

[2139] Kai Hu. Terminal twist-induced writhe of DNA with intrinsic curvature. Bulletin of Mathematical Biology, 69(3):1019–1030, April 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-

006-9156-y;   http://link.springer.com/content/pdf/10.1007/
s11538-006-9156-y.pdf.

Bampfylde:2007:BCT

[2140] C. J. Bampfylde and M. A. Lewis. Biological control through intraguild predation: Case studies in pest control, invasive species and range expansion. Bulletin of Mathematical Biology, 69(3):1031–1066, April 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
006-9158-9;   http://link.springer.com/content/pdf/10.1007/
s11538-006-9158-9.pdf.

Bacaer:2007:ABR

[2141] Nicolas Bacaër. Approximation of the basic reproduction number R_{0} for vector-borne diseases with a periodic vector population. Bulletin of Mathematical Biology, 69(3):1067–1091, April 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
006-9166-9;   http://link.springer.com/content/pdf/10.1007/
s11538-006-9166-9.pdf.

Mitchener:2007:GDL

[2142] W. Garrett Mitchener. Game dynamics with learning and evolution of universal grammar. Bulletin of Mathematical Biology, 69(3):1093–1118, April 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
006-9165-x;   http://link.springer.com/content/pdf/10.1007/
s11538-006-9165-x.pdf.

Rueda:2007:BRB

[2143] Raul Rueda. Book review: Bayesian Logical Data Analysis for the Physical Sciences: a Comparative Approach with Mathematica Support by P. C. Gregory. Hardcover: 486 pages. Cambridge University Press. ISBN 0-521-84150-X, $75.00. Bulletin of Mathematical Biology, 69(3):1119–1120, April 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
006-9144-2;   http://link.springer.com/content/pdf/10.1007/
s11538-006-9144-2.pdf.

Luther:2007:FWF

[2144] Roger M. Luther, Mark Broom, and Graeme D. Ruxton. Is food worth fighting for? ESS’s in mixed populations of kleptoparasites and foragers. Bulletin of Mathematical Biology, 69(4):1121–1146, May 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
005-9052-x;   http://link.springer.com/content/pdf/10.1007/
s11538-005-9052-x.pdf.

Tindall:2007:MCC

[2145] M. J. Tindall and C. P. Please. Modelling the cell cycle and cell movement in multicellular tumour spheroids. Bulletin of Mathematical Biology, 69 (4):1147–1165, May 2007. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-006-9110-z;   http://link.springer.com/content/
pdf/10.1007/s11538-006-9110-z.pdf.

Witting:2007:BIS

[2146] Lars Witting. Behavioural interactions selecting for symmetry and asymmetry in sexual reproductive systems of eusocial species. Bulletin of Mathematical Biology, 69(4):1167–1198, May 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-006-9112-x;
 http://link.springer.com/content/pdf/10.1007/s11538-006-9112-
x.pdf.

Dittrich:2007:COT

[2147] Peter Dittrich and Pietro Speroni di Fenizio. Chemical organisation theory. Bulletin of Mathematical Biology, 69(4):1199–1231, May 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
006-9130-8;   http://link.springer.com/content/pdf/10.1007/
s11538-006-9130-8.pdf.

August:2007:DML

[2148] E. August, K. H. Parker, and M. Barahona. A dynamical model of lipoprotein metabolism. Bulletin of Mathematical Biology, 69(4): 1233–1254, May 2007. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-006-9132-6;   http://link.springer.com/content/
pdf/10.1007/s11538-006-9132-6.pdf.

Mesterton-Gibbons:2007:SEG

[2149] Mike Mesterton-Gibbons and Tom N. Sherratt. Social eavesdropping: a game-theoretic analysis. Bulletin of Mathematical Biology, 69(4): 1255–1276, May 2007. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-006-9151-3;   http://link.springer.com/content/
pdf/10.1007/s11538-006-9151-3.pdf.

Bhattacharyya:2007:MRA

[2150] S. Bhattacharyya and D. K. Bhattacharya. A more realistic approach to pest-management problem. Bulletin of Mathematical Biology, 69(4): 1277–1310, May 2007. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-006-9159-8;   http://link.springer.com/content/
pdf/10.1007/s11538-006-9159-8.pdf.

Nyquist:2007:MRS

[2151] B. Nyquist, R. Tyson, and K. Larsen. Modelling recolonization of secondgrowth forest stands by the North American red squirrel Tamiasciurus hudsonicus. Bulletin of Mathematical Biology, 69(4):1311–1339, May 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
006-9160-2;   http://link.springer.com/content/pdf/10.1007/
s11538-006-9160-2.pdf.

de-Camino-Beck:2007:NMC

[2152] T. de Camino-Beck and M. A. Lewis. A new method for calculating net reproductive rate from graph reduction with applications to the control of invasive species. Bulletin of Mathematical Biology, 69(4):1341–1354, May 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
006-9162-0;   http://link.springer.com/content/pdf/10.1007/
s11538-006-9162-0.pdf.

Hsieh:2007:ITB

[2153] Ying-Hen Hsieh, P. van den Driessche, and Lin Wang. Impact of travel between patches for spatial spread of disease. Bulletin of Mathematical Biology, 69(4):1355–1375, May 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-006-9169-6;
 http://link.springer.com/content/pdf/10.1007/s11538-006-9169-
6.pdf.

Tao:2007:SFT

[2154] Yi Tao and Ross Cressman. Stochastic fluctuations through intrinsic noise in evolutionary game dynamics. Bulletin of Mathematical Biology, 69(4):1377–1399, May 2007. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-006-9170-0;   http://link.springer.com/content/
pdf/10.1007/s11538-006-9170-0.pdf.

Guirey:2007:EFD

[2155] E. J. Guirey, M. A. Bees, A. P. Martin, M. A. Srokosz, and M. J. R. Fasham. Emergent features due to grid-cell biology: Synchronisation in biophysical models. Bulletin of Mathematical Biology, 69(4):1401–1422, May 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
006-9180-y;   http://link.springer.com/content/pdf/10.1007/
s11538-006-9180-y.pdf.

Olasagasti:2007:EPM

[2156] Felix Olasagasti, Alvaro Moreno, Juli Peretó, and Federico Morán. Energetically plausible model of a self-maintaining protocellular system. Bulletin of Mathematical Biology, 69(4):1423–1445, May 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
006-9171-z;   http://link.springer.com/content/pdf/10.1007/
s11538-006-9171-z.pdf.

Forest:2007:ECM

[2157] Loïc Forest and Jacques Demongeot. Erratum: Cellular Modelling of Secondary Radial Growth in Conifer Trees: Application to Pinus radiata (D. Don). Bulletin of Mathematical Biology, 69(4):1447, May 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
006-9157-x;   http://link.springer.com/content/pdf/10.1007/
s11538-006-9157-x.pdf. See [2025].

Smith:2007:VTL

[2158] D. J. Smith, E. A. Gaffney, and J. R. Blake. A viscoelastic traction layer model of muco-ciliary transport. Bulletin of Mathematical Biology, 69(4):1449, May 2007. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-006-9177-6;   http://link.springer.com/content/
pdf/10.1007/s11538-006-9177-6.pdf.

Skjondal-Bara:2007:EDM

[2159] Nadav Skjøndal-Bara and David R. Morris. Erratum: Dynamic Model of the Process of Protein Synthesis in Eukaryotic Cells. Bulletin of Mathematical Biology, 69(4):1451, May 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-006-9179-4;
 http://link.springer.com/content/pdf/10.1007/s11538-006-9179-

4.pdf. See [2110].

Cojocaru:2007:DVS

[2160] Monica-Gabriela Cojocaru, Chris T. Bauch, and Matthew D. Johnston. Dynamics of vaccination strategies via projected dynamical systems. Bulletin of Mathematical Biology, 69(5):1453–1476, July 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-006-9173-x;   https://link.springer.com/content/pdf/
10.1007/s11538-006-9173-x.pdf.

Smith:2007:DCM

[2161] D. J. Smith, E. A. Gaffney, and J. R. Blake. Discrete cilia modelling with singularity distributions: Application to the embryonic node and the airway surface liquid. Bulletin of Mathematical Biology, 69(5):1477–1510, July 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-006-9172-y;   https://link.springer.
com/content/pdf/10.1007/s11538-006-9172-y.pdf.

Feng:2007:EMN

[2162] Zhilan Feng, Dashun Xu, and Haiyun Zhao. Epidemiological models with non–exponentially distributed disease stages and applications to disease control. Bulletin of Mathematical Biology, 69(5):1511–1536, July 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-006-9174-9;   https://link.springer.com/content/pdf/
10.1007/s11538-006-9174-9.pdf.

Eftimie:2007:MGF

[2163] R. Eftimie, G. de Vries, M. A. Lewis, and F. Lutscher. Modeling group formation and activity patterns in self-organizing collectives of individuals. Bulletin of Mathematical Biology, 69(5):1537–1565, July 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-006-9175-8;   https://link.springer.com/content/pdf/
10.1007/s11538-006-9175-8.pdf.

Paszek:2007:MSG

[2164] Pawel Paszek. Modeling stochasticity in gene regulation: Characterization in the terms of the underlying distribution function. Bulletin of Mathematical Biology, 69(5):1567–1601, July 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-006-9176-7;
 https://link.springer.com/content/pdf/10.1007/s11538-006-9176-
7.pdf.

Ioslovich:2007:EEE

[2165] Ilya Ioslovich and Per-Olof Gutman. Evaluation of experiments for estimation of dynamical crop model parameters. Bulletin of Mathematical Biology, 69(5):1603–1614, July 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-006-9181-x;   https://link.springer.
com/content/pdf/10.1007/s11538-006-9181-x.pdf.

Lutscher:2007:SNS

[2166] Frithjof Lutscher. A short note on short dispersal events. Bulletin of Mathematical Biology, 69(5):1615–1630, July 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-006-9182-9;
 https://link.springer.com/content/pdf/10.1007/s11538-006-9182-
9.pdf.

Hernandez:2007:GMD

[2167] Julio A. Hernández. A general model for the dynamics of the cell volume. Bulletin of Mathematical Biology, 69(5):1631–1648, July 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-006-9183-8;   https://link.springer.com/content/pdf/
10.1007/s11538-006-9183-8.pdf.

Utz:2007:QLC

[2168] Margarete Utz, Éva Kisdi, and Michael Doebeli. Quasi-local competition in stage–structured metapopulations: a new mechanism of pattern formation. Bulletin of Mathematical Biology, 69(5):1649–1672, July 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-006-9184-7;   https://link.springer.com/content/pdf/
10.1007/s11538-006-9184-7.pdf.

Basse:2007:GAP

[2169] Britta Basse and Paolo Ubezio. A generalised age- and phase–structured model of human tumour cell populations both unperturbed and exposed to a range of cancer therapies. Bulletin of Mathematical Biology, 69(5):1673–1690, July 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-006-9185-6;   https://link.springer.
com/content/pdf/10.1007/s11538-006-9185-6.pdf.

Kozusko:2007:MBM

[2170] Frank Kozusko, Michele Bourdeau, Zeljko Bajzer, and David Dingli. A microenvironment based model of antimitotic therapy of Gompertzian tumor growth. Bulletin of Mathematical Biology, 69(5):1691–1708, July 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-006-9186-5;   https://link.springer.com/content/pdf/
10.1007/s11538-006-9186-5.pdf.

Willson:2007:UDS

[2171] Stephen J. Willson. Unique determination of some homoplasies at hybridization events. Bulletin of Mathematical Biology, 69(5):1709–1725, July 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-006-9187-4;   https://link.springer.com/content/pdf/
10.1007/s11538-006-9187-4.pdf.

Vellela:2007:QAS

[2172] Melissa Vellela and Hong Qian. A quasistationary analysis of a stochastic chemical reaction: Keizer’s paradox. Bulletin of Mathematical Biology, 69(5):1727–1746, July 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-006-9188-3;   https://link.springer.
com/content/pdf/10.1007/s11538-006-9188-3.pdf.

Webb:2007:MBA

[2173] Steven D. Webb, Markus R. Owen, Helen M. Byrne, Craig Murdoch, and Claire E. Lewis. Macrophage-based anti-cancer therapy: Modelling different modes of tumour targeting. Bulletin of Mathematical Biology, 69(5):1747–1776, July 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-006-9189-2;   https://link.springer.
com/content/pdf/10.1007/s11538-006-9189-2.pdf.

Bergman:2007:DBH

[2174] David J. Bergman and Jacob S. Ishay. Do bees and hornets use acoustic resonance in order to monitor and coordinate comb construction? Bulletin of Mathematical Biology, 69(5):1777–1790, July 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-006-9190-9;   https://link.springer.com/content/pdf/
10.1007/s11538-006-9190-9.pdf.

Anderson:2007:PFB

[2175] David F. Anderson, Jonathan C. Mattingly, H. Frederik Nijhout, and Michael C. Reed. Propagation of fluctuations in biochemical systems, I: Linear SSC networks. Bulletin of Mathematical Biology, 69(6):1791– 1813, August 2007. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-007-9192-2;   http://link.springer.com/content/
pdf/10.1007/s11538-007-9192-2.pdf.

Torres:2007:NMS

[2176] J.-L. Torres, D. Castillo-Guajardo, M. Equihua, O. Pérez-Maqueo, V. Sosa, and L. Torres. On a new method to study drastic change in ecosystems. Bulletin of Mathematical Biology, 69(6):1815–1826, August 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
007-9193-1;   http://link.springer.com/content/pdf/10.1007/
s11538-007-9193-1.pdf.

Edwards:2007:PSV

[2177] Helen J. Edwards, Calvin Dytham, Jonathan W. Pitchford, and David Righton. Prey selection, vertical migrations and the impacts of harvesting upon the population dynamics of a predator-prey system. Bulletin of Mathematical Biology, 69(6):1827–1846, August 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
007-9194-0;   http://link.springer.com/content/pdf/10.1007/
s11538-007-9194-0.pdf.

Bunimovich-Mendrazitsky:2007:MMB

[2178] Svetlana Bunimovich-Mendrazitsky, Eliezer Shochat, and Lewi Stone. Mathematical model of BCG immunotherapy in superficial bladder cancer. Bulletin of Mathematical Biology, 69(6):1847–1870, August 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
007-9195-z;   http://link.springer.com/content/pdf/10.1007/
s11538-007-9195-z.pdf.

Korobeinikov:2007:GPI

[2179] Andrei Korobeinikov. Global properties of infectious disease models with nonlinear incidence. Bulletin of Mathematical Biology, 69(6): 1871–1886, August 2007. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-007-9196-y;   http://link.springer.com/content/
pdf/10.1007/s11538-007-9196-y.pdf.

Mollee:2007:MST

[2180] T. R. Mollee and A. J. Bracken. A model of solute transport through Stratum Corneum using solute capture and release. Bulletin of Mathematical Biology, 69(6):1887–1907, August 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
007-9197-x;   http://link.springer.com/content/pdf/10.1007/
s11538-007-9197-x.pdf.

Singh:2007:DMA

[2181] Abhyudai Singh and João Pedro Hespanha. A derivative matching approach to moment closure for the stochastic logistic model. Bulletin of Mathematical Biology, 69(6):1909–1925, August 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
007-9198-9;   http://link.springer.com/content/pdf/10.1007/
s11538-007-9198-9.pdf.

Edwards:2007:BMC

[2182] Carina M. Edwards and S. Jonathan Chapman. Biomechanical modelling of colorectal crypt budding and fission. Bulletin of Mathematical Biology, 69(6):1927–1942, August 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-007-9199-8;
 http://link.springer.com/content/pdf/10.1007/s11538-007-9199-
8.pdf.

Jilkine:2007:MMS

[2183] Alexandra Jilkine, Athanasius F. M. Marée, and Leah Edelstein-Keshet. Mathematical model for spatial segregation of the rho-family GTPases based on inhibitory crosstalk. Bulletin of Mathematical Biology, 69(6): 1943–1978, August 2007. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-007-9200-6;   http://link.springer.com/content/
pdf/10.1007/s11538-007-9200-6.pdf.

Schumaker:2007:SOB

[2184] Mark F. Schumaker. Single-occupancy binding in simple bounded and unbounded systems. Bulletin of Mathematical Biology, 69(6):1979– 2003, August 2007. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-007-9201-5;   http://link.springer.com/content/
pdf/10.1007/s11538-007-9201-5.pdf.

Cook:2007:BIS

[2185] Alex Cook, Glenn Marion, Adam Butler, and Gavin Gibson. Bayesian inference for the spatio–temporal invasion of alien species. Bulletin of Mathematical Biology, 69(6):2005–2025, August 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-007-9202-4;
 http://link.springer.com/content/pdf/10.1007/s11538-007-9202-
4.pdf.

Rong:2007:EHD

[2186] Libin Rong, Zhilan Feng, and Alan S. Perelson. Emergence of HIV-1 drug resistance during antiretroviral treatment. Bulletin of Mathematical Biology, 69(6):2027–2060, August 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-007-9203-3;
 http://link.springer.com/content/pdf/10.1007/s11538-007-9203-
3.pdf.

Mukandavire:2007:ASS

[2187] Z. Mukandavire and W. Garira. Age and sex structured model for assessing the demographic impact of mother-to-child transmission of HIV/AIDS. Bulletin of Mathematical Biology, 69(6):2061–2092, August 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
007-9204-2;   http://link.springer.com/content/pdf/10.1007/
s11538-007-9204-2.pdf.

Huxel:2007:ASI

[2188] Gary R. Huxel. Antagonistic and synergistic interactions among predators. Bulletin of Mathematical Biology, 69(6):2093–2104, August 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
007-9214-0;   http://link.springer.com/content/pdf/10.1007/
s11538-007-9214-0.pdf.

Deng:2007:OTS

[2189] Bo Deng. The origin of two sexes through optimization of recombination entropy against time and energy. Bulletin of Mathematical Biology, 69(6): 2105–2114, August 2007. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-007-9215-z;   http://link.springer.com/content/
pdf/10.1007/s11538-007-9215-z.pdf.

Hui:2007:SPP

[2190] C. Hui and M. A. McGeoch. Spatial patterns of Prisoner’s dilemma game in metapopulations. Bulletin of Mathematical Biology, 69(6):2115, August 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
007-9209-x;   http://link.springer.com/content/pdf/10.1007/
s11538-007-9209-x.pdf.

Coutinho:2007:TCN

[2191] F. A. B. Coutinho, M. N. Burattini, L. F. Lopez, and E. Massad. Threshold conditions for a non-autonomous epidemic system describing the population dynamics of dengue. Bulletin of Mathematical Biology, 69 (6):2117, August 2007. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-007-9210-4;   http://link.springer.com/content/
pdf/10.1007/s11538-007-9210-4.pdf.

Landman:2007:TWA

[2192] Kerry A. Landman, Anna Q. Cai, and Barry D. Hughes. Travelling waves of attached and detached cells in a wound–healing cell migration assay. Bulletin of Mathematical Biology, 69(7):2119–2138, October 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
007-9206-0;   http://link.springer.com/content/pdf/10.1007/
s11538-007-9206-0.pdf.

Banks:2007:EDR

[2193] H. T. Banks, John E. Banks, Lara K. Dick, and John D. Stark. Estimation of dynamic rate parameters in insect populations undergoing sublethal exposure to pesticides. Bulletin of Mathematical Biology, 69(7): 2139–2180, October 2007. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-007-9207-z;   http://link.springer.com/content/
pdf/10.1007/s11538-007-9207-z.pdf.

Jung:2007:MMV

[2194] Eunok Jung. A mathematical model of valveless pumping: a lumped model with time-dependent compliance, resistance, and inertia. Bulletin of Mathematical Biology, 69(7):2181–2198, October 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
007-9208-y;   http://link.springer.com/content/pdf/10.1007/
s11538-007-9208-y.pdf.

Whiteley:2007:STM

[2195] Jonathan P. Whiteley, Martin J. Bishop, and David J. Gavaghan. Soft tissue modelling of cardiac fibres for use in coupled mechano-electric simulations. Bulletin of Mathematical Biology, 69(7):2199–2225, October 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
007-9213-1;   http://link.springer.com/content/pdf/10.1007/
s11538-007-9213-1.pdf.

Boldin:2007:REB

[2196] B. Boldin, M. J. M. Bonten, and O. Diekmann. Relative effects of barrier precautions and topical antibiotics on nosocomial bacterial transmission: Results of multi-compartment models. Bulletin of Mathematical Biology, 69(7):2227–2248, October 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-007-9205-1;
 http://link.springer.com/content/pdf/10.1007/s11538-007-9205-
1.pdf.

Campbell:2007:CSV

[2197] R. B. Campbell. Coalescent size versus coalescent time with strong selection. Bulletin of Mathematical Biology, 69(7):2249–2259, October 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-

007-9218-9;   http://link.springer.com/content/pdf/10.1007/
s11538-007-9218-9.pdf.

Yerushalmi:2007:EEF

[2198] Uri Yerushalmi and Mina Teicher. Examining emergence of functional gene clustering in a simulated evolution. Bulletin of Mathematical Biology, 69(7):2261–2280, October 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-007-9219-8;   http://link.springer.com/
content/pdf/10.1007/s11538-007-9219-8.pdf.

Baeumer:2007:FRD

[2199] Boris Baeumer, Mihály Kovács, and Mark M. Meerschaert. Fractional reproduction–dispersal equations and heavy tail dispersal kernels. Bulletin of Mathematical Biology, 69(7):2281–2297, October 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
007-9220-2;   http://link.springer.com/content/pdf/10.1007/
s11538-007-9220-2.pdf.

Shochat:2007:GCC

[2200] E. Shochat, V. Rom-Kedar, and L. A. Segel. G-CSF control of neutrophils dynamics in the blood. Bulletin of Mathematical Biology, 69(7): 2299–2338, October 2007. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-007-9221-1;   http://link.springer.com/content/
pdf/10.1007/s11538-007-9221-1.pdf.

Cantrell:2007:DDB

[2201] Robert Stephen Cantrell and Chris Cosner. Density dependent behavior at habitat boundaries and the Allee effect. Bulletin of Mathematical Biology, 69(7):2339–2360, October 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-007-9222-0;
 http://link.springer.com/content/pdf/10.1007/s11538-007-9222-
0.pdf.

Ball:2007:MWH

[2202] Colleen L. Ball, Michael A. Gilchrist, and Daniel Coombs. Modeling within-host evolution of HIV: Mutation, competition and strain replacement. Bulletin of Mathematical Biology, 69(7):2361–2385, October 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-

007-9223-z;   http://link.springer.com/content/pdf/10.1007/
s11538-007-9223-z.pdf.

Nicolis:2007:KAF

[2203] S. C. Nicolis. Kinetics of aggregate formation in social insects. Bulletin of Mathematical Biology, 69(7):2387–2403, October 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
007-9224-y;   http://link.springer.com/content/pdf/10.1007/
s11538-007-9224-y.pdf.

Landman:2007:CPO

[2204] Kerry A. Landman and Anna Q. Cai. Cell proliferation and oxygen diffusion in a vascularising scaffold. Bulletin of Mathematical Biology, 69(7): 2405–2428, October 2007. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-007-9225-x;   http://link.springer.com/content/
pdf/10.1007/s11538-007-9225-x.pdf.

Gorban:2007:MTS

[2205] A. N. Gorban and A. Y. Zinovyev. The mystery of two straight lines in bacterial genome statistics. Bulletin of Mathematical Biology, 69(7): 2429–2442, October 2007. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-007-9229-6;   http://link.springer.com/content/
pdf/10.1007/s11538-007-9229-6.pdf.

Shpak:2007:RLC

[2206] Max Shpak and Stephen R. Proulx. The role of life cycle and migration in selection for variance in offspring number. Bulletin of Mathematical Biology, 69(7):2443, October 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-007-9211-3;
 http://link.springer.com/content/pdf/10.1007/s11538-007-9211-
3.pdf.

Alpkvist:2007:MMC

[2207] Erik Alpkvist and Isaac Klapper. A multidimensional multispecies continuum model for heterogeneous biofilm development. Bulletin of Mathematical Biology, 69(7):2445, October 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-007-9212-2;

 http://link.springer.com/content/pdf/10.1007/s11538-007-9212-
2.pdf.

Podder:2007:CCM

[2208] C. N. Podder, O. Sharomi, A. B. Gumel, and S. Moses. To cut or not to cut: a modeling approach for assessing the role of male circumcision in HIV control. Bulletin of Mathematical Biology, 69(8): 2447–2466, November 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-007-9226-9;   https://link.springer.
com/content/pdf/10.1007/s11538-007-9226-9.pdf.

Coscoy:2007:SAS

[2209] Sylvie Coscoy, Etienne Huguet, and François Amblard. Statistical analysis of sets of random walks: How to resolve their generating mechanism. Bulletin of Mathematical Biology, 69(8):2467–2492, November 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-007-9227-8;   https://link.springer.com/content/pdf/
10.1007/s11538-007-9227-8.pdf.

Guedj:2007:PIH

[2210] J. Guedj, R. Thiébaut, and D. Commenges. Practical identifiability of HIV dynamics models. Bulletin of Mathematical Biology, 69 (8):2493–2513, November 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-007-9228-7;   https://link.springer.
com/content/pdf/10.1007/s11538-007-9228-7.pdf.

Uriu:2007:TPF

[2211] Koichiro Uriu and Yoh Iwasa. Turing pattern formation with two kinds of cells and a diffusive chemical. Bulletin of Mathematical Biology, 69(8):2515–2536, November 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-007-9230-0;   https://link.springer.
com/content/pdf/10.1007/s11538-007-9230-0.pdf.

Zhang:2007:NSM

[2212] Tailei Zhang and Zhidong Teng. On a nonautonomous SEIRS model in epidemiology. Bulletin of Mathematical Biology, 69(8):2537–2559, November 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/

s11538-007-9231-z;   https://link.springer.com/content/pdf/
10.1007/s11538-007-9231-z.pdf.

Willson:2007:RSH

[2213] Stephen J. Willson. Reconstruction of some hybrid phylogenetic networks with homoplasies from distances. Bulletin of Mathematical Biology, 69(8):2561–2590, November 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-007-9232-y;   https://link.springer.
com/content/pdf/10.1007/s11538-007-9232-y.pdf.

Yokley:2007:EMS

[2214] Karen A. Yokley and Marina V. Evans. An example of model structure differences using sensitivity analyses in physiologically based pharmacokinetic models of trichloroethylene in humans. Bulletin of Mathematical Biology, 69(8):2591–2625, November 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-007-9233-x;
 https://link.springer.com/content/pdf/10.1007/s11538-007-9233-
x.pdf.

Krakovska:2007:DSR

[2215] O. Krakovska and L. M. Wahl. Drug-sparing regimens for HIV combination therapy: Benefits predicted for “drug coasting”. Bulletin of Mathematical Biology, 69(8):2627–2647, November 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-007-9234-9;
 https://link.springer.com/content/pdf/10.1007/s11538-007-9234-
9.pdf.

Ghosh:2007:RVI

[2216] S. Ghosh, S. Bhattacharyya, and D. K. Bhattacharya. The role of viral infection in pest control: a mathematical study. Bulletin of Mathematical Biology, 69(8):2649–2691, November 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-007-9235-8;
 https://link.springer.com/content/pdf/10.1007/s11538-007-9235-
8.pdf.

Wright:2007:SPH

[2217] R. W. Wright and Alan Hastings. Spontaneous patchiness in a host– parasitoid integrodifference model. Bulletin of Mathematical Biology, 69(8):2693–2709, November 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-007-9236-7;   https://link.springer.
com/content/pdf/10.1007/s11538-007-9236-7.pdf.

Reluga:2007:OTD

[2218] Timothy C. Reluga, Jan Medlock, Eric Poolman, and Alison P. Galvani. Optimal timing of disease transmission in an age-structured population. Bulletin of Mathematical Biology, 69(8):2711–2722, November 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-007-9238-5;   https://link.springer.com/content/pdf/
10.1007/s11538-007-9238-5.pdf.

Huggins:2007:THG

[2219] Peter Huggins, Lior Pachter, and Bernd Sturmfels. Toward the human genotope. Bulletin of Mathematical Biology, 69(8):2723–2735, November 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-007-9244-7;   https://link.springer.com/content/pdf/
10.1007/s11538-007-9244-7.pdf.

Stamper:2007:MRA

[2220] I. J. Stamper, H. M. Byrne, M. R. Owen, and P. K. Maini. Modelling the role of angiogenesis and vasculogenesis in solid tumour growth. Bulletin of Mathematical Biology, 69(8):2737–2772, November 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-007-9253-6;   https://link.springer.com/content/pdf/
10.1007/s11538-007-9253-6.pdf.

Romero:2007:BRI

[2221] Martín Romero. Book review: Introduction to Computer–Intensive Methods of Data Analysis in Biology by Derek A. Roff. Bulletin of Mathematical Biology, 69(8):2773–2774, November 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-007-9216-y;
 https://link.springer.com/content/pdf/10.1007/s11538-007-9216-
y.pdf.

Rojo-Dominguez:2007:BRS

[2222] Arturo Rojo-Domínguez. Book review: Srinivas Aluru (ed): Handbook of Computational Molecular Biology. Bulletin of Mathematical Biology, 69(8):2775–2776, November 2007. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-007-9217-x;   https://link.springer.
com/content/pdf/10.1007/s11538-007-9217-x.pdf.

Logan:2008:PSP

[2223] J. David Logan. Phenologically-structured predator-prey dynamics with temperature dependence. Bulletin of Mathematical Biology, 70 (1):1–20, January 2008. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-007-9237-6;   http://link.springer.com/content/
pdf/10.1007/s11538-007-9237-6.pdf.

Lee:2008:MCP

[2224] Ha Youn Lee and Alan S. Perelson. Modeling T cell proliferation and death in vitro based on labeling data: Generalizations of the Smith–Martin cell cycle model. Bulletin of Mathematical Biology, 70 (1):21–44, January 2008. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-007-9239-4;   http://link.springer.com/content/
pdf/10.1007/s11538-007-9239-4.pdf.

Jin:2008:CRS

[2225] Emma Y. Jin, Jing Qin, and Christian M. Reidys. Combinatorics of RNA structures with pseudoknots. Bulletin of Mathematical Biology, 70 (1):45–67, January 2008. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-007-9240-y;   http://link.springer.com/content/
pdf/10.1007/s11538-007-9240-y.pdf.

Stern:2008:RBM

[2226] Julie V. Stern, Hinke M. Osinga, Andrew LeBeau, and Arthur Sherman. Resetting behavior in a model of bursting in secretory pituitary cells: Distinguishing plateaus from pseudo–plateaus. Bulletin of Mathematical Biology, 70(1):68–88, January 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-007-9241-x;
 http://link.springer.com/content/pdf/10.1007/s11538-007-9241-
x.pdf.

Jain:2008:MVB

[2227] Harsh V. Jain, Jacques E. Nör, and Trachette L. Jackson. Modeling the VEGF–Bcl-2–CXCL8 pathway in intratumoral agiogenesis. Bulletin of Mathematical Biology, 70(1):89–117, January 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
007-9242-9;   http://link.springer.com/content/pdf/10.1007/
s11538-007-9242-9.pdf.

Ma:2008:CFP

[2228] Junling Ma and P. van den Driessche. Case fatality proportion. Bulletin of Mathematical Biology, 70(1):118–133, January 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
007-9243-8;   http://link.springer.com/content/pdf/10.1007/
s11538-007-9243-8.pdf.

Hsu:2008:RAI

[2229] Sze-Bi Hsu and Ying-Hen Hsieh. On the role of asymptomatic infection in transmission dynamics of infectious diseases. Bulletin of Mathematical Biology, 70(1):134–155, January 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-007-9245-6;
 http://link.springer.com/content/pdf/10.1007/s11538-007-9245-
6.pdf.

Kasatkin:2008:MGS

[2230] Victor Kasatkin, Alain Prochiantz, and David Holcman. Morphogenetic gradients and the stability of boundaries between neighboring morphogenetic regions. Bulletin of Mathematical Biology, 70(1):156–178, January 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
007-9246-5;   http://link.springer.com/content/pdf/10.1007/
s11538-007-9246-5.pdf.

Camelo-Neto:2008:EPJ

[2231] G. Camelo-Neto, Ana T. C. Silva, L. Giuggioli, and V. M. Kenkre. Effect of predators of juvenile rodents on the spread of the hantavirus epidemic. Bulletin of Mathematical Biology, 70(1):179–188, January 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
007-9247-4;   http://link.springer.com/content/pdf/10.1007/
s11538-007-9247-4.pdf.

Cameron:2008:ORS

[2232] David A. Cameron, David J. Ivers, Theodore A. Evans, and Mary R. Myerscough. Optimal reproduction strategies in two species of moundbuilding termites. Bulletin of Mathematical Biology, 70(1):189–209, January 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
007-9248-3;   http://link.springer.com/content/pdf/10.1007/
s11538-007-9248-3.pdf.

Pang:2008:LTF

[2233] Ning-Ning Pang and Wen-Jer Tzeng. On the long–term fitness of cells in periodically switching environments. Bulletin of Mathematical Biology, 70(1):210–235, January 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-007-9250-9;   http://link.springer.com/
content/pdf/10.1007/s11538-007-9250-9.pdf.

Adler:2008:HHP

[2234] Frederick R. Adler, Jessica M. C. Pearce-Duvet, and M. Denise Dearing. How host population dynamics translate into time-lagged prevalence: an investigation of Sin Nombre virus in deer mice. Bulletin of Mathematical Biology, 70(1):236–252, January 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
007-9251-8;   http://link.springer.com/content/pdf/10.1007/
s11538-007-9251-8.pdf.

Wagner:2008:CVD

[2235] Bradley G. Wagner and David J. D. Earn. Circulating vaccine derived polio viruses and their impact on global polio eradication. Bulletin of Mathematical Biology, 70(1):253–280, January 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
007-9255-4;   http://link.springer.com/content/pdf/10.1007/
s11538-007-9255-4.pdf.

Cattoni:2008:VEC

[2236] D. I. Cattoni and O. Chara. Vacuum effects over the closing of enterocutaneous fistulae: a mathematical modeling approach. Bulletin of Mathematical Biology, 70(1):281–296, January 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-

007-9258-1;   http://link.springer.com/content/pdf/10.1007/
s11538-007-9258-1.pdf.

Zhang:2008:MRF

[2237] Jun Zhang, Stanly L. Steinberg, Bridget S. Wilson, Janet M. Oliver, and Lance R. Williams. Markov random field modeling of the spatial distribution of proteins on cell membranes. Bulletin of Mathematical Biology, 70 (1):297–321, January 2008. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-007-9259-0;   http://link.springer.com/content/
pdf/10.1007/s11538-007-9259-0.pdf.

Jones:2008:EEW

[2238] M. B. Jones, G. R. Fulford, C. P. Please, D. L. S. McElwain, and M. J. Collins. Elastohydrodynamics of the eyelid wiper. Bulletin of Mathematical Biology, 70(2):323–343, February 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
007-9252-7;   http://link.springer.com/content/pdf/10.1007/
s11538-007-9252-7.pdf.

Lemesle:2008:SUO

[2239] V. Lemesle and J. L. Gouzé. A simple unforced oscillatory growth model in the chemostat. Bulletin of Mathematical Biology, 70(2):344– 357, February 2008. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-007-9254-5;   http://link.springer.com/content/
pdf/10.1007/s11538-007-9254-5.pdf.

Gilioli:2008:BIF

[2240] Gianni Gilioli, Sara Pasquali, and Fabrizio Ruggeri. Bayesian inference for functional response in a stochastic predator–prey system. Bulletin of Mathematical Biology, 70(2):358–381, February 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
007-9256-3;   http://link.springer.com/content/pdf/10.1007/
s11538-007-9256-3.pdf.

Alexander:2008:DDM

[2241] Murray E. Alexander, Seyed M. Moghadas, Gergely Röst, and Jianhong Wu. A delay differential model for pandemic influenza with antiviral treatment. Bulletin of Mathematical Biology, 70(2):382–397, February 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
007-9257-2;   http://link.springer.com/content/pdf/10.1007/
s11538-007-9257-2.pdf.

Whittle:2008:OCG

[2242] Andrew Whittle, Suzanne Lenhart, and K. A. J. White. Optimal control of gypsy moth populations. Bulletin of Mathematical Biology, 70(2): 398–411, February 2008. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-007-9260-7;   http://link.springer.com/content/
pdf/10.1007/s11538-007-9260-7.pdf.

Petrovskii:2008:CAE

[2243] Sergei Petrovskii, Rod Blackshaw, and Bai-Lian Li. Consequences of the Allee effect and intraspecific competition on population persistence under adverse environmental conditions. Bulletin of Mathematical Biology, 70 (2):412–437, February 2008. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-007-9262-5;   http://link.springer.com/content/
pdf/10.1007/s11538-007-9262-5.pdf.

Hirashima:2008:DBA

[2244] Tsuyoshi Hirashima, Yoh Iwasa, and Yoshihiro Morishita. Distance between AER and ZPA is defined by feed-forward loop and is stabilized by their feedback loop in vertebrate limb bud. Bulletin of Mathematical Biology, 70(2):438–459, February 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
007-9263-4;   http://link.springer.com/content/pdf/10.1007/
s11538-007-9263-4.pdf.

Alber:2008:MLM

[2245] Mark Alber, Tilmann Glimm, H. G. E. Hentschel, Bogdan Kazmierczak, Yong-Tao Zhang, Jianfeng Zhu, and Stuart A. Newman. The morphostatic limit for a model of skeletal pattern formation in the vertebrate limb. Bulletin of Mathematical Biology, 70(2):460–483, February 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
007-9264-3;   http://link.springer.com/content/pdf/10.1007/
s11538-007-9264-3.pdf.

DeVille:2008:RSM

[2246] R. E. Lee DeVille and Eric Vanden-Eijnden. Regularity and synchrony in motor proteins. Bulletin of Mathematical Biology, 70(2):484–516, February 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
007-9266-1;   http://link.springer.com/content/pdf/10.1007/
s11538-007-9266-1.pdf.

Biktashev:2008:AAA

[2247] V. N. Biktashev, R. Suckley, Y. E. Elkin, and R. D. Simitev. Asymptotic analysis and analytical solutions of a model of cardiac excitation. Bulletin of Mathematical Biology, 70(2):517–554, February 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
007-9267-0;   http://link.springer.com/content/pdf/10.1007/
s11538-007-9267-0.pdf.

Yokley:2008:SIR

[2248] Karen A. Yokley, Hien Tran, and Paul M. Schlosser. Sensory irritation response in rats: Modeling, analysis and validation. Bulletin of Mathematical Biology, 70(2):555–588, February 2008. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-007-9268-z;
 http://link.springer.com/content/pdf/10.1007/s11538-007-9268-
z.pdf.

Gordillo:2008:BES

[2249] Luis F. Gordillo, Stephen A. Marion, Anders Martin-Löf, and Priscilla E. Greenwood. Bimodal epidemic size distributions for near-critical SIR with vaccination. Bulletin of Mathematical Biology, 70(2):589–602, February 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-

9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-007-9269-y;   http://link.springer.com/content/
pdf/10.1007/s11538-007-9269-y.pdf.

Baake:2008:SCD

[2250] Ellen Baake and Inke Herms. Single-crossover dynamics: Finite versus infinite populations. Bulletin of Mathematical Biology, 70(2):603– 624, February 2008. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-007-9270-5;   http://link.springer.com/content/
pdf/10.1007/s11538-007-9270-5.pdf.

Lopez-Sauceda:2008:ESU

[2251] J. López-Sauceda and J. L. Aragón. Eutacticity in sea urchin evolution. Bulletin of Mathematical Biology, 70(2):625–634, February 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
007-9273-2;   http://link.springer.com/content/pdf/10.1007/
s11538-007-9273-2.pdf.

Nadarajah:2008:EDD

[2252] Saralees Nadarajah and Samuel Kotz. The exact distribution of divergence times. Bulletin of Mathematical Biology, 70(2):635–640, February 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
007-9272-3;   http://link.springer.com/content/pdf/10.1007/
s11538-007-9272-3.pdf.

Miramontes:2008:BRC

[2253] Octavio Miramontes. Book review: Complex Systems in Biomedicine. A. Quarteroni, L. Formaggia and A. Veneziani (eds). Bulletin of Mathematical Biology, 70(2):641–642, February 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-007-9249-2;
 http://link.springer.com/content/pdf/10.1007/s11538-007-9249-
2.pdf.

Yamagishi:2008:NFH

[2254] Michel E. Beleza Yamagishi and Alex Itiro Shimabukuro. Nucleotide frequencies in human genome and Fibonacci numbers. Bulletin of Mathematical Biology, 70(3):643–653, April 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-007-9261-6;
 http://link.springer.com/content/pdf/10.1007/s11538-007-9261-
6.pdf.

Lee:2008:CTW

[2255] J. M. Lee, T. Hillen, and M. A. Lewis. Continuous traveling waves for prey–taxis. Bulletin of Mathematical Biology, 70(3):654–676, April 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
007-9271-4;   http://link.springer.com/content/pdf/10.1007/
s11538-007-9271-4.pdf.

Rejniak:2008:CSDa

[2256] Katarzyna A. Rejniak and Alexander R. A. Anderson. A computational study of the development of epithelial acini: I. Sufficient conditions for the formation of a hollow structure. Bulletin of Mathematical Biology, 70(3):677–712, April 2008. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-007-9274-1;   http://link.springer.com/content/
pdf/10.1007/s11538-007-9274-1.pdf.

Floyd:2008:SEP

[2257] William Floyd, Leslie Kay, and Michael Shapiro. Some elementary properties of SIR networks or, can I get sick because you got vaccinated? Bulletin of Mathematical Biology, 70(3):713–727, April 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
007-9275-0;   http://link.springer.com/content/pdf/10.1007/
s11538-007-9275-0.pdf.

Kim:2008:MIT

[2258] Peter S. Kim, Peter P. Lee, and Doron Levy. Modeling imatinib– treated chronic myelogenous leukemia: Reducing the complexity of agent-based models. Bulletin of Mathematical Biology, 70(3):728–744, April 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
007-9276-z;   http://link.springer.com/content/pdf/10.1007/
s11538-007-9276-z.pdf.

Bortz:2008:KPF

[2259] D. M. Bortz, T. L. Jackson, K. A. Taylor, A. P. Thompson, and J. G. Younger. Klebsiella pneumoniae flocculation dynamics. Bulletin of Mathematical Biology, 70(3):745–768, April 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
007-9277-y;   http://link.springer.com/content/pdf/10.1007/
s11538-007-9277-y.pdf.

Vaidya:2008:MSB

[2260] Naveen K. Vaidya and Jianhong Wu. Modeling spruce budworm population revisited: Impact of physiological structure on outbreak control. Bulletin of Mathematical Biology, 70(3):769–784, April 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-

007-9278-x;   http://link.springer.com/content/pdf/10.1007/
s11538-007-9278-x.pdf.

Wu:2008:PIE

[2261] Hulin Wu, Haihong Zhu, Hongyu Miao, and Alan S. Perelson. Parameter identifiability and estimation of HIV/ AIDS dynamic models. Bulletin of Mathematical Biology, 70(3):785–799, April 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
007-9279-9;   http://link.springer.com/content/pdf/10.1007/
s11538-007-9279-9.pdf.

Cogan:2008:TFM

[2262] N. G. Cogan. Two-fluid model of biofilm disinfection. Bulletin of Mathematical Biology, 70(3):800–819, April 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-007-9280-3;
 http://link.springer.com/content/pdf/10.1007/s11538-007-9280-
3.pdf.

Atkinson:2008:QRT

[2263] Michael P. Atkinson and Lawrence M. Wein. Quantifying the routes of transmission for pandemic influenza. Bulletin of Mathematical Biology, 70(3):820–867, April 2008. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-007-9281-2;   http://link.springer.com/content/
pdf/10.1007/s11538-007-9281-2.pdf.

Bessonov:2008:MBP

[2264] N. Bessonov, N. Morozova, and V. Volpert. Modeling of branching patterns in plants. Bulletin of Mathematical Biology, 70(3):868–893, April 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
007-9282-1;   http://link.springer.com/content/pdf/10.1007/
s11538-007-9282-1.pdf.

Elbasha:2008:GSE

[2265] Elamin H. Elbasha. Global stability of equilibria in a two-sex HPV vaccination model. Bulletin of Mathematical Biology, 70(3):894–909, April 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
007-9283-0;   http://link.springer.com/content/pdf/10.1007/
s11538-007-9283-0.pdf.

Sulman:2008:MMM

[2266] Tatiana Sulman, Leonid B. Katsnelson, Olga Solovyova, and Vladimir S. Markhasin. Mathematical modeling of mechanically modulated rhythm disturbances in homogeneous and heterogeneous myocardium with attenuated activity of Na^{+}–K^{+} pump. Bulletin of Mathematical Biology, 70(3):910–949, April 2008. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-007-9285-y;   http://link.springer.com/content/
pdf/10.1007/s11538-007-9285-y.pdf.

Jin:2008:AER

[2267] Emma Y. Jin and Christian M. Reidys. Asymptotic enumeration of RNA structures with pseudoknots. Bulletin of Mathematical Biology, 70(4):951–970, May 2008. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-007-9265-2;   http://link.springer.com/content/
pdf/10.1007/s11538-007-9265-2.pdf.

Rudge:2008:EIE

[2268] Tim Rudge and Kevin Burrage. Effects of intrinsic and extrinsic noise can accelerate juxtacrine pattern formation. Bulletin of Mathematical Biology, 70(4):971–991, May 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-007-9286-x;
 http://link.springer.com/content/pdf/10.1007/s11538-007-9286-
x.pdf.

Bertuzzi:2008:RSD

[2269] A. Bertuzzi, A. Fasano, A. Gandolfi, and C. Sinisgalli. Reoxygenation and split-dose response to radiation in a tumour model with Krogh-type vascular geometry. Bulletin of Mathematical Biology, 70(4):992–1012, May 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
007-9287-9;   http://link.springer.com/content/pdf/10.1007/
s11538-007-9287-9.pdf.

Wysham:2008:SSE

[2270] Derin B. Wysham and Alan Hastings. Sudden shifts in ecological systems: Intermittency and transients in the coupled Ricker population model. Bulletin of Mathematical Biology, 70(4):1013–1031, May 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-

007-9288-8;   http://link.springer.com/content/pdf/10.1007/
s11538-007-9288-8.pdf.

Narang:2008:BLO

[2271] Atul Narang and Sergei S. Pilyugin. Bistability of the lac operon during growth of Escherichia coli on lactose and lactose + glucose. Bulletin of Mathematical Biology, 70(4):1032–1064, May 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
007-9289-7;   http://link.springer.com/content/pdf/10.1007/
s11538-007-9289-7.pdf.

Lawson:2008:UCT

[2272] Daniel John Lawson and Henrik Jeldtoft Jensen. Understanding clustering in type space using field theoretic techniques. Bulletin of Mathematical Biology, 70(4):1065–1081, May 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-007-9290-1;
 http://link.springer.com/content/pdf/10.1007/s11538-007-9290-
1.pdf.

Gernhard:2008:NAR

[2273] Tanja Gernhard. New analytic results for speciation times in neutral models. Bulletin of Mathematical Biology, 70(4):1082–1097, May 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
007-9291-0;   http://link.springer.com/content/pdf/10.1007/
s11538-007-9291-0.pdf.

Ruan:2008:DRM

[2274] Shigui Ruan, Dongmei Xiao, and John C. Beier. On the delayed Ross–Macdonald model for malaria transmission. Bulletin of Mathematical Biology, 70(4):1098–1114, May 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-007-9292-z;
 http://link.springer.com/content/pdf/10.1007/s11538-007-9292-
z.pdf.

Matsen:2008:MTS

[2275] Frederick A. Matsen, Elchanan Mossel, and Mike Steel. Mixed-up trees: the structure of phylogenetic mixtures. Bulletin of Mathematical Biology, 70(4):1115–1139, May 2008. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-007-9293-y;   http://link.springer.com/content/
pdf/10.1007/s11538-007-9293-y.pdf.

Malik:2008:DDI

[2276] Tufail Malik and Hal L. Smith. Does dormancy increase fitness of bacterial populations in time-varying environments? Bulletin of Mathematical Biology, 70(4):1140–1162, May 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-008-9294-5;
 http://link.springer.com/content/pdf/10.1007/s11538-008-9294-
5.pdf.

Bhunu:2008:TTM

[2277] C. P. Bhunu, W. Garira, Z. Mukandavire, and M. Zimba. Tuberculosis transmission model with chemoprophylaxis and treatment. Bulletin of Mathematical Biology, 70(4):1163–1191, May 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
008-9295-4;   http://link.springer.com/content/pdf/10.1007/
s11538-008-9295-4.pdf.

Yang:2008:ICM

[2278] Xingzhou Yang, Robert H. Dillon, and Lisa J. Fauci. An integrative computational model of multiciliary beating. Bulletin of Mathematical Biology, 70(4):1192–1215, May 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-008-9296-3;
 http://link.springer.com/content/pdf/10.1007/s11538-008-9296-
3.pdf.

Maidana:2008:GSM

[2279] Norberto Aníbal Maidana and Wilson Castro Ferreira, Jr. The geographic spread of “El mal de las caderas” in capybaras (Hydrochaeris hydrochaeris). Bulletin of Mathematical Biology, 70(4):1216–1234, May 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
008-9297-2;   http://link.springer.com/content/pdf/10.1007/
s11538-008-9297-2.pdf.

Spasic:2008:SFA

[2280] Sladjana Spasic, Milka Culic, Gordana Grbic, Ljiljana Martac, Slobodan Sekulic, and Dragosav Mutavdzic. Spectral and fractal analysis of cerebellar activity after single and repeated brain injury. Bulletin of Mathematical Biology, 70(4):1235–1249, May 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
008-9306-5;   http://link.springer.com/content/pdf/10.1007/
s11538-008-9306-5.pdf.

Santillan:2008:ASM

[2281] Moisés Santillán and Eduardo S. Zeron. Analytical study of the multiplicity of regulatory mechanisms in the tryptophan operon. Bulletin of Mathematical Biology, 70(4):1250, May 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-007-9284-z;
 http://link.springer.com/content/pdf/10.1007/s11538-007-9284-
z.pdf.

Bertram:2008:MSD

[2282] Richard Bertram and Rudy C. Arceo II. A mathematical study of the differential effects of two SERCA isoforms on Ca^{2+} oscillations in pancreatic islets. Bulletin of Mathematical Biology, 70(5):1251–1271, July 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
008-9298-1;   http://link.springer.com/content/pdf/10.1007/
s11538-008-9298-1.pdf.

Chitnis:2008:DIP

[2283] Nakul Chitnis, James M. Hyman, and Jim M. Cushing. Determining important parameters in the spread of malaria through the sensitivity analysis of a mathematical model. Bulletin of Mathematical Biology, 70 (5):1272–1296, July 2008. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-008-9299-0;   http://link.springer.com/content/
pdf/10.1007/s11538-008-9299-0.pdf.

Otero:2008:SSD

[2284] Marcelo Otero, Nicolás Schweigmann, and Hernán G. Solari. A stochastic spatial dynamical model for aedes aegypti. Bulletin of Mathematical Biology, 70(5):1297–1325, July 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-008-9300-y;
 http://link.springer.com/content/pdf/10.1007/s11538-008-9300-
y.pdf.

Mase:2008:MME

[2285] Michela Masé, Leon Glass, and Flavia Ravelli. A model for mechano–electrical feedback effects on atrial flutter interval variability. Bulletin of Mathematical Biology, 70(5):1326–1347, July 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
008-9301-x;   http://link.springer.com/content/pdf/10.1007/
s11538-008-9301-x.pdf.

Stirling:2008:MAE

[2286] J. R. Stirling, M. S. Zakynthinaki, and V. Billat. Modeling and analysis of the effect of training on V̇ O_{2} kinetics and anaerobic capacity. Bulletin of Mathematical Biology, 70(5):1348–1370, July 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
008-9302-9;   http://link.springer.com/content/pdf/10.1007/
s11538-008-9302-9.pdf.

Potapov:2008:AEC

[2287] A. B. Potapov and M. A. Lewis. Allee effect and control of lake system invasion. Bulletin of Mathematical Biology, 70(5):1371–1397, July 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
008-9303-8;   http://link.springer.com/content/pdf/10.1007/
s11538-008-9303-8.pdf.

Aracena:2008:MNF

[2288] Julio Aracena. Maximum number of fixed points in regulatory Boolean networks. Bulletin of Mathematical Biology, 70(5):1398–1409, July 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
008-9304-7;   http://link.springer.com/content/pdf/10.1007/
s11538-008-9304-7.pdf.

Traulsen:2008:ARI

[2289] Arne Traulsen, Noam Shoresh, and Martin A. Nowak. Analytical results for individual and group selection of any intensity. Bulletin of Mathematical Biology, 70(5):1410–1424, July 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
008-9305-6;   http://link.springer.com/content/pdf/10.1007/
s11538-008-9305-6.pdf.

Arribas:2008:MLF

[2290] E. Arribas, A. Muñoz-Lopez, M. J. Garcia-Meseguer, A. Lopez-Najera, L. Avalos, F. Garcia-Molina, M. Garcia-Moreno, and R. Varon. Mean lifetime and first-passage time of the enzyme species involved in an enzyme reaction. Application to unstable enzyme systems. Bulletin of Mathematical Biology, 70(5):1425–1449, July 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
008-9307-4;   http://link.springer.com/content/pdf/10.1007/
s11538-008-9307-4.pdf.

Rejniak:2008:CSDb

[2291] Katarzyna A. Rejniak and Alexander R. A. Anderson. A computational study of the development of epithelial acini: II. Necessary conditions for structure and lumen stability. Bulletin of Mathematical Biology, 70 (5):1450–1479, July 2008. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-008-9308-3;   http://link.springer.com/content/
pdf/10.1007/s11538-008-9308-3.pdf.

Anderson:2008:TRS

[2292] Kurt E. Anderson, Roger M. Nisbet, and Edward McCauley. Transient responses to spatial perturbations in advective systems. Bulletin of Mathematical Biology, 70(5):1480–1502, July 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-008-9309-2;
 http://link.springer.com/content/pdf/10.1007/s11538-008-9309-
2.pdf.

Willms:2008:GCS

[2293] Allan R. Willms and Dominic Nelson. A geometric comparison of single chain multi-state models of ion channel gating. Bulletin of Mathematical Biology, 70(5):1503–1524, July 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-008-9310-9;
 http://link.springer.com/content/pdf/10.1007/s11538-008-9310-
9.pdf.

Tindall:2008:OMAa

[2294] M. J. Tindall, S. L. Porter, P. K. Maini, G. Gaglia, and J. P. Armitage. Overview of mathematical approaches used to model bacterial chemotaxis I: The single cell. Bulletin of Mathematical Biology, 70(6):

1525–1569, August 2008. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-008-9321-6;   http://link.springer.com/content/
pdf/10.1007/s11538-008-9321-6.pdf.

Tindall:2008:OMAb

[2295] M. J. Tindall, P. K. Maini, S. L. Porter, and J. P. Armitage. Overview of mathematical approaches used to model bacterial chemotaxis II: Bacterial populations. Bulletin of Mathematical Biology, 70(6):1570–1607, August 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
008-9322-5;   http://link.springer.com/content/pdf/10.1007/
s11538-008-9322-5.pdf.

DeVille:2008:SAF

[2296] R. E. Lee DeVille and Charles S. Peskin. Synchrony and asynchrony in a fully stochastic neural network. Bulletin of Mathematical Biology, 70(6): 1608–1633, August 2008. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-008-9311-8;   http://link.springer.com/content/
pdf/10.1007/s11538-008-9311-8.pdf.

Verdy:2008:SAR

[2297] Ariane Verdy and Hal Caswell. Sensitivity analysis of reactive ecological dynamics. Bulletin of Mathematical Biology, 70(6):1634–1659, August 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
008-9312-7;   http://link.springer.com/content/pdf/10.1007/
s11538-008-9312-7.pdf.

Gedeon:2008:WAR

[2298] Tomáš Gedeon, Konstantin Mischaikow, Kate Patterson, and Eliane Traldi. When activators repress and repressors activate: a qualitative analysis of the Shea–Ackers model. Bulletin of Mathematical Biology, 70(6):1660–1683, August 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-008-9313-6;   http://link.springer.com/
content/pdf/10.1007/s11538-008-9313-6.pdf.

Levy:2008:SMP

[2299] Doron Levy and Tiago Requeijo. Stochastic models for phototaxis. Bulletin of Mathematical Biology, 70(6):1684–1706, August 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
008-9314-5;   http://link.springer.com/content/pdf/10.1007/
s11538-008-9314-5.pdf.

Eladdadi:2008:MME

[2300] Amina Eladdadi and David Isaacson. A mathematical model for the effects of HER2 overexpression on cell proliferation in breast cancer. Bulletin of Mathematical Biology, 70(6):1707–1729, August 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
008-9315-4;   http://link.springer.com/content/pdf/10.1007/
s11538-008-9315-4.pdf.

Haseltine:2008:IGM

[2301] Eric L. Haseltine, Vy Lam, John Yin, and James B. Rawlings. Image-guided modeling of virus growth and spread. Bulletin of Mathematical Biology, 70(6):1730–1748, August 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-008-9316-3;
 http://link.springer.com/content/pdf/10.1007/s11538-008-9316-
3.pdf.

Miao:2008:MEK

[2302] Hongyu Miao, Carrie Dykes, Lisa M. Demeter, James Cavenaugh, Sung Yong Park, Alan S. Perelson, and Hulin Wu. Modeling and estimation of kinetic parameters and replicative fitness of HIV-1 from flow-cytometry-based growth Competition Experiments. Bulletin of Mathematical Biology, 70(6):1749–1771, August 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
008-9323-4;   http://link.springer.com/content/pdf/10.1007/
s11538-008-9323-4.pdf.

Roose:2008:NDB

[2303] Tiina Roose and Andrew C. Fowler. Network development in biological gels: Role in lymphatic vessel development. Bulletin of Mathematical Biology, 70(6):1772–1789, August 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-008-9324-3;
 http://link.springer.com/content/pdf/10.1007/s11538-008-9324-
3.pdf.

Kot:2008:SPA

[2304] Mark Kot and Michael G. Neubert. Saddle-point approximations, integrodifference equations, and invasions. Bulletin of Mathematical Biology, 70(6):1790–1826, August 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-008-9325-2;   http://link.springer.com/
content/pdf/10.1007/s11538-008-9325-2.pdf.

Lof:2008:ECI

[2305] Marjolein E. Lof, Rampal S. Etienne, James Powell, Maarten de Gee, and Lia Hemerik. The effect of chemical information on the spatial distribution of fruit flies: I. Model results. Bulletin of Mathematical Biology, 70(7):1827–1849, October 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-008-9327-0;
 http://link.springer.com/content/pdf/10.1007/s11538-008-9327-
0.pdf.

deGee:2008:ECI

[2306] Maarten de Gee, Marjolein E. Lof, and Lia Hemerik. The effect of chemical information on the spatial distribution of fruit flies: II. Parameterization, calibration, and sensitivity. Bulletin of Mathematical Biology, 70(7): 1850–1868, October 2008. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-008-9329-y;   http://link.springer.com/content/
pdf/10.1007/s11538-008-9329-y.pdf.

Brauer:2008:EMH

[2307] Fred Brauer. Epidemic models with heterogeneous mixing and treatment. Bulletin of Mathematical Biology, 70(7):1869–1885, October 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
008-9326-1;   http://link.springer.com/content/pdf/10.1007/
s11538-008-9326-1.pdf.

Zhao:2008:MMS

[2308] Ruijun Zhao and Fabio Augusto Milner. A mathematical model of Schistosoma mansoni in Biomphalaria glabrata with control strategies. Bulletin of Mathematical Biology, 70(7):1886–1905, October 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
008-9330-5;   http://link.springer.com/content/pdf/10.1007/
s11538-008-9330-5.pdf.

Grunewald:2008:TNC

[2309] S. Grünewald, K. T. Huber, and Q. Wu. Two novel closure rules for constructing phylogenetic super-networks. Bulletin of Mathematical Biology, 70(7):1906–1924, October 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-008-9331-4;
 http://link.springer.com/content/pdf/10.1007/s11538-008-9331-
4.pdf.

Ortega:2008:FRF

[2310] Neli R. S. Ortega, Fabiano S. Santos, Dirce M. T. Zanetta, and Eduardo Massad. A fuzzy Reed–Frost Model for epidemic spreading. Bulletin of Mathematical Biology, 70(7):1925–1936, October 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
008-9332-3;   http://link.springer.com/content/pdf/10.1007/
s11538-008-9332-3.pdf.

Campos:2008:LMI

[2311] Daniel Campos, Vicenç Méndez, and Vicente Ortega-Cejas. Lattice models for invasions through patchy environments. Bulletin of Mathematical Biology, 70(7):1937–1956, October 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-008-9333-2;
 http://link.springer.com/content/pdf/10.1007/s11538-008-9333-
2.pdf.

Morishita:2008:GBM

[2312] Yoshihiro Morishita and Yoh Iwasa. Growth based morphogenesis of vertebrate limb bud. Bulletin of Mathematical Biology, 70(7):1957–1978, October 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
008-9334-1;   http://link.springer.com/content/pdf/10.1007/
s11538-008-9334-1.pdf.

Bertram:2008:PBM

[2313] Richard Bertram, Joseph Rhoads, and Wendy P. Cimbora. A phantom bursting mechanism for episodic bursting. Bulletin of Mathematical Biology, 70(7):1979–1993, October 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-008-9335-0;
 http://link.springer.com/content/pdf/10.1007/s11538-008-9335-
0.pdf.

Kim:2008:PMI

[2314] Peter S. Kim, Peter P. Lee, and Doron Levy. A PDE model for imatinib-treated chronic myelogenous leukemia. Bulletin of Mathematical Biology, 70(7):1994–2016, October 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-008-9336-z;
 http://link.springer.com/content/pdf/10.1007/s11538-008-9336-
z.pdf.

Bairagi:2008:IID

[2315] N. Bairagi, R. R. Sarkar, and J. Chattopadhyay. Impacts of incubation delay on the dynamics of an eco-epidemiological system — a theoretical study. Bulletin of Mathematical Biology, 70(7):2017–2038, October 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
008-9337-y;   http://link.springer.com/content/pdf/10.1007/
s11538-008-9337-y.pdf.

Bode:2008:ODA

[2316] Michael Bode, Kerrie Wilson, Marissa McBride, and Hugh Possingham. Optimal dynamic allocation of conservation funding among priority regions. Bulletin of Mathematical Biology, 70(7):2039–2054, October 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
008-9343-0;   http://link.springer.com/content/pdf/10.1007/
s11538-008-9343-0.pdf.

Bunimovich-Mendrazitsky:2008:MMP

[2317] Svetlana Bunimovich-Mendrazitsky, Helen Byrne, and Lewi Stone. Mathematical model of pulsed immunotherapy for superficial bladder cancer. Bulletin of Mathematical Biology, 70(7):2055–2076, October 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
008-9344-z;   http://link.springer.com/content/pdf/10.1007/
s11538-008-9344-z.pdf.

Boldin:2008:PSG

[2318] Barbara Boldin. Persistence and spread of gastro-intestinal infections: the case of enterotoxigenic Escherichia coli in piglets. Bulletin of Mathematical Biology, 70(7):2077–2101, October 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-

008-9348-8;   http://link.springer.com/content/pdf/10.1007/
s11538-008-9348-8.pdf.

Li:2008:RSR

[2319] Yue-Xian Li and Anmar Khadra. Robust synchrony and rhythmogenesis in endocrine neurons via autocrine regulations in vitro and in vivo. Bulletin of Mathematical Biology, 70(8):2103–2125, November 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
008-9328-z;   http://link.springer.com/content/pdf/10.1007/
s11538-008-9328-z.pdf.

Elbasha:2008:MTH

[2320] Elamin H. Elbasha, Erik J. Dasbach, and Ralph P. Insinga. A multitype HPV transmission model. Bulletin of Mathematical Biology, 70(8): 2126–2176, November 2008. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-008-9338-x;   http://link.springer.com/content/
pdf/10.1007/s11538-008-9338-x.pdf.

Larsson:2008:ETR

[2321] Sara Larsson, Tobias Rydén, Ulla Holst, Stina Oredsson, and Maria Johansson. Estimating the total rate of DNA replication using branching processes. Bulletin of Mathematical Biology, 70(8):2177–2194, November 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
008-9339-9;   http://link.springer.com/content/pdf/10.1007/
s11538-008-9339-9.pdf.

Su:2008:SDE

[2322] Min Su, Cang Hui, YanYu Zhang, and Zizhen Li. Spatiotemporal dynamics of the epidemic transmission in a predator–prey system. Bulletin of Mathematical Biology, 70(8):2195–2210, November 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
008-9340-3;   http://link.springer.com/content/pdf/10.1007/
s11538-008-9340-3.pdf.

Bertram:2008:MMA

[2323] Richard Bertram and Yue-Xian Li. A mathematical model for the actions of activin, inhibin, and follistatin on pituitary gonadotrophs. Bulletin of Mathematical Biology, 70(8):2211–2228, November 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
008-9341-2;   http://link.springer.com/content/pdf/10.1007/
s11538-008-9341-2.pdf.

Gaffney:2008:MAM

[2324] E. A. Gaffney, J. K. Heath, and M. Z. Kwiatkowska. A mass action model of a fibroblast growth factor signaling pathway and its simplification. Bulletin of Mathematical Biology, 70(8):2229–2263, November 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
008-9342-1;   http://link.springer.com/content/pdf/10.1007/
s11538-008-9342-1.pdf.

Hurtado:2008:PID

[2325] Paul Hurtado. The potential impact of disease on the migratory structure of a partially migratory passerine population. Bulletin of Mathematical Biology, 70(8):2264–2282, November 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
008-9345-y;   http://link.springer.com/content/pdf/10.1007/
s11538-008-9345-y.pdf.

Jahnke:2008:DLR

[2326] Tobias Jahnke and Wilhelm Huisinga. A dynamical low-rank approach to the chemical master equation. Bulletin of Mathematical Biology, 70(8): 2283–2302, November 2008. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-008-9346-x;   http://link.springer.com/content/
pdf/10.1007/s11538-008-9346-x.pdf.

Wattis:2008:MML

[2327] J. A. D. Wattis, B. O’Malley, H. Blackburn, L. Pickersgill, J. Panovska, H. M. Byrne, and K. G. Jackson. Mathematical model for low density lipoprotein (LDL) endocytosis by hepatocytes. Bulletin of Mathematical Biology, 70(8):2303–2333, November 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-008-9347-9;
 http://link.springer.com/content/pdf/10.1007/s11538-008-9347-
9.pdf.

Chapman:2008:MMF

[2328] S. Jonathan Chapman, Rebecca J. Shipley, and Rossa Jawad. Multiscale modeling of fluid transport in tumors. Bulletin of Mathematical Biology, 70(8):2334–2357, November 2008. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-008-9349-7;
 http://link.springer.com/content/pdf/10.1007/s11538-008-9349-
7.pdf.

Armstrong:2009:AAC

[2329] Nicola J. Armstrong, Kevin J. Painter, and Jonathan A. Sherratt. Adding adhesion to a chemical signaling model for somite formation. Bulletin of Mathematical Biology, 71(1):1–24, January 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
008-9350-1;   http://link.springer.com/content/pdf/10.1007/
s11538-008-9350-1.pdf.

Burkhead:2009:DSC

[2330] E. G. Burkhead, J. M. Hawkins, and D. K. Molinek. A dynamical study of a cellular automata model of the spread of HIV in a lymph node. Bulletin of Mathematical Biology, 71(1):25–74, January 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
008-9351-0;   http://link.springer.com/content/pdf/10.1007/
s11538-008-9351-0.pdf.

Korobeinikov:2009:GPS

[2331] Andrei Korobeinikov. Global properties of SIR and SEIR epidemic models with multiple parallel infectious stages. Bulletin of Mathematical Biology, 71(1):75–83, January 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-008-9352-z;
 http://link.springer.com/content/pdf/10.1007/s11538-008-9352-
z.pdf.

Bakhtin:2009:LDR

[2332] Yuri Bakhtin and Christine E. Heitsch. Large deviations for random trees and the branching of RNA Secondary structures. Bulletin of Mathematical Biology, 71(1):84–106, January 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-008-9353-y;
 http://link.springer.com/content/pdf/10.1007/s11538-008-9353-
y.pdf.

McKenzie:2009:FPT

[2333] Hannah W. McKenzie, Mark A. Lewis, and Evelyn H. Merrill. First passage time analysis of animal movement and insights into the functional response. Bulletin of Mathematical Biology, 71(1):107–129, January 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
008-9354-x;   http://link.springer.com/content/pdf/10.1007/
s11538-008-9354-x.pdf.

Goldwyn:2009:SHL

[2334] Eli E. Goldwyn and Alan Hastings. Small heterogeneity has large effects on synchronization of ecological oscillators. Bulletin of Mathematical Biology, 71(1):130–144, January 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-008-9355-9;
 http://link.springer.com/content/pdf/10.1007/s11538-008-9355-
9.pdf.

White:2009:PDM

[2335] Michael C. White and Xiao-Qiang Zhao. A periodic droop model for two species competition in A Chemostat. Bulletin of Mathematical Biology, 71(1):145–161, January 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-008-9357-7;   http://link.springer.com/
content/pdf/10.1007/s11538-008-9357-7.pdf.

Latulippe:2009:NPM

[2336] J. Latulippe and M. Pernarowski. A nonautonomous phenomenological model for On and Off Responses of cells in sensory system. Bulletin of Mathematical Biology, 71(1):162–188, January 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
008-9358-6;   http://link.springer.com/content/pdf/10.1007/
s11538-008-9358-6.pdf.

DeLeenheer:2009:WHV

[2337] Patrick De Leenheer. Within-host virus models with periodic antiviral therapy. Bulletin of Mathematical Biology, 71(1):189–210, January 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
008-9359-5;   http://link.springer.com/content/pdf/10.1007/
s11538-008-9359-5.pdf.

Thackham:2009:CAS

[2338] Jennifer A. Thackham, D. L. Sean McElwain, and Ian W. Turner. Computational approaches to solving equations arising from wound healing. Bulletin of Mathematical Biology, 71(1):211–246, January 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
008-9360-z;   http://link.springer.com/content/pdf/10.1007/
s11538-008-9360-z.pdf.

Colijn:2009:LCM

[2339] Caroline Colijn, Ted Cohen, and Megan Murray. Latent coinfection and the maintenance of strain diversity. Bulletin of Mathematical Biology, 71 (1):247–263, January 2009. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-008-9361-y;   http://link.springer.com/content/
pdf/10.1007/s11538-008-9361-y.pdf.

Reidys:2009:LCN

[2340] Christian M. Reidys. Local connectivity of neutral networks. Bulletin of Mathematical Biology, 71(2):265–290, February 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
008-9356-8;   http://link.springer.com/content/pdf/10.1007/
s11538-008-9356-8.pdf.

Trewenack:2009:TWM

[2341] Abbey J. Trewenack and Kerry A. Landman. A traveling wave model for invasion by precursor and differentiated cells. Bulletin of Mathematical Biology, 71(2):291–317, February 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-008-9362-x;
 http://link.springer.com/content/pdf/10.1007/s11538-008-9362-
x.pdf.

Soubeyrand:2009:MSM

[2342] S. Soubeyrand, S. Neuvonen, and A. Penttinen. Mechanical-statistical modeling in ecology: From outbreak detections to pest dynamics. Bulletin of Mathematical Biology, 71(2):318–338, February 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
008-9363-9;   http://link.springer.com/content/pdf/10.1007/
s11538-008-9363-9.pdf.

Bryant:2009:HPM

[2343] David Bryant. Hadamard phylogenetic methods and the n-taxon process. Bulletin of Mathematical Biology, 71(2):339–351, February 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
008-9364-8;   http://link.springer.com/content/pdf/10.1007/
s11538-008-9364-8.pdf.

Lukeman:2009:CMM

[2344] Ryan Lukeman, Yue-Xian Li, and Leah Edelstein-Keshet. A conceptual model for milling formations in biological aggregates. Bulletin of Mathematical Biology, 71(2):352–382, February 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
008-9365-7;   http://link.springer.com/content/pdf/10.1007/
s11538-008-9365-7.pdf.

Melkikh:2009:MAT

[2345] A. V. Melkikh and V. D. Seleznev. Model of active transport of ions in archaea cells. Bulletin of Mathematical Biology, 71(2):383–398, February 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
008-9366-6;   http://link.springer.com/content/pdf/10.1007/
s11538-008-9366-6.pdf.

Berestycki:2009:CSK

[2346] H. Berestycki, O. Diekmann, C. J. Nagelkerke, and P. A. Zegeling. Can a species keep pace with a shifting climate? Bulletin of Mathematical Biology, 71(2):399–429, February 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-008-9367-5;
 http://link.springer.com/content/pdf/10.1007/s11538-008-9367-
5.pdf.

Haque:2009:RDP

[2347] Mainul Haque. Ratio-dependent predator–prey models of interacting populations. Bulletin of Mathematical Biology, 71(2):430–452, February 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
008-9368-4;   http://link.springer.com/content/pdf/10.1007/
s11538-008-9368-4.pdf.

Noel:2009:GRC

[2348] Jason T. Noel and Atul Narang. Gene regulation in continuous cultures: a unified theory for bacteria and yeasts. Bulletin of Mathematical Biology, 71(2):453–514, February 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-008-9369-3;
 http://link.springer.com/content/pdf/10.1007/s11538-008-9369-
3.pdf.

Layton:2009:MDM

[2349] Anita T. Layton, Leon C. Moore, and Harold E. Layton. Multistable dynamics mediated by tubuloglomerular feedback in a model of coupled nephrons. Bulletin of Mathematical Biology, 71(3):515–555, April 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
008-9370-x;   http://link.springer.com/content/pdf/10.1007/
s11538-008-9370-x.pdf.

Lindquist:2009:DBA

[2350] W. Brent Lindquist and Ivan D. Chase. Data-based analysis of winner–loser models of hierarchy formation in animals. Bulletin of Mathematical Biology, 71(3):556–584, April 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-008-9371-9;
 http://link.springer.com/content/pdf/10.1007/s11538-008-9371-
9.pdf.

Thalhauser:2009:ESG

[2351] Craig J. Thalhauser, Tejas Sankar, Mark C. Preul, and Yang Kuang. Explicit separation of growth and motility in a new tumor cord model. Bulletin of Mathematical Biology, 71(3):585–601, April 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
008-9372-8;   http://link.springer.com/content/pdf/10.1007/
s11538-008-9372-8.pdf.

Roeder:2009:ASM

[2352] Ingo Roeder, Maria Herberg, and Matthias Horn. An “age” structured model of hematopoietic stem cell organization with application to chronic myeloid leukemia. Bulletin of Mathematical Biology, 71(3):602–626, April 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-

008-9373-7;   http://link.springer.com/content/pdf/10.1007/
s11538-008-9373-7.pdf.

Jiang:2009:TCW

[2353] Jifa Jiang, Zhipeng Qiu, Jianhong Wu, and Huaiping Zhu. Threshold conditions for West Nile virus outbreaks. Bulletin of Mathematical Biology, 71(3):627–647, April 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-008-9374-6;   http://link.springer.com/
content/pdf/10.1007/s11538-008-9374-6.pdf.

Rida:2009:MPL

[2354] Wasima Rida and Sonja Sandberg. Modeling the population level effects of an HIV-1 vaccine in an era of highly active antiretroviral therapy. Bulletin of Mathematical Biology, 71(3):648–680, April 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
008-9375-5;   http://link.springer.com/content/pdf/10.1007/
s11538-008-9375-5.pdf.

Bahi:2009:SMG

[2355] Jacques M. Bahi and Christian J. Michel. A stochastic model of gene evolution with time dependent pseudochaotic mutations. Bulletin of Mathematical Biology, 71(3):681–700, April 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-008-9376-4;
 http://link.springer.com/content/pdf/10.1007/s11538-008-9376-
4.pdf.

Vogel:2009:GFP

[2356] Richard M. Vogel, Jonathan R. M. Hosking, Chris S. Elphick, David L. Roberts, and J. Michael Reed. Goodness of fit of probability distributions for sightings as species approach extinction. Bulletin of Mathematical Biology, 71(3):701–719, April 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
008-9377-3;   http://link.springer.com/content/pdf/10.1007/
s11538-008-9377-3.pdf.

Foss:2009:DMH

[2357] Anna M. Foss, Peter T. Vickerman, Zaid Chalabi, Philippe Mayaud, Michel Alary, and Charlotte H. Watts. Dynamic modeling of herpes simplex virus Type-2 (HSV-2) transmission: Issues in structural uncertainty. Bulletin of Mathematical Biology, 71(3):720–749, April 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
008-9379-1;   http://link.springer.com/content/pdf/10.1007/
s11538-008-9379-1.pdf.

Felix:2009:ERM

[2358] Liliana Félix, Francesc Rosselló, and Gabriel Valiente. Efficient reconstruction of metabolic pathways by bidirectional chemical search. Bulletin of Mathematical Biology, 71(3):750–769, April 2009. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-008-9380-8;
 http://link.springer.com/content/pdf/10.1007/s11538-008-9380-
8.pdf.

Roux:2009:AOL

[2359] C. Z. Roux. The applicability of ordinary least squares to consistently short distances between taxa in phylogenetic tree construction and the normal distribution test consequences. Bulletin of Mathematical Biology, 71(4):771–780, May 2009. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-008-9378-2;   http://link.springer.com/content/
pdf/10.1007/s11538-008-9378-2.pdf.

Simpson:2009:DBD

[2360] Matthew J. Simpson, Kerry A. Landman, and Barry D. Hughes. Distinguishing between directed and undirected cell motility within an invading cell population. Bulletin of Mathematical Biology, 71(4):781–799, May 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
008-9381-7;   http://link.springer.com/content/pdf/10.1007/
s11538-008-9381-7.pdf.

Demetrius:2009:EEP

[2361] Lloyd Demetrius, Stéphane Legendre, and Peter Harremöes. Evolutionary entropy: a predictor of body size, metabolic rate and maximal life span. Bulletin of Mathematical Biology, 71(4):800–818, May 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
008-9382-6;   http://link.springer.com/content/pdf/10.1007/
s11538-008-9382-6.pdf.

Martcheva:2009:ECP

[2362] Maia Martcheva. Evolutionary consequences of predation for pathogens in prey. Bulletin of Mathematical Biology, 71(4):819–844, May 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
008-9383-5;   http://link.springer.com/content/pdf/10.1007/
s11538-008-9383-5.pdf.

Joh:2009:DIT

[2363] Richard I. Joh, Hao Wang, Howard Weiss, and Joshua S. Weitz. Dynamics of indirectly transmitted infectious diseases with immunological threshold. Bulletin of Mathematical Biology, 71(4):845–862, May 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
008-9384-4;   http://link.springer.com/content/pdf/10.1007/
s11538-008-9384-4.pdf.

Morozov:2009:EPD

[2364] Andrew Morozov and Sergei Petrovskii. Excitable population dynamics, biological control failure, and spatiotemporal pattern formation in a model ecosystem. Bulletin of Mathematical Biology, 71(4):863–887, May 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
008-9385-3;   http://link.springer.com/content/pdf/10.1007/
s11538-008-9385-3.pdf.

Kiss:2009:CNB

[2365] Istvan Z. Kiss, Péter L. Simon, and Rowland R. Kao. A contact– network-based formulation of a preferential mixing model. Bulletin of Mathematical Biology, 71(4):888–905, May 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-008-9386-2;
 http://link.springer.com/content/pdf/10.1007/s11538-008-9386-
2.pdf.

Green:2009:MML

[2366] J. E. F. Green, S. L. Waters, K. M. Shakesheff, and H. M. Byrne. A mathematical model of liver cell aggregation in vitro. Bulletin of Mathematical Biology, 71(4):906–930, May 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-008-9387-1;
 http://link.springer.com/content/pdf/10.1007/s11538-008-9387-
1.pdf.

Takeuchi:2009:DAP

[2367] Yasuhiro Takeuchi, Wendi Wang, Shinji Nakaoka, and Shingo Iwami. Dynamical adaptation of parental care. Bulletin of Mathematical Biology, 71(4):931–951, May 2009. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-008-9388-0;   http://link.springer.com/content/
pdf/10.1007/s11538-008-9388-0.pdf.

Yurk:2009:MEI

[2368] Brian P. Yurk and James A. Powell. Modeling the evolution of insect phenology. Bulletin of Mathematical Biology, 71(4):952–979, May 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
008-9389-z;   http://link.springer.com/content/pdf/10.1007/
s11538-008-9389-z.pdf.

Drew:2009:AMF

[2369] Donald A. Drew, Gretchen A. Koch, Heather Vellante, Ronak Talati, and Oswaldo Sanchez. Analyses of mechanisms for force generation during cell septation in Escherichia coli. Bulletin of Mathematical Biology, 71(4):980–1005, May 2009. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-008-9390-6;   http://link.springer.com/content/
pdf/10.1007/s11538-008-9390-6.pdf.

Zeiser:2009:HMN

[2370] Stefan Zeiser, Uwe Franz, Johannes Müller, and Volkmar Liebscher. Hybrid modeling of noise reduction by a negatively autoregulated system. Bulletin of Mathematical Biology, 71(4):1006–1024, May 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
008-9391-5;   http://link.springer.com/content/pdf/10.1007/
s11538-008-9391-5.pdf.

Chara:2009:KEA

[2371] Osvaldo Chara, Diego E. Pafundo, and Pablo J. Schwarzbaum. Kinetics of extracellular ATP from goldfish hepatocytes: A Lesson from mathematical modeling. Bulletin of Mathematical Biology, 71(5):1025–1047, July 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
008-9392-4;   http://link.springer.com/content/pdf/10.1007/
s11538-008-9392-4.pdf.

Verzi:2009:CMA

[2372] D. W. Verzi and O. Y. Noris. A compartmental model for activity– dependent dendritic spine branching. Bulletin of Mathematical Biology, 71(5):1048–1072, July 2009. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-009-9393-y;   http://link.springer.com/content/
pdf/10.1007/s11538-009-9393-y.pdf.

Basu:2009:ETV

[2373] Sanjay Basu and Alison P. Galvani. The evolution of tuberculosis virulence. Bulletin of Mathematical Biology, 71(5):1073–1088, July 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
009-9394-x;   http://link.springer.com/content/pdf/10.1007/
s11538-009-9394-x.pdf.

Locsei:2009:RTC

[2374] J. T. Locsei and T. J. Pedley. Run and tumble chemotaxis in a shear flow: The effect of temporal comparisons, persistence, rotational Diffusion, and cell shape. Bulletin of Mathematical Biology, 71(5):1089–1116, July 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
009-9395-9;   http://link.springer.com/content/pdf/10.1007/
s11538-009-9395-9.pdf.

Painter:2009:CMC

[2375] Kevin J. Painter. Continuous models for cell migration in tissues and applications to cell sorting via differential chemotaxis. Bulletin of Mathematical Biology, 71(5):1117–1147, July 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-009-9396-8;
 http://link.springer.com/content/pdf/10.1007/s11538-009-9396-
8.pdf.

Gao:2009:CPA

[2376] Meng Gao. Continuous probabilistic analysis to evolutionary game dynamics in finite populations. Bulletin of Mathematical Biology, 71(5): 1148–1159, July 2009. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-009-9397-7;   http://link.springer.com/content/
pdf/10.1007/s11538-009-9397-7.pdf.

OMalley:2009:EIR

[2377] Lauren O’Malley, G. Korniss, and Thomas Caraco. Ecological invasion, roughened fronts, and a Competitor’s extreme advance: Integrating stochastic spatial–growth models. Bulletin of Mathematical Biology, 71 (5):1160–1188, July 2009. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-009-9398-6;   http://link.springer.com/content/
pdf/10.1007/s11538-009-9398-6.pdf.

Poplawski:2009:FII

[2378] Nikodem J. Popławski, Ubirajara Agero, J. Scott Gens, Maciej Swat, James A. Glazier, and Alexander R. A. Anderson. Front instabilities and invasiveness of simulated avascular tumors. Bulletin of Mathematical Biology, 71(5):1189–1227, July 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-009-9399-5;
 http://link.springer.com/content/pdf/10.1007/s11538-009-9399-
5.pdf.

Hsu:2009:MRS

[2379] Chia-Yu Hsu and Robert Dillon. A 3D motile rod-shaped monotrichous bacterial model. Bulletin of Mathematical Biology, 71(5):1228–1263, July 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
009-9400-3;   http://link.springer.com/content/pdf/10.1007/
s11538-009-9400-3.pdf.

Altenberg:2009:ERP

[2380] Lee Altenberg. The evolutionary reduction principle for linear variation in genetic transmission. Bulletin of Mathematical Biology, 71(5): 1264–1284, July 2009. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-009-9401-2;   http://link.springer.com/content/
pdf/10.1007/s11538-009-9401-2.pdf.

Weinstein:2009:MPT

[2381] Alan M. Weinstein and Eduardo D. Sontag. Modeling proximal tubule cell homeostasis: Tracking Changes in Luminal Flow. Bulletin of Mathematical Biology, 71(6):1285–1322, August 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
009-9402-1;   http://link.springer.com/content/pdf/10.1007/
s11538-009-9402-1.pdf.

Strohm:2009:EHF

[2382] S. Strohm and R. Tyson. The effect of habitat fragmentation on cyclic population dynamics: a numerical study. Bulletin of Mathematical Biology, 71(6):1323–1348, August 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-009-9403-0;
 http://link.springer.com/content/pdf/10.1007/s11538-009-9403-
0.pdf.

Atamas:2009:DDS

[2383] Sergei P. Atamas and Jonathan Bell. Degeneracy-driven self-structuring dynamics in Selective Repertoires. Bulletin of Mathematical Biology, 71(6):1349–1365, August 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-009-9404-z;   http://link.springer.com/
content/pdf/10.1007/s11538-009-9404-z.pdf.

Beard:2009:ADT

[2384] Daniel A. Beard and Fan Wu. Apparent diffusivity and Taylor dispersion of water and solutes in capillary beds. Bulletin of Mathematical Biology, 71(6):1366–1377, August 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-009-9405-y;
 http://link.springer.com/content/pdf/10.1007/s11538-009-9405-
y.pdf.

Cruz-Pacheco:2009:SOW

[2385] Gustavo Cruz-Pacheco, Lourdes Esteva, and Cristobal Vargas. Seasonality and outbreaks in West Nile virus infection. Bulletin of Mathematical Biology, 71(6):1378–1393, August 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-009-9406-x;
 http://link.springer.com/content/pdf/10.1007/s11538-009-9406-
x.pdf.

Fournier:2009:SMN

[2386] Thomas Fournier, Jean-Pierre Gabriel, Christian Mazza, Jerôme Pasquier, José Galbete, and Nicolas Mermod. Stochastic models and numerical algorithms for a class of regulatory gene networks. Bulletin of Mathematical Biology, 71(6):1394–1431, August 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-

009-9407-9;   http://link.springer.com/content/pdf/10.1007/
s11538-009-9407-9.pdf.

Fenwick:2009:EFA

[2387] Michael K. Fenwick and Fernando A. Escobedo. Exploration of factors affecting the onset and maturation course of follicular lymphoma through simulations of the germinal center. Bulletin of Mathematical Biology, 71(6):1432–1462, August 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-009-9408-8;
 http://link.springer.com/content/pdf/10.1007/s11538-009-9408-
8.pdf.

Cardoso:2009:MOE

[2388] Rodrigo T. N. Cardoso, André R. da Cruz, Elizabeth F. Wanner, and Ricardo H. C. Takahashi. Multi-objective evolutionary optimization of biological pest control with impulsive dynamics in soybean crops. Bulletin of Mathematical Biology, 71(6):1463–1481, August 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
009-9409-7;   http://link.springer.com/content/pdf/10.1007/
s11538-009-9409-7.pdf.

Bayram:2009:TMD

[2389] Saziye Bayram, Tracy L. Stepien, and E. Bruce Pitman. TGF-Mediated dynamics in a system of many coupled nephrons. Bulletin of Mathematical Biology, 71(6):1482–1506, August 2009. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-009-9410-1;
 http://link.springer.com/content/pdf/10.1007/s11538-009-9410-
1.pdf.

Haegeman:2009:NMG

[2390] Bart Haegeman and Rampal S. Etienne. Neutral models with generalised speciation. Bulletin of Mathematical Biology, 71(6):1507–1519, August 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
009-9411-0;   http://link.springer.com/content/pdf/10.1007/
s11538-009-9411-0.pdf.

Pozrikidis:2009:NSB

[2391] C. Pozrikidis. Numerical simulation of blood flow through microvascular capillary networks. Bulletin of Mathematical Biology, 71(6):1520– 1541, August 2009. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-009-9412-z;   http://link.springer.com/content/
pdf/10.1007/s11538-009-9412-z.pdf.

Cooper:2009:CAL

[2392] Zachary Cooper, Michael Greenwood, and Borbala Mazzag. A computational analysis of localized ca^{2+}-dynamics generated by heterogeneous release sites. Bulletin of Mathematical Biology, 71(7):1543–1579, October 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
009-9413-y;   http://link.springer.com/content/pdf/10.1007/
s11538-009-9413-y.pdf.

Grosberg:2009:EFG

[2393] Anna Grosberg, Morteza Gharib, and Arash Kheradvar. Effect of fiber geometry on pulsatile pumping and energy expenditure. Bulletin of Mathematical Biology, 71(7):1580–1598, October 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
009-9414-x;   http://link.springer.com/content/pdf/10.1007/
s11538-009-9414-x.pdf.

Gadgil:2009:SID

[2394] Chetan J. Gadgil. Size-independent differences between the mean of discrete stochastic systems and the corresponding continuous deterministic systems. Bulletin of Mathematical Biology, 71(7):1599–1611, October 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
009-9415-9;   http://link.springer.com/content/pdf/10.1007/
s11538-009-9415-9.pdf.

Carrillo:2009:DCS

[2395] Claudia Carrillo, Kirill Cherednichenko, Nicholas Britton, and Michael Mogie. Dynamic coexistence of sexual and asexual invasion fronts in a system of integro–difference equations. Bulletin of Mathematical Biology, 71(7):1612–1625, October 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-009-9416-8;   http://link.springer.com/
content/pdf/10.1007/s11538-009-9416-8.pdf.

Haario:2009:RMA

[2396] Heikki Haario, Leonid Kalachev, and Marko Laine. Reduced models of algae growth. Bulletin of Mathematical Biology, 71(7):1626–1648, October 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
009-9417-7;   http://link.springer.com/content/pdf/10.1007/
s11538-009-9417-7.pdf.

Lee:2009:ICO

[2397] Ha Youn Lee, Edwin Hawkins, Martin S. Zand, Tim Mosmann, Hulin Wu, Philip D. Hodgkin, and Alan S. Perelson. Interpreting CFSE obtained division histories of b cells in vitro with Smith–Martin and cyton type models. Bulletin of Mathematical Biology, 71(7):1649–1670, October 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
009-9418-6;   http://link.springer.com/content/pdf/10.1007/
s11538-009-9418-6.pdf.

Hensel:2009:SKM

[2398] Sebastian C. Hensel, James B. Rawlings, and John Yin. Stochastic kinetic modeling of vesicular stomatitis virus intracellular growth. Bulletin of Mathematical Biology, 71(7):1671–1692, October 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
009-9419-5;   http://link.springer.com/content/pdf/10.1007/
s11538-009-9419-5.pdf.

House:2009:MBA

[2399] Thomas House, Geoffrey Davies, Leon Danon, and Matt J. Keeling. A motif-based approach to network epidemics. Bulletin of Mathematical Biology, 71(7):1693–1706, October 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-009-9420-z;
 http://link.springer.com/content/pdf/10.1007/s11538-009-9420-
z.pdf.

Hand:2009:DMM

[2400] Paul E. Hand, Boyce E. Griffith, and Charles S. Peskin. Deriving macroscopic myocardial conductivities by homogenization of microscopic models. Bulletin of Mathematical Biology, 71(7):1707–1726, October 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-

009-9421-y;   http://link.springer.com/content/pdf/10.1007/
s11538-009-9421-y.pdf.

Mendez:2009:MPI

[2401] Vicenç Méndez, Daniel Campos, and Andy W. Sheppard. A model for plant invasions: the role of distributed generation times. Bulletin of Mathematical Biology, 71(7):1727–1744, October 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
009-9422-x;   http://link.springer.com/content/pdf/10.1007/
s11538-009-9422-x.pdf.

Bhunu:2009:MHA

[2402] C. P. Bhunu, W. Garira, and Z. Mukandavire. Modeling HIV/ AIDS and tuberculosis coinfection. Bulletin of Mathematical Biology, 71(7): 1745–1780, October 2009. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-009-9423-9;   http://link.springer.com/content/
pdf/10.1007/s11538-009-9423-9.pdf.

Bacaer:2009:PMP

[2403] Nicolas Bacaër. Periodic matrix population models: Growth rate, basic reproduction number, and entropy. Bulletin of Mathematical Biology, 71(7):1781–1792, October 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-009-9426-6;   http://link.springer.com/
content/pdf/10.1007/s11538-009-9426-6.pdf.

Hambrock:2009:ECD

[2404] R. Hambrock and Y. Lou. The evolution of conditional dispersal strategies in spatially heterogeneous habitats. Bulletin of Mathematical Biology, 71(8):1793–1817, November 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-009-9425-7;
 http://link.springer.com/content/pdf/10.1007/s11538-009-9425-
7.pdf.

Masuda:2009:TAT

[2405] Naoki Masuda and Hisashi Ohtsuki. A theoretical analysis of temporal difference learning in the iterated Prisoner’s dilemma game. Bulletin of Mathematical Biology, 71(8):1818–1850, November 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-

009-9424-8;   http://link.springer.com/content/pdf/10.1007/
s11538-009-9424-8.pdf.

Oyarzun:2009:SAM

[2406] Diego A. Oyarzún, Brian P. Ingalls, Richard H. Middleton, and Dimitrios Kalamatianos. Sequential activation of metabolic pathways: a dynamic optimization approach. Bulletin of Mathematical Biology, 71(8): 1851–1872, November 2009. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-009-9427-5;   http://link.springer.com/content/
pdf/10.1007/s11538-009-9427-5.pdf.

Tian:2009:MPC

[2407] Jianjun Paul Tian and Xiao-Song Lin. The mutation process in colored coalescent theory. Bulletin of Mathematical Biology, 71(8):1873– 1889, November 2009. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-009-9428-4;   http://link.springer.com/content/
pdf/10.1007/s11538-009-9428-4.pdf.

Capistran:2009:PES

[2408] Marcos A. Capistrán, Miguel A. Moreles, and Bruno Lara. Parameter estimation of some epidemic models. The case of recurrent epidemics caused by respiratory syncytial virus. Bulletin of Mathematical Biology, 71(8):1890–1901, November 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
009-9429-3;   http://link.springer.com/content/pdf/10.1007/
s11538-009-9429-3.pdf.

Camitz:2009:ETD

[2409] Martin Camitz and Åke Svensson. The effect of time distribution shape on a complex epidemic model. Bulletin of Mathematical Biology, 71(8): 1902–1913, November 2009. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-009-9430-x;   http://link.springer.com/content/
pdf/10.1007/s11538-009-9430-x.pdf.

Sit:2009:GBA

[2410] Atilla Sit, Zhijun Wu, and Yaxiang Yuan. A geometric buildup algorithm for the solution of the distance geometry problem using least–squares approximation. Bulletin of Mathematical Biology, 71(8):1914–1933, November 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
009-9431-9;   http://link.springer.com/content/pdf/10.1007/
s11538-009-9431-9.pdf.

Mistro:2009:RSR

[2411] Diomar Cristina Mistro, Luiz Alberto Díaz Rodrigues, and Maria Cristina Varriale. The role of spatial refuges in coupled map lattice model for host–parasitoid systems. Bulletin of Mathematical Biology, 71(8):1934– 1953, November 2009. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-009-9432-8;   http://link.springer.com/content/
pdf/10.1007/s11538-009-9432-8.pdf.

Bacaer:2009:FSE

[2412] Nicolas Bacaër and M. Gabriela M. Gomes. On the final size of epidemics with seasonality. Bulletin of Mathematical Biology, 71(8):1954– 1966, November 2009. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-009-9433-7;   http://link.springer.com/content/
pdf/10.1007/s11538-009-9433-7.pdf.

Ying:2009:TSA

[2413] Wenxia Ying, Gabriel Huerta, Stanly Steinberg, and Martha Zúñiga. Time series analysis of particle tracking data for molecular motion on the cell membrane. Bulletin of Mathematical Biology, 71(8):1967– 2024, November 2009. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-009-9434-6;   http://link.springer.com/content/
pdf/10.1007/s11538-009-9434-6.pdf.

Tang:2009:EPP

[2414] Sanyi Tang, Robert A. Cheke, and Yanni Xiao. Effects of predator and prey dispersal on success or failure of biological control. Bulletin of Mathematical Biology, 71(8):2025–2047, November 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
009-9438-2;   http://link.springer.com/content/pdf/10.1007/
s11538-009-9438-2.pdf.

Li:2009:MSS

[2415] Jing Li and Xingfu Zou. Modeling spatial spread of infectious diseases with a fixed latent period in a spatially continuous domain. Bulletin of Mathematical Biology, 71(8):2048–2079, November 2009. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
009-9457-z;   http://link.springer.com/content/pdf/10.1007/
s11538-009-9457-z.pdf.

Qiu:2010:TDI

[2416] Zhipeng Qiu and Zhilan Feng. Transmission dynamics of an influenza model with vaccination and antiviral treatment. Bulletin of Mathematical Biology, 72(1):1–33, January 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-009-9435-5;
 http://link.springer.com/content/pdf/10.1007/s11538-009-9435-
5.pdf.

Layton:2010:TFS

[2417] Anita T. Layton and Aurélie Edwards. Tubuloglomerular feedback signal transduction in a short loop of Henle. Bulletin of Mathematical Biology, 72(1):34–62, January 2010. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-009-9436-4;   http://link.springer.com/content/
pdf/10.1007/s11538-009-9436-4.pdf.

Teboh-Ewungkem:2010:MSR

[2418] Miranda I. Teboh-Ewungkem, Chandra N. Podder, and Abba B. Gumel. Mathematical study of the role of gametocytes and an imperfect vaccine on malaria transmission dynamics. Bulletin of Mathematical Biology, 72 (1):63–93, January 2010. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-009-9437-3;   http://link.springer.com/content/
pdf/10.1007/s11538-009-9437-3.pdf.

Berec:2010:IFF

[2419] Luděk Berec. Impacts of foraging facilitation among predators on predator-prey dynamics. Bulletin of Mathematical Biology, 72(1):94– 121, January 2010. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-009-9439-1;   http://link.springer.com/content/
pdf/10.1007/s11538-009-9439-1.pdf.

Hsieh:2010:CGA

[2420] Ying-Hen Hsieh, Cathy W. S. Chen, Shu-Fang Hsu Schmitz, Chwan-Chuan King, Wei-Ju Chen, Yi-Chun Wu, and Mei-Shang Ho. Candidate genes associated with susceptibility for SARS-coronavirus.

Bulletin of Mathematical Biology, 72(1):122–132, January 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
009-9440-8;   http://link.springer.com/content/pdf/10.1007/
s11538-009-9440-8.pdf. See erratum [2447].

Noori:2010:NOB

[2421] Hamid Reza Noori and Willi Jäger. Neurochemical oscillations in the basal ganglia. Bulletin of Mathematical Biology, 72(1):133–147, January 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
009-9441-7;   http://link.springer.com/content/pdf/10.1007/
s11538-009-9441-7.pdf.

Gyrya:2010:MHI

[2422] Vitaliy Gyrya, Igor S. Aranson, Leonid V. Berlyand, and Dmitry Karpeev. A model of hydrodynamic interaction between swimming bacteria. Bulletin of Mathematical Biology, 72(1):148–183, January 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
009-9442-6;   http://link.springer.com/content/pdf/10.1007/
s11538-009-9442-6.pdf.

Eskola:2010:AEM

[2423] Hanna T. M. Eskola and Kalle Parvinen. The Allee effect in mechanistic models based on inter-individual interaction processes. Bulletin of Mathematical Biology, 72(1):184–207, January 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
009-9443-5;   http://link.springer.com/content/pdf/10.1007/
s11538-009-9443-5.pdf.

Fischer:2010:REB

[2424] Mareike Fischer and Bhalchandra Thatte. Revisiting an equivalence between maximum parsimony and maximum likelihood methods in phylogenetics. Bulletin of Mathematical Biology, 72(1):208–220, January 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
009-9446-2;   http://link.springer.com/content/pdf/10.1007/
s11538-009-9446-2.pdf.

Johnson:2010:TCF

[2425] Stewart D. Johnson. Theoretical considerations of Fisher’s theorem in small at-risk populations. Bulletin of Mathematical Biology, 72(1): 221–229, January 2010. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-009-9455-1;   http://link.springer.com/content/
pdf/10.1007/s11538-009-9455-1.pdf.

Kohler:2010:LSI

[2426] Brynja R. Kohler, Rebecca J. Swank, James W. Haefner, and James A. Powell. Leading students to investigate diffusion as a model of brine shrimp movement. Bulletin of Mathematical Biology, 72(1):230–257, January 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
009-9444-4;   http://link.springer.com/content/pdf/10.1007/
s11538-009-9444-4.pdf.

deRoos:2010:NEA

[2427] A. M. de Roos, O. Diekmann, P. Getto, and M. A. Kirkilionis. Numerical equilibrium analysis for structured consumer resource models. Bulletin of Mathematical Biology, 72(2):259–297, February 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
009-9445-3;   http://link.springer.com/content/pdf/10.1007/
s11538-009-9445-3.pdf. See erratum [3120].

Norwich:2010:MEM

[2428] Kenneth H. Norwich. A mathematical exploration of the mystery of loudness adaptation. Bulletin of Mathematical Biology, 72(2):298–313, February 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
009-9447-1;   http://link.springer.com/content/pdf/10.1007/
s11538-009-9447-1.pdf.

Marcano:2010:MUC

[2429] Mariano Marcano, Anita T. Layton, and Harold E. Layton. Maximum urine concentrating capability in a mathematical model of the inner medulla of the Rat Kidney. Bulletin of Mathematical Biology, 72(2): 314–339, February 2010. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-009-9448-0;   http://link.springer.com/content/
pdf/10.1007/s11538-009-9448-0.pdf.

Willson:2010:PNP

[2430] Stephen J. Willson. Properties of normal phylogenetic networks. Bulletin of Mathematical Biology, 72(2):340–358, February 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
009-9449-z;   http://link.springer.com/content/pdf/10.1007/
s11538-009-9449-z.pdf.

Sachs:2010:RMA

[2431] Rainer K. Sachs and Lynn Hlatky. A rapid-mutation approximation for cell population dynamics. Bulletin of Mathematical Biology, 72(2): 359–374, February 2010. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-009-9450-6;   http://link.springer.com/content/
pdf/10.1007/s11538-009-9450-6.pdf.

Jin:2010:IRS

[2432] Emma Y. Jin and Christian M. Reidys. Irreducibility in RNA structures. Bulletin of Mathematical Biology, 72(2):375–399, February 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
009-9451-5;   http://link.springer.com/content/pdf/10.1007/
s11538-009-9451-5.pdf.

Moreo:2010:MBP

[2433] P. Moreo, E. A. Gaffney, J. M. García-Aznar, and M. Doblaré. On the modelling of biological patterns with mechanochemical models: Insights from analysis and computation. Bulletin of Mathematical Biology, 72(2): 400–431, February 2010. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-009-9452-4;   http://link.springer.com/content/
pdf/10.1007/s11538-009-9452-4.pdf.

Mendez:2010:ECP

[2434] Vicenç Méndez, Daniel Campos, Isaac Llopis, and Werner Horsthemke. Extinction and chaotic patterns in map lattices under hostile conditions. Bulletin of Mathematical Biology, 72(2):432–443, February 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
009-9453-3;   http://link.springer.com/content/pdf/10.1007/
s11538-009-9453-3.pdf.

Oliveira:2010:MDI

[2435] Nuno M. Oliveira and Frank M. Hilker. Modelling disease introduction as biological control of invasive predators to preserve endangered prey. Bulletin of Mathematical Biology, 72(2):444–468, February 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
009-9454-2;   http://link.springer.com/content/pdf/10.1007/
s11538-009-9454-2.pdf.

Biesecker:2010:OVC

[2436] Matt Biesecker, Jung-Han Kimn, Huitian Lu, David Dingli, and Željko Bajzer. Optimization of virotherapy for cancer. Bulletin of Mathematical Biology, 72(2):469–489, February 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-009-9456-0;
 http://link.springer.com/content/pdf/10.1007/s11538-009-9456-
0.pdf.

Blyuss:2010:SBE

[2437] Konstantin B. Blyuss and Yuliya N. Kyrychko. Stability and bifurcations in an epidemic model with varying immunity period. Bulletin of Mathematical Biology, 72(2):490–505, February 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
009-9458-y;   http://link.springer.com/content/pdf/10.1007/
s11538-009-9458-y.pdf.

Manning:2010:MLB

[2438] Cammey C. Manning, Paul M. Schlosser, and Hien T. Tran. A multicompartment liver-based pharmacokinetic model for benzene and its metabolites in mice. Bulletin of Mathematical Biology, 72(3):507–540, April 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
009-9459-x;   http://link.springer.com/content/pdf/10.1007/
s11538-009-9459-x.pdf.

Gedeon:2010:HTS

[2439] Tomáš Gedeon, Clara Bodelón, and Amy Kuenzi. Hantavirus transmission in Sylvan and peridomestic environments. Bulletin of Mathematical Biology, 72(3):541–564, April 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-009-9460-4;   http://link.springer.com/
content/pdf/10.1007/s11538-009-9460-4.pdf.

Lewis:2010:STM

[2440] Mark A. Lewis, William Nelson, and Cailin Xu. A structured threshold model for mountain pine beetle outbreak. Bulletin of Mathematical Biology, 72(3):565–589, April 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-009-9461-3;   http://link.springer.com/
content/pdf/10.1007/s11538-009-9461-3.pdf.

Mitchell:2010:OIH

[2441] Jonathan L. Mitchell and Thomas W. Carr. Oscillations in an intrahost model of Plasmodium falciparum malaria due to cross-reactive immune response. Bulletin of Mathematical Biology, 72(3):590–610, April 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
009-9462-2;   http://link.springer.com/content/pdf/10.1007/
s11538-009-9462-2.pdf.

Kim:2010:EGD

[2442] Peter S. Kim, Peter P. Lee, and Doron Levy. Emergent group dynamics governed by regulatory cells produce a robust primary T cell response. Bulletin of Mathematical Biology, 72(3):611–644, April 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
009-9463-1;   http://link.springer.com/content/pdf/10.1007/
s11538-009-9463-1.pdf.

Schamberg:2010:MAP

[2443] S. Schamberg, P. Houston, N. A. M. Monk, and M. R. Owen. Modelling and analysis of planar cell polarity. Bulletin of Mathematical Biology, 72 (3):645–680, April 2010. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-009-9464-0;   http://link.springer.com/content/
pdf/10.1007/s11538-009-9464-0.pdf.

Gomez-Acevedo:2010:MMC

[2444] Horacio Gómez-Acevedo, Michael Y. Li, and Steven Jacobson. Multistability in a model for CTL response to HTLV-I infection and its implications to HAM/TSP development and prevention. Bulletin of Mathematical Biology, 72(3):681–696, April 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
009-9465-z;   http://link.springer.com/content/pdf/10.1007/
s11538-009-9465-z.pdf.

Brown:2010:PPP

[2445] Martin Brown, Fei He, and Stephen J. Wilkinson. Properties of the proximate parameter tuning regularization algorithm. Bulletin of Mathematical Biology, 72(3):697–718, April 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-009-9466-y;
 http://link.springer.com/content/pdf/10.1007/s11538-009-9466-
y.pdf.

Baker:2010:MMD

[2446] Ruth E. Baker, Christian A. Yates, and Radek Erban. From microscopic to macroscopic descriptions of cell migration on growing domains. Bulletin of Mathematical Biology, 72(3):719–762, April 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
009-9467-x;   http://link.springer.com/content/pdf/10.1007/
s11538-009-9467-x.pdf.

Hsieh:2010:ECG

[2447] Ying-Hen Hsieh, Cathy W. S. Chen, Shu-Fang Hsu Schmitz, Chwan-Chuan King, Wei-Ju Chen, Yi-Chun Wu, and Mei-Shang Ho. Erratum to: Candidate Genes Associated with Susceptibility for SARS-Coronavirus. Bulletin of Mathematical Biology, 72(3):763, April 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9512-9;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9512-9.pdf. See [2420].

Herald:2010:GMI

[2448] Meagan C. Herald. General model of inflammation. Bulletin of Mathematical Biology, 72(4):765–779, May 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/s11538-009-9468-9;   http://link.springer.
com/content/pdf/10.1007/s11538-009-9468-9.pdf.

Bani-Yaghoub:2010:DNA

[2449] Majid Bani-Yaghoub and David E. Amundsen. Dynamics of notch activity in a model of interacting signaling pathways. Bulletin of Mathematical Biology, 72(4):780–804, May 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-009-9469-8;   http://link.springer.com/
content/pdf/10.1007/s11538-009-9469-8.pdf.

Lei:2010:MMF

[2450] Jinzhi Lei and You Song. Mathematical model of the formation of morphogen gradients through membrane-associated nonreceptors. Bulletin of Mathematical Biology, 72(4):805–829, May 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
009-9470-2;   http://link.springer.com/content/pdf/10.1007/
s11538-009-9470-2.pdf.

Jackson:2010:CBM

[2451] Trachette Jackson and Xiaoming Zheng. A cell-based model of endothelial cell migration, proliferation and maturation during corneal angiogenesis. Bulletin of Mathematical Biology, 72(4):830–868, May 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
009-9471-1;   http://link.springer.com/content/pdf/10.1007/
s11538-009-9471-1.pdf.

Diotallevi:2010:RGM

[2452] F. Diotallevi, B. M. Mulder, and J. Grasman. On the robustness of the geometrical model for cell wall deposition. Bulletin of Mathematical Biology, 72(4):869–895, May 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-009-9472-0;
 http://link.springer.com/content/pdf/10.1007/s11538-009-9472-
0.pdf.

James:2010:EIA

[2453] Alex James, Jonathan W. Pitchford, and Michael J. Plank. Efficient or inaccurate? Analytical and numerical modelling of random search strategies. Bulletin of Mathematical Biology, 72(4):896–913, May 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
009-9473-z;   http://link.springer.com/content/pdf/10.1007/
s11538-009-9473-z.pdf.

Dembele:2010:MMO

[2454] Bassidy Dembele, Avner Friedman, and Abdul-Aziz Yakubu. Mathematical model for optimal use of sulfadoxine-pyrimethamine as a temporary malaria vaccine. Bulletin of Mathematical Biology, 72(4):914–930, May 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-

009-9476-9;   http://link.springer.com/content/pdf/10.1007/
s11538-009-9476-9.pdf.

Liu:2010:TMS

[2455] Luju Liu, Xiao-Qiang Zhao, and Yicang Zhou. A tuberculosis model with seasonality. Bulletin of Mathematical Biology, 72(4):931–952, May 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
009-9477-8;   http://link.springer.com/content/pdf/10.1007/
s11538-009-9477-8.pdf.

Crowder:2010:SGC

[2456] Tanner Crowder and Chi-Kwong Li. Studying genetic code by a matrix approach. Bulletin of Mathematical Biology, 72(4):953–972, May 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
009-9478-7;   http://link.springer.com/content/pdf/10.1007/
s11538-009-9478-7.pdf.

Michelin:2010:LTD

[2457] Sébastien Michelin and Eric Lauga. The long-time dynamics of two hydrodynamically-coupled swimming cells. Bulletin of Mathematical Biology, 72(4):973–1005, May 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-009-9479-6;   http://link.springer.com/
content/pdf/10.1007/s11538-009-9479-6.pdf.

Blayneh:2010:BBO

[2458] Kbenesh W. Blayneh, Abba B. Gumel, Suzanne Lenhart, and Tim Clayton. Backward bifurcation and optimal control in transmission dynamics of West Nile virus. Bulletin of Mathematical Biology, 72(4):1006–1028, May 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
009-9480-0;   http://link.springer.com/content/pdf/10.1007/
s11538-009-9480-0.pdf.

Kim:2010:ITM

[2459] Yangjin Kim and Avner Friedman. Interaction of tumor with its microenvironment: A Mathematical model. Bulletin of Mathematical Biology, 72(5):1029–1068, July 2010. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-009-9481-z;   http://link.springer.com/content/
pdf/10.1007/s11538-009-9481-z.pdf.

Bertuzzi:2010:RTS

[2460] A. Bertuzzi, C. Bruni, A. Fasano, A. Gandolfi, F. Papa, and C. Sinisgalli. Response of tumor spheroids to radiation: Modeling and parameter estimation. Bulletin of Mathematical Biology, 72(5):1069–1091, July 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
009-9482-y;   http://link.springer.com/content/pdf/10.1007/
s11538-009-9482-y.pdf.

Walton:2010:OSE

[2461] D. Brian Walton, Benjamin P. Leard, and Anthony Tongen. An optimal strategy for energy allocation in a multiple resource environment. Bulletin of Mathematical Biology, 72(5):1092–1123, July 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
009-9483-x;   http://link.springer.com/content/pdf/10.1007/
s11538-009-9483-x.pdf.

Karev:2010:REP

[2462] G. P. Karev. Replicator equations and the principle of minimal production of information. Bulletin of Mathematical Biology, 72(5):1124–1142, July 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
009-9484-9;   http://link.springer.com/content/pdf/10.1007/
s11538-009-9484-9.pdf.

Vaughan:2010:IFF

[2463] Benjamin L. Vaughan, Jr., Bryan G. Smith, and David L. Chopp. The influence of fluid flow on modeling quorum sensing in bacterial biofilms. Bulletin of Mathematical Biology, 72(5):1143–1165, July 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
009-9485-8;   http://link.springer.com/content/pdf/10.1007/
s11538-009-9485-8.pdf.

Hamel:2010:SSS

[2464] François Hamel, Julien Fayard, and Lionel Roques. Spreading speeds in slowly oscillating environments. Bulletin of Mathematical Biology, 72 (5):1166–1191, July 2010. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-009-9486-7;   http://link.springer.com/content/
pdf/10.1007/s11538-009-9486-7.pdf.

Huang:2010:GSD

[2465] Gang Huang, Yasuhiro Takeuchi, Wanbiao Ma, and Daijun Wei. Global stability for delay SIR and SEIR epidemic models with nonlinear incidence rate. Bulletin of Mathematical Biology, 72(5):1192–1207, July 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
009-9487-6;   http://link.springer.com/content/pdf/10.1007/
s11538-009-9487-6.pdf.

Wasserstein-Robbins:2010:MMH

[2466] Freda Wasserstein-Robbins. A mathematical model of HIV infection: Simulating T4, T8, macrophages, antibody, and virus via specific anti-HIV response in the presence of adaptation and tropism. Bulletin of Mathematical Biology, 72(5):1208–1253, July 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
009-9488-5;   http://link.springer.com/content/pdf/10.1007/
s11538-009-9488-5.pdf.

Kolpas:2010:EDS

[2467] Allison Kolpas and Roger M. Nisbet. Effects of demographic stochasticity on Population Persistence in advective media. Bulletin of Mathematical Biology, 72(5):1254–1270, July 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-009-9489-4;
 http://link.springer.com/content/pdf/10.1007/s11538-009-9489-
4.pdf.

Gurney:2010:MDS

[2468] William S. C. Gurney, Eddie McKenzie, and Philip J. Bacon. Maturity dispersion, stock auto-correlation, and management strategy in exploited populations. Bulletin of Mathematical Biology, 72(5):1271–1293, July 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
009-9490-y;   http://link.springer.com/content/pdf/10.1007/
s11538-009-9490-y.pdf.

Amaku:2010:MDV

[2469] Marcos Amaku, Marcelo Nascimento Burattini, Francisco Antonio Bezerra Coutinho, and Eduardo Massad. Modeling the dynamics of viral evolution Considering Competition within individual hosts and at population level: The effects of treatment. Bulletin of Mathematical Biology, 72(5):1294–1314, July 2010. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-009-9495-6;   http://link.springer.com/content/
pdf/10.1007/s11538-009-9495-6.pdf.

Jackson:2010:HS

[2470] J. H. Jackson and C. R. MacCluer. Hyperbolic saturation. Bulletin of Mathematical Biology, 72(5):1315–1322, July 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-009-9475-x;
 http://link.springer.com/content/pdf/10.1007/s11538-009-9475-
x.pdf.

Larkin:2010:MMN

[2471] Timothy J. Larkin and Philip W. Kuchel. Mathematical models of naturally “morphed” human erythrocytes: Stomatocytes and echinocytes. Bulletin of Mathematical Biology, 72(6):1323–1333, August 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
009-9493-8;   http://link.springer.com/content/pdf/10.1007/
s11538-009-9493-8.pdf.

Yurk:2010:MED

[2472] Brian P. Yurk and James A. Powell. Modeling the effects of developmental variation on insect phenology. Bulletin of Mathematical Biology, 72(6):1334–1360, August 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-009-9494-7;   http://link.springer.com/
content/pdf/10.1007/s11538-009-9494-7.pdf.

Datta:2010:JGM

[2473] Samik Datta, Gustav W. Delius, and Richard Law. A jump-growth model for predator–prey dynamics: Derivation and application to marine ecosystems. Bulletin of Mathematical Biology, 72(6):1361–1382, August 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
009-9496-5;   http://link.springer.com/content/pdf/10.1007/
s11538-009-9496-5.pdf.

Middleton:2010:MMA

[2474] A. M. Middleton, J. R. King, M. J. Bennett, and M. R. Owen. Mathematical modelling of the Aux/IAA negative feedback loop. Bulletin of Mathematical Biology, 72(6):1383–1407, August 2010.

CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
009-9497-4;   http://link.springer.com/content/pdf/10.1007/
s11538-009-9497-4.pdf.

Hand:2010:HEM

[2475] Paul E. Hand and Charles S. Peskin. Homogenization of an electrophysiological model for a strand of cardiac myocytes with gap-junctional and electric-field coupling. Bulletin of Mathematical Biology, 72(6): 1408–1424, August 2010. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-009-9499-2;   http://link.springer.com/content/
pdf/10.1007/s11538-009-9499-2.pdf.

Jarrah:2010:DCD

[2476] Abdul Salam Jarrah, Reinhard Laubenbacher, and Alan Veliz-Cuba. The dynamics of conjunctive and disjunctive Boolean network models. Bulletin of Mathematical Biology, 72(6):1425–1447, August 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9501-z;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9501-z.pdf.

Shiu:2010:SCR

[2477] Anne Shiu and Bernd Sturmfels. Siphons in chemical reaction networks. Bulletin of Mathematical Biology, 72(6):1448–1463, August 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9502-y;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9502-y.pdf.

Shipley:2010:MMF

[2478] Rebecca J. Shipley and S. Jonathan Chapman. Multiscale modelling of fluid and drug transport in vascular tumours. Bulletin of Mathematical Biology, 72(6):1464–1491, August 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-010-9504-9;
 http://link.springer.com/content/pdf/10.1007/s11538-010-9504-
9.pdf.

Li:2010:GDH

[2479] Michael Y. Li and Hongying Shu. Global dynamics of an in-host viral model with intracellular delay. Bulletin of Mathematical Biology, 72(6):

1492–1505, August 2010. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-010-9503-x;   http://link.springer.com/content/
pdf/10.1007/s11538-010-9503-x.pdf.

Tien:2010:MTP

[2480] Joseph H. Tien and David J. D. Earn. Multiple transmission pathways and disease dynamics in a waterborne pathogen model. Bulletin of Mathematical Biology, 72(6):1506–1533, August 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9507-6;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9507-6.pdf.

Ougrinovskaia:2010:OME

[2481] Anna Ougrinovskaia, Rosemary S. Thompson, and Mary R. Myerscough. An ODE model of early stages of atherosclerosis: Mechanisms of the inflammatory response. Bulletin of Mathematical Biology, 72(6): 1534–1561, August 2010. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-010-9509-4;   http://link.springer.com/content/
pdf/10.1007/s11538-010-9509-4.pdf.

Wei:2010:DGB

[2482] Guo-Wei Wei. Differential geometry based multiscale models. Bulletin of Mathematical Biology, 72(6):1562–1622, August 2010. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-010-9511-x;
 http://link.springer.com/content/pdf/10.1007/s11538-010-9511-
x.pdf.

Seo:2010:MAM

[2483] Yeon-Jung Seo, Shiliang Chen, Marit Nilsen-Hamilton, and Howard A. Levine. A mathematical analysis of multiple-target SELEX. Bulletin of Mathematical Biology, 72(7):1623–1665, October 2010. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-009-9491-x;
 http://link.springer.com/content/pdf/10.1007/s11538-009-9491-
x.pdf.

Terry:2010:PCI

[2484] Alan J. Terry and Stephen A. Gourley. Perverse consequences of infrequently culling a pest. Bulletin of Mathematical Biology, 72(7):1666– 1695, October 2010. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-009-9492-9;   http://link.springer.com/content/
pdf/10.1007/s11538-009-9492-9.pdf.

Bock:2010:GVT

[2485] Martin Bock, Amit Kumar Tyagi, Jan-Ulrich Kreft, and Wolfgang Alt. Generalized Voronoi tessellation as a model of two-dimensional cell tissue dynamics. Bulletin of Mathematical Biology, 72(7):1696–1731, October 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
009-9498-3;   http://link.springer.com/content/pdf/10.1007/
s11538-009-9498-3.pdf.

Doumic-Jauffret:2010:SAS

[2486] Marie Doumic-Jauffret, Peter S. Kim, and Benoît Perthame. Stability analysis of a simplified yet complete model for chronic myelogenous leukemia. Bulletin of Mathematical Biology, 72(7):1732–1759, October 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
009-9500-0;   http://link.springer.com/content/pdf/10.1007/
s11538-009-9500-0.pdf.

Zhang:2010:AFM

[2487] Louxin Zhang, Jian Shen, Jialiang Yang, and Guoliang Li. Analyzing the Fitch method for reconstructing ancestral states on ultrametric phylogenetic trees. Bulletin of Mathematical Biology, 72(7):1760–1782, October 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9505-8;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9505-8.pdf.

vanIersel:2010:QEL

[2488] Leo van Iersel, Charles Semple, and Mike Steel. Quantifying the extent of lateral gene transfer required to avert a ‘Genome of Eden’. Bulletin of Mathematical Biology, 72(7):1783–1798, October 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9506-7;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9506-7.pdf.

Luke:2010:DQM

[2489] Nicholas S. Luke, Michael J. DeVito, Imran Shah, and Hisham A. El-Masri. Development of a quantitative model of pregnane X receptor (PXR) mediated xenobiotic metabolizing enzyme induction. Bulletin of Mathematical Biology, 72(7):1799–1819, October 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9508-5;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9508-5.pdf.

Pardi:2010:RPI

[2490] Fabio Pardi, Sylvain Guillemot, and Olivier Gascuel. Robustness of phylogenetic inference based on minimum evolution. Bulletin of Mathematical Biology, 72(7):1820–1839, October 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-010-9510-y;
 http://link.springer.com/content/pdf/10.1007/s11538-010-9510-
y.pdf.

Newby:2010:QSS

[2491] Jay M. Newby and Paul C. Bressloff. Quasi-steady state reduction of molecular motor-based models of directed intermittent search. Bulletin of Mathematical Biology, 72(7):1840–1866, October 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9513-8;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9513-8.pdf.

Flegg:2010:MMH

[2492] Jennifer A. Flegg, Helen M. Byrne, and D. L. Sean McElwain. Mathematical model of hyperbaric oxygen therapy applied to chronic diabetic wounds. Bulletin of Mathematical Biology, 72(7):1867–1891, October 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9514-7;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9514-7.pdf.

Xiao:2010:IHM

[2493] Dongmei Xiao and William H. Bossert. An intra-host mathematical model on interaction between HIV and malaria. Bulletin of Mathematical Biology, 72(7):1892–1911, October 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-010-9515-6;
 http://link.springer.com/content/pdf/10.1007/s11538-010-9515-
6.pdf.

Melkikh:2010:MAT

[2494] A. V. Melkikh and D. S. Bessarab. Model of active transport of ions through diatom cell biomembrane. Bulletin of Mathematical Biology, 72(7):1912–1924, October 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-010-9520-9;   http://link.springer.com/
content/pdf/10.1007/s11538-010-9520-9.pdf.

Olson:2010:MCC

[2495] Sarah D. Olson, Susan S. Suarez, and Lisa J. Fauci. A model of CatSper channel mediated calcium dynamics in mammalian spermatozoa. Bulletin of Mathematical Biology, 72(8):1925–1946, November 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9516-5;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9516-5.pdf.

Anderson:2010:PFS

[2496] David F. Anderson, Gheorghe Craciun, and Thomas G. Kurtz. Productform stationary distributions for deficiency zero chemical reaction networks. Bulletin of Mathematical Biology, 72(8):1947–1970, November 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9517-4;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9517-4.pdf.

Seymour:2010:CGD

[2497] Robert M. Seymour, Gwenan Knight, and Tak Fung. Conditions for global dynamic stability of a class of resource-bounded model ecosystems. Bulletin of Mathematical Biology, 72(8):1971–2003, November 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9518-3;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9518-3.pdf.

MillerNeilan:2010:MOI

[2498] Rachael L. Miller Neilan, Elsa Schaefer, Holly Gaff, K. Renee Fister, and Suzanne Lenhart. Modeling optimal intervention strategies for cholera. Bulletin of Mathematical Biology, 72(8):2004–2018, November 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9521-8;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9521-8.pdf.

Ramakrishnan:2010:SST

[2499] Rajasekhar Ramakrishnan and Janak D. Ramakrishnan. A state space transformation can yield identifiable models for tracer kinetic studies with enrichment data. Bulletin of Mathematical Biology, 72(8):2019– 2046, November 2010. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-010-9522-7;   http://link.springer.com/content/
pdf/10.1007/s11538-010-9522-7.pdf.

Ghosal:2010:NWC

[2500] Sandip Ghosal and Zhen Chen. Nonlinear waves in capillary electrophoresis. Bulletin of Mathematical Biology, 72(8):2047–2066, November 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9527-2;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9527-2.pdf.

Farkas:2010:SUC

[2501] József Z. Farkas and Peter Hinow. Structured and unstructured continuous models for Wolbachia infections. Bulletin of Mathematical Biology, 72(8):2067–2088, November 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-010-9528-1;   http://link.springer.com/
content/pdf/10.1007/s11538-010-9528-1.pdf.

Samia:2010:CSC

[2502] Yasmine Samia and Frithjof Lutscher. Coexistence and spread of competitors in heterogeneous landscapes. Bulletin of Mathematical Biology, 72(8):2089–2112, November 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-010-9529-0;   http://link.springer.com/
content/pdf/10.1007/s11538-010-9529-0.pdf.

Nundloll:2010:III

[2503] Sapna Nundloll, Ludovic Mailleret, and Frédéric Grognard. Influence of intrapredatory interferences on impulsive biological control efficiency. Bulletin of Mathematical Biology, 72(8):2113–2138, November 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9531-6;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9531-6.pdf.

Lee:2010:IGE

[2504] S. Seirin Lee, E. A. Gaffney, and N. A. M. Monk. The influence of gene expression time delays on Gierer–Meinhardt pattern formation systems. Bulletin of Mathematical Biology, 72(8):2139–2160, November 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9532-5;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9532-5.pdf.

SeirinLee:2010:ABR

[2505] S. Seirin Lee and E. A. Gaffney. Aberrant behaviours of reaction diffusion self-organisation models on growing domains in the presence of gene expression time delays. Bulletin of Mathematical Biology, 72(8): 2161–2179, November 2010. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-010-9533-4;   http://link.springer.com/content/
pdf/10.1007/s11538-010-9533-4.pdf.

Bacaer:2010:ASM

[2506] Nicolas Bacaër, Carel Pretorius, and Bertran Auvert. An agestructured model for the potential impact of generalized access to antiretrovirals on the South African HIV epidemic. Bulletin of Mathematical Biology, 72(8):2180–2198, November 2010. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-010-9535-2;
 http://link.springer.com/content/pdf/10.1007/s11538-010-9535-
2.pdf.

Maini:2011:E

[2507] Philip K. Maini. Editorial. Bulletin of Mathematical Biology, 73 (1):1, January 2011. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-011-9627-7;   http://link.springer.com/content/
pdf/10.1007/s11538-011-9627-7.pdf.

Eftimie:2011:IBI

[2508] Raluca Eftimie, Jonathan L. Bramson, and David J. D. Earn. Interactions between the immune system and cancer: A Brief review of non-spatial mathematical models. Bulletin of Mathematical Biology, 73(1):2–32, January 2011. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-010-9526-3;   http://link.springer.com/content/
pdf/10.1007/s11538-010-9526-3.pdf.

Alexander:2011:STA

[2509] H. K. Alexander and L. M. Wahl. Self-tolerance and autoimmunity in a regulatory T cell model. Bulletin of Mathematical Biology, 73(1): 33–71, January 2011. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-010-9519-2;   http://link.springer.com/content/
pdf/10.1007/s11538-010-9519-2.pdf.

Simitev:2011:ACV

[2510] R. D. Simitev and V. N. Biktashev. Asymptotics of conduction velocity restitution in models of electrical excitation in the heart. Bulletin of Mathematical Biology, 73(1):72–115, January 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9523-6;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9523-6.pdf.

Banks:2011:ECP

[2511] H. T. Banks, Karyn L. Sutton, W. Clayton Thompson, Gennady Bocharov, Dirk Roose, Tim Schenkel, and Andreas Meyerhans. Estimation of cell proliferation dynamics using CFSE data. Bulletin of Mathematical Biology, 73(1):116–150, January 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9524-5;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9524-5.pdf.

Musso:2011:SVE

[2512] Fabio Musso. A stochastic version of the Eigen model. Bulletin of Mathematical Biology, 73(1):151–180, January 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-010-9525-4;
 http://link.springer.com/content/pdf/10.1007/s11538-010-9525-
4.pdf.

Jabbari:2011:MMS

[2513] Sara Jabbari, John T. Heap, and John R. King. Mathematical modelling of the sporulation-initiation network in Bacillus subtilis revealing the dual role of the putative quorum-sensing signal molecule PhrA. Bulletin of Mathematical Biology, 73(1):181–211, January 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9530-7;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9530-7.pdf.

Cogan:2011:TDP

[2514] N. G. Cogan and C. W. Wolgemuth. Two-dimensional patterns in bacterial veils arise from self-generated, three-dimensional fluid flows. Bulletin of Mathematical Biology, 73(1):212–229, January 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9536-1;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9536-1.pdf.

Das:2011:MES

[2515] Raibatak Das, Robert B. Nachbar, Leah Edelstein-Keshet, Jeffrey S. Saltzman, Matthew C. Wiener, Ansuman Bagchi, James Bailey, Daniel Coombs, Adam J. Simon, Richard J. Hargreaves, and Jacquelynn J. Cook. Modeling effect of a γ-secretase inhibitor on amyloidβ dynamics reveals significant role of an amyloid clearance mechanism. Bulletin of Mathematical Biology, 73(1):230–247, January 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9540-5;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9540-5.pdf.

Bianco:2011:APM

[2516] Simone Bianco and Leah B. Shaw. Asymmetry in the presence of migration stabilizes multistrain disease outbreaks. Bulletin of Mathematical Biology, 73(1):248–260, January 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-010-9541-4;
 http://link.springer.com/content/pdf/10.1007/s11538-010-9541-
4.pdf.

beimGraben:2011:MPN

[2517] Peter beim Graben and James Wright. From McCulloch–Pitts neurons toward biology. Bulletin of Mathematical Biology, 73(2):261–265, February 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
011-9629-5;   http://link.springer.com/content/pdf/10.1007/
s11538-011-9629-5.pdf.

Afraimovich:2011:NDE

[2518] Valentin Afraimovich, Todd Young, Mehmet K. Muezzinoğlu, and Mikhail I. Rabinovich. Nonlinear dynamics of emotion-cognition interaction: When emotion does not destroy cognition? Bulletin of Mathematical Biology, 73(2):266–284, February 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-010-9572-x;
 http://link.springer.com/content/pdf/10.1007/s11538-010-9572-
x.pdf.

Galka:2011:DNM

[2519] Andreas Galka, Kin Foon Kevin Wong, Tohru Ozaki, Hiltrud Muhle, Ulrich Stephani, and Michael Siniatchkin. Decomposition of neurological multivariate time series by state space modelling. Bulletin of Mathematical Biology, 73(2):285–324, February 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9563-y;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9563-y.pdf.

Jirsa:2011:NPM

[2520] Viktor K. Jirsa and Roxana A. Stefanescu. Neural population modes capture biologically realistic large scale network dynamics. Bulletin of Mathematical Biology, 73(2):325–343, February 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9573-9;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9573-9.pdf.

Kay:2011:CIC

[2521] Jim W. Kay and W. A. Phillips. Coherent Infomax as a computational goal for neural systems. Bulletin of Mathematical Biology, 73(2): 344–372, February 2011. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-010-9564-x;   http://link.springer.com/content/
pdf/10.1007/s11538-010-9564-x.pdf.

Mizraji:2011:LDB

[2522] Eduardo Mizraji and Juan Lin. Logic in a dynamic brain. Bulletin of Mathematical Biology, 73(2):373–397, February 2011. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-010-9561-0;
 http://link.springer.com/content/pdf/10.1007/s11538-010-9561-
0.pdf.

Steyn-Ross:2011:MUS

[2523] Moira L. Steyn-Ross, D. A. Steyn-Ross, J. W. Sleigh, and M. T. Wilson. A mechanism for ultra-slow oscillations in the cortical default network. Bulletin of Mathematical Biology, 73(2):398–416, February 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9565-9;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9565-9.pdf.

Tsukada:2011:CSM

[2524] Minoru Tsukada and Yasuhiro Fukushima. A context-sensitive mechanism in hippocampal CA1 networks. Bulletin of Mathematical Biology, 73 (2):417–435, February 2011. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-010-9566-8;   http://link.springer.com/content/
pdf/10.1007/s11538-010-9566-8.pdf.

Wright:2011:ADT

[2525] J. J. Wright. Attractor dynamics and thermodynamic analogies in the cerebral cortex: Synchronous oscillation, the background EEG, and the regulation of attention. Bulletin of Mathematical Biology, 73(2): 436–457, February 2011. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-010-9562-z;   http://link.springer.com/content/
pdf/10.1007/s11538-010-9562-z.pdf. See erratum [2526].

Wright:2011:EAD

[2526] J. J. Wright. Erratum to: Attractor Dynamics and Thermodynamic Analogies in the Cerebral Cortex: Synchronous Oscillation, the Background EEG, and the Regulation of Attention. Bulletin of Mathematical Biology, 73(2):458, February 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-011-9639-3;
 http://link.springer.com/content/pdf/10.1007/s11538-011-9639-

3.pdf. See [2525].

Xie:2011:SFS

[2527] Xiaohui Xie. The site-frequency spectrum of linked sites. Bulletin of Mathematical Biology, 73(3):459–494, March 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-010-9534-3;
 http://link.springer.com/content/pdf/10.1007/s11538-010-9534-
3.pdf.

Forgoston:2011:MSD

[2528] Eric Forgoston, Simone Bianco, Leah B. Shaw, and Ira B. Schwartz. Maximal sensitive dependence and the optimal path to epidemic extinction. Bulletin of Mathematical Biology, 73(3):495–514, March 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9537-0;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9537-0.pdf.

Sharomi:2011:MTD

[2529] O. Sharomi, C. N. Podder, A. B. Gumel, S. M. Mahmud, and E. Rubinstein. Modelling the transmission dynamics and control of the novel 2009 swine influenza (H1N1) pandemic. Bulletin of Mathematical Biology, 73 (3):515–548, March 2011. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-010-9538-z;   http://link.springer.com/content/
pdf/10.1007/s11538-010-9538-z.pdf.

Katouli:2011:WDR

[2530] Allen A. Katouli and Natalia L. Komarova. The worst drug rule revisited: Mathematical modeling of cyclic cancer treatments. Bulletin of Mathematical Biology, 73(3):549–584, March 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-010-9539-y;
 http://link.springer.com/content/pdf/10.1007/s11538-010-9539-
y.pdf.

Roe-Dale:2011:MMB

[2531] Rachel Roe-Dale, David Isaacson, and Michael Kupferschmid. A mathematical model of breast cancer treatment with CMF and doxorubicin. Bulletin of Mathematical Biology, 73(3):585–608, March 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9549-9;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9549-9.pdf.

Souza:2011:GSC

[2532] Max O. Souza and Jorge P. Zubelli. Global stability for a class of virus models with cytotoxic T lymphocyte immune response and antigenic variation. Bulletin of Mathematical Biology, 73(3):609–625, March 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9543-2;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9543-2.pdf.

Wilhelm:2011:GTA

[2533] M. Wilhelm, M. Chhetri, J. Rychtář, and O. Rueppell. A game theoretical analysis of the mating sign behavior in the honey bee. Bulletin of Mathematical Biology, 73(3):626–638, March 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9544-1;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9544-1.pdf.

Chamchod:2011:AVB

[2534] Farida Chamchod and Nicholas F. Britton. Analysis of a vector-bias model on malaria transmission. Bulletin of Mathematical Biology, 73 (3):639–657, March 2011. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-010-9545-0;   http://link.springer.com/content/
pdf/10.1007/s11538-010-9545-0.pdf.

Coskun:2011:CPR

[2535] Hasan Coskun and Huseyin Coskun. Cell physician: Reading cell motion. Bulletin of Mathematical Biology, 73(3):658–682, March 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9580-x;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9580-x.pdf.

Broom:2011:KMM

[2536] M. Broom and J. Rychtář. Kleptoparasitic melees — modelling food stealing featuring contests with multiple individuals. Bulletin of Mathematical Biology, 73(3):683–699, March 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-010-9546-z;
 http://link.springer.com/content/pdf/10.1007/s11538-010-9546-
z.pdf.

Laubenbacher:2011:AMM

[2537] Reinhard Laubenbacher. Algebraic methods in mathematical biology. Bulletin of Mathematical Biology, 73(4):701–705, April 2011. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-011-9643-7;
 https://link.springer.com/content/pdf/10.1007/s11538-011-9643-
7.pdf.

Blochl:2011:EPD

[2538] Florian Blöchl, Dominik M. Wittmann, and Fabian J. Theis. Effective parameters determining the information flow in hierarchical biological systems. Bulletin of Mathematical Biology, 73(4):706–725, April 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-010-9604-6;   https://link.springer.com/content/pdf/
10.1007/s11538-010-9604-6.pdf.

Bona:2011:EVC

[2539] Miklós Bóna, Meera Sitharam, and Andrew Vince. Enumeration of viral capsid assembly pathways: Tree orbits under permutation group action. Bulletin of Mathematical Biology, 73(4):726–753, April 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-010-9606-4;   https://link.springer.com/content/pdf/
10.1007/s11538-010-9606-4.pdf.

Hower:2011:PAR

[2540] Valerie Hower and Christine E. Heitsch. Parametric analysis of RNA branching configurations. Bulletin of Mathematical Biology, 73(4):754– 776, April 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-010-9607-3;   https://link.springer.com/content/pdf/
10.1007/s11538-010-9607-3.pdf.

McCaig:2011:SIS

[2541] Chris McCaig, Mike Begon, Rachel Norman, and Carron Shankland. A symbolic investigation of superspreaders. Bulletin of Mathematical Biology, 73(4):777–794, April 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-010-9603-7;   https://link.springer.
com/content/pdf/10.1007/s11538-010-9603-7.pdf.

Malaspinas:2011:PAA

[2542] Anna-Sapfo Malaspinas, Nicholas Eriksson, and Peter Huggins. Parametric analysis of alignment and phylogenetic uncertainty. Bulletin of Mathematical Biology, 73(4):795–810, April 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-010-9610-8;
 https://link.springer.com/content/pdf/10.1007/s11538-010-9610-
8.pdf.

Dickenstein:2011:HFC

[2543] Alicia Dickenstein and Mercedes Pérez Millán. How far is complex balancing from detailed balancing? Bulletin of Mathematical Biology, 73(4):811–828, April 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-010-9611-7;   https://link.springer.
com/content/pdf/10.1007/s11538-010-9611-7.pdf.

Sainudiin:2011:ESF

[2544] Raazesh Sainudiin, Kevin Thornton, Jennifer Harlow, James Booth, Michael Stillman, Ruriko Yoshida, Robert Griffiths, Gil McVean, and Peter Donnelly. Experiments with the site frequency spectrum. Bulletin of Mathematical Biology, 73(4):829–872, April 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-010-9605-5;
 https://link.springer.com/content/pdf/10.1007/s11538-010-9605-
5.pdf.

Siebert:2011:ADB

[2545] Heike Siebert. Analysis of discrete bioregulatory networks using symbolic steady states. Bulletin of Mathematical Biology, 73(4):873–898, April 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-010-9609-1;   https://link.springer.com/content/pdf/
10.1007/s11538-010-9609-1.pdf.

Weber:2011:AGC

[2546] Andreas Weber, Thomas Sturm, and Essam O. Abdel-Rahman. Algorithmic global criteria for excluding oscillations. Bulletin of Mathematical Biology, 73(4):899–916, April 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-010-9618-0;   https://link.springer.
com/content/pdf/10.1007/s11538-010-9618-0.pdf.

Venturino:2011:SMM

[2547] Ezio Venturino. Simple metaecoepidemic models. Bulletin of Mathematical Biology, 73(5):917–950, May 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-010-9542-3;   http://link.springer.com/
content/pdf/10.1007/s11538-010-9542-3.pdf.

Pfuhl:2011:PRA

[2548] G. Pfuhl, H. Tjelmeland, and R. Biegler. Precision and reliability in animal navigation. Bulletin of Mathematical Biology, 73(5):951–977, May 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9547-y;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9547-y.pdf.

Theis:2011:TAT

[2549] Fabian J. Theis, Sebastian Bohl, and Ursula Klingmüller. Theoretical analysis of time-to-peak responses in biological reaction networks. Bulletin of Mathematical Biology, 73(5):978–1003, May 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9548-x;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9548-x.pdf.

Christensen:2011:GCM

[2550] Asbjorn Christensen and Ken Haste Andersen. General classification of maturation reaction-norm shape from size-based processes. Bulletin of Mathematical Biology, 73(5):1004–1027, May 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9550-3;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9550-3.pdf.

Wild:2011:IFM

[2551] Geoff Wild. Inclusive fitness from multitype branching processes. Bulletin of Mathematical Biology, 73(5):1028–1051, May 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-010-9551-2;
 http://link.springer.com/content/pdf/10.1007/s11538-010-9551-
2.pdf.

Cobbold:2011:ESS

[2552] C. A. Cobbold and J. A. Powell. Evolution stabilises the synchronising dynamics of poikilotherm life cycles. Bulletin of Mathematical Biology, 73(5):1052–1081, May 2011. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-010-9552-1;   http://link.springer.com/content/
pdf/10.1007/s11538-010-9552-1.pdf.

Paquin:2011:STI

[2553] Dana Paquin, Peter S. Kim, Peter P. Lee, and Doron Levy. Strategic treatment interruptions during imatinib treatment of chronic myelogenous leukemia. Bulletin of Mathematical Biology, 73(5):1082–1100, May 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9553-0;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9553-0.pdf.

Lavi:2011:EVE

[2554] Orit Lavi, Eyal Klement, and Yoram Louzoun. Effect of vaccination in environmentally induced diseases. Bulletin of Mathematical Biology, 73 (5):1101–1117, May 2011. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-010-9554-z;   http://link.springer.com/content/
pdf/10.1007/s11538-010-9554-z.pdf.

Sanchez-Garduno:2011:SSP

[2555] Faustino Sánchez-Garduño and Víctor F. Breña-Medina. Searching for spatial patterns in a pollinator–plant–herbivore mathematical model. Bulletin of Mathematical Biology, 73(5):1118–1153, May 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9599-z;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9599-z.pdf.

Hoyle:2011:EBT

[2556] Andy Hoyle, Roger G. Bowers, and Andy White. Evolutionary behaviour, trade-offs and cyclic and chaotic population dynamics. Bulletin of Mathematical Biology, 73(5):1154–1169, May 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9567-7;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9567-7.pdf.

Aydemir:2011:EPL

[2557] E. Aydemir, C. J. W. Breward, and T. P. Witelski. The effect of polar lipids on tear film dynamics. Bulletin of Mathematical Biology, 73 (6):1171–1201, June 2011. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-010-9555-y;   http://link.springer.com/content/
pdf/10.1007/s11538-010-9555-y.pdf.

Cueto:2011:PGP

[2558] María Angélica Cueto and Frederick A. Matsen. Polyhedral geometry of phylogenetic rogue taxa. Bulletin of Mathematical Biology, 73 (6):1202–1226, June 2011. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-010-9556-x;   http://link.springer.com/content/
pdf/10.1007/s11538-010-9556-x.pdf.

Altenberg:2011:ERP

[2559] Lee Altenberg. An evolutionary reduction principle for mutation rates at multiple loci. Bulletin of Mathematical Biology, 73(6):1227–1270, June 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9557-9;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9557-9.pdf.

Drew:2011:MMT

[2560] Donald A. Drew, Gretchen A. Koch, Shannon Hitchcock, Jill Kowalski, Ronak Talati, and Vera Valakh. A mathematical model for timing the release from sequestration and the resultant Brownian migration of SeqA clusters in E. coli. Bulletin of Mathematical Biology, 73(6):1271–1291, June 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9558-8;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9558-8.pdf.

Teka:2011:PPP

[2561] Wondimu Teka, Krasimira Tsaneva-Atanasova, Richard Bertram, and Joël Tabak. From plateau to pseudo-plateau bursting: Making the transition. Bulletin of Mathematical Biology, 73(6):1292–1311, June 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9559-7;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9559-7.pdf.

Eskola:2011:ETR

[2562] Hanna T. M. Eskola, Stefan A. H. Geritz, and Mats Gyllenberg. On the evolution of the timing of reproduction with non-equilibrium resident dynamics. Bulletin of Mathematical Biology, 73(6):1312–1332, June 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9560-1;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9560-1.pdf.

Bhunu:2011:TAE

[2563] C. P. Bhunu, S. Mushayabasa, and J. M. Tchuenche. A theoretical assessment of the effects of smoking on the transmission dynamics of tuberculosis. Bulletin of Mathematical Biology, 73(6):1333–1357, June 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9568-6;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9568-6.pdf.

Gurarie:2011:EMO

[2564] Eliezer Gurarie, Daniel Grünbaum, and Michael T. Nishizaki. Estimating 3D movements from 2D observations using a continuous model of helical swimming. Bulletin of Mathematical Biology, 73(6):1358–1377, June 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9575-7;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9575-7.pdf.

Gonzalez-Olivares:2011:MLC

[2565] Eduardo González-Olivares and Alejandro Rojas-Palma. Multiple limit cycles in a Gause type predator–prey model with Holling Type III functional response and Allee effect on prey. Bulletin of Mathematical Biology, 73(6):1378–1397, June 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-010-9577-5;
 http://link.springer.com/content/pdf/10.1007/s11538-010-9577-
5.pdf.

Chai:2011:NBC

[2566] Juanjuan Chai and Elizabeth Ann Housworth. On the number of binary characters needed to recover a phylogeny using maximum parsimony. Bulletin of Mathematical Biology, 73(6):1398–1411, June 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9579-3;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9579-3.pdf.

Kribs-Zaleta:2011:MNT

[2567] Christopher M. Kribs-Zaleta, Jean-François Jusot, Philippe Vanhems, and Sandrine Charles. Modeling nosocomial transmission of rotavirus in pediatric wards. Bulletin of Mathematical Biology, 73(7):1413–1442, July 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9570-z;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9570-z.pdf.

Jose:2011:GHS

[2568] Marco V. José, Eberto R. Morgado, and Tzipe Govezensky. Genetic hotels for the standard genetic code: Evolutionary analysis based upon novel three-dimensional algebraic models. Bulletin of Mathematical Biology, 73(7):1443–1476, July 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-010-9571-y;
 http://link.springer.com/content/pdf/10.1007/s11538-010-9571-
y.pdf.

Didier:2011:PMP

[2569] Gilles Didier. Parametric maximum parsimonious reconstruction on trees. Bulletin of Mathematical Biology, 73(7):1477–1502, July 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9574-8;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9574-8.pdf.

Schley:2011:MFM

[2570] David Schley, John Ward, and Zhidong Zhang. Modelling foot-and-mouth disease virus dynamics in oral epithelium to help identify the determinants of lysis. Bulletin of Mathematical Biology, 73(7):1503–1528, July 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9576-6;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9576-6.pdf.

Bloomfield:2011:HDC

[2571] J. M. Bloomfield, K. J. Painter, and J. A. Sherratt. How does cellular contact affect differentiation mediated pattern formation? Bulletin of Mathematical Biology, 73(7):1529–1558, July 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-010-9578-4;
 http://link.springer.com/content/pdf/10.1007/s11538-010-9578-
4.pdf.

Gotzen:2011:CDP

[2572] Bernd Gotzen, Volkmar Liebscher, and Sebastian Walcher. On a class of deterministic population models with stochastic foundation. Bulletin of Mathematical Biology, 73(7):1559–1582, July 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9581-9;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9581-9.pdf.

Hinkelmann:2011:MFA

[2573] Franziska Hinkelmann, David Murrugarra, Abdul Salam Jarrah, and Reinhard Laubenbacher. A mathematical framework for agent based models of complex biological networks. Bulletin of Mathematical Biology, 73(7):1583–1602, July 2011. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-010-9582-8;   http://link.springer.com/content/
pdf/10.1007/s11538-010-9582-8.pdf.

Dallon:2011:USD

[2574] J. C. Dallon, Brittany Dalton, and Chelsea Malani. Understanding streaming in Dictyostelium discoideum: Theory versus experiments. Bulletin of Mathematical Biology, 73(7):1603–1626, July 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9583-7;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9583-7.pdf.

Mossel:2011:ILP

[2575] Elchanan Mossel, Sébastien Roch, and Allan Sly. On the inference of large phylogenies with long branches: How long is too long? Bulletin of Mathematical Biology, 73(7):1627–1644, July 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-

010-9584-6;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9584-6.pdf.

Kim:2011:TIA

[2576] Peter S. Kim, Peter P. Lee, and Doron Levy. A theory of immunodominance and adaptive regulation. Bulletin of Mathematical Biology, 73 (7):1645–1665, July 2011. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-010-9585-5;   http://link.springer.com/content/
pdf/10.1007/s11538-010-9585-5.pdf.

Bjorklund:2011:EFL

[2577] M. Björklund, E. Ranta, V. Kaitala, L. A. Bach, and P. Lundberg. Environmental fluctuations and level of density-compensation strongly affects the Probability of Fixation and fixation times. Bulletin of Mathematical Biology, 73(7):1666–1681, July 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-010-9587-3;
 http://link.springer.com/content/pdf/10.1007/s11538-010-9587-
3.pdf.

Huang:2011:RDC

[2578] Huaxiong Huang, Jonathan J. Wylie, and Robert M. Miura. Restricted diffusion in cellular media: (1 + 1)-dimensional model. Bulletin of Mathematical Biology, 73(7):1682–1694, July 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-010-9589-1;
 http://link.springer.com/content/pdf/10.1007/s11538-010-9589-
1.pdf.

Xue:2011:TWH

[2579] Chuan Xue, Hyung Ju Hwang, Kevin J. Painter, and Radek Erban. Travelling waves in hyperbolic chemotaxis equations. Bulletin of Mathematical Biology, 73(8):1695–1733, August 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-010-9586-4;
 http://link.springer.com/content/pdf/10.1007/s11538-010-9586-
4.pdf.

Saccomani:2011:EAP

[2580] Maria Pia Saccomani. An effective automatic procedure for testing parameter identifiability of HIV/AIDS models. Bulletin of Mathematical Biology, 73(8):1734–1753, August 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-010-9588-2;
 http://link.springer.com/content/pdf/10.1007/s11538-010-9588-
2.pdf.

Huynh:2011:AHC

[2581] Giao T. Huynh and Frederick R. Adler. Alternating host cell tropism shapes the persistence, evolution and coexistence of Epstein–Barr virus infections in human. Bulletin of Mathematical Biology, 73(8):1754– 1773, August 2011. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-010-9590-8;   http://link.springer.com/content/
pdf/10.1007/s11538-010-9590-8.pdf.

Li:2011:MSP

[2582] Michael Y. Li and Hongying Shu. Multiple stable periodic oscillations in a mathematical model of CTL response to HTLV-I infection. Bulletin of Mathematical Biology, 73(8):1774–1793, August 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9591-7;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9591-7.pdf.

Liu:2011:SVM

[2583] Rongsong Liu, Carlos Martinez del Rio, and Jianhong Wu. Spatiotemporal variation of mistletoes: a dynamic modeling approach. Bulletin of Mathematical Biology, 73(8):1794–1811, August 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9592-6;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9592-6.pdf.

Rodrigues:2011:PFL

[2584] Luiz Alberto Díaz Rodrigues, Diomar Cristina Mistro, and Sergei Petrovskii. Pattern formation, long-term transients, and the Turing– Hopf bifurcation in a space- and time-discrete predator–prey system. Bulletin of Mathematical Biology, 73(8):1812–1840, August 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9593-5;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9593-5.pdf.

Sapsis:2011:IPD

[2585] Themistoklis Sapsis, Jifeng Peng, and George Haller. Instabilities on prey dynamics in jellyfish feeding. Bulletin of Mathematical Biology, 73(8): 1841–1856, August 2011. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-010-9594-4;   http://link.springer.com/content/
pdf/10.1007/s11538-010-9594-4.pdf.

Davis:2011:NSU

[2586] J. M. Davis and C. Pozrikidis. Numerical simulation of unsteady blood flow through capillary networks. Bulletin of Mathematical Biology, 73(8): 1857–1880, August 2011. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-010-9595-3;   http://link.springer.com/content/
pdf/10.1007/s11538-010-9595-3.pdf.

Delgado-Eckert:2011:EVR

[2587] Edgar Delgado-Eckert, Samuel Ojosnegros, and Niko Beerenwinkel. The evolution of virulence in RNA viruses under a competition-colonization trade-off. Bulletin of Mathematical Biology, 73(8):1881–1908, August 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9596-2;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9596-2.pdf.

Goor:2011:MAA

[2588] Robert M. Goor, Lisa Forman Neall, Douglas Hoffman, and Stephen T. Sherry. A mathematical approach to the analysis of multiplex DNA profiles. Bulletin of Mathematical Biology, 73(8):1909–1931, August 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9598-0;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9598-0.pdf.

Ben-Ari:2011:SAM

[2589] Iddo Ben-Ari, Khalid Boushaba, Anastasios Matzavinos, and Alexander Roitershtein. Stochastic analysis of the motion of DNA nanomechanical bipeds. Bulletin of Mathematical Biology, 73(8):1932–1951, August 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9600-x;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9600-x.pdf.

Sen:2011:BAA

[2590] S. Murat Sen, E. Zerrin Bagci, and Mehmet C. Camurdan. Bistability analysis of an apoptosis model in the presence of nitric oxide. Bulletin of Mathematical Biology, 73(8):1952–1968, August 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9613-5;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9613-5.pdf.

Liu:2011:SAS

[2591] Meng Liu, Ke Wang, and Qiong Wu. Survival analysis of stochastic competitive models in a polluted environment and stochastic competitive exclusion principle. Bulletin of Mathematical Biology, 73(9):1969– 2012, September 2011. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-010-9569-5;   http://link.springer.com/content/
pdf/10.1007/s11538-010-9569-5.pdf.

Gorban:2011:LMP

[2592] Alexander N. Gorban, Lyudmila I. Pokidysheva, Elena V. Smirnova, and Tatiana A. Tyukina. Law of the minimum paradoxes. Bulletin of Mathematical Biology, 73(9):2013–2044, September 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9597-1;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9597-1.pdf.

Salinas:2011:LME

[2593] Dino G. Salinas, Juan G. Reyes, and Milton De La Fuente. Lipid metabolizing enzyme activities modulated by phospholipid substrate lateral distribution. Bulletin of Mathematical Biology, 73(9):2045–2067, September 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9602-8;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9602-8.pdf.

Hilbe:2011:LRD

[2594] Christian Hilbe. Local replicator dynamics: a simple link between deterministic and stochastic models of evolutionary game theory. Bulletin of Mathematical Biology, 73(9):2068–2087, September 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-

010-9608-2;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9608-2.pdf.

Garlick:2011:HLS

[2595] Martha J. Garlick, James A. Powell, Mevin B. Hooten, and Leslie R. McFarlane. Homogenization of large-scale movement models in ecology. Bulletin of Mathematical Biology, 73(9):2088–2108, September 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9612-6;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9612-6.pdf.

Bennett:2011:MNR

[2596] Max R. Bennett, Les Farnell, and William G. Gibson. A model of NMDA receptor control of F-actin treadmilling in synaptic spines and their growth. Bulletin of Mathematical Biology, 73(9):2109–2131, September 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9614-4;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9614-4.pdf.

Michaels:2011:DSP

[2597] Jay L. Michaels, Vincent Naudot, and Larry S. Liebovitch. Dynamic stabilization in the PU1-GATA1 circuit Using a model with timedependent kinetic change. Bulletin of Mathematical Biology, 73(9):2132– 2151, September 2011. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-010-9615-3;   http://link.springer.com/content/
pdf/10.1007/s11538-010-9615-3.pdf.

Schumaker:2011:CMC

[2598] Mark F. Schumaker and David M. Kramer. Comparison of Monte Carlo simulations of cytochrome b_{6} f with experiment using Latin hypercube sampling. Bulletin of Mathematical Biology, 73(9):2152–2174, September 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9616-2;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9616-2.pdf.

Schnepf:2011:MNU

[2599] Andrea Schnepf, Davey Jones, and Tiina Roose. Modelling nutrient uptake by individual hyphae of arbuscular mycorrhizal fungi:

Temporal and spatial scales for an experimental design. Bulletin of Mathematical Biology, 73(9):2175–2200, September 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9617-1;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9617-1.pdf.

Cartwright:2011:BPC

[2600] Reed A. Cartwright. Bards, poets, and cliques: Frequency-dependent selection and the evolution of language genes. Bulletin of Mathematical Biology, 73(9):2201–2212, September 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-010-9619-z;   http://link.springer.com/
content/pdf/10.1007/s11538-010-9619-z.pdf.

Klapper:2011:EPI

[2601] Isaac Klapper and Barbara Szomolay. An exclusion principle and the importance of mobility for a class of biofilm models. Bulletin of Mathematical Biology, 73(9):2213–2230, September 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9621-5;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9621-5.pdf.

Miekisz:2011:SMG

[2602] Jacek Miekisz, Jan Poleszczuk, Marek Bodnar, and Urszula Foryś. Stochastic models of gene expression with delayed degradation. Bulletin of Mathematical Biology, 73(9):2231–2247, September 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9622-4;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9622-4.pdf.

Srinivasu:2011:RQA

[2603] P. D. N. Srinivasu and B. S. R. V. Prasad. Role of quantity of additional food to predators as a control in predator–prey systems with relevance to pest management and biological conservation. Bulletin of Mathematical Biology, 73(10):2249–2276, October 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9601-9;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9601-9.pdf.

Mincheva:2011:OBR

[2604] Maya Mincheva. Oscillations in biochemical reaction networks arising from pairs of subnetworks. Bulletin of Mathematical Biology, 73(10): 2277–2304, October 2011. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-010-9620-6;   http://link.springer.com/content/
pdf/10.1007/s11538-010-9620-6.pdf.

Andreasen:2011:FSE

[2605] Viggo Andreasen. The final size of an epidemic and its relation to the Basic Reproduction Number. Bulletin of Mathematical Biology, 73(10): 2305–2321, October 2011. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-010-9623-3;   http://link.springer.com/content/
pdf/10.1007/s11538-010-9623-3.pdf.

Willson:2011:RTS

[2606] Stephen J. Willson. Restricted trees: Simplifying networks with bottlenecks. Bulletin of Mathematical Biology, 73(10):2322–2338, October 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9624-2;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9624-2.pdf.

Bewick:2011:RXC

[2607] Sharon Bewick, Jing Wu, Scott C. Lenaghan, Ruoting Yang, Mingjun Zhang, and William Hamel. The R5 to X4 coreceptor switch: a dead-end path, or a strategic maneuver? Bulletin of Mathematical Biology, 73(10): 2339–2356, October 2011. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-010-9625-1;   http://link.springer.com/content/
pdf/10.1007/s11538-010-9625-1.pdf.

Jones:2011:BBF

[2608] Don A. Jones and Hal L. Smith. Bacteriophage and bacteria in a flow reactor. Bulletin of Mathematical Biology, 73(10):2357–2383, October 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
010-9626-0;   http://link.springer.com/content/pdf/10.1007/
s11538-010-9626-0.pdf.

Lou:2011:MMC

[2609] Yijun Lou and Xiao-Qiang Zhao. Modelling malaria control by introduction of larvivorous fish. Bulletin of Mathematical Biology, 73(10): 2384–2407, October 2011. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-011-9628-6;   http://link.springer.com/content/
pdf/10.1007/s11538-011-9628-6.pdf.

Jang:2011:FTT

[2610] Taesoo Jang, Hee-Dae Kwon, and Jeehyun Lee. Free terminal time optimal control problem of an HIV model based on a conjugate gradient method. Bulletin of Mathematical Biology, 73(10):2408–2429, October 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
011-9630-z;   http://link.springer.com/content/pdf/10.1007/
s11538-011-9630-z.pdf.

Aubert:2011:CMM

[2611] M. Aubert, M. A. J. Chaplain, S. R. McDougall, A. Devlin, and C. A. Mitchell. A continuum mathematical model of the developing murine retinal vasculature. Bulletin of Mathematical Biology, 73(10):2430– 2451, October 2011. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-011-9631-y;   http://link.springer.com/content/
pdf/10.1007/s11538-011-9631-y.pdf.

Krishnan:2011:RRP

[2612] Arjun Krishnan and Bogdan I. Epureanu. Renewal-reward process formulation of motor protein dynamics. Bulletin of Mathematical Biology, 73(10):2452–2482, October 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-011-9632-x;   http://link.springer.com/
content/pdf/10.1007/s11538-011-9632-x.pdf.

Gopalakrishnan:2011:FPT

[2613] Manoj Gopalakrishnan and Bindu S. Govindan. A first-passage-time theory for search and capture of chromosomes by microtubules in mitosis. Bulletin of Mathematical Biology, 73(10):2483–2506, October 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
011-9633-9;   http://link.springer.com/content/pdf/10.1007/
s11538-011-9633-9.pdf.

Postnov:2011:FMS

[2614] D. E. Postnov, J. C. Brings Jacobsen, N.-H. Holstein-Rathlou, and O. V. Sosnovtseva. Functional modeling of the shift in cellular calcium dynamics at the onset of synchronization in smooth muscle cells. Bulletin of Mathematical Biology, 73(10):2507–2525, October 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
011-9636-6;   http://link.springer.com/content/pdf/10.1007/
s11538-011-9636-6.pdf.

Lee:2011:DTP

[2615] S. Seirin Lee, E. A. Gaffney, and R. E. Baker. The dynamics of Turing patterns for morphogen-regulated growing domains with cellular response delays. Bulletin of Mathematical Biology, 73(11):2527– 2551, November 2011. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-011-9634-8;   http://link.springer.com/content/
pdf/10.1007/s11538-011-9634-8.pdf.

Chaffee:2011:ELI

[2616] J. Chaffee and R. Kuske. The effect of loss of immunity on noise-induced sustained oscillations in epidemics. Bulletin of Mathematical Biology, 73(11):2552–2574, November 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-011-9635-7;   http://link.springer.com/
content/pdf/10.1007/s11538-011-9635-7.pdf.

Childs:2011:IWH

[2617] Lauren M. Childs, Michael Paskow, Sidney M. Morris, Jr., Matthias Hesse, and Steven Strogatz. From inflammation to wound healing: Using a simple model to understand the functional versatility of murine macrophages. Bulletin of Mathematical Biology, 73(11):2575– 2604, November 2011. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-011-9637-5;   http://link.springer.com/content/
pdf/10.1007/s11538-011-9637-5.pdf.

Parvinen:2011:ADA

[2618] Kalle Parvinen. Adaptive dynamics of altruistic cooperation in a metapopulation: Evolutionary emergence of cooperators and defectors or evolutionary suicide? Bulletin of Mathematical Biology, 73(11):2605– 2626, November 2011. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-011-9638-4;   http://link.springer.com/content/
pdf/10.1007/s11538-011-9638-4.pdf.

Haws:2011:ONJ

[2619] David C. Haws, Terrell L. Hodge, and Ruriko Yoshida. Optimality of the neighbor joining algorithm and faces of the balanced minimum evolution polytope. Bulletin of Mathematical Biology, 73(11):2627–2648, November 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
011-9640-x;   http://link.springer.com/content/pdf/10.1007/
s11538-011-9640-x.pdf.

Lemaire:2011:MTI

[2620] T. Lemaire, E. Capiez-Lernout, J. Kaiser, S. Naili, E. Rohan, and V. Sansalone. A multiscale theoretical investigation of electric measurements in living bone. Bulletin of Mathematical Biology, 73(11): 2649–2677, November 2011. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-011-9641-9;   http://link.springer.com/content/
pdf/10.1007/s11538-011-9641-9.pdf.

Lang:2011:ASC

[2621] Moritz Lang, Tatiana T. Marquez-Lago, Jörg Stelling, and Steffen Waldherr. Autonomous synchronization of chemically coupled synthetic oscillators. Bulletin of Mathematical Biology, 73(11):2678–2706, November 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
011-9642-8;   http://link.springer.com/content/pdf/10.1007/
s11538-011-9642-8.pdf.

Moore:2011:SMB

[2622] S. M. Moore, C. A. Manore, V. A. Bokil, E. T. Borer, and P. R. Hosseini. Spatiotemporal model of barley and cereal yellow dwarf virus transmission dynamics with seasonality and plant competition. Bulletin of Mathematical Biology, 73(11):2707–2730, November 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
011-9654-4;   http://link.springer.com/content/pdf/10.1007/
s11538-011-9654-4.pdf.

Sotero:2011:BOL

[2623] Roberto C. Sotero and Yasser Iturria-Medina. From blood oxygenation level dependent (BOLD) signals to brain temperature maps. Bulletin of Mathematical Biology, 73(11):2731–2747, November 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
011-9645-5;   http://link.springer.com/content/pdf/10.1007/
s11538-011-9645-5.pdf.

Wells:2011:IIP

[2624] Chad R. Wells, Jean M. Tchuenche, Lauren Ancel Meyers, Alison P. Galvani, and Chris T. Bauch. Impact of imitation processes on the effectiveness of ring vaccination. Bulletin of Mathematical Biology, 73(11): 2748–2772, November 2011. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-011-9646-4;   http://link.springer.com/content/
pdf/10.1007/s11538-011-9646-4.pdf. See erratum [2916].

Yao:2011:CNM

[2625] Wei Yao, Huaxiong Huang, and Robert M. Miura. A continuum neuronal model for the instigation and propagation of cortical spreading depression. Bulletin of Mathematical Biology, 73(11):2773–2790, November 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
011-9647-3;   http://link.springer.com/content/pdf/10.1007/
s11538-011-9647-3.pdf.

Thibodeaux:2011:OTS

[2626] Jeremy J. Thibodeaux and Timothy P. Schlittenhardt. Optimal treatment strategies for malaria infection. Bulletin of Mathematical Biology, 73(11):2791–2808, November 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-011-9650-8;   http://link.springer.com/
content/pdf/10.1007/s11538-011-9650-8.pdf.

Sit:2011:SGD

[2627] Atilla Sit and Zhijun Wu. Solving a generalized distance geometry problem for protein structure determination. Bulletin of Mathematical Biology, 73(12):2809–2836, December 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-011-9644-6;
 http://link.springer.com/content/pdf/10.1007/s11538-011-9644-
6.pdf.

Pearson:2011:MMA

[2628] Yanthe E. Pearson, Emilio Castronovo, Tara A. Lindsley, and Donald A. Drew. Mathematical modeling of axonal formation. Part I: Geometry. Bulletin of Mathematical Biology, 73(12):2837–2864, December 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
011-9648-2;   http://link.springer.com/content/pdf/10.1007/
s11538-011-9648-2.pdf.

Eladdadi:2011:MME

[2629] Amina Eladdadi and David Isaacson. A mathematical model for the effects of HER2 over-expression on cell cycle progression in breast cancer. Bulletin of Mathematical Biology, 73(12):2865–2887, December 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
011-9663-3;   http://link.springer.com/content/pdf/10.1007/
s11538-011-9663-3.pdf.

Hove-Musekwa:2011:MHH

[2630] S. D. Hove-Musekwa, F. Nyabadza, and H. Mambili-Mamboundou. Modelling hospitalization, home-based care, and individual withdrawal for people living with HIV/AIDS in high prevalence settings. Bulletin of Mathematical Biology, 73(12):2888–2915, December 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
011-9651-7;   http://link.springer.com/content/pdf/10.1007/
s11538-011-9651-7.pdf.

Sanches:2011:RIS

[2631] Rosângela P. Sanches, Claudia P. Ferreira, and Roberto A. Kraenkel. The role of immunity and seasonality in cholera epidemics. Bulletin of Mathematical Biology, 73(12):2916–2931, December 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
011-9652-6;   http://link.springer.com/content/pdf/10.1007/
s11538-011-9652-6.pdf.

Eftimie:2011:MSM

[2632] Raluca Eftimie, Jonathan Dushoff, Byram W. Bridle, Jonathan L. Bramson, and David J. D. Earn. Multi-stability and multi-instability phenomena in a mathematical model of tumor–immune-virus interactions. Bulletin of Mathematical Biology, 73(12):2932–2961, December 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
011-9653-5;   http://link.springer.com/content/pdf/10.1007/
s11538-011-9653-5.pdf.

Gopalkrishnan:2011:CRN

[2633] Manoj Gopalkrishnan. Catalysis in reaction networks. Bulletin of Mathematical Biology, 73(12):2962–2982, December 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-011-9655-3;
 http://link.springer.com/content/pdf/10.1007/s11538-011-9655-
3.pdf.

Drobnjak:2011:MOB

[2634] Ivana Drobnjak and A. C. Fowler. A model of oscillatory blood cell counts in chronic myelogenous leukaemia. Bulletin of Mathematical Biology, 73(12):2983–3007, December 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-011-9656-2;   http://link.springer.com/
content/pdf/10.1007/s11538-011-9656-2.pdf.

Li:2011:MRT

[2635] Michael Y. Li and Aaron G. Lim. Modelling the role of tax expression in HTLV-I persistence in vivo. Bulletin of Mathematical Biology, 73(12): 3008–3029, December 2011. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-011-9657-1;   http://link.springer.com/content/
pdf/10.1007/s11538-011-9657-1.pdf.

Hu:2011:TAE

[2636] Guang Hu, Ze Wang, and Wen-Yuan Qiu. Topological analysis of enzymatic actions on DNA polyhedral links. Bulletin of Mathematical Biology, 73(12):3030–3046, December 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-011-9659-z;
 http://link.springer.com/content/pdf/10.1007/s11538-011-9659-
z.pdf.

Hiebeler:2011:MRR

[2637] David E. Hiebeler, Isaac J. Michaud, Hamilton Hoxie Ackerman, Shannon Reed Iosevich, and Andre Robinson. Multigeneration reproduction ratios and the effects of clustered unvaccinated individuals on epidemic outbreak. Bulletin of Mathematical Biology, 73(12):3047–3070, December 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
011-9660-6;   http://link.springer.com/content/pdf/10.1007/
s11538-011-9660-6.pdf.

Hand:2011:ESA

[2638] Paul E. Hand and Boyce E. Griffith. Empirical study of an adaptive multiscale model for simulating cardiac conduction. Bulletin of Mathematical Biology, 73(12):3071–3089, December 2011. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
011-9661-5;   http://link.springer.com/content/pdf/10.1007/
s11538-011-9661-5.pdf.

Kang:2012:RSE

[2639] Y. Kang and N. Lanchier. The role of space in the exploitation of resources. Bulletin of Mathematical Biology, 74(1):1–44, January 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
011-9649-1;   http://link.springer.com/content/pdf/10.1007/
s11538-011-9649-1.pdf.

Feliu:2012:AAS

[2640] Elisenda Feliu, Michael Knudsen, Lars N. Andersen, and Carsten Wiuf. An algebraic approach to signaling cascades with n layers. Bulletin of Mathematical Biology, 74(1):45–72, January 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
011-9658-0;   http://link.springer.com/content/pdf/10.1007/
s11538-011-9658-0.pdf.

Lee:2012:MAG

[2641] Yeonok Lee and Hulin Wu. MARS approach for global sensitivity analysis of differential equation models with applications to dynamics of influenza infection. Bulletin of Mathematical Biology, 74(1):73–90, January 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
011-9664-2;   http://link.springer.com/content/pdf/10.1007/
s11538-011-9664-2.pdf.

Galante:2012:BHM

[2642] Amanda Galante, Koji Tamada, and Doron Levy. B7–H1 and a mathematical model for cytotoxic T cell and tumor cell interaction. Bulletin of Mathematical Biology, 74(1):91–102, January 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
011-9665-1;   http://link.springer.com/content/pdf/10.1007/
s11538-011-9665-1.pdf.

Musso:2012:RBE

[2643] Fabio Musso. On the relation between the Eigen model and the asexual Wright–Fisher model. Bulletin of Mathematical Biology, 74(1): 103–115, January 2012. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-011-9666-0;   http://link.springer.com/content/
pdf/10.1007/s11538-011-9666-0.pdf.

Guiro:2012:SEH

[2644] Aboudramane Guiro, Abderrahman Iggidr, and Diène Ngom. On the stock estimation for a harvested fish population. Bulletin of Mathematical Biology, 74(1):116–142, January 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-011-9667-z;
 http://link.springer.com/content/pdf/10.1007/s11538-011-9667-
z.pdf.

Maggio:2012:MMM

[2645] Charles D. Maggio, Scott R. Jennings, Jennifer L. Robichaux, Peter C. Stapor, and James M. Hyman. A modified Hai–Murphy model of uterine smooth muscle contraction. Bulletin of Mathematical Biology, 74 (1):143–158, January 2012. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-011-9681-1;   http://link.springer.com/content/
pdf/10.1007/s11538-011-9681-1.pdf.

Roe-Dale:2012:MMC

[2646] Rachel Roe-Dale, David Isaacson, and Michael Kupferschmid. A mathematical model of cell cycle effects in gastric cancer chemotherapy. Bulletin of Mathematical Biology, 74(1):159–174, January 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
011-9669-x;   http://link.springer.com/content/pdf/10.1007/
s11538-011-9669-x.pdf.

Floyd:2012:CGA

[2647] William Floyd, Leslie Kay, and Michael Shapiro. A covering– graph approach to epidemics on SIS and SIS-like networks. Bulletin of Mathematical Biology, 74(1):175–189, January 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-011-9684-y;
 http://link.springer.com/content/pdf/10.1007/s11538-011-9684-
y.pdf.

Espinoza:2012:UHC

[2648] Flor A. Espinoza, Janet M. Oliver, Bridget S. Wilson, and Stanly L. Steinberg. Using hierarchical clustering and dendrograms to quantify the clustering of membrane proteins. Bulletin of Mathematical Biology, 74 (1):190–211, January 2012. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-011-9671-3;   http://link.springer.com/content/
pdf/10.1007/s11538-011-9671-3.pdf.

Rhodes:2012:ILP

[2649] John A. Rhodes and Seth Sullivant. Identifiability of large phylogenetic mixture models. Bulletin of Mathematical Biology, 74(1):212–231, January 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
011-9672-2;   http://link.springer.com/content/pdf/10.1007/
s11538-011-9672-2.pdf.

OMalley:2012:OSB

[2650] Stephen O’Malley and Martin A. Bees. The orientation of swimming biflagellates in shear flows. Bulletin of Mathematical Biology, 74(1): 232–255, January 2012. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-011-9673-1;   http://link.springer.com/content/
pdf/10.1007/s11538-011-9673-1.pdf.

Gejji:2012:ECI

[2651] Richard Gejji, Yuan Lou, Daniel Munther, and Justin Peyton. Evolutionary convergence to ideal free dispersal strategies and coexistence. Bulletin of Mathematical Biology, 74(2):257–299, February 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
011-9662-4;   http://link.springer.com/content/pdf/10.1007/
s11538-011-9662-4.pdf.

Miao:2012:EMM

[2652] Hongyu Miao, Xia Jin, Alan S. Perelson, and Hulin Wu. Evaluation of multitype mathematical models for CFSE-labeling experiment data. Bulletin of Mathematical Biology, 74(2):300–326, February 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
011-9668-y;   http://link.springer.com/content/pdf/10.1007/
s11538-011-9668-y.pdf.

Zhang:2012:EGC

[2653] Lai Zhang, Zhigui Lin, and Michael Pedersen. Effects of growth curve plasticity on size-structured population dynamics. Bulletin of Mathematical Biology, 74(2):327–345, February 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-011-9675-z;
 http://link.springer.com/content/pdf/10.1007/s11538-011-9675-
z.pdf.

Chen:2012:EDC

[2654] Zhen Chen and Sandip Ghosal. Electromigration dispersion in capillary electrophoresis. Bulletin of Mathematical Biology, 74(2):346–355, February 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
011-9708-7;   http://link.springer.com/content/pdf/10.1007/
s11538-011-9708-7.pdf.

Kelly:2012:ACP

[2655] W. P. Kelly and M. P. H. Stumpf. Assessing coverage of protein interaction data using capture–recapture models. Bulletin of Mathematical Biology, 74(2):356–374, February 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-011-9680-2;
 http://link.springer.com/content/pdf/10.1007/s11538-011-9680-
2.pdf.

Sun:2012:TPG

[2656] Qiwen Sun, Moxun Tang, and Jianshe Yu. Temporal profile of gene transcription noise modulated by cross-talking signal transduction pathways. Bulletin of Mathematical Biology, 74(2):375–398, February 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
011-9683-z;   http://link.springer.com/content/pdf/10.1007/
s11538-011-9683-z.pdf.

Fudenberg:2012:PSM

[2657] Drew Fudenberg and Lorens A. Imhof. Phenotype switching and mutations in random environments. Bulletin of Mathematical Biology, 74(2): 399–421, February 2012. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-011-9687-8;   http://link.springer.com/content/
pdf/10.1007/s11538-011-9687-8.pdf.

Layne:2012:NCD

[2658] Lori Layne, Elena Dimitrova, and Matthew Macauley. Nested canalyzing depth and network stability. Bulletin of Mathematical Biology, 74(2): 422–433, February 2012. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-011-9692-y;   http://link.springer.com/content/
pdf/10.1007/s11538-011-9692-y.pdf.

Hu:2012:SSA

[2659] Wen-Yong Hu, Wei-Rong Zhong, Feng-Hua Wang, Li Li, and Yuan-Zhi Shao. In silico synergism and antagonism of an anti-tumour system intervened by coupling immunotherapy and chemotherapy: a mathematical modelling approach. Bulletin of Mathematical Biology, 74(2): 434–452, February 2012. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-011-9693-x;   http://link.springer.com/content/
pdf/10.1007/s11538-011-9693-x.pdf.

Garnier:2012:SRB

[2660] Jimmy Garnier, Lionel Roques, and François Hamel. Success rate of a biological invasion in terms of the spatial distribution of the founding population. Bulletin of Mathematical Biology, 74(2):453–473, February 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
011-9694-9;   http://link.springer.com/content/pdf/10.1007/
s11538-011-9694-9.pdf.

Binder:2012:SAM

[2661] Benjamin J. Binder, Kerry A. Landman, Donald F. Newgreen, Johanna E. Simkin, Yoshiko Takahashi, and Dongcheng Zhang. Spatial analysis of multi-species exclusion processes: Application to neural crest cell migration in the embryonic gut. Bulletin of Mathematical Biology, 74 (2):474–490, February 2012. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-011-9703-z;   http://link.springer.com/content/
pdf/10.1007/s11538-011-9703-z.pdf.

Robertson:2012:LSI

[2662] Suzanne L. Robertson, J. M. Cushing, and R. F. Costantino. Life stages: Interactions and spatial patterns. Bulletin of Mathematical Biology, 74 (2):491–508, February 2012. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-011-9705-x;   http://link.springer.com/content/
pdf/10.1007/s11538-011-9705-x.pdf.

Leviyang:2012:SHI

[2663] Sivan Leviyang. Sampling HIV intrahost genealogies based on a model of acute stage CTL response. Bulletin of Mathematical Biology, 74(3): 509–535, March 2012. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-011-9670-4;   http://link.springer.com/content/
pdf/10.1007/s11538-011-9670-4.pdf.

Miller:2012:CTZ

[2664] Jennifer M. Miller, Allison Kolpas, Joao Plinio Juchem Neto, and Louis F. Rossi. A continuum three-zone model for swarms. Bulletin of Mathematical Biology, 74(3):536–561, March 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-011-9676-y;
 http://link.springer.com/content/pdf/10.1007/s11538-011-9676-
y.pdf.

Browne:2012:PMT

[2665] Cameron J. Browne and Sergei S. Pilyugin. Periodic multidrug therapy in a within-host virus model. Bulletin of Mathematical Biology, 74 (3):562–589, March 2012. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-011-9677-x;   http://link.springer.com/content/
pdf/10.1007/s11538-011-9677-x.pdf.

Curto:2012:FMN

[2666] Carina Curto, Anda Degeratu, and Vladimir Itskov. Flexible memory networks. Bulletin of Mathematical Biology, 74(3):590–614, March 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-

011-9678-9;   http://link.springer.com/content/pdf/10.1007/
s11538-011-9678-9.pdf.

Gurney:2012:SAV

[2667] William S. C. Gurney, Philip J. Bacon, Douglas C. Speirs, Philip McGinnity, and Eric Verspoor. Sea-age variation in maiden Atlantic Salmon spawners: Phenotypic plasticity or genetic polymorphism? Bulletin of Mathematical Biology, 74(3):615–640, March 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
011-9679-8;   http://link.springer.com/content/pdf/10.1007/
s11538-011-9679-8.pdf.

Ackleh:2012:FSJ

[2668] Azmy S. Ackleh, Jacoby Carter, Keng Deng, Qihua Huang, Nabendu Pal, and Xing Yang. Fitting a structured juvenile–adult model for green tree frogs to population estimates from capture–mark–recapture field data. Bulletin of Mathematical Biology, 74(3):641–665, March 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
011-9682-0;   http://link.springer.com/content/pdf/10.1007/
s11538-011-9682-0.pdf.

Glimm:2012:RDS

[2669] Tilmann Glimm, Jianying Zhang, Yun-Qiu Shen, and Stuart A. Newman. Reaction–diffusion systems and external morphogen gradients: The two-dimensional case, with an application to skeletal pattern formation. Bulletin of Mathematical Biology, 74(3):666–687, March 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
011-9689-6;   http://link.springer.com/content/pdf/10.1007/
s11538-011-9689-6.pdf.

Bazil:2012:GPM

[2670] Jason N. Bazil, Gregory T. Buzzard, and Ann E. Rundell. A global parallel model based design of experiments method to minimize model output uncertainty. Bulletin of Mathematical Biology, 74(3):688–716, March 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
011-9686-9;   http://link.springer.com/content/pdf/10.1007/
s11538-011-9686-9.pdf.

Bennett:2012:MNC

[2671] Max R. Bennett, Les Farnell, and William G. Gibson. A model of neuregulin control of NMDA receptors on synaptic spines. Bulletin of Mathematical Biology, 74(3):717–735, March 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-011-9706-9;
 http://link.springer.com/content/pdf/10.1007/s11538-011-9706-
9.pdf.

Simms:2012:MMC

[2672] Kate Simms, Nigel Bean, and Adrian Koerber. A mathematical model of cell cycle progression applied to the MCF-7 breast cancer cell line. Bulletin of Mathematical Biology, 74(3):736–767, March 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
011-9700-2;   http://link.springer.com/content/pdf/10.1007/
s11538-011-9700-2.pdf.

DeVille:2012:SAN

[2673] R. E. Lee DeVille and Charles S. Peskin. Synchrony and asynchrony for neuronal dynamics defined on complex networks. Bulletin of Mathematical Biology, 74(4):769–802, April 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-011-9674-0;
 http://link.springer.com/content/pdf/10.1007/s11538-011-9674-
0.pdf.

Kealy:2012:NSA

[2674] Bonni J. Kealy and David J. Wollkind. A nonlinear stability analysis of vegetative Turing pattern formation for an interaction– diffusion plant–surface water model system in an arid flat environment. Bulletin of Mathematical Biology, 74(4):803–833, April 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
011-9688-7;   http://link.springer.com/content/pdf/10.1007/
s11538-011-9688-7.pdf.

Merkey:2012:PMF

[2675] Brian V. Merkey and David L. Chopp. The performance of a microbial fuel cell depends strongly on anode geometry: a multidimensional modeling study. Bulletin of Mathematical Biology, 74(4):834–857, April 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
011-9690-0;   http://link.springer.com/content/pdf/10.1007/
s11538-011-9690-0.pdf.

Sumner:2012:AGM

[2676] J. G. Sumner, B. R. Holland, and P. D. Jarvis. The algebra of the general Markov model on phylogenetic trees and networks. Bulletin of Mathematical Biology, 74(4):858–880, April 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
011-9691-z;   http://link.springer.com/content/pdf/10.1007/
s11538-011-9691-z.pdf.

Smith:2012:ITF

[2677] Joshua H. Smith, Kathleen A. Starkweather, and José Jaime García. Implications of transvascular fluid exchange in nonlinear, biphasic analyses of flow-controlled infusion in brain. Bulletin of Mathematical Biology, 74 (4):881–907, April 2012. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-011-9696-7;   http://link.springer.com/content/
pdf/10.1007/s11538-011-9696-7.pdf.

Pena-Miller:2012:SAA

[2678] Rafael Peña-Miller, David Lähnemann, Hinrich Schulenburg, Martin Ackermann, and Robert Beardmore. Selecting against antibioticresistant pathogens: Optimal treatments in the presence of commensal bacteria. Bulletin of Mathematical Biology, 74(4):908–934, April 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
011-9698-5;   http://link.springer.com/content/pdf/10.1007/
s11538-011-9698-5.pdf.

Klika:2012:IRM

[2679] Václav Klika, Ruth E. Baker, Denis Headon, and Eamonn A. Gaffney. The influence of receptor–mediated interactions on reaction–diffusion mechanisms of cellular self-organisation. Bulletin of Mathematical Biology, 74(4):935–957, April 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-011-9699-4;   http://link.springer.com/
content/pdf/10.1007/s11538-011-9699-4.pdf.

Lee:2012:MOA

[2680] Sunmi Lee, Michael Golinski, and Gerardo Chowell. Modeling optimal age-specific vaccination strategies against pandemic influenza. Bulletin of Mathematical Biology, 74(4):958–980, April 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-011-9704-y;
 http://link.springer.com/content/pdf/10.1007/s11538-011-9704-
y.pdf.

Cheer:2012:CFD

[2681] Angela Cheer, Samson Cheung, Tien-Chieh Hung, Raul H. Piedrahita, and S. Laurie Sanderson. Computational fluid dynamics of fish Gill rakers during crossflow filtration. Bulletin of Mathematical Biology, 74 (4):981–1000, April 2012. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-011-9709-6;   http://link.springer.com/content/
pdf/10.1007/s11538-011-9709-6.pdf.

Masic:2012:PSS

[2682] Alma Mašić and Hermann J. Eberl. Persistence in a single species CSTR model with suspended flocs and wall attached biofilms. Bulletin of Mathematical Biology, 74(4):1001–1026, April 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-011-9707-8;
 http://link.springer.com/content/pdf/10.1007/s11538-011-9707-
8.pdf.

PerezMillan:2012:CRS

[2683] Mercedes Pérez Millán, Alicia Dickenstein, Anne Shiu, and Carsten Conradi. Chemical reaction systems with toric steady states. Bulletin of Mathematical Biology, 74(5):1027–1065, May 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-011-9685-x;
 http://link.springer.com/content/pdf/10.1007/s11538-011-9685-
x.pdf.

Hughes:2012:KEN

[2684] John Hughes, William O. Hancock, and John Fricks. Kinesins with extended neck linkers: a chemomechanical model for variable-length stepping. Bulletin of Mathematical Biology, 74(5):1066–1097, May 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-

011-9697-6;   http://link.springer.com/content/pdf/10.1007/
s11538-011-9697-6.pdf.

Chitnis:2012:PFM

[2685] Nakul Chitnis, Diggory Hardy, and Thomas Smith. A periodicallyforced mathematical model for the seasonal dynamics of malaria in mosquitoes. Bulletin of Mathematical Biology, 74(5):1098–1124, May 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
011-9710-0;   http://link.springer.com/content/pdf/10.1007/
s11538-011-9710-0.pdf.

Mendoza-Juez:2012:MMG

[2686] Berta Mendoza-Juez, Alicia Martínez-González, Gabriel F. Calvo, and Víctor M. Pérez-García. A mathematical model for the glucose–lactate metabolism of in vitro cancer cells. Bulletin of Mathematical Biology, 74 (5):1125–1142, May 2012. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-011-9711-z;   http://link.springer.com/content/
pdf/10.1007/s11538-011-9711-z.pdf.

Murphy:2012:FMM

[2687] Kelly E. Murphy, Cameron L. Hall, Philip K. Maini, Scott W. Mc-Cue, and D. L. Sean McElwain. A fibrocontractive mechanochemical model of dermal wound closure incorporating realistic growth factor kinetics. Bulletin of Mathematical Biology, 74(5):1143–1170, May 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
011-9712-y;   http://link.springer.com/content/pdf/10.1007/
s11538-011-9712-y.pdf.

Cummings:2012:ABE

[2688] L. J. Cummings, R. Perez-Castillejos, and E. T. Mack. Analysis of biochemical equilibria relevant to the immune response: Finding the dissociation constants. Bulletin of Mathematical Biology, 74(5):1171–1206, May 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
012-9716-2;   http://link.springer.com/content/pdf/10.1007/
s11538-012-9716-2.pdf.

Tian:2012:MMT

[2689] Jianjun Paul Tian, Zhigang Jin, and Ting Xie. Mathematical model for two germline stem cells competing for niche occupancy. Bulletin of Mathematical Biology, 74(5):1207–1225, May 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
011-9713-x;   http://link.springer.com/content/pdf/10.1007/
s11538-011-9713-x.pdf.

Zhang:2012:MSR

[2690] Juan Zhang, Zhen Jin, Gui-Quan Sun, Xiang-Dong Sun, and Shigui Ruan. Modeling seasonal rabies epidemics in China. Bulletin of Mathematical Biology, 74(5):1226–1251, May 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-012-9720-6;
 http://link.springer.com/content/pdf/10.1007/s11538-012-9720-

6.pdf. See erratum [2771].

Scianna:2012:MHM

[2691] Marco Scianna. A multiscale hybrid model for Pro-angiogenic calcium signals in a vascular endothelial cell. Bulletin of Mathematical Biology, 74 (6):1253–1291, June 2012. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-011-9695-8;   http://link.springer.com/content/
pdf/10.1007/s11538-011-9695-8.pdf.

Jabbari:2012:CSQa

[2692] Sara Jabbari, John R. King, and Paul Williams. Cross-strain quorum sensing inhibition by Staphylococcus aureus. Part 1: a spatially homogeneous model. Bulletin of Mathematical Biology, 74(6):1292–1325, June 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
011-9701-1;   http://link.springer.com/content/pdf/10.1007/
s11538-011-9701-1.pdf.

Jabbari:2012:CSQb

[2693] Sara Jabbari, John R. King, and Paul Williams. Cross-strain quorum sensing inhibition by Staphylococcus aureus. Part 2: a spatially inhomogeneous model. Bulletin of Mathematical Biology, 74(6):1326–1353, June 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
011-9702-0;   http://link.springer.com/content/pdf/10.1007/
s11538-011-9702-0.pdf.

Wikan:2012:ASS

[2694] Arild Wikan. Age or stage structure? Bulletin of Mathematical Biology, 74(6):1354–1378, June 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-012-9715-3;   http://link.springer.com/
content/pdf/10.1007/s11538-012-9715-3.pdf.

Tomasetti:2012:PRG

[2695] Cristian Tomasetti. On the probability of random genetic mutations for various types of tumor growth. Bulletin of Mathematical Biology, 74 (6):1379–1395, June 2012. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-012-9717-1;   http://link.springer.com/content/
pdf/10.1007/s11538-012-9717-1.pdf.

Jayasuriya:2012:ETD

[2696] Suren Jayasuriya and Zachary P. Kilpatrick. Effects of timedependent stimuli in a competitive neural network model of perceptual rivalry. Bulletin of Mathematical Biology, 74(6):1396–1426, June 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
012-9718-0;   http://link.springer.com/content/pdf/10.1007/
s11538-012-9718-0.pdf.

Zhang:2012:MBA

[2697] Tianyu Zhang. Modeling of biocide action against biofilm. Bulletin of Mathematical Biology, 74(6):1427–1447, June 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-012-9719-z;
 http://link.springer.com/content/pdf/10.1007/s11538-012-9719-
z.pdf.

Kawasaki:2012:IDM

[2698] Kohkichi Kawasaki, Keisuke Asano, and Nanako Shigesada. Impact of directed movement on invasive spread in periodic patchy environments. Bulletin of Mathematical Biology, 74(6):1448–1467, June 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
011-9714-9;   http://link.springer.com/content/pdf/10.1007/
s11538-011-9714-9.pdf.

Huang:2012:GDM

[2699] Yuanyuan Huang and Zhijun Wu. Game dynamic model for yeast development. Bulletin of Mathematical Biology, 74(7):1469–1484, July 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
012-9721-5;   http://link.springer.com/content/pdf/10.1007/
s11538-012-9721-5.pdf.

Wilson:2012:MME

[2700] Shelby Wilson and Doron Levy. A mathematical model of the enhancement of tumor vaccine efficacy by immunotherapy. Bulletin of Mathematical Biology, 74(7):1485–1500, July 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-012-9722-4;
 http://link.springer.com/content/pdf/10.1007/s11538-012-9722-
4.pdf.

Kiss:2012:NMC

[2701] Istvan Z. Kiss and Péter L. Simon. New moment closures based on a priori distributions with applications to epidemic dynamics. Bulletin of Mathematical Biology, 74(7):1501–1515, July 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
012-9723-3;   http://link.springer.com/content/pdf/10.1007/
s11538-012-9723-3.pdf.

Benson:2012:AOC

[2702] James D. Benson, Carmen C. Chicone, and John K. Critser. Analytical optimal controls for the state constrained addition and removal of cryoprotective agents. Bulletin of Mathematical Biology, 74(7):1516–1530, July 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
012-9724-2;   http://link.springer.com/content/pdf/10.1007/
s11538-012-9724-2.pdf.

Sturrock:2012:INM

[2703] Marc Sturrock, Alan J. Terry, Dimitris P. Xirodimas, Alastair M. Thompson, and Mark A. J. Chaplain. Influence of the nuclear membrane, active transport, and cell shape on the Hes1 and p53– Mdm2 pathways: Insights from spatio-temporal modelling. Bulletin of Mathematical Biology, 74(7):1531–1579, July 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-012-9725-1;
 http://link.springer.com/content/pdf/10.1007/s11538-012-9725-
1.pdf.

Anazawa:2012:ICM

[2704] Masahiro Anazawa. Interspecific competition models derived from competition among individuals. Bulletin of Mathematical Biology, 74(7): 1580–1605, July 2012. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-012-9726-0;   http://link.springer.com/content/
pdf/10.1007/s11538-012-9726-0.pdf.

Drury:2012:CPP

[2705] Kevin L. S. Drury and David M. Lodge. Combining perturbations and parameter variation to influence mean first passage times. Bulletin of Mathematical Biology, 74(7):1606–1628, July 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
012-9727-z;   http://link.springer.com/content/pdf/10.1007/
s11538-012-9727-z.pdf.

Samia:2012:PPS

[2706] Yasmine Samia and Frithjof Lutscher. Persistence probabilities for stream populations. Bulletin of Mathematical Biology, 74(7):1629–1650, July 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
012-9728-y;   http://link.springer.com/content/pdf/10.1007/
s11538-012-9728-y.pdf.

Forde:2012:LIC

[2707] Jonathan Forde, Joseph M. Volpe, and Stanca M. Ciupe. Latently infected cell activation: a way to reduce the size of the HIV reservoir? Bulletin of Mathematical Biology, 74(7):1651–1672, July 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
012-9729-x;   http://link.springer.com/content/pdf/10.1007/
s11538-012-9729-x.pdf.

Yokley:2012:SIR

[2708] Karen A. Yokley. Sensory irritation response in rats II: Recovery and dose-dependence. Bulletin of Mathematical Biology, 74(7):1673–1690, July 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-

012-9730-4;   http://link.springer.com/content/pdf/10.1007/
s11538-012-9730-4.pdf.

Bootsma:2012:MNI

[2709] M. C. J. Bootsma, M. A. van der Horst, T. Guryeva, B. H. ter Kuile, and O. Diekmann. Modeling non-inherited antibiotic resistance. Bulletin of Mathematical Biology, 74(8):1691–1705, August 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
012-9731-3;   http://link.springer.com/content/pdf/10.1007/
s11538-012-9731-3.pdf.

Gedeon:2012:EMS

[2710] Tomáš Gedeon, Graham Cummins, and Jeffrey J. Heys. Effect of model selection on prediction of periodic behavior in gene regulatory networks. Bulletin of Mathematical Biology, 74(8):1706–1726, August 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
012-9732-2;   http://link.springer.com/content/pdf/10.1007/
s11538-012-9732-2.pdf.

Hu:2012:MHG

[2711] J. Hu, K. R. Qin, C. Xiang, and T. H. Lee. Modeling of hysteresis in gene regulatory networks. Bulletin of Mathematical Biology, 74(8): 1727–1753, August 2012. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-012-9733-1;   http://link.springer.com/content/
pdf/10.1007/s11538-012-9733-1.pdf.

Chisholm:2012:LDI

[2712] Ryan A. Chisholm and Simon A. Levin. Linking dispersal and immigration in multidimensional environments. Bulletin of Mathematical Biology, 74(8):1754–1763, August 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-012-9734-0;   http://link.springer.com/
content/pdf/10.1007/s11538-012-9734-0.pdf.

Lebre:2012:EMS

[2713] Sophie Lèbre and Christian J. Michel. An evolution model for sequence length based on residue insertion–deletion independent of substitution: an application to the GC content in bacterial genomes. Bulletin of Mathematical Biology, 74(8):1764–1788, August 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
012-9735-z;   http://link.springer.com/content/pdf/10.1007/
s11538-012-9735-z.pdf.

Rong:2012:MQD

[2714] Libin Rong, Ruy M. Ribeiro, and Alan S. Perelson. Modeling quasispecies and drug resistance in hepatitis c patients treated with a protease inhibitor. Bulletin of Mathematical Biology, 74(8):1789–1817, August 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
012-9736-y;   http://link.springer.com/content/pdf/10.1007/
s11538-012-9736-y.pdf.

Leviyang:2012:CIH

[2715] Sivan Leviyang. The coalescence of intrahost HIV lineages under symmetric CTL attack. Bulletin of Mathematical Biology, 74(8):1818– 1856, August 2012. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-012-9737-x;   http://link.springer.com/content/
pdf/10.1007/s11538-012-9737-x.pdf.

Espinoza:2012:ICM

[2716] Flor A. Espinoza, Michael J. Wester, Janet M. Oliver, Bridget S. Wilson, Nicholas L. Andrews, Diane S. Lidke, and Stanly L. Steinberg. Insights into cell membrane microdomain organization from live cell single particle tracking of the IgE high affinity receptor FcϵRI of mast cells. Bulletin of Mathematical Biology, 74(8):1857–1911, August 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
012-9738-9;   http://link.springer.com/content/pdf/10.1007/
s11538-012-9738-9.pdf.

Kwong:2012:LFI

[2717] Grace P. S. Kwong and Rob Deardon. Linearized forms of individuallevel models for large-scale spatial infectious disease systems. Bulletin of Mathematical Biology, 74(8):1912–1937, August 2012. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-012-9739-8;
 http://link.springer.com/content/pdf/10.1007/s11538-012-9739-
8.pdf.

Rempala:2012:LSE

[2718] Grzegorz A. Rempala. Least squares estimation in stochastic biochemical networks. Bulletin of Mathematical Biology, 74(8):1938–1955, August 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
012-9744-y;   http://link.springer.com/content/pdf/10.1007/
s11538-012-9744-y.pdf.

Zabalo:2012:PIP

[2719] Joaquin Zabalo. Permanence in an intraguild predation model with prey switching. Bulletin of Mathematical Biology, 74(9):1957–1984, September 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
012-9740-2;   http://link.springer.com/content/pdf/10.1007/
s11538-012-9740-2.pdf.

Ward:2012:ILI

[2720] Zoë Ward and Jane White. Impact of latently infected cells on strain archiving within HIV hosts. Bulletin of Mathematical Biology, 74(9): 1985–2003, September 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-012-9742-0;   http://link.springer.com/
content/pdf/10.1007/s11538-012-9742-0.pdf.

Adamson:2012:RRS

[2721] M. W. Adamson and A. Y. Morozov. Revising the role of species mobility in maintaining biodiversity in communities with cyclic competition. Bulletin of Mathematical Biology, 74(9):2004–2031, September 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
012-9743-z;   http://link.springer.com/content/pdf/10.1007/
s11538-012-9743-z.pdf.

Crossley:2012:CMN

[2722] Laurie Crossley, Caroline A. Garrett, Majid Hafezparast, and Anotida Madzvamuse. From the cell membrane to the nucleus: Unearthing transport mechanisms for dynein. Bulletin of Mathematical Biology, 74(9):2032–2061, September 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-012-9745-x;   http://link.springer.com/
content/pdf/10.1007/s11538-012-9745-x.pdf.

Zubkov:2012:CFS

[2723] V. S. Zubkov, C. J. W. Breward, and E. A. Gaffney. Coupling fluid and solute dynamics within the ocular surface tear film: a modelling study of black line osmolarity. Bulletin of Mathematical Biology, 74(9):2062– 2093, September 2012. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-012-9746-9;   http://link.springer.com/content/
pdf/10.1007/s11538-012-9746-9.pdf.

Malunguza:2012:IAC

[2724] N. J. Malunguza, S. D. Hove-Musekwa, G. Musuka, and Z. Mukandavire. Investigating alcohol consumption as a risk factor for HIV transmission in heterosexual settings in Sub-Saharan African communities. Bulletin of Mathematical Biology, 74(9):2094–2124, September 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
012-9747-8;   http://link.springer.com/content/pdf/10.1007/
s11538-012-9747-8.pdf.

Miller:2012:NDE

[2725] Joel C. Miller. A note on the derivation of epidemic final sizes. Bulletin of Mathematical Biology, 74(9):2125–2141, September 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
012-9749-6;   http://link.springer.com/content/pdf/10.1007/
s11538-012-9749-6.pdf.

Wang:2012:UDI

[2726] Yuanshi Wang, Donald L. DeAngelis, and J. Nathaniel Holland. Uni-directional interaction and plant–pollinator–robber coexistence. Bulletin of Mathematical Biology, 74(9):2142–2164, September 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
012-9750-0;   http://link.springer.com/content/pdf/10.1007/
s11538-012-9750-0.pdf.

Segal:2012:DEM

[2727] Rebecca A. Segal, Robert F. Diegelmann, Kevin R. Ward, and Angela Reynolds. A differential equation model of collagen accumulation in a healing wound. Bulletin of Mathematical Biology, 74(9):2165– 2182, September 2012. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-012-9751-z;   http://link.springer.com/content/
pdf/10.1007/s11538-012-9751-z.pdf.

Rashkov:2012:MOP

[2728] Peter Rashkov, Bernhard A. Schmitt, Lotte Søgaard-Andersen, Peter Lenz, and Stephan Dahlke. A model of oscillatory protein dynamics in bacteria. Bulletin of Mathematical Biology, 74(9):2183–2203, September 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
012-9752-y;   http://link.springer.com/content/pdf/10.1007/
s11538-012-9752-y.pdf.

Al-Shammari:2012:REU

[2729] A. A. Al-Shammari, E. A. Gaffney, and S. Egginton. Re-evaluating the use of Voronoi tessellations in the assessment of oxygen supply from capillaries in muscle. Bulletin of Mathematical Biology, 74(9):2204– 2231, September 2012. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-012-9753-x;   http://link.springer.com/content/
pdf/10.1007/s11538-012-9753-x.pdf.

Powell:2012:CBE

[2730] James A. Powell, Brynja R. Kohler, James W. Haefner, and Janice Bodily. Carrying BioMath education in a leaky bucket. Bulletin of Mathematical Biology, 74(9):2232–2264, September 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
012-9741-1;   http://link.springer.com/content/pdf/10.1007/
s11538-012-9741-1.pdf.

Meinhardt:2012:NOR

[2731] Hans Meinhardt. No oscillations in real activator–inhibitor systems in accomplishing pattern formation. Bulletin of Mathematical Biology, 74(10): 2265–2267, October 2012. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-012-9767-4;   http://link.springer.com/content/
pdf/10.1007/s11538-012-9767-4.pdf. See reply [2732].

Gaffney:2012:RCN

[2732] Eamonn A. Gaffney, Nick A. M. Monk, Ruth E. Baker, and S. Seirin Lee. Reply to correspondence: No Oscillations in Real Activator–Inhibitor Systems in Accomplishing Pattern Formation.

Bulletin of Mathematical Biology, 74(10):2268–2271, October 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
012-9768-3;   http://link.springer.com/content/pdf/10.1007/
s11538-012-9768-3.pdf. See [2731].

McDougall:2012:HDC

[2733] S. R. McDougall, M. G. Watson, A. H. Devlin, C. A. Mitchell, and M. A. J. Chaplain. A hybrid discrete–continuum mathematical model of pattern prediction in the developing retinal vasculature. Bulletin of Mathematical Biology, 74(10):2272–2314, October 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
012-9754-9;   http://link.springer.com/content/pdf/10.1007/
s11538-012-9754-9.pdf.

Nattrass:2012:QLC

[2734] Stuart Nattrass, Stephen Baigent, and David J. Murrell. Quantifying the likelihood of co-existence for communities with asymmetric competition. Bulletin of Mathematical Biology, 74(10):2315–2338, October 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
012-9755-8;   http://link.springer.com/content/pdf/10.1007/
s11538-012-9755-8.pdf.

Xin:2012:TDB

[2735] Xiangrong Xin and Hans G. Othmer. A “trimer of dimers”-based model for the chemotactic signal transduction network in bacterial chemotaxis. Bulletin of Mathematical Biology, 74(10):2339–2382, October 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
012-9756-7;   http://link.springer.com/content/pdf/10.1007/
s11538-012-9756-7.pdf.

Lewis:2012:SST

[2736] Mark A. Lewis and Bingtuan Li. Spreading speed, traveling waves, and minimal domain size in impulsive reaction–diffusion models. Bulletin of Mathematical Biology, 74(10):2383–2402, October 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
012-9757-6;   http://link.springer.com/content/pdf/10.1007/
s11538-012-9757-6.pdf.

Xiao:2012:SMC

[2737] Yanni Xiao, Xiaxia Xu, and Sanyi Tang. Sliding mode control of outbreaks of emerging infectious diseases. Bulletin of Mathematical Biology, 74(10):2403–2422, October 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-012-9758-5;
 http://link.springer.com/content/pdf/10.1007/s11538-012-9758-
5.pdf.

Shuai:2012:CMH

[2738] Zhisheng Shuai, Joseph H. Tien, and P. van den Driessche. Cholera models with hyperinfectivity and temporary immunity. Bulletin of Mathematical Biology, 74(10):2423–2445, October 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-012-9759-4;
 http://link.springer.com/content/pdf/10.1007/s11538-012-9759-
4.pdf.

Li:2012:NDM

[2739] Yongfeng Li, Scott C. Lenaghan, and Mingjun Zhang. Nonlinear dynamics of the movement of the Venus flytrap. Bulletin of Mathematical Biology, 74(10):2446–2473, October 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-012-9760-y;
 http://link.springer.com/content/pdf/10.1007/s11538-012-9760-
y.pdf.

Lachiany:2012:VMM

[2740] Menachem Lachiany and Lewi Stone. A vaccination model for a multicity system. Bulletin of Mathematical Biology, 74(10):2474–2487, October 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
012-9762-9;   http://link.springer.com/content/pdf/10.1007/
s11538-012-9762-9.pdf.

Blyuss:2012:SBM

[2741] Konstantin B. Blyuss and Yuliya N. Kyrychko. Symmetry breaking in a model of antigenic variation with immune delay. Bulletin of Mathematical Biology, 74(10):2488–2509, October 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-012-9763-8;
 http://link.springer.com/content/pdf/10.1007/s11538-012-9763-
8.pdf.

Daukste:2012:MDC

[2742] Liene Daukste, Britta Basse, Bruce C. Baguley, and David J. N. Wall. Mathematical determination of cell population doubling times for multiple cell lines. Bulletin of Mathematical Biology, 74(10):2510–2534, October 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
012-9764-7;   http://link.springer.com/content/pdf/10.1007/
s11538-012-9764-7.pdf.

Bakalis:2012:PTC

[2743] Evangelos Bakalis, Marios Kosmas, and Emmanouel M. Papamichael. Perturbation theory in the catalytic rate constant of the Henri– Michaelis–Menten enzymatic reaction. Bulletin of Mathematical Biology, 74(11):2535–2546, November 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-012-9761-x;   http://link.springer.com/
content/pdf/10.1007/s11538-012-9761-x.pdf.

Hamlet:2012:FCU

[2744] Christina L. Hamlet and Laura A. Miller. Feeding currents of the upside down jellyfish in the presence of background flow. Bulletin of Mathematical Biology, 74(11):2547–2569, November 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
012-9765-6;   http://link.springer.com/content/pdf/10.1007/
s11538-012-9765-6.pdf.

Walther:2012:DVS

[2745] Georg R. Walther, Athanasius F. M. Marée, Leah Edelstein-Keshet, and Verônica A. Grieneisen. Deterministic versus stochastic cell polarisation through wave-pinning. Bulletin of Mathematical Biology, 74(11): 2570–2599, November 2012. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-012-9766-5;   http://link.springer.com/content/
pdf/10.1007/s11538-012-9766-5.pdf.

Fetecau:2012:MMF

[2746] R. C. Fetecau and A. Guo. A mathematical model for flight guidance in honeybee swarms. Bulletin of Mathematical Biology, 74(11):2600– 2621, November 2012. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-012-9769-2;   http://link.springer.com/content/
pdf/10.1007/s11538-012-9769-2.pdf.

Parvinen:2012:ECD

[2747] Kalle Parvinen, Anne Seppänen, and John D. Nagy. Evolution of complex density-dependent dispersal strategies. Bulletin of Mathematical Biology, 74(11):2622–2649, November 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-012-9770-9;
 http://link.springer.com/content/pdf/10.1007/s11538-012-9770-
9.pdf.

Allen:2012:MRE

[2748] Benjamin Allen and Daniel I. Scholes Rosenbloom. Mutation rate evolution in replicator dynamics. Bulletin of Mathematical Biology, 74(11): 2650–2675, November 2012. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-012-9771-8;   http://link.springer.com/content/
pdf/10.1007/s11538-012-9771-8.pdf.

Garay:2012:MFA

[2749] József Garay and Tamás F. Móri. Monogamy has a fixation advantage based on fitness variance in an ideal promiscuity group. Bulletin of Mathematical Biology, 74(11):2676–2691, November 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
012-9772-7;   http://link.springer.com/content/pdf/10.1007/
s11538-012-9772-7.pdf.

Hasenauer:2012:ASD

[2750] J. Hasenauer, D. Schittler, and F. Allgöwer. Analysis and simulation of division- and label-structured population models. Bulletin of Mathematical Biology, 74(11):2692–2732, November 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
012-9774-5;   http://link.springer.com/content/pdf/10.1007/
s11538-012-9774-5.pdf.

Akbari:2012:TDS

[2751] Samira Akbari, Naveen K. Vaidya, and Lindi M. Wahl. The time distribution of sulfadoxine–pyrimethamine protection from malaria. Bulletin of Mathematical Biology, 74(11):2733–2751, November 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
012-9775-4;   http://link.springer.com/content/pdf/10.1007/
s11538-012-9775-4.pdf.

Chen:2012:HIF

[2752] Duan Chen, Julie M. Roda, Clay B. Marsh, Timothy D. Eubank, and Avner Friedman. Hypoxia inducible factors-mediated inhibition of cancer by GM–CSF: a mathematical model. Bulletin of Mathematical Biology, 74(11):2752–2777, November 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
012-9776-3;   http://link.springer.com/content/pdf/10.1007/
s11538-012-9776-3.pdf.

Veliz-Cuba:2012:RSS

[2753] Alan Veliz-Cuba, Joseph Arthur, Laura Hochstetler, Victoria Klomps, and Erikka Korpi. On the relationship of steady states of continuous and discrete models arising from biology. Bulletin of Mathematical Biology, 74(12):2779–2792, December 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-012-9778-1;
 http://link.springer.com/content/pdf/10.1007/s11538-012-9778-
1.pdf.

Thompson:2012:MCM

[2754] Robin N. Thompson, Christian A. Yates, and Ruth E. Baker. Modelling cell migration and adhesion during development. Bulletin of Mathematical Biology, 74(12):2793–2809, December 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
012-9779-0;   http://link.springer.com/content/pdf/10.1007/
s11538-012-9779-0.pdf.

Bhattacharya:2012:TSR

[2755] Subhra Bhattacharya and Frederick R. Adler. A time since recovery model with varying rates of loss of immunity. Bulletin of Mathematical Biology, 74(12):2810–2819, December 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-012-9780-7;
 http://link.springer.com/content/pdf/10.1007/s11538-012-9780-
7.pdf.

Postnov:2012:DNV

[2756] D. D. Postnov, D. E. Postnov, D. J. Marsh, N.-H. Holstein-Rathlou, and O. V. Sosnovtseva. Dynamics of nephron–vascular network. Bulletin of Mathematical Biology, 74(12):2820–2841, December 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
012-9781-6;   http://link.springer.com/content/pdf/10.1007/
s11538-012-9781-6.pdf.

Bodin:2012:SOH

[2757] M. Bodin, Å. Brännström, and U. Dieckmann. A systematic overview of harvesting-induced maturation evolution in predator– prey systems with three different life-history tradeoffs. Bulletin of Mathematical Biology, 74(12):2842–2860, December 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-012-9783-4;
 http://link.springer.com/content/pdf/10.1007/s11538-012-9783-
4.pdf.

Singh:2012:SSM

[2758] Aditya P. Singh, Erik Andries, Jeremy S. Edwards, and Stanly Steinberg. The shuttling scaffold model for prevention of yeast pheromone pathway misactivation. Bulletin of Mathematical Biology, 74(12):2861– 2874, December 2012. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-012-9785-2;   http://link.springer.com/content/
pdf/10.1007/s11538-012-9785-2.pdf.

Martinez-Gonzalez:2012:HCW

[2759] Alicia Martínez-González, Gabriel F. Calvo, Luis A. Pérez Romasanta, and Víctor M. Pérez-García. Hypoxic cell waves around necrotic cores in glioblastoma: a biomathematical model and its therapeutic implications. Bulletin of Mathematical Biology, 74(12):2875–2896, December 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
012-9786-1;   http://link.springer.com/content/pdf/10.1007/
s11538-012-9786-1.pdf.

Wei:2012:ESM

[2760] Yi Wei and Alexei A. Koulakov. An exactly solvable model of random site-specific recombinations. Bulletin of Mathematical Biology, 74(12): 2897–2916, December 2012. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-012-9788-z;   http://link.springer.com/content/
pdf/10.1007/s11538-012-9788-z.pdf.

Klapper:2012:PES

[2761] Isaac Klapper. Productivity and equilibrium in simple biofilm models. Bulletin of Mathematical Biology, 74(12):2917–2934, December 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
012-9791-4;   http://link.springer.com/content/pdf/10.1007/
s11538-012-9791-4.pdf.

Vasilyeva:2012:HFS

[2762] Olga Vasilyeva and Frithjof Lutscher. How flow speed alters competitive outcome in advective environments. Bulletin of Mathematical Biology, 74(12):2935–2958, December 2012. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-012-9792-3;   http://link.springer.com/
content/pdf/10.1007/s11538-012-9792-3.pdf.

Meng:2013:DMS

[2763] Xinzhu Meng, Niklas L. P. Lundström, Mats Bodin, and Åke Brännström. Dynamics and management of stage-structured fish stocks. Bulletin of Mathematical Biology, 75(1):1–23, January 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
012-9789-y;   http://link.springer.com/content/pdf/10.1007/
s11538-012-9789-y.pdf.

Nyabadza:2013:MDC

[2764] Farai Nyabadza, John B. H. Njagarah, and Robert J. Smith. Modelling the dynamics of crystal meth (‘tik’) abuse in the presence of drugsupply chains in South Africa. Bulletin of Mathematical Biology, 75 (1):24–48, January 2013. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-012-9790-5;   http://link.springer.com/content/
pdf/10.1007/s11538-012-9790-5.pdf.

Heppell:2013:MFD

[2765] Charles Heppell, Giles Richardson, and Tiina Roose. A model for fluid drainage by the lymphatic system. Bulletin of Mathematical Biology, 75 (1):49–81, January 2013. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-012-9793-2;   http://link.springer.com/content/
pdf/10.1007/s11538-012-9793-2.pdf.

Amaku:2013:IHV

[2766] Marcos Amaku, Francisco Antonio Bezerra Coutinho, Eleazar Chaib, and Eduardo Massad. The impact of hepatitis a virus infection on hepatitis c virus infection: a competitive exclusion hypothesis. Bulletin of Mathematical Biology, 75(1):82–93, January 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
012-9795-0;   http://link.springer.com/content/pdf/10.1007/
s11538-012-9795-0.pdf.

Cogan:2013:EPD

[2767] N. G. Cogan, Barbara Szomolay, and Martin Dindos. Effect of periodic disinfection on persisters in a one-dimensional biofilm model. Bulletin of Mathematical Biology, 75(1):94–123, January 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
012-9796-z;   http://link.springer.com/content/pdf/10.1007/
s11538-012-9796-z.pdf.

Park:2013:CGN

[2768] Jun-Koo Park, Robert Jernigan, and Zhijun Wu. Coarse grained normal mode analysis vs. refined Gaussian network model for protein residuelevel structural fluctuations. Bulletin of Mathematical Biology, 75(1): 124–160, January 2013. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-012-9797-y;   http://link.springer.com/content/
pdf/10.1007/s11538-012-9797-y.pdf.

Hillen:2013:TGP

[2769] Thomas Hillen, Heiko Enderling, and Philip Hahnfeldt. The tumor growth paradox and immune system-mediated selection for cancer stem cells. Bulletin of Mathematical Biology, 75(1):161–184, January 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
012-9798-x;   http://link.springer.com/content/pdf/10.1007/
s11538-012-9798-x.pdf.

Chan:2013:MMB

[2770] Cliburn Chan, Matthew Billard, Samuel A. Ramirez, Harald Schmidl, Eric Monson, and Thomas B. Kepler. A model for migratory B cell oscillations from receptor down-regulation induced by external chemokine fields. Bulletin of Mathematical Biology, 75(1):185–205, January 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
012-9799-9;   http://link.springer.com/content/pdf/10.1007/
s11538-012-9799-9.pdf.

Zhang:2013:EMS

[2771] Juan Zhang, Zhen Jin, Gui-Quan Sun, Xiang-Dong Sun, and Shigui Ruan. Erratum to: Modeling Seasonal Rabies Epidemics in China. Bulletin of Mathematical Biology, 75(1):206–211, January 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
012-9803-4;   http://link.springer.com/content/pdf/10.1007/
s11538-012-9803-4.pdf. See [2690].

Xiao:2013:CNT

[2772] Sa Xiao, Liang Zhao, Jia-Lin Zhang, Xiang-Tai Wang, and Shu-Yan Chen. Comparing neutral and trade-off community models in shaping the community biomass–diversity relationship under different disturbance levels. Bulletin of Mathematical Biology, 75(2):213–222, February 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
012-9800-7;   http://link.springer.com/content/pdf/10.1007/
s11538-012-9800-7.pdf.

Weir:2013:IEE

[2773] Brad Weir, Robert N. Miller, and Yvette H. Spitz. Implicit estimation of ecological model parameters. Bulletin of Mathematical Biology, 75(2): 223–257, February 2013. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-012-9801-6;   http://link.springer.com/content/
pdf/10.1007/s11538-012-9801-6.pdf.

Berec:2013:FHE

[2774] Luděk Berec and Daniel Maxin. Fatal or harmless: Extreme bistability induced by sterilizing, sexually transmitted pathogens. Bulletin of Mathematical Biology, 75(2):258–273, February 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
012-9802-5;   http://link.springer.com/content/pdf/10.1007/
s11538-012-9802-5.pdf.

Zhao:2013:NSI

[2775] Gaiping Zhao, Wentao Yan, Eryun Chen, Xiaoli Yu, and Wenjie Cai. Numerical simulation of the inhibitory effect of angiostatin on metastatic tumor angiogenesis and microenvironment. Bulletin of Mathematical Biology, 75(2):274–287, February 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
012-9805-2;   http://link.springer.com/content/pdf/10.1007/
s11538-012-9805-2.pdf.

Allena:2013:CMM

[2776] Rachele Allena. Cell migration with multiple pseudopodia: Temporal and spatial sensing models. Bulletin of Mathematical Biology, 75(2): 288–316, February 2013. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-012-9806-1;   http://link.springer.com/content/
pdf/10.1007/s11538-012-9806-1.pdf.

Miekisz:2013:GES

[2777] Jacek Miekisz and Paulina Szymańska. Gene expression in self-repressing system with multiple gene copies. Bulletin of Mathematical Biology, 75 (2):317–330, February 2013. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-013-9808-7;   http://link.springer.com/content/
pdf/10.1007/s11538-013-9808-7.pdf.

Saini:2013:PCB

[2778] Ashish Saini and Jingyu Hou. Progressive clustering based method for protein function prediction. Bulletin of Mathematical Biology, 75(2): 331–350, February 2013. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-013-9809-6;   http://link.springer.com/content/
pdf/10.1007/s11538-013-9809-6.pdf.

Bokes:2013:TBD

[2779] Pavol Bokes, John R. King, Andrew T. A. Wood, and Matthew Loose. Transcriptional bursting diversifies the behaviour of a toggle switch: Hybrid simulation of stochastic gene expression. Bulletin of Mathematical Biology, 75(2):351–371, February 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9811-z;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9811-z.pdf.

Zakharova:2013:DRB

[2780] A. Zakharova, Z. Nikoloski, and A. Koseska. Dimensionality reduction of bistable biological systems. Bulletin of Mathematical Biology, 75 (3):373–392, March 2013. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-013-9807-8;   http://link.springer.com/content/
pdf/10.1007/s11538-013-9807-8.pdf.

Pohlmeyer:2013:MMG

[2781] J. V. Pohlmeyer, S. L. Waters, and L. J. Cummings. Mathematical model of growth factor driven haptotaxis and proliferation in a tissue engineering scaffold. Bulletin of Mathematical Biology, 75(3):393–427, March 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9810-0;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9810-0.pdf.

Hadfield:2013:MSM

[2782] James Hadfield, Michael J. Plank, and Tim David. Modeling secondary messenger pathways in neurovascular coupling. Bulletin of Mathematical Biology, 75(3):428–443, March 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-013-9813-x;
 http://link.springer.com/content/pdf/10.1007/s11538-013-9813-
x.pdf.

Huber:2013:LCR

[2783] Katharina T. Huber and Andrei-Alin Popescu. Lassoing and corralling rooted phylogenetic trees. Bulletin of Mathematical Biology, 75(3): 444–465, March 2013. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-013-9815-8;   http://link.springer.com/content/
pdf/10.1007/s11538-013-9815-8.pdf.

Rattana:2013:CPM

[2784] Prapanporn Rattana, Konstantin B. Blyuss, Ken T. D. Eames, and Istvan Z. Kiss. A class of pairwise models for epidemic dynamics on weighted networks. Bulletin of Mathematical Biology, 75(3):466–490, March 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9816-7;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9816-7.pdf.

Davis:2013:MMM

[2785] Courtney L. Davis and Frederick R. Adler. Mathematical models of memory CD8^{+} T-cell repertoire dynamics in response to viral infections. Bulletin of Mathematical Biology, 75(3):491–522, March 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9817-6;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9817-6.pdf.

Gao:2013:MSS

[2786] Daozhou Gao, Chris Cosner, Robert Stephen Cantrell, John C. Beier, and Shigui Ruan. Modeling the spatial spread of Rift valley fever in Egypt. Bulletin of Mathematical Biology, 75(3):523–542, March 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9818-5;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9818-5.pdf.

Wang:2013:WHV

[2787] Zhen Wang and Xiao-Qiang Zhao. A within-host virus model with periodic multidrug therapy. Bulletin of Mathematical Biology, 75(3): 543–563, March 2013. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-013-9820-y;   http://link.springer.com/content/
pdf/10.1007/s11538-013-9820-y.pdf.

Kareva:2013:PTC

[2788] Irina Kareva, Benjamin Morin, and Georgy Karev. Preventing the tragedy of the commons through punishment of over-consumers and encouragement of under-consumers. Bulletin of Mathematical Biology, 75 (4):565–588, April 2013. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-012-9804-3;   http://link.springer.com/content/
pdf/10.1007/s11538-012-9804-3.pdf.

Gatenby:2013:CRI

[2789] Robert A. Gatenby and B. Roy Frieden. The critical roles of information and nonequilibrium thermodynamics in evolution of living systems. Bulletin of Mathematical Biology, 75(4):589–601, April 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9821-x;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9821-x.pdf.

Antoneli:2013:VRP

[2790] Fernando Antoneli, Francisco Bosco, Diogo Castro, and Luiz Mario Janini. Virus replication as a phenotypic version of polynucleotide evolution. Bulletin of Mathematical Biology, 75(4):602–628, April 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9822-9;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9822-9.pdf.

Dumont:2013:SEI

[2791] Grégory Dumont and Jacques Henry. Synchronization of an excitatory integrate-and-fire neural network. Bulletin of Mathematical Biology, 75 (4):629–648, April 2013. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-013-9823-8;   http://link.springer.com/content/
pdf/10.1007/s11538-013-9823-8.pdf.

Zhao:2013:MIE

[2792] Yuqin Zhao, Dobromir T. Dimitrov, Hao Liu, and Yang Kuang. Mathematical insights in evaluating state dependent effectiveness of HIV prevention interventions. Bulletin of Mathematical Biology, 75(4):649–675, April 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9824-7;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9824-7.pdf.

Tawfik:2013:MNI

[2793] Yasmine Tawfik and Robert G. Owens. A mathematical and numerical investigation of the hemodynamical origins of oscillations in microvascular networks. Bulletin of Mathematical Biology, 75(4):676–707, April 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9825-6;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9825-6.pdf.

Lundh:2013:CFD

[2794] Torbjörn Lundh and Philip Gerlee. Cross-feeding dynamics described by a series expansion of the replicator equation. Bulletin of Mathematical Biology, 75(5):709–724, May 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-013-9828-3;
 http://link.springer.com/content/pdf/10.1007/s11538-013-9828-
3.pdf.

Yang:2013:PHV

[2795] Youping Yang, Yanni Xiao, and Jianhong Wu. Pulse HIV vaccination: Feasibility for virus eradication and optimal vaccination schedule. Bulletin of Mathematical Biology, 75(5):725–751, May 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9831-8;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9831-8.pdf.

Raba:2013:ECA

[2796] Ashley E. Raba, Jonathan M. Cordeiro, Charles Antzelevitch, and Jacques Beaumont. Extending the conditions of application of an inversion of the Hodgkin–Huxley gating model. Bulletin of Mathematical Biology, 75(5):752–773, May 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-013-9832-7;   http://link.springer.com/
content/pdf/10.1007/s11538-013-9832-7.pdf.

Stancevic:2013:TPD

[2797] O. Stancevic, C. N. Angstmann, J. M. Murray, and B. I. Henry. Turing patterns from dynamics of early HIV infection. Bulletin of Mathematical Biology, 75(5):774–795, May 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-013-9834-5;
 http://link.springer.com/content/pdf/10.1007/s11538-013-9834-
5.pdf.

Hughes:2013:MUW

[2798] Harriet Hughes and N. F. Britton. Modelling the use of Wolbachia to control dengue fever transmission. Bulletin of Mathematical Biology, 75(5):796–818, May 2013. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-013-9835-4;   http://link.springer.com/content/
pdf/10.1007/s11538-013-9835-4.pdf.

Bodaker:2013:MCI

[2799] Matan Bodaker, Yoram Louzoun, and Eduardo Mitrani. Mathematical conditions for induced cell differentiation and trans-differentiation in adult cells. Bulletin of Mathematical Biology, 75(5):819–844, May 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9837-2;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9837-2.pdf.

Cho:2013:SDD

[2800] Eunjoo Cho and Yong-Jung Kim. Starvation driven diffusion as a survival strategy of biological organisms. Bulletin of Mathematical Biology, 75(5):845–870, May 2013. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-013-9838-1;   http://link.springer.com/content/
pdf/10.1007/s11538-013-9838-1.pdf.

Simpson:2013:EMI

[2801] Matthew J. Simpson, Benjamin J. Binder, Parvathi Haridas, Benjamin K. Wood, Katrina K. Treloar, D. L. Sean McElwain, and Ruth E. Baker. Experimental and modelling investigation of monolayer development with clustering. Bulletin of Mathematical Biology, 75(5):871–889, May 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9839-0;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9839-0.pdf.

Chaouiya:2013:LMR

[2802] Claudine Chaouiya and Elisabeth Remy. Logical modelling of regulatory networks, methods and applications. Bulletin of Mathematical Biology, 75(6):891–895, June 2013. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-013-9863-0;   http://link.springer.com/content/
pdf/10.1007/s11538-013-9863-0.pdf.

Thomas:2013:RRR

[2803] R. Thomas. Remarks on the respective roles of logical parameters and time delays in asynchronous logic: an homage to El Houssine Snoussi. Bulletin of Mathematical Biology, 75(6):896–904, June 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9830-9;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9830-9.pdf. See erratum [2804].

Thomas:2013:ERR

[2804] R. Thomas. Erratum to: Remarks on the Respective Roles of Logical Parameters and Time Delays in Asynchronous Logic: an Homage to El Houssine Snoussi. Bulletin of Mathematical Biology, 75(6):905, June 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9843-4;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9843-4.pdf. See [2803].

Comet:2013:CFB

[2805] Jean-Paul Comet, Mathilde Noual, Adrien Richard, Julio Aracena, Laurence Calzone, Jacques Demongeot, Marcelle Kaufman, Aurélien Naldi, El Houssine Snoussi, and Denis Thieffry. On circuit functionality in Boolean networks. Bulletin of Mathematical Biology, 75(6):906–919, June 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9829-2;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9829-2.pdf.

Lorenz:2013:ACA

[2806] Therese Lorenz, Heike Siebert, and Alexander Bockmayr. Analysis and characterization of asynchronous state transition graphs using extremal states. Bulletin of Mathematical Biology, 75(6):920–938, June 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
012-9782-5;   http://link.springer.com/content/pdf/10.1007/
s11538-012-9782-5.pdf.

Goles:2013:DDR

[2807] Eric Goles, Marco Montalva, and Gonzalo A. Ruz. Deconstruction and dynamical robustness of regulatory networks: Application to the yeast cell cycle networks. Bulletin of Mathematical Biology, 75(6):939–966, June 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
012-9794-1;   http://link.springer.com/content/pdf/10.1007/
s11538-012-9794-1.pdf.

Chaves:2013:PAP

[2808] Madalena Chaves, Etienne Farcot, and Jean-Luc Gouzé. Probabilistic approach for predicting periodic orbits in piecewise affine differential models. Bulletin of Mathematical Biology, 75(6):967–987, June 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
012-9773-6;   http://link.springer.com/content/pdf/10.1007/
s11538-012-9773-6.pdf.

Madrahimov:2013:DIV

[2809] Alex Madrahimov, Tomáš Helikar, Bryan Kowal, Guoqing Lu, and Jim Rogers. Dynamics of influenza virus and human host interactions during infection and replication cycle. Bulletin of Mathematical Biology, 75 (6):988–1011, June 2013. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-012-9777-2;   http://link.springer.com/content/
pdf/10.1007/s11538-012-9777-2.pdf.

Mendoza:2013:VCC

[2810] Luis Mendoza. A virtual culture of CD4+ T lymphocytes. Bulletin of Mathematical Biology, 75(6):1012–1029, June 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-013-9814-9;
 http://link.springer.com/content/pdf/10.1007/s11538-013-9814-
9.pdf.

Artalejo:2013:EMD

[2811] J. R. Artalejo and M. J. Lopez-Herrero. On the exact measure of disease spread in stochastic epidemic models. Bulletin of Mathematical Biology, 75(7):1031–1050, July 2013. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-013-9836-3;   http://link.springer.com/content/
pdf/10.1007/s11538-013-9836-3.pdf.

Crawford:2013:ISC

[2812] Britnee A. Crawford, Christopher M. Kribs-Zaleta, and Gaik Ambartsoumian. Invasion speed in cellular automaton models for T. cruzi vector migration. Bulletin of Mathematical Biology, 75(7):1051–1081, July 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9840-7;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9840-7.pdf.

Gallo:2013:PGG

[2813] Ignacio Gallo. Population genetics of gene function. Bulletin of Mathematical Biology, 75(7):1082–1103, July 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-013-9841-6;
 http://link.springer.com/content/pdf/10.1007/s11538-013-9841-
6.pdf.

Safi:2013:DAM

[2814] Mohammad A. Safi, Dessalegn Y. Melesse, and Abba B. Gumel. Dynamics analysis of a multi-strain cholera model with an imperfect vaccine. Bulletin of Mathematical Biology, 75(7):1104–1137, July 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-

013-9845-2;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9845-2.pdf.

Wang:2013:INF

[2815] Yuanshi Wang and Hong Wu. Invasibility of nectarless flowers in plant– pollinator systems. Bulletin of Mathematical Biology, 75(7):1138–1156, July 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9846-1;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9846-1.pdf.

Lahodny:2013:PDO

[2816] Glenn E. Lahodny, Jr. and Linda J. S. Allen. Probability of a disease outbreak in stochastic multipatch epidemic models. Bulletin of Mathematical Biology, 75(7):1157–1180, July 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-013-9848-z;
 http://link.springer.com/content/pdf/10.1007/s11538-013-9848-
z.pdf.

Hamann:2013:TDI

[2817] Heiko Hamann, Istvan Karsai, and Thomas Schmickl. Time delay implies cost on task switching: a model to investigate the efficiency of task partitioning. Bulletin of Mathematical Biology, 75(7):1181–1206, July 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9851-4;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9851-4.pdf.

Hsu:2013:DPC

[2818] Sze-Bi Hsu, Chiu-Ju Lin, Chih-Hao Hsieh, and Kohei Yoshiyama. Dynamics of phytoplankton communities under photoinhibition. Bulletin of Mathematical Biology, 75(7):1207–1232, July 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-013-9852-3;
 http://link.springer.com/content/pdf/10.1007/s11538-013-9852-
3.pdf.

Alber:2013:BMR

[2819] Mark Alber, Philip K. Maini, and Glen Niebur. Biomedical modeling: The role of transport and mechanics. Bulletin of Mathematical Biology, 75(8):1233–1237, August 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-013-9873-y;
 http://link.springer.com/content/pdf/10.1007/s11538-013-9873-
y.pdf.

Thompson:2013:TME

[2820] Mark S. Thompson. Tendon mechanobiology: Experimental models require mathematical underpinning. Bulletin of Mathematical Biology, 75(8):1238–1254, August 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-013-9850-5;   http://link.springer.com/
content/pdf/10.1007/s11538-013-9850-5.pdf.

Leiderman:2013:IHT

[2821] Karin Leiderman and Aaron L. Fogelson. The influence of hindered transport on the development of platelet thrombi under flow. Bulletin of Mathematical Biology, 75(8):1255–1283, August 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
012-9784-3;   http://link.springer.com/content/pdf/10.1007/
s11538-012-9784-3.pdf.

Weafer:2013:ECI

[2822] P. P. Weafer, W. Ronan, S. P. Jarvis, and J. P. McGarry. Experimental and computational investigation of the role of stress fiber contractility in the resistance of osteoblasts to compression. Bulletin of Mathematical Biology, 75(8):1284–1303, August 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9812-y;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9812-y.pdf.

Kim:2013:HMT

[2823] Yangjin Kim and Hans G. Othmer. A hybrid model of tumor–stromal interactions in breast cancer. Bulletin of Mathematical Biology, 75(8): 1304–1350, August 2013. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-012-9787-0;   http://link.springer.com/content/
pdf/10.1007/s11538-012-9787-0.pdf.

Alonso:2013:NTS

[2824] Sergio Alonso, Markus Bär, and Alexander V. Panfilov. Negative tension of scroll wave filaments and turbulence in threedimensional excitable media and application in cardiac dynamics. Bulletin of Mathematical Biology, 75(8):1351–1376, August 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
012-9748-7;   http://link.springer.com/content/pdf/10.1007/
s11538-012-9748-7.pdf.

Daub:2013:CBM

[2825] Josephine T. Daub and Roeland M. H. Merks. A cell-based model of extracellular-matrix-guided endothelial cell migration during angiogenesis. Bulletin of Mathematical Biology, 75(8):1377–1399, August 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9826-5;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9826-5.pdf.

Suhail:2013:MIT

[2826] Yasir Suhail, Kshitiz, Justin Lee, Mark Walker, Deok-Ho Kim, Matthew D. Brennan, Joel S. Bader, and Andre Levchenko. Modeling intercellular transfer of biomolecules through tunneling nanotubes. Bulletin of Mathematical Biology, 75(8):1400–1416, August 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9819-4;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9819-4.pdf.

Lo:2013:MMR

[2827] Wing-Cheong Lo, Razvan I. Arsenescu, and Avner Friedman. Mathematical model of the roles of T cells in inflammatory bowel disease. Bulletin of Mathematical Biology, 75(9):1417–1433, September 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9853-2;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9853-2.pdf.

Caudill:2013:SPM

[2828] Lester F. Caudill, Jr. A single-parameter model of the immune response to bacterial invasion. Bulletin of Mathematical Biology, 75(9):1434– 1449, September 2013. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-013-9854-1;   http://link.springer.com/content/
pdf/10.1007/s11538-013-9854-1.pdf.

Forgoston:2013:PUE

[2829] Eric Forgoston and Ira B. Schwartz. Predicting unobserved exposures from seasonal epidemic data. Bulletin of Mathematical Biology, 75(9):1450–1471, September 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-013-9855-0;   http://link.springer.com/
content/pdf/10.1007/s11538-013-9855-0.pdf.

Lamb:2013:MMT

[2830] A. Lamb, G. Evans, and J. R. King. Mathematical modelling of toxicity associated with intracellular chromium reduction. Bulletin of Mathematical Biology, 75(9):1472–1500, September 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9856-z;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9856-z.pdf.

Chan:2013:MWI

[2831] Matthew H. T. Chan and Peter S. Kim. Modelling a Wolbachia invasion using a slow–fast dispersal reaction–diffusion approach. Bulletin of Mathematical Biology, 75(9):1501–1523, September 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9857-y;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9857-y.pdf.

Zubkov:2013:MTF

[2832] V. S. Zubkov, C. J. W. Breward, and E. A. Gaffney. Meniscal tear film fluid dynamics near Marx’s line. Bulletin of Mathematical Biology, 75(9):1524–1543, September 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-013-9858-x;   http://link.springer.com/
content/pdf/10.1007/s11538-013-9858-x.pdf.

Cattani:2013:FGD

[2833] Carlo Cattani and Gaetano Pierro. On the fractal geometry of DNA by the binary image analysis. Bulletin of Mathematical Biology, 75(9):1544– 1570, September 2013. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-013-9859-9;   http://link.springer.com/content/
pdf/10.1007/s11538-013-9859-9.pdf.

Curto:2013:NRA

[2834] Carina Curto, Vladimir Itskov, Alan Veliz-Cuba, and Nora Youngs. The neural ring: an algebraic tool for analyzing the intrinsic structure of neural codes. Bulletin of Mathematical Biology, 75(9):1571–1611, September 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9860-3;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9860-3.pdf.

Buchanan:2013:KVD

[2835] James L. Buchanan, Robert Gilbert, Yvonne Ou, Anja Nohe, and Rachel Schaefer. The kinetics of vitamin D_{3} in the osteoblastic cell. Bulletin of Mathematical Biology, 75(9):1612–1635, September 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9861-2;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9861-2.pdf.

Saa:2013:MAP

[2836] Alberto Saa and Kellen M. Siqueira. Modeling the ATP production in mitochondria. Bulletin of Mathematical Biology, 75(9):1636–1651, September 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9862-1;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9862-1.pdf.

Basile:2013:GBA

[2837] Raffaele Basile, Ramon Grima, and Nikola Popović. A graph-based approach for the approximate solution of the chemical master equation. Bulletin of Mathematical Biology, 75(10):1653–1696, October 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9864-z;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9864-z.pdf.

Nie:2013:SDP

[2838] Lin-Fei Nie, Zhi-Dong Teng, and Bao-Zhu Guo. A state dependent pulse control strategy for a SIRS epidemic system. Bulletin of Mathematical Biology, 75(10):1697–1715, October 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-013-9865-y;
 http://link.springer.com/content/pdf/10.1007/s11538-013-9865-
y.pdf.

Hernandez-Ceron:2013:DEM

[2839] Nancy Hernandez-Ceron, Zhilan Feng, and Carlos Castillo-Chavez. Discrete epidemic models with arbitrary stage distributions and applications to disease control. Bulletin of Mathematical Biology, 75(10): 1716–1746, October 2013. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-013-9866-x;   http://link.springer.com/content/
pdf/10.1007/s11538-013-9866-x.pdf.

Clarke:2013:AOC

[2840] James Clarke, K. A. Jane White, and Katy Turner. Approximating optimal controls for networks when there are combinations of populationlevel and targeted measures available: Chlamydia infection as a casestudy. Bulletin of Mathematical Biology, 75(10):1747–1777, October 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9867-9;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9867-9.pdf.

Strohm:2013:PFM

[2841] S. Strohm, R. C. Tyson, and J. A. Powell. Pattern formation in a model for mountain pine beetle dispersal: Linking model predictions to data. Bulletin of Mathematical Biology, 75(10):1778–1797, October 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9868-8;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9868-8.pdf.

Safuan:2013:IBR

[2842] H. M. Safuan, H. S. Sidhu, Z. Jovanoski, and I. N. Towers. Impacts of biotic resource enrichment on a predator–prey population. Bulletin of Mathematical Biology, 75(10):1798–1812, October 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9869-7;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9869-7.pdf.

Brown:2013:LMP

[2843] David Brown. Linking molecular and population processes in mathematical models of quorum sensing. Bulletin of Mathematical Biology, 75(10): 1813–1839, October 2013. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-013-9870-1;   http://link.springer.com/content/
pdf/10.1007/s11538-013-9870-1.pdf.

Willson:2013:RCP

[2844] Stephen J. Willson. Reconstruction of certain phylogenetic networks from their tree-average distances. Bulletin of Mathematical Biology, 75(10): 1840–1878, October 2013. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-013-9872-z;   http://link.springer.com/content/
pdf/10.1007/s11538-013-9872-z.pdf.

Humphries:2013:CPC

[2845] Peter J. Humphries, Simone Linz, and Charles Semple. Cherry picking: a characterization of the temporal hybridization number for a set of phylogenies. Bulletin of Mathematical Biology, 75(10):1879–1890, October 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9874-x;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9874-x.pdf.

Wilson:2013:FSS

[2846] Shelby Wilson and Doron Levy. Functional switching and stability of regulatory t cells. Bulletin of Mathematical Biology, 75(10):1891–1911, October 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9875-9;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9875-9.pdf.

Planque:2013:IBS

[2847] Robert Planqué, Jan Bouwe van den Berg, and Nigel R. Franks. The interplay between scent trails and group-mass recruitment systems in ants. Bulletin of Mathematical Biology, 75(10):1912–1940, October 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9876-8;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9876-8.pdf.

Chen:2013:REB

[2848] Xingjiang Chen, Karl Iremonger, Allan Herbison, Vivien Kirk, and James Sneyd. Regulation of electrical bursting in a spatiotemporal model of a GnRH neuron. Bulletin of Mathematical Biology, 75(10):1941–1960, October 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9877-7;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9877-7.pdf.

Reluga:2013:EEG

[2849] Timothy C. Reluga. Equilibria of an epidemic game with piecewise linear social distancing cost. Bulletin of Mathematical Biology, 75(10): 1961–1984, October 2013. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-013-9879-5;   http://link.springer.com/content/
pdf/10.1007/s11538-013-9879-5.pdf.

Winstanley:2013:BDS

[2850] H. F. Winstanley and A. C. Fowler. Biomat development in soil treatment units for on-site wastewater treatment. Bulletin of Mathematical Biology, 75(10):1985–2001, October 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-013-9881-y;
 http://link.springer.com/content/pdf/10.1007/s11538-013-9881-
y.pdf.

Kluth:2013:CAP

[2851] Sandra Kluth, Thiemo Hustedt, and Ellen Baake. The common ancestor process revisited. Bulletin of Mathematical Biology, 75(11):2003– 2027, November 2013. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-013-9871-0;   http://link.springer.com/content/
pdf/10.1007/s11538-013-9871-0.pdf.

Holstein:2013:MSD

[2852] Katharina Holstein, Dietrich Flockerzi, and Carsten Conradi. Multistationarity in sequential distributed multisite phosphorylation networks. Bulletin of Mathematical Biology, 75(11):2028–2058, November 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9878-6;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9878-6.pdf.

Bate:2013:CDE

[2853] Andrew M. Bate and Frank M. Hilker. Complex dynamics in an ecoepidemiological model. Bulletin of Mathematical Biology, 75(11):2059– 2078, November 2013. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-013-9880-z;   http://link.springer.com/content/
pdf/10.1007/s11538-013-9880-z.pdf.

Naasell:2013:IIL

[2854] Ingemar Nåsell. The influence of immunity loss on persistence and recurrence of endemic infections. Bulletin of Mathematical Biology, 75(11): 2079–2092, November 2013. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-013-9882-x;   http://link.springer.com/content/
pdf/10.1007/s11538-013-9882-x.pdf.

Isaacson:2013:ISV

[2855] Samuel A. Isaacson, Carolyn A. Larabell, Mark A. Le Gros, David M. McQueen, and Charles S. Peskin. The influence of spatial variation in chromatin density determined by X-ray tomograms on the time to find DNA binding sites. Bulletin of Mathematical Biology, 75(11):2093– 2117, November 2013. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-013-9883-9;   http://link.springer.com/content/
pdf/10.1007/s11538-013-9883-9.pdf.

Mirzaev:2013:LDG

[2856] Inomzhon Mirzaev and Jeremy Gunawardena. Laplacian dynamics on general graphs. Bulletin of Mathematical Biology, 75(11):2118– 2149, November 2013. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-013-9884-8;   http://link.springer.com/content/
pdf/10.1007/s11538-013-9884-8.pdf.

Plank:2013:LFM

[2857] Michael J. Plank and Matthew J. Simpson. Lattice-free models of cell invasion: Discrete simulations and travelling waves. Bulletin of Mathematical Biology, 75(11):2150–2166, November 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9885-7;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9885-7.pdf.

Liang:2013:ARN

[2858] Juhua Liang, Sanyi Tang, Robert A. Cheke, and Jianhong Wu. Adaptive release of natural enemies in a pest–natural enemy system with pesticide resistance. Bulletin of Mathematical Biology, 75(11):2167–2195, November 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9886-6;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9886-6.pdf.

McQuaid:2013:TSS

[2859] C. Finn McQuaid and Nicholas F. Britton. Trophic structure, stability, and parasite persistence threshold in food Webs. Bulletin of Mathematical Biology, 75(11):2196–2207, November 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9887-5;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9887-5.pdf.

Stefanescu:2013:CMC

[2860] Roxana A. Stefanescu, R. G. Shivakeshavan, Pramod P. Khargonekar, and Sachin S. Talathi. Computational modeling of channelrhodopsin-2 photocurrent characteristics in relation to neural signaling. Bulletin of Mathematical Biology, 75(11):2208–2240, November 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9888-4;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9888-4.pdf.

Best:2013:ESF

[2861] Alex Best. The effects of seasonal forcing on invertebrate-disease interactions with immune priming. Bulletin of Mathematical Biology, 75(11): 2241–2256, November 2013. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-013-9889-3;   http://link.springer.com/content/
pdf/10.1007/s11538-013-9889-3.pdf.

Mallet:2013:MMC

[2862] Dann G. Mallet, Masoumeh Bagher-Oskouei, A. Charisse Farr, Daniel P. Simpson, and Kelly-Jean Sutton. A mathematical model of chlamydial infection incorporating movement of chlamydial particles. Bulletin of Mathematical Biology, 75(11):2257–2270, November 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9891-9;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9891-9.pdf.

Daub:2013:ABI

[2863] Markus Daub. An appropriate bounded invariant region for a bistable reaction–diffusion model of the Caspase–3/8 feedback loop. Bulletin of Mathematical Biology, 75(11):2271–2288, November 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9892-8;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9892-8.pdf.

Soliman:2013:SNC

[2864] Sylvain Soliman. A stronger necessary condition for the multistationarity of chemical reaction networks. Bulletin of Mathematical Biology, 75(11): 2289–2303, November 2013. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-013-9893-7;   http://link.springer.com/content/
pdf/10.1007/s11538-013-9893-7.pdf.

Cohen:2013:EDC

[2865] Yosef Cohen and Gonzalo Galiano. Evolutionary distributions and competition by way of reaction–diffusion and by way of convolution. Bulletin of Mathematical Biology, 75(12):2305–2323, December 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9890-x;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9890-x.pdf.

Lucas:2013:PMC

[2866] Timothy A. Lucas, Garrett Johns, Wancen Jiang, and Lucie Yang. A population model of chaparral vegetation response to frequent wildfires. Bulletin of Mathematical Biology, 75(12):2324–2345, December 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9894-6;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9894-6.pdf.

Nagata:2013:RDP

[2867] Wayne Nagata, Hamid R. Z. Zangeneh, and David M. Holloway. Reaction–diffusion patterns in plant tip morphogenesis: Bifurcations on spherical caps. Bulletin of Mathematical Biology, 75(12):2346– 2371, December 2013. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-013-9895-5;   http://link.springer.com/content/
pdf/10.1007/s11538-013-9895-5.pdf.

McQuaid:2013:NDC

[2868] C. Finn McQuaid and Nicholas F. Britton. Network dynamics contribute to structure: Nestedness in mutualistic networks. Bulletin of Mathematical Biology, 75(12):2372–2388, December 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9896-4;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9896-4.pdf.

Yun:2013:PFM

[2869] Ana Yun, Soon-Hyuck Lee, and Junseok Kim. A phase-field model for articular cartilage regeneration in degradable scaffolds. Bulletin of Mathematical Biology, 75(12):2389–2409, December 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9897-3;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9897-3.pdf.

Clote:2013:ANH

[2870] Peter Clote, Evangelos Kranakis, and Danny Krizanc. Asymptotic number of hairpins of saturated RNA secondary structures. Bulletin of Mathematical Biology, 75(12):2410–2430, December 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9899-1;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9899-1.pdf.

Vey:2013:GCH

[2871] Gregory Vey. Gene coexpression as Hebbian learning in prokaryotic genomes. Bulletin of Mathematical Biology, 75(12):2431–2449, December 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9900-z;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9900-z.pdf.

Pohlmeyer:2013:CLG

[2872] J. V. Pohlmeyer and L. J. Cummings. Cyclic loading of growing tissue in a bioreactor: Mathematical model and asymptotic analysis. Bulletin of Mathematical Biology, 75(12):2450–2473, December 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9902-x;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9902-x.pdf.

Kim:2013:EDT

[2873] Yong-Jung Kim, Ohsang Kwon, and Fang Li. Evolution of dispersal toward fitness. Bulletin of Mathematical Biology, 75(12):2474– 2498, December 2013. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-013-9904-8;   http://link.springer.com/content/
pdf/10.1007/s11538-013-9904-8.pdf.

Mandal:2013:RHV

[2874] Sandip Mandal, Somdatta Sinha, and Ram Rup Sarkar. A realistic host–vector transmission model for describing malaria prevalence pattern. Bulletin of Mathematical Biology, 75(12):2499–2528, December 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9905-7;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9905-7.pdf.

Bouchard-Cote:2013:NPM

[2875] Alexandre Bouchard-Côté. A note on probabilistic models over strings: The linear algebra approach. Bulletin of Mathematical Biology, 75(12): 2529–2550, December 2013. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-013-9906-6;   http://link.springer.com/content/
pdf/10.1007/s11538-013-9906-6.pdf.

Sgouralis:2013:CMF

[2876] Ioannis Sgouralis and Anita T. Layton. Control and modulation of fluid flow in the rat kidney. Bulletin of Mathematical Biology, 75(12): 2551–2574, December 2013. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-013-9907-5;   http://link.springer.com/content/
pdf/10.1007/s11538-013-9907-5.pdf.

Allena:2013:MIB

[2877] R. Allena, D. Aubry, and J. Sharpe. On the mechanical interplay between intra- and inter-synchronization during collective cell migration: A numerical investigation. Bulletin of Mathematical Biology, 75(12): 2575–2599, December 2013. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-013-9908-4;   http://link.springer.com/content/
pdf/10.1007/s11538-013-9908-4.pdf.

Innocentini:2013:MFS

[2878] Guilherme da Costa Pereira Innocentini, Michael Forger, Alexandre Ferreira Ramos, Ovidiu Radulescu, and José Eduardo Martinho Hornos. Multimodality and flexibility of stochastic gene expression. Bulletin of Mathematical Biology, 75(12):2600–2630, December 2013. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9909-3;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9909-3.pdf.

Sherratt:2014:MBG

[2879] Jonathan A. Sherratt, Ayawoa S. Dagbovie, and Frank M. Hilker. A mathematical Biologist’s guide to absolute and convective instability. Bulletin of Mathematical Biology, 76(1):1–26, January 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9911-9;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9911-9.pdf.

Masic:2014:MSS

[2880] Alma Mašić and Hermann J. Eberl. A modeling and simulation study of the role of suspended microbial populations in nitrification in a biofilm reactor. Bulletin of Mathematical Biology, 76(1):27–58, January 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9898-2;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9898-2.pdf.

Miron:2014:RPI

[2881] Rachelle E. Miron and Robert J. Smith. Resistance to protease inhibitors in a model of HIV-1 infection with impulsive drug effects. Bulletin of Mathematical Biology, 76(1):59–97, January 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9903-9;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9903-9.pdf.

Friedman:2014:BBM

[2882] Avner Friedman and Abdul-Aziz Yakubu. A Bovine Babesiosis model with dispersion. Bulletin of Mathematical Biology, 76(1):98–135, January 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-

013-9912-8;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9912-8.pdf.

Hendrix:2014:MEC

[2883] Angelean O. Hendrix, Claude L. Hughes, and James F. Selgrade. Modeling endocrine control of the pituitary–ovarian axis: Androgenic influence and chaotic dynamics. Bulletin of Mathematical Biology, 76(1): 136–156, January 2014. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-013-9913-7;   http://link.springer.com/content/
pdf/10.1007/s11538-013-9913-7.pdf.

Holmes:2014:ENS

[2884] William R. Holmes. An efficient, nonlinear stability analysis for detecting pattern formation in reaction diffusion systems. Bulletin of Mathematical Biology, 76(1):157–183, January 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-013-9914-6;
 http://link.springer.com/content/pdf/10.1007/s11538-013-9914-
6.pdf.

Jost:2014:OPA

[2885] Jürgen Jost and Ying Wang. Optimization and phenotype allocation. Bulletin of Mathematical Biology, 76(1):184–200, January 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9915-5;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9915-5.pdf.

Hordijk:2014:ASB

[2886] Wim Hordijk, Peter R. Wills, and Mike Steel. Autocatalytic sets and biological specificity. Bulletin of Mathematical Biology, 76(1):201–224, January 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9916-4;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9916-4.pdf.

Aviram:2014:BLP

[2887] Ira Aviram and Avinoam Rabinovitch. Bactria and lytic phage coexistence in a chemostat with periodic nutrient supply. Bulletin of Mathematical Biology, 76(1):225–244, January 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-013-9917-3;

 http://link.springer.com/content/pdf/10.1007/s11538-013-9917-
3.pdf.

Ma:2014:EIE

[2888] Junling Ma, Jonathan Dushoff, Benjamin M. Bolker, and David J. D. Earn. Estimating initial epidemic growth rates. Bulletin of Mathematical Biology, 76(1):245–260, January 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-013-9918-2;
 http://link.springer.com/content/pdf/10.1007/s11538-013-9918-
2.pdf.

Lam:2014:ESC

[2889] King-Yeung Lam and Yuan Lou. Evolutionarily stable and convergent stable strategies in reaction–diffusion models for conditional dispersal. Bulletin of Mathematical Biology, 76(2):261–291, February 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9901-y;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9901-y.pdf.

Camacho:2014:OCT

[2890] E. T. Camacho, L. A. Melara, M. C. Villalobos, and S. Wirkus. Optimal control in the treatment of Retinitis pigmentosa. Bulletin of Mathematical Biology, 76(2):292–313, February 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-013-9919-1;
 http://link.springer.com/content/pdf/10.1007/s11538-013-9919-
1.pdf.

Salazar-Cavazos:2014:OPT

[2891] Emanuel Salazar-Cavazos and Moisés Santillán. Optimal performance of the Tryptophan operon of E. coli: a stochastic, dynamical, mathematicalmodeling approach. Bulletin of Mathematical Biology, 76(2):314– 334, February 2014. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-013-9920-8;   http://link.springer.com/content/
pdf/10.1007/s11538-013-9920-8.pdf.

Liao:2014:ETD

[2892] Kang-Ling Liao and Yuan Lou. The effect of time delay in a two-patch model with random dispersal. Bulletin of Mathematical Biology, 76(2): 335–376, February 2014. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-013-9921-7;   http://link.springer.com/content/
pdf/10.1007/s11538-013-9921-7.pdf.

Franz:2014:TWH

[2893] Benjamin Franz, Chuan Xue, Kevin J. Painter, and Radek Erban. Travelling waves in hybrid chemotaxis models. Bulletin of Mathematical Biology, 76(2):377–400, February 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-013-9924-4;
 http://link.springer.com/content/pdf/10.1007/s11538-013-9924-
4.pdf.

Roesch:2014:MLT

[2894] Katja Roesch, Dirk Hasenclever, and Markus Scholz. Modelling lymphoma therapy and outcome. Bulletin of Mathematical Biology, 76(2): 401–430, February 2014. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-013-9925-3;   http://link.springer.com/content/
pdf/10.1007/s11538-013-9925-3.pdf.

Bruce:2014:MEG

[2895] Doug Bruce, Pras Pathmanathan, and Jonathan P. Whiteley. Modelling the effect of gap junctions on tissue-level cardiac electrophysiology. Bulletin of Mathematical Biology, 76(2):431–454, February 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9927-1;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9927-1.pdf.

Bratby:2014:BMA

[2896] Peter Bratby, John Montgomery, and James Sneyd. A biophysical model of adaptive noise filtering in the shark brain. Bulletin of Mathematical Biology, 76(2):455–475, February 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-013-9928-0;
 http://link.springer.com/content/pdf/10.1007/s11538-013-9928-
0.pdf.

Chen:2014:RSP

[2897] Shu-Yan Chen, Zi long Chen, Peng Guo, Chen-Chen Ding, Yu xin Wang, Xiang tai Wang, Jia-Lin Zhang, Peng Jia, Gang Wang, and Sa Xiao.

The roles of spatial pattern and size variation in shaping height inequality of plant population. Bulletin of Mathematical Biology, 76(2): 476–485, February 2014. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-014-9933-y;   http://link.springer.com/content/
pdf/10.1007/s11538-014-9933-y.pdf.

Immonen:2014:MHD

[2898] Taina Immonen, Erkki Somersalo, and Daniela Calvetti. Modeling HIV-1 dynamics and fitness in cell culture across scales. Bulletin of Mathematical Biology, 76(2):486–514, February 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9926-2;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9926-2.pdf.

Lacitignola:2014:PRD

[2899] Deborah Lacitignola and Giuseppe Saccomandi. Parametric resonance in DNA. Bulletin of Mathematical Biology, 76(3):515–540, March 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9930-6;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9930-6.pdf.

Lu:2014:SMI

[2900] Zhao Lu, Yrjö T. Gröhn, Rebecca L. Smith, Jeffrey S. Karns, Ernest Hovingh, and Ynte H. Schukken. Stochastic modeling of imperfect Salmonella vaccines in an adult dairy herd. Bulletin of Mathematical Biology, 76 (3):541–565, March 2014. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-013-9931-5;   http://link.springer.com/content/
pdf/10.1007/s11538-013-9931-5.pdf.

Payvandi:2014:MMW

[2901] S. Payvandi, K. R. Daly, D. L. Jones, P. Talboys, K. C. Zygalakis, and T. Roose. A mathematical model of water and nutrient transport in xylem vessels of a wheat plant. Bulletin of Mathematical Biology, 76 (3):566–596, March 2014. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-013-9932-4;   http://link.springer.com/content/
pdf/10.1007/s11538-013-9932-4.pdf.

Dinh:2014:EDD

[2902] Vu Dinh, Ann E. Rundell, and Gregery T. Buzzard. Experimental design for dynamics identification of cellular processes. Bulletin of Mathematical Biology, 76(3):597–626, March 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-014-9935-9;
 http://link.springer.com/content/pdf/10.1007/s11538-014-9935-
9.pdf.

Greene:2014:ICD

[2903] James Greene, Orit Lavi, Michael M. Gottesman, and Doron Levy. The impact of cell density and mutations in a model of multidrug resistance in solid tumors. Bulletin of Mathematical Biology, 76(3):627–653, March 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-9936-8;   http://link.springer.com/content/pdf/10.1007/
s11538-014-9936-8.pdf.

Fontanari:2014:EMD

[2904] José F. Fontanari and Maurizio Serva. Effect of migration in a diffusion model for template coexistence in protocells. Bulletin of Mathematical Biology, 76(3):654–672, March 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-014-9937-7;
 http://link.springer.com/content/pdf/10.1007/s11538-014-9937-
7.pdf.

Kajunguri:2014:MUI

[2905] Damian Kajunguri, John W. Hargrove, Rachid Ouifki, J. Y. T. Mugisha, Paul G. Coleman, and Susan C. Welburn. Modelling the use of insecticide-treated cattle to control tsetse and Trypanosoma brucei rhodesiense in a multi-host population. Bulletin of Mathematical Biology, 76 (3):673–696, March 2014. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-014-9938-6;   http://link.springer.com/content/
pdf/10.1007/s11538-014-9938-6.pdf.

Amaku:2014:CAR

[2906] Marcos Amaku, Francisco Antonio Bezerra Coutinho, Silvia Martorano Raimundo, Luis Fernandez Lopez, Marcelo Nascimento Burattini, and Eduardo Massad. A comparative analysis of the relative efficacy of vector-control strategies against dengue fever. Bulletin of Mathematical Biology, 76(3):697–717, March 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-014-9939-5;
 http://link.springer.com/content/pdf/10.1007/s11538-014-9939-
5.pdf.

Embleton:2014:NAE

[2907] Nina Embleton and Natalia Petrovskaya. A novel approach to evaluation of pest insect abundance in the presence of noise. Bulletin of Mathematical Biology, 76(3):718–743, March 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-014-9940-z;
 http://link.springer.com/content/pdf/10.1007/s11538-014-9940-
z.pdf.

GhasemiV:2014:ASA

[2908] A. Ghasemi V., B. Firoozabadi, and M. S. Saidi. Analytical solutions of actin–retrograde-flow in a circular stationary cell: a mechanical point of view. Bulletin of Mathematical Biology, 76(3):744–760, March 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-9941-y;   http://link.springer.com/content/pdf/10.1007/
s11538-014-9941-y.pdf.

Erban:2014:ESI

[2909] Radek Erban and Hans G. Othmer. Editorial: Special issue on stochastic modelling of reaction–diffusion processes in biology. Bulletin of Mathematical Biology, 76(4):761–765, April 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-013-9929-z;
 http://link.springer.com/content/pdf/10.1007/s11538-013-9929-
z.pdf.

Sturrock:2014:RDN

[2910] Marc Sturrock, Andreas Hellander, Sahar Aldakheel, Linda Petzold, and Mark A. J. Chaplain. The role of dimerisation and nuclear transport in the Hes1 gene regulatory network. Bulletin of Mathematical Biology, 76 (4):766–798, April 2014. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-013-9842-5;   http://link.springer.com/content/
pdf/10.1007/s11538-013-9842-5.pdf.

Erban:2014:MSR

[2911] Radek Erban, Mark B. Flegg, and Garegin A. Papoian. Multiscale stochastic reaction–diffusion modeling: Application to actin dynamics in Filopodia. Bulletin of Mathematical Biology, 76(4):799–818, April 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9844-3;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9844-3.pdf.

Wang:2014:SRD

[2912] Siyang Wang, Johan Elf, Stefan Hellander, and Per Lötstedt. Stochastic reaction–diffusion processes with embedded lower-dimensional structures. Bulletin of Mathematical Biology, 76(4):819–853, April 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9910-x;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9910-x.pdf.

Hu:2014:SAR

[2913] Jifeng Hu, Hye-Won Kang, and Hans G. Othmer. Stochastic analysis of reaction–diffusion processes. Bulletin of Mathematical Biology, 76 (4):854–894, April 2014. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-013-9849-y;   http://link.springer.com/content/
pdf/10.1007/s11538-013-9849-y.pdf.

McKane:2014:SPF

[2914] Alan J. McKane, Tommaso Biancalani, and Tim Rogers. Stochastic pattern formation and spontaneous polarisation: The linear noise approximation and beyond. Bulletin of Mathematical Biology, 76(4):895–921, April 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9827-4;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9827-4.pdf.

Agbanusi:2014:CBR

[2915] I. C. Agbanusi and S. A. Isaacson. A comparison of bimolecular reaction models for stochastic reaction–diffusion systems. Bulletin of Mathematical Biology, 76(4):922–946, April 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-013-9833-6;
 http://link.springer.com/content/pdf/10.1007/s11538-013-9833-
6.pdf.

Bruna:2014:DFS

[2916] Maria Bruna and S. Jonathan Chapman. Diffusion of finite-size particles in confined geometries. Bulletin of Mathematical Biology, 76(4):947–982, April 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
013-9847-0;   http://link.springer.com/content/pdf/10.1007/
s11538-013-9847-0.pdf. See erratum [2917].

Bruna:2014:EDF

[2917] Maria Bruna and S. Jonathan Chapman. Erratum to: Diffusion of Finite-Size Particles in Confined Geometries. Bulletin of Mathematical Biology, 76(4):983, April 2014. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-014-9953-7;   http://link.springer.com/content/
pdf/10.1007/s11538-014-9953-7.pdf. See [2916].

Britton:2014:SEG

[2918] Tom Britton and Pieter Trapman. Stochastic epidemics in growing populations. Bulletin of Mathematical Biology, 76(5):985–996, May 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-9942-x;   http://link.springer.com/content/pdf/10.1007/
s11538-014-9942-x.pdf.

Gavina:2014:MPM

[2919] Maica Krizna A. Gavina, Jomar F. Rabajante, and Cleofas R. Cervancia. Mathematical programming models for determining the optimal location of beehives. Bulletin of Mathematical Biology, 76(5):997–1016, May 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-9943-9;   http://link.springer.com/content/pdf/10.1007/
s11538-014-9943-9.pdf.

Alfonso:2014:SRD

[2920] J. C. L. Alfonso, G. Buttazzo, B. García-Archilla, M. A. Herrero, and L. Núñez. Selecting radiotherapy dose distributions by means of constrained optimization problems. Bulletin of Mathematical Biology, 76 (5):1017–1044, May 2014. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-014-9945-7;   http://link.springer.com/content/
pdf/10.1007/s11538-014-9945-7.pdf.

Lee:2014:MRT

[2921] Seongwon Lee, Hyung Ju Hwang, and Yangjin Kim. Modeling the role of TGF-β in regulation of the Th17 phenotype in the LPS-driven immune system. Bulletin of Mathematical Biology, 76(5):1045–1080, May 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-9946-6;   http://link.springer.com/content/pdf/10.1007/
s11538-014-9946-6.pdf.

Johnston:2014:TCR

[2922] Matthew D. Johnston. Translated chemical reaction networks. Bulletin of Mathematical Biology, 76(5):1081–1116, May 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-014-9947-5;
 http://link.springer.com/content/pdf/10.1007/s11538-014-9947-
5.pdf.

Cohen:2014:APE

[2923] Anna Cohen, Mary R. Myerscough, and Rosemary S. Thompson. Atheroprotective effects of high density lipoproteins (HDL): an ODE model of the early stages of atherosclerosis. Bulletin of Mathematical Biology, 76 (5):1117–1142, May 2014. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-014-9948-4;   http://link.springer.com/content/
pdf/10.1007/s11538-014-9948-4.pdf.

Keegan:2014:ACF

[2924] Lindsay Keegan and Jonathan Dushoff. Analytic calculation of finitepopulation reproductive numbers for direct- and vector-transmitted diseases with homogeneous mixing. Bulletin of Mathematical Biology, 76 (5):1143–1154, May 2014. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-014-9950-x;   http://link.springer.com/content/
pdf/10.1007/s11538-014-9950-x.pdf.

Tonnelier:2014:CNF

[2925] Arnaud Tonnelier. Cyclic negative feedback systems: What is the chance of oscillation? Bulletin of Mathematical Biology, 76(5):1155–1193, May 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-9959-1;   http://link.springer.com/content/pdf/10.1007/
s11538-014-9959-1.pdf.

Shan:2014:DGI

[2926] Chunhua Shan, Xiaonong Zhou, and Huaiping Zhu. The dynamics of growing islets and transmission of Schistosomiasis japonica in the Yangtze River. Bulletin of Mathematical Biology, 76(5):1194–1217, May 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-9961-7;   http://link.springer.com/content/pdf/10.1007/
s11538-014-9961-7.pdf.

Adamson:2014:BAM

[2927] M. W. Adamson and A. Yu. Morozov. Bifurcation analysis of models with uncertain function specification: How should we proceed? Bulletin of Mathematical Biology, 76(5):1218–1240, May 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-9951-9;   http://link.springer.com/content/pdf/10.1007/
s11538-014-9951-9.pdf.

Qi:2014:SSN

[2928] Qi Qi, Jonathan A. D. Wattis, and Helen M. Byrne. Stochastic simulations of normal aging and Werner’s syndrome. Bulletin of Mathematical Biology, 76(6):1241–1269, June 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-014-9952-8;
 http://link.springer.com/content/pdf/10.1007/s11538-014-9952-
8.pdf.

Goede:2014:HPT

[2929] Simon L. Goede, Melvin Khee-Shing Leow, Jan W. A. Smit, Harald H. Klein, and Johannes W. Dietrich. Hypothalamus–pituitary– thyroid feedback control: Implications of mathematical modeling and consequences for thyrotropin (TSH) and free thyroxine (FT4) reference ranges. Bulletin of Mathematical Biology, 76(6):1270–1287, June 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-9955-5;   http://link.springer.com/content/pdf/10.1007/
s11538-014-9955-5.pdf.

Mello:2014:SDM

[2930] Renato Francisco Lopes Mello and César Castilho. A structured discrete model for dengue fever infections and the determination of R_{0} from age-stratified serological data. Bulletin of Mathematical Biology, 76 (6):1288–1305, June 2014. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-014-9956-4;   http://link.springer.com/content/
pdf/10.1007/s11538-014-9956-4.pdf.

delosReyesV:2014:SCP

[2931] Aurelio A. de los Reyes V., Eunok Jung, and Franz Kappel. Stabilizing control for a pulsatile cardiovascular mathematical model. Bulletin of Mathematical Biology, 76(6):1306–1332, June 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-014-9958-2;
 http://link.springer.com/content/pdf/10.1007/s11538-014-9958-
2.pdf.

Sriyudthsak:2014:PSM

[2932] Kansuporn Sriyudthsak, Michio Iwata, Masami Yokota Hirai, and Fumihide Shiraishi. PENDISC: a simple method for constructing a mathematical model from time-series data of metabolite concentrations. Bulletin of Mathematical Biology, 76(6):1333–1351, June 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-9960-8;   http://link.springer.com/content/pdf/10.1007/
s11538-014-9960-8.pdf.

Leander:2014:MCR

[2933] R. Leander and A. Friedman. Modulation of the cAMP response by Gα_{i} and Gβγ: a computational study of G protein signaling in immune cells. Bulletin of Mathematical Biology, 76(6):1352–1375, June 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-9964-4;   http://link.springer.com/content/pdf/10.1007/
s11538-014-9964-4.pdf. See correction [3625].

Edgington:2014:FCD

[2934] Matthew P. Edgington and Marcus J. Tindall. Fold-change detection in a whole-pathway model of Escherichia coli chemotaxis. Bulletin of Mathematical Biology, 76(6):1376–1395, June 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-014-9965-3;
 http://link.springer.com/content/pdf/10.1007/s11538-014-9965-
3.pdf.

Wang:2014:CST

[2935] Yuanshi Wang and Hong Wu. A cooperative system of two species with bidirectional interactions. Bulletin of Mathematical Biology, 76 (6):1396–1415, June 2014. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-014-9966-2;   http://link.springer.com/content/
pdf/10.1007/s11538-014-9966-2.pdf.

Li:2014:SSM

[2936] Wei-Shen Li and Yuan-Zhi Shao. The scaling and shift of morphogen gene expression boundary in a nonlinear reaction diffusion system. Bulletin of Mathematical Biology, 76(6):1416–1428, June 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-9967-1;   http://link.springer.com/content/pdf/10.1007/
s11538-014-9967-1.pdf.

Merkey:2014:MII

[2937] Brian V. Merkey and David L. Chopp. Modeling the impact of interspecies competition on performance of a microbial fuel cell. Bulletin of Mathematical Biology, 76(6):1429–1453, June 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-9968-0;   http://link.springer.com/content/pdf/10.1007/
s11538-014-9968-0.pdf.

Seo:2014:CSA

[2938] Yeon-Jung Seo, Marit Nilsen-Hamilton, and Howard A. Levine. A computational study of alternate SELEX. Bulletin of Mathematical Biology, 76(7):1455–1521, July 2014. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-014-9954-6;   http://link.springer.com/content/
pdf/10.1007/s11538-014-9954-6.pdf.

Jin:2014:SID

[2939] Yu Jin, Frank M. Hilker, Peter M. Steffler, and Mark A. Lewis. Seasonal invasion dynamics in a spatially heterogeneous river with fluctuating flows. Bulletin of Mathematical Biology, 76(7):1522–1565, July 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-9957-3;   http://link.springer.com/content/pdf/10.1007/
s11538-014-9957-3.pdf.

Choi:2014:OTP

[2940] Sunhwa Choi and Eunok Jung. Optimal tuberculosis prevention and control strategy from a mathematical model based on real data. Bulletin of Mathematical Biology, 76(7):1566–1589, July 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-9962-6;   http://link.springer.com/content/pdf/10.1007/
s11538-014-9962-6.pdf.

Matzavinos:2014:SMC

[2941] Anastasios Matzavinos, Blerta Shtylla, Zachary Voller, Sijia Liu, and Mark A. J. Chaplain. Stochastic modelling of chromosomal segregation: Errors can introduce correction. Bulletin of Mathematical Biology, 76 (7):1590–1606, July 2014. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-014-9969-z;   http://link.springer.com/content/
pdf/10.1007/s11538-014-9969-z.pdf.

Agusto:2014:MDR

[2942] F. B. Agusto. Malaria drug resistance: The impact of human movement and spatial heterogeneity. Bulletin of Mathematical Biology, 76 (7):1607–1641, July 2014. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-014-9970-6;   http://link.springer.com/content/
pdf/10.1007/s11538-014-9970-6.pdf.

Pazdziorek:2014:MMS

[2943] Przemysław Rafał Paździorek. Mathematical model of stem cell differentiation and tissue regeneration with stochastic noise. Bulletin of Mathematical Biology, 76(7):1642–1669, July 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-014-9971-5;
 http://link.springer.com/content/pdf/10.1007/s11538-014-9971-
5.pdf.

Alsaleh:2014:ARS

[2944] Aliya A. Alsaleh and Abba B. Gumel. Analysis of risk-structured vaccination model for the dynamics of oncogenic and warts-causing HPV types. Bulletin of Mathematical Biology, 76(7):1670–1726, July 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-9972-4;   http://link.springer.com/content/pdf/10.1007/
s11538-014-9972-4.pdf.

Uecke:2014:IBM

[2945] Hannes Uecke, Johannes Müller, and Burkhard A. Hense. Individualbased model for quorum sensing with background flow. Bulletin of Mathematical Biology, 76(7):1727–1746, July 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-014-9974-2;
 http://link.springer.com/content/pdf/10.1007/s11538-014-9974-

2.pdf. See erratum [3032].

Ferreira:2014:LDI

[2946] Claudia P. Ferreira, Lourdes Esteva, Wesley A. C. Godoy, and Fernando L. Cônsoli. Landscape diversity influences dispersal and establishment of pest with complex nutritional ecology. Bulletin of Mathematical Biology, 76(7):1747–1761, July 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-014-9975-1;
 http://link.springer.com/content/pdf/10.1007/s11538-014-9975-
1.pdf.

Weekes:2014:MMM

[2947] Suzanne L. Weekes, Brian Barker, Sarah Bober, Karina Cisneros, Justina Cline, Amanda Thompson, Lynn Hlatky, Philip Hahnfeldt, and Heiko Enderling. A multicompartment mathematical model of cancer stem cell-driven tumor growth dynamics. Bulletin of Mathematical Biology, 76(7):1762–1782, July 2014. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-014-9976-0;   http://link.springer.com/content/
pdf/10.1007/s11538-014-9976-0.pdf.

Zumbrum:2014:MSR

[2948] Matthew E. Zumbrum and David A. Edwards. Multiple surface reactions in arrays with applications to optical biosensors. Bulletin of Mathematical Biology, 76(7):1783–1808, July 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-014-9977-z;
 http://link.springer.com/content/pdf/10.1007/s11538-014-9977-
z.pdf.

Eager:2014:MAD

[2949] Eric Alan Eager, Richard Rebarber, and Brigitte Tenhumberg. Modeling and analysis of a density-dependent stochastic integral projection model for a disturbance specialist plant and its seed bank. Bulletin of Mathematical Biology, 76(7):1809–1834, July 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-9978-y;   http://link.springer.com/content/pdf/10.1007/
s11538-014-9978-y.pdf.

Lo:2014:MAS

[2950] Wing-Cheong Lo, Hay-Oak Park, and Ching-Shan Chou. Mathematical analysis of spontaneous emergence of cell polarity. Bulletin of Mathematical Biology, 76(8):1835–1865, August 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-014-9982-2;
 http://link.springer.com/content/pdf/10.1007/s11538-014-9982-
2.pdf.

Anazawa:2014:IBC

[2951] Masahiro Anazawa. Individual-based competition between species with spatial correlation and aggregation. Bulletin of Mathematical Biology, 76(8):1866–1891, August 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-014-9980-4;   http://link.springer.com/
content/pdf/10.1007/s11538-014-9980-4.pdf.

Flockerzi:2014:SPS

[2952] Dietrich Flockerzi, Katharina Holstein, and Carsten Conradi. Nsite phosphorylation systems with 2N-1 steady states. Bulletin of Mathematical Biology, 76(8):1892–1916, August 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-014-9984-0;
 http://link.springer.com/content/pdf/10.1007/s11538-014-9984-
0.pdf.

White:2014:MPP

[2953] Diana White, Gerda de Vries, and Adriana Dawes. Microtubule patterning in the presence of stationary motor distributions. Bulletin of Mathematical Biology, 76(8):1917–1940, August 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-014-9991-1;
 http://link.springer.com/content/pdf/10.1007/s11538-014-9991-
1.pdf.

Lopez-Fenner:2014:BAM

[2954] Julio López-Fenner, Soledad Berríos, Catalina Manieu, Jesús Page, and Raúl Fernández-Donoso. Bivalent associations in Mus domesticus 2n = 40 spermatocytes. Are they random? Bulletin of Mathematical Biology, 76(8):1941–1952, August 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-014-9992-0;
 http://link.springer.com/content/pdf/10.1007/s11538-014-9992-
0.pdf.

Dunster:2014:RIM

[2955] J. L. Dunster, H. M. Byrne, and J. R. King. The resolution of inflammation: a mathematical model of neutrophil and macrophage interactions. Bulletin of Mathematical Biology, 76(8):1953–1980, August 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-9987-x;   http://link.springer.com/content/pdf/10.1007/
s11538-014-9987-x.pdf.

OFarrell:2014:MDB

[2956] Hayley O’Farrell and Stephen A. Gourley. Modelling the dynamics of bluetongue disease and the effect of seasonality. Bulletin of Mathematical Biology, 76(8):1981–2009, August 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-014-9989-8;
 http://link.springer.com/content/pdf/10.1007/s11538-014-9989-
8.pdf.

Sarapata:2014:CCI

[2957] E. A. Sarapata and L. G. de Pillis. A comparison and catalog of intrinsic tumor growth models. Bulletin of Mathematical Biology, 76(8): 2010–2024, August 2014. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-014-9986-y;   http://link.springer.com/content/
pdf/10.1007/s11538-014-9986-y.pdf.

Kong:2014:DCT

[2958] Jude D. Kong, William Davis, and Hao Wang. Dynamics of a cholera transmission model with immunological threshold and natural phage control in reservoir. Bulletin of Mathematical Biology, 76(8):2025–2051, August 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-

014-9996-9;   http://link.springer.com/content/pdf/10.1007/
s11538-014-9996-9.pdf.

Chamchod:2014:MAI

[2959] Farida Chamchod, Robert Stephen Cantrell, Chris Cosner, Ali N. Hassan, John C. Beier, and Shigui Ruan. A modeling approach to investigate epizootic outbreaks and enzootic maintenance of Rift valley fever virus. Bulletin of Mathematical Biology, 76(8):2052–2072, August 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-9998-7;   http://link.springer.com/content/pdf/10.1007/
s11538-014-9998-7.pdf.

Cruz-Pacheco:2014:VSS

[2960] Gustavo Cruz-Pacheco, Lourdes Esteva, and Cristobal Vargas. Vaccination strategies for SIR vector-transmitted diseases. Bulletin of Mathematical Biology, 76(8):2073–2090, August 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-014-9999-6;
 http://link.springer.com/content/pdf/10.1007/s11538-014-9999-
6.pdf.

Pearson:2014:MMH

[2961] T. Pearson, J. A. D. Wattis, J. R. King, I. A. MacDonald, and D. J. Mazzatti. A mathematical model of the human metabolic system and metabolic flexibility. Bulletin of Mathematical Biology, 76(9):2091– 2121, September 2014. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-014-0001-4;   http://link.springer.com/content/
pdf/10.1007/s11538-014-0001-4.pdf.

Chan:2014:MIM

[2962] Matthew H. Chan and Peter S. Kim. Modelling the impact of marine reserves on a population with depensatory dynamics. Bulletin of Mathematical Biology, 76(9):2122–2143, September 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-0003-2;   http://link.springer.com/content/pdf/10.1007/
s11538-014-0003-2.pdf.

Dolejs:2014:ATM

[2963] Josef Dolejs. Age trajectories of mortality from all diseases in the six most populated countries of the South America during the last decades. Bulletin of Mathematical Biology, 76(9):2144–2174, September 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-0005-0;   http://link.springer.com/content/pdf/10.1007/
s11538-014-0005-0.pdf.

Peace:2014:DPG

[2964] Angela Peace, Hao Wang, and Yang Kuang. Dynamics of a producer– grazer model incorporating the effects of excess food nutrient content on grazer’s growth. Bulletin of Mathematical Biology, 76(9):2175– 2197, September 2014. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-014-0006-z;   http://link.springer.com/content/
pdf/10.1007/s11538-014-0006-z.pdf.

Mandelshtam:2014:WLS

[2965] Yelena Mandelshtam and Natalia L. Komarova. When learners surpass their models: Mathematical modeling of learning from an inconsistent source. Bulletin of Mathematical Biology, 76(9):2198–2216, September 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-9990-2;   http://link.springer.com/content/pdf/10.1007/
s11538-014-9990-2.pdf.

Davis:2014:SSO

[2966] J. M. Davis and C. Pozrikidis. Self-sustained oscillations in blood flow through a honeycomb capillary network. Bulletin of Mathematical Biology, 76(9):2217–2237, September 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-014-0002-3;   http://link.springer.com/
content/pdf/10.1007/s11538-014-0002-3.pdf.

Evstigneev:2014:TDM

[2967] Vladyslav P. Evstigneev, Marina G. Holyavka, Sergii V. Khrapatiy, and Maxim P. Evstigneev. Theoretical description of metabolism using queueing theory. Bulletin of Mathematical Biology, 76(9):2238– 2248, September 2014. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-014-0004-1;   http://link.springer.com/content/
pdf/10.1007/s11538-014-0004-1.pdf.

Fischer:2014:MGS

[2968] Stephan Fischer, Samuel Bernard, Guillaume Beslon, and Carole Knibbe. A model for genome size evolution. Bulletin of Mathematical Biology, 76(9):2249–2291, September 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-014-9997-8;
 http://link.springer.com/content/pdf/10.1007/s11538-014-9997-
8.pdf.

Liddell:2014:MMS

[2969] Chelsea Liddell, Nicci Owusu-Brackett, and Dorothy Wallace. A mathematical model of sickle cell genome frequency in response to selective pressure from malaria. Bulletin of Mathematical Biology, 76(9):2292– 2305, September 2014. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-014-9993-z;   http://link.springer.com/content/
pdf/10.1007/s11538-014-9993-z.pdf.

Saut:2014:MGG

[2970] Olivier Saut, Jean-Baptiste Lagaert, Thierry Colin, and Hassan M. Fathallah-Shaykh. A multilayer grow-or-go model for GBM: Effects of invasive cells and anti-angiogenesis on growth. Bulletin of Mathematical Biology, 76(9):2306–2333, September 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-014-0007-y;
 http://link.springer.com/content/pdf/10.1007/s11538-014-0007-
y.pdf.

Lopez-Caamal:2014:EPD

[2971] Fernando López-Caamal and Tatiana T. Marquez-Lago. Exact probability distributions of selected species in stochastic chemical reaction networks. Bulletin of Mathematical Biology, 76(9):2334–2361, September 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-9985-z;   http://link.springer.com/content/pdf/10.1007/
s11538-014-9985-z.pdf.

Terry:2014:MST

[2972] Alan J. Terry. A minimal spatio-temporal model of the NF-κ B signalling pathway exhibits a range of behaviours. Bulletin of Mathematical Biology, 76(10):2363–2388, October 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-014-0011-2;
 http://link.springer.com/content/pdf/10.1007/s11538-014-0011-
2.pdf.

Gulbudak:2014:SAI

[2973] Hayriye Gulbudak and Maia Martcheva. A structured avian influenza model with imperfect vaccination and vaccine-induced asymptomatic infection. Bulletin of Mathematical Biology, 76(10):2389–2425, October 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-0012-1;   http://link.springer.com/content/pdf/10.1007/
s11538-014-0012-1.pdf.

Kravtsova:2014:ARS

[2974] Natalia Kravtsova and Adriana T. Dawes. Actomyosin regulation and symmetry breaking in a model of polarization in the early Caenorhabditis elegans embryo. Bulletin of Mathematical Biology, 76(10):2426–2448, October 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-0016-x;   http://link.springer.com/content/pdf/10.1007/
s11538-014-0016-x.pdf.

Yan:2014:MRT

[2975] Zhifeng Yan and Markus Hilpert. A multiple-relaxation-time lattice– Boltzmann model for bacterial chemotaxis: Effects of initial concentration, diffusion, and hydrodynamic dispersion on traveling bacterial bands. Bulletin of Mathematical Biology, 76(10):2449–2475, October 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-0020-1;   http://link.springer.com/content/pdf/10.1007/
s11538-014-0020-1.pdf.

Ngwa:2014:RSS

[2976] Gideon A. Ngwa, Terence T. Wankah, Mary Y. Fomboh-Nforba, Calsitus N. Ngonghala, and Miranda I. Teboh-Ewungkem. On a reproductive stage-structured model for the population dynamics of the malaria vector. Bulletin of Mathematical Biology, 76(10):2476–2516, October 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-0021-0;   http://link.springer.com/content/pdf/10.1007/
s11538-014-0021-0.pdf.

Keijsper:2014:RPL

[2977] J. C. M. Keijsper and R. A. Pendavingh. Reconstructing a phylogenetic level-1 network from quartets. Bulletin of Mathematical Biology, 76(10): 2517–2541, October 2014. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-014-0022-z;   http://link.springer.com/content/
pdf/10.1007/s11538-014-0022-z.pdf.

Blanchini:2014:SCC

[2978] Franco Blanchini, Elisa Franco, and Giulia Giordano. A structural classification of candidate oscillatory and multistationary biochemical systems. Bulletin of Mathematical Biology, 76(10):2542–2569, October 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-0023-y;   http://link.springer.com/content/pdf/10.1007/
s11538-014-0023-y.pdf.

Deshpande:2014:ARN

[2979] Abhishek Deshpande and Manoj Gopalkrishnan. Autocatalysis in reaction networks. Bulletin of Mathematical Biology, 76(10):2570–2595, October 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-0024-x;   http://link.springer.com/content/pdf/10.1007/
s11538-014-0024-x.pdf.

Heinrich:2014:ITD

[2980] Philippe Heinrich, Mariano Gonzalez Pisfil, Jonas Kahn, Laurent Héliot, and Aymeric Leray. Implementation of transportation distance for analyzing FLIM and FRET experiments. Bulletin of Mathematical Biology, 76(10):2596–2626, October 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-014-0025-9;
 http://link.springer.com/content/pdf/10.1007/s11538-014-0025-
9.pdf.

Rodrigues:2014:CEA

[2981] Paula Rodrigues, Cristiana J. Silva, and Delfim F. M. Torres. Cost-effectiveness analysis of optimal control measures for tuberculosis. Bulletin of Mathematical Biology, 76(10):2627–2645, October 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-0028-6;   http://link.springer.com/content/pdf/10.1007/
s11538-014-0028-6.pdf.

Martinez-Rodriguez:2014:NIW

[2982] P. Martínez-Rodríguez, R. Granero-Belinchón, F. Arroyo-Yebras, and J. L. Bella. New insight into Wolbachia epidemiology: Its varying incidence during the host life cycle can alter bacteria spread. Bulletin of Mathematical Biology, 76(10):2646–2663, October 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-0029-5;   http://link.springer.com/content/pdf/10.1007/
s11538-014-0029-5.pdf.

Cordue:2014:PND

[2983] Paul Cordue, Simone Linz, and Charles Semple. Phylogenetic networks that display a tree twice. Bulletin of Mathematical Biology, 76(10): 2664–2679, October 2014. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-014-0032-x;   http://link.springer.com/content/
pdf/10.1007/s11538-014-0032-x.pdf.

Stover:2014:CDH

[2984] Joseph P. Stover, Bruce E. Kendall, and Roger M. Nisbet. Consequences of dispersal heterogeneity for population spread and persistence. Bulletin of Mathematical Biology, 76(11):2681–2710, November 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-0014-z;   http://link.springer.com/content/pdf/10.1007/
s11538-014-0014-z.pdf.

Spina:2014:SMC

[2985] Serena Spina, Virginia Giorno, Patricia Román-Román, and Francisco Torres-Ruiz. A stochastic model of cancer growth subject to an intermittent treatment with combined effects: Reduction in tumor size and rise in growth rate. Bulletin of Mathematical Biology, 76(11):2711– 2736, November 2014. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-014-0026-8;   http://link.springer.com/content/
pdf/10.1007/s11538-014-0026-8.pdf.

Sanchez-Tapia:2014:FTC

[2986] Cynthia Sanchez-Tapia and Frederic Y. M. Wan. Fastest time to cancer by loss of tumor suppressor genes. Bulletin of Mathematical Biology, 76(11):2737–2784, November 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-014-0027-7;   http://link.springer.com/
content/pdf/10.1007/s11538-014-0027-7.pdf.

Tan:2014:SSE

[2987] Shaolin Tan, Shasha Feng, Pei Wang, and Yao Chen. Strategy selection in evolutionary game dynamics on group interaction networks. Bulletin of Mathematical Biology, 76(11):2785–2805, November 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-0031-y;   http://link.springer.com/content/pdf/10.1007/
s11538-014-0031-y.pdf.

Lai:2014:RES

[2988] Xiulan Lai and Xingfu Zou. Repulsion effect on superinfecting virions by infected cells. Bulletin of Mathematical Biology, 76(11):2806– 2833, November 2014. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-014-0033-9;   http://link.springer.com/content/
pdf/10.1007/s11538-014-0033-9.pdf.

Payvandi:2014:MMP

[2989] S. Payvandi, K. R. Daly, K. C. Zygalakis, and T. Roose. Mathematical modelling of the phloem: The importance of diffusion on sugar transport at osmotic equilibrium. Bulletin of Mathematical Biology, 76(11): 2834–2865, November 2014. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-014-0035-7;   http://link.springer.com/content/
pdf/10.1007/s11538-014-0035-7.pdf.

Marasco:2014:VPF

[2990] Addolorata Marasco, Annalisa Iuorio, Fabrizio Cartení, Giuliano Bonanomi, Daniel M. Tartakovsky, Stefano Mazzoleni, and Francesco Giannino. Vegetation pattern formation due to interactions between water availability and toxicity in plant–soil feedback. Bulletin of Mathematical Biology, 76(11):2866–2883, November 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-0036-6;   http://link.springer.com/content/pdf/10.1007/
s11538-014-0036-6.pdf.

Lopez:2014:VMM

[2991] Álvaro G. López, Jesús M. Seoane, and Miguel A. F. Sanjuán. A validated mathematical model of tumor growth including tumor– host interaction, cell-mediated immune response and chemotherapy. Bulletin of Mathematical Biology, 76(11):2884–2906, November 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-0037-5;   http://link.springer.com/content/pdf/10.1007/
s11538-014-0037-5.pdf.

Foteinopoulos:2014:EAM

[2992] Panayiotis Foteinopoulos and Bela M. Mulder. The effect of anisotropic microtubule-bound nucleations on ordering in the plant cortical array. Bulletin of Mathematical Biology, 76(11):2907–2922, November 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-0039-3;   http://link.springer.com/content/pdf/10.1007/
s11538-014-0039-3.pdf.

Dimitrova:2014:DII

[2993] Elena Dimitrova and Brandilyn Stigler. Data identification for improving gene network inference using computational algebra. Bulletin of Mathematical Biology, 76(11):2923–2940, November 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-9979-x;   http://link.springer.com/content/pdf/10.1007/
s11538-014-9979-x.pdf.

Wells:2014:EII

[2994] Chad R. Wells, Jean M. Tchuenche, Lauren Ancel Meyers, Alison P. Galvani, and Chris T. Bauch. Erratum to: Impact of Imitation Processes on the Effectiveness of Ring Vaccination. Bulletin of Mathematical Biology, 76(11):2941–2944, November 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-0008-x;   http://link.springer.com/content/pdf/10.1007/
s11538-014-0008-x.pdf. See [2624].

Veliz-Cuba:2014:PLB

[2995] Alan Veliz-Cuba, Ajit Kumar, and Krešimir Josić. Piecewise linear and Boolean models of chemical reaction networks. Bulletin of Mathematical Biology, 76(12):2945–2984, December 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-0040-x;   http://link.springer.com/content/pdf/10.1007/
s11538-014-0040-x.pdf.

Davis:2014:LSB

[2996] Jeffrey M. Davis. On the linear stability of blood flow through model capillary networks. Bulletin of Mathematical Biology, 76(12):2985– 3015, December 2014. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-014-0041-9;   http://link.springer.com/content/
pdf/10.1007/s11538-014-0041-9.pdf.

Kenkre:2014:TTI

[2997] V. M. Kenkre and S. Sugaya. Theory of the transmission of infection in the spread of epidemics: Interacting random walkers with and without confinement. Bulletin of Mathematical Biology, 76(12):3016– 3027, December 2014. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-014-0042-8;   http://link.springer.com/content/
pdf/10.1007/s11538-014-0042-8.pdf.

Allena:2014:RDF

[2998] R. Allena and P. K. Maini. Reaction–diffusion finite element model of lateral line primordium migration to explore cell leadership. Bulletin of Mathematical Biology, 76(12):3028–3050, December 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-0043-7;   http://link.springer.com/content/pdf/10.1007/
s11538-014-0043-7.pdf.

Cao:2014:STP

[2999] Yang Cao and Radek Erban. Stochastic Turing patterns: Analysis of compartment-based approaches. Bulletin of Mathematical Biology, 76(12):3051–3069, December 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-014-0044-6;   http://link.springer.com/
content/pdf/10.1007/s11538-014-0044-6.pdf.

Seppanen:2014:EDD

[3000] Anne Seppänen and Kalle Parvinen. Evolution of density-dependent cooperation. Bulletin of Mathematical Biology, 76(12):3070–3087, December 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-9994-y;   http://link.springer.com/content/pdf/10.1007/
s11538-014-9994-y.pdf.

Chen:2014:MMG

[3001] C. Y. Chen and J. P. Ward. A mathematical model of the growth of uterine myomas. Bulletin of Mathematical Biology, 76(12):3088– 3121, December 2014. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-014-0045-5;   http://link.springer.com/content/
pdf/10.1007/s11538-014-0045-5.pdf.

Nieddu:2014:ACP

[3002] Garrett Nieddu, Lora Billings, and Eric Forgoston. Analysis and control of pre-extinction dynamics in stochastic populations. Bulletin of Mathematical Biology, 76(12):3122–3137, December 2014. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-0047-3;   http://link.springer.com/content/pdf/10.1007/
s11538-014-0047-3.pdf.

Lorz:2015:MES

[3003] Alexander Lorz, Tommaso Lorenzi, Jean Clairambault, Alexandre Escargueil, and Benoît Perthame. Modeling the effects of space structure and combination therapies on phenotypic heterogeneity and drug resistance in solid tumors. Bulletin of Mathematical Biology, 77(1):1–22, January 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-0046-4;   http://link.springer.com/content/pdf/10.1007/
s11538-014-0046-4.pdf.

Alexander:2015:ACI

[3004] Samuel A. Alexander, Arie de Bruin, and D. J. Kornet. An alternative construction of internodons: The emergence of a multi-level tree of life. Bulletin of Mathematical Biology, 77(1):23–45, January 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-0048-2;   http://link.springer.com/content/pdf/10.1007/
s11538-014-0048-2.pdf.

Baskowski:2015:NTC

[3005] Sarah Baskowski, Vincent Moulton, Andreas Spillner, and Taoyang Wu. Neighborhoods of trees in circular orderings. Bulletin of Mathematical Biology, 77(1):46–70, January 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-014-0049-1;   http://link.springer.com/
content/pdf/10.1007/s11538-014-0049-1.pdf. See erratum [3031].

Frieden:2015:CSC

[3006] B. Roy Frieden and Robert A. Gatenby. Cancer suppression by compression. Bulletin of Mathematical Biology, 77(1):71–82, January 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-0051-7;   http://link.springer.com/content/pdf/10.1007/
s11538-014-0051-7.pdf.

Kurella:2015:AAF

[3007] Venu Kurella, Justin C. Tzou, Daniel Coombs, and Michael J. Ward. Asymptotic analysis of first passage time problems inspired by ecology. Bulletin of Mathematical Biology, 77(1):83–125, January 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-0053-5;   http://link.springer.com/content/pdf/10.1007/
s11538-014-0053-5.pdf.

Conradi:2015:GCR

[3008] Carsten Conradi and Anne Shiu. A global convergence result for processive multisite phosphorylation systems. Bulletin of Mathematical Biology, 77(1):126–155, January 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-014-0054-4;   http://link.springer.com/
content/pdf/10.1007/s11538-014-0054-4.pdf.

Echeverri:2015:ESB

[3009] L. F. Echeverri, M. A. Herrero, J. M. Lopez, and G. Oleaga. Early stages of bone fracture healing: Formation of a fibrin–collagen scaffold in the fracture hematoma. Bulletin of Mathematical Biology, 77 (1):156–183, January 2015. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-014-0055-3;   http://link.springer.com/content/
pdf/10.1007/s11538-014-0055-3.pdf.

AlGaralleh:2015:MIB

[3010] Hakim Al Garalleh, Ngamta Thamwattana, Barry J. Cox, and James M. Hill. Modeling interactions between c _{60} antiviral compounds and HIV protease. Bulletin of Mathematical Biology, 77(1):184–201, January 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-0056-2;   http://link.springer.com/content/pdf/10.1007/
s11538-014-0056-2.pdf.

Wang:2015:PPM

[3011] Yuanshi Wang and Shikun Wang. Persistence of pollination mutualisms in the presence of ants. Bulletin of Mathematical Biology, 77 (1):202–229, January 2015. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-014-0057-1;   http://link.springer.com/content/
pdf/10.1007/s11538-014-0057-1.pdf.

Postnov:2015:CVD

[3012] D. E. Postnov, A. Y. Neganova, O. V. Sosnovtseva, N.-H. Holstein-Rathlou, and J. C. Brings Jacobsen. Conducted vasoreactivity: the dynamical point of view. Bulletin of Mathematical Biology, 77(1): 230–249, January 2015. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-014-0058-0;   http://link.springer.com/content/
pdf/10.1007/s11538-014-0058-0.pdf.

Levin:2015:WMC

[3013] Simon Levin. What mathematics can do for sustainability. Bulletin of Mathematical Biology, 77(2):251–253, February 2015. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-014-0038-4;
 http://link.springer.com/content/pdf/10.1007/s11538-014-0038-
4.pdf.

Muneepeerakul:2015:TQS

[3014] Rachata Muneepeerakul and Carlos Castillo-Chavez. Towards a quantitative science of sustainability. Bulletin of Mathematical Biology, 77(2): 254–258, February 2015. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-014-0034-8;   http://link.springer.com/content/
pdf/10.1007/s11538-014-0034-8.pdf.

Anderies:2015:UDS

[3015] John M. Anderies. Understanding the dynamics of sustainable social– ecological systems: Human behavior, institutions, and regulatory feedback networks. Bulletin of Mathematical Biology, 77(2):259–280, February 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-0030-z;   http://link.springer.com/content/pdf/10.1007/
s11538-014-0030-z.pdf.

DeLara:2015:SVO

[3016] Michel De Lara, Vincent Martinet, and Luc Doyen. Satisficing versus optimality: Criteria for sustainability. Bulletin of Mathematical Biology, 77 (2):281–297, February 2015. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-014-9944-8;   http://link.springer.com/content/
pdf/10.1007/s11538-014-9944-8.pdf.

Vico:2015:EAQ

[3017] Giulia Vico and Amilcare Porporato. Ecohydrology of agroecosystems: Quantitative approaches towards sustainable irrigation. Bulletin of Mathematical Biology, 77(2):298–318, February 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-9988-9;   http://link.springer.com/content/pdf/10.1007/
s11538-014-9988-9.pdf.

Kareva:2015:RCS

[3018] Irina Kareva, Benjamin Morin, and Carlos Castillo-Chavez. Resource consumption, sustainability, and cancer. Bulletin of Mathematical Biology, 77(2):319–338, February 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-014-9983-1;   http://link.springer.com/
content/pdf/10.1007/s11538-014-9983-1.pdf.

Ridolfi:2015:ICS

[3019] Luca Ridolfi, Paolo D’Odorico, and Francesco Laio. Indicators of collapse in systems undergoing unsustainable growth. Bulletin of Mathematical Biology, 77(2):339–347, February 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-013-9922-6;
 http://link.springer.com/content/pdf/10.1007/s11538-013-9922-
6.pdf.

Fenichel:2015:SS

[3020] Eli P. Fenichel and Jinhua Zhao. Sustainability and substitutability. Bulletin of Mathematical Biology, 77(2):348–367, February 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-9963-5;   http://link.springer.com/content/pdf/10.1007/
s11538-014-9963-5.pdf.

Angulo:2015:DFS

[3021] David Angulo, Fabiola Angulo, and Gerard Olivar. Dynamics and forecast in a simple model of sustainable development for rural populations. Bulletin of Mathematical Biology, 77(2):368–389, February 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-9934-x;   http://link.springer.com/content/pdf/10.1007/
s11538-014-9934-x.pdf.

Qubbaj:2015:LNS

[3022] Murad R. Qubbaj, Shade T. Shutters, and Rachata Muneepeerakul. Living in a network of scaling cities and finite resources. Bulletin of Mathematical Biology, 77(2):390–407, February 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-9949-3;   http://link.springer.com/content/pdf/10.1007/
s11538-014-9949-3.pdf.

Ribeiro:2015:NPM

[3023] Fabiano L. Ribeiro. A non-phenomenological model of competition and cooperation to explain population growth behaviors. Bulletin of Mathematical Biology, 77(3):409–433, March 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-014-0059-z;
 http://link.springer.com/content/pdf/10.1007/s11538-014-0059-
z.pdf.

Oremland:2015:OHP

[3024] Matthew Oremland and Reinhard Laubenbacher. Optimal harvesting for a predator–prey agent-based model using difference equations. Bulletin of Mathematical Biology, 77(3):434–459, March 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-0060-6;   http://link.springer.com/content/pdf/10.1007/
s11538-014-0060-6.pdf.

Brauer:2015:SSN

[3025] Fred Brauer. Some simple nosocomial disease transmission models. Bulletin of Mathematical Biology, 77(3):460–469, March 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
015-0061-0;   http://link.springer.com/content/pdf/10.1007/
s11538-015-0061-0.pdf.

Picardo:2015:USA

[3026] Jason R. Picardo and S. Pushpavanam. Understanding the shape of ant craters: a continuum model. Bulletin of Mathematical Biology, 77 (3):470–487, March 2015. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-015-0063-y;   http://link.springer.com/content/
pdf/10.1007/s11538-015-0063-y.pdf.

Wilcox:2015:MEF

[3027] Marianne Wilcox, Allan R. Willms, and William F. Langford. A model of the effects of fluid variation due to body position on Cheyne–Stokes respiration. Bulletin of Mathematical Biology, 77(3):488–498, March 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
015-0064-x;   http://link.springer.com/content/pdf/10.1007/
s11538-015-0064-x.pdf.

DeLegge:2015:EMM

[3028] Anthony DeLegge, Katie Hunzinger, Reema Khatri, and Kiran Munir. An epidemic model with a multistage vaccine. Bulletin of Mathematical Biology, 77(3):499–513, March 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-015-0069-5;
 http://link.springer.com/content/pdf/10.1007/s11538-015-0069-
5.pdf.

Saad-Roy:2015:MBB

[3029] C. M. Saad-Roy, Zhisheng Shuai, and P. van den Driessche. Models of Bovine Babesiosis including juvenile cattle. Bulletin of Mathematical Biology, 77(3):514–547, March 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-015-0068-6;   http://link.springer.com/
content/pdf/10.1007/s11538-015-0068-6.pdf.

Garvie:2015:SFE

[3030] Marcus R. Garvie, John Burkardt, and Jeff Morgan. Simple finite element methods for approximating predator–prey dynamics in two dimensions using Matlab. Bulletin of Mathematical Biology, 77(3):548–578, March 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
015-0062-z;   http://link.springer.com/content/pdf/10.1007/
s11538-015-0062-z.pdf.

Bastkowski:2015:ENT

[3031] Sarah Bastkowski, Vincent Moulton, Andreas Spillner, and Taoyang Wu. Erratum to: Neighborhoods of Trees in Circular Orderings. Bulletin of Mathematical Biology, 77(3):579, March 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
015-0071-y;   http://link.springer.com/content/pdf/10.1007/
s11538-015-0071-y.pdf. See [3005].

Uecker:2015:EIB

[3032] Hannes Uecker, Johannes Müller, and Burkhard A. Hense. Erratum to: Individual-Based Model for Quorum Sensing with Background Flow. Bulletin of Mathematical Biology, 77(3):580, March 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
015-0072-x;   http://link.springer.com/content/pdf/10.1007/
s11538-015-0072-x.pdf. See [2945].

Simpson:2015:SIS

[3033] Matthew J. Simpson and Ruth E. Baker. Special issue on spatial moment techniques for modelling biological processes. Bulletin of Mathematical Biology, 77(4):581–585, April 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-015-0066-8;
 http://link.springer.com/content/pdf/10.1007/s11538-015-0066-
8.pdf.

Plank:2015:SPP

[3034] Michael J. Plank and Richard Law. Spatial point processes and moment dynamics in the life sciences: a parsimonious derivation and some extensions. Bulletin of Mathematical Biology, 77(4):586–613, April 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-0018-8;   http://link.springer.com/content/pdf/10.1007/
s11538-014-0018-8.pdf.

Sharkey:2015:EES

[3035] K. J. Sharkey, I. Z. Kiss, R. R. Wilkinson, and P. L. Simon. Exact equations for SIR epidemics on tree graphs. Bulletin of Mathematical Biology, 77(4):614–645, April 2015. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-013-9923-5;   http://link.springer.com/content/
pdf/10.1007/s11538-013-9923-5.pdf.

House:2015:AMC

[3036] Thomas House. Algebraic moment closure for population dynamics on discrete structures. Bulletin of Mathematical Biology, 77(4):646–659, April 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-9981-3;   http://link.springer.com/content/pdf/10.1007/
s11538-014-9981-3.pdf.

Mente:2015:AIC

[3037] Carsten Mente, Anja Voss-Böhme, and Andreas Deutsch. Analysis of individual cell trajectories in lattice–gas cellular automaton models for migrating cell populations. Bulletin of Mathematical Biology, 77(4): 660–697, April 2015. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-015-0079-3;   http://link.springer.com/content/
pdf/10.1007/s11538-015-0079-3.pdf.

Hiebeler:2015:VCS

[3038] David E. Hiebeler, Rachel M. Rier, Josh Audibert, Phillip J. LeClair, and Anna Webber. Variability in a community-structured SIS epidemiological model. Bulletin of Mathematical Biology, 77(4):698–712, April 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-0017-9;   http://link.springer.com/content/pdf/10.1007/
s11538-014-0017-9.pdf.

Markham:2015:CAM

[3039] Deborah C. Markham, Matthew J. Simpson, and Ruth E. Baker. Choosing an appropriate modelling framework for analysing multispecies coculture cell biology experiments. Bulletin of Mathematical Biology, 77 (4):713–734, April 2015. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-014-0050-8;   http://link.springer.com/content/
pdf/10.1007/s11538-014-0050-8.pdf.

Mackey:2015:WMD

[3040] Michael C. Mackey and Philip K. Maini. What has mathematics done for biology? Bulletin of Mathematical Biology, 77(5):735–738, May 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
015-0065-9;   http://link.springer.com/content/pdf/10.1007/
s11538-015-0065-9.pdf.

Dale:2015:UTA

[3041] David C. Dale and Michael C. Mackey. Understanding, treating and avoiding hematological disease: Better medicine through mathematics? Bulletin of Mathematical Biology, 77(5):739–757, May 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-9995-x;   http://link.springer.com/content/pdf/10.1007/
s11538-014-9995-x.pdf.

Friedman:2015:MMA

[3042] Avner Friedman and Wenrui Hao. A mathematical model of atherosclerosis with reverse cholesterol transport and associated risk factors. Bulletin of Mathematical Biology, 77(5):758–781, May 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-0010-3;   http://link.springer.com/content/pdf/10.1007/
s11538-014-0010-3.pdf.

Glass:2015:FCO

[3043] Leon Glass and Alvin Shrier. Functional characterization of oscillatory and excitable media. Bulletin of Mathematical Biology, 77(5):782–795, May 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-0015-y;   http://link.springer.com/content/pdf/10.1007/
s11538-014-0015-y.pdf.

Tyson:2015:BOT

[3044] John J. Tyson and Bela Novak. Bistability, oscillations, and traveling waves in frog egg extracts. Bulletin of Mathematical Biology, 77 (5):796–816, May 2015. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-014-0009-9;   http://link.springer.com/content/
pdf/10.1007/s11538-014-0009-9.pdf.

Umulis:2015:RMM

[3045] David M. Umulis and Hans G. Othmer. The role of mathematical models in understanding pattern formation in developmental biology. Bulletin of Mathematical Biology, 77(5):817–845, May 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-0019-7;   http://link.springer.com/content/pdf/10.1007/
s11538-014-0019-7.pdf.

Jackson:2015:PSM

[3046] Pamela R. Jackson, Joseph Juliano, Andrea Hawkins-Daarud, Russell C. Rockne, and Kristin R. Swanson. Patient-specific mathematical neurooncology: Using a simple proliferation and invasion tumor model to inform clinical practice. Bulletin of Mathematical Biology, 77(5):846–856, May 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
015-0067-7;   http://link.springer.com/content/pdf/10.1007/
s11538-015-0067-7.pdf.

Mangel:2015:SDP

[3047] Marc Mangel. Stochastic dynamic programming illuminates the link between environment, physiology, and evolution. Bulletin of Mathematical Biology, 77(5):857–877, May 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-014-9973-3;   http://link.springer.com/
content/pdf/10.1007/s11538-014-9973-3.pdf.

Wallace:2015:MFE

[3048] Rodrick Wallace. Metabolic free energy and biological codes: a ‘data rate theorem’ aging model. Bulletin of Mathematical Biology, 77(6):879–903, June 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-0013-0;   http://link.springer.com/content/pdf/10.1007/
s11538-014-0013-0.pdf.

Chen:2015:CMD

[3049] Xingjiang Chen and James Sneyd. A computational model of the dendron of the GnRH neuron. Bulletin of Mathematical Biology, 77(6):904–926, June 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
014-0052-6;   http://link.springer.com/content/pdf/10.1007/
s11538-014-0052-6.pdf.

Kura:2015:MDH

[3050] Klodeta Kura, Mark Broom, and Anne Kandler. Modelling dominance hierarchies under winner and loser effects. Bulletin of Mathematical Biology, 77(6):927–952, June 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-015-0070-z;   http://link.springer.com/
content/pdf/10.1007/s11538-015-0070-z.pdf.

Edwards:2015:MFG

[3051] R. Edwards, A. Machina, G. McGregor, and P. van den Driessche. A modelling framework for gene regulatory networks including transcription and translation. Bulletin of Mathematical Biology, 77(6):953–983, June 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
015-0073-9;   http://link.springer.com/content/pdf/10.1007/
s11538-015-0073-9.pdf.

Jacobsen:2015:ECM

[3052] Karly Jacobsen, Luke Russell, Balveen Kaur, and Avner Friedman. Effects of CCN1 and macrophage content on glioma virotherapy: a mathematical model. Bulletin of Mathematical Biology, 77(6):984–1012, June 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
015-0074-8;   http://link.springer.com/content/pdf/10.1007/
s11538-015-0074-8.pdf.

Mirzaev:2015:LDS

[3053] Inom Mirzaev and David M. Bortz. Laplacian dynamics with synthesis and degradation. Bulletin of Mathematical Biology, 77(6):1013–1045, June 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
015-0075-7;   http://link.springer.com/content/pdf/10.1007/
s11538-015-0075-7.pdf.

Hong:2015:MFU

[3054] Tian Hong, Cihan Oguz, and John J. Tyson. A mathematical framework for understanding four-dimensional heterogeneous differentiation of CD4^{+} T cells. Bulletin of Mathematical Biology, 77(6):1046–1064, June 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
015-0076-6;   http://link.springer.com/content/pdf/10.1007/
s11538-015-0076-6.pdf.

Johnston:2015:CAS

[3055] Matthew D. Johnston. A computational approach to steady state correspondence of regular and generalized mass action systems. Bulletin of Mathematical Biology, 77(6):1065–1100, June 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-015-0077-5;
 http://link.springer.com/content/pdf/10.1007/s11538-015-0077-
5.pdf.

Heppell:2015:MID

[3056] Charles Heppell, Tiina Roose, and Giles Richardson. A model for interstitial drainage through a sliding lymphatic valve. Bulletin of Mathematical Biology, 77(6):1101–1131, June 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-015-0078-4;
 http://link.springer.com/content/pdf/10.1007/s11538-015-0078-
4.pdf.

Painter:2015:NMC

[3057] K. J. Painter, J. M. Bloomfield, J. A. Sherratt, and A. Gerisch. A nonlocal model for contact attraction and repulsion in heterogeneous cell populations. Bulletin of Mathematical Biology, 77(6):1132–1165, June 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
015-0080-x;   http://link.springer.com/content/pdf/10.1007/
s11538-015-0080-x.pdf.

Ghorai:2015:WSG

[3058] S. Ghorai, R. Singh, and N. A. Hill. Wavelength selection in gyrotactic bioconvection. Bulletin of Mathematical Biology, 77(6):1166–1184, June 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
015-0081-9;   http://link.springer.com/content/pdf/10.1007/
s11538-015-0081-9.pdf.

Mavelli:2015:SPS

[3059] Fabio Mavelli, Roberto Marangoni, and Pasquale Stano. A simple protein synthesis model for the PURE system operation. Bulletin of Mathematical Biology, 77(6):1185–1212, June 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-015-0082-8;
 http://link.springer.com/content/pdf/10.1007/s11538-015-0082-
8.pdf.

Taylor-King:2015:BBG

[3060] Jake P. Taylor-King, E. Emiel van Loon, Gabriel Rosser, and S. Jon Chapman. From birds to bacteria: Generalised velocity jump processes with resting states. Bulletin of Mathematical Biology, 77(7):1213–1236, July 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
015-0083-7;   http://link.springer.com/content/pdf/10.1007/
s11538-015-0083-7.pdf.

Wang:2015:EPM

[3061] Xinxin Wang, Shengqiang Liu, Lin Wang, and Weiwei Zhang. An epidemic patchy model with entry–exit screening. Bulletin of Mathematical Biology, 77(7):1237–1255, July 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-015-0084-6;
 http://link.springer.com/content/pdf/10.1007/s11538-015-0084-
6.pdf.

Duncan:2015:MMP

[3062] Jacob P. Duncan, James A. Powell, Luis F. Gordillo, and Joseph Eason. A model for mountain pine beetle outbreaks in an agestructured forest: Predicting severity and outbreak-recovery cycle period. Bulletin of Mathematical Biology, 77(7):1256–1284, July 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
015-0085-5;   http://link.springer.com/content/pdf/10.1007/
s11538-015-0085-5.pdf.

Zhao:2015:SSD

[3063] Yu Zhao, Sanling Yuan, and Junling Ma. Survival and stationary distribution analysis of a stochastic competitive model of three species in a polluted environment. Bulletin of Mathematical Biology, 77(7):1285–1326, July 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
015-0086-4;   http://link.springer.com/content/pdf/10.1007/
s11538-015-0086-4.pdf.

Young:2015:BVA

[3064] Glenn Young, Bard Ermentrout, and Jonathan E. Rubin. A boundary value approach to optimization with an application to Salmonella competition. Bulletin of Mathematical Biology, 77(7):1327–1348, July 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
015-0087-3;   http://link.springer.com/content/pdf/10.1007/
s11538-015-0087-3.pdf.

Lee:2015:MMC

[3065] Wanho Lee and Eunok Jung. A multiscale model of cardiovascular system including an immersed whole heart in the cases of normal and ventricular septal defect (VSD). Bulletin of Mathematical Biology, 77(7):1349–1376, July 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
015-0088-2;   http://link.springer.com/content/pdf/10.1007/
s11538-015-0088-2.pdf.

Karst:2015:OME

[3066] Nathaniel J. Karst, Brian D. Storey, and John B. Geddes. Oscillations and multiple equilibria in microvascular blood flow. Bulletin of Mathematical Biology, 77(7):1377–1400, July 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-015-0089-1;
 http://link.springer.com/content/pdf/10.1007/s11538-015-0089-
1.pdf.

Kollar:2015:EJQ

[3067] Richard Kollár and Katarína Šišková. Extension and justification of quasi–steady-state approximation for reversible bimolecular binding. Bulletin of Mathematical Biology, 77(7):1401–1436, July 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
015-0090-8;   http://link.springer.com/content/pdf/10.1007/
s11538-015-0090-8.pdf.

Forgoston:2015:FIU

[3068] Eric Forgoston, Leah B. Shaw, and Ira B. Schwartz. A framework for inferring unobserved multistrain epidemic subpopulations using synchronization dynamics. Bulletin of Mathematical Biology, 77(7):1437–1455, July 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
015-0091-7;   http://link.springer.com/content/pdf/10.1007/
s11538-015-0091-7.pdf.

Dahlquist:2015:PEG

[3069] Kam D. Dahlquist, Ben G. Fitzpatrick, Erika T. Camacho, Stephanie D. Entzminger, and Nathan C. Wanner. Parameter estimation for gene regulatory networks from microarray data: Cold shock response in Saccharomyces cerevisiae. Bulletin of Mathematical Biology, 77(8):1457– 1492, August 2015. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-015-0092-6;   http://link.springer.com/content/
pdf/10.1007/s11538-015-0092-6.pdf.

Ratti:2015:MMH

[3070] Vardayani Ratti, Peter G. Kevan, and Hermann J. Eberl. A mathematical model of the honeybee–Varroa destructor–acute bee paralysis virus system with seasonal effects. Bulletin of Mathematical Biology, 77(8): 1493–1520, August 2015. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-015-0093-5;   http://link.springer.com/content/
pdf/10.1007/s11538-015-0093-5.pdf.

Rabinovitch:2015:BMR

[3071] A. Rabinovitch, M. Friedman, D. Braunstein, Y. Biton, and I. Aviram. The baroreflex mechanism revisited. Bulletin of Mathematical Biology, 77(8):1521–1538, August 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-015-0094-4;   http://link.springer.com/
content/pdf/10.1007/s11538-015-0094-4.pdf.

Fowler:2015:SAE

[3072] A. C. Fowler and T. Déirdre Hollingsworth. Simple approximations for epidemics with exponential and fixed infectious periods. Bulletin of Mathematical Biology, 77(8):1539–1555, August 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
015-0095-3;   http://link.springer.com/content/pdf/10.1007/
s11538-015-0095-3.pdf.

Warsinske:2015:IMH

[3073] Hayley C. Warsinske, Shanna L. Ashley, Jennifer J. Linderman, Bethany B. Moore, and Denise E. Kirschner. Identifying mechanisms of homeostatic signaling in fibroblast differentiation. Bulletin of Mathematical Biology, 77(8):1556–1582, August 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-015-0096-2;
 http://link.springer.com/content/pdf/10.1007/s11538-015-0096-
2.pdf.

Rodrigues:2015:PIS

[3074] Luiz Alberto Díaz Rodrigues, Diomar Cristina Mistro, Elisa Regina Cara, Natalia Petrovskaya, and Sergei Petrovskii. Patchy invasion of stage-structured alien species with short-distance and long-distance dispersal. Bulletin of Mathematical Biology, 77(8):1583–1619, August 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
015-0097-1;   http://link.springer.com/content/pdf/10.1007/
s11538-015-0097-1.pdf.

Meshkat:2015:IRS

[3075] Nicolette Meshkat, Seth Sullivant, and Marisa Eisenberg. Identifiability results for several classes of linear compartment models. Bulletin of Mathematical Biology, 77(8):1620–1651, August 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
015-0098-0;   http://link.springer.com/content/pdf/10.1007/
s11538-015-0098-0.pdf.

Paton:2015:MCG

[3076] Kelly M. Paton, Lisa Anderson, Pauline Flottat, and Eric N. Cytrynbaum. A model of chloroplast growth regulation in mesophyll cells. Bulletin of Mathematical Biology, 77(9):1653–1667, September 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
015-0099-z;   http://link.springer.com/content/pdf/10.1007/
s11538-015-0099-z.pdf.

Ivorra:2015:CMM

[3077] Benjamin Ivorra, Diène Ngom, and Ángel M. Ramos. Be-CoDiS: a mathematical model to predict the risk of human diseases spread between countries — validation and application to the 2014–2015 Ebola virus disease epidemic. Bulletin of Mathematical Biology, 77(9):1668– 1704, September 2015. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-015-0100-x;   http://link.springer.com/content/
pdf/10.1007/s11538-015-0100-x.pdf.

Li:2015:EDS

[3078] Dan Li, Jing’an Cui, Meng Liu, and Shengqiang Liu. The evolutionary dynamics of stochastic epidemic model with nonlinear incidence rate. Bulletin of Mathematical Biology, 77(9):1705–1743, September 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
015-0101-9;   http://link.springer.com/content/pdf/10.1007/
s11538-015-0101-9.pdf.

Anderson:2015:LFS

[3079] David F. Anderson, Gheorghe Craciun, Manoj Gopalkrishnan, and Carsten Wiuf. Lyapunov functions, stationary distributions, and nonequilibrium potential for reaction networks. Bulletin of Mathematical Biology, 77(9):1744–1767, September 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-015-0102-8;   http://link.springer.com/
content/pdf/10.1007/s11538-015-0102-8.pdf.

vanWieringen:2015:THC

[3080] Wessel N. van Wieringen and Aad W. van der Vaart. Transcriptomic heterogeneity in cancer as a consequence of dysregulation of the gene– gene interaction network. Bulletin of Mathematical Biology, 77(9):1768– 1786, September 2015. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-015-0103-7;   http://link.springer.com/content/
pdf/10.1007/s11538-015-0103-7.pdf.

Jarrett:2015:MMM

[3081] Angela M. Jarrett, N. G. Cogan, and M. Y. Hussaini. Mathematical model for MRSA nasal carriage. Bulletin of Mathematical Biology, 77(9):1787–1812, September 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-015-0104-6;   http://link.springer.com/
content/pdf/10.1007/s11538-015-0104-6.pdf.

Nikmaneshi:2015:TPA

[3082] M. R. Nikmaneshi, B. Firoozabadi, and M. S. Saidi. Two-phase acto-cytosolic fluid flow in a moving keratocyte: a 2D continuum model. Bulletin of Mathematical Biology, 77(9):1813–1832, September 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
015-0105-5;   http://link.springer.com/content/pdf/10.1007/
s11538-015-0105-5.pdf.

Gedeon:2015:DSF

[3083] Tomas Gedeon and Patrick Murphy. Dynamics of simple food webs. Bulletin of Mathematical Biology, 77(10):1833–1853, October 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
015-0106-4;   http://link.springer.com/content/pdf/10.1007/
s11538-015-0106-4.pdf. See erratum [3121].

Young:2015:QEM

[3084] Glenn Young, Eunha Shim, and G. Bard Ermentrout. Qualitative effects of monovalent vaccination against rotavirus: a comparison of North America and South America. Bulletin of Mathematical Biology, 77(10): 1854–1885, October 2015. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-015-0107-3;   http://link.springer.com/content/
pdf/10.1007/s11538-015-0107-3.pdf.

Farkas:2015:RSS

[3085] J. Z. Farkas, A. Yu. Morozov, E. G. Arashkevich, and A. Nikishina. Revisiting the stability of spatially heterogeneous predator–prey systems under eutrophication. Bulletin of Mathematical Biology, 77(10): 1886–1908, October 2015. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-015-0108-2;   http://link.springer.com/content/
pdf/10.1007/s11538-015-0108-2.pdf.

Sherborne:2015:DMS

[3086] N. Sherborne, K. B. Blyuss, and I. Z. Kiss. Dynamics of multi-stage infections on networks. Bulletin of Mathematical Biology, 77(10):1909– 1933, October 2015. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-015-0109-1;   http://link.springer.com/content/
pdf/10.1007/s11538-015-0109-1.pdf.

Talkington:2015:ETG

[3087] Anne Talkington and Rick Durrett. Estimating tumor growth rates In Vivo. Bulletin of Mathematical Biology, 77(10):1934–1954, October 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
015-0110-8;   http://link.springer.com/content/pdf/10.1007/
s11538-015-0110-8.pdf.

Laguzet:2015:IVN

[3088] Laetitia Laguzet and Gabriel Turinici. Individual vaccination as Nash equilibrium in a SIR model with application to the 2009–2010 influenza A (H1N1) epidemic in France. Bulletin of Mathematical Biology, 77(10): 1955–1984, October 2015. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-015-0111-7;   http://link.springer.com/content/
pdf/10.1007/s11538-015-0111-7.pdf.

Birch:2015:ESM

[3089] Michael Birch and Benjamin M. Bolker. Evolutionary stability of minimal mutation rates in an evo-epidemiological model. Bulletin of Mathematical Biology, 77(11):1985–2003, November 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
015-0112-6;   http://link.springer.com/content/pdf/10.1007/
s11538-015-0112-6.pdf.

Bichara:2015:SSE

[3090] Derdei Bichara, Yun Kang, Carlos Castillo-Chavez, Richard Horan, and Charles Perrings. SIS and SIR epidemic models under virtual dispersal. Bulletin of Mathematical Biology, 77(11):2004–2034, November 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
015-0113-5;   http://link.springer.com/content/pdf/10.1007/
s11538-015-0113-5.pdf.

Croicu:2015:SLT

[3091] Ana-Maria Croicu. Short- and long–term optimal control of a mathematical model for HIV infection of CD4^{+}T cells. Bulletin of Mathematical Biology, 77(11):2035–2071, November 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
015-0114-4;   http://link.springer.com/content/pdf/10.1007/
s11538-015-0114-4.pdf.

Tian:2015:TPF

[3092] Canrong Tian. Turing pattern formation in a semiarid vegetation model with fractional-in-space diffusion. Bulletin of Mathematical Biology, 77(11):2072–2085, November 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-015-0116-2;   http://link.springer.com/
content/pdf/10.1007/s11538-015-0116-2.pdf.

Rahman:2015:MCB

[3093] Kazi A. Rahman, Rangarajan Sudarsan, and Hermann J. Eberl. A mixed-culture biofilm model with cross-diffusion. Bulletin of Mathematical Biology, 77(11):2086–2124, November 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-015-0117-1;
 http://link.springer.com/content/pdf/10.1007/s11538-015-0117-
1.pdf.

Phillips:2015:PTD

[3094] Austin Phillips and Mark Kot. Persistence in a two-dimensional moving-habitat model. Bulletin of Mathematical Biology, 77(11):2125– 2159, November 2015. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-015-0119-z;   http://link.springer.com/content/
pdf/10.1007/s11538-015-0119-z.pdf.

Whidden:2015:TDM

[3095] Mark Whidden, Nick Cogan, Matt Donahue, Fernando Navarrete, and Leonardo De La Fuente. A two-dimensional multiphase model of biofilm formation in microfluidic chambers. Bulletin of Mathematical Biology, 77(12):2161–2179, December 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
015-0115-3;   http://link.springer.com/content/pdf/10.1007/
s11538-015-0115-3.pdf.

Samal:2015:GMM

[3096] Satya Swarup Samal, Dima Grigoriev, Holger Fröhlich, Andreas Weber, and Ovidiu Radulescu. A geometric method for model reduction of biochemical networks with polynomial rate functions. Bulletin of Mathematical Biology, 77(12):2180–2211, December 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
015-0118-0;   http://link.springer.com/content/pdf/10.1007/
s11538-015-0118-0.pdf.

Knipl:2015:PIV

[3097] Diána Knipl and Seyed M. Moghadas. The potential impact of vaccination on the dynamics of dengue infections. Bulletin of Mathematical Biology, 77(12):2212–2230, December 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-015-0120-6;
 http://link.springer.com/content/pdf/10.1007/s11538-015-0120-
6.pdf.

Kong:2015:IMC

[3098] Jude D. Kong, Chaochao Jin, and Hao Wang. The inverse method for a childhood infectious disease model with its application to pre-vaccination and post-vaccination measles data. Bulletin of Mathematical Biology, 77(12):2231–2263, December 2015.

CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
015-0121-5;   http://link.springer.com/content/pdf/10.1007/
s11538-015-0121-5.pdf.

Wang:2015:MMC

[3099] Qiming Wang, Nathan Gold, Martin G. Frasch, Huaxiong Huang, Marc Thiriet, and Xiaogang Wang. Mathematical model of cardiovascular and metabolic responses to umbilical cord occlusions in fetal sheep. Bulletin of Mathematical Biology, 77(12):2264–2293, December 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
015-0122-4;   http://link.springer.com/content/pdf/10.1007/
s11538-015-0122-4.pdf.

Krishna:2015:CLG

[3100] Nitin A. Krishna, Hannah M. Pennington, Canaan D. Coppola, Marisa C. Eisenberg, and Richard C. Schugart. Connecting local and global sensitivities in a mathematical model for wound healing. Bulletin of Mathematical Biology, 77(12):2294–2324, December 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
015-0123-3;   http://link.springer.com/content/pdf/10.1007/
s11538-015-0123-3.pdf.

Sekerci:2015:MMP

[3101] Yadigar Sekerci and Sergei Petrovskii. Mathematical modelling of plankton–oxygen dynamics under the climate change. Bulletin of Mathematical Biology, 77(12):2325–2353, December 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
015-0126-0;   http://link.springer.com/content/pdf/10.1007/
s11538-015-0126-0.pdf.

Serva:2015:SMI

[3102] Maurizio Serva. A stochastic model for the interbreeding of two populations continuously sharing the same habitat. Bulletin of Mathematical Biology, 77(12):2354–2365, December 2015. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-015-0127-z;
 http://link.springer.com/content/pdf/10.1007/s11538-015-0127-
z.pdf.

Mesterton-Gibbons:2015:ELT

[3103] Mike Mesterton-Gibbons and Yao Dai. An effect of landmarks on territory shape in a convict cichlid. Bulletin of Mathematical Biology, 77(12): 2366–2378, December 2015. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-015-0130-4;   http://link.springer.com/content/
pdf/10.1007/s11538-015-0130-4.pdf.

Laubenbacher:2016:E

[3104] Reinhard Laubenbacher and Alan Hastings. Editorial. Bulletin of Mathematical Biology, 78(1):1–3, January 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/s11538-015-0134-0;   http://link.springer.
com/content/pdf/10.1007/s11538-015-0134-0.pdf.

Asih:2016:DHI

[3105] Tri Sri Noor Asih, Suzanne Lenhart, Steven Wise, Lina Aryati, F. Adi-Kusumo, Mardiah S. Hardianti, and Jonathan Forde. The dynamics of HPV infection and cervical cancer cells. Bulletin of Mathematical Biology, 78(1):4–20, January 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-015-0124-2;
 http://link.springer.com/content/pdf/10.1007/s11538-015-0124-
2.pdf.

Gross:2016:ASB

[3106] Elizabeth Gross, Heather A. Harrington, Zvi Rosen, and Bernd Sturmfels. Algebraic systems biology: a case study for the Wnt pathway. Bulletin of Mathematical Biology, 78(1):21–51, January 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
015-0125-1;   http://link.springer.com/content/pdf/10.1007/
s11538-015-0125-1.pdf.

Cooper:2016:IBM

[3107] Laura J. Cooper, James P. Heppell, Geraldine F. Clough, Bharathram Ganapathisubramani, and Tiina Roose. An image-based model of fluid flow through lymph nodes. Bulletin of Mathematical Biology, 78(1): 52–71, January 2016. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-015-0128-y;   http://link.springer.com/content/
pdf/10.1007/s11538-015-0128-y.pdf.

Reimer:2016:ACD

[3108] Jody R. Reimer, Michael B. Bonsall, and Philip K. Maini. Approximating the critical domain size of integrodifference equations. Bulletin of Mathematical Biology, 78(1):72–109, January 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-015-0129-x;
 http://link.springer.com/content/pdf/10.1007/s11538-015-0129-
x.pdf.

Innocentini:2016:PSD

[3109] Guilherme C. P. Innocentini, Michael Forger, Ovidiu Radulescu, and Fernando Antoneli. Protein synthesis driven by dynamical stochastic transcription. Bulletin of Mathematical Biology, 78(1):110–131, January 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
015-0131-3;   http://link.springer.com/content/pdf/10.1007/
s11538-015-0131-3.pdf.

Semple:2016:PNE

[3110] Charles Semple. Phylogenetic networks with every embedded phylogenetic tree a base tree. Bulletin of Mathematical Biology, 78(1): 132–137, January 2016. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-015-0132-2;   http://link.springer.com/content/
pdf/10.1007/s11538-015-0132-2.pdf.

Wu:2016:ISB

[3111] Xiaotian Wu, Gergely Röst, and Xingfu Zou. Impact of spring bird migration on the range expansion of Ixodes scapularis tick population. Bulletin of Mathematical Biology, 78(1):138–168, January 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
015-0133-1;   http://link.springer.com/content/pdf/10.1007/
s11538-015-0133-1.pdf.

Waters:2016:ZTW

[3112] Edward K. Waters, Andrew J. Hamilton, Harvinder S. Sidhu, Leesa A. Sidhu, and Michelle Dunbar. Zoonotic transmission of waterborne disease: a mathematical model. Bulletin of Mathematical Biology, 78(1): 169–183, January 2016. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-015-0136-y;   http://link.springer.com/content/
pdf/10.1007/s11538-015-0136-y.pdf.

Lopez:2016:MIE

[3113] Luis Fernandez Lopez, Marcos Amaku, Francisco Antonio Bezerra Coutinho, Mikkel Quam, Marcelo Nascimento Burattini, Claudio José Struchiner, Annelies Wilder-Smith, and Eduardo Massad. Modeling importations and exportations of infectious diseases via travelers. Bulletin of Mathematical Biology, 78(2):185–209, February 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
015-0135-z;   http://link.springer.com/content/pdf/10.1007/
s11538-015-0135-z.pdf.

Mellor:2016:GMA

[3114] Nathan Mellor, Malcolm J. Bennett, and John R. King. GH3-Mediated auxin conjugation can result in either transient or oscillatory transcriptional auxin responses. Bulletin of Mathematical Biology, 78(2): 210–234, February 2016. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-015-0137-x;   http://link.springer.com/content/
pdf/10.1007/s11538-015-0137-x.pdf.

Rapti:2016:EIE

[3115] Z. Rapti and C. E. Cáceres. Effects of intrinsic and extrinsic host mortality on disease spread. Bulletin of Mathematical Biology, 78(2): 235–253, February 2016. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-016-0141-9;   http://link.springer.com/content/
pdf/10.1007/s11538-016-0141-9.pdf.

Kim:2016:EDS

[3116] Yong-Jung Kim and Ohsang Kwon. Evolution of dispersal with starvation measure and coexistence. Bulletin of Mathematical Biology, 78(2): 254–279, February 2016. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-016-0142-8;   http://link.springer.com/content/
pdf/10.1007/s11538-016-0142-8.pdf.

Plank:2016:ESS

[3117] Michael J. Plank, Jonathan W. Pitchford, and Alex James. Evolutionarily stable strategies for fecundity and swimming speed of fish. Bulletin of Mathematical Biology, 78(2):280–292, February 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-

016-0143-7;   http://link.springer.com/content/pdf/10.1007/
s11538-016-0143-7.pdf.

Lydeamore:2016:CAA

[3118] Michael Lydeamore, Nigel Bean, Andrew J. Black, and Joshua V. Ross. Choice of antiviral allocation scheme for pandemic influenza depends on strain transmissibility, delivery delay and stockpile size. Bulletin of Mathematical Biology, 78(2):293–321, February 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
016-0144-6;   http://link.springer.com/content/pdf/10.1007/
s11538-016-0144-6.pdf.

Wang:2016:DHM

[3119] Xia Wang, Xinyu Song, Sanyi Tang, and Libin Rong. Dynamics of an HIV model with multiple infection stages and treatment with different drug classes. Bulletin of Mathematical Biology, 78(2):322–349, February 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
016-0145-5;   http://link.springer.com/content/pdf/10.1007/
s11538-016-0145-5.pdf.

deRoos:2016:ENE

[3120] A. M. de Roos, O. Diekmann, P. Getto, and M. A. Kirkilionis. Erratum to: Numerical Equilibrium Analysis for Structured Consumer Resource Models. Bulletin of Mathematical Biology, 78(2):350–351, February 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
015-0138-9;   http://link.springer.com/content/pdf/10.1007/
s11538-015-0138-9.pdf. See [2427].

Gedeon:2016:EDS

[3121] Tomas Gedeon and Patrick Murphy. Erratum to: Dynamics of Simple Food Webs. Bulletin of Mathematical Biology, 78(2):352, February 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
015-0139-8;   http://link.springer.com/content/pdf/10.1007/
s11538-015-0139-8.pdf. See [3083].

Huang:2016:TTD

[3122] Qihua Huang, Hao Wang, Anthony Ricciardi, and Mark A. Lewis. Temperature- and turbidity-dependent competitive interactions between invasive freshwater mussels. Bulletin of Mathematical Biology, 78(3):

353–380, March 2016. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-016-0146-4;   http://link.springer.com/content/
pdf/10.1007/s11538-016-0146-4.pdf.

Zhao:2016:MMP

[3123] Jiantao Zhao, Jianjun Paul Tian, and Junjie Wei. Minimal model of plankton systems revisited with spatial diffusion and maturation delay. Bulletin of Mathematical Biology, 78(3):381–412, March 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
016-0147-3;   http://link.springer.com/content/pdf/10.1007/
s11538-016-0147-3.pdf.

Ippolito:2016:ASS

[3124] Stephen Ippolito, Vincent Naudot, and Erik G. Noonburg. Alternative stable states, coral reefs, and smooth dynamics with a Kick. Bulletin of Mathematical Biology, 78(3):413–435, March 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
016-0148-2;   http://link.springer.com/content/pdf/10.1007/
s11538-016-0148-2.pdf.

Brown:2016:MMM

[3125] L. E. Brown, A. M. Middleton, J. R. King, and M. Loose. Multicellular mathematical modelling of mesendoderm formation in amphibians. Bulletin of Mathematical Biology, 78(3):436–467, March 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
016-0150-8;   http://link.springer.com/content/pdf/10.1007/
s11538-016-0150-8.pdf.

Angstmann:2016:FOR

[3126] C. N. Angstmann, B. I. Henry, and A. V. McGann. A fractional order recovery SIR model from a stochastic process. Bulletin of Mathematical Biology, 78(3):468–499, March 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-016-0151-7;
 http://link.springer.com/content/pdf/10.1007/s11538-016-0151-
7.pdf.

Moore:2016:MMT

[3127] James R. Moore and Fred Adler. A mathematical model of T1D acceleration and delay by viral infection. Bulletin of Mathematical Biology, 78 (3):500–530, March 2016. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-016-0152-6;   http://link.springer.com/content/
pdf/10.1007/s11538-016-0152-6.pdf.

Wang:2016:SYC

[3128] M. Wang, Y. Huang, and Z. Wu. Simulation of yeast cooperation in 2D. Bulletin of Mathematical Biology, 78(3):531–555, March 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
016-0153-5;   http://link.springer.com/content/pdf/10.1007/
s11538-016-0153-5.pdf.

Thygesen:2016:DAB

[3129] Uffe Høgsbro Thygesen. A diffusion approximation based on renewal processes with applications to strongly biased run-tumble motion. Bulletin of Mathematical Biology, 78(3):556–579, March 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
016-0155-3;   http://link.springer.com/content/pdf/10.1007/
s11538-016-0155-3.pdf.

Potomkin:2016:ERP

[3130] Mykhailo Potomkin, Shawn D. Ryan, and Leonid Berlyand. Effective rheological properties in semi-dilute bacterial suspensions. Bulletin of Mathematical Biology, 78(3):580–615, March 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-016-0156-2;
 http://link.springer.com/content/pdf/10.1007/s11538-016-0156-
2.pdf.

Cao:2016:SST

[3131] Youfang Cao, Anna Terebus, and Jie Liang. State space truncation with quantified errors for accurate solutions to discrete chemical master equation. Bulletin of Mathematical Biology, 78(4):617–661, April 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
016-0149-1;   http://link.springer.com/content/pdf/10.1007/
s11538-016-0149-1.pdf.

Lv:2016:MRE

[3132] Dongyu Lv, Meng Fan, Yun Kang, and Krystal Blanco. Modeling refuge effect of submerged macrophytes in lake system. Bulletin of Mathematical Biology, 78(4):662–694, April 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-016-0154-4;
 http://link.springer.com/content/pdf/10.1007/s11538-016-0154-
4.pdf.

Hamelin:2016:MFS

[3133] Frédéric M. Hamelin, François Castella, Valentin Doli, Benoît Marçais, Virginie Ravigné, and Mark A. Lewis. Mate finding, sexual spore production, and the spread of fungal plant parasites. Bulletin of Mathematical Biology, 78(4):695–712, April 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-016-0157-1;
 http://link.springer.com/content/pdf/10.1007/s11538-016-0157-
1.pdf.

Bodo:2016:SEP

[3134] Ágnes Bodó, Gyula Y. Katona, and Péter L. Simon. SIS epidemic propagation on hypergraphs. Bulletin of Mathematical Biology, 78(4):713–735, April 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
016-0158-0;   http://link.springer.com/content/pdf/10.1007/
s11538-016-0158-0.pdf.

Sadovsky:2016:MPP

[3135] Michael Sadovsky and Mariya Senashova. Model of prey–predator dynamics with reflexive spatial behaviour of species based on optimal migration. Bulletin of Mathematical Biology, 78(4):736–753, April 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
016-0159-z;   http://link.springer.com/content/pdf/10.1007/
s11538-016-0159-z.pdf.

Konstorum:2016:FRC

[3136] Anna Konstorum, Thomas Hillen, and John Lowengrub. Feedback regulation in a cancer stem cell model can cause an Allee effect. Bulletin of Mathematical Biology, 78(4):754–785, April 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-

016-0161-5;   http://link.springer.com/content/pdf/10.1007/
s11538-016-0161-5.pdf.

Zanetti:2016:MAG

[3137] Joao Paulo Pereira Zanetti, Priscila Biller, and Joao Meidanis. Median approximations for genomes modeled as matrices. Bulletin of Mathematical Biology, 78(4):786–814, April 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-016-0162-4;
 http://link.springer.com/content/pdf/10.1007/s11538-016-0162-
4.pdf.

Fowler:2016:DAL

[3138] A. C. Fowler and T. Déirdre Hollingsworth. The dynamics of Ascaris lumbricoides infections. Bulletin of Mathematical Biology, 78(4):815–833, April 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
016-0164-2;   http://link.springer.com/content/pdf/10.1007/
s11538-016-0164-2.pdf. See correction [3427].

Karev:2016:MME

[3139] G. P. Karev and I. Kareva. Mathematical modeling of extinction of inhomogeneous populations. Bulletin of Mathematical Biology, 78(4): 834–858, April 2016. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-016-0166-0;   http://link.springer.com/content/
pdf/10.1007/s11538-016-0166-0.pdf.

Leffler:2016:MMG

[3140] Shoshana R. Leffler, Emilie Legué, Orlando Aristizábal, Alexandra L. Joyner, Charles S. Peskin, and Daniel H. Turnbull. A mathematical model of granule cell generation during mouse cerebellum development. Bulletin of Mathematical Biology, 78(5):859–878, May 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
016-0163-3;   http://link.springer.com/content/pdf/10.1007/
s11538-016-0163-3.pdf.

Kang:2016:DMT

[3141] Yun Kang and Guy Theraulaz. Dynamical models of task organization in social insect colonies. Bulletin of Mathematical Biology, 78(5):879–915, May 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-

016-0165-1;   http://link.springer.com/content/pdf/10.1007/
s11538-016-0165-1.pdf.

Dorsett:2016:ORU

[3142] Chasity Dorsett, Hyunju Oh, Marie Laura Paulemond, and Jan Rychtář. Optimal repellent usage to combat dengue fever. Bulletin of Mathematical Biology, 78(5):916–922, May 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-016-0167-z;
 http://link.springer.com/content/pdf/10.1007/s11538-016-0167-
z.pdf.

Sgouralis:2016:TFA

[3143] Ioannis Sgouralis, Vasileios Maroulas, and Anita T. Layton. Transfer function analysis of dynamic blood flow control in the rat kidney. Bulletin of Mathematical Biology, 78(5):923–960, May 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
016-0168-y;   http://link.springer.com/content/pdf/10.1007/
s11538-016-0168-y.pdf.

Anaya:2016:DIT

[3144] Maria Anaya, Olga Anipchenko-Ulaj, Aisha Ashfaq, Joyce Chiu, Mahedi Kaiser, Max Shoji Ohsawa, Megan Owen, Ella Pavlechko, Katherine St. John, Shivam Suleria, Keith Thompson, and Corrine Yap. On determining if tree-based networks contain fixed trees. Bulletin of Mathematical Biology, 78(5):961–969, May 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
016-0169-x;   http://link.springer.com/content/pdf/10.1007/
s11538-016-0169-x.pdf.

Amaku:2016:ESH

[3145] Marcos Amaku, Marcelo Nascimento Burattini, Francisco Antonio Bezerra Coutinho, Luis Fernandez Lopez, Fabio Mesquita, Marcelo Contardo Moscoso Naveira, Gerson Fernando Mendes Pereira, Melina Érica Santos, and Eduardo Massad. Estimating the size of the HCV infection prevalence: A modeling approach using the incidence of cases reported to an official notification system. Bulletin of Mathematical Biology, 78 (5):970–990, May 2016. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-016-0170-4;   http://link.springer.com/content/
pdf/10.1007/s11538-016-0170-4.pdf.

Semenov:2016:EQM

[3146] Yuri S. Semenov and Artem S. Novozhilov. On Eigen’s quasispecies model, two-valued fitness landscapes, and isometry groups acting on finite metric spaces. Bulletin of Mathematical Biology, 78(5):991–1038, May 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
016-0172-2;   http://link.springer.com/content/pdf/10.1007/
s11538-016-0172-2.pdf.

Martini:2016:MCM

[3147] Johannes W. R. Martini, Martin Schlather, and Stefan Schütz. A model for carrier-mediated biological signal transduction based on equilibrium ligand binding theory. Bulletin of Mathematical Biology, 78 (5):1039–1057, May 2016. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-016-0173-1;   http://link.springer.com/content/
pdf/10.1007/s11538-016-0173-1.pdf.

Urheim:2016:CLO

[3148] Ellen Urheim, Eric Ford, and Katherine St. John. Characterizing local optima for maximum parsimony. Bulletin of Mathematical Biology, 78 (5):1058–1075, May 2016. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-016-0174-0;   http://link.springer.com/content/
pdf/10.1007/s11538-016-0174-0.pdf.

Fan:2016:CPP

[3149] Gaoyang Fan, Bree Cummins, and Tomáš Gedeon. Convergence properties of posttranslationally modified protein–protein switching networks with fast decay rates. Bulletin of Mathematical Biology, 78(6):1077–1120, June 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
016-0175-z;   http://link.springer.com/content/pdf/10.1007/
s11538-016-0175-z.pdf.

Patsatzis:2016:AAT

[3150] Dimitris G. Patsatzis, Dimitris T. Maris, and Dimitris A. Goussis. Asymptotic analysis of a target-mediated drug disposition model: Algorithmic and traditional approaches. Bulletin of Mathematical Biology, 78 (6):1121–1161, June 2016. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-016-0176-y;   http://link.springer.com/content/
pdf/10.1007/s11538-016-0176-y.pdf.

Rhodes:2016:MMR

[3151] Adam Rhodes and Thomas Hillen. Mathematical modeling of the role of Survivin on dedifferentiation and radioresistance in cancer. Bulletin of Mathematical Biology, 78(6):1162–1188, June 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
016-0177-x;   http://link.springer.com/content/pdf/10.1007/
s11538-016-0177-x.pdf.

Pearson:2016:EIR

[3152] Taliesin Pearson, Jonathan A. D. Wattis, John R. King, Ian A. Mac-Donald, and Dawn J. Mazzatti. The effects of insulin resistance on individual tissues: an application of a mathematical model of metabolism in humans. Bulletin of Mathematical Biology, 78(6):1189–1217, June 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
016-0181-1;   http://link.springer.com/content/pdf/10.1007/
s11538-016-0181-1.pdf.

Duran:2016:TDR

[3153] María Rosa Durán, Ana Podolski-Renić, Arturo Álvarez-Arenas, Jelena Dinić, Juan Belmonte-Beitia, Milica Pešić, and Víctor M. Pérez-García. Transfer of drug resistance characteristics between cancer cell subpopulations: a study using simple mathematical models. Bulletin of Mathematical Biology, 78(6):1218–1237, June 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
016-0182-0;   http://link.springer.com/content/pdf/10.1007/
s11538-016-0182-0.pdf.

Cerf:2016:QCD

[3154] Raphaël Cerf and Joseba Dalmau. Quasispecies on class-dependent fitness landscapes. Bulletin of Mathematical Biology, 78(6):1238–1258, June 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
016-0184-y;   http://link.springer.com/content/pdf/10.1007/
s11538-016-0184-y.pdf.

Kura:2016:GTW

[3155] Klodeta Kura, Mark Broom, and Anne Kandler. A game-theoretical winner and loser model of dominance hierarchy formation. Bulletin of Mathematical Biology, 78(6):1259–1290, June 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-016-0186-9;
 http://link.springer.com/content/pdf/10.1007/s11538-016-0186-
9.pdf.

Gedeon:2016:CFS

[3156] Tomas Gedeon, Mark Pernarowski, and Adam Wilander. Cyclic feedback systems with quorum sensing coupling. Bulletin of Mathematical Biology, 78(6):1291–1317, June 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-016-0187-8;   http://link.springer.com/
content/pdf/10.1007/s11538-016-0187-8.pdf.

Chen:2016:MGM

[3157] Ying Chen, Brendan C. Fry, and Anita T. Layton. Modeling glucose metabolism in the kidney. Bulletin of Mathematical Biology, 78(6): 1318–1336, June 2016. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-016-0188-7;   http://link.springer.com/content/
pdf/10.1007/s11538-016-0188-7.pdf.

Li:2016:PSS

[3158] Bingtuan Li, Sharon Bewick, Michael R. Barnard, and William F. Fagan. Persistence and spreading speeds of integro–difference equations with an expanding or contracting habitat. Bulletin of Mathematical Biology, 78 (7):1337–1379, July 2016. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-016-0180-2;   http://link.springer.com/content/
pdf/10.1007/s11538-016-0180-2.pdf.

Boronovsky:2016:CSG

[3159] S. E. Boronovsky and Y. R. Nartsissov. Computer simulator of glycine receptor activity: a new window into a virtual world. Bulletin of Mathematical Biology, 78(7):1380–1393, July 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-016-0183-z;
 http://link.springer.com/content/pdf/10.1007/s11538-016-0183-
z.pdf.

Camacho:2016:MMR

[3160] Erika T. Camacho, Thierry Léveillard, José-Alain Sahel, and Stephen Wirkus. Mathematical model of the role of RdCVF in the coexistence of rods and cones in a healthy eye. Bulletin of Mathematical Biology, 78 (7):1394–1409, July 2016. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-016-0185-x;   http://link.springer.com/content/
pdf/10.1007/s11538-016-0185-x.pdf.

Castillo:2016:THB

[3161] Jorge A. Castillo, Faustino Sánchez-Garduño, and Pablo Padilla. A Turing–Hopf bifurcation scenario for pattern formation on growing domains. Bulletin of Mathematical Biology, 78(7):1410–1449, July 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
016-0189-6;   http://link.springer.com/content/pdf/10.1007/
s11538-016-0189-6.pdf.

Dera:2016:ARI

[3162] Dimah Dera, Nidhal Bouaynaya, and Hassan M. Fathallah-Shaykh. Automated robust image segmentation: Level set method using nonnegative matrix factorization with application to brain MRI. Bulletin of Mathematical Biology, 78(7):1450–1476, July 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-016-0190-0;
 http://link.springer.com/content/pdf/10.1007/s11538-016-0190-
0.pdf.

Alter:2016:PGC

[3163] Idan Alter and Yoram Louzoun. Population growth combined with wide offspring distributions can increase fixation rate and reduce genetic diversity. Bulletin of Mathematical Biology, 78(7):1477–1492, July 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
016-0191-z;   http://link.springer.com/content/pdf/10.1007/
s11538-016-0191-z.pdf.

Pal:2016:SBA

[3164] D. Pal, G. S. Mahapatra, and G. P. Samanta. Stability and bionomic analysis of fuzzy prey–predator harvesting model in presence of toxicity: a dynamic approach. Bulletin of Mathematical Biology, 78(7):1493–1519, July 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
016-0192-y;   http://link.springer.com/content/pdf/10.1007/
s11538-016-0192-y.pdf.

Matache:2016:LRB

[3165] Mihaela T. Matache and Valentin Matache. Logical reduction of biological networks to their most determinative components. Bulletin of Mathematical Biology, 78(7):1520–1545, July 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-016-0193-x;
 http://link.springer.com/content/pdf/10.1007/s11538-016-0193-
x.pdf.

Sanz:2016:NBA

[3166] Julia Sánchez Sanz and Philipp Getto. Numerical bifurcation analysis of physiologically structured populations: Consumer–resource, cannibalistic and trophic models. Bulletin of Mathematical Biology, 78(7): 1546–1584, July 2016. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-016-0194-9;   http://link.springer.com/content/
pdf/10.1007/s11538-016-0194-9.pdf.

Perez-Velazquez:2016:MMB

[3167] Judith Pérez-Velázquez, Meltem Gölgeli, and Rodolfo García-Contreras. Mathematical modelling of bacterial quorum sensing: a review. Bulletin of Mathematical Biology, 78(8):1585–1639, August 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
016-0160-6;   http://link.springer.com/content/pdf/10.1007/
s11538-016-0160-6.pdf.

Lester:2016:EML

[3168] Christopher Lester, Ruth E. Baker, Michael B. Giles, and Christian A. Yates. Extending the multi-level method for the simulation of stochastic biological systems. Bulletin of Mathematical Biology, 78(8):1640– 1677, August 2016. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-016-0178-9;   http://link.springer.com/content/
pdf/10.1007/s11538-016-0178-9.pdf.

Gourley:2016:MMB

[3169] Stephen A. Gourley, Jing Li, and Xingfu Zou. A mathematical model for biocontrol of the invasive weed Fallopia japonica. Bulletin of Mathematical Biology, 78(8):1678–1702, August 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-

016-0195-8;   http://link.springer.com/content/pdf/10.1007/
s11538-016-0195-8.pdf.

Chen:2016:NPN

[3170] Duan Chen. A new Poisson–Nernst–Planck model with ion–water interactions for charge transport in ion channels. Bulletin of Mathematical Biology, 78(8):1703–1726, August 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-016-0196-7;
 http://link.springer.com/content/pdf/10.1007/s11538-016-0196-
7.pdf.

Fan:2016:DPP

[3171] Meng Fan, Ping Wu, Zhilan Feng, and Robert K. Swihart. Dynamics of predator–prey metapopulations with Allee effects. Bulletin of Mathematical Biology, 78(8):1727–1748, August 2016. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-016-0197-6;
 http://link.springer.com/content/pdf/10.1007/s11538-016-0197-
6.pdf.

Parvinen:2016:ESS

[3172] Kalle Parvinen and Åke Brännström. Evolution of site-selection stabilizes population dynamics, promotes even distribution of individuals, and occasionally causes evolutionary suicide. Bulletin of Mathematical Biology, 78(8):1749–1772, August 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-016-0198-5;
 http://link.springer.com/content/pdf/10.1007/s11538-016-0198-
5.pdf.

Gambette:2016:DBL

[3173] Philippe Gambette, Leo van Iersel, Steven Kelk, Fabio Pardi, and Celine Scornavacca. Do branch lengths help to locate a tree in a phylogenetic network? Bulletin of Mathematical Biology, 78(9):1773–1795, September 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
016-0199-4;   http://link.springer.com/content/pdf/10.1007/
s11538-016-0199-4.pdf.

Tuncer:2016:SPI

[3174] Necibe Tuncer, Hayriye Gulbudak, Vincent L. Cannataro, and Maia Martcheva. Structural and practical identifiability issues of immuno– epidemiological vector–host models with application to Rift Valley fever. Bulletin of Mathematical Biology, 78(9):1796–1827, September 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
016-0200-2;   http://link.springer.com/content/pdf/10.1007/
s11538-016-0200-2.pdf.

Beams:2016:HIH

[3175] Alexander B. Beams, Damon J. A. Toth, Karim Khader, and Frederick R. Adler. Harnessing intra-host strain competition to limit antibiotic resistance: Mathematical model results. Bulletin of Mathematical Biology, 78(9):1828–1846, September 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
016-0201-1;   http://link.springer.com/content/pdf/10.1007/
s11538-016-0201-1.pdf.

Pomi:2016:PNR

[3176] Andrés Pomi. A possible neural representation of mathematical group structures. Bulletin of Mathematical Biology, 78(9):1847–1865, September 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
016-0202-0;   http://link.springer.com/content/pdf/10.1007/
s11538-016-0202-0.pdf.

Bouhours:2016:CCI

[3177] Juliette Bouhours and Mark A. Lewis. Climate change and integrodifference equations in a stochastic environment. Bulletin of Mathematical Biology, 78(9):1866–1903, September 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-016-0203-z;
 http://link.springer.com/content/pdf/10.1007/s11538-016-0203-
z.pdf.

Bianchi:2016:STM

[3178] Arianna Bianchi, Kevin J. Painter, and Jonathan A. Sherratt. Spatiotemporal models of lymphangiogenesis in wound healing. Bulletin of Mathematical Biology, 78(9):1904–1941, September 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
016-0205-x;   http://link.springer.com/content/pdf/10.1007/
s11538-016-0205-x.pdf.

OConnell:2016:EGT

[3179] Rosemary O’Connell and Yoichiro Mori. Effects of glia in a triphasic continuum model of cortical spreading depression. Bulletin of Mathematical Biology, 78(10):1943–1967, October 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-016-0206-9;
 http://link.springer.com/content/pdf/10.1007/s11538-016-0206-
9.pdf.

Zhang:2016:MEA

[3180] Xianghong Zhang, Sanyi Tang, Robert A. Cheke, and Huaiping Zhu. Modeling the effects of augmentation strategies on the control of dengue fever with an impulsive differential equation. Bulletin of Mathematical Biology, 78(10):1968–2010, October 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
016-0208-7;   http://link.springer.com/content/pdf/10.1007/
s11538-016-0208-7.pdf.

Falcon-Lezama:2016:DDP

[3181] Jorge A. Falcón-Lezama, Ruth A. Martínez-Vega, Pablo A. Kuri-Morales, José Ramos-Castañeda, and Ben Adams. Day-to-day population movement and the management of dengue epidemics. Bulletin of Mathematical Biology, 78(10):2011–2033, October 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
016-0209-6;   http://link.springer.com/content/pdf/10.1007/
s11538-016-0209-6.pdf.

Yurk:2016:HDD

[3182] Brian P. Yurk. Homogenization of a directed dispersal model for animal movement in a heterogeneous environment. Bulletin of Mathematical Biology, 78(10):2034–2056, October 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-016-0210-0;
 http://link.springer.com/content/pdf/10.1007/s11538-016-0210-
0.pdf.

Zhao:2016:IPR

[3183] Yuqin Zhao, Daniel T. Wood, Hristo V. Kojouharov, Yang Kuang, and Dobromir T. Dimitrov. Impact of population recruitment on the HIV epidemics and the effectiveness of HIV prevention interventions. Bulletin of Mathematical Biology, 78(10):2057–2090, October 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
016-0211-z;   http://link.springer.com/content/pdf/10.1007/
s11538-016-0211-z.pdf.

Eftimie:2016:MMI

[3184] Raluca Eftimie, Joseph J. Gillard, and Doreen A. Cantrell. Mathematical models for immunology: Current state of the art and future research directions. Bulletin of Mathematical Biology, 78(10):2091–2134, October 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
016-0214-9;   http://link.springer.com/content/pdf/10.1007/
s11538-016-0214-9.pdf.

Ptashnyk:2016:IMO

[3185] Mariya Ptashnyk and Brian Seguin. The impact of microfibril orientations on the biomechanics of plant cell walls and tissues. Bulletin of Mathematical Biology, 78(11):2135–2164, November 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
016-0207-8;   http://link.springer.com/content/pdf/10.1007/
s11538-016-0207-8.pdf.

Garnier:2016:EUC

[3186] Jimmy Garnier and Mark A. Lewis. Expansion under climate change: The genetic consequences. Bulletin of Mathematical Biology, 78(11): 2165–2185, November 2016. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-016-0213-x;   http://link.springer.com/content/
pdf/10.1007/s11538-016-0213-x.pdf.

Kozlov:2016:SEU

[3187] Vladimir Kozlov, Sergey Vakulenko, and Uno Wennergren. Stability of ecosystems under invasions. Bulletin of Mathematical Biology, 78(11): 2186–2211, November 2016. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-016-0216-7;   http://link.springer.com/content/
pdf/10.1007/s11538-016-0216-7.pdf.

Vyska:2016:CDB

[3188] Martin Vyska and Christopher Gilligan. Complex dynamical behaviour in an epidemic model with control. Bulletin of Mathematical Biology, 78(11):2212–2227, November 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-016-0217-6;   http://link.springer.com/
content/pdf/10.1007/s11538-016-0217-6.pdf.

Baca-Carrasco:2016:SME

[3189] David Baca-Carrasco and Jorge X. Velasco-Hernández. Sex, mosquitoes and epidemics: an evaluation of Zika disease dynamics. Bulletin of Mathematical Biology, 78(11):2228–2242, November 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
016-0219-4;   http://link.springer.com/content/pdf/10.1007/
s11538-016-0219-4.pdf.

Nicholson:2016:UAC

[3190] Michael D. Nicholson and Tibor Antal. Universal asymptotic clone size distribution for general population growth. Bulletin of Mathematical Biology, 78(11):2243–2276, November 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-016-0221-x;
 http://link.springer.com/content/pdf/10.1007/s11538-016-0221-
x.pdf.

Binny:2016:CCB

[3191] Rachelle N. Binny, Alex James, and Michael J. Plank. Collective cell behaviour with neighbour-dependent proliferation, death and directional bias. Bulletin of Mathematical Biology, 78(11):2277–2301, November 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
016-0222-9;   http://link.springer.com/content/pdf/10.1007/
s11538-016-0222-9.pdf.

Hastings:2016:E

[3192] Alan Hastings and Reinhard Laubenbacher. Editorial. Bulletin of Mathematical Biology, 78(12):2303, December 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-016-0223-8;
 http://link.springer.com/content/pdf/10.1007/s11538-016-0223-
8.pdf.

Craig:2016:MMG

[3193] M. Craig, A. R. Humphries, and M. C. Mackey. A mathematical model of granulopoiesis incorporating the negative feedback dynamics and kinetics of G-CSF/ neutrophil binding and internalization. Bulletin of Mathematical Biology, 78(12):2304–2357, December 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
016-0179-8;   http://link.springer.com/content/pdf/10.1007/
s11538-016-0179-8.pdf.

Bernard:2016:MBG

[3194] Samuel Bernard. Moving the boundaries of granulopoiesis modelling. Bulletin of Mathematical Biology, 78(12):2358–2363, December 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
016-0215-8;   http://link.springer.com/content/pdf/10.1007/
s11538-016-0215-8.pdf.

Mushanyu:2016:MDA

[3195] J. Mushanyu, F. Nyabadza, G. Muchatibaya, and A. G. R. Stewart. Modelling drug abuse epidemics in the presence of limited rehabilitation capacity. Bulletin of Mathematical Biology, 78(12):2364–2389, December 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
016-0218-5;   http://link.springer.com/content/pdf/10.1007/
s11538-016-0218-5.pdf.

Anderson:2016:PFS

[3196] David F. Anderson and Simon L. Cotter. Product-form stationary distributions for deficiency zero networks with non-mass action kinetics. Bulletin of Mathematical Biology, 78(12):2390–2407, December 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
016-0220-y;   http://link.springer.com/content/pdf/10.1007/
s11538-016-0220-y.pdf.

Wallace:2016:EIN

[3197] Rodrick Wallace. Environmental induction of neurodevelopmental disorders. Bulletin of Mathematical Biology, 78(12):2408–2426, December 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
016-0226-5;   http://link.springer.com/content/pdf/10.1007/
s11538-016-0226-5.pdf.

Britton:2016:NEM

[3198] Tom Britton, David Juher, and Joan Saldaña. A network epidemic model with preventive rewiring: Comparative analysis of the initial phase. Bulletin of Mathematical Biology, 78(12):2427–2454, December 2016. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
016-0227-4;   http://link.springer.com/content/pdf/10.1007/
s11538-016-0227-4.pdf. See erratum [3264].

Lawson:2017:SLM

[3199] Brodie A. J. Lawson and Graeme J. Pettet. Space-limited mitosis in the Glazier–Graner–Hogeweg model. Bulletin of Mathematical Biology, 79(1):1–20, January 2017. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-016-0204-y;   http://link.springer.com/content/
pdf/10.1007/s11538-016-0204-y.pdf.

Guo:2017:RSQ

[3200] Qiang-Hui Guo, Lisa Hui Sun, and Jian Wang. Regular simple queues of protein contact maps. Bulletin of Mathematical Biology, 79(1): 21–35, January 2017. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-016-0212-y;   http://link.springer.com/content/
pdf/10.1007/s11538-016-0212-y.pdf.

Bintz:2017:ASE

[3201] Jason Bintz, Suzanne Lenhart, and Cristina Lanzas. Antimicrobial stewardship and environmental decontamination for the control of Clostridium difficile transmission in healthcare settings. Bulletin of Mathematical Biology, 79(1):36–62, January 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-016-0224-7;
 http://link.springer.com/content/pdf/10.1007/s11538-016-0224-
7.pdf.

An:2017:OCA

[3202] G. An, B. G. Fitzpatrick, S. Christley, P. Federico, A. Kanarek, R. Miller Neilan, M. Oremland, R. Salinas, R. Laubenbacher, and S. Lenhart. Optimization and control of agent-based models in biology: a perspective. Bulletin of Mathematical Biology, 79(1):63–87, January 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-

016-0225-6;   http://link.springer.com/content/pdf/10.1007/
s11538-016-0225-6.pdf.

Numfor:2017:OCB

[3203] Eric Numfor, Frank M. Hilker, and Suzanne Lenhart. Optimal culling and biocontrol in a predator–prey model. Bulletin of Mathematical Biology, 79(1):88–116, January 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-016-0228-3;
 http://link.springer.com/content/pdf/10.1007/s11538-016-0228-
3.pdf.

Ross:2017:MMB

[3204] David S. Ross, Khamir Mehta, and Antonio Cabal. Mathematical model of bone remodeling captures the antiresorptive and anabolic actions of various therapies. Bulletin of Mathematical Biology, 79(1):117–142, January 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
016-0229-2;   http://link.springer.com/content/pdf/10.1007/
s11538-016-0229-2.pdf.

Fiorini:2017:ORS

[3205] Camilla Fiorini. Optimization of running strategies according to the physiological parameters for a two-runners model. Bulletin of Mathematical Biology, 79(1):143–162, January 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-016-0230-9;
 http://link.springer.com/content/pdf/10.1007/s11538-016-0230-
9.pdf.

Li:2017:PPH

[3206] Jing Li, Darla V. Lindberg, Rachel A. Smith, and Timothy C. Reluga. Provisioning of public health can be designed to anticipate public policy responses. Bulletin of Mathematical Biology, 79(1):163–190, January 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
016-0231-8;   http://link.springer.com/content/pdf/10.1007/
s11538-016-0231-8.pdf.

Seigal:2017:DAR

[3207] Anna Seigal, Portia Mira, Bernd Sturmfels, and Miriam Barlow. Does antibiotic resistance evolve in hospitals? Bulletin of Mathematical Biology, 79(1):191–208, January 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-016-0232-7;
 http://link.springer.com/content/pdf/10.1007/s11538-016-0232-
7.pdf.

Sfakianakis:2017:MAM

[3208] Nikolaos Sfakianakis, Niklas Kolbe, Nadja Hellmann, and Mária Lukáčová-Medvid’ová. A multiscale approach to the migration of cancer stem cells: Mathematical modelling and simulations. Bulletin of Mathematical Biology, 79(1):209–235, January 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
016-0233-6;   http://link.springer.com/content/pdf/10.1007/
s11538-016-0233-6.pdf.

Storey:2017:SMG

[3209] K. Storey, M. D. Ryser, K. Leder, and J. Foo. Spatial measures of genetic heterogeneity during carcinogenesis. Bulletin of Mathematical Biology, 79 (2):237–276, February 2017. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-016-0234-5;   http://link.springer.com/content/
pdf/10.1007/s11538-016-0234-5.pdf.

Choi:2017:DVK

[3210] Sun-Ho Choi and Yong-Jung Kim. A discrete velocity kinetic model with food metric: Chemotaxis traveling waves. Bulletin of Mathematical Biology, 79(2):277–302, February 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-016-0235-4;
 http://link.springer.com/content/pdf/10.1007/s11538-016-0235-
4.pdf.

Saad-Roy:2017:MMA

[3211] C. M. Saad-Roy, P. van den Driessche, and Abdul-Aziz Yakubu. A mathematical model of anthrax transmission in animal populations. Bulletin of Mathematical Biology, 79(2):303–324, February 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
016-0238-1;   http://link.springer.com/content/pdf/10.1007/
s11538-016-0238-1.pdf.

Gulbudak:2017:VBP

[3212] Hayriye Gulbudak, Vincent L. Cannataro, Necibe Tuncer, and Maia Martcheva. Vector-borne pathogen and host evolution in a structured immuno–epidemiological system. Bulletin of Mathematical Biology, 79 (2):325–355, February 2017. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-016-0239-0;   http://link.springer.com/content/
pdf/10.1007/s11538-016-0239-0.pdf.

Lewis:2017:YMF

[3213] Matthew Lewis and James Powell. Yeast for mathematicians: a ferment of discovery and model competition to describe data. Bulletin of Mathematical Biology, 79(2):356–382, February 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
016-0236-3;   http://link.springer.com/content/pdf/10.1007/
s11538-016-0236-3.pdf.

Maini:2017:OHM

[3214] Philip K. Maini, Hans G. Othmer, and Andreas Deutsch. Obituary: Hans Meinhardt (1938–2016). Bulletin of Mathematical Biology, 79(2): 383–388, February 2017. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-016-0243-4;   http://link.springer.com/content/
pdf/10.1007/s11538-016-0243-4.pdf.

Peng:2017:MMM

[3215] Lu Peng, Dumitru Trucu, Ping Lin, Alastair Thompson, and Mark A. J. Chaplain. A multiscale mathematical model of tumour invasive growth. Bulletin of Mathematical Biology, 79(3):389–429, March 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
016-0237-2;   http://link.springer.com/content/pdf/10.1007/
s11538-016-0237-2.pdf.

Ravigne:2017:MLF

[3216] Virginie Ravigné, Valérie Lemesle, Alicia Walter, Ludovic Mailleret, and Frédéric M. Hamelin. Mate limitation in fungal plant parasites can lead to cyclic epidemics in perennial host populations. Bulletin of Mathematical Biology, 79(3):430–447, March 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-016-0240-7;

 http://link.springer.com/content/pdf/10.1007/s11538-016-0240-
7.pdf.

Lin:2017:IPI

[3217] Lin Lin and Hans G. Othmer. Improving parameter inference from FRAP data: an analysis motivated by pattern formation in the Drosophila wing disc. Bulletin of Mathematical Biology, 79(3):448–497, March 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
016-0241-6;   http://link.springer.com/content/pdf/10.1007/
s11538-016-0241-6.pdf.

Humphries:2017:MCC

[3218] D. L. Humphries, J. A. Grogan, and E. A. Gaffney. Mechanical cell– cell communication in fibrous networks: The importance of network geometry. Bulletin of Mathematical Biology, 79(3):498–524, March 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
016-0242-5;   http://link.springer.com/content/pdf/10.1007/
s11538-016-0242-5.pdf.

Kollar:2017:ETW

[3219] Richard Kollár and Sebastian Novak. Existence of traveling waves for the generalized F–KPP equation. Bulletin of Mathematical Biology, 79 (3):525–559, March 2017. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-016-0244-3;   http://link.springer.com/content/
pdf/10.1007/s11538-016-0244-3.pdf.

Zhang:2017:LWN

[3220] Li Zhang, Fan Zhang, and Shigui Ruan. Linear and weakly nonlinear stability analyses of Turing patterns for diffusive predator–prey systems in freshwater Marsh landscapes. Bulletin of Mathematical Biology, 79 (3):560–593, March 2017. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-017-0245-x;   http://link.springer.com/content/
pdf/10.1007/s11538-017-0245-x.pdf.

Machineni:2017:INA

[3221] Lakshmi Machineni, Anil Rajapantul, Vandana Nandamuri, and Parag D. Pawar. Influence of nutrient availability and quorum sensing on the formation of metabolically inactive microcolonies within structurally heterogeneous bacterial biofilms: an individual-based 3D cellular automata model. Bulletin of Mathematical Biology, 79(3):594–618, March 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0246-9;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0246-9.pdf.

Sumner:2017:DRG

[3222] Jeremy G. Sumner. Dimensional reduction for the general Markov model on phylogenetic trees. Bulletin of Mathematical Biology, 79(3): 619–634, March 2017. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-017-0249-6;   http://link.springer.com/content/
pdf/10.1007/s11538-017-0249-6.pdf.

Ma:2017:MML

[3223] Timmy Ma and Natalia L. Komarova. Mathematical modeling of learning from an inconsistent source: a nonlinear approach. Bulletin of Mathematical Biology, 79(3):635–661, March 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-017-0250-0;
 http://link.springer.com/content/pdf/10.1007/s11538-017-0250-
0.pdf.

Karst:2017:MMN

[3224] Nathaniel J. Karst, John B. Geddes, and Russell T. Carr. Model microvascular networks can have many equilibria. Bulletin of Mathematical Biology, 79(3):662–681, March 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-017-0251-z;   http://link.springer.com/
content/pdf/10.1007/s11538-017-0251-z.pdf.

Zheng:2017:TUD

[3225] Qi Zheng. Toward a unique definition of the mutation rate. Bulletin of Mathematical Biology, 79(4):683–692, April 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-017-0247-8;
 http://link.springer.com/content/pdf/10.1007/s11538-017-0247-
8.pdf.

Wertheim:2017:MML

[3226] Kenneth Y. Wertheim and Tiina Roose. A mathematical model of lymphangiogenesis in a zebrafish embryo. Bulletin of Mathematical Biology, 79(4):693–737, April 2017. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-017-0248-7;   http://link.springer.com/content/
pdf/10.1007/s11538-017-0248-7.pdf.

Malunguza:2017:API

[3227] Noble J. Malunguza, Senelani D. Hove-Musekwa, and Zindoga Mukandavire. Assessing the potential impact of hormonal-based contraceptives on HIV transmission dynamics among heterosexuals. Bulletin of Mathematical Biology, 79(4):738–771, April 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0252-y;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0252-y.pdf.

Holland:2017:HHB

[3228] E. Penelope Holland, Julie Mugford, Rachelle N. Binny, and Alex James. How herbivore browsing strategy affects whole-plant photosynthetic capacity. Bulletin of Mathematical Biology, 79(4):772–787, April 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0253-x;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0253-x.pdf.

Dhillon:2017:BAR

[3229] Daljit Singh J. Dhillon, Michel C. Milinkovitch, and Matthias Zwicker. Bifurcation analysis of reaction diffusion systems on arbitrary surfaces. Bulletin of Mathematical Biology, 79(4):788–827, April 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0255-8;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0255-8.pdf.

Marculis:2017:NGP

[3230] Nathan G. Marculis, Roger Lui, and Mark A. Lewis. Neutral genetic patterns for expanding populations with nonoverlapping generations. Bulletin of Mathematical Biology, 79(4):828–852, April 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0256-7;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0256-7.pdf.

Martins:2017:BES

[3231] José Martins and Alberto Pinto. Bistability of evolutionary stable vaccination strategies in the reinfection SIRI model. Bulletin of Mathematical Biology, 79(4):853–883, April 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-017-0257-6;
 http://link.springer.com/content/pdf/10.1007/s11538-017-0257-
6.pdf.

Zhao:2017:TDN

[3232] Jia Zhao and Qi Wang. Three-dimensional numerical simulations of biofilm dynamics with quorum sensing in a flow cell. Bulletin of Mathematical Biology, 79(4):884–919, April 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-017-0259-4;
 http://link.springer.com/content/pdf/10.1007/s11538-017-0259-
4.pdf.

Baste:2017:EFA

[3233] Julien Baste, Christophe Paul, Ignasi Sau, and Celine Scornavacca. Efficient FPT algorithms for (strict) compatibility of unrooted phylogenetic trees. Bulletin of Mathematical Biology, 79(4):920–938, April 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0260-y;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0260-y.pdf.

Collis:2017:BCV

[3234] Joe Collis, Anthony J. Connor, Marcin Paczkowski, Pavitra Kannan, Joe Pitt-Francis, Helen M. Byrne, and Matthew E. Hubbard. Bayesian calibration, validation and uncertainty quantification for predictive modelling of tumour growth: a tutorial. Bulletin of Mathematical Biology, 79(4):939–974, April 2017. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-017-0258-5;   http://link.springer.com/content/
pdf/10.1007/s11538-017-0258-5.pdf.

Forcey:2017:SFB

[3235] Stefan Forcey, Logan Keefe, and William Sands. Split-facets for balanced minimal evolution polytopes and the permutoassociahedron. Bulletin of Mathematical Biology, 79(5):975–994, May 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0264-7;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0264-7.pdf.

Korytowski:2017:PSK

[3236] Daniel A. Korytowski and Hal Smith. Permanence and stability of a kill the winner model in marine ecology. Bulletin of Mathematical Biology, 79(5):995–1004, May 2017. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-017-0265-6;   http://link.springer.com/content/
pdf/10.1007/s11538-017-0265-6.pdf.

Dutta:2017:GMM

[3237] Annwesha Dutta and Debashish Chowdhury. A generalized Michaelis– Menten equation in protein synthesis: Effects of mis-charged cognate tRNA and mis-reading of codon. Bulletin of Mathematical Biology, 79 (5):1005–1027, May 2017. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-017-0266-5;   http://link.springer.com/content/
pdf/10.1007/s11538-017-0266-5.pdf.

Jin:2017:LPC

[3238] Wang Jin, Esha T. Shah, Catherine J. Penington, Scott W. McCue, Philip K. Maini, and Matthew J. Simpson. Logistic proliferation of cells in scratch assays is delayed. Bulletin of Mathematical Biology, 79 (5):1028–1050, May 2017. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-017-0267-4;   http://link.springer.com/content/
pdf/10.1007/s11538-017-0267-4.pdf.

Golubitsky:2017:DST

[3239] Martin Golubitsky, Wenrui Hao, King-Yeung Lam, and Yuan Lou. Dimorphism by singularity theory in a model for river ecology. Bulletin of Mathematical Biology, 79(5):1051–1069, May 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-017-0268-3;
 http://link.springer.com/content/pdf/10.1007/s11538-017-0268-
3.pdf.

Rossa:2017:ESU

[3240] Fabio Della Rossa, Fabio Dercole, and Cristina Vicini. Extreme selection unifies evolutionary game dynamics in finite and infinite populations. Bulletin of Mathematical Biology, 79(5):1070–1099, May 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0269-2;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0269-2.pdf.

Ghosh:2017:MSC

[3241] Indrajit Ghosh, Tridip Sardar, and Joydev Chattopadhyay. A mathematical study to control visceral leishmaniasis: an application to South Sudan. Bulletin of Mathematical Biology, 79(5):1100–1134, May 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0274-5;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0274-5.pdf. See comments [3352].

vanIersel:2017:BFB

[3242] Leo van Iersel, Vincent Moulton, Eveline de Swart, and Taoyang Wu. Binets: Fundamental building blocks for phylogenetic networks. Bulletin of Mathematical Biology, 79(5):1135–1154, May 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0275-4;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0275-4.pdf.

Wang:2017:MTM

[3243] Xiunan Wang and Xiao-Qiang Zhao. A malaria transmission model with temperature-dependent incubation period. Bulletin of Mathematical Biology, 79(5):1155–1182, May 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-017-0276-3;
 http://link.springer.com/content/pdf/10.1007/s11538-017-0276-
3.pdf.

Ibrahim:2017:MFK

[3244] Bashar Ibrahim. A mathematical framework for kinetochore-driven activation feedback in the mitotic checkpoint. Bulletin of Mathematical Biology, 79(5):1183–1200, May 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-017-0278-1;   http://link.springer.com/
content/pdf/10.1007/s11538-017-0278-1.pdf.

Hogan:2017:MSI

[3245] Alexandra B. Hogan and Mary R. Myerscough. A model for the spread of an invasive weed, tradescantia fluminensis. Bulletin of Mathematical Biology, 79(6):1201–1217, June 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-017-0280-7;
 http://link.springer.com/content/pdf/10.1007/s11538-017-0280-
7.pdf.

Ratti:2017:MMF

[3246] Vardayani Ratti, Peter G. Kevan, and Hermann J. Eberl. A mathematical model of forager loss in honeybee colonies infested with varroa destructor and the acute bee paralysis virus. Bulletin of Mathematical Biology, 79(6):1218–1253, June 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-017-0281-6;
 http://link.springer.com/content/pdf/10.1007/s11538-017-0281-
6.pdf.

Banks:2017:DMP

[3247] H. T. Banks, Kidist Bekele-Maxwell, R. A. Everett, Lyric Stephenson, Sijing Shao, and Jon Morgenstern. Dynamic modeling of problem drinkers undergoing behavioral treatment. Bulletin of Mathematical Biology, 79 (6):1254–1273, June 2017. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-017-0282-5;   http://link.springer.com/content/
pdf/10.1007/s11538-017-0282-5.pdf.

Vasilyeva:2017:CMS

[3248] Olga Vasilyeva. Competition of multiple species in advective environments. Bulletin of Mathematical Biology, 79(6):1274–1294, June 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0285-2;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0285-2.pdf.

Yildirim:2017:COF

[3249] Vehpi Yildirim and Richard Bertram. Calcium oscillation frequencysensitive gene regulation and homeostatic compensation in pancreatic β-cells. Bulletin of Mathematical Biology, 79(6):1295–1324, June 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-

017-0286-1;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0286-1.pdf.

Wang:2017:MFE

[3250] Xiaoying Wang and Xingfu Zou. Modeling the fear effect in predator– prey interactions with adaptive avoidance of predators. Bulletin of Mathematical Biology, 79(6):1325–1359, June 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-017-0287-0;
 http://link.springer.com/content/pdf/10.1007/s11538-017-0287-
0.pdf.

Alfiniyah:2017:PGQ

[3251] Cicik Alfiniyah, Martin A. Bees, and A. Jamie Wood. Pulse generation in the quorum machinery of Pseudomonas aeruginosa. Bulletin of Mathematical Biology, 79(6):1360–1389, June 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-017-0288-z;
 http://link.springer.com/content/pdf/10.1007/s11538-017-0288-
z.pdf.

Deng:2017:AMH

[3252] Bo Deng. Alternative models to Hodgkin–Huxley equations. Bulletin of Mathematical Biology, 79(6):1390–1411, June 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-017-0289-y;
 http://link.springer.com/content/pdf/10.1007/s11538-017-0289-
y.pdf.

Adi-Kusumo:2017:DSS

[3253] Fajar Adi-Kusumo. The dynamics of a SEIR–SIRC antigenic drift influenza model. Bulletin of Mathematical Biology, 79(6):1412–1425, June 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0290-5;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0290-5.pdf.

Wilkie:2017:MDI

[3254] Kathleen P. Wilkie and Philip Hahnfeldt. Modeling the dichotomy of the immune response to cancer: Cytotoxic effects and tumor-promoting inflammation. Bulletin of Mathematical Biology, 79(6):1426–1448, June 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-

017-0291-4;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0291-4.pdf.

Snowden:2017:MMR

[3255] Thomas J. Snowden, Piet H. van der Graaf, and Marcus J. Tindall. Methods of model reduction for large-scale biological systems: a survey of current methods and trends. Bulletin of Mathematical Biology, 79 (7):1449–1486, July 2017. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-017-0277-2;   http://link.springer.com/content/
pdf/10.1007/s11538-017-0277-2.pdf.

Bangsgaard:2017:IIS

[3256] Elisabeth O. Bangsgaard, Poul G. Hjorth, Mette S. Olufsen, Jesper Mehlsen, and Johnny T. Ottesen. Integrated Inflammatory Stress (ITIS) model. Bulletin of Mathematical Biology, 79(7):1487–1509, July 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0293-2;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0293-2.pdf.

Cortez:2017:EPE

[3257] Michael H. Cortez and Swati Patel. The effects of predator evolution and genetic variation on predator–prey population-level dynamics. Bulletin of Mathematical Biology, 79(7):1510–1538, July 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0297-y;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0297-y.pdf.

Ingalls:2017:PSA

[3258] Brian Ingalls, Maya Mincheva, and Marc R. Roussel. Parametric sensitivity analysis of oscillatory delay systems with an application to gene regulation. Bulletin of Mathematical Biology, 79(7):1539–1563, July 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0298-x;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0298-x.pdf.

Miller:2017:MMC

[3259] Anna K. Miller, Karl Munger, and Frederick R. Adler. A mathematical model of cell cycle dysregulation due to human papillomavirus infection. Bulletin of Mathematical Biology, 79(7):1564–1585, July 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0299-9;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0299-9.pdf.

Betti:2017:RNA

[3260] M. I. Betti, L. M. Wahl, and M. Zamir. Reproduction number and asymptotic stability for the dynamics of a honey bee colony with continuous age structure. Bulletin of Mathematical Biology, 79(7):1586–1611, July 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0300-7;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0300-7.pdf.

Erdem:2017:MAS

[3261] Mustafa Erdem, Muntaser Safan, and Carlos Castillo-Chavez. Mathematical analysis of an SIQR influenza model with imperfect quarantine. Bulletin of Mathematical Biology, 79(7):1612–1636, July 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0301-6;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0301-6.pdf.

Delboni:2017:MMIa

[3262] Roberta Regina Delboni and Hyun Mo Yang. Mathematical model of interaction between bacteriocin-producing lactic acid bacteria and Listeria. Part 1: Steady states and thresholds. Bulletin of Mathematical Biology, 79(7):1637–1661, July 2017. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-017-0302-5;   http://link.springer.com/content/
pdf/10.1007/s11538-017-0302-5.pdf.

deFreitas:2017:IGC

[3263] Michael Marcondes de Freitas, Carsten Wiuf, and Elisenda Feliu. Intermediates and generic convergence to equilibria. Bulletin of Mathematical Biology, 79(7):1662–1686, July 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-017-0303-4;
 http://link.springer.com/content/pdf/10.1007/s11538-017-0303-
4.pdf.

Britton:2017:ENE

[3264] Tom Britton, David Juher, and Joan Saldaña. Erratum to: a network epidemic model with preventive rewiring: Comparative analysis of the initial phase. Bulletin of Mathematical Biology, 79(7):1687–1689, July 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0295-0;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0295-0.pdf. See [3198].

Catozzi:2017:DDS

[3265] S. Catozzi and J.-A. Sepulchre. A discrete dynamical system approach to pathway activation profiles of signaling cascades. Bulletin of Mathematical Biology, 79(8):1691–1735, August 2017. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-017-0296-z;
 http://link.springer.com/content/pdf/10.1007/s11538-017-0296-
z.pdf.

Timalsina:2017:MCM

[3266] Asim Timalsina, Jianjun Paul Tian, and Jin Wang. Mathematical and computational modeling for tumor virotherapy with mediated immunity. Bulletin of Mathematical Biology, 79(8):1736–1758, August 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0304-3;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0304-3.pdf.

Braverman:2017:SSP

[3267] Elena Braverman and Daniel Franco. Stabilization of structured populations via vector target-oriented control. Bulletin of Mathematical Biology, 79(8):1759–1777, August 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-017-0305-2;   http://link.springer.com/
content/pdf/10.1007/s11538-017-0305-2.pdf.

Jenkins:2017:BAB

[3268] Andy Jenkins and Matthew Macauley. Bistability and asynchrony in a Boolean model of the l-arabinose Operon in Escherichia coli. Bulletin of Mathematical Biology, 79(8):1778–1795, August 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0306-1;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0306-1.pdf.

Fimmel:2017:SCF

[3269] Elena Fimmel, Christian J. Michel, and Lutz Strüngmann. Strong comma-free codes in genetic information. Bulletin of Mathematical Biology, 79(8):1796–1819, August 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-017-0307-0;   http://link.springer.com/
content/pdf/10.1007/s11538-017-0307-0.pdf.

Pulwicki:2017:DRG

[3270] Julia Pulwicki and David Hobill. The dynamics of root growth: a geometric model. Bulletin of Mathematical Biology, 79(8):1820–1845, August 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0308-z;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0308-z.pdf.

Mitchell:2017:CMC

[3271] Christopher Mitchell and Christopher Kribs. A comparison of methods for calculating the basic reproductive number for periodic epidemic systems. Bulletin of Mathematical Biology, 79(8):1846–1869, August 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0309-y;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0309-y.pdf.

Hampson:2017:SCM

[3272] Karen M. Hampson, Matthew P. Cufflin, and Edward A. H. Mallen. Sensitivity of chaos measures in detecting stress in the focusing control mechanism of the short-sighted eye. Bulletin of Mathematical Biology, 79(8): 1870–1887, August 2017. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-017-0310-5;   http://link.springer.com/content/
pdf/10.1007/s11538-017-0310-5.pdf.

Browning:2017:BCA

[3273] Alexander P. Browning, Scott W. McCue, and Matthew J. Simpson. A Bayesian computational approach to explore the optimal duration of a cell proliferation assay. Bulletin of Mathematical Biology, 79(8): 1888–1906, August 2017. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-017-0311-4;   http://link.springer.com/content/
pdf/10.1007/s11538-017-0311-4.pdf.

Loo:2017:FMM

[3274] Sara L. Loo, Matthew H. Chan, Kristen Hawkes, and Peter S. Kim. Further mathematical modelling of mating sex ratios & male strategies with special relevance to human life history. Bulletin of Mathematical Biology, 79(8):1907–1922, August 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-017-0313-2;
 http://link.springer.com/content/pdf/10.1007/s11538-017-0313-
2.pdf.

Zmurchok:2017:AQS

[3275] Cole Zmurchok, Tim Small, Michael J. Ward, and Leah Edelstein-Keshet. Application of quasi–steady-state methods to nonlinear models of intracellular transport by molecular motors. Bulletin of Mathematical Biology, 79(9):1923–1978, September 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-017-0314-1;
 http://link.springer.com/content/pdf/10.1007/s11538-017-0314-

1.pdf. See correction [3337].

Lee:2017:MMM

[3276] Junehyuk Lee, Frederick R. Adler, and Peter S. Kim. A mathematical model for the macrophage response to respiratory viral infection in normal and asthmatic conditions. Bulletin of Mathematical Biology, 79(9):1979–1998, September 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-017-0315-0;
 http://link.springer.com/content/pdf/10.1007/s11538-017-0315-
0.pdf.

Maliyoni:2017:STB

[3277] Milliward Maliyoni, Faraimunashe Chirove, Holly D. Gaff, and Keshlan S. Govinder. A stochastic tick-borne disease model: Exploring the probability of pathogen persistence. Bulletin of Mathematical Biology, 79(9):1999–2021, September 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-017-0317-y;   http://link.springer.com/
content/pdf/10.1007/s11538-017-0317-y.pdf.

Gambette:2017:UPN

[3278] P. Gambette, K. T. Huber, and G. E. Scholz. Uprooted phylogenetic networks. Bulletin of Mathematical Biology, 79(9):2022–2048, September 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0318-x;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0318-x.pdf.

Magal:2017:MTP

[3279] P. Magal, A. Noussair, J. Pasquier, P. Zongo, and F. Le Foll. A model for transfer of P-glycoproteins in MCF-7 breast cancer cell line with multiple transfer rules. Bulletin of Mathematical Biology, 79(9):2049– 2067, September 2017. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-017-0319-9;   http://link.springer.com/content/
pdf/10.1007/s11538-017-0319-9.pdf.

Gaivao:2017:TFC

[3280] Maria Gaivão, Francisco Dionisio, and Erida Gjini. Transmission fitness in co-colonization and the persistence of bacterial pathogens. Bulletin of Mathematical Biology, 79(9):2068–2087, September 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0320-3;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0320-3.pdf.

Naasell:2017:AMC

[3281] Ingemar Nåsell. An alternative to moment closure. Bulletin of Mathematical Biology, 79(9):2088–2108, September 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-017-0321-2;
 http://link.springer.com/content/pdf/10.1007/s11538-017-0321-
2.pdf.

Yang:2017:IAP

[3282] Chayu Yang, Xueying Wang, Daozhou Gao, and Jin Wang. Impact of awareness programs on cholera dynamics: Two modeling approaches. Bulletin of Mathematical Biology, 79(9):2109–2131, September 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0322-1;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0322-1.pdf.

Chan:2017:MET

[3283] Matthew H. Chan, Kristen Hawkes, and Peter S. Kim. Modelling the evolution of traits in a two-sex population, with an application to grandmothering. Bulletin of Mathematical Biology, 79(9):2132–2148, September 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0323-0;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0323-0.pdf.

Wang:2017:EIU

[3284] Xiaojing Wang, Yangyang Shi, Zhilan Feng, and Jingan Cui. Evaluations of interventions using mathematical models with exponential and non-exponential distributions for disease stages: The case of Ebola. Bulletin of Mathematical Biology, 79(9):2149–2173, September 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0324-z;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0324-z.pdf.

Hilker:2017:DSP

[3285] Frank M. Hilker, Marta Paliga, and Ezio Venturino. Diseased social predators. Bulletin of Mathematical Biology, 79(10):2175–2196, October 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0325-y;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0325-y.pdf. See correction [3428].

Minors:2017:ISC

[3286] K. Minors and J. H. P. Dawes. Invasions slow down or collapse in the presence of reactive boundaries. Bulletin of Mathematical Biology, 79(10): 2197–2214, October 2017. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-017-0326-x;   http://link.springer.com/content/
pdf/10.1007/s11538-017-0326-x.pdf.

Evans:2017:TEM

[3287] Ryan M. Evans and David A. Edwards. Transport effects on multiple-component reactions in optical biosensors. Bulletin of Mathematical Biology, 79(10):2215–2241, October 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-017-0327-9;
 http://link.springer.com/content/pdf/10.1007/s11538-017-0327-
9.pdf.

McLure:2017:HAC

[3288] Angus McLure, Archie C. A. Clements, Martyn Kirk, and Kathryn Glass. Healthcare-associated clostridium difficile infections are sustained by disease from the community. Bulletin of Mathematical Biology, 79(10): 2242–2257, October 2017. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-017-0328-8;   http://link.springer.com/content/
pdf/10.1007/s11538-017-0328-8.pdf.

Jarrett:2017:CTM

[3289] Angela M. Jarrett, N. G. Cogan, and M. Y. Hussaini. Combining two methods of global sensitivity analysis to investigate MRSA nasal carriage model. Bulletin of Mathematical Biology, 79(10):2258–2272, October 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0329-7;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0329-7.pdf.

Delboni:2017:MMIb

[3290] Roberta Regina Delboni and Hyun Mo Yang. Mathematical model of interaction between bacteriocin-producing lactic acid bacteria and Listeria. Part 2: Bifurcations and applications. Bulletin of Mathematical Biology, 79(10):2273–2301, October 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-017-0330-1;
 http://link.springer.com/content/pdf/10.1007/s11538-017-0330-
1.pdf.

Kurowski:2017:TSM

[3291] Lawrence Kurowski, Andrew L. Krause, Hanako Mizuguchi, Peter Grindrod, and Robert A. Van Gorder. Two-species migration and clustering in two-dimensional domains. Bulletin of Mathematical Biology, 79(10):2302–2333, October 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-017-0331-0;   http://link.springer.com/
content/pdf/10.1007/s11538-017-0331-0.pdf.

Didier:2017:TDA

[3292] Gilles Didier. Time-dependent-asymmetric-linear-parsimonious ancestral state reconstruction. Bulletin of Mathematical Biology, 79(10): 2334–2355, October 2017. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-017-0332-z;   http://link.springer.com/content/
pdf/10.1007/s11538-017-0332-z.pdf.

Leroy-Leretre:2017:TCL

[3293] Mathieu Leroy-Lerêtre, Giacomo Dimarco, Martine Cazales, Marie-Laure Boizeau, Bernard Ducommun, Valérie Lobjois, and Pierre Degond. Are tumor cell lineages solely shaped by mechanical forces? Bulletin of Mathematical Biology, 79(10):2356–2393, October 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0333-y;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0333-y.pdf.

Becker:2017:MTD

[3294] Matthew Becker and Doron Levy. Modeling the transfer of drug resistance in solid tumors. Bulletin of Mathematical Biology, 79(10): 2394–2412, October 2017. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-017-0334-x;   http://link.springer.com/content/
pdf/10.1007/s11538-017-0334-x.pdf.

Mesterton-Gibbons:2017:VCE

[3295] Mike Mesterton-Gibbons, Yao Dai, Marlène Goubault, and Ian C. W. Hardy. Volatile chemical emission as a weapon of rearguard action: a game-theoretic model of contest behavior. Bulletin of Mathematical Biology, 79(11):2413–2449, November 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0335-9;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0335-9.pdf.

Shafaghati:2017:SBA

[3296] Leila Shafaghati, Zahra Razaghi-Moghadam, and Javad Mohammadnejad. A systems biology approach to understanding alcoholic liver disease molecular mechanism: The development of static and dynamic models. Bulletin of Mathematical Biology, 79(11):2450–2473, November 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0336-8;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0336-8.pdf.

Fong:2017:MMI

[3297] D. Fong and L. J. Cummings. Mathematical modeling of ischemia– reperfusion injury and postconditioning therapy. Bulletin of Mathematical Biology, 79(11):2474–2511, November 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/article/10.1007/s11538-017-0337-7;   http://link.
springer.com/content/pdf/10.1007/s11538-017-0337-7.pdf.

Edwards:2017:CVR

[3298] Aurélie Edwards and Anita T. Layton. Cell volume regulation in the proximal tubule of rat kidney. Bulletin of Mathematical Biology, 79(11): 2512–2533, November 2017. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-017-0338-6;   http://link.springer.com/content/
pdf/10.1007/s11538-017-0338-6.pdf.

Reed:2017:AHM

[3299] Michael Reed, Janet Best, Martin Golubitsky, Ian Stewart, and H. Frederik Nijhout. Analysis of homeostatic mechanisms in biochemical networks. Bulletin of Mathematical Biology, 79(11):2534–2557, November 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0340-z;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0340-z.pdf.

Dolejs:2017:MHM

[3300] Josef Dolejs. Modeling human mortality from all diseases in the five most populated countries of the European Union. Bulletin of Mathematical Biology, 79(11):2558–2598, November 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0341-y;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0341-y.pdf.

Fan:2017:PMQ

[3301] Gaoyang Fan and Paul C. Bressloff. Population model of quorum sensing with multiple parallel pathways. Bulletin of Mathematical Biology, 79(11):2599–2626, November 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-017-0343-9;   http://link.springer.com/
content/pdf/10.1007/s11538-017-0343-9.pdf.

Rutter:2017:CSP

[3302] Erica M. Rutter, H. T. Banks, Gerald A. LeBlanc, and Kevin B. Flores. Continuous structured population models for Daphnia magna. Bulletin of Mathematical Biology, 79(11):2627–2648, November 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0344-8;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0344-8.pdf.

Croicu:2017:STA

[3303] Ana-Maria Croicu, Angela M. Jarrett, N. G. Cogan, and M. Yousuff Hussaini. Short-term antiretroviral treatment recommendations based on sensitivity analysis of a mathematical model for HIV infection of CD4^{+}T cells. Bulletin of Mathematical Biology, 79(11):2649–2671, November 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0345-7;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0345-7.pdf.

Meinecke:2017:MMD

[3304] Lina Meinecke. Multiscale modeling of diffusion in a crowded environment. Bulletin of Mathematical Biology, 79(11):2672–2695, November 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0346-6;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0346-6.pdf.

Chen:2017:FPN

[3305] Duan Chen. Fractional Poisson–Nernst–Planck model for ion channels I: Basic formulations and algorithms. Bulletin of Mathematical Biology, 79(11):2696–2726, November 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-017-0349-3;   http://link.springer.com/
content/pdf/10.1007/s11538-017-0349-3.pdf.

Barril:2017:RNG

[3306] Carles Barril, Àngel Calsina, and Jordi Ripoll. On the reproduction number of a gut microbiota model. Bulletin of Mathematical Biology, 79(11): 2727–2746, November 2017. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-017-0352-8;   http://link.springer.com/content/
pdf/10.1007/s11538-017-0352-8.pdf.

Ledder:2017:SDS

[3307] Glenn Ledder. Scaling for dynamical systems in biology. Bulletin of Mathematical Biology, 79(11):2747–2772, November 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0339-5;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0339-5.pdf.

Severino:2017:EHS

[3308] Gerardo Severino, Francesco Giannino, Fabrizio Cartení, Stefano Mazzoleni, and Daniel M. Tartakovsky. Effects of hydraulic soil properties on vegetation pattern formation in sloping landscapes. Bulletin of Mathematical Biology, 79(12):2773–2784, December 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0348-4;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0348-4.pdf.

Keyes:2017:ESM

[3309] Samuel David Keyes, Konstantinos C. Zygalakis, and Tiina Roose. An explicit structural model of root hair and soil interactions parameterised by synchrotron X-ray computed tomography. Bulletin of Mathematical Biology, 79(12):2785–2813, December 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0350-x;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0350-x.pdf.

Stapf:2017:DTF

[3310] M. R. Stapf, R. J. Braun, and P. E. King-Smith. Duplex tear film evaporation analysis. Bulletin of Mathematical Biology, 79(12):2814– 2846, December 2017. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-017-0351-9;   http://link.springer.com/content/
pdf/10.1007/s11538-017-0351-9.pdf.

Bernal:2017:MSS

[3311] Esteban Vargas Bernal and Camilo Sanabria Malagon. Modeling sexual selection in Túngara frog and rationality of mate choice. Bulletin of Mathematical Biology, 79(12):2847–2864, December 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-

017-0353-7;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0353-7.pdf.

Herbst:2017:ASR

[3312] Lina Herbst and Mareike Fischer. Ancestral sequence reconstruction with maximum parsimony. Bulletin of Mathematical Biology, 79(12): 2865–2886, December 2017. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-017-0354-6;   http://link.springer.com/content/
pdf/10.1007/s11538-017-0354-6.pdf.

Milliken:2017:PEI

[3313] Evan Milliken. The probability of extinction of infectious salmon anemia virus in one and two patches. Bulletin of Mathematical Biology, 79(12): 2887–2904, December 2017. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-017-0355-5;   http://link.springer.com/content/
pdf/10.1007/s11538-017-0355-5.pdf.

Yates:2017:MSR

[3314] Christian A. Yates, Matthew J. Ford, and Richard L. Mort. A multi-stage representation of cell proliferation as a Markov process. Bulletin of Mathematical Biology, 79(12):2905–2928, December 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0356-4;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0356-4.pdf.

Mazoyer:2017:TIM

[3315] Adrien Mazoyer. Time inhomogeneous mutation models with birth date dependence. Bulletin of Mathematical Biology, 79(12):2929– 2953, December 2017. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-017-0357-3;   http://link.springer.com/content/
pdf/10.1007/s11538-017-0357-3.pdf.

Jin:2017:MRH

[3316] Yu Jin, Frithjof Lutscher, and Yuan Pei. Meandering rivers: How important is lateral variability for species persistence? Bulletin of Mathematical Biology, 79(12):2954–2985, December 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-

017-0358-2;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0358-2.pdf.

Cho:2017:MDH

[3317] Heyrim Cho and Doron Levy. Modeling the dynamics of heterogeneity of solid tumors in response to chemotherapy. Bulletin of Mathematical Biology, 79(12):2986–3012, December 2017. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0359-1;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0359-1.pdf.

Xia:2018:MPF

[3318] Kelin Xia, Zhiming Li, and Lin Mu. Multiscale persistent functions for biomolecular structure characterization. Bulletin of Mathematical Biology, 80(1):1–31, January 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-017-0362-6;   http://link.springer.com/
content/pdf/10.1007/s11538-017-0362-6.pdf.

Flondor:2018:QAO

[3319] Paul Flondor, Mircea Olteanu, and Radu Ştefan. Qualitative analysis of an ODE model of a class of enzymatic reactions. Bulletin of Mathematical Biology, 80(1):32–45, January 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-017-0360-8;
 http://link.springer.com/content/pdf/10.1007/s11538-017-0360-
8.pdf.

Moore:2018:WCA

[3320] James Moore, Hasan Ahmed, Jonathan Jia, Rama Akondy, Rafi Ahmed, and Rustom Antia. What controls the acute viral infection following yellow fever vaccination? Bulletin of Mathematical Biology, 80(1): 46–63, January 2018. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-017-0365-3;   http://link.springer.com/content/
pdf/10.1007/s11538-017-0365-3.pdf.

Allman:2018:SPS

[3321] Elizabeth S. Allman, James H. Degnan, and John A. Rhodes. Split probabilities and species tree inference under the multispecies coalescent model. Bulletin of Mathematical Biology, 80(1):64–103, January 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0363-5;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0363-5.pdf.

Elias:2018:TER

[3322] Ján Eliaš. Trend to equilibrium for a reaction–diffusion system modelling reversible enzyme reaction. Bulletin of Mathematical Biology, 80 (1):104–129, January 2018. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-017-0364-4;   http://link.springer.com/content/
pdf/10.1007/s11538-017-0364-4.pdf.

Bilinsky:2018:SPT

[3323] L. M. Bilinsky and S. M. Baer. Slow passage through a Hopf bifurcation in excitable nerve cables: Spatial delays and spatial memory effects. Bulletin of Mathematical Biology, 80(1):130–150, January 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0366-2;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0366-2.pdf.

Kareva:2018:ETW

[3324] Irina Kareva and Georgy Karev. From experiment to theory: What can we learn from growth curves? Bulletin of Mathematical Biology, 80 (1):151–174, January 2018. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-017-0347-5;   http://link.springer.com/content/
pdf/10.1007/s11538-017-0347-5.pdf.

Thon:2018:QME

[3325] Moritz P. Thon, Hugh Z. Ford, Michael W. Gee, and Mary R. Myerscough. A quantitative model of early atherosclerotic plaques parameterized using in vitro experiments. Bulletin of Mathematical Biology, 80 (1):175–214, January 2018. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-017-0367-1;   http://link.springer.com/content/
pdf/10.1007/s11538-017-0367-1.pdf.

Takagi:2018:DHS

[3326] Daisuke Takagi and Daniel K. Hartline. Directional hydrodynamic sensing by free-swimming organisms. Bulletin of Mathematical Biology, 80 (1):215–227, January 2018. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-017-0368-0;   http://link.springer.com/content/
pdf/10.1007/s11538-017-0368-0.pdf.

Aydogmus:2018:PTL

[3327] Ozgur Aydogmus. Phase transitions in a logistic metapopulation model with nonlocal interactions. Bulletin of Mathematical Biology, 80(1): 228–253, January 2018. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-017-0373-3;   http://link.springer.com/content/
pdf/10.1007/s11538-017-0373-3.pdf.

Vera-Siguenza:2018:MMS

[3328] Elías Vera-Sigüenza, Marcelo A. Catalán, Gaspar Peña-Münzenmayer, James E. Melvin, and James Sneyd. A mathematical model supports a key role for Ae4 (Slc4a9) in salivary gland secretion. Bulletin of Mathematical Biology, 80(2):255–282, February 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0370-6;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0370-6.pdf.

Forouzannia:2018:MME

[3329] Farinaz Forouzannia, Heiko Enderling, and Mohammad Kohandel. Mathematical modeling of the effects of tumor heterogeneity on the efficiency of radiation treatment schedule. Bulletin of Mathematical Biology, 80(2): 283–293, February 2018. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-017-0371-5;   http://link.springer.com/content/
pdf/10.1007/s11538-017-0371-5.pdf.

Belgacem:2018:RSA

[3330] Ismail Belgacem, Stefano Casagranda, Edith Grac, Delphine Ropers, and Jean-Luc Gouzé. Reduction and stability analysis of a transcription–translation model of RNA polymerase. Bulletin of Mathematical Biology, 80(2):294–318, February 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-017-0372-4;
 http://link.springer.com/content/pdf/10.1007/s11538-017-0372-
4.pdf.

Jamilov:2018:VCS

[3331] U. U. Jamilov, A. Yu. Khamraev, and M. Ladra. On a Volterra cubic stochastic operator. Bulletin of Mathematical Biology, 80(2):319– 334, February 2018. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-017-0376-0;   http://link.springer.com/content/
pdf/10.1007/s11538-017-0376-0.pdf.

Almet:2018:MMI

[3332] Axel A. Almet, Barry D. Hughes, and Kerry A. Landman. A multicellular model of intestinal crypt buckling and fission. Bulletin of Mathematical Biology, 80(2):335–359, February 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-017-0377-z;
 http://link.springer.com/content/pdf/10.1007/s11538-017-0377-
z.pdf.

Nikolov:2018:MBP

[3333] Svetoslav Nikolov, Guido Santos, Olaf Wolkenhauer, and Julio Vera. Model-based phenotypic signatures governing the dynamics of the stem and semi-differentiated cell populations in dysplastic colonic crypts. Bulletin of Mathematical Biology, 80(2):360–384, February 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0378-y;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0378-y.pdf.

Guo:2018:CCP

[3334] Qiang-Hui Guo and Lisa H. Sun. Combinatorics of contacts in protein contact maps. Bulletin of Mathematical Biology, 80(2):385–403, February 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0380-4;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0380-4.pdf.

Francis:2018:TBU

[3335] A. Francis, K. T. Huber, and V. Moulton. Tree-based unrooted phylogenetic networks. Bulletin of Mathematical Biology, 80(2):404–416, February 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0381-3;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0381-3.pdf. See correction [3485].

Liz:2018:GPG

[3336] Eduardo Liz. A global picture of the gamma–Ricker map: a flexible discrete-time model with factors of positive and negative density dependence. Bulletin of Mathematical Biology, 80(2):417–434, February 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0382-2;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0382-2.pdf.

Zmurchok:2018:CAQ

[3337] Cole Zmurchok, Tim Small, Michael J. Ward, and Leah Edelstein-Keshet. Correction to: Application of Quasi-Steady-State Methods to Nonlinear Models of Intracellular Transport by Molecular Motors. Bulletin of Mathematical Biology, 80(2):435, February 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0386-y;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0386-y.pdf. See [3275].

Nwankwo:2018:MAT

[3338] A. Nwankwo and D. Okuonghae. Mathematical analysis of the transmission dynamics of HIV syphilis co-infection in the presence of treatment for syphilis. Bulletin of Mathematical Biology, 80(3):437–492, March 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0384-0;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0384-0.pdf.

Liebscher:2018:NGI

[3339] Volkmar Liebscher. New Gromov–inspired metrics on phylogenetic tree space. Bulletin of Mathematical Biology, 80(3):493–518, March 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0385-z;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0385-z.pdf.

Hanin:2018:SMP

[3340] Leonid Hanin and Jason Rose. Suppression of metastasis by primary tumor and acceleration of metastasis following primary tumor resection: a natural law? Bulletin of Mathematical Biology, 80(3):519–539, March 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-

017-0388-9;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0388-9.pdf.

Wang:2018:CMP

[3341] Wei Wang and Tongqian Zhang. Caspase-1-mediated pyroptosis of the predominance for driving CD_{4}^{+} T cells death: a nonlocal spatial mathematical model. Bulletin of Mathematical Biology, 80(3):540–582, March 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0389-8;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0389-8.pdf.

Lee:2018:MMC

[3342] Seunggyu Lee. Mathematical model of contractile ring-driven cytokinesis in a three-dimensional domain. Bulletin of Mathematical Biology, 80 (3):583–597, March 2018. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-018-0390-x;   http://link.springer.com/content/
pdf/10.1007/s11538-018-0390-x.pdf.

Valega-Mackenzie:2018:CVS

[3343] Wencel Valega-Mackenzie and Karen R. Ríos-Soto. Can vaccination save a Zika virus epidemic? Bulletin of Mathematical Biology, 80(3): 598–625, March 2018. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-018-0393-7;   http://link.springer.com/content/
pdf/10.1007/s11538-018-0393-7.pdf.

Verma:2018:MEP

[3344] Maitri Verma and A. K. Misra. Modeling the effect of prey refuge on a ratio-dependent predator–prey system with the Allee effect. Bulletin of Mathematical Biology, 80(3):626–656, March 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0394-6;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0394-6.pdf.

Asaduzzaman:2018:ECI

[3345] Sarder Mohammed Asaduzzaman, Junling Ma, and P. van den Driessche. Estimation of cross-immunity between drifted strains of influenza A/H3N2. Bulletin of Mathematical Biology, 80(3):657–669, March 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-

018-0395-5;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0395-5.pdf.

Hiller:2018:ARR

[3346] Josh Hiller and James Keesling. Asymptotic relative risk results from a simplified Armitage and Doll model of carcinogenesis. Bulletin of Mathematical Biology, 80(3):670–686, March 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0397-3;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0397-3.pdf.

Nishimura:2018:EII

[3347] Joel Nishimura, Rebecca Smith, Kathleen Jensen, Gerald Ankley, and Karen Watanabe. Estimating intermittent individual spawning behavior via disaggregating group data. Bulletin of Mathematical Biology, 80 (3):687–700, March 2018. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-017-0379-x;   http://link.springer.com/content/
pdf/10.1007/s11538-017-0379-x.pdf.

Hodgkinson:2018:CAA

[3348] Arran Hodgkinson, Mark A. J. Chaplain, Pia Domschke, and Dumitru Trucu. Computational approaches and analysis for a spatio–structuraltemporal invasive carcinoma model. Bulletin of Mathematical Biology, 80(4):701–737, April 2018. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-018-0396-4;   http://link.springer.com/content/
pdf/10.1007/s11538-018-0396-4.pdf.

Matsiaka:2018:DCA

[3349] Oleksii M. Matsiaka, Catherine J. Penington, Ruth E. Baker, and Matthew J. Simpson. Discrete and continuum approximations for collective cell migration in a scratch assay with cell size dynamics. Bulletin of Mathematical Biology, 80(4):738–757, April 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0398-2;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0398-2.pdf.

Edgington:2018:MAE

[3350] Matthew P. Edgington and Marcus J. Tindall. Mathematical analysis of the Escherichia coli chemotaxis signalling pathway. Bulletin of Mathematical Biology, 80(4):758–787, April 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-018-0400-z;
 http://link.springer.com/content/pdf/10.1007/s11538-018-0400-
z.pdf.

Oduro:2018:MDV

[3351] Bismark Oduro, Mario J. Grijalva, and Winfried Just. Models of disease vector control: When can aggressive initial intervention lower longterm cost? Bulletin of Mathematical Biology, 80(4):788–824, April 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0401-y;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0401-y.pdf.

Iboi:2018:CMS

[3352] E. Iboi, K. Okuneye, O. Sharomi, and A. B. Gumel. Comments on “A Mathematical Study to Control Visceral Leishmaniasis: an Application to South Sudan”. Bulletin of Mathematical Biology, 80(4):825–839, April 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0403-9;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0403-9.pdf. See [3241].

Zhang:2018:TPM

[3353] Juping Zhang, Chris Cosner, and Huaiping Zhu. Two-patch model for the spread of West Nile virus. Bulletin of Mathematical Biology, 80 (4):840–863, April 2018. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-018-0404-8;   http://link.springer.com/content/
pdf/10.1007/s11538-018-0404-8.pdf.

Herbst:2018:AAS

[3354] Lina Herbst and Mareike Fischer. On the accuracy of ancestral sequence reconstruction for ultrametric trees with parsimony. Bulletin of Mathematical Biology, 80(4):864–879, April 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-018-0407-5;
 http://link.springer.com/content/pdf/10.1007/s11538-018-0407-
5.pdf.

Pilvankar:2018:MMG

[3355] Minu R. Pilvankar, Michele A. Higgins, and Ashlee N. Ford Versypt. Mathematical model for glucose dependence of the local renin– angiotensin system in podocytes. Bulletin of Mathematical Biology, 80 (4):880–905, April 2018. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-018-0408-4;   http://link.springer.com/content/
pdf/10.1007/s11538-018-0408-4.pdf.

Pal:2018:APP

[3356] S. Pal, S. Ghorai, and M. Banerjee. Analysis of a prey–predator model with non-local interaction in the prey population. Bulletin of Mathematical Biology, 80(4):906–925, April 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-018-0410-x;
 http://link.springer.com/content/pdf/10.1007/s11538-018-0410-
x.pdf.

Bennie:2018:USR

[3357] Barbara Bennie, Eric Alan Eager, James P. Peirce, and Gregory J. Sandland. Using a Summer REU to help develop the next generation of mathematical ecologists. Bulletin of Mathematical Biology, 80 (4):926–944, April 2018. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-018-0405-7;   http://link.springer.com/content/
pdf/10.1007/s11538-018-0405-7.pdf.

Anderson:2018:MO

[3358] Alexander R. A. Anderson and Philip K. Maini. Mathematical oncology. Bulletin of Mathematical Biology, 80(5):945–953, May 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0423-5;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0423-5.pdf.

Gravenmier:2018:AST

[3359] Curtis A. Gravenmier, Miriam Siddique, and Robert A. Gatenby. Adaptation to stochastic temporal variations in intratumoral blood flow: The Warburg effect as a bet hedging strategy. Bulletin of Mathematical Biology, 80(5):954–970, May 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-017-0261-x;   http://link.springer.com/
content/pdf/10.1007/s11538-017-0261-x.pdf.

Jain:2018:MMC

[3360] Harsh Jain and Trachette Jackson. Mathematical modeling of cellular cross-talk between endothelial and tumor cells highlights counterintuitive effects of VEGF-targeted therapies. Bulletin of Mathematical Biology, 80(5):971–1016, May 2018. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-017-0273-6;   http://link.springer.com/content/
pdf/10.1007/s11538-017-0273-6.pdf.

Giverso:2018:HNM

[3361] Chiara Giverso, Alessandro Arduino, and Luigi Preziosi. How nucleus mechanics and ECM microstructure influence the invasion of single cells and multicellular aggregates. Bulletin of Mathematical Biology, 80(5): 1017–1045, May 2018. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-017-0262-9;   http://link.springer.com/content/
pdf/10.1007/s11538-017-0262-9.pdf.

Araujo:2018:SMM

[3362] Arturo Araujo, Leah M. Cook, Conor C. Lynch, and David Basanta. Size matters: Metastatic cluster size and stromal recruitment in the establishment of successful prostate cancer to bone metastases. Bulletin of Mathematical Biology, 80(5):1046–1058, May 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0416-4;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0416-4.pdf.

Talkington:2018:ODE

[3363] Anne Talkington, Claudia Dantoin, and Rick Durrett. Ordinary differential equation models for adoptive immunotherapy. Bulletin of Mathematical Biology, 80(5):1059–1083, May 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-017-0263-8;
 http://link.springer.com/content/pdf/10.1007/s11538-017-0263-
8.pdf.

Besse:2018:SAM

[3364] Apollos Besse, Geoffrey D. Clapp, Samuel Bernard, Franck E. Nicolini, Doron Levy, and Thomas Lepoutre. Stability analysis of a model of interaction between the immune system and cancer cells in chronic myelogenous leukemia. Bulletin of Mathematical Biology, 80(5):1084–1110, May 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0272-7;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0272-7.pdf.

Friedman:2018:REP

[3365] Avner Friedman and Wenrui Hao. The role of exosomes in pancreatic cancer microenvironment. Bulletin of Mathematical Biology, 80 (5):1111–1133, May 2018. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-017-0254-9;   http://link.springer.com/content/
pdf/10.1007/s11538-017-0254-9.pdf.

Hoehme:2018:MPV

[3366] Stefan Hoehme, François Bertaux, William Weens, Bettina Grasl-Kraupp, Jan G. Hengstler, and Dirk Drasdo. Model prediction and validation of an order mechanism controlling the spatiotemporal phenotype of early hepatocellular carcinoma. Bulletin of Mathematical Biology, 80(5):1134–1171, May 2018. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-017-0375-1;   http://link.springer.com/content/
pdf/10.1007/s11538-017-0375-1.pdf.

Burazin:2018:PTA

[3367] Andrijana Burazin, Corina S. Drapaca, Giuseppe Tenti, and Siv Sivaloganathan. A poroelasticity theory approach to study the mechanisms leading to elevated interstitial fluid pressure in solid tumours. Bulletin of Mathematical Biology, 80(5):1172–1194, May 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0383-1;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0383-1.pdf.

Poleszczuk:2018:PPS

[3368] Jan Poleszczuk, Rachel Walker, Eduardo G. Moros, Kujtim Latifi, Jimmy J. Caudell, and Heiko Enderling. Predicting patient-specific radiotherapy protocols based on mathematical model choice for proliferation saturation index. Bulletin of Mathematical Biology, 80(5):1195–1206, May 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0279-0;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0279-0.pdf.

Lewin:2018:ETC

[3369] Thomas D. Lewin, Philip K. Maini, Eduardo G. Moros, Heiko Enderling, and Helen M. Byrne. The evolution of tumour composition during fractionated radiotherapy: Implications for outcome. Bulletin of Mathematical Biology, 80(5):1207–1235, May 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0391-9;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0391-9.pdf.

Williams:2018:ADM

[3370] Katherine S. Williams, Timothy W. Secomb, and Ardith W. El-Kareh. Additive damage models for cellular pharmacodynamics of radiation– chemotherapy combinations. Bulletin of Mathematical Biology, 80(5): 1236–1258, May 2018. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-017-0316-z;   http://link.springer.com/content/
pdf/10.1007/s11538-017-0316-z.pdf.

Swan:2018:PSA

[3371] Amanda Swan, Thomas Hillen, John C. Bowman, and Albert D. Murtha. A patient-specific anisotropic diffusion model for brain tumour spread. Bulletin of Mathematical Biology, 80(5):1259–1291, May 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0271-8;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0271-8.pdf.

Massey:2018:SPD

[3372] Susan Christine Massey, Russell C. Rockne, Andrea Hawkins-Daarud, Jill Gallaher, Alexander R. A. Anderson, Peter Canoll, and Kristin R. Swanson. Simulating PDGF-driven glioma growth and invasion in an anatomically accurate brain domain. Bulletin of Mathematical Biology, 80(5):1292–1309, May 2018. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-017-0312-3;   http://link.springer.com/content/
pdf/10.1007/s11538-017-0312-3.pdf.

Wynn:2018:IIS

[3373] Michelle L. Wynn, Megan Egbert, Nikita Consul, Jungsoo Chang, Zhi-Fen Wu, Sofia D. Meravjer, and Santiago Schnell. Inferring intracellular signal transduction circuitry from molecular perturbation experiments. Bulletin of Mathematical Biology, 80(5):1310–1344, May 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0270-9;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0270-9.pdf.

Komarova:2018:SCN

[3374] Natalia L. Komarova and P. van den Driessche. Stability of control networks in autonomous homeostatic regulation of stem cell lineages. Bulletin of Mathematical Biology, 80(5):1345–1365, May 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0283-4;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0283-4.pdf.

Szymanska:2018:CMC

[3375] Zuzanna Szymańska, Maciej Cytowski, Elaine Mitchell, Cicely K. Macnamara, and Mark A. J. Chaplain. Computational modelling of cancer development and growth: Modelling at multiple scales and multiscale modelling. Bulletin of Mathematical Biology, 80(5):1366–1403, May 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0292-3;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0292-3.pdf.

Yan:2018:TDS

[3376] Huaming Yan, Anna Konstorum, and John S. Lowengrub. Threedimensional spatiotemporal modeling of colon cancer organoids reveals that multimodal control of stem cell self-renewal is a critical determinant of size and shape in early stages of tumor growth. Bulletin of Mathematical Biology, 80(5):1404–1433, May 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-017-0294-1;
 http://link.springer.com/content/pdf/10.1007/s11538-017-0294-
1.pdf.

Joshi:2018:CDH

[3377] Tanvi V. Joshi, Daniele Avitabile, and Markus R. Owen. Capturing the dynamics of a hybrid multiscale cancer model with a continuum model. Bulletin of Mathematical Biology, 80(6):1435–1475, June 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0406-6;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0406-6.pdf.

Otto:2018:HPV

[3378] Garrett Otto, Sharon Bewick, Bingtuan Li, and William F. Fagan. How phenological variation affects species spreading speeds. Bulletin of Mathematical Biology, 80(6):1476–1513, June 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-018-0409-3;
 http://link.springer.com/content/pdf/10.1007/s11538-018-0409-
3.pdf.

Li:2018:RSR

[3379] Thomas J. X. Li and Christian M. Reidys. The rainbow spectrum of RNA secondary structures. Bulletin of Mathematical Biology, 80(6): 1514–1538, June 2018. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-018-0411-9;   http://link.springer.com/content/
pdf/10.1007/s11538-018-0411-9.pdf.

Macfarlane:2018:MIR

[3380] Fiona R. Macfarlane, Tommaso Lorenzi, and Mark A. J. Chaplain. Modelling the immune response to cancer: an individual-based approach accounting for the difference in movement between inactive and activated T cells. Bulletin of Mathematical Biology, 80(6):1539–1562, June 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0412-8;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0412-8.pdf.

Grunewald:2018:RUP

[3381] Stefan Grünewald, Yangjing Long, and Yaokun Wu. Reconstructing unrooted phylogenetic trees from symbolic ternary metrics. Bulletin of Mathematical Biology, 80(6):1563–1577, June 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-018-0413-7;
 http://link.springer.com/content/pdf/10.1007/s11538-018-0413-
7.pdf.

Lagergren:2018:FUQ

[3382] John Lagergren, Amanda Reeder, Franz Hamilton, Ralph C. Smith, and Kevin B. Flores. Forecasting and uncertainty quantification using a hybrid of mechanistic and non-mechanistic models for an age-structured population model. Bulletin of Mathematical Biology, 80(6):1578–1595, June 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0421-7;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0421-7.pdf.

Bashkirtseva:2018:SSA

[3383] Irina Bashkirtseva and Lev Ryashko. Stochastic sensitivity analysis of noise-induced extinction in the Ricker model with delay and Allee effect. Bulletin of Mathematical Biology, 80(6):1596–1614, June 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0422-6;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0422-6.pdf.

Jenner:2018:MMI

[3384] Adrianne L. Jenner, Chae-Ok Yun, Peter S. Kim, and Adelle C. F. Coster. Mathematical modelling of the interaction between cancer cells and an oncolytic virus: Insights into the effects of treatment protocols. Bulletin of Mathematical Biology, 80(6):1615–1629, June 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0424-4;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0424-4.pdf.

ORegan:2018:HSI

[3385] Suzanne M. O’Regan and Danielle L. Burton. How stochasticity influences leading indicators of critical transitions. Bulletin of Mathematical Biology, 80(6):1630–1654, June 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-018-0429-z;
 http://link.springer.com/content/pdf/10.1007/s11538-018-0429-
z.pdf.

Rodriguez:2018:PDB

[3386] N. Rodríguez and G. Malanson. Plant dynamics, birth–jump processes, and sharp traveling waves. Bulletin of Mathematical Biology, 80(6): 1655–1687, June 2018. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-018-0431-5;   http://link.springer.com/content/
pdf/10.1007/s11538-018-0431-5.pdf.

Dalmau:2018:ABE

[3387] Joseba Dalmau. Asymptotic behavior of Eigen’s quasispecies model. Bulletin of Mathematical Biology, 80(7):1689–1712, July 2018. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-018-0420-8;
 http://link.springer.com/content/pdf/10.1007/s11538-018-0420-
8.pdf.

Lin:2018:AFD

[3388] Chiu-Ju Lin, Lin Wang, and Gail S. K. Wolkowicz. An alternative formulation for a distributed delayed logistic equation. Bulletin of Mathematical Biology, 80(7):1713–1735, July 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-018-0432-4;
 http://link.springer.com/content/pdf/10.1007/s11538-018-0432-
4.pdf.

Ghasemi:2018:SBE

[3389] Maryam Ghasemi, Burkhard A. Hense, Hermann J. Eberl, and Christina Kuttler. Simulation-based exploration of quorum sensing triggered resistance of biofilms to antibiotics. Bulletin of Mathematical Biology, 80 (7):1736–1775, July 2018. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-018-0433-3;   http://link.springer.com/content/
pdf/10.1007/s11538-018-0433-3.pdf.

Yoon:2018:OTS

[3390] Nara Yoon, Robert Vander Velde, Andriy Marusyk, and Jacob G. Scott. Optimal therapy scheduling based on a pair of collaterally sensitive drugs. Bulletin of Mathematical Biology, 80(7):1776–1809, July 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0434-2;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0434-2.pdf.

Bichara:2018:MSV

[3391] Derdei Bichara, Abderrahman Iggidr, and Laura Smith. Multistage vector-borne zoonoses models: a global analysis. Bulletin of Mathematical Biology, 80(7):1810–1848, July 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-018-0435-1;
 http://link.springer.com/content/pdf/10.1007/s11538-018-0435-
1.pdf.

Dipierro:2018:SMM

[3392] Serena Dipierro and Enrico Valdinoci. A simple mathematical model inspired by the Purkinje cells: From delayed travelling waves to fractional diffusion. Bulletin of Mathematical Biology, 80(7):1849–1870, July 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0437-z;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0437-z.pdf.

Bokes:2018:HCN

[3393] Pavol Bokes, Yen Ting Lin, and Abhyudai Singh. High cooperativity in negative feedback can amplify noisy gene expression. Bulletin of Mathematical Biology, 80(7):1871–1899, July 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0438-y;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0438-y.pdf.

Hodgkinson:2018:SPS

[3394] Arran Hodgkinson, Gilles Uzé, Ovidiu Radulescu, and Dumitru Trucu. Signal propagation in sensing and reciprocating cellular systems with spatial and structural heterogeneity. Bulletin of Mathematical Biology, 80(7):1900–1936, July 2018. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-018-0439-x;   http://link.springer.com/content/
pdf/10.1007/s11538-018-0439-x.pdf.

Farrell:2018:TID

[3395] Alex P. Farrell, James P. Collins, Amy L. Greer, and Horst R. Thieme. Times from infection to disease-induced death and their influence on final population sizes after epidemic outbreaks. Bulletin of Mathematical Biology, 80(7):1937–1961, July 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-018-0446-y;
 http://link.springer.com/content/pdf/10.1007/s11538-018-0446-
y.pdf.

Wang:2018:TDT

[3396] Xiunan Wang and Xingfu Zou. Threshold dynamics of a temperaturedependent stage-structured mosquito population model with nested delays. Bulletin of Mathematical Biology, 80(7):1962–1987, July 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0447-x;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0447-x.pdf.

Chelle:2018:ECS

[3397] Pierre Chelle, Claire Morin, Aurélie Montmartin, Michèle Piot, Michel Cournil, and Brigitte Tardy-Poncet. Evaluation and calibration of in silico models of thrombin generation using experimental data from healthy and haemophilic subjects. Bulletin of Mathematical Biology, 80(8): 1989–2025, August 2018. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-018-0440-4;   http://link.springer.com/content/
pdf/10.1007/s11538-018-0440-4.pdf.

He:2018:PMN

[3398] Yixuan He, Anita Kodali, and Dorothy I. Wallace. Predictive modeling of neuroblastoma growth dynamics in Xenograft model after Bevacizumab anti–VEGF therapy. Bulletin of Mathematical Biology, 80(8): 2026–2048, August 2018. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-018-0441-3;   http://link.springer.com/content/
pdf/10.1007/s11538-018-0441-3.pdf.

Chen:2018:TDS

[3399] Shanshan Chen, Michael Small, Yizhou Tao, and Xinchu Fu. Transmission dynamics of an SIS model with age structure on heterogeneous networks. Bulletin of Mathematical Biology, 80(8):2049–2087, August 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0445-z;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0445-z.pdf.

Smith:2018:AIE

[3400] J. Darby Smith and Scott A. McKinley. Assessing the impact of electrostatic drag on processive molecular motor transport. Bulletin of Mathematical Biology, 80(8):2088–2123, August 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-018-0448-9;
 http://link.springer.com/content/pdf/10.1007/s11538-018-0448-
9.pdf.

Huang:2018:SRB

[3401] Bin Huang, Ning Zhong, Lili Xia, Guiping Yu, and Hongbao Cao. Sparse representation-based patient-specific diagnosis and treatment for esophageal squamous cell carcinoma. Bulletin of Mathematical Biology, 80(8):2124–2136, August 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-018-0449-8;
 http://link.springer.com/content/pdf/10.1007/s11538-018-0449-
8.pdf.

Huber:2018:QIR

[3402] Katharina T. Huber, Vincent Moulton, Charles Semple, and Taoyang Wu. Quarnet inference rules for level-1 networks. Bulletin of Mathematical Biology, 80(8):2137–2153, August 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-018-0450-2;
 http://link.springer.com/content/pdf/10.1007/s11538-018-0450-
2.pdf.

Klemm:2018:CEF

[3403] Konstantin Klemm, Anita Mehta, and Peter F. Stadler. Cover-encodings of fitness landscapes. Bulletin of Mathematical Biology, 80(8):2154– 2176, August 2018. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-018-0451-1;   http://link.springer.com/content/
pdf/10.1007/s11538-018-0451-1.pdf.

Janssen:2018:ETR

[3404] Remie Janssen, Mark Jones, Péter L. Erdős, Leo van Iersel, and Celine Scornavacca. Exploring the tiers of rooted phylogenetic network space using tail moves. Bulletin of Mathematical Biology, 80(8):2177–2208, August 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0452-0;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0452-0.pdf.

Tuncer:2018:SPI

[3405] Necibe Tuncer, Maia Marctheva, Brian LaBarre, and Sabrina Payoute. Structural and practical identifiability analysis of Zika epidemiological models. Bulletin of Mathematical Biology, 80(8):2209–2241, August 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-

018-0453-z;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0453-z.pdf.

Caudill:2018:MMI

[3406] Lester Caudill and Fiona Lynch. A mathematical model of the inflammatory response to pathogen challenge. Bulletin of Mathematical Biology, 80(8):2242–2271, August 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-018-0459-6;   http://link.springer.com/
content/pdf/10.1007/s11538-018-0459-6.pdf.

Mahdipour-Shirayeh:2018:MCD

[3407] Ali Mahdipour-Shirayeh and Leili Shahriyari. Modeling cell dynamics in colon and intestinal crypts: The significance of central stem cells in tumorigenesis. Bulletin of Mathematical Biology, 80(9):2273–2305, September 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0457-8;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0457-8.pdf.

Tonello:2018:NTS

[3408] Elisa Tonello and Matthew D. Johnston. Network translation and steady-state properties of chemical reaction systems. Bulletin of Mathematical Biology, 80(9):2306–2337, September 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0458-7;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0458-7.pdf.

Semple:2018:WPN

[3409] Charles Semple and Jack Simpson. When is a phylogenetic network simply an amalgamation of two trees? Bulletin of Mathematical Biology, 80(9):2338–2348, September 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-018-0463-x;   http://link.springer.com/
content/pdf/10.1007/s11538-018-0463-x.pdf.

Pasin:2018:CIH

[3410] Chloé Pasin, François Dufour, Laura Villain, Huilong Zhang, and Rodolphe Thiébaut. Controlling IL-7 injections in HIV-infected patients. Bulletin of Mathematical Biology, 80(9):2349–2377, September 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-

018-0465-8;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0465-8.pdf.

deAssis:2018:EHI

[3411] Raul Abreu de Assis, Luca Pietro Casacci, Simona Bonelli, Francesca Barbero, Luciana Mafalda Elias de Assis, and Ezio Venturino. Effects of host interspecific interaction in the maculinea–myrmica parasite–host system. Bulletin of Mathematical Biology, 80(9):2378–2407, September 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0466-7;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0466-7.pdf.

Wang:2018:DIP

[3412] Yuanshi Wang, Hong Wu, Shikun Wang, and Wen Shi. Dynamics of intraguild predation systems with intraspecific competition. Bulletin of Mathematical Biology, 80(9):2408–2434, September 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0467-6;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0467-6.pdf.

Qi:2018:STM

[3413] Longxing Qi, Meng Xue, Jing an Cui, Qizhi Wang, and Tianping Wang. Schistosomiasis transmission model and its control in Anhui Province. Bulletin of Mathematical Biology, 80(9):2435–2451, September 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0474-7;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0474-7.pdf.

Kabir:2018:SWF

[3414] M. H. Kabir, M. Mimura, and J. C. Tsai. Spreading waves in a farmers and hunter-gatherers model of the Neolithic transition in Europe. Bulletin of Mathematical Biology, 80(9):2452–2480, September 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0475-6;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0475-6.pdf.

Simons:2018:MAR

[3415] Julie Simons and Lisa Fauci. A model for the acrosome reaction in mammalian sperm. Bulletin of Mathematical Biology, 80(9):2481– 2501, September 2018. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-018-0478-3;   http://link.springer.com/content/
pdf/10.1007/s11538-018-0478-3.pdf.

Leman:2018:SMR

[3416] Hélène Leman. A stochastic model for reproductive isolation under asymmetrical mating preferences. Bulletin of Mathematical Biology, 80(9):2502–2525, September 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-018-0479-2;   http://link.springer.com/
content/pdf/10.1007/s11538-018-0479-2.pdf.

Hening:2018:PSL

[3417] Alexandru Hening and Dang H. Nguyen. Persistence in stochastic Lotka–Volterra food chains with intraspecific competition. Bulletin of Mathematical Biology, 80(10):2527–2560, October 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-018-0468-5;
 http://link.springer.com/content/pdf/10.1007/s11538-018-0468-
5.pdf.

Anderson:2018:NES

[3418] David F. Anderson, Daniele Cappelletti, Masanori Koyama, and Thomas G. Kurtz. Non-explosivity of stochastically modeled reaction networks that are complex balanced. Bulletin of Mathematical Biology, 80(10):2561–2579, October 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-018-0473-8;   http://link.springer.com/
content/pdf/10.1007/s11538-018-0473-8.pdf.

Kobe:2018:GTM

[3419] Julia Kobe, Neil Pritchard, Ziaqueria Short, Igor V. Erovenko, Jan Rychtář, and Jonathan T. Rowell. A game-theoretic model of cholera with optimal personal protection strategies. Bulletin of Mathematical Biology, 80(10):2580–2599, October 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-018-0476-5;
 http://link.springer.com/content/pdf/10.1007/s11538-018-0476-
5.pdf.

Bitsouni:2018:NLP

[3420] Vasiliki Bitsouni and Raluca Eftimie. Non-local parabolic and hyperbolic models for cell polarisation in heterogeneous cancer cell populations. Bulletin of Mathematical Biology, 80(10):2600–2632, October 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0477-4;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0477-4.pdf.

Zou:2018:ADM

[3421] Lan Zou, Jing Chen, Xiaomei Feng, and Shigui Ruan. Analysis of a dengue model with vertical transmission and application to the 2014 dengue outbreak in Guangdong Province, China. Bulletin of Mathematical Biology, 80(10):2633–2651, October 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0480-9;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0480-9.pdf.

Stratton:2018:OMS

[3422] Corin Stratton, Allison Kolpas, and Josh R. Auld. Optimal mating strategies for preferentially outcrossing simultaneous hermaphrodites in the presence of predators. Bulletin of Mathematical Biology, 80(10): 2652–2668, October 2018. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-018-0481-8;   http://link.springer.com/content/
pdf/10.1007/s11538-018-0481-8.pdf.

Wentz:2018:PFL

[3423] J. M. Wentz, A. R. Mendenhall, and D. M. Bortz. Pattern formation in the longevity-related expression of heat shock protein-16.2 in Caenorhabditis elegans. Bulletin of Mathematical Biology, 80(10):2669–2697, October 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0482-7;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0482-7.pdf.

Barnard:2018:EBC

[3424] Rosanna C. Barnard, Istvan Z. Kiss, Luc Berthouze, and Joel C. Miller. Edge-based compartmental modelling of an SIR epidemic on a dual-layer static–dynamic multiplex network with tunable clustering. Bulletin of Mathematical Biology, 80(10):2698–2733, October 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0484-5;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0484-5.pdf.

Kasman:2018:DGC

[3425] Alex Kasman. The duplexing of the genetic code and sequence-dependent DNA geometry. Bulletin of Mathematical Biology, 80(10):2734–2760, October 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0486-3;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0486-3.pdf.

Tanaka:2018:PLD

[3426] Cinthia M. Tanaka, Joung-Hun Lee, and Yoh Iwasa. The persistence of a local dialect when a national standard language is present: an evolutionary dynamics model of cultural diversity. Bulletin of Mathematical Biology, 80(10):2761–2786, October 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0487-2;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0487-2.pdf.

Fowler:2018:CDA

[3427] A. C. Fowler and T. Déirdre Hollingsworth. Correction to: The Dynamics of Ascaris lumbricoides Infections. Bulletin of Mathematical Biology, 80(10):2787, October 2018. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-018-0470-y;   http://link.springer.com/content/
pdf/10.1007/s11538-018-0470-y.pdf.

Hilker:2018:CDS

[3428] Frank M. Hilker, Marta Paliaga, and Ezio Venturino. Correction to: Diseased Social Predators. Bulletin of Mathematical Biology, 80(10): 2788, October 2018. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-018-0471-x;   http://link.springer.com/content/
pdf/10.1007/s11538-018-0471-x.pdf. See [3285].

Sfakianakis:2018:SCS

[3429] N. Sfakianakis and A. Brunk. Stability, convergence, and sensitivity analysis of the FBLM and the corresponding FEM. Bulletin of Mathematical Biology, 80(11):2789–2827, November 2018.

CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0460-0;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0460-0.pdf.

Surendran:2018:SMD

[3430] Anudeep Surendran, Michael J. Plank, and Matthew J. Simpson. Spatial moment description of birth–death-movement processes incorporating the effects of crowding and obstacles. Bulletin of Mathematical Biology, 80(11):2828–2855, November 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0488-1;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0488-1.pdf.

Talu:2018:MCB

[3431] Ştefan Ţălu, Ilya A. Morozov, Dinara Sobola, and Pavel Škarvada. Multifractal characterization of butterfly wings scales. Bulletin of Mathematical Biology, 80(11):2856–2870, November 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-018-0490-7;
 http://link.springer.com/content/pdf/10.1007/s11538-018-0490-
7.pdf.

Clancy:2018:PEP

[3432] Damian Clancy. Precise estimates of persistence time for SIS infections in heterogeneous populations. Bulletin of Mathematical Biology, 80(11): 2871–2896, November 2018. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-018-0491-6;   http://link.springer.com/content/
pdf/10.1007/s11538-018-0491-6.pdf.

Amster:2018:CPA

[3433] Pablo Amster and Carlos Alliera. Control of pyragas applied to a coupled system with unstable periodic orbits. Bulletin of Mathematical Biology, 80(11):2897–2916, November 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-018-0492-5;
 http://link.springer.com/content/pdf/10.1007/s11538-018-0492-
5.pdf.

Latif:2018:MAB

[3434] Majid Latif, Jr. and Elebeoba E. May. A multiscale agent-based model for the investigation of E. coli K12 metabolic response during biofilm formation. Bulletin of Mathematical Biology, 80(11):2917–2956, November 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0494-3;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0494-3.pdf.

Andreasen:2018:ECP

[3435] Viggo Andreasen. Epidemics in competition: Partial cross–immunity. Bulletin of Mathematical Biology, 80(11):2957–2977, November 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0495-2;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0495-2.pdf.

Saldana:2018:CSM

[3436] Fernando Saldaña and Ignacio Barradas. Control strategies in multigroup models: The case of the star network topology. Bulletin of Mathematical Biology, 80(11):2978–3001, November 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0503-6;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0503-6.pdf.

Pearce:2018:CSK

[3437] Simon P. Pearce, Matthias Heil, Oliver E. Jensen, Gareth Wyn Jones, and Andreas Prokop. Curvature-sensitive kinesin binding can explain microtubule ring formation and reveals chaotic dynamics in a mathematical model. Bulletin of Mathematical Biology, 80(11):3002–3022, November 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0505-4;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0505-4.pdf.

Rabinovitch:2018:ROR

[3438] A. Rabinovitch, I. Aviram, Y. Biton, and D. Braunstein. Reentry as an origin for rotors. Bulletin of Mathematical Biology, 80(11):3023– 3037, November 2018. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-018-0506-3;   http://link.springer.com/content/
pdf/10.1007/s11538-018-0506-3.pdf.

Sasmal:2018:MSZ

[3439] Sourav Kumar Sasmal, Indrajit Ghosh, Amit Huppert, and Joydev Chattopadhyay. Modeling the spread of Zika virus in a stage-structured population: Effect of sexual transmission. Bulletin of Mathematical Biology, 80(11):3038–3067, November 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0510-7;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0510-7.pdf.

Laubenbacher:2018:E

[3440] Reinhard Laubenbacher and Alan Hastings. Editorial. Bulletin of Mathematical Biology, 80(12):3069–3070, December 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-018-0501-8;
 http://link.springer.com/content/pdf/10.1007/s11538-018-0501-
8.pdf.

An:2018:CRD

[3441] Gary An. The crisis of reproducibility, the denominator problem and the scientific role of multi-scale modeling. Bulletin of Mathematical Biology, 80(12):3071–3080, December 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-018-0497-0;
 http://link.springer.com/content/pdf/10.1007/s11538-018-0497-
0.pdf.

Mendes:2018:RRU

[3442] Pedro Mendes. Reproducible research using biomodels. Bulletin of Mathematical Biology, 80(12):3081–3087, December 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0498-z;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0498-z.pdf.

Gavaghan:2018:PCA

[3443] David Gavaghan. Problems with the current approach to the dissemination of computational science research and its implications for research integrity. Bulletin of Mathematical Biology, 80(12):3088–3094, December 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0499-y;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0499-y.pdf.

Schnell:2018:RRM

[3444] Santiago Schnell. “Reproducible” research in mathematical sciences requires changes in our peer review culture and modernization of our current publication approach. Bulletin of Mathematical Biology, 80(12): 3095–3105, December 2018. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-018-0500-9;   http://link.springer.com/content/
pdf/10.1007/s11538-018-0500-9.pdf.

Sugaya:2018:ATI

[3445] S. Sugaya and V. M. Kenkre. Analysis of transmission of infection in epidemics: Confined random walkers in dimensions higher than one. Bulletin of Mathematical Biology, 80(12):3106–3126, December 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0507-2;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0507-2.pdf.

He:2018:SMA

[3446] Sha He, Sanyi Tang, Yanni Xiao, and Robert A. Cheke. Stochastic modelling of air pollution impacts on respiratory infection risk. Bulletin of Mathematical Biology, 80(12):3127–3153, December 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0512-5;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0512-5.pdf.

Eilertsen:2018:KAC

[3447] Justin Eilertsen and Santiago Schnell. A kinetic analysis of coupled (or auxiliary) enzyme reactions. Bulletin of Mathematical Biology, 80(12): 3154–3183, December 2018. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-018-0513-4;   http://link.springer.com/content/
pdf/10.1007/s11538-018-0513-4.pdf.

Deville:2018:CMM

[3448] Manon Deville, Roberto Natalini, and Clair Poignard. A continuum mechanics model of enzyme-based tissue degradation in cancer therapies. Bulletin of Mathematical Biology, 80(12):3184–3226, December 2018. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0515-2;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0515-2.pdf.

Bhatia:2018:PCP

[3449] Sangeeta Bhatia, Pedro Feijão, and Andrew R. Francis. Position and content paradigms in genome rearrangements: The wild and crazy world of permutations in genomics. Bulletin of Mathematical Biology, 80(12): 3227–3246, December 2018. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-018-0514-3;   http://link.springer.com/content/
pdf/10.1007/s11538-018-0514-3.pdf.

Fitzpatrick:2019:IRS

[3450] B. G. Fitzpatrick. Issues in reproducible simulation research. Bulletin of Mathematical Biology, 81(1):1–6, January 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-018-0496-1;
 http://link.springer.com/content/pdf/10.1007/s11538-018-0496-
1.pdf.

Bowler:2019:RMC

[3451] Louise A. Bowler, David J. Gavaghan, Gary R. Mirams, and Jonathan P. Whiteley. Representation of multiple cellular phenotypes within tissue-level simulations of cardiac electrophysiology. Bulletin of Mathematical Biology, 81(1):7–38, January 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0516-1;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0516-1.pdf.

Zhong:2019:DFI

[3452] L. Zhong, R. J. Braun, C. G. Begley, and P. E. King-Smith. Dynamics of fluorescent imaging for rapid tear thinning. Bulletin of Mathematical Biology, 81(1):39–80, January 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-018-0517-0;
 http://link.springer.com/content/pdf/10.1007/s11538-018-0517-
0.pdf.

Campillo-Funollet:2019:BPI

[3453] Eduard Campillo-Funollet, Chandrasekhar Venkataraman, and Anotida Madzvamuse. Bayesian parameter identification for Turing systems on stationary and evolving domains. Bulletin of Mathematical Biology, 81 (1):81–104, January 2019. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-018-0518-z;   http://link.springer.com/content/
pdf/10.1007/s11538-018-0518-z.pdf.

Chakravarty:2019:NMM

[3454] Koyel Chakravarty and D. C. Dalal. A nonlinear mathematical model of drug delivery from polymeric matrix. Bulletin of Mathematical Biology, 81(1):105–130, January 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-018-0519-y;
 http://link.springer.com/content/pdf/10.1007/s11538-018-0519-
y.pdf.

Azoz:2019:SDL

[3455] Shaimaa A. Azoz and Daniel Coombs. Stochastic dynamics of the latently infected cell reservoir during HIV infection. Bulletin of Mathematical Biology, 81(1):131–154, January 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-018-0520-5;
 http://link.springer.com/content/pdf/10.1007/s11538-018-0520-
5.pdf.

Veron:2019:SMD

[3456] Alain R. Véron and Assis F. Martins. Stochastic mechanochemical description of a bioinspired polymerization process. Bulletin of Mathematical Biology, 81(1):155–192, January 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-018-0522-3;
 http://link.springer.com/content/pdf/10.1007/s11538-018-0522-
3.pdf.

Manore:2019:IPT

[3457] Carrie A. Manore, Miranda I. Teboh-Ewungkem, Olivia Prosper, Angela Peace, Katharine Gurski, and Zhilan Feng. Intermittent Preventive Treatment (IPT): Its role in averting disease-induced mortality in children and in promoting the spread of antimalarial drug resistance. Bulletin of Mathematical Biology, 81(1):193–234, January 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0524-1;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0524-1.pdf.

Croicu:2019:OCM

[3458] Ana-Maria Croicu. An optimal control model to reduce and eradicate anthrax disease in herbivorous animals. Bulletin of Mathematical Biology, 81(1):235–255, January 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-018-0525-0;   http://link.springer.com/
content/pdf/10.1007/s11538-018-0525-0.pdf.

Kadelka:2019:MBD

[3459] Sarah Kadelka, Brittany P. Boribong, Liwu Li, and Stanca M. Ciupe. Modeling the bistable dynamics of the innate immune system. Bulletin of Mathematical Biology, 81(1):256–276, January 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0527-y;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0527-y.pdf.

Haspolat:2019:DSM

[3460] E. Haspolat, B. Huard, and M. Angelova. Deterministic and stochastic models of Arabidopsis thaliana flowering. Bulletin of Mathematical Biology, 81(1):277–311, January 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-018-0528-x;   http://link.springer.com/
content/pdf/10.1007/s11538-018-0528-x.pdf.

Casanellas:2019:AMP

[3461] Marta Casanellas and John A. Rhodes. Algebraic methods in phylogenetics. Bulletin of Mathematical Biology, 81(2):313–315, February 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-018-00551-3;   http://link.springer.com/content/pdf/
10.1007/s11538-018-00551-3.pdf.

Banos:2019:DGB

[3462] Hector Baños, Nathaniel Bushek, Ruth Davidson, Elizabeth Gross, Pamela E. Harris, Robert Krone, Colby Long, Allen Stewart, and Robert Walker. Dimensions of group-based phylogenetic mixtures. Bulletin of Mathematical Biology, 81(2):316–336, February 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0489-0;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0489-0.pdf.

Kosta:2019:MLE

[3463] Dimitra Kosta and Kaie Kubjas. Maximum likelihood estimation of symmetric group-based models via numerical algebraic geometry. Bulletin of Mathematical Biology, 81(2):337–360, February 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0523-2;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0523-2.pdf.

Sumner:2019:LMM

[3464] Jeremy G. Sumner and Michael D. Woodhams. Lie–Markov models derived from finite semigroups. Bulletin of Mathematical Biology, 81(2): 361–383, February 2019. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-018-0455-x;   http://link.springer.com/content/
pdf/10.1007/s11538-018-0455-x.pdf.

Disanto:2019:NNE

[3465] Filippo Disanto and Noah A. Rosenberg. On the number of nonequivalent ancestral configurations for matching gene trees and species trees. Bulletin of Mathematical Biology, 81(2):384–407, February 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
017-0342-x;   http://link.springer.com/content/pdf/10.1007/
s11538-017-0342-x.pdf.

Long:2019:IRS

[3466] Colby Long and Laura Kubatko. Identifiability and reconstructibility of species phylogenies under a modified coalescent. Bulletin of Mathematical Biology, 81(2):408–430, February 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-018-0456-9;
 http://link.springer.com/content/pdf/10.1007/s11538-018-0456-
9.pdf.

Durden:2019:IPP

[3467] Chris Durden and Seth Sullivant. Identifiability of phylogenetic parameters from k-mer data under the coalescent. Bulletin of Mathematical Biology, 81(2):431–451, February 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-018-0399-1;
 http://link.springer.com/content/pdf/10.1007/s11538-018-0399-
1.pdf.

Kim:2019:MSB

[3468] Jaehee Kim, Filippo Disanto, Naama M. Kopelman, and Noah A. Rosenberg. Mathematical and simulation-based analysis of the behavior of admixed taxa in the neighbor-joining algorithm. Bulletin of Mathematical Biology, 81(2):452–493, February 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-018-0444-0;
 http://link.springer.com/content/pdf/10.1007/s11538-018-0444-
0.pdf.

Banos:2019:ISN

[3469] Hector Baños. Identifying species network features from gene tree quartets under the coalescent model. Bulletin of Mathematical Biology, 81(2): 494–534, February 2019. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-018-0485-4;   http://link.springer.com/content/
pdf/10.1007/s11538-018-0485-4.pdf.

Terauds:2019:MLE

[3470] Venta Terauds and Jeremy Sumner. Maximum likelihood estimates of rearrangement distance: Implementing a representation-theoretic approach. Bulletin of Mathematical Biology, 81(2):535–567, February 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0511-6;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0511-6.pdf.

Yoshida:2019:TPC

[3471] Ruriko Yoshida, Leon Zhang, and Xu Zhang. Tropical principal component analysis and its application to phylogenetics. Bulletin of Mathematical Biology, 81(2):568–597, February 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-018-0493-4;
 http://link.springer.com/content/pdf/10.1007/s11538-018-0493-
4.pdf.

Huber:2019:PFH

[3472] Katharina T. Huber, Vincent Moulton, and Mike Steel. Phylogenetic flexibility via Hall-type inequalities and submodularity. Bulletin of Mathematical Biology, 81(2):598–617, February 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-

018-0419-1;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0419-1.pdf.

Wicke:2019:SVU

[3473] Kristina Wicke and Mareike Fischer. On the Shapley value of unrooted phylogenetic trees. Bulletin of Mathematical Biology, 81(2): 618–638, February 2019. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-018-0392-8;   http://link.springer.com/content/
pdf/10.1007/s11538-018-0392-8.pdf.

Bayliss:2019:MSB

[3474] C. D. Bayliss, C. Fallaize, R. Howitt, and M. V. Tretyakov. Mutation and selection in bacteria: Modelling and calibration. Bulletin of Mathematical Biology, 81(3):639–675, March 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-018-0529-9;
 http://link.springer.com/content/pdf/10.1007/s11538-018-0529-
9.pdf.

Browning:2019:BSL

[3475] Alexander P. Browning, Parvathi Haridas, and Matthew J. Simpson. A Bayesian sequential learning framework to parameterise continuum models of melanoma invasion into human skin. Bulletin of Mathematical Biology, 81(3):676–698, March 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-018-0532-1;
 http://link.springer.com/content/pdf/10.1007/s11538-018-0532-
1.pdf.

Vera-Siguenza:2019:MMF

[3476] Elías Vera-Sigüenza, Nathan Pages, John Rugis, David I. Yule, and James Sneyd. A mathematical model of fluid transport in an accurate reconstruction of parotid acinar cells. Bulletin of Mathematical Biology, 81(3):699–721, March 2019. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-018-0534-z;   http://link.springer.com/content/
pdf/10.1007/s11538-018-0534-z.pdf.

Honore:2019:GFA

[3477] Stéphane Honoré, Florence Hubert, Magali Tournus, and Diana White. A growth–fragmentation approach for modeling microtubule dynamic instability. Bulletin of Mathematical Biology, 81(3):722–758, March 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0531-2;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0531-2.pdf.

Krause:2019:ICG

[3478] Andrew L. Krause, Meredith A. Ellis, and Robert A. Van Gorder. Influence of curvature, growth, and anisotropy on the evolution of Turing patterns on growing manifolds. Bulletin of Mathematical Biology, 81 (3):759–799, March 2019. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-018-0535-y;   http://link.springer.com/content/
pdf/10.1007/s11538-018-0535-y.pdf.

Bressloff:2019:PEF

[3479] P. C. Bressloff and E. Levien. Propagation of extrinsic fluctuations in biochemical birth–death processes. Bulletin of Mathematical Biology, 81(3):800–829, March 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-018-00538-0;   http://link.springer.
com/content/pdf/10.1007/s11538-018-00538-0.pdf.

Balelli:2019:MTG

[3480] Irene Balelli, Vuk Milišić, and Gilles Wainrib. Multi-type Galton– Watson processes with affinity-dependent selection applied to antibody affinity maturation. Bulletin of Mathematical Biology, 81(3):830–868, March 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-018-00548-y;   http://link.springer.com/content/pdf/
10.1007/s11538-018-00548-y.pdf.

Brauer:2019:FSS

[3481] Fred Brauer. The final size of a serious epidemic. Bulletin of Mathematical Biology, 81(3):869–877, March 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-018-00549-x;   http://link.springer.
com/content/pdf/10.1007/s11538-018-00549-x.pdf.

Manceau:2019:SPM

[3482] Marc Manceau and Amaury Lambert. The species problem from the Modeler’s point of view. Bulletin of Mathematical Biology, 81(3):878– 898, March 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-018-00536-2;   http://link.springer.com/content/pdf/
10.1007/s11538-018-00536-2.pdf.

Peyravi:2019:CAP

[3483] Farzad Peyravi, Alimohammad Latif, and Seyed Mohammad Moshtaghioun. A composite approach to protein tertiary structure prediction: Hidden Markov model based on lattice. Bulletin of Mathematical Biology, 81(3):899–918, March 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-018-00542-4;   http://link.springer.
com/content/pdf/10.1007/s11538-018-00542-4.pdf.

Katauskis:2019:MEC

[3484] Pranas Katauskis, Feliksas Ivanauskas, and Aidas Alaburda. The “memory” effect in a chain of biochemical reactions with a positive feedback is enhanced by substrate saturation described by Michaelis– Menten kinetics. Bulletin of Mathematical Biology, 81(3):919–935, March 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-018-00541-5;   http://link.springer.com/content/pdf/
10.1007/s11538-018-00541-5.pdf.

Francis:2019:CTB

[3485] A. Francis, K. T. Huber, and V. Moulton. Correction to: Tree-Based Unrooted Phylogenetic Networks. Bulletin of Mathematical Biology, 81 (3):936–937, March 2019. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-018-0530-3;   http://link.springer.com/content/
pdf/10.1007/s11538-018-0530-3.pdf. See [3335].

Huang:2019:MTD

[3486] Jicai Huang, Shigui Ruan, Yaqin Shu, and Xiao Wu. Modeling the transmission dynamics of rabies for dog, Chinese ferret badger and human interactions in Zhejiang Province, China. Bulletin of Mathematical Biology, 81(4):939–962, April 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-018-00537-1;   http://link.springer.
com/content/pdf/10.1007/s11538-018-00537-1.pdf.

Wang:2019:MEN

[3487] Aili Wang, Yanni Xiao, and Robert Smith. Multiple equilibria in a non-smooth epidemic model with medical-resource constraints. Bulletin of Mathematical Biology, 81(4):963–994, April 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-018-00544-2;   http://link.springer.com/content/pdf/
10.1007/s11538-018-00544-2.pdf.

Bate:2019:PTW

[3488] Andrew M. Bate and Frank M. Hilker. Preytaxis and travelling waves in an eco-epidemiological model. Bulletin of Mathematical Biology, 81(4):995–1030, April 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-018-00546-0;   http://link.springer.
com/content/pdf/10.1007/s11538-018-00546-0.pdf.

Fabreti:2019:SMS

[3489] Luiza Guimarães Fabreti, Diogo Castro, Bruno Gorzoni, Luiz Mario Ramos Janini, and Fernando Antoneli. Stochastic modeling and simulation of viral evolution. Bulletin of Mathematical Biology, 81(4):1031–1069, April 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-018-00550-4;   http://link.springer.com/content/pdf/
10.1007/s11538-018-00550-4.pdf.

Bakhtin:2019:USI

[3490] Yuri Bakhtin. Universal statistics of incubation periods and other detection times via diffusion models. Bulletin of Mathematical Biology, 81(4):1070–1088, April 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-018-00558-w;   http://link.springer.
com/content/pdf/10.1007/s11538-018-00558-w.pdf.

Feliu:2019:MBF

[3491] Elisenda Feliu and Martin Helmer. Multistationarity and bistability for fewnomial chemical reaction networks. Bulletin of Mathematical Biology, 81(4):1089–1121, April 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-018-00555-z;   http://link.springer.
com/content/pdf/10.1007/s11538-018-00555-z.pdf.

Machta:2019:HMP

[3492] Jonathan Machta, Julie C. Blackwood, Andrew Noble, Andrew M. Liebhold, and Alan Hastings. A hybrid model for the population dynamics of periodical cicadas. Bulletin of Mathematical Biology, 81(4):1122–1142, April 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-018-00554-0;   http://link.springer.
com/content/pdf/10.1007/s11538-018-00554-0.pdf.

Johnston:2019:DBA

[3493] Matthew D. Johnston, Stefan Müller, and Casian Pantea. A deficiencybased approach to parametrizing positive equilibria of biochemical reaction systems. Bulletin of Mathematical Biology, 81(4):1143–1172, April 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-018-00562-0;   http://link.springer.com/content/pdf/
10.1007/s11538-018-00562-0.pdf.

Galla:2019:SIM

[3494] Michelle Galla, Kristina Wicke, and Mareike Fischer. Statistical inconsistency of maximum parsimony for k-tuple-site data. Bulletin of Mathematical Biology, 81(4):1173–1200, April 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/s11538-018-00552-2;   http://link.springer.
com/content/pdf/10.1007/s11538-018-00552-2.pdf.

Wertheim:2019:CVF

[3495] Kenneth Y. Wertheim and Tiina Roose. Can VEGFC form Turing patterns in the zebrafish embryo? Bulletin of Mathematical Biology, 81(4):1201–1237, April 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-018-00560-2;   http://link.springer.
com/content/pdf/10.1007/s11538-018-00560-2.pdf.

Barber:2019:TDF

[3496] Jared Barber and Luoding Zhu. Two-dimensional finite element model of breast cancer cell motion through a microfluidic channel. Bulletin of Mathematical Biology, 81(4):1238–1259, April 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-018-00557-x;   http://link.springer.com/content/pdf/
10.1007/s11538-018-00557-x.pdf.

Enciso:2019:ENC

[3497] German Enciso and Jinsu Kim. Embracing noise in chemical reaction networks. Bulletin of Mathematical Biology, 81(5):1261–1267, May 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-019-00575-3;   http://link.springer.com/content/pdf/
10.1007/s11538-019-00575-3.pdf.

Xu:2019:CDS

[3498] Bin Xu, Hye-Won Kang, and Alexandra Jilkine. Comparison of deterministic and stochastic regime in a model for Cdc42 oscillations in fission yeast. Bulletin of Mathematical Biology, 81(5):1268–1302, May 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-019-00573-5;   http://link.springer.com/content/pdf/
10.1007/s11538-019-00573-5.pdf.

Kang:2019:QSS

[3499] Hye-Won Kang, Wasiur R. KhudaBukhsh, Heinz Koeppl, and Grzegorz A. Rempała. Quasi-steady-state approximations derived from the stochastic model of enzyme kinetics. Bulletin of Mathematical Biology, 81(5):1303–1336, May 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-019-00574-4;   http://link.springer.
com/content/pdf/10.1007/s11538-019-00574-4.pdf.

Mai:2019:IDD

[3500] Ali Mai, Guowei Sun, and Lin Wang. Impacts of the dispersal delay on the stability of the coexistence equilibrium of a two-patch predator– prey model with random predator dispersal. Bulletin of Mathematical Biology, 81(5):1337–1351, May 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-018-00568-8;   http://link.springer.
com/content/pdf/10.1007/s11538-018-00568-8.pdf.

Asik:2019:DPG

[3501] Lale Asik and Angela Peace. Dynamics of a producer–grazer model incorporating the effects of phosphorus loading on grazer’s growth. Bulletin of Mathematical Biology, 81(5):1352–1368, May 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-018-00567-9;   http://link.springer.com/content/pdf/
10.1007/s11538-018-00567-9.pdf.

Marques:2019:ESP

[3502] Joice Chaves Marques, Horst Malchow, Luiz Alberto Díaz Rodrigues, and Diomar Cristina Mistro. The effects of stochasticity on pattern formation in a space- and time–discrete predator–prey system with strong Allee effect in the prey. Bulletin of Mathematical Biology, 81(5):1369–1393, May 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-018-00539-z;   http://link.springer.
com/content/pdf/10.1007/s11538-018-00539-z.pdf.

Pages:2019:MDA

[3503] Nathan Pages, Elías Vera-Sigüenza, John Rugis, Vivien Kirk, David I. Yule, and James Sneyd. A model of Ca^{2+} dynamics in an accurate reconstruction of parotid acinar cells. Bulletin of Mathematical Biology, 81(5):1394–1426, May 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-018-00563-z;   http://link.springer.
com/content/pdf/10.1007/s11538-018-00563-z.pdf.

Kitagawa:2019:MAT

[3504] Kosaku Kitagawa, Toshikazu Kuniya, Shinji Nakaoka, Yusuke Asai, Koichi Watashi, and Shingo Iwami. Mathematical analysis of a transformed ODE from a PDE multiscale model of hepatitis C virus infection. Bulletin of Mathematical Biology, 81(5):1427–1441, May 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-018-00564-y;   http://link.springer.com/content/pdf/
10.1007/s11538-018-00564-y.pdf.

Mortveit:2019:ASF

[3505] Henning S. Mortveit and Ryan D. Pederson. Attractor stability in finite asynchronous biological system models. Bulletin of Mathematical Biology, 81(5):1442–1460, May 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-018-00565-x;   http://link.springer.
com/content/pdf/10.1007/s11538-018-00565-x.pdf.

Dougoud:2019:EMP

[3506] Michaël Dougoud, Christian Mazza, Beat Schwaller, and László Pecze. Extending the mathematical palette for developmental pattern formation: Piebaldism. Bulletin of Mathematical Biology, 81(5):1461–1478, May 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-019-00569-1;   http://link.springer.com/content/pdf/
10.1007/s11538-019-00569-1.pdf.

Fan:2019:MRF

[3507] Gaoyang Fan and Paul C. Bressloff. Modeling the role of feedback in the adaptive response of bacterial quorum sensing. Bulletin of Mathematical Biology, 81(5):1479–1505, May 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-019-00570-8;   http://link.springer.
com/content/pdf/10.1007/s11538-019-00570-8.pdf.

Liu:2019:NMS

[3508] Yuewu Liu and Xiufen Zou. A new model system for exploring assembly mechanisms of the HIV-1 immature capsid in vivo. Bulletin of Mathematical Biology, 81(5):1506–1526, May 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-019-00571-7;   http://link.springer.
com/content/pdf/10.1007/s11538-019-00571-7.pdf.

Dickenstein:2019:MSR

[3509] Alicia Dickenstein, Mercedes Pérez Millán, Anne Shiu, and Xiaoxian Tang. Multistationarity in structured reaction networks. Bulletin of Mathematical Biology, 81(5):1527–1581, May 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.

springer.com/article/10.1007/s11538-019-00572-6;   http://link.
springer.com/content/pdf/10.1007/s11538-019-00572-6.pdf.

Song:2019:AES

[3510] Pengfei Song and Yanni Xiao. Analysis of an epidemic system with two response delays in media impact function. Bulletin of Mathematical Biology, 81(5):1582–1612, May 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-019-00586-0;   http://link.springer.
com/content/pdf/10.1007/s11538-019-00586-0.pdf.

Johnston:2019:CWR

[3511] Matthew D. Johnston and Evan Burton. Computing weakly reversible deficiency zero network translations using elementary flux modes. Bulletin of Mathematical Biology, 81(5):1613–1644, May 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-019-00579-z;   http://link.springer.com/content/pdf/
10.1007/s11538-019-00579-z.pdf.

Massey:2019:LDU

[3512] Susan Christine Massey, Andrea Hawkins-Daarud, Jill Gallaher, Alexander R. A. Anderson, Peter Canoll, and Kristin R. Swanson. Lesion dynamics under varying paracrine PDGF signaling in brain tissue. Bulletin of Mathematical Biology, 81(6):1645–1664, June 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-019-00587-z;   http://link.springer.com/content/pdf/
10.1007/s11538-019-00587-z.pdf.

Wang:2019:AST

[3513] Yuanshi Wang. Asymptotic state of a two-patch system with infinite diffusion. Bulletin of Mathematical Biology, 81(6):1665–1686, June 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-019-00582-4;   http://link.springer.com/content/pdf/
10.1007/s11538-019-00582-4.pdf.

Steindorf:2019:MSS

[3514] Vanessa Steindorf and Norberto Aníbal Maidana. Modeling the spatial spread of Chagas disease. Bulletin of Mathematical Biology, 81(6):1687– 1730, June 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-019-00581-5;   http://link.springer.com/content/pdf/
10.1007/s11538-019-00581-5.pdf.

Dangerfield:2019:RAE

[3515] Ciara E. Dangerfield, Martin Vyska, and Christopher A. Gilligan. Resource allocation for epidemic control across multiple subpopulations. Bulletin of Mathematical Biology, 81(6):1731–1759, June 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-019-00584-2;   http://link.springer.com/content/pdf/
10.1007/s11538-019-00584-2.pdf.

Warne:2019:UED

[3516] David J. Warne, Ruth E. Baker, and Matthew J. Simpson. Using experimental data and information criteria to guide model selection for reaction–diffusion problems in mathematical biology. Bulletin of Mathematical Biology, 81(6):1760–1804, June 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/s11538-019-00589-x;   http://link.springer.
com/content/pdf/10.1007/s11538-019-00589-x.pdf.

Hart:2019:GSA

[3517] J. L. Hart, P. A. Gremaud, and T. David. Global sensitivity analysis of high-dimensional neuroscience models: an example of neurovascular coupling. Bulletin of Mathematical Biology, 81(6):1805–1828, June 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-019-00578-0;   http://link.springer.com/content/pdf/
10.1007/s11538-019-00578-0.pdf.

Conradi:2019:EOM

[3518] Carsten Conradi, Maya Mincheva, and Anne Shiu. Emergence of oscillations in a mixed-mechanism phosphorylation system. Bulletin of Mathematical Biology, 81(6):1829–1852, June 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/s11538-019-00580-6;   http://link.springer.
com/content/pdf/10.1007/s11538-019-00580-6.pdf.

Renardy:2019:DDM

[3519] Marissa Renardy, Timothy Wessler, Silvia Blemker, Jennifer Linderman, Shayn Peirce, and Denise Kirschner. Data-driven model validation across dimensions. Bulletin of Mathematical Biology, 81(6):1853–1866, June 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-019-00590-4;   http://link.springer.com/content/pdf/
10.1007/s11538-019-00590-4.pdf.

Matabuena:2019:IVC

[3520] Marcos Matabuena and Rosana Rodríguez-López. An improved version of the classical banister model to predict changes in physical condition. Bulletin of Mathematical Biology, 81(6):1867–1884, June 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-019-00588-y;   http://link.springer.com/content/pdf/
10.1007/s11538-019-00588-y.pdf.

Hu:2019:MPC

[3521] Xiaochuan Hu, Guoyi Ke, and Sophia R.-J. Jang. Modeling pancreatic cancer dynamics with immunotherapy. Bulletin of Mathematical Biology, 81(6):1885–1915, June 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-019-00591-3;   http://link.springer.
com/content/pdf/10.1007/s11538-019-00591-3.pdf.

Arino:2019:NSP

[3522] Julien Arino, Nicolas Bajeux, and Steve Kirkland. Number of source patches required for population persistence in a source–sink metapopulation with explicit movement. Bulletin of Mathematical Biology, 81(6):1916–1942, June 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-019-00593-1;   http://link.springer.
com/content/pdf/10.1007/s11538-019-00593-1.pdf.

Asili:2019:MMA

[3523] Eda Asili, Shantia Yarahmadian, Hadi Khani, and Meisam Sharify. A mathematical model for amyloid-β aggregation in the presence of metal ions: a timescale analysis for the progress of Alzheimer disease. Bulletin of Mathematical Biology, 81(6):1943–1964, June 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-019-00583-3;   http://link.springer.com/content/pdf/
10.1007/s11538-019-00583-3.pdf.

Franssen:2019:MFM

[3524] Linnea C. Franssen, Tommaso Lorenzi, Andrew E. F. Burgess, and Mark A. J. Chaplain. A mathematical framework for modelling the metastatic spread of cancer. Bulletin of Mathematical Biology, 81(6):1965–2010, June 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-019-00597-x;   http://link.springer.com/content/pdf/
10.1007/s11538-019-00597-x.pdf.

Shaw:2019:MAI

[3525] Allison K. Shaw, Morganne Igoe, Alison G. Power, Nilsa A. Bosque-Pérez, and Angela Peace. Modeling approach influences dynamics of a vector-borne pathogen system. Bulletin of Mathematical Biology, 81(6):2011–2028, June 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-019-00595-z;   http://link.springer.
com/content/pdf/10.1007/s11538-019-00595-z.pdf.

Barradas:2019:BBD

[3526] Ignacio Barradas and Virgilio Vázquez. Backward bifurcation as a desirable phenomenon: Increased fecundity through infection. Bulletin of Mathematical Biology, 81(6):2029–2050, June 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-019-00604-1;   http://link.springer.com/content/pdf/
10.1007/s11538-019-00604-1.pdf.

Laubenbacher:2019:TDA

[3527] Reinhard Laubenbacher and Alan Hastings. Topological data analysis. Bulletin of Mathematical Biology, 81(7):2051, July 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
019-00610-3;   http://link.springer.com/content/pdf/10.1007/
s11538-019-00610-3.pdf.

Cho:2019:ESS

[3528] Min Ho Cho, Amir Asiaee, and Sebastian Kurtek. Elastic statistical shape analysis of biological structures with case studies: a tutorial. Bulletin of Mathematical Biology, 81(7):2052–2073, July 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-019-00609-w;   http://link.springer.com/content/pdf/
10.1007/s11538-019-00609-w.pdf.

Memoli:2019:PPH

[3529] Facundo Mémoli and Kritika Singhal. A primer on persistent homology of finite metric spaces. Bulletin of Mathematical Biology, 81(7):2074–2116, July 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-019-00614-z;   http://link.springer.
com/content/pdf/10.1007/s11538-019-00614-z.pdf.

Kareva:2019:NSB

[3530] Irina Kareva and Georgy Karev. Natural selection between two games with applications to game theoretical models of cancer. Bulletin of Mathematical Biology, 81(7):2117–2132, July 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/s11538-019-00592-2;   http://link.springer.
com/content/pdf/10.1007/s11538-019-00592-2.pdf.

Jeronimo:2019:IFS

[3531] Gabriela Jeronimo, Mercedes Pérez Millán, and Pablo Solernó. Identifiability from a few species for a class of biochemical reaction networks. Bulletin of Mathematical Biology, 81(7):2133–2175, July 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-019-00594-0;   http://link.springer.com/content/pdf/
10.1007/s11538-019-00594-0.pdf.

Shuttleworth:2019:MMF

[3532] Robyn Shuttleworth and Dumitru Trucu. Multiscale modelling of fibres dynamics and cell adhesion within moving boundary cancer invasion. Bulletin of Mathematical Biology, 81(7):2176–2219, July 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-019-00598-w;   http://link.springer.com/content/pdf/
10.1007/s11538-019-00598-w.pdf.

LaiDeOliveira:2019:MUN

[3533] Alexander Lai De Oliveira and Benjamin J. Binder. Modeling uniaxial nonuniform cell proliferation. Bulletin of Mathematical Biology, 81(7):2220–2238, July 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-019-00601-4;   http://link.springer.
com/content/pdf/10.1007/s11538-019-00601-4.pdf.

Blazej:2019:QGC

[3534] P. Błażej, E. Fimmel, and M. Gumbel. The quality of genetic code models in terms of their robustness against point mutations. Bulletin of Mathematical Biology, 81(7):2239–2257, July 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-019-00603-2;   http://link.springer.com/content/pdf/
10.1007/s11538-019-00603-2.pdf.

Battista:2019:MPO

[3535] Nicholas A. Battista, Leigh B. Pearcy, and W. Christopher Strickland. Modeling the prescription opioid epidemic. Bulletin of Mathematical Biology, 81(7):2258–2289, July 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-019-00605-0;   http://link.springer.
com/content/pdf/10.1007/s11538-019-00605-0.pdf.

Eigentler:2019:MCM

[3536] Lukas Eigentler and Jonathan A. Sherratt. Metastability as a coexistence mechanism in a model for dryland vegetation patterns. Bulletin of Mathematical Biology, 81(7):2290–2322, July 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-019-00606-z;   http://link.springer.com/content/pdf/
10.1007/s11538-019-00606-z.pdf.

Bouchnita:2019:MME

[3537] Anass Bouchnita, Stefan Hellander, and Andreas Hellander. A 3D multiscale model to explore the role of EGFR overexpression in tumourigenesis. Bulletin of Mathematical Biology, 81(7):2323–2344, July 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-019-00607-y;   http://link.springer.com/content/pdf/
10.1007/s11538-019-00607-y.pdf.

Fassoni:2019:CCC

[3538] Artur César Fassoni, Ingo Roeder, and Ingmar Glauche. To cure or not to cure: Consequences of immunological interactions in CML treatment. Bulletin of Mathematical Biology, 81(7):2345–2395, July 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-019-00608-x;   http://link.springer.com/content/pdf/
10.1007/s11538-019-00608-x.pdf.

Zhao:2019:SMO

[3539] Jiantao Zhao and Jianjun Paul Tian. Spatial model for oncolytic virotherapy with lytic cycle delay. Bulletin of Mathematical Biology, 81(7):2396–2427, July 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-019-00611-2;   http://link.springer.
com/content/pdf/10.1007/s11538-019-00611-2.pdf.

Sadeghimanesh:2019:MSN

[3540] AmirHosein Sadeghimanesh and Elisenda Feliu. The multistationarity structure of networks with intermediates and a binomial core network. Bulletin of Mathematical Biology, 81(7):2428–2462, July 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-019-00612-1;   http://link.springer.com/content/pdf/
10.1007/s11538-019-00612-1.pdf.

Barabanschikov:2019:MMT

[3541] Alexander Barabanschikov and Jeremy Gunawardena. Monostationarity and multistationarity in tree networks of Goldbeter–Koshland loops. Bulletin of Mathematical Biology, 81(7):2463–2509, July 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-019-00615-y;   http://link.springer.com/content/pdf/
10.1007/s11538-019-00615-y.pdf.

Vitale:2019:ISP

[3542] Caterina Vitale and Alex Best. The impact of selective predation on host–parasite SIS dynamics. Bulletin of Mathematical Biology, 81(7):2510–2528, July 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-019-00616-x;   http://link.springer.
com/content/pdf/10.1007/s11538-019-00616-x.pdf.

Chakraborty:2019:NUM

[3543] Biman Chakraborty, Amiya Ranjan Bhowmick, Joydev Chattopadhyay, and Sabyasachi Bhattacharya. A novel unification method to characterize a broad class of growth curve models using relative growth rate. Bulletin of Mathematical Biology, 81(7):2529–2552, July 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-019-00617-w;   http://link.springer.com/content/pdf/
10.1007/s11538-019-00617-w.pdf.

Moore:2019:DCC

[3544] James R. Moore, Hasan Ahmed, Don McGuire, Rama Akondy, Rafi Ahmed, and Rustom Antia. Dependence of CD8 T cell response upon antigen load during primary infection. Bulletin of Mathematical Biology, 81(7):2553–2568, July 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-019-00618-9;   http://link.springer.
com/content/pdf/10.1007/s11538-019-00618-9.pdf.

Wang:2019:MPR

[3545] Xiunan Wang and Xingfu Zou. Modeling the potential role of engineered symbiotic bacteria in malaria control. Bulletin of Mathematical Biology, 81(7):2569–2595, July 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-019-00619-8;   http://link.springer.
com/content/pdf/10.1007/s11538-019-00619-8.pdf.

Li:2019:TIL

[3546] Dan Li and Hui Wan. The threshold infection level for Wolbachia invasion in a two-sex mosquito population model. Bulletin of Mathematical Biology, 81(7):2596–2624, July 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-019-00620-1;   http://link.springer.
com/content/pdf/10.1007/s11538-019-00620-1.pdf.

Cao:2019:SSB

[3547] Yang Cao, Petzold Linda, and Effrosyni Seitaridou. Stochastic simulation of biochemical systems: In memory of Dan T. Gillespie’s contributions. Bulletin of Mathematical Biology, 81(8):2819–2821, August 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-019-00633-w;   http://link.springer.com/content/pdf/
10.1007/s11538-019-00633-w.pdf.

Suderman:2019:GGD

[3548] Ryan Suderman, Eshan D. Mitra, Yen Ting Lin, Keesha E. Erickson, Song Feng, and William S. Hlavacek. Generalizing Gillespie’s direct method to enable network–free simulations. Bulletin of Mathematical Biology, 81(8):2822–2848, August 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-018-0418-2;
 http://link.springer.com/content/pdf/10.1007/s11538-018-0418-
2.pdf.

Rathinam:2019:SDE

[3549] Muruhan Rathinam and Yuriy Sverchkov. Stochastic dynamics of eukaryotic flagellar growth. Bulletin of Mathematical Biology, 81(8): 2849–2872, August 2019. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-018-0427-1;   http://link.springer.com/content/
pdf/10.1007/s11538-018-0427-1.pdf.

Lotstedt:2019:LNA

[3550] Per Lötstedt. The linear noise approximation for spatially dependent biochemical networks. Bulletin of Mathematical Biology, 81(8): 2873–2901, August 2019. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-018-0428-0;   http://link.springer.com/content/
pdf/10.1007/s11538-018-0428-0.pdf.

Anderson:2019:LVC

[3551] David F. Anderson and Chaojie Yuan. Low variance couplings for stochastic models of intracellular processes with time-dependent rate functions. Bulletin of Mathematical Biology, 81(8):2902–2930, August 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0430-6;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0430-6.pdf.

Beentjes:2019:QMC

[3552] Casper H. L. Beentjes and Ruth E. Baker. Quasi-Monte Carlo methods applied to tau-leaping in stochastic biological systems. Bulletin of Mathematical Biology, 81(8):2931–2959, August 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0442-2;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0442-2.pdf.

Smith:2019:SSI

[3553] Stephen Smith and Ramon Grima. Spatial stochastic intracellular kinetics: a review of modelling approaches. Bulletin of Mathematical Biology, 81(8):2960–3009, August 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-018-0443-1;
 http://link.springer.com/content/pdf/10.1007/s11538-018-0443-
1.pdf.

Engblom:2019:SSP

[3554] Stefan Engblom. Stochastic simulation of pattern formation in growing tissue: a multilevel approach. Bulletin of Mathematical Biology, 81(8): 3010–3023, August 2019. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-018-0454-y;   http://link.springer.com/content/
pdf/10.1007/s11538-018-0454-y.pdf.

Chen:2019:AAH

[3555] Minghan Chen, Shuo Wang, and Yang Cao. Accuracy analysis of hybrid stochastic simulation algorithm on linear chain reaction systems. Bulletin of Mathematical Biology, 81(8):3024–3052, August 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0461-z;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0461-z.pdf.

Thanh:2019:CCR

[3556] Vo Hong Thanh. A critical comparison of rejection-based algorithms for simulation of large biochemical reaction networks. Bulletin of Mathematical Biology, 81(8):3053–3073, August 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0462-y;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0462-y.pdf.

Lipkova:2019:LAA

[3557] Jana Lipková, Georgios Arampatzis, Philippe Chatelain, Bjoern Menze, and Petros Koumoutsakos. S-leaping: an adaptive, accelerated stochastic simulation algorithm, bridging τ-leaping and Rleaping. Bulletin of Mathematical Biology, 81(8):3074–3096, August 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0464-9;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0464-9.pdf.

Roh:2019:DDM

[3558] Min K. Roh. Data-driven method for efficient characterization of rare event probabilities in biochemical systems. Bulletin of Mathematical Biology, 81(8):3097–3120, August 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-018-0509-0;
 http://link.springer.com/content/pdf/10.1007/s11538-018-0509-
0.pdf.

Gupta:2019:SAM

[3559] Ankit Gupta and Mustafa Khammash. Sensitivity analysis for multiscale stochastic reaction networks using hybrid approximations. Bulletin of Mathematical Biology, 81(8):3121–3158, August 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0521-4;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0521-4.pdf.

Maginnis:2019:EVR

[3560] P. A. Maginnis, M. West, and G. E. Dullerud. Exact variancereduced simulation of lattice continuous–time Markov chains with applications in reaction networks. Bulletin of Mathematical Biology, 81 (8):3159–3184, August 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-019-00576-2;   http://link.springer.
com/content/pdf/10.1007/s11538-019-00576-2.pdf.

Kang:2019:MSR

[3561] Hye-Won Kang and Radek Erban. Multiscale stochastic reaction– diffusion algorithms combining Markov chain models with stochastic partial differential equations. Bulletin of Mathematical Biology, 81 (8):3185–3213, August 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-019-00613-0;   http://link.springer.
com/content/pdf/10.1007/s11538-019-00613-0.pdf.

Alber:2019:I

[3562] Mark Alber, Christophe Godin, Philip K. Maini, Roeland Merks, and Eric Mjolsness. Introduction. Bulletin of Mathematical Biology, 81 (8):3214–3218, August 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-019-00649-2;   http://link.springer.
com/content/pdf/10.1007/s11538-019-00649-2.pdf.

Erlich:2019:HSS

[3563] Alexander Erlich, Derek E. Moulton, and Alain Goriely. Are homeostatic states stable? Dynamical stability in morphoelasticity. Bulletin of Mathematical Biology, 81(8):3219–3244, August 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0502-7;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0502-7.pdf.

Banwarth-Kuhn:2019:CBM

[3564] Mikahl Banwarth-Kuhn, Ali Nematbakhsh, Kevin W. Rodriguez, Stephen Snipes, Carolyn G. Rasmussen, G. Venugopala Reddy, and Mark Alber. Cell-based model of the generation and maintenance of the shape and structure of the multilayered shoot apical meristem of Arabidopsis thaliana. Bulletin of Mathematical Biology, 81 (8):3245–3281, August 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-018-00547-z;   http://link.springer.
com/content/pdf/10.1007/s11538-018-00547-z.pdf.

Jiang:2019:DIF

[3565] J. Jiang, K. Garikipati, and S. Rudraraju. A diffuse interface framework for modeling the evolution of multi-cell aggregates as a soft packing problem driven by the growth and division of cells. Bulletin of Mathematical Biology, 81(8):3282–3300, August 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-019-00577-1;   http://link.springer.
com/content/pdf/10.1007/s11538-019-00577-1.pdf.

Hassan:2019:IBP

[3566] Abdel-Rahman Hassan, Thomas Biel, David M. Umulis, and Taeyoon Kim. Interplay between the persistent random walk and the contact inhibition of locomotion leads to collective cell behaviors. Bulletin of Mathematical Biology, 81(8):3301–3321, August 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-019-00585-1;   http://link.springer.com/content/pdf/
10.1007/s11538-019-00585-1.pdf.

Moore:2019:WCM

[3567] Helen Moore and Richard Allen. What can mathematics do for drug development? Bulletin of Mathematical Biology, 81(9):3421–3424, September 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-019-00632-x;   http://link.springer.com/content/pdf/
10.1007/s11538-019-00632-x.pdf.

Allen:2019:PRM

[3568] Richard Allen and Helen Moore. Perspectives on the role of mathematics in drug discovery and development. Bulletin of Mathematical Biology, 81(9):3425–3435, September 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-018-00556-y;   http://link.springer.
com/content/pdf/10.1007/s11538-018-00556-y.pdf.

Wu:2019:ASE

[3569] Xiaotian Wu, Fahima Nekka, and Jun Li. Analytical solution and exposure analysis of a pharmacokinetic model with simultaneous elimination pathways and endogenous production: The case of multiple dosing administration. Bulletin of Mathematical Biology, 81(9): 3436–3459, September 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-019-00651-8;   http://link.springer.
com/content/pdf/10.1007/s11538-019-00651-8.pdf.

Saltzman:2019:MEM

[3570] Jeffrey Saltzman and Claus Bendtsen. Modeling the effect of mucin binding in the gut on drug delivery. Bulletin of Mathematical Biology, 81(9):3460–3476, September 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-018-0417-3;   http://link.springer.com/
content/pdf/10.1007/s11538-018-0417-3.pdf.

Vendel:2019:IPL

[3571] E. Vendel, V. Rottschäfer, and E. C. M. de Lange. Improving the prediction of local drug distribution profiles in the brain with a new 2D mathematical model. Bulletin of Mathematical Biology, 81(9):3477– 3507, September 2019. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-018-0469-4;   http://link.springer.com/content/
pdf/10.1007/s11538-018-0469-4.pdf.

Naganawa:2019:QPI

[3572] Mika Naganawa, Jean-Dominique Gallezot, Samantha Rossano, and Richard E. Carson. Quantitative PET imaging in drug development: Estimation of target occupancy. Bulletin of Mathematical Biology, 81(9):3508–3541, September 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-017-0374-2;   http://link.springer.com/
content/pdf/10.1007/s11538-017-0374-2.pdf.

White:2019:LBD

[3573] Carla White and Lloyd J. Bridge. Ligand binding dynamics for predimerised G protein-coupled receptor homodimers: Linear models and analytical solutions. Bulletin of Mathematical Biology, 81(9):3542– 3574, September 2019. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-017-0387-x;   http://link.springer.com/content/
pdf/10.1007/s11538-017-0387-x.pdf.

Lemaire:2019:DBC

[3574] Vincent Lemaire and David R. Cox. Dynamics of bone cell interactions and differential responses to PTH and antibody-based therapies. Bulletin of Mathematical Biology, 81(9):3575–3622, September 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0533-0;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0533-0.pdf.

Karolak:2019:MSA

[3575] Aleksandra Karolak and Katarzyna A. Rejniak. Micropharmacology: An in silico approach for assessing drug efficacy within a tumor tissue. Bulletin of Mathematical Biology, 81(9):3623–3641, September 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0402-x;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0402-x.pdf.

Gnacadja:2019:IP

[3576] Gilles Gnacadja. An invitation to pharmacostatics. Bulletin of Mathematical Biology, 81(9):3642–3654, September 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-017-0369-z;
 http://link.springer.com/content/pdf/10.1007/s11538-017-0369-
z.pdf.

Duron:2019:VBC

[3577] Christina Durón, Yuan Pan, David H. Gutmann, Johanna Hardin, and Ami Radunskaya. Variability of betweenness centrality and its effect on identifying essential genes. Bulletin of Mathematical Biology, 81(9):3655– 3673, September 2019. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-018-0526-z;   http://link.springer.com/content/
pdf/10.1007/s11538-018-0526-z.pdf.

Reinharz:2019:PEM

[3578] Vladimir Reinharz, Alexander Churkin, Stephanie Lewkiewicz, Harel Dahari, and Danny Barash. A parameter estimation method for multiscale models of hepatitis C virus dynamics. Bulletin of Mathematical Biology, 81(10):3675–3721, October 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-019-00644-7;   http://link.springer.
com/content/pdf/10.1007/s11538-019-00644-7.pdf.

Brady:2019:MMC

[3579] Renee Brady and Heiko Enderling. Mathematical models of cancer: When to predict novel therapies, and when not to. Bulletin of Mathematical Biology, 81(10):3722–3731, October 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/s11538-019-00640-x;   http://link.springer.
com/content/pdf/10.1007/s11538-019-00640-x.pdf.

Callahan:2019:ALS

[3580] Jason Callahan, Eric Eager, Richard Rebarber, Eva Strawbridge, and Shenglan Yuan. Analysis of a length-structured density-dependent model for fish. Bulletin of Mathematical Biology, 81(10):3732–3753, October 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-019-00648-3;   http://link.springer.com/content/pdf/
10.1007/s11538-019-00648-3.pdf. See correction [3626].

Kohnke:2019:IDI

[3581] M. C. Köhnke. Invasion dynamics in an intraguild predation system with predator-induced defense. Bulletin of Mathematical Biology, 81 (10):3754–3777, October 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-019-00655-4;   http://link.springer.
com/content/pdf/10.1007/s11538-019-00655-4.pdf.

Duncan:2019:MSH

[3582] Simon J. Duncan, Keith R. Daly, Daniel M. McKay Fletcher, Siul Ruiz, Paul Sweeney, and Tiina Roose. Multiple scale homogenisation of nutrient movement and crop growth in partially saturated soil. Bulletin of Mathematical Biology, 81(10):3778–3802, October 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-019-00656-3;   http://link.springer.com/content/pdf/
10.1007/s11538-019-00656-3.pdf.

Davis:2019:LDA

[3583] Alexander L. Davis, Alexander P. Hoover, and Laura A. Miller. Lift and drag acting on the Shell of the American horseshoe crab (Limulus polyphemus). Bulletin of Mathematical Biology, 81(10):3803–3822, October 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-019-00657-2;   http://link.springer.com/content/pdf/
10.1007/s11538-019-00657-2.pdf.

Murakami:2019:RTC

[3584] Yukihiro Murakami, Leo van Iersel, Remie Janssen, Mark Jones, and Vincent Moulton. Reconstructing tree–child networks from reticulate– edge-deleted subnetworks. Bulletin of Mathematical Biology, 81(10): 3823–3863, October 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-019-00641-w;   http://link.springer.
com/content/pdf/10.1007/s11538-019-00641-w.pdf.

Fassoni:2019:RAC

[3585] Artur César Fassoni and Denis Carvalho Braga. Resilience analysis for competing populations. Bulletin of Mathematical Biology, 81 (10):3864–3888, October 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-019-00660-7;   http://link.springer.
com/content/pdf/10.1007/s11538-019-00660-7.pdf.

Wang:2019:HSH

[3586] Xiaoying Wang, Messoud Efendiev, and Frithjof Lutscher. How spatial heterogeneity affects transient behavior in reaction–diffusion systems for ecological interactions? Bulletin of Mathematical Biology, 81(10):3889–3917, October 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-019-00659-0;   http://link.springer.
com/content/pdf/10.1007/s11538-019-00659-0.pdf.

Plank:2019:LES

[3587] Michael J. Plank, Nick Stringer, Shona L. Lamoureaux, Graeme W. Bourdôt, and Alex James. Limiting effect of self-shading on the height of Tradescantia fluminensis mats. Bulletin of Mathematical Biology, 81(10):3918–3932, October 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-019-00631-y;   http://link.springer.
com/content/pdf/10.1007/s11538-019-00631-y.pdf.

Cronin:2019:EPM

[3588] James T. Cronin, Jerome Goddard, and Ratnasingham Shivaji. Effects of patch-matrix composition and individual movement response on population persistence at the patch level. Bulletin of Mathematical Biology, 81(10):3933–3975, October 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-019-00634-9;   http://link.springer.
com/content/pdf/10.1007/s11538-019-00634-9.pdf.

Houston:2019:FMM

[3589] Matthew T. Houston and Juan B. Gutierrez. The FRiND model: A mathematical model for representing macrophage plasticity in muscular dystrophy pathogenesis. Bulletin of Mathematical Biology, 81 (10):3976–3997, October 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-019-00635-8;   http://link.springer.
com/content/pdf/10.1007/s11538-019-00635-8.pdf.

Contreras:2019:MME

[3590] Carlos Contreras, Gustavo Carrero, and Gerda de Vries. A mathematical model for the effect of low-dose radiation on the G2/M transition. Bulletin of Mathematical Biology, 81(10):3998–4021, October 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-019-00645-6;   http://link.springer.com/content/pdf/
10.1007/s11538-019-00645-6.pdf.

Thon:2019:SRQ

[3591] Moritz P. Thon, Mary R. Myerscough, and Michael W. Gee. A spatially resolved and quantitative model of early atherosclerosis. Bulletin of Mathematical Biology, 81(10):4022–4068, October 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-019-00646-5;   http://link.springer.com/content/pdf/
10.1007/s11538-019-00646-5.pdf.

Jensen:2019:AMI

[3592] Melanie A. Jensen, Ying-Ying Wang, Samuel K. Lai, M. Gregory Forest, and Scott A. McKinley. Antibody-mediated immobilization of virions in mucus. Bulletin of Mathematical Biology, 81(10):4069–4099, October 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-019-00653-6;   http://link.springer.com/content/pdf/
10.1007/s11538-019-00653-6.pdf.

Camacho:2019:MPH

[3593] Ariel Camacho, Fernando Saldaña, Ignacio Barradas, and Silvia Jerez. Modeling public health campaigns for sexually transmitted infections via optimal and feedback control. Bulletin of Mathematical Biology, 81(10):4100–4123, October 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-019-00642-9;   http://link.springer.
com/content/pdf/10.1007/s11538-019-00642-9.pdf.

Chizhov:2019:CBR

[3594] Anton Chizhov, Fabien Campillo, Mathieu Desroches, Antoni Guillamon, and Serafim Rodrigues. Conductance-based refractory density approach for a population of bursting neurons. Bulletin of Mathematical Biology, 81(10):4124–4143, October 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-019-00643-8;   http://link.springer.
com/content/pdf/10.1007/s11538-019-00643-8.pdf.

Valle:2019:TCA

[3595] Paul A. Valle, Luis N. Coria, and Yolocuauhtli Salazar. Tumor clearance analysis on a cancer chemo–immunotherapy mathematical model. Bulletin of Mathematical Biology, 81(10):4144–4173, October 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-019-00636-7;   http://link.springer.com/content/pdf/
10.1007/s11538-019-00636-7.pdf.

Conradi:2019:MST

[3596] Carsten Conradi, Alexandru Iosif, and Thomas Kahle. Multistationarity in the space of total concentrations for systems that admit a monomial parametrization. Bulletin of Mathematical Biology, 81 (10):4174–4209, October 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-019-00639-4;   http://link.springer.
com/content/pdf/10.1007/s11538-019-00639-4.pdf.

Smirnova:2019:SPE

[3597] Alexandra Smirnova, Benjamin Sirb, and Gerardo Chowell. On stable parameter estimation and forecasting in epidemiology by the Levenberg– Marquardt algorithm with Broyden’s rank-one updates for the Jacobian operator. Bulletin of Mathematical Biology, 81(10):4210–4232, October 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-019-00650-9;   http://link.springer.com/content/pdf/
10.1007/s11538-019-00650-9.pdf.

Arino:2019:CTM

[3598] Julien Arino and James Watmough. Current trends in mathematical epidemiology. Bulletin of Mathematical Biology, 81(11):4311–4312, November 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-019-00672-3;   http://link.springer.com/content/pdf/
10.1007/s11538-019-00672-3.pdf.

Milwid:2019:HSV

[3599] Rachael M. Milwid, Federico Frascoli, Marc Steben, and Jane M. Heffernan. HPV screening and vaccination strategies in an unscreened population: A mathematical modeling study. Bulletin of Mathematical Biology, 81(11):4313–4342, November 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL

http://link.springer.com/article/10.1007/s11538-018-0425-3;
 http://link.springer.com/content/pdf/10.1007/s11538-018-0425-
3.pdf.

Smirnova:2019:FET

[3600] Alexandra Smirnova, Linda deCamp, and Gerardo Chowell. Forecasting epidemics through nonparametric estimation of time–dependent transmission rates using the SEIR model. Bulletin of Mathematical Biology, 81(11):4343–4365, November 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-017-0284-3;   http://link.springer.com/
content/pdf/10.1007/s11538-017-0284-3.pdf.

Kamgang:2019:AMC

[3601] Jean Claude Kamgang and Christopher Penniman Thron. Analysis of malaria control measures’ effectiveness using multistage vector model. Bulletin of Mathematical Biology, 81(11):4366–4411, November 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-019-00637-6;   http://link.springer.com/content/pdf/
10.1007/s11538-019-00637-6.pdf.

vandenDriessche:2019:DEV

[3602] P. van den Driessche and Abdul-Aziz Yakubu. Disease extinction versus persistence in discrete-time epidemic models. Bulletin of Mathematical Biology, 81(11):4412–4446, November 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0426-2;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0426-2.pdf.

Rivera:2019:VPA

[3603] Rocio Caja Rivera and Ignacio Barradas. Vector preference annihilates backward bifurcation and reduces endemicity. Bulletin of Mathematical Biology, 81(11):4447–4469, November 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.

com/article/10.1007/s11538-018-00561-1;   http://link.springer.
com/content/pdf/10.1007/s11538-018-00561-1.pdf.

Malagon:2019:CHE

[3604] Camilo Sanabria Malagón and Esteban Vargas Bernal. Competent hosts and endemicity of multi-host vector-borne diseases. Bulletin of Mathematical Biology, 81(11):4470–4483, November 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-00543-3;   http://link.springer.com/content/pdf/10.1007/
s11538-018-00543-3.pdf.

Mata:2019:RCD

[3605] May Anne Mata, Priscilla Greenwood, and Rebecca Tyson. The relative contribution of direct and environmental transmission routes in stochastic avian flu epidemic recurrence: An approximate analysis. Bulletin of Mathematical Biology, 81(11):4484–4517, November 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0414-6;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0414-6.pdf.

Harvim:2019:TDC

[3606] Prince Harvim, Hong Zhang, Paul Georgescu, and Lai Zhang. Transmission dynamics and control mechanisms of vector-borne diseases with active and passive movements between urban and satellite cities. Bulletin of Mathematical Biology, 81(11):4518–4563, November 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-019-00671-4;   http://link.springer.com/content/pdf/
10.1007/s11538-019-00671-4.pdf.

Woldegerima:2019:IRD

[3607] Woldegebriel A. Woldegerima, Miranda I. Teboh-Ewungkem, and Gideon A. Ngwa. The impact of recruitment on the dynamics of an immune-suppressed within-human–host model of the Plasmodium falciparum parasite. Bulletin of Mathematical Biology, 81(11):4564– 4619, November 2019. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-018-0436-0;   http://link.springer.com/content/
pdf/10.1007/s11538-018-0436-0.pdf.

Morozov:2019:MBE

[3608] Andrew Morozov. Modelling biological evolution: Developing novel approaches. Bulletin of Mathematical Biology, 81(11):4620–4624, November 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-019-00670-5;   http://link.springer.com/content/pdf/
10.1007/s11538-019-00670-5.pdf.

Chalub:2019:FPP

[3609] Fabio A. C. C. Chalub and Max O. Souza. From fixation probabilities to d-player games: An inverse problem in evolutionary dynamics. Bulletin of Mathematical Biology, 81(11):4625–4642, November 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-018-00566-w;   http://link.springer.com/content/pdf/
10.1007/s11538-018-00566-w.pdf.

Broom:2019:GSD

[3610] Mark Broom, Karan Pattni, and Jan Rychtář. Generalized social dilemmas: The evolution of cooperation in populations with variable group size. Bulletin of Mathematical Biology, 81(11):4643–4674, November 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-018-00545-1;   http://link.springer.com/content/pdf/
10.1007/s11538-018-00545-1.pdf.

Kuzenkov:2019:TCM

[3611] Oleg Kuzenkov and Andrew Morozov. Towards the construction of a mathematically rigorous framework for the modelling of evolutionary fitness. Bulletin of Mathematical Biology, 81(11):4675–4700, November 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-019-00602-3;   http://link.springer.com/content/pdf/
10.1007/s11538-019-00602-3.pdf.

Sandhu:2019:REO

[3612] Simran Kaur Sandhu, Andrew Morozov, and Oleg Kuzenkov. Revealing evolutionarily optimal strategies in self-reproducing systems via a new computational approach. Bulletin of Mathematical Biology, 81 (11):4701–4725, November 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-019-00663-4;   http://link.springer.
com/content/pdf/10.1007/s11538-019-00663-4.pdf.

Dissanayake:2019:GSS

[3613] Chandani Dissanayake, Lourdes Juan, Kevin R. Long, Angela Peace, and Md Masud Rana. Genotypic selection in spatially heterogeneous producer–grazer systems subject to stoichiometric constraints. Bulletin of Mathematical Biology, 81(11):4726–4742, November 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-00559-9;   http://link.springer.com/content/pdf/10.1007/
s11538-018-00559-9.pdf.

Reisch:2019:MCT

[3614] Cordula Reisch and Dirk Langemann. Modeling the chronification tendency of liver infections as evolutionary advantage. Bulletin of Mathematical Biology, 81(11):4743–4760, November 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
019-00596-y;   http://link.springer.com/content/pdf/10.1007/
s11538-019-00596-y.pdf.

Gidoin:2019:RET

[3615] Cindy Gidoin and Stephan Peischl. Range expansion theories could shed light on the spatial structure of intra-tumour heterogeneity. Bulletin of Mathematical Biology, 81(11):4761–4777, November 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-018-00540-6;   http://link.springer.com/content/pdf/
10.1007/s11538-018-00540-6.pdf.

Vitale:2019:ESP

[3616] Caterina Vitale and Eva Kisdi. Evolutionary suicide of prey: Matsuda and Abrams’ model revisited. Bulletin of Mathematical Biology, 81(11): 4778–4802, November 2019. CODEN BMTBAP. ISSN 0092-8240 (print),

1522-9602 (electronic). URL http://link.springer.com/article/10.
1007/s11538-018-0472-9;   http://link.springer.com/content/
pdf/10.1007/s11538-018-0472-9.pdf.

Cropp:2019:SEE

[3617] Roger Cropp and John Norbury. Simulating eco-evolutionary processes in an obligate pollination model with a genetic algorithm. Bulletin of Mathematical Biology, 81(11):4803–4820, November 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
018-0508-1;   http://link.springer.com/content/pdf/10.1007/
s11538-018-0508-1.pdf.

Cao:2019:ADR

[3618] Hui Cao, Dongxue Yan, Suxia Zhang, and Xiaoqin Wang. Analysis of dynamics of recurrent epidemics: Periodic or non-periodic. Bulletin of Mathematical Biology, 81(12):4889–4907, December 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
019-00638-5;   http://link.springer.com/content/pdf/10.1007/
s11538-019-00638-5.pdf.

Magal:2019:EMH

[3619] P. Magal, G. F. Webb, and Yixiang Wu. An environmental model of honey bee colony collapse due to pesticide contamination. Bulletin of Mathematical Biology, 81(12):4908–4931, December 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-019-00662-5;   http://link.springer.com/content/pdf/
10.1007/s11538-019-00662-5.pdf.

Peace:2019:CFS

[3620] Angela Peace and Hao Wang. Compensatory foraging in stoichiometric producer–grazer models. Bulletin of Mathematical Biology, 81 (12):4932–4950, December 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-019-00665-2;   http://link.springer.
com/content/pdf/10.1007/s11538-019-00665-2.pdf.

Borisov:2019:MHI

[3621] Milen Borisov, Gabriel Dimitriu, and Peter Rashkov. Modelling the host immune response to mature and immature dengue viruses. Bulletin of Mathematical Biology, 81(12):4951–4976, December 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (elec-

tronic). URL http://link.springer.com/article/10.1007/s11538-
019-00664-3;   http://link.springer.com/content/pdf/10.1007/
s11538-019-00664-3.pdf.

Aldila:2019:PDM

[3622] Dipo Aldila and Hiromi Seno. A population dynamics model of mosquitoborne disease transmission, focusing on mosquitoes’ biased distribution and mosquito repellent use. Bulletin of Mathematical Biology, 81(12):4977–5008, December 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-019-00666-1;   http://link.springer.
com/content/pdf/10.1007/s11538-019-00666-1.pdf.

Zhou:2019:TIC

[3623] Zhi Zhou and Robert A. Van Gorder. Turing instability and colony formation in spatially extended Rosenzweig–MacArthur predator–prey models with allochthonous resources. Bulletin of Mathematical Biology, 81(12):5009–5053, December 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL http://link.springer.com/

article/10.1007/s11538-019-00667-0;   http://link.springer.
com/content/pdf/10.1007/s11538-019-00667-0.pdf.

Girardin:2019:CMI

[3624] Léo Girardin, Vincent Calvez, and Florence Débarre. Catch me if you can: A spatial model for a brake-driven gene drive reversal. Bulletin of Mathematical Biology, 81(12):5054–5088, December 2019. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL http://link.springer.com/article/10.1007/
s11538-019-00668-z;   http://link.springer.com/content/pdf/
10.1007/s11538-019-00668-z.pdf.

Leander:2020:CMC

[3625] R. Leander and A. Friedman. Correction to: Modulation of the cAMP Response by Gα_{i} and Gβγ: a Computational Study of G Protein Signaling in Immune Cells. Bulletin of Mathematical Biology, 82(6):??, June 2020. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602

(electronic). URL https://link.springer.com/article/10.1007/
s11538-020-00753-8. See [2933].

Callahan:2021:CAL

[3626] Jason Callahan, Eric Eager, and Shenglan Yuan. Correction to: Analysis of a Length-Structured Density-Dependent Model for Fish. Bulletin of Mathematical Biology, 83(8):??, August 2021. CODEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic). URL https://link.

springer.com/article/10.1007/s11538-021-00916-1. See [3580].

Turing:1952:CBM

[3627] A. M. Turing. The chemical basis of morphogenesis. Philosophical transactions of the Royal Society of London Series B, Biological sciences, B 237(641):37–72, August 14, 1952. CODEN PTRBAE. ISSN 0962-8436 (print), 1471-2970 (electronic). URL http://turing.ecs.soton.ac. uk/browse.php/B/22. Reprinted in [3628, pp. 153–197] and [1069].

Mangel:1990:CTB

[3628] Marc Mangel, editor. Classics of theoretical biology: from material presented at a meeting held on 5 July 1988 in Oxford, UK. Bulletin of mathematical biology. Pergamon, New York, NY, USA, 1990. ISBN ???? 334 (vol. 1), 3326 (vol. 2) pp. LCCN ???? Two volumes. Printed in Bulletin of mathematical biology (ISSN 0092-8240) 52(1/2) 1990 and 53(1/2) 1991.